A Complete Bibliography of Publications in the International Statistical Review

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 August 2024 Version 1.26

Title word cross-reference

$100 [Blu20]. $104 [Ana21]. $105 [Liu22a, Liu22b]. $115 [Liu20b]. $115.00 [Blu21]. $120 [Fis22, Mai22]. $128.95 [Sun22]. $129.00 [Ric18]. $140.00 [Mer20a]. $154.40 [Mai19]. $165 [Pod21b]. $170 [Sen23]. $174.95 [Mer20c, Mer20b]. $190.00 [Ng21]. $199.99 [Dur22]. 2 × 2 [HH93]. $39.95 [Slo23]. $44.21 [Sen23]. $45.95 [McC22]. $46.75 [Tso20]. $54.99 [Hau18]. $59.95 [Slo24]. $59.99 [Cox16, Pod22a]. $68 [Ric17a]. $69.95 [Gho22]. $74.96 [Fis24]. $74.99 [Sun19]. $79.99 [Mar21]. $89.99 [Liu17]. $90 [Zha21]. $92.99 [Liu20b]. $95.00/£ [Dek22]. $96 [Pod22b]. $99.00 [Cox19]. $99.95 [Fis24, Lan18, Liu18, Var20]. ^{(}R) [Ros15b]. C_{p} [BCF^{+}14]. χ^{2} [CR89, NL75]. F [NL75]. f100 [Koe79]. g [Qui21]. G^{2} [CR89]. H [LK16]. K [MPR12]. L [Rém84]. L_{1} [CG12, VBL13]. M [DC19, La 16]. n [Kin75a]. p [CK19, KX09, OH19, Tho06]. P [at Least t Events] [Hop93]. P [Exactly t Events] [Hop93]. R [SBB^{+}12, SS17]. R^{+} [SDT^{+}23]. R^{d} [Cas91]. t [AG08, BBS77, DM05, FS08, LN74, NK06, NK08, PWKO24]. X [BS92]. z [BNKS82].

-Banknotes [Koe79]. -Distribution [LN74]. -estimators [La 16]. -GARCH [PWKO24]. -Geographic [Pod24]. -Indicators [SBB^{+}12, SS17]. -likelihood [LK16]. -Median [CG12]. -Quantiles [DC19]. -Random [BS92]. -Server [Kin75a]. -statistic [BBS77]. -sufficiency [Rém84]. -Value [KX09]. -Values [OH19, Tho06, CK19].

/$55 [Liu22b].

0 [Ana21, Cox16, Els20, Fis19, Liu19, Mar21, Mer17, O’B20, Pod22a, Qiw20, Zha21]. 0-19-289787-X [Liu22b].

1 [Ana21, Cox19, Dur17, Dur18, Dur23a, Els18, Fis18, Lan18, Mai21, Pod18, Rod21, San17, Slo24, Tso16, Unw21, Zha17b]. 1-119-68988-0 [Ana21]. 10.1111/insr.12432 [Ano21]. 100.00/EUR [Hee16]. 100th [Sch84a]. 115/$150 [Sun24]. 11th [Deb78]. 125/$160 [Kha24]. 139-2 [Sen23]. 173-6 [Sen23]. 189 [Daw02a]. 19 [HRR20, SLY23]. 1920s [Pea74]. 1966 [BBS07]. 1990 [WG97].

2 [BPV21, Gho23, Hau18, Hee16, Kha24, Mer18, Nor73, Num20, Sun19, Zha16b]. 2-01-801062-X [You19]. 2004 [Ano05a, CM06, FL05]. 2009 [Dur13b, Han13g]. 2021 [RBB^{+}21]. 20th [Bor09]. 21 [Num23]. 21st [Mar21]. 23.99/$29.99 [Kha20]. 2nd [Cox07, Emb08, Gho08a, Han08e, Lis07a, Lis07b, Sta07a, Tha07b, Tra07, Tra09, Val08, Veh07a].

3 [Gho22, Mai22, Ng21, Pod21b, Ric17b, Ros19, Rui17, Tso17, Zha19b, San77]. 39.19/$89.95 [Pod24]. 39ième [Ano72a]. 39th [Ano72b]. 3rd [Liu20b, Pun07a].

4 [Cap19, Fis22, Kha17, McE17, Pod17a, Ric18, Rob19, Unw24]. 41 [Hoy76]. 43 [BT76, Ott76]. 437.50/$229.00 [Ros19]. 44 [And77a]. 45 [Goo78, Moo78, Sch78]. 4th [Eva07a, Man07, Pos78b].

5 [Dur22, Kha20, Liu20b, Liu22a, Pod21a, Sun22, Unw20]. 50th [Ano76a]. 51 [dBPR86]. 54 [And87, GS87]. 56 [Fis24]. 56.99/$74.95 [Dur23a].

6 [Blu18, Blu20, Gho17, Gho20, Liu20a, Liu21, Mai17, Ric17a, Var20, Zha18]. 60th [Liu07]. 65/$100 [O’B20].

7 [Blu21, Liu18, Mai16b, Mer20c, Mer20b, Num19, Slo23, Sun17, Sun18, Tso20, Var19, Zha17a]. 70 [Daw02a]. 70/$85 [Gho20]. 74.99/$99.95 [Qiw20]. 75/x2007 [Ano07e]. 75th [Csi08]. 76/2008 [Ano08b]. 77/2009 [Ano09b]. 78/2010 [Ano10b]. 79/2011 [Ano11b]. 7th [Dur13b, Ios79b].

8 [Dek22, Dur23b, Liu17, McC22, Mer20a, O’B21, Sun16a]. 80.99/$108 [Liu23]. 80/2012 [Ano12b]. 82.20EUR [Fis18]. 85k [dBPR86]. 8th [Sun19].

9 [Alb19, Dur21, Mai18, Mai19, Pod17b, Pod17c, Tso18, You20]. 95th [Ada80]. 978 [Alb19, BPV21, Blu18, Blu20, Blu21, Cap19, Cox16, Cox19, Dek22, Dur17, Dur18, Dur21, Dur22, Dur23a, Dur23b, Els18, Els20, Fis18, Fis19, Fis22, Fis24, Gho17, Gho20, Gho22, Gho23, Hau18, Hee16, Kha17, Kha20, Kha24, Lan18, Liu17, Liu18, Liu19, Liu20a, Liu20b, Liu21, Liu22a, Mai16b, Mai17, Mai18, Mai19, Mai21, Mai22, Mar21, McC22, McE17, Mer17, Mer18, Mer20c, Mer20a, Mer20b, Ng21, Num20, Num23, Num19, O’B20, O’B21, Pod17a, Pod17b, Pod17c, Pod18, Pod21a, Pod21b, Pod22a, Qiw20, Ric17a, Ric17b, Ric18, Rob19, Rod21, Ros19, Rui17, San17, Slo23, Slo24, Sun16a, Sun17, Sun18, Sun19, Sun22, Tso16, Tso17, Tso18, Tso20, Unw20, Unw21, Unw24, Var19, Var20, You20, Zha16b, Zha17a, Zha17b, Zha18, Zha19b, Zha21]. 978-0-19-289787-9 [Liu22b]. 978-0-19-289788-6 [Liu22b]. 978-0-19-884129-6 [Blu20]. 978-0-19-885998-7 [Blu21]. 978-0-19-934953-1 [Cox19]. 978-0-1988-6275-8 [McC22]. 978-0-367-13559-1 [Unw21]. 978-0-367-17387-6 [Liu23]. 978-0-367-18678-4 [Sun24]. 978-0-367-34585-3 [Ng21]. 978-0-367-54 [Sen23]. 978-0-367-61970-1 [Slo24]. 978-0-367-76344-2 [Kha24]. 978-0-3677-0505-3 [Gho22]. 978-0-387-87810-2 [Zha16b]. 978-0-429-05646-8 [Liu23]. 978-0-470-68205-0 [O’B20]. 978-0-471-68717-7 [Var19]. 978-0-674-97557-6 [Blu18]. 978-0-691-19729-6 [Gho20]. 978-0-691-21436-8 [Dek22]. 978-0-8153-8410-6 [Liu20a]. 978-036748-388-3 [Mai22]. 978-1-003-08806-6 [Sen23]. 978-1-032-12578-7 [Slo23]. 978-1-032-20067-5 [Pod22b]. 978-1-032-36215-1 [Dur23a]. 978-1-032-37674-5 [Liu23]. 978-1-032-44595-3 [Pod24]. 978-1-0323-5758-4 [Unw24]. 978-1-107-05321-2 [Mer18]. 978-1-107-14989-2 [Sun19]. 978-1-107-15075-1 [Pod18]. 978-1-107-18592-0 [Liu19]. 978-1-108-47241-8 [Dur23b]. 978-1-108-49420-5 [Unw20]. 978-1-108-49433-5 [Liu22a]. 978-1-108-49802-9 [You20]. 978-1-1084-8295-0 [Pod22a]. 978-1-118-00232-2 [Hee16]. 978-1-118-02755-4 [Ric18]. 978-1-118-07858-7 [Num19]. 978-1-118-12334-8 [Mer20a]. 978-1-118-77104-4 [You19]. 978-1-118-79619-1 [Els18]. 978-1-119-07582-0 [Mer17]. 978-1-119-18684-7 [Tso20]. 978-1-119-24687-9 [Alb19]. 978-1-119-27204-5 [Sun22]. 978-1-119-43919-6 [Var20]. 978-1-119-51493-0 [Els20]. 978-1-119-52307-9 [Dur21]. 978-1-119-53697-0 [Zha21].

978-1-119-54116-5 [Liu20b]. 978-1-119-68988-1 [Ana21]. 978-1-119-82118-2 [Gho23]. 978-1-119-95324-1 [Fis18]. 978-1-119-97966-1 [Lan18]. 978-1-138-19637-7 [Sun17]. 978-1-138-35109-7 [Mer20c, Mer20b]. 978-1-138-48469-6 [Liu21]. 978-1-138-49998-0 [Fis19]. 978-1-138-57827-2 [BPV21]. 978-1-138-72698-7 [Sun18]. 978-1-1386-0818-4 [Fis22]. 978-1-316-60105-1 [Dur18]. 978-1-316-64221-4 [Rob19]. 978-1-4200-9087-1 [Zha17b]. 978-1-4398-7651-0 [Qiw20]. 978-1-4398-8442-3 [Tso17]. 978-1-4398-9311-1 [San17]. 978-1-4665-6149-6 [Zha18]. 978-1-4665-6973-7 [Mai16b]. 978-1-4665-7320-8 [Sun16a]. 978-1-4822-4406-9 [Pod17c]. 978-1-4822-4944-6 [Ric17a]. 978-1-4822-5344-3 [Rui17]. 978-1-4939-8853-2 [Num20]. 978-1-4987-0127-3 [Ric17b]. 978-1-4987-0129-7 [Zha17a]. 978-1-4987-1917-9 [Mai18, Mai19]. 978-1-4987-2087-8 [O’B21]. 978-1-4987-2448-7 [Liu18]. 978-1-4987-4622-9 [Pod17b, Tso18]. 978-1-4987-8185-5 [Pod21a]. 978-1-63526-152-3 [Zha19b]. 978-1-64295-806-5 [Kha20]. 978-1-78630-141-3 [Pod21b]. 978-10322032-56 [Fis24]. 978-3-030-31491-0 [Mar21]. 978-3-030-52680-1 [Rod21]. 978-3-030-56484-1 [Mai21]. 978-3-030-74941-5 [Dur22]. 978-3-319-16064-1 [Tso16]. 978-3-319-28156-8 [Liu17]. 978-3-319-32773-0 [Cox16]. 978-3-319-32861-4 [McE17]. 978-3-319-38764-2 [Hau18]. 978-3-319-45795-6 [Mai17]. 978-3-319-45807-6 [Gho17]. 978-3-319-50016-4 [Pod17a]. 978-3-319-66193-3 [Ros19]. 978-3-319-72346-4 [Cap19]. 978-981-10-2913-4 [Kha17]. 978-981-10-3300-1 [Dur17]. 978-981-19-9485-21 [Num23].

A. [Hal84a, Sen84, Tam84, Tho84]. Aalen [Gho09h]. Ababa [Yim90]. ABC [PSKL^{+}23]. Abhaya [Num09b]. Aboul [O’B09b]. Abowd [SV22]. Abraham [Sto76]. Absence [Smi20]. Absolute [Gow79, NW82]. Abstract [Dur13a, Loy76]. Abstracts [Knü80]. Abundance [Seb92, Cox75]. Abuses [LLP01]. Academia [FMS95, GL19]. Academy [ANK03]. Accelerated [ME93]. Acceleration [Wil07]. acceptance [WC75]. Access [MDS94]. Accessible [WPRP17]. Accidental [Dra13a]. Accidents [And73b, LH73, CP80]. Accomplish [O’B12e]. Account [LvR74]. Accountability [ZZN^{+}22]. Accountants [Buc77c]. Accounting [AH17, FW04, Has72, Keu97, MP23, OS73, Rad99]. Accounts [AH17, LA96, Bje73]. Accreditation [Fis05, Sti05]. Accuracy [Fel73, FM95, Ken97]. Accurate [Koe79]. Achim [Sco09b]. Acknowledgement [Ano99a, Ano00a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a]. Acquiring [Mac16]. across [DK87]. Act [CKK^{+}11b]. Action [Cle01, Dra14d, Wel92, Tip80]. Activity [OA91]. Actually [Men16c]. Actuarial [Bin07, Cha15, Dur15a, Kos11, Sta12b].

Acute [KMM^{+}17]. Acyclic [Cas20]. Ad [MDG^{+}92]. Adam [Han11a]. Adams [Spe10e]. Adamu [Som77b]. Adaptive [BK16, Cha08, Dem18, SS17, Zha16c, FMS12a, Bra15, O’B07a, O’B08a, O’B12a]. Adarsh [Sen23]. Addenda [Ano72c]. Addendum [Dav72, Har75a, Som79a]. Addis [Yim90]. Additional [Piz10, Joh80]. Additive [TT20, TB22]. Additivity [SKK18]. Address [Wyc03, Mal97]. Adelchi [Dur14c, Mai13b]. Adequacy [Mas73, McL93]. Adhikari [Sen84, Tam84, Tho84]. Adi [Roh76a]. Adjusting [JSW11, SH96, AM93b]. Adjustment [MR22, RJ91, KT93]. Adjustments [Ano83a, GP11a, Lit86, PF82]. Adkins [Pos78d]. ADMB [Liu12b]. Administratien [Ano74i]. Administration [Ano72k, Ano72l, Ano72m, Ano73i, Ano73j, Ano73k, Ano74j, Ano74k, Ano74i, Ano75m, Ano75n, Ano75l, Ano76j]. Administrative [Chi22, Hof97, LTS98, TH98, Sun22]. Admissibility [Ruk95]. ADMM [YL17]. Adnan [Han12i]. Ado [LSS05]. Adriaan [Cro07b]. Adrian [DS20, Unw20, AMR23]. Adult [Ahl02, Gal02a, Wel02]. Adulthood [Wat02]. Advanced [And78, Bar76, Han13k, Has08d, Hen78, JLS20, Liu10a, Pos79c, All79a, Dur14a, Pat75a, Pun12a]. Advances [MMC13, Rod21, ZgZ89, Fil11, Ios76b, Pod14a]. Adversarial [SAG^{+}16]. Aeilko [Ros09c, Sta12e]. Aerospace [Lee79]. Affecting [HBK05]. Afriat [Cra80]. Africa [EB92, HRR07]. African [CN98, Kir82, NSS73, SDGD^{+}11b]. After [HY01, Kab09, AS72, Chi22]. against [Jag86, Tip75a]. Age [Ale94, Cha20a, KMM^{+}17, NSS73, O’B15c, Ram03, Roy94, Smi94a, Smi94b, Sun19, Tam94, VJ95, Dav09, Mer12a, San15]. Age-Period-Cohort [O’B15c, San15]. Agencies [Bin00, Cla00a, Dil00, Lyb00, Tre00b]. Agent [HWS20]. Agent-Based [HWS20]. Ages [Sti88]. Aggrawal [Sen23]. Aggregate [O’B15c, San15, Rui72]. Aggregation [Abr82, Kle76, Men73, Pro11, SH96]. Agility [Nem73]. Agoston [Bre79a]. Agreement [Han13j, Num19, Han13o, O’B10b]. Agreement-Models [Num19]. Agresti [Lis07a, Lis07b]. Agricultural [Ayh01, CG05, ZBA76]. Agriculture [Som78d, Kha73]. Ahmed [Kha24]. AIC [DNC06]. Aided [Sch07, Yan73]. Aidenoff [Sto76]. Air [Doc22, DSC03]. Aitkin [Sta12f, Han11k, Sta12f]. Aiyi [Han13m]. Ajit [Dra10c]. Ajustement [AM93b]. Akaike [BCF^{+}14]. Akalin [Pod21a]. Alain [O’B07b]. Alan [Bat08b, Han09c, Kha80, Lis07a, Lis07b, Pun08c, Sen14d, Sta12d, Tra13]. Albert [Cro08b, Ken76a, Pun11b, Sta08c, Bau14]. Alberta [Liu10b]. Albino [Col72]. Alcohol [WD84]. Alden [Mai11b]. Aldo [Ros10b]. aléatoires [Ios77b]. Alessandra [Fil11]. Alessandro [Liu22b]. Alessio [Pyn14b]. Alex [Sta07e]. Alexander [Cro10e, Dur21, Han10j, Lis14, Puk07]. Alexandre [Sen11a]. Alexandru [Blu21]. Alexey [Han10j]. Alfréd [Kin75b, Für76, Sch72a]. Algebra [Dra14a, Fil09b, Han12b, Pun07b, Pun12c, Spr77a]. Algebraic [Dra12a, Bre79a, Gho11a]. Algorithm [BR09, CD18, HT92, Jør84, Mor84]. Algorithmic [FHTV93, Gho23, Han10d]. Algorithms [CKL14, Don12, Dur13c, HL90, Han08d, HGKJ20, LZ22, Lar15, O’B10d, O’B16b, RG23, RDMP16, Sun19, ZZN^{+}22, dWC05, SLY23, Kha17, Nor13a, Mai17, O’B11a]. Ali [Kos13a, O’B13b, Pod14a, Pun13c]. Alice [Gho20]. Aligning [BK20, Han08g]. Allan [Cro10a, Sco07]. Alleged [Goo78]. Allerhand [Sen14c]. allocating [FMS12a]. Allocation [Arm85a, BCL^{+}85, Bat85b, Che85b, Lou85, Pet85, Whi85, Bat85a, Rao77]. Allocations [Gal80a]. Alquier [Han13g]. Alsina [Han11c, O’B10e]. Alternative [CNS85, Mac14, NSW72, NWE16, OS73, Ott75a, Smi72, SD11, BG83, Bic13, LHK78, Ott76]. Alternatives [Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, Tan92, FX16, Mai16b, Pun07a]. Altuna [Pod21a]. Alvin [Pun08b, Pun13b]. Alwin [Veh08f]. Alyson [Gho09d]. Amendment [And87, Ano83a, Ano84a, Ano86a, Ano88a, Ano89a, GS87, dBPR86, FT88]. Amendments [Ano87a]. America [HRR07]. American [Pos78d, Fis09]. Amiable [SAG^{+}16]. Amicorum [Spe08a]. Among [DMMG22, FGFBGM22, KMM^{+}17, KY91, dBP77]. Amos [Cox19]. Amount [RW09]. Ana [Bat08b, Sco13c]. Analisis [Tin78a]. Analyse [Dag75, Lin75a, Pre79b, Lin80a, Mar07a]. Analyses [Eva07a, MBW^{+}23, O’B15c, Rob11c, San15, She00, Dru11, Cox07]. Analysing [EM19, NAL76, SH96, lRG05, O’B07b]. Analysis [Alv11, AJS81, Ano76m, Aur10, Aus17, Bar80b, Bar80a, Bar77c, BS89, Beh04, BL24, Bel88, BJO^{+}98, Bil77a, Bil77b, BD94, Bre75, Buc79, CD18, CHXY19, CK19, Cro08e, Cro10f, Cro11g, DK01, DG90, Dra09b, Dra10c, Dra11a, Dra12c, DK87, Dur13b, Dur14a, Ela73, FM95, Fil09b, Fil09a, Fin77c, Fis18, Fis24, FH89, FMR99, FFBK76, Gan78b, Gan80b, Gel03, Gho09i, Gho09f, Gho09h, Gho10a, Gho11b, Gho12f, Gho12b, Gho17, GR90, Goo86b, GP11a, GW14a, GW14b, Hai72, HL19, Han08n, Han08l, Han08h, Han09a, Han10f, Han10e, Han11a, Han12n, Han12v, Han13a, Han13h, HLC97, HM04, Har96, Hau01, He14, HR83b, HR83a, HK77, Hil15, Hir93, Hoe87, Hub14, IKS^{+}81, IKS^{+}82, JK78, Ken76b, Kha17, KW19, KFL76, Koc14, KC14b, Law98, LW13, Lep14, LNA75, Liu08, Liu10a]. Analysis [Liu14b, Liu23, LC20, Loy77, Loy80, Mai08, Mai09c, Mai11c, Mai13a, Mai13b, Mai14b, Mai16a, Mak78, MM96, MM01, MW81, MBW^{+}23, MSW76, NY19, Nan10a, New81, New84, New88, Nor15, Nov73, Num13a, O’B08d, O’B12d, O’B12c, O’B14c, O’B16a, OH13, OY92, Ott75a, Ott75b, PZ21, PSI19, Pla88, PCA^{+}19, Pol11b, Pol11a, Pol13d, Pol13b, Pos73, Pos76a, Pos78c, Pos78b, Pos79d, Pos80d, Pre78a, Pre79a, Pre80a, Pre89, Pun10b, Pun12b, Pun13c, RM93, Rob11d, Rob11f, Ros15b, San17, SFG10, STI06, Sco10b, ŞA09, SJ89, SAG^{+}16, Ski19, Sko12, Sma88, Spr75a, Sta07a, Sta08a, Sta12d, Sta12e, Stu80, Sun15, Sun16b, SM89, Szc11a, Tal93, Tam87, Tho14, Tin78e, Tju84, TBSW84, Tsa24, Tut21, UT10, Unw11b, Unw11a, Unw12a, Unw21]. Analysis [Veh08b, Veh08c, Veh08d, Veh10d, VL96, VF00, Vid15, Wil14a, Wil14b, WB97, WY02, ZHW23, ZgZ89, Zha21, Zhu14, Zur91, dBV82, And87, BH80, DY16, GS87, Goo86a, Hal86, Har79, KS85, KFF75, KASG80, Lai22, Lau81, LL87, LOD^{+}85, Mor84, OC13, Ott76, Sax78, Sch83, SSJ77, Sha88, Cro09b, Cro10c, Dek22, Dur21, Fil11, Gho08b, Gir73, Gow72, Han08i, Han10h, Han11h, Han12r, Han12w, Han13m, Ken73a, Khu13a, Lis07a, Lis07b, Muk14b, Num14, Oga73, Pos76c, Pos78b, Pre80b, Pun13b, Pun13a, Ros09b, Sai08, Sen14d, Sen23, Smi75, Smi77, Spr80, Str73, Sun17, Szc11b, Unw10b, Veh08a, Veh09a, Veh10a, Veh10b, Wet80, Zha16b, Zha17a, Cro10b, Der75a, Dra09a, Gil15, Han08b, Kha15, Liu16, Nel72b, O’B12b, Ric18, Rui16, Sco09a, Sun16c, Cro11d]. analysis [O’B09e]. Analysis-Theory [Fis18]. Analytic [Han08c]. Analytics [Cap19, Han13i, O’B12f, Pod18, Rod21, Sco13a, Sta07b, Unw24]. Analyzing [NKH91, PGN^{+}96, Bau14]. Anand [Sen23]. Anatoli [Gho20]. Anatoly [Ng21]. Ancestry [Tho98]. ancillary [BG83]. Anders [Gho08b, Liu22b, Sch80a]. Andersen [Aal82, CV82, Dra11c, Duf82, Har82, Kos10b, LT82]. Anderson [Dra09d, Ken73a, Pre79a, Wim75, Ano87a, Mai09a, Ord80a, Tin78e, Unw12b]. Ando [Rob11e]. André [BPV21, Unw13b, Dra10a, Han11d]. Andrea [Fis10a, Liu08]. Andres [Liu08]. Andrew [Gho09a, Gho13c, Häg12a, Han13d, Mai10a, Mai16b, O’B09c, O’B12f, Pol12a]. Andrews [EH91]. Andrey [Han10j]. Andriy [Han13m]. Andrzej [Mai14c]. Andy [Liu10b, Unw13a]. Angel [Alv11]. Angelini [Bil80a]. Angewandte [Dei76]. Angiolo [Tin78b]. Anh [Bla07]. Anil [Mit79]. Animal [Cox75, Seb92]. Anirban [Gho12d, Han13k, Oja09]. Ann [Dek10, Mai11a]. Anna [Man08]. Anne [Han09b, vdL80c]. Anneliese [Han13l]. Anniversary [Ano76a, Sch84a]. Annotated [AG76, Dix73, Fis83, Joh77, Joh80, Sax78, Sha80, Sri80]. Announcement [Ano75a, Ano76b, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano80a, Pin10a]. Announcements [Had10a]. Announcing [Had10c]. Annual [Ano07e, Ano08b, Ano09b, Ano10b, Ano11b, Ano12b, HL01]. Anomalies [GPdMLMGGSG20, JR08, JR10, KG10]. Anomaly [PGV22]. ANOVA [Ros10b, NAL76]. ans [Fou80b]. Answers [Pun11c, Sti14]. Anthony [Dek22, Fil11, Gho13c, Han10h, Puk07]. Antieigenvalue [Pun12b]. Antonella [Cam07]. Antonio [Jer78, Lis10b]. Antony [Veh08e]. Antti [Han08l]. Anwendungen [Kri74]. Aoyama [Han10a]. Apart [Dra13e]. apercu [EMR88]. Applebaum [Han08j]. Applicability [Fal72]. Applicable [LH73, CP80]. Application [Ano76m, BCM19, BJO^{+}98, BCO82, CFP22, CG12, DMS06, FFBK76, GP73, Hoe87, HzB91, Ios78, LQ04, Mat73, MFAG10, NSS23, O’N04, Pyn14a, She00, VT19, Vin73, WC98, MMC13, NSZH22, RLS23, SSJ77, SLY23, Han09c, Lis14, Man72, Nor14b]. Applications [ABGK82a, ABGK82b, Ano76f, Ano84a, Aus17, BT75, BT76, BFST18, Cap19, CKN^{+}24, CKL14, Cor75b, Cro08c, Cro10g, Der75a, Dur22, EFP19, Gho10a, Gri15, Han07a, Han10j, Hir93, HCV^{+}21, Ios76f, JT13, Lef79, Liu20b, LC20, Luo07, ME89, Mai09c, Meh15, Mor80a, Mul76, MW76, Nor14a, O’B11c, O’B13a, O’B14c, Oja08b, Pol13d, Pos78e, Pun07b, Pun12b, RYG^{+}21, Ros10b, RC93, Sär11, Sen11b, Sri73, Sta13c, Tra73, Veh07b, Veh14, Ver13, WS01, Wan19, Yas93, DY16, Har79, Mor84, SS87, Shi81, dF74, Tip77, Bar75d, Bil77b, Cro09d, Cro09c, Cro10h, Dek14, Dra10c, Dra12c, Dur13c, Gho12g, GP11b, Haj77, Han08d, Han08k, Han10g, Han11d, Han12g, Han13c, Ios77d, Kos11, Kos12a, Lis13, Liu11, Mai10b, Mai11f, Mat07, Mer20a, Nor13b, Num19, O’B09b]. Applications [O’B12i, Pun10c, Pun11b, Rad77, Roh76a, RR79b, Roy14, Szc11a, Vos75, You14, Zha19b, vdL80a, Esp11, Cor80, Lar15, O’B15a, Pre77a, Zha16a]. Applied [Coh75, Dem77, Dur17, Dur18, Gho08d, Gho21, Gow79, Han10h, Han12k, Hen78, HK77, Huz80, Liu23, Mai10c, PGV22, Ros09c, Sco13a, Som79c, Som80c, Sta09, Unw12a, Veh09a, Bil77a, Cro07a, Cro10c, Elf72, Eva07b, Fis10a, Han08c, Lis09a, Mai13a, Num14, O’B13b, Oja08a, Ric13a, Ros09a, Sco09b, Sen14d, Spr75a, Veh07a, Veh08a]. Applying [lRG05]. Appraisal [MM04, Tri09]. Approach [BK20, BR02, BCG07, Bow04a, BMM08, Cor14, CM06, DS99, DK01, DW15, EM19, FMR^{+}20, FS94, Fou80a, GP73, Goo86b, Gue90, Ham78, Han08a, HS88, Ich09, Ios78, KS89a, KS89b, KX09, Lav14, LHSD15, Liu22a, Mai10c, MGMRPA90, Nee09, Nov73, O’B15b, OH13, RW09, RN92, Ros15b, SPC73, SB04, And87, GS87, Goo86a, Blu20, Buc79, Dra11d, Dra12d, Fis24, Gho09c, Han07c, Han08j, Han09b, Han11k, Han13n, Han13h, Pol13b, Ric10c, Rob11g, Som79c, Sta07a, Tin78e, Unw11a, Unw12b, Wim75, Ric15, Sun16c]. Approaches [FW04, FST19, HPHdAM22, Hoo05, KMT03, NKH91, O’B14a, O’B15c, PZ21, Rob11b, San15, SFG10, Sik18, SW15, SD11, TB22]. Approximate [HK18, Jas15, NN20a, NN20b, Kin75a]. Approximation [BD09, CW91, Cro09e, Háj72, Hou88, Ios77a, Tan92]. Approximations [Cro11g, Dan87, LR91, Pun10d]. approximator [SLY23]. Arab [Fer94]. Arbuthnot [Bel89]. Archaeologists [Dra11d]. Area [BCM19, BSS18, BFST18, CDM19, CL19, CMH19, DC19, Dor18, Eid07, GW23, MSS20, NY19, Num23, Pfe02, PBH19, SH21, SST20, WOP20, Hed20, RLS23]. Areas [Ano80e, Cle01, Con09, Mas73, PK80, Tam82]. Argument [Ano96d, Bar95, FB96, Sho85, Fer73]. Ariel [Dur14a]. ARIMA [Gue90, MN98]. ARIMA-Based [Gue90]. Arising [GMZ03]. Aristides [Dur13a]. Aristotle [Mis08]. Arjun [Gho12c]. Arkadi [Gho20]. ARMA [God97, MN98]. Armitage [Bat85b, BCL^{+}85, Che85b, Lou85, Pet85, Whi85]. Arnaud [Unw13b]. Arner [Dra14c]. Arnholt [Bat08b]. Arnold [Szc11b, Pun11b, dO79a]. Arribas [Pod24]. Arribas-Bel [Pod24]. Ars [Syl14]. Art [Mai12a, McC22, RR79c, Rus10, Iac84, NW82]. Arthur [Luo07, Mer09a]. Article [gXZ20]. Articles [LSS80b, LSS80a]. Artificial [Mai11a, Yoo21, ZZN^{+}22]. Artigas [Jer78]. Arup [Mer20c, Mer20b]. Ary [Pos80a]. ASA [Sun18]. ASA-CRC [Sun18]. Asaph [Sun22]. Asghar [Dek10]. Asher [Sta08b]. Ashok [Han10g, Pod14a]. Asia [Hug06]. aspect [WC75]. Aspects [And73b, BNS94, Cox86, DER03, Ell91, FW04, Kis94, Oak81a, Oak81b, Pur72, ZW19, vBS73, Zha16c]. Assess [CM06]. Assessing [BO15, CFP22, GMZ03, Han12a, JRL10, Kaf15, Ney75, Rao93]. Assessment [EIRS18, FGC03, GG99, Sta12f, dWC05, Han09a, O’B14c]. Assessments [Mai18]. Asset [Kos09a, TS06]. Assignment [SS88]. Assimilation [BEW03, Han07c, Pod21b]. Assist [FB73]. Assisted [CM12, Nov73, WQS19]. Associate [Pat75b]. Association [Bic00, CS88, Deb78, Fis10a, LHK78]. Assumptions [AdB04, SRN17, SAA22]. Assurance [Car93]. Assuring [FL05, OMGT05]. Astronomer [SS09b]. astronomical [Hal86]. Astronomy [Pod14a]. Aswani [Mai11b]. Asymmetric [Alv20, GKV19, GKV20, HK16, Koe16, Men16c, MFAG10, Smi16, WS16, YWH16a, YWH16b]. Asymmetry [Tsi03]. Asymptotic [BNS94, BS89, EGS72, Hou88, Ios77a, Le 90b, Lis10b, dO79b, vdV91, McC85, You20, Oja09]. Asymptotically [FMS12a, Loy76, Sen76, Bin83]. Asymptotics [Cro07a, GP12, La 16, Lia20, NK14, O’B16b]. Atallah [O’B11a]. Athens [Mis10]. Athreya [Rad75]. Atkinson [Fil11, Puk07]. Atreus [LKC^{+}04]. Atta [Gan80d]. attack [OC13]. Attainment [Nov73]. Attention [BP12]. Attribute [RW09]. Attribution [CS20, Efr20a, FHT20, gXZ20]. Audio [God97]. Auditing [Ano93d, Bic92]. Aufgaben [Lin75b]. Augmentation [Ken10]. August [Ano05a, FL05, Han13g]. Austin [Kró80a]. Australia [TG94, Tre05b]. Australian [BCM19, CS00, Pow03, TC15]. Author [Bar72c, Har76]. Autobiography [Spe11b]. Autocorrelated [CD94]. Autocorrelation [FFS11, Smi73]. Autocovariances [NK14]. Automated [BPV21, OC13]. Automatic [BP92, Tan92, dWC05]. Automatische [Lin75a]. Autoregressive [Bha91, KY91, LD24, AL86, LL87, dGAGR85]. autoregressive-moving [dGAGR85]. aux [Ano75a, Ano76b, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano80a, Dag76, Dam74]. Auxiliary [AT15, ES06, Mer24, Sär11, Sha23, JOT85, ZTK13]. Available [Mas73, Pat88, Sun96]. Average [Bha91, KY91, LHK78, LG08, MFAG10, AL86, dGAGR85]. Averaged [NKH91]. Averages [Sch16b]. Averaging [FGDS18, FBL18, Lis09c]. Avis [Ano75a, Ano76b, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano80a]. Avoidance [NSS73]. Award [RBB^{+}21]. Awareness [Fis05, Smi01, Sti05]. Axial [Ell91]. Axiomatic [Bal95]. Ayanendranath [Han13n]. Azzalini [Mai13b, Dur14c].

B [Ana21, Bar72b, Bar80b, Bar72c, Bre79c, Bre80, Buc76a, Buc76b, Buc77b, Buc78, CS20, Dra11e, Fin80c, Fis72, Fon13, Gan72a, Gan80c, Gho08b, Gho09a, Han08o, Han09b, Han11c, Imb07, Ios79a, Ken76a, Ken76c, Kha80, Lin75b, Lis12, Liu12b, Liu14a, Mai11a, O’B10e, O’B12d, Pet07, Pod24, Pol13b, Pos78a, Pos79d, Pos80e, Pre78b, Pun10c, Pun12c, Rad75, RR79c, Sen11a, Som76a, Som78f, Spe10c, Sta07f, Sta10a, Tha07b, Tin80b, Tip77, Unw11c, BBS07, Lia20]. B-splines-based [Lia20]. B/W [Pod24]. Babak [O’B13c]. Babette [Gho22]. Bacchieri [Cam07]. Back [Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano74a, Ano74b, Ano74c, Ano75b, Ano75c, Ano75d, Ano76c, Ano76d, Ano76e, Ano77d, Ano77e, Ano77f, Ano78d, Ano78e, Ano78f, Ano79b, Ano79c, Ano79d, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano82a, Ano82b, Ano82c, Ano83b, Ano83c, Ano83d, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano86b, Ano86c, Ano86d, Ano87b, Ano87c, Ano87d, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96c, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano98c, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano01a]. Back [Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano05b, Ano05d, Ano05c, Ano06b, Ano06c, Ano06d, Ano08c, Ano08d, Ano08e, Han12b]. Backer [Auk78]. Backward [AET04]. Bacterial [GK01]. Bagchi [Ios76d]. Baggett [Far13]. Bahadir [Kha24]. Bahadur [GO81]. Bailar [Sen11c]. Bailer [Rob19]. Bailer-Jones [Rob19]. Bailey [Dow72, Gan80a, Kin77b]. Balakrishnan [Mer20a, Sta13c]. Balanced [BPP17, Bre99b, Kag72a, Kag72b, Pat75b]. Balding [Gho13b]. Balkrishna [Kem80]. Bandos [Han13m]. Bands [LLP08]. Bandwidth [KSS14, Wan19]. Banerjee [Der77a, Fou80a, Pre77b]. Bani [Gho12a]. Bank [MA05]. Banking [HCV^{+}21]. Banknotes [Koe79]. Banks [Gho10b, Han12s, Sta08b]. Bapat [Liu14a, Pun12c]. Barbara [Veh14]. Barberi [Gua77]. Barbu [Cro09d]. Barker [Tha07a]. Barlow [Str76, Vin73]. Barndorff [Bil80b]. Barndorff-Nielsen [Bil80b]. Barnett [Der77c, Wet75]. Baron [Mai07, Mai14b, Der75a]. Barra [Pre78b]. Barrie [Sco77a]. Barry [Pun11b]. Bart [Gho17]. Bartel [Gho13c]. Bartelmus [Hau08b]. Bartholomew [Han08b, Han13h, Vin73, Sto75]. Bartlett [Buc77a]. Bartolucci [Pyn14b]. Basawa [Gay11]. Baseball [Bau14]. Based [AG08, BK88, BMM08, Bre99b, Das13, DC83, DW15, Fon11, Gue90, Had10c, HH93, Han13c, HM19, HWS20, Ich09, JSW11, KP04, KLTG21, Mai09a, NPT08, Pau19, Rob11g, ST07, SB17, SKK18, Sne99b, Sri73, SS07, TB22, Yim90, You19, ZL22, Zha19a, vdL03, Alv20, BH77, BD20, Fis10a, GKV19, GKV20, GTF18, GT17, HK18, Lia20, SGM21, Sta07a, Unw11a, Wan19, WGKL22, You86, O’B12d, Unw20]. Basic [Ani08, FB73, HH07, O’B10e, Veh14, Zha19b, dWvDS20, Tip77, Cor77a, Cro10d, Ric10a, Som77a, Som79b, vdL80a]. Basics [Haw97a]. Basis [Cox10, Hof97]. Bass [Han12o]. Basso [Ros10b]. Basu [Han13n, Gho12d]. Bather [Arm85a, BCL^{+}85, Che85b, Lou85, Pet85, Whi85]. Baum [Buc77d]. Baumer [Liu18]. Baxi [Ano72c]. Baye [Rob12b]. Bayes [Spe11a, Bic14, CDM19, DS99, DK01, Fal72, HY01, HM19, Hoi13, LdBP90, NT72, O’N09, TWRB20, vdHK99]. Bayesian [Alb19, BCZ12, BW12b, Dat12, Gho08f, Gho08e, Gho08g, Gho08c, Gho08d, Gho08h, Han10h, Han11k, Han12i, Kos10a, Liu23, O’B13e, Pal72, Pol11a, Pos78e, Rob11b, Rob19, Rui17, Sch78, Zha16a, Ald02, Bar87, BS89, BR02, BW12a, CFP22, Cas20, CKL14, CP96, Con09, DS99, DW15, FMR^{+}20, FMR99, FGDS18, For22, FBL18, Gel03, God97, GPW05, Hig77, HK16, Jas15, Koe16, LLP01, MYA95, MN98, Men16c, NY19, NC10, Nov73, OH19, PW91, PGV22, Red01, RDMP16, SFG10, SCW08, Sha23, SGLS02, Smi92, Smi16, Tho14, Tsi03, TC72, WS16, WS21, YWH16a, YWH16b, Blu20, Cha14, Hil15, Lar15, Rob11d, Smi75, Zha14, Bas16, Bla07, Bol07, Cro11a, Gho08a, Gho08b, Gho09c, Gho09b, Gho09a, Gho09d, Gho11b, Gho12a, GP11b, Han11a, Ken10, Mai11a, Nan10a, O’B10b]. Bayesian [O’B12c, Pod15a, Rob11e, Rui16]. Bayesian/Likelihood [Han11k]. Bayesianism [Goo78, Moo78, dF74, Ber77, Ham77a, Moo78]. Baylac [Mai09b]. Be [God97, LLP01, Spe10a, O’B13d]. Beck [dO79a]. Becoming [PR10]. before [Pla88, Wat02]. Beggs [Unw24]. Beginner [Veh11a]. Beginners [Blu20, Moo99, Som77b, Som77b]. Beginning [Zar90, Lin80c]. Beh [Fis18]. Behavior [O’B14a]. Behavioral [Bin00, Cla00a, Dek16, Dil00, Gho09c, Gho10a, Lyb00, Som78c, Sta12c, Sta13c, Tre00b, Veh08g, Veh10d, Mai14b]. Behaviour [Kin75a, OS73, Fis72]. Behavioural [Bar80b]. Behrens [MPR12]. Being [Rus10, MSS20]. Bel [Pod24]. Belgian [LvR74]. Bell [Ket12]. Belz [Tip75b, Wil75]. Ben [Roh76a, Ros19, Unw09]. Ben-Israel [Roh76a]. Ben-Zvi [Ros19]. Benchmarked [SH21]. Benchmarking [CW06, CD94, DCC98, DQ97, SST20]. Benedetto [Gua77]. Benefit [Ast94]. Benefits [Kró80b]. Bengt [Han08e]. Benjamin [And07, Liu18, Mer09a]. Bennet [Pod18]. Benoît [Ios79a]. Benyaklef [Rao80a]. Bereanu [Gan80c]. Berenblut [Lin76]. Berger [Pol11a]. Berglund [Cro10c]. Bergomi [Pod17c]. Bergsma [Kos10c]. Bergstrom [Tin80b, Rao80b, Kin75c]. Berk [O’B09h]. Berkeley [Spe07, Ric17a]. Berlin [Lid80]. Bernard [Cor77a, Fil11, Fin78, Som79b]. Bernardara [Dur22]. Bernardinelli [Roy14]. Bernhard [Fil09b, Gho13a, Sco09a]. Bernoulli [Ios76f, Sen14a, Hal84b, Sho85, Sti14, Syl14]. Bernstein [Sch05]. Bertka [Sta10b]. Bertrand [Mai11d]. Bertsch [Rob12b]. Berzuini [Roy14]. Bessiere [Fou80b]. best [FMS12a]. Beta [BGS24]. Bethea [Tip77]. better [PWKO24]. Between [GL19, HG21, ANK03, BG83, DG90, FMS95, KMT03, Lec76, LSD^{+}11]. Bevoelkerungen [Kro80c]. Beyond [Ber14, Hir93, Hop93, Ros15a]. Bharucha [Gan80b]. Bharucha-Reid [Gan80b]. Bhat [Coh75]. Bhattacharya [Han08k, Ios77a]. Bhattacharyya [Pos80b]. Bhimasankaram [Pun10c]. Bhopal [Rao93]. Bianca [Bol07]. Bias [Ayh01, Bet10, Jag86, NWE16, Sär11, ZZN^{+}22, Hed20]. Biases [Whi99]. Bibby [Som79e]. Bibliographic [Beh04]. bibliographical [Col81]. Bibliographies [Lan83, Lan85, Lan72, Lan73, Lan74a, Lan74b, Lan75, Lan76a, Lan78, Lan79, Lan81, Lan82, Lan84, Lan86, Lan87]. Bibliographique [Col81]. Bibliography [AG76, AA78, Ano73d, Bak73, Dal77a, Dal77b, Dal77c, DM73, Gan72b, Har78, Hoy73, Ken73a, Lan83, Lan85, LSS80b, MM04, Ovu80, PS80, dBP77, Sah79, Sow72, Sow78, WS79, BH77, Chm81, Fed80, Fed81a, Fed81b, Fis83, Hoy76, Joh77, Joh80, KZ76, KASG80, Lan72, Lan73, Lan74a, Lan74b, Lan75, Lan76a, Lan78, Lan79, Lan81, Lan82, Lan84, Lan86, Lan87, LSS80a, Nau79, Sax78, Sha80, Sri80, Fed81a, Fed81b]. Bicentenary [Sen14a]. Bickel [Gho10b]. Bid [MCH13]. Biecek [Unw21]. Bienaymé [BJS92, Sen98]. Big [FPR^{+}16, KW19, KT21, TC15, YL17]. Biggeri [Big05]. Bikas [Dra14b]. Bilinear [Ord80a]. Bill [Han07g]. Bin [Csi08, Gho12c]. Binary [DG90, FRS08, NKH91, O’B09e, PGN^{+}96, SLWH22, SCC14, DFM^{+}22]. Binomial [BD09, Ham72b, O’B08b, O’B11b, WHT24, HS12]. Bioassay [Fin79]. Bioassays [Sen72]. Bioavailability [Gho09f]. Biobehavioral [Pol13b]. Bioconductor [O’B09a]. Bioengineering [O’B13h]. Bioequivalence [Gho09f, Mat07]. Bioinformatics [Gho12a, Gho13a, Man08]. Biological [O’B13c, Rad77, Ric10b]. Biologists [Rou07]. Biology [Cor75b, Wil73, Gho13b, Pos79a]. Biomarker [Ric18]. Biomedical [Blu16, Cap19, Gho13c, Kró80d, Mai12c, Ric10a]. Biometric [Ald03]. Biometrika [Ken73b]. Biophysics [Gho13c]. Biostatics [Gri77]. Biostatistical [Num13b, Ric16]. Biostatistics [Doc22, Mas73, Sun24, Zha17b, dO13, Cor75a, Gho08c, Han12j, Num09b, Gho09e, Gun15, O’B13c, Unw12b]. Biosurveillance [Esp14b, Han12d]. Biplot [SAP97, lRG05]. Biplots [Veh11b]. Birger [Lin13]. Birnbaum [Fin80c]. Birth [Sen96]. birthday [Ada80, Hol86, Csi08, Liu07]. Biscuits [Mer09a]. Bisgaard [Han12v]. Bivariate [CW91, Dou95, Feu93, FS10, Ham72b, HH93, Huz73, LH73, Mod08, Wan19, CP80]. Black [Veh13b]. Blackwellisation [RR21]. Blalock [Bil77a]. Blanton [O’B11a]. Blink [Dra13e]. Bliss [Gre80, Wil73]. Block [Kag72a, Kag72b, Pat88]. Blocking [Bai85b]. Bloomfield [Gan78b]. BLUEs [Tia07]. Blunders [Sun18]. Boat [Tip75a]. Bock [Lin75a]. Bodin [Sap15]. Boelen [Gho09b]. Boes [Cro10b]. Bohdana [Muk14a]. Böhning [O’B09e]. Boisvert [Cro11e]. Boland [Bin07]. Boldrini [Par73]. Bolfarine [RBB^{+}21]. Bolla [Slo24]. Bolt [Amu73]. Bonferroni [HS90, HS12]. Bonferroni-Type [HS90]. Bonita [Dur13a]. Book [Alb19, All79a, Alv11, Amu73, Ana21, And78, And07, Aoy72, Aur10, Bal76a, Bal76b, Ban14, Bar77a, BN07, Bar75c, Bar75b, Bar72b, Bar75d, Bar80b, Bar80a, Bar72c, Bar76, Bar77c, Bar77b, Bas16, Bat07, Bat08a, Bat08b, Bau14, Bec73, BPV21, Bel07, Bel16, Ben75, Ber78, Bha07, Bil77a, Bil77b, Bil80a, Bil80b, Bin07, Bje73, Bla07, Blu07, Blu16, Blu18, Blu20, Blu21, Bol07, Boo79, Bra14, Bra15, Bre79a, Bre79b, Bre79c, Bre80, Buc79, Buc74, Buc76a, Buc76b, Buc77d, Buc77a, Buc77b, Buc77c, Buc78, Cac74, Cam07, Cap19, Cau80, Cha14, Cha80, Coc73, Coh75, vC75, Coo07, Cop77, Cor80, Cor75a, Cor75b, Cor77c, Cor77b, Cor77a, Cox75, Cox80, Cox07, Cox10, Cox12, Cox16, Cox19, Cra73, Cra80, Cro07a, Cro07b, Cro08a, Cro08b]. Book [Cro08c, Cro08e, Cro08d, Cro09b, Cro09a, Cro09e, Cro09d, Cro09c, Cro10j, Cro10e, Cro10b, Cro10d, Cro10a, Cro10h, Cro10f, Cro10c, Cro10g, Cro10i, Cro11e, Cro11b, Cro11d, Cro11f, Cro11a, Cro11g, Cro11h, Cro11c, Csi80, Csi08, Dag75, Dal72, Dei76, Dek10, Dek14, Dek16, Dek22, Dem73, Der75b, Der75a, Der77c, Der77a, Der77b, Don12, Dor73, Dow72, Dra08a, Dra08b, Dra09b, Dra09c, Dra09d, Dra09a, Dra10c, Dra10a, Dra10b, Dra11b, Dra11a, Dra11d, Dra11e, Dra11c, Dra12f, Dra12a, Dra12b, Dra12d, Dra12e, Dra12c, Dra13d, Dra13f, Dra13g, Dra13a, Dra13e, Dra13c, Dra13b, Dra14a, Dra14b, Dra14c, Dra14d, Dri78, Dru11, Dur13c, Dur13a, Dur13b, Dur14b, Dur14a, Dur14c, Dur15a, Dur17, Dur18, Dur21, Dur22, Dur23a, Dur23b, Eag74, EV73, Elf72, Els18]. Book [Els20, Emb08, Eng78, Esp11, Esp14a, Esp14b, Esp15a, Esp15b, Esp15c, Eva07b, Eva07a, Far13, Feh07, Fer73, Fer80b, Fer80a, Fil09b, Fil09a, Fil11, Fin75, Fin77b, Fin77a, Fin77c, Fin77d, Fin78, Fin80a, Fin80d, Fin80c, Fin80b, Fis10a, Fis18, Fis19, Fis22, Fis24, Fis72, Fon11, Fon13, Fou80b, Fou80a, Fra72, Gal80a, Gal80b, Gan72a, Gan76, Gan78b, Gan78c, Gan78a, Gan80c, Gan80a, Gan80d, Gan80b, Gay11, Gho08f, Gho08e, Gho08g, Gho08a, Gho08c, Gho08d, Gho08h, Gho08b, Gho09c, Gho09g, Gho09b, Gho09e, Gho09i, Gho09a, Gho09f, Gho09h, Gho09d, Gho10b, Gho10a, Gho11d, Gho11b, Gho11a, Gho11c, Gho12a, Gho12c, Gho12f, Gho12h, Gho12d, Gho12g, Gho12b, Gho12e, Gho13b, Gho13c, Gho13a, Gho13d, Gho17, Gho20, Gho22, Gho23, Gil15, Gir73, Goo73, Gow72, Gow79, Gri77, Gun15, GP11b]. Book [GP12, Haa12, Häg07, Häg12b, Häg12a, Háj72, Haj77, HN13b, Ham72a, Ham79, Han07g, Han07d, Han07a, Han07h, Han07b, Han07c, Han07f, Han07e, Han08a, Han08d, Han08o, Han08n, Han08k, Han08e, Han08f, Han08i, Han08c, Han08j, Han08m, Han08b, Han08g, Han08l, Han08h, Han09a, Han09d, Han09b, Han09c, Han10g, Han10f, Han10i, Han10d, Han10c, Han10a, Han10j, Han10e, Han10h, Han10b, Han11k, Han11c, Han11d, Han11h, Han11a, Han11g, Han11j, Han11e, Han11f, Han11i, Han11l, Han11b, Han12b, Han12h, Han12u, Han12q, Han12i, Han12f, Han12d, Han12o, Han12e, Han12j, Han12n, Han12k, Han12r, Han12w, Han12s, Han12t, Han12m, Han12v, Han12x, Han12p, Han12g, Han12c, Han12l, Han13k, Han13j, Han13o, Han13n, Han13c, Han13g, Han13l, Han13b, Han13e, Han13a, Han13d, Han13i, Han13h, Han13f, Han13m, Här08, Has08b, Has08d, Has08a, Has08c, Hau08b]. Book [Hau08a, Hau18, He14, Hee16, Hil15, Hoe74, Hoe83, Hol08, Hor79, Høy72, Huz73, Huz80, Imb07, Ios72, Ios76e, Ios76f, Ios76a, Ios76c, Ios76b, Ios76d, Ios77a, Ios77b, Ios77d, Ios77c, Ios78, Ios79b, Ios79c, Ios79a, Ios80a, Ios80b, Jau14, Kau12, Kau13, Ken80, Ken73b, Ken73a, Ken74, Ken75, Ken76a, Ken76b, Ken76c, Ken10, Kha73, Kha80, Kha74, Kha15, Kha17, Kha20, Kha24, Khu13b, Khu13a, Kim14, Kim15, Kin75a, Kin75b, Kin75c, Kin77b, Kin77a, Kle76, Knü80, Kok07, Kos09a, Kos09b, Kos10b, Kos10a, Kos10c, Kos11, Kos12a, Kos12b, Kos12c, Kos13d, Kos13c, Kos13a, Kos13b, Kri74, Kró80d, Kró80a, Kro80c, Kró80b, Lan18, Lar15, Law07, Lep14, Lid80, Lin13, Lin14b, Lin75a, Lin75b, Lin80a, Lin80b, Lin80c, Lis07c, Lis07a]. Book [Lis07b, Lis08b, Lis08a, Lis09c, Lis09b, Lis09a, Lis10b, Lis10c, Lis10a, Lis12, Lis13, Lis14, Liu07, Liu08, Liu10a, Liu10b, Liu11, Liu12a, Liu12b, Liu14a, Liu14b, Liu15, Liu16, Liu17, Liu18, Liu19, Liu20a, Liu20b, Liu21, Liu22a, Liu22b, Liu23, Loy77, Loy80, Luo07, Luo09, Mai07, Mai08, Mai09c, Mai09b, Mai09a, Mai10c, Mai10b, Mai10a, Mai11d, Mai11b, Mai11a, Mai11c, Mai11e, Mai11f, Mai12a, Mai12b, Mai12c, Mai13c, Mai13a, Mai13b, Mai14b, Mai14a, Mai14c, Mai15, Mai16a, Mai16b, Mai17, Mai18, Mai19, Mai21, Mai22, MM77, Man72, Man75, Man07, Man08, Mar07a, Mar73, Mar07b, Mar75, Mar21, Mat07, Mat73, McC22, McE17, Meh75, Meh15, Men73, Mer07, Mer09c, Mer09d, Mer09b, Mer09a, Mer12a, Mer12b, Mer14, Mer15, Mer17]. Book [Mer18, Mer20c, Mer20a, Mer20b, Moo75, Mor73, Mor80a, Mui73, Muk14a, Muk14b, Muk15, Mus75, Nad74, Nan10b, Nan10a, Nee09, Nee15, Nel72b, Nel80, Ng21, Nor08, Nor09b, Nor09a, Nor09c, Nor13b, Nor13a, Nor14b, Nor14a, Nor15, Num09a, Num09b, Num13a, Num13b, Num13c, Num14, Num15, Num20, Num23, Num19, O’B07a, O’B07b, O’B07c, O’B08a, O’B08d, O’B08c, O’B08b, O’B09h, O’B09e, O’B09a, O’B09i, O’B09c, O’B09d, O’B09f, O’B09g, O’B09b, O’B10e, O’B10b, O’B10a, O’B10c, O’B10d, O’B11a, O’B11c, O’B11b, O’B12h, O’B12g, O’B12e, O’B12d, O’B12f, O’B12a, O’B12b, O’B12i, O’B12c, O’B13a, O’B13d, O’B13g, O’B13e, O’B13h, O’B13c, O’B13f, O’B13b, O’B14a, O’B14b, O’B14c, O’B15a, O’B15b, O’B16a, O’B16b, O’B16c, O’B20, O’B21, Oga73, Oja07, Oja08b, Oja08a, Oja09, Oja11]. Book [Oni79, Oni80, Ord72, Ord80a, Ord80b, Pal72, Par73, Pat73, Pat75b, Pet07, Pin75, Pla75, Ple15, Pod11, Pod14a, Pod14b, Pod15a, Pod15b, Pod16, Pod17a, Pod17b, Pod17c, Pod18, Pod21a, Pod21b, Pod22a, Pod22b, Pod24, Pod72, Pol11b, Pol11a, Pol12a, Pol12b, Pol13a, PS13, Pol13d, Pol13c, Pol13b, Pos76e, Pos76d, Pos76a, Pos76b, Pos76c, Pos78d, Pos78a, Pos78c, Pos78b, Pos78e, Pos79b, Pos79c, Pos79a, Pos79d, Pos80a, Pos80d, Pos80e, Pos80b, Pos80c, Pre77b, Pre77a, Pre78a, Pre78b, Pre79a, Pre79b, Pre80a, Pre80b, Puk07, Pun07b, Pun07a, Pun08d, Pun08b, Pun08a, Pun08c, Pun09, Pun10b, Pun10a, Pun10d, Pun10c, Pun11a, Pun11b, Pun11d, Pun11c, Pun12a, Pun12b, Pun12c, Pun13c, Pun13b, Pun13a, Pyn14b]. Book [Pyn14a, Qiw20, Rad75, Rad77, Rao80a, Rao80b, Rao80c, Rév73, Rév80, Ric09, Ric10b, Ric10c, Ric10a, Ric11c, Ric11a, Ric11b, Ric13c, Ric13a, Ric13b, Ric14, Ric15, Ric16, Ric17a, Ric17b, Ric18, Rob11b, Rob11e, Rob11d, Rob11c, Rob11h, Rob11g, Rob11f, Rob12a, Rob12b, Rob19, Rod21, Roh76a, Roh76b, Roh78, Roh80, Ros09a, Ros09c, Ros09b, Ros10b, Ros10c, Ros10a, Ros14, Ros15a, Ros15b, Ros19, RR79b, RR79c, RR79a, Rou07, Roy14, Rui72, Rui16, Rui17, Rus10, Rus74, Sai08, San15, San17, San77, Sar73, Sär72b, Sat15, Sch80a, Sch80b, Sch16a, Sch72b, Sch72c, Sco77b, Sco77a, Sco07, Sco09b, Sco09a, Sco10b, Sco10a, Sco12b, Sco13c, Sco13b, Sco13a, Sen11c, Sen11a, Sen11b, Sen14d, Sen14c, Sen14b, Sen15, Sen23, Sil73]. Book [Sko12, Slo23, Slo24, Smi75, Smi77, Smi80, Som76c, Som76a, Som76b, Som77b, Som77a, Som78g, Som78d, Som78e, Som78f, Som78a, Som78b, Som78c, Som79e, Som79a, Som79b, Som79d, Som79c, Som80a, Som80c, Som80b, Spe07, Spe08b, Spe08a, Spe10e, Spe10a, Spe10b, Spe10c, Spe10d, Spe11b, Spe11a, Spr75a, Spr75b, Spr77b, Spr77a, Spr80, Sta07a, Sta07g, Sta07f, Sta07b, Sta07e, Sta07c, Sta07d, Sta08a, Sta08c, Sta08b, Sta09, Sta10a, Sta10c, Sta10b, Sta11a, Sta11b, Sta12a, Sta12d, Sta12e, Sta12f, Sta12c, Sta12b, Sta13a, Sta13b, Sta13c, Sto75, Str76, Str73, Stu72, Stu80, Sun15, Sun16a, Sun16b, Sun16c, Sun17, Sun18, Sun19, Sun22, Sun24, Szc11a, Szc11b, Tan07, Tha07a, Tha07b, Tin78e, Tin78b, Tin78c, Tin78d, Tin78a, Tin80a, Tin80b, Tip77, Tip75b, Tip75a, Tip80, Tra07]. Book [Tra09, Tra13, Tra73, Tso16, Tso17, Tso18, Tso20, Unw09, Unw10a, Unw10b, Unw11b, Unw11a, Unw11c, Unw12c, Unw12a, Unw12b, Unw12d, Unw13b, Unw13a, Unw20, Unw21, Unw24, Val08, Var14, Var19, Var20, Veh07a, Veh07b, Veh08b, Veh08g, Veh08c, Veh08f, Veh08d, Veh08a, Veh08e, Veh08h, Veh09a, Veh09b, Veh10a, Veh10d, Veh10b, Veh10c, Veh11b, Veh11a, Veh13b, Veh13a, Veh14, Vin73, Voe09, Vos75, Wet75, Wet80, Wil73, Wim75, Yan73, You14, You15, You19, You20, Zha14, Zha16a, Zha16b, Zha16c, Zha17a, Zha17b, Zha18, Zha19b, Zha21, dO79a, dO79b, vdL80c, vdL80a, vdL80b, vdL80d, Som80a, Liu10b, Sta07g]. Books [Ano75k]. Bootstrap [BP12, Han13d, HHK03, O’B13a, Qiw20, SH23, dAY92]. Bootstraps [WPRP17]. Borcard [O’B12h]. Borenstein [Ros09b]. Borgan [Aal82, CV82, Duf82, Gho09h, Har82, LT82]. Borin [Veh10c]. Boris [Gal80a]. Borkar [Cro09e]. Bortkiewicz [HV15, QS87]. Bos [Cro07b]. Bose [Mer20c, Mer20b, Ald99]. both [dF74]. Boullion [Tip77]. Bound [Ano93d, Bic92]. Bounding [GS02]. Bounds [HS90, Nai93]. Bousquet [Dur22]. Bouveyron [Unw20]. Bouza [Esp14a]. Bouza-Herrera [Esp14a]. Box [Cox07, Smi75, Tin78e, BS85, BCO82, Dra13a, NSW72, San93, Dra13a]. Boyd [Kok07]. Brad [Cox20]. Bradford [Kró80a]. Bradley [Dra12d, FHT20, Spe10d, Spe11a, Sun19, gXZ20, Nar20a]. Brady [Cro10c]. Brain [Dra12f, LEGW01]. Branchements [RR90]. Branching [Dur15b, Rad75, Rad77, Tso16]. Brandeis [Gho23]. Brani [O’B13h, Oja08b]. Branscum [Han11a]. Brasov [Gan80c]. Braun [Nor09a, Sta07a]. Brazzale [Cro07a]. Break [Kir05]. Bremner [Vin73]. Brendan [Unw20]. Bretz [Ric11b]. Brian [Dra13b, Eva07a, Han09a, Kau13, Mai17, Num13b, Pod22a, Pun10a, Ric13a, Rob11c, Unw12a, Veh10d]. Bridge [Dur13a]. Bridging [Cam07, KMT03]. Brief [AM14, Hun14, LCZ14a, Rob14, Sik18, Tsa24]. Brillinger [Cox12, Cox12]. British [All72, MDS94, WB97]. Brodeau [Pre78b]. Broe [Rob11g]. Broemeling [O’B10b, Zha14]. Brooks [Pol12a, Gho11b]. Brown [Bar72c, Fis24, Kró80d, Pol13b, Veh07b, Hal73, Mer09a, Tip77]. Brownian [IK86]. Bruce [Gho13c, Pol13b, Pun08b, Han07a]. Brumback [Gho22]. Brunk [Vin73]. Bruno [Mai13b, Mar75, O’B12b]. Bruti [PS13]. Bruti-Liberati [PS13]. Bryan [Els18]. Bryman [Sta12d]. Bubbles [Han12s]. Buchen [Jau14]. Bueno [Dek22]. Buffers [Gan78c]. Bühlmann [Gho12g]. Build [Til22]. Builder [OZW23]. Building [Cra73, Eng78, Han12f, SDGD^{+}11b]. Bulgaru [Oni78]. Bulk [Ios76d]. Buonaccorsi [Cro10h]. Burden [Ast94, Sun77, dVKW96]. Bureau [Pow03, TC15, LH04]. Burger [Spe10e]. Burkhardt [Ano73e]. Burr [Tad80]. Burrington [Der75b]. Burton [Mai11f]. Burzykowski [Unw21, Mai14c]. Business [Abr05, CKS17, Dra08b, FPD05, HHDG12, Han07a, Han10i, Han10a, Has72, HL01, Ich09, KRR^{+}11, KCB00, NP97, Pod15a, Pol13c, Pos79d, Riv02, Sco13a, Tip77, Tsi03, dWC05, Ana21]. Business-Pragmatic [Pod15a]. Byrne [Veh14]. Byron [Gho11b, Kró80d].

C [All79a, Aoy72, Bar75d, Boo79, Cau80, Cha80, Cox15, Cro07a, Cro10j, Cro10g, Dal72, Dra08b, Dra10c, Els18, Fil11, Fin77b, FvZ15, Gan80d, Gay11, Gho09d, Goo78, Han07a, Han12r, HR21, Ios76d, Khu13a, Law07, Lis10b, Liu20a, Loy77, MM77, Moo78, O’B13f, Oja08a, Ord80a, Ord80b, Pod22a, Pos76b, Pos78e, Pos79b, Pos80a, Pos80d, Pre77a, Pre78a, Pun07a, Pun08b, Pun11b, Pun13b, Pyn14a, Ric11c, Roh76b, Ros14, Sch73, Sco77b, Sco12b, Sco13c, Sen11c, Sen14d, Smi75, Som78a, Spe07, Spe11b, Sta13a, Tip77, Tip75a, Veh08d, Wil73, vdL80c]. C. [Hir05, KK21]. C48.14 [Pod17a]. Cadusch [Gho13c]. Caithamer [Pod11]. Calculating [wB09, Han11b]. Calculations [Bel00, Nee15, O’B08c, AHZK15]. Calculator [Sta12e, Sta12e]. Calculus [Han12b, Liu20b, Mer12b]. Calibrated [GT17, SS07]. Calibrating [Han08h]. Calibration [DTP^{+}20, ES06, Ham72a, HS88, KP10, LSD^{+}11, Osb91, OH19, WL16, ZHW23]. Call [Had10a, Had10b]. Calvin [Spe07]. Cambridge [Dra11e, Dur18, Dur23b, Liu19, Liu22a, Mer18, Pod18, Pod22a, Rob19, Sun19, Unw20, You20, And77b]. Camouflaging [KMNS05]. Campagnoli [O’B10c]. Campolongo [Liu08]. Can [Dra13e, God97, HMW21]. Canada [SHGK01]. Canadian [Dra14d, Liu10b, Fel73, GP03, HL01]. Canadiana [Cop77]. Cancer [Dur15b, Kim14, LHSD15, MM96, KT93, Tso16]. Candidates [OS93]. Cani [Tip77]. Canonical [Ham72b]. Canossa [Liu22b]. Cantelli [BBS07]. Capability [Ani08, Uni78]. Capacity [SDGD^{+}11b]. Capital [Har99, Has72]. Capture [HH11, LdBP90]. Capture-Recapture [HH11]. Capture/Recapture [LdBP90]. Car [And73b]. Card [Mai13a]. Carey [Han07h]. Cariboni [Liu08]. Carl [AK95, Für76, Lid80, O’B09f, Spe10d, Kul78]. Carl-Erik [Kul78]. Carlo [vC75, Don12, Han12t, Nan10b, Pod17b, Roy14, Sco10a, Tra09, Tso17, Tso18, BO15, Gob16, JLS20, KLTG21, RR21, She00, Pol12a]. Carlos [Esp14a]. Carolina [NGC^{+}78]. Carolyn [Spe08a]. Carpenter [Nor14b]. Carr [Unw11c]. Carter [Bil77a, LLT04, LQ04]. Carton [Ios77b]. Case [BCG07, BS92, Cro07a, Don12, Dra12d, Gar97, Gho97, GL19, Han12d, Haw97b, JSW11, KC14a, Ket97b, LvR74, Mai18, MA05, Mer07, Moo97a, Moo97b, O’B09a, OMGT05, PPMS11, Sch97, Wil95, lRG05, Aoy72, O’B07c, Sat15, Som78a]. Case-Control [JSW11, KC14a, Sat15]. Casebook [Buc77d]. Casella [Cro08a, Pun08d, Sco10a]. Cases [GMZ03]. Cassel [Pos79b]. Castillo [Bol07]. Casuistry [Bel88]. Catalogue [CSW93]. Catania [Dur13b]. Catastrophe [Han11b]. Categorical [AN01, Bar77c, Dar81, HS88, HK77, IKS^{+}81, IKS^{+}82, Kos10c, Lin14b, Lis07a, Lis07b, Liu16, Liu23, Num14, Num15, PCA^{+}19, PPMS11, Sär11, Sim98, Ski19, SM14, Stu80, Unw10b, Unw12d, Sch83, WLC82]. Categorized [AP81]. Catt [Tin80b]. Cattell [Bar80b]. Causal [Blu18, BD20, BC20, Daw02a, Daw02b, DTP^{+}20, Fro06, Gho22, Han11f, Imb07, Pea02a, RYG^{+}21, Sin02a, Spe10b, ZMSL19, ZHW23, ZZLZ15, Blu16, Tso20]. Causality [CW04, Fro06, Hec08, Pea02a, Sin02a, Slo23, Gho11c, Roy14]. Causation [Lin02, Sin02a]. Cecrops [Mis10]. Celebrate [Spe08a]. Celebrating [GG17]. Celebration [CWCT19]. Celeux [Unw20]. Cell [LNA75, Tal93]. Censored [ABGK82a, ABGK82b, Ano84a, DS22, Sch84b]. Censoring [Nai93]. Census [BCM19, CT11, Che85a, Dek94, Kir82, LH04, MDS94, Mis10, Red01, SCZZ99, Shl07, Som78g, Tam82, TG94, Yim90, BKV13]. Censuses [Ano78k, Kis79, LTS98, NSS73, Red74, Red01]. Central [AP94, Han12s, LN74, NL75]. Centrality [SG77]. Centre [Hen78]. Centuries [Bel88, Rob12b, Sen05]. Century [Bor09, KK21, Sti88, Dra13b, Mar21]. Cernat [Blu21]. Certain [BKBP21, Nel72b]. Cesar [Dur14a]. Château [Han13g]. Chacko [Elf72, Yan73]. Chain [KLTG21, Pol12a, RR21, Sen96, She00, Ney75]. Chains [Cro09d, Haj77, Ios76d, Mai14a]. Challenge [Mal05]. Challenges [BM16, Car05, Fel04, GG99, Gho10b, Mac97, Mul76, Wyc03, Mal97, Rod21, Veh07b, Zha18]. Cham [Dur22]. Chambers [Han13c, Tra13, Cha16, Wet80]. Chan [Mai09c, Ros14]. Chance [Bel89, BF97, Bel02, Cox02, Ios79a, Sin02c, Sin02d, Spi02, dM84, Hal84a, Pun10a]. Chancy [Rao19]. Chandler [Han12n, Mai17]. Chandra [Ano72j, Rao73]. Chang [Cro08a, Don12, Dra13c, Han12j, O’B07a, O’B08a, O’B12a]. Chang-Xing [Cro08a]. Change [BE17, Gal02b, GT99, Hoe99, HS89, O’N04, Sha80, Sie88, Han10b, Tin78d]. Change-Point [Sie88]. change-points [Han10b]. Changements [Mal97]. Changepoint [Sha23]. changes [Mal97]. Changhua [Kha17]. Channel [Ovu80]. Chanseok [Han13n]. Chaoqun [Han08d]. chaos [GT12]. Chapman [BPV21, Dur23a, Fis22, Fis24, Gho22, Kha24, Liu18, Liu21, Liu23, Mai18, Mai22, O’B21, Pod17b, Pod17c, Pod21a, Pod22b, Ric17b, Slo23, Sun16a, Sun18, Sun24, Tso18, Unw21, Unw24, Zha17a, Zha17b, Zha18, Zha19b]. Character [DW84]. Characterisations [Ken80]. Characteristic [CMH19, MS87]. Characteristics [Bon96, DPCM21, Fel96, HG94, MDS^{+}19, NSS23, Nor96, Smi92, Wal96]. Characterizations [Kot74]. Characters [RR79a]. Charles [Cro11e, Dra13e, O’B12e, Pun09, Som76b, Unw20]. Charlesworth [Ken76a, Ken76a]. Charlier [Hal00]. Charmed [Pun11a]. Charming [Han11c]. Chart [EPP08]. Chatfield [Loy77, Som80c]. Chatterjee [Pun13c, Pun13a]. Chaudhuri [Sen84, Tam84, Tho84]. Chaumont [Pod14b]. Chemistry [Pre77b]. Chen [Num13a, Pol11a, Ric10b, Veh08e]. Chernick [O’B13a]. Chester [Gre80]. Chi [HK77, Ord72, Pla83]. Chi-Square [HK77]. Chi-Squared [Ord72, Pla83]. Chief [Cha16, PA15, Pin10a, SZ09]. Chihara [Sco12b]. Child [Meh15, Ver13]. Childhood [Gar93]. Children [KMM^{+}17]. Chin [Han12g]. China [Che85a, SCZZ99, Wei05, ZgZ89]. Chinese [She06]. Chipman [Pun12a]. Chistyakov [Gal80a]. Chiswick [Bre80]. Chitturi [O’B12d]. Choice [DO73, LLP01, O’B12d]. Choice-Based [O’B12d]. Choix [Lef79]. Cholesky [KDTP20]. Cholesterol [GOZV94]. Choose [Blu05, CD18]. Choosing [GS02, Lef79]. Chosen [Gue08, Tho78b]. Choudhary [Num19]. Chow [Gho09f, O’B07a, O’B08c, O’B12a, Han12e, O’B08d]. Chris [Cro11a, Ric13c, HS19]. Christensen [Han11a, AK95, Pun11c, Pun13b].

Christian [Cha14, Emb08, Gho12h, Liu11, Mar07b, O’B09f, Pun07a, Sco09b, Sco10a]. Christie [Häg12a]. Christl [Han12k]. Christodoulopoulos [Val73]. Christoph [Cro11h, Sch80b]. Christopher [GP12, Som80c]. Christos [Pod21b]. Christy [Sta07e]. Chronicle [Ber14, Ros15a]. Chronological [Sow72]. Chuan [Kos13c]. Chuang [Sta07e]. Chuang-Stein [Sta07e]. Chun [Sun22, Veh08e]. Chun-houh [Veh08e]. Chung [Gho09f, Han12e, O’B07a, O’B08d, O’B08c, O’B12a]. cinq [Fou80b]. Cioppa [Cam07]. Ciorciari [Gho13c]. Cira [Sta12b]. Circuit [Fri93]. Circulant [Mer20c, Mer20b]. Circular [PNR13, Liu15]. Circulation [Koe79]. Citizen [Ole03]. Citizens [Ott02]. Cízek [Kos12c]. Claassen [Zha19b]. Claeskens [Lis09c]. Clarifications [KG10]. Clark [Ano21, Cro11e, Han13c, Ric10a, Spr75a]. Clarke [Mai11d]. Class [Gho10a, Jas15, Pat75b]. Classes [Gou94, HB07]. Classical [Mer18, Pal72, SFG10, Hol86]. Classification [Ahn14, Cia14, Dur13b, FBL18, GM19, Han10g, Han12a, Loh14a, Loh14b, NAL76, Pun08c, RZ14, SFVP^{+}24, SZ14, Str14, Sun77, TRRL23, Unw20]. Classifications [Pos78c]. Classified [SR92, Sow72, Sow78, Stu80]. Clatworthy [Pat75b]. Claude [Mai09b]. Claudi [Han11c, O’B10e]. Claudia [Dur13c]. Claudio [Aur10]. Claus [Häg12b, Sko12]. Clearing [CS00]. Clearly [Dek22]. Clelland [Tip77]. Cleophas [Ros09c, Sta12e, Ros09c]. Cliffe [Häg12a]. Climate [Dur23b, O’N04]. Clinical [Arm85a, Arm85b, BCL^{+}85, Bat85b, Bra14, Cam07, Che85b, Cro10f, ET16, FWC^{+}21a, FWC^{+}21b, Han08e, Han12e, Han13l, Hir93, Kim14, Lou85, Nee15, O’B07a, O’B08c, O’B12a, Pet85, Ros09c, Ros10c, Sch16a, Sta12e, Til22, Tuk91, Whi85, Zha19a, Zha16c, Zha17b, Zha18, Zha21, AHZK15, Muk14a, Nee09, Num13a]. Clinicians [Tha07b]. Clive [Unw24]. clumps [Nau79]. Cluster [Ber72, EUC09, Häg07, Lin75a, NKH91, TSR96, Cro09a]. Cluster- [Ber72]. Cluster-Analyse [Lin75a]. Cluster-Specific [NKH91]. Clustered [AN01, Dem07, Kos10c, Nee15, PGN^{+}96, TO17, YK20, AHZK15]. Clustering [FMR19, Han08d, Han10g, HT92, Har73, Men16a, Oja07, Unw20, Cro11f, Men73]. clusters [Nau79, O’B09g]. Co [O’B14a, Cha16, PA15, Pin10a, Pin10b]. co-Editor-in-Chief [Cha16, PA15]. Co-Editors-in-Chief [Pin10a]. Co-EditorsinChief [Pin10b]. Co-evolution [O’B14a]. Coale [Hoe74]. Coarsened [PCA^{+}19, ZZLZ15]. Cochran [Dem83]. Code [Dem72, Rob12b]. Coding [Lee01]. Cody [Khu13b, Sta07d]. Coefficient [AS72, BR07, EUC09, FZ15, Hon15, HS15, PMLL15a, PMLL15b, WY02, Don86]. Coefficients [MNS77, NW06]. Cognitive [GW14a, GW14b, Hub14, Wil14a]. Cohen [Mai10c, Han08c, Mai10c]. Cohesion [Teu05b]. Cohort [O’B15c, San15]. Cohorts [MBW^{+}23, WGKL22]. coincidences [Nau79]. Coincident [Nie04]. Cointegrated [LD24, Sco09a]. Coke [RM93]. Cole [O’B21]. Coleman [Dra08b, Ios76e]. Colin [Spe10e, Sta12a, DL13]. Collaboration [Fis05, Sti05]. Collapse [Dra13e]. Collapsibility [GP11a]. Collection [Fin75, Hau01, JRL10, Knü80, WD84]. collector [Hol86]. Collier [Spe10c]. Collinearity [GPdMLMGGSG20, GSGdMLM16]. Collins [Gho10a]. Color [Pod24]. Colosimo [Bol07]. Combination [HS88]. combinations [PMM87]. Combinatorial [Pat73, Liu14a]. Combinatorics [Smi80]. Combined [MBW^{+}23]. Combining [Cro11d, FMR^{+}20, GL14, Han08h, KT21, KMS14, SKK18, Tho19, TH98]. Comes [Mer12a]. Coming [VJ95]. Comment [Bel02, Cox02, Cre20, Hoy76, Hum77, Mal97, Ott76, PP98, Spi02, Sch78, Sin02a]. Commentary [SJ89, And77a]. Comments [Gan72b, Pea02a, Pea02b, RvA72, SB04]. Committee [DD80, FMS95, MDG^{+}92]. Common [Dra11d, Mer07, Pro11, Smi20, Tit85]. Commonwealth [CS00]. Communicating [McC22]. Communication [GP03]. Communications [Pow03, San77, vdL80d]. Community [Dem77, LLP01]. comonotonic [SGM21]. Comorbidity [KMM^{+}17]. compact [AC21]. Companies [Han10a]. Companion [Kau13]. Company [Spe08b]. Comparability [Dal98]. Comparative [BDS04, Cor77c, CR89, Dra10b, FBGGM17, HM04, PSI19, Wan19, Wet75]. Comparing [BP92, BBvdB18, SBB^{+}12, SG77, Han13d]. Comparison [ABGK82a, ABGK82b, Ano84a, Bic14, CFL^{+}21, CW06, CKS17, DH19, DS99, DMS06, DS78, FWS14, FST19, GNM21, HK18, HGKJ20, Hum77, KSS14, KY91, LNA75, LCL^{+}23, MN98, Mod20, NAL76, NKH91, RSG06, Rob11b, SFG10, SCW08, Smi92, Som78g, SAA22, WD84, YYY20]. Comparisons [Kha80, LG08, Ric11b, SC05]. Compartment [TZW^{+}20]. Competing [Cox80, Cro94, JK78, O’B16a]. Competitions [Liu10b]. Competitive [Vov01]. Complete [Ano76m, HMA^{+}18]. Complex [BDLP20, BD94, BO15, CG12, Cro10i, ES06, FFBK76, Han08o, Han10a, MSS01, Pun11c, dMRAMR18, Sax78, She00, Ski19, WS21, WL16, ZW19, Bin83, KFF75]. Complexity [CQOR^{+}22, Häg12a, Lis07c]. Component [CD18, CD00, FBGGM17, Kha17, Ott75a, Ott76, Vid15, Zha16b]. Components [Ahl02, Gal02a, HS89, Khu00, Sah79, Wel02, KS85]. composite [Bic13, O’B12e]. Compstat [Cha80, Cha80]. Computation [Jas15, O’B11a, Pos76d, Rob11h, Tri09, Dal72, Han09d, Ken10, Ros14]. Computational [Cha80, Cha15, FG72, FST19, Gho08g, Kos13c, RDMP16, ZYW23, Dur15a, Kos13a, Kos13b, O’B09b, O’B14a, Pod21a, Wet80]. Computations [Lee79, Pun07b, Tho78b]. Computer [Cle05, Dek94, Fin77b, Ken76b, Mai07, Mai10b, Mor80a, Nov73, O’B10a, Rip83, ST07, Sch07, Vos75, WD84, Lin75b, Sun19, Yan73]. Computer-Aided [Sch07, Yan73]. Computer-Assisted [Nov73]. Computer-Based [ST07]. Computers [FB73, Mul76, Nel80]. Computing [And73c, Cam97, Esp15b, Gow15, Sco13c, Unw12b, Lon15, Pun10b, Unw09]. Concentration [LV14]. Concept [Hor72, Lin02, Ruk95, Sin02a, KM80, Rém84]. Conceptions [WPRP17]. Concepts [Cro10d, Don12, ET16, Ham72a, Han07a, Han12g, Han12l, Har99, Lis13, Lis14, Mai18, NP97, Pos80b, TC72, Veh14, Zha17b, Han12c, Som78b, Tip77]. Conceptual [FBL18, Fro06]. concernant [Ios80b]. Concerning [KSS^{+}03b, Ber03, Cox03, KSS^{+}03a, LS86, LN77, LN78, Sch03]. Concise [Liu20a, Ric14, Var20, Veh08h]. Concordance [Lov00]. Condition [MV22]. Conditional [BK16, DGS18, HH11, LHSD15, O’N04, SSOL04, Sha85a, Sha85b, Til98, WC98, BH77]. Conditioned [You86]. conditioning [Kal84, Pod14b]. Conduct [Dem72]. Conference [Ano05a, Deb78, Dur13b, FL05, Gan72a, Gan80c, Ios79b]. Conferences [And77b]. Confidence [Bic14, CMH19, Cox13, DR90, Efr13, EGS72, Fra13, Han09c, Han12w, Kab09, Le 90b, LLP08, NN20a, NN20b, PS20, Par13, Rob13, SH13, SB17, Sie88, gXS13, gX13, Roy86, Sco12a, Lis09b]. Confidential [Rei09]. Confidentiality [Han12m, Mar74, OC13]. Conformity [HTW01]. Confounding [NSP21]. confronting [Dag76]. Congruential [EH92]. Conjectandi [Syl14]. conjecture [LS86]. Conjoint [O’B12d, RW09]. Connecting [Jon07]. Connections [LSD^{+}11, RR90]. Conover [Stu72]. Conscience [Gho23]. Consequences [GP11a, HMW21]. Consideration [LS04]. Considerations [Coh05, Fri93]. Consistency [FH89]. Consistent [Bha91, Feu93, KBM21, WvdH06]. Constance [Sta10b]. Constrained [GS19, Pan07, Piz10, Hoi13, Sha88]. Constraints [O’B16b, Pro11, RW09, Sri80, Man75]. Construct [Mon01]. Constructing [NN20a, NN20b, SLWH22]. Construction [Gho08b, HB07, LLP08, RR90, RR90]. constructions [Dru11, Pat73]. Constructive [SDT^{+}23]. Consumer [Dal98, Fis72, Smi21, Whi99]. Consumption [BCG07, Has72, OS73, WD84]. Contact [SD11]. Containing [Bel00]. Contemporary [CWCT19]. Content [Gar97, Gho97, Han07b, Haw97b, Ket97b, Moo97a, Moo97b, Sch97]. Context [Fer94, Red74, Smi72, Sti14, Ros10a]. Contexts [Sen09]. Contiguous [Tan92]. Contingence [HL74]. Contingency [Csi80, GP73, Goo86b, And87, GS87, Goo86a, HL74, KZ76, LHK78]. Continuation [Ano75k]. Continuity [Hab82]. Continuous [Buc74, CW73, EM19, HH95, Mer24, Num15, Rao80b, SM14, Wil95, VLM12]. Contrasts [GP73]. Contributed [BBG^{+}76]. Contribution [Dam74, Hol84, SJ89, Dam74]. Contributions [CWCT19, FT96, For82, Hal81, Pun07a, Dur13c, Gho13c, Lau81, Hor79, Tin78c, Huz80, Veh07b]. Control [Bol07, Che85a, EPP08, FHTV93, Fri93, Frö08, Han12p, Har96, Ich09, JSW11, Mus75, NA16, NZY18, O’B11c, Roh16, Shl07, Yas93, vdHvdH02, KC14a, Els20, Kró80b, Sat15]. Controlled [AS73, IMG22, RN92, Sun24]. Controlling [Pan07]. Controversial [Han12e]. Controversies [Kró80b]. Controversy [LSS05, LS06, NLSS08, Rob12b]. Conurbations [BJO^{+}98]. Convergence [BF91, Hal89, Mat73, Sch72b]. Conversation [Can95, Cit16, Cox15, Fis10b, GN14, GN20, Gen24, He17, Sap15, Sch95, Sin95, Spe10a, Wil16b, DL13]. Convex [Gho20]. Convolution [RDM^{+}19, vdHK99]. Cook [Han08n, Mai08]. Cooley [Gow72]. Cooper [Han12s]. Cooperation [ANK03]. Coordinate [IOS12]. Coordinating [NPT08]. Cope [Bie99, LA96]. Copper [Pos73]. Copula [DM05, Tso17, Kos12a]. Copulas [DM05, Dur13c, HCR18]. Copyright [BS03]. Corbelled [LKC^{+}04]. Corder [Ric10c]. Core [Han12u]. Cormack [Fer73]. Cornell [Dra12b]. Cornillon [Unw13b]. Cornuejols [Feh07]. Coronavirus [LC20]. Corrado [GG17]. Corrected [DNC06]. Correction [Ano93d, CP08, Eid07, Hab82, NN20a]. Corrections [Ano80e]. Correlated [AJS81, Dur18, FRS08, MR10, NKH91, SCC14]. Correlates [NSS73]. Correlation [Bly94, EFP19, EUC09, Ma17, McL93, Mor73, O’N09, RC93, Don86, LL87]. correlations [LN78, TTB86]. Correspondence [Beh04, BL24, DG90, GR90, Goo86b, Sch78, And87, GS87, Goo86a, Sch83, Fis18, Sta08a, Veh08b]. Correspondent [ABGK82a, Ano84a, Atk82, Bar83, Dar81, GH87, Hey81, JK82, Kei87, New81, New84, New88, PS85, Rip84, SS88]. Corrigenda [Ano96d, Daw02a]. Corrigendum [Ano79e]. Corsten [Cha80]. Corten [O’B14a]. Coryn [Rob19]. Cost [CR93, Fou80a, Der77a]. Costs [LTS98]. could [Mal97]. Count [CHXY19, HPHdAM22, He14, MR10, Num14]. Counterterrorism [Unw10a]. Counting [ABGK82a, ABGK82b, Ano84a, Ham78]. Countries [Azo01, DH19, Fer94, Jab96, KP04, Kis96, Mul76, OS73, Ver96, Hen78]. Country [MA05]. Counts [BDS04, SP11, SR92]. Course [Cro10a, Has08d, Ios76a, Ios80a, Kos10a, Mai14a, Mar73, Moo05, Nor09a, Qiw20, Rui17, Sco13b, Sen14b, Som80c, Sta10a, Wil95, Bre79a, Han12u, Han12l, Pos79c, Rus74]. Courtroom [Kay03]. Courts [Ize03]. Covariance [DPCM21, GNM21, Pou02, ŞA09, KDTP20, PWKO24]. Covariate [IMG22, KT03, Kuk20, RJ91]. Covariates [DMMG22, FGFBGM22, GTF18, Sar08, Tuk91]. Coverage [Kab09, KL09, Sco12a]. COVID [HRR20, SLY23]. COVID-19 [HRR20, SLY23]. Cowell [Gho08g]. Cowpertwait [Mai10a]. Cox [Han12k, Pos76e, Atk84, BG83, BS85, Box84, BCO82, Coc79, Cox03, Dar84, Joh83, Mic19, Spj84, Wah84]. Cp [DNC06]. Cracked [Rob12b]. Cracks [Bel02, Cox02, Sin02c, Sin02d, Spi02]. Cramer [Sta12d, Lea88, Sch84b]. Crawley [Dur14b, Sta07g]. CRC [Dur23a, Fis19, Fis22, Fis24, Gho22, Kha24, Liu18, Liu20a, Liu21, Liu23, Mai16b, Mai18, Mai19, Mai22, Mer20c, Mer20b, Ng21, O’B21, Pod17b, Pod17c, Pod24, Qiw20, Ric17a, Ric17b, Rui17, San17, Sen23, Slo23, Sun16a, Sun17, Sun18, Sun24, Tso17, Tso18, Unw21, Unw24, Zha17a, Zha17b, Zha18, Zha19b, BPV21, Pod21a, Pod22b]. Create [Sne99b]. Creation [PP98, Pul97]. Creative [Mer07, Mer09b, Pun11a]. Credible [PW91]. Credit [Han12s]. Cressie [BL24]. Crewther [Gho13c]. Crises [Cro94, Han12s]. crisis [Ney75]. Cristina [Dek14]. Criteria [Lis08b, WLC82]. Criterion [BCF^{+}14, FWS14, Ros87, BG83]. Critical [CH19, FGFBGM22, Iac82, MP21]. Criticality [BJS92, Sen98]. Critique [Tho06]. Croon [Kos10c]. Crop [MW81]. Cross [Bar80a, HK77, PS85, Pos78c, Pro11, RC93, SR92, Stu80]. Cross-Classified [SR92, Stu80]. Cross-Correlation [RC93]. Cross-Section [PS85].

Cross-Sectional [HK77, Pro11]. Cross-Tabulated [Bar80a]. Crossing [Haa12, LR91]. Crossover [Mat88]. Crowded [Hor16]. Crowdsourcing [SFVP^{+}24]. Crucial [Fis01]. Cryer [Mai09c]. Cryptic [Ber03, Cox03, KSS^{+}03b, KSS^{+}03a, Sch03]. Csicsery [Spe10a]. CUB [PSI19]. Cube [Mer09d]. Cudeck [Veh08d]. Cui [Ng21]. Culture [YB20]. Cumberland [Sti88]. Cumming [Han12w]. Cumulants [Hal00, wB09, Sko86a]. Cumulative [PSI19]. Currency [Koe79]. Current [Cam97, Cox97a, Cox97b, ESTTD05, GG99, GBZ07, KM79c, ME93, Tam82, Teu97, Pyn14a, Zha16a]. Currently [Mas73]. Curriculum [Tam07]. Curt [Han08e]. Curve [Bly94, CMH19, GMZ03, O’B15a]. Curves [BM16, CG12, MW10, Dra12a]. Customer [Sco13a]. Cutsem [Pre78b]. Cyber [PGV22]. Cycle [BJPT99, Kir05, Tsi03]. Czado [Dur13c]. Czech [MA05].

D [Ana21, Bar77b, Buc76a, Buc76b, Buc77b, Buc78, Cap19, Cop77, Csi80, Dow72, Dra11d, Dur18, Esp14b, Fis19, Fon11, Gan72a, Gho08a, Gho08h, Gho10b, Haj77, Ham79, Han08e, Han09b, Han10f, Han12k, Han13d, Ios77b, Ios77d, Ios80b, Ken76a, Ken76c, Kha73, Kos10a, Liu22b, Mai09c, Mai12c, Men73, O’B10b, O’B12e, O’B12f, Pat73, Pod14a, Pod22a, Pos76e, Pos76c, Pos78c, Pos79d, Pos80a, Pos80e, Pos80c, Pre78a, Sco13b, Som78g, Som78c, Sta07c, Sta12a, Sto75, Sun17, Sun22, Szc11b, Tha07a, Tin78e, Tin78d, Tin80b, Vin73, Zha14, Zha19b, BBS07, MCF12]. D. [AG81, Atk84, Bar82, Ber82, Box84, BC81, Cox81, Cox03, Dar84, Dei83, Din82, Eng83, Gra83, Joh81, Kei81, Klo83, Miz83, New83, Sha82, Sil82, Smi82, Spj84, Wah84, Zad82]. D’Addario [Bar75a]. Daejon [Ano05a, FL05]. Daeyoung [Dur21]. Dagnelie [Lin80a]. D’Agostino [Sta07e]. Daily [MFAG10]. Dale [Ric10c]. Dalenius [Sch80a]. Daley [Val08]. Dalgaard [Nor09b]. Dallas [Dra09a]. Damages [BS03]. Damaraju [O’B12d]. Damon [Rod21]. Dan [Ano21]. Dani [Pod24, Ros19]. Danica [Han13l]. Daniel [Cro11e, Han07b, Has08c, Liu18, O’B12h, Sco13a, Unw10b, Unw11c, Pre77a]. Daniela [Nor14a]. Dankmar [O’B09e]. dans [HL74, Lef79, Sau76]. Dario [Ros10b]. Dark [Far13, Sti88]. Darlene [Cro11d]. Darwiche [Han12i]. Daryl [Gho09e, Val08]. DasGupta [Gho12d, Han13k, Oja09, Som76a]. Data [AP81, AN01, Alv11, Alv14, ABGK82a, ABGK82b, And73c, Ano79e, Ano84a, BCM19, Bau14, Ber17, BH14, BEW03, BRST76, BCD^{+}94, Bre75, CB00, CHXY19, CMH19, Cle01, Cle05, Cre20, Cro09b, Cro10c, DMS06, Dek94, Dek22, Dem07, DGS18, Dra09a, DS22, Dur13b, EM93, EKIKW77, Ell91, EPP08, Esp14a, FRS08, Fin75, Fis24, FMR19, FPR^{+}16, FFBK76, Fri01, Gel03, Gho12f, Gho12g, Gho17, Gil15, GDSN18, GTF18, GHK20, GNM21, GC95, Gow72, Gre06, GW14a, GW14b, HL19, Han07c, Han07e, Han08o, Han08c, Han10f, Han11a, Han11l, Han12h, Han12f, Han13a, HM04, HM19, HPHdAM22, Hau01, HK77, Hjo92, Hub14, IKS^{+}81, IKS^{+}82, JRL10, Kau12, Ken76b, Ken10, Kha20, KW19, KT21, KFL76, Koc14, KC14b, Law98, LW13, LdBP90, Lep14, LA96, LNA75, LLT04, Lin14b, Lis07a, Lis07b]. Data [Liu16, Liu18, Liu19, Liu22b, Loy80, LSD^{+}11, LQ04, Mac97, MYA95, Mai08, Mai10c, Mai11d, Mai12a, Mai13b, Mai14b, Mai16a, Mai17, MDS94, Mar11, McC22, MBW^{+}23, Mis08, MSS01, Muk14a, Muk15, NAL76, NKH91, Nor08, Nor09c, Nor15, NWS23, Nov73, Num13a, Num14, O’B09e, O’B10d, O’B12b, O’B15c, OS73, PZ21, PGN^{+}96, Per14, Pfe93, PSI19, Pla88, Pod14a, Pod15b, Pod17a, Pod21b, Pod24, Pol11b, PPMS11, PAF18, Pre80a, QS87, Red01, Ric10b, Rid16, dMRAMR18, Rui16, Ryd00, SFVP^{+}24, Sch80a, SLWH22, Sco10b, Sco13a, SCZZ99, Sim98, Ski19, Sko12, Sne99b, Som78g, Sta12d, Sta12e, Sun16b, Sun17, Sun19, Sun96, TC15, Tho19, TH98, Unw09, Unw10b, Unw11b, Unw11c, Unw20, Utt21, Veh08a, Veh08e, WQS19, WWY23, WD84, WC98, Wet80]. Data [Wil14a, WY02, XK15, Yim90, You14, YK20, YL17, YB20, Zha10, ZHW23, ZZLZ15, Zhu14, lRG05, HGA76, KFF75, Sch84b, SL13, WLC82, You86, ZT13, Gho10b, Bal76a, Bal76b, Bar80a, Bar77c, Blu21, Coo07, Cro10i, Dra08a, Dur18, Dur23a, Gho09i, Gho13d, Háj72, Ham79, Han08b, Han12x, Han13i, He14, Kha24, Kos10c, Liu17, Mai12c, O’B07b, O’B13g, O’B13e, O’B13c, Oja07, Pos80d, Pre79a, Pyn14b, Ric18, Stu80, Unw12d, Bat08a, Cap19, Dra09b, Esp15a, Fis19, Gho23, Han07a, Han07b, Han08d, Has08a, Kha15, Liu22a, Mai12a, Mai13b, Mus75, O’B16a, San15, Sta07a, Sta07b, Unw11a]. Data-Based [Sne99b, You86]. Data-Driven [NWS23]. data-gathering [HGA76]. Database [KRR^{+}11]. Datasets [Han08o, Han12p]. Daten [Lin75a]. Dates [Ano76m]. Dato [Veh08g]. Datta [Gho17]. Dave [Dra08b]. David [All79a, Cox80, Gho08g, Gho10b, Gho12f, Gho13b, Gho13c, Mai09a, Mer12b, Num20, Pet07, Pol13d, Ric11a, Sen11c, Spe10b, Spe10c, Spe11b, Sta08b, Sun18, Tha07a, Tra13, Unw10a, Val08, vdL80c, All79b, Huz80, Mar73, Sar73]. Davidian [Cro09b]. Davino [Dek14]. Davis [Kos10b, Bat07, Mer07, Sta07f]. Davison [Cro07a]. Dawid [Gho08g, Häg12a, Roy14]. Dawson [Pol13b, Daw05]. Day [DQS92]. Days [Ano83a, PF82]. Deal [LG04]. Dealing [CH19, ZTK13, Veh10a]. Death [Han10a, KT93]. Deaths [Kei87]. Debabrata [Gho12d]. Debate [Spe10e]. Debbie [Han12c, Han12l]. Deblurring [Mar07b]. Debora [Liu08]. Deborah [Mai08, McC22]. Decades [Ano74h, Fis01, NC21]. December [BBS07, WG97]. Decision [Buc79, Cop77, Gho08a, GPW05, Han08m, Ios79b, Pol11a, Pos80c, Rob11d, Sco13a, Sta09, Var14, Yan73, Elf72, Pos76b]. Decision-Making [Buc79, Elf72]. decisiones [Tin78a]. Decisions [Fal72]. Deck [AL10]. Declaration [Ano86h, Ole03]. Decoded [dO13, Gun15]. Decomposition [Rad10, KDTP20, Pun11d]. Decompositions [Tia07, Han08o]. Deconfounding [BC20]. d’Économie [Fou80b]. Deconvolution [Cro10e]. Deep [BDLP20, For22]. Deepayan [Nor08]. Deeper [Men16b]. Default [DS99]. Definite [Sch16b]. Definition [Dam74, Dam74]. Definitions [EUC09]. Défis [Mal97]. Degree [Lan76b, Pos78d]. del [Bol07, Tin78b]. Della [Cam07]. Dell’Oro [Tin78b]. DelSole [Dur23b]. Delta [Jør84]. Demand [BS85, BCO82]. Democracy [HRR07, Mal05]. Democratic [Ole03]. Demographic [Ben75, GW23, HWS20, SAA22, VL96]. Démographie [San77]. Demographische [San77]. Demographisches [Kro80c]. Demometrie [Sch72c]. Demyanov [Roh78]. Denison [Aoy72]. Dennis [Kau13, Sin02a, Tre05a]. Density [BT80, BK16, Ich09, LN74, SJ89, Sim98, Sri73, WS79, NL75]. Density-Based [Ich09]. d’Entrées [HL74]. Departures [SRN17]. Dependence [Dur21, DMMG22, Feu93, FGFBGM22, Sen96, Kos12b]. Dependent [AS72, AP94, Kos10c, Nai93, Num15, PAF18, SM14, Sch83]. depending [dM84]. Depends [CG23]. Derivation [GR90, IK86]. Derivatives [FGFBGM22, Sko86a]. Derrick [Ken75]. Descent [Tho98]. Describing [MW76]. Description [Lee01, Pre79b]. Descriptive [BO15]. Deshpande [Des05]. Design [AP81, Ald07, Atk82, Bic00, Bje73, BMM08, Bra15, Bre99b, CW73, Cox09, DPS07, Dra12d, Dra14c, DK87, Els18, Fil11, Gho21, GT17, Han08i, Lin14b, Luo07, MM96, Mas73, Nel72b, O’B07a, O’B08a, O’B08d, O’B12a, Pat73, RM93, RDMP16, STI06, SS17, SGLS02, Slo23, Spr80, SS07, Szc11b, Cha08, DM12, Fed80, Fed81a, Fed81b, Puk87, TTB86, Bil77a, Dra09b, Fil09a, Mai11b, Pun08d, Som78a, Tan07, Aur10, Cro10f, Dra11a, Dra12c, Fil09a, Fon11, Gho09f, Lis10a, Mai19, Pre78a, Veh08c, Wim75]. Design-Based [Bre99b, SS07, GT17]. Designed [AM93a, Dra09a, Dra10c]. Designing [Han12x]. Designs [BD20, BKBP21, CSW93, Dem18, Dra12c, DW15, Kag72a, Kag72b, Khu00, Kis94, Mat88, Pat88, Pre77b, Puk07, dMRAMR18, Tal93, Tju84, TBSW84, Atk88, FMS12a, NSZH22, Fin77c, O’B12d, Pat75b]. despite [Hed20]. d’été [Van16]. Detecting [SRN17]. Detection [Han12n, MW81, O’B09g, PGV22, RCA18, Sha23]. Detective [Kru89]. Determinacy [Ald99]. Determination [Kim15, Rya13]. determining [dGAGR85]. Deterministic [CHR18]. Deuxième [Esp11]. Develop [Sne99b]. Developed [Hoo05]. Developing [And73c, Azo01, DH19, Gal02b, Hen78, Jab96, JS03, Kis96, LH04, Mul76, Ram03, SCZZ99, Ver96]. Development [Cra81, DM12, Gow15, Han07f, Han08i, Han13i, Kir05, Mat07, Nee09, Pow03, Rad99, Ram03, SDGD^{+}11b, Sne99a, Som78g, Wat98, Yoo21, MCF12, SB75, Lan18]. Developments [Atk82, Bar83, BT80, CH19, CP96, FW04, FZ15, Haw97a, Hon15, HS15, JK82, Kal19, Mat88, Nay09, New81, New84, New88, PMLL15a, PMLL15b, Pfe02, Pre78b, Rip84, Tsa24, Veh08d, dVKW96, Atk88, Mal97, MMC13, Shi81]. Deviance [FWS14]. Deviations [Gow79]. device [HGA76, MDS^{+}19]. Devices [MDS^{+}19]. d’expérience [Dru11]. Dey [Gho12a, Pol11a, Zha17a]. Dhall [Han07c]. Dharmendra [Veh07b]. Diagnosis [SAG^{+}16]. Diagnostic [GSGdMLM16, Han13m, WWY23, Han11d]. Diagnostics [DT92, MR22, Pou02, SCLR16]. Diagrams [Daw02a, Daw02b, Gho08b]. Dialectic [Fér72]. Dialectique [Fér72]. Dialogue [Spe10b]. Diana [O’B21]. Diane [Cro10f]. Dianne [Mai08, Mar07b]. Diarrhoea [KMM^{+}17]. Diary [Dra12e]. Dictionary [Dra11e]. Dies [Ste95]. Dieter [Cro11c, Mat07, Veh13a]. Dietrich [Han08l]. Dietz [Han10c]. Difference [Kag72a, Sun18]. Differences [HG21, Nem73]. Different [HM04, KMT03, MDS^{+}19, Red74]. Differential [AET04, Dem18, Liu20b, Luo09, Pin75, PS13, San21, BNCR86]. Differentially [FWC^{+}21a, FWC^{+}21b]. Differentiated [MR80]. differentiation [Sko86a]. Diffusion [GP03, JS03, Sør04, Sør83, Pos79a]. diffusion-type [Sør83]. Digit [NSS73]. Digital [Nel80]. Dilemmas [O’B14a]. d’Images [AM93b]. Dimension [CD18, MZ13, Slo24, YAA20, vdL03, Ios79a]. Dimensional [Gho12g, GM19, Han13g, Kha15, Kha24, Koc14, LD24, Mod20, Oja07, Pun10d, BCM19, Gho10b, Gir15, RDM^{+}19, You15, ZT13, You20]. Dimensionality [Hol08]. Dimensions [LKC^{+}04, ZYW23]. Dimitri [vdL80d]. Dimitris [Qiw20]. Din [Num13a]. Ding [Num13a]. Ding-Geng [Num13a]. Dinov [Cap19]. Dipak [Gho12a, Pol11a, Zha17a]. Direct [Bat07]. Directed [Cas20]. Direction [ANK03, Dur21]. Directional [AW95, Ell91, JM89]. Directionality [Law91]. Directions [GG99, KS22, Nic01, Pfe02, SHLW22, Veh08d]. Director [Spe10a]. Dirichlet [AAF18, FMR99, GR01]. Dirichlet-multinomial [AAF18]. Dirk [Pod18, Ros14, Tra09]. Disadvantage [BCM19]. Disaggregation [Gue90]. Disaster [Rao93]. Discipline [Ott05]. Disclosure [CFP22, CO14, CKK^{+}11b, Han12p, Ich09, NA16, Shl07, vdHvdH02, O’K12, Rei12, Ski12a, Ski12b, Wil12]. Disclosure-protected [CO14]. Discontinuities [vdBZT20]. Discoveries [Hoe83]. Discovering [CC08, Unw13a]. Discovery [DTP^{+}20, Han12h, NWS23, Unw10a, Mai12a]. Discrete [Dou95, GM19, Ios78, JK82, Liu17, MYA95, Sør04, GT12]. Discriminant [SJ89, vdL03, Smi77]. Discriminating [dBP77, BG83]. Discrimination [FH89, Ott75b]. Discriminatory [FH89]. Discussant [Big05, Daw05, Des05, Hir05, Huh05, Kal05, Lee05, MG05, Par05, Ric05, Teu05a, Tre05a, Tru05, Van05]. Discussion [Aal82, Abe15, ACV84a, Ahl02, Ahn14, Ale94, AG99, AG81, ABGK82b, And86, Arm85a, AM14, Atk84, AG15, Bai95, Bai84, BBG^{+}76, Bar82, BCL^{+}85, Bat02, Bat85b, BE99, Ber82, Ber03, Bie99, Bie95, Bin00, Bon96, BCZ12, Box84, BC81, Bre99a, Cam97, CS20, Can95, Cau86, Che85b, Cia14, Cla00a, Clo86, Cox81, Cox84b, Cox13, Cox20, CV82, Dar84, Dat12, Dav20, Dei83, Din82, Duf82, DG17, Efr13, Eng83, EJ86, FZ15, Fra13, FHT20, Gar97, Gho97, GDV20, GS86, GSW17, Goo86b, Gra20, Gra83, HHDG12, Har82, Haw97a, Haw97b, HR83a, HK16, HHD97, Hoe99, Hon15, Hor16, HS15, Hub14, Hun14, Jab96, Joh81, Kei81, Ken86, Ket97a, Ket97b, Klo83, Koe16, LT82, LP15, Lin82a, Lou85, Lyb00, Ma15, Mac16, MR03, Men16b, Men16c, Miz83, Moo99, Nar20b]. Discussion [New83, Nor96, Oak81b, OCRC17, Ott02, PJK17, Par13, Pet85, Pew15, Rei12, Rob13, Rob14, Roy94, Rub03, Rum02, RZ14, San93, Sch97, Sch03, SH13, Sen84, Sha82, Sha85b, Sil82, Sin95, Smi82, Smi94b, Smi99, Smi16, Sne99a, Sne99b, Sne02, SZ14, Spe84, Spe85, Spj84, SR15, Str14, Taa11, Tam84, Tam94, Tan97, Teu97, Tho84, TC99, TBSW84, Ver96, Wah84, Wal96, WS16, Wan19, Wat02, Wel02, Whi85, Wil16a, Wil14a, Wil12, Wyn84, gXZ20, Zad82, ZM17, And87, GS87, LHK78, Lau81, LN78]. Discussions [Abe15, Ahn14, AM14, Cia14, CKK^{+}11b, Cox13, DF11, Dun11, DP20, DG17, Efr13, FZ15, Fra13, GSW17, Han11m, HK16, Hon15, Hor16, HS15, Hub14, Hun14, Kir11, Koe16, LK11, LSD^{+}11, Mac16, Men16b, Men16c, O’K11, O’K12, OCRC17, PJK17, Par13, Rob13, Rob14, RZ14, SDGD^{+}11b, SH13, SW11, Shl11, Smi16, SZ14, Str14, TD11, WS16, Wil16a, Wil14a, ZM17]. Disease [BR00, BD94, Cla00b, Cro11f, Gho09a, GDSN18, LC20, TZW^{+}20, WOP20, MMC13]. Diseases [CD00, Kin77b]. Disequilibrium [Cla00b]. Disparate [GMZ03]. Dispersion [BK88, CPB94, Jør92, SB87]. Displaying [Per14, Pod15b]. Disproportionate [LNA75, NAL76]. Disseminate [Rei09]. Distance [AES18, DC83, EFP19, FMR19, GM19, Han13n, SM19]. Distance-Based [DC83]. Distinguish [GL19]. Distributed [Smi73]. Distribution [And78, Cox13, DW84, Efr13, FL83, Fra13, Ham72b, Ham75, HHMR18, Hou88, LN74, NK06, NK08, NN20b, Par13, Pod72, Rob13, SH13, VF00, BBS77, Han10b, McC85, gXS13, gX13, NN20a, Ord72]. Distribution-free [NN20b, NN20a]. Distributional [DS78]. Distributions [Abe15, Alv20, AG08, AG15, Ber78, BCF^{+}14, CMW17, Cro08b, Dou95, Eng78, GKV19, GKV20, GL14, GR01, HH14, JK82, Jon07, Jon15a, Jon15b, KX05, Lee79, LV14, LP15, Ma15, Mod20, NT72, Pat75a, Pew15, Rob12a, Sch16b, SDT^{+}23, SR15, Tad80, Tan92, BNKS82, Chm81, Kot74, Mit88, NL75, Rao77, Buc74, Huz73, Kha74, Pet07, Buc74, Mar75]. Disturbances [Cau80]. Divergence [ADE^{+}22, BL24]. Diverging [Ano88a, FT87, FT88]. Diversions [Mer09d]. Divide [FPD05]. Divine [Sho85]. Dixon [Ken76b]. Dmitrienko [Sta07e]. Dmytro [Han10j]. DNA [Gho09g, Han09d, FCE^{+}03]. Do [Bla07, Sne02, Ize03, Sco12a]. Documentation [Dag76, SAA22]. Documents [NL07]. Dodge [Veh08h]. Dodson [Els18]. Does [AAF18]. Doganaksoy [Han10i]. doi [Ano21]. Doing [Ald98, Bie97, Lin02, Pea02b, Sin02a]. Dolores [Bat08b]. Domains [Ano80e, PK80]. Domenica [Rod21]. Domenico [Dek14]. Domes [LKC^{+}04]. Dominique [Han13i]. Donald [Fin80c, Imb07, Kin77a, Pre80a, Som76b, Daw05]. Donatas [Kos12b]. Done [Lee79]. Donev [Puk07]. Dongchu [Pol11a]. Donna [Spr77b]. données [Mar07a]. Donnelly [Han12k]. Donovan [Blu20]. Don’t [Dra13g]. Door [MBW^{+}23]. Dora [Dra12e]. Dorans [Han08g]. Dose [FMP20]. Dose-Finding [FMP20]. Double [Ber72, TWRB20]. Douglas [Cro10g, Dra08b, Khu13a, Pos78d]. Doukhan [Kos12b]. Down [Rob12b]. Drachen [Liu22b]. Draper [Cox07]. Drawn [HRR07]. Dream [CD00]. Drechsler [Han12p]. Drennan [Dra11d]. Driven [NWS23]. Droesbeke [Mar07a]. Drogin [Som76c]. Dropout [WQS19]. Drosg [Veh10a]. Drug [Han08i, Ize03, Mat07, Yoo21]. Dual [CNS85, CT11]. Duan [Kha17, Kos13c]. Duane [Veh08f]. Duanli [BPV21]. Dubbelman [Cau80]. Duchateau [Här08]. Dudewicz [Lin80b, Rob12a]. Dudoit [Cro08c, Dra13d]. Due [LHSD15, BH80]. Dufrenoy [Ves73]. Dumontier [Mal75]. d’un [Dam74]. Duncan [Cro10j, Han12m, Nor09a, Sta12d]. d’une [Lef79]. Dunn [Ric10a, Spr75a]. Dunne [Han08a]. Duran [Tip77]. Durand [Som78g]. Durante [Kos12a, Tso17]. Duration [FMR99, HJ83]. Durrant [Sun22]. Durrett [Tso16, BN07, Gho09g]. Duties [Ole03]. Dworsky [Hau08a]. Dwyer [Ken75]. Dynamic [EM93, HM04, Mai08, Men16b, NL07, NWS23, RSG06, Roh16, SAP97, Tri09, Tam87, Han07c, Ios72, Liu10a, O’B10c]. Dynamical [Cro09e, Han08k, Han11i, Ric16]. Dynamics [BCM19, Hol84, BN07].

e-book [Kha20, Sen23, Zha19b]. E-Learning [HKZ07]. E. [SJ89]. Early [Bor09, Hal00, Hoe83, Har79]. Earthquake [Mis08]. ebk [Liu23]. eBook [Rod21, Fis24, Liu20b, Num20, Num19, You19]. Eckhard [PS13]. École [Van16]. Ecological [GTF18, KMM^{+}17, O’B07b, O’B13e, SH96]. Ecologist [O’B13d]. Ecology [O’B12h, O’B14b, Pat79, SFVP^{+}24, Gho11b]. Econometric [ABC11, Bar75b, DO73, Fil09b, Hec08, HR83b, HR83a, MSW76, MNS77, NSW72, Pun12a, Smi72]. Econometrics [Bre80, Gow79, Liu10a, Sco09b, Tin78d, Neu16, Bre80, Liu20b, MM77, McE17, O’B09d]. Economic [All72, And73a, Ano83a, Bje73, CW73, HR83b, HR83a, Ken97, Keu97, MSS20, Pan07, PF82, Pos76d, Rao80b, Rui72, Som78g, Sun77, vdL80c, WC75]. Economics [Ald07, Cra73, Hai72, Hau08b, Knü80, MM77, Pre77b, Men73, O’B11c]. economies [Hau08b]. Economist [HV15]. Economists [Alb19, Pos76a]. Economy [Gol97, Nie04, O’B12e]. Econophysics [Han10a]. Edgeworth [Ano89a, Sko86b]. Edgington [Man07]. Edited [Dur22, Sun22, Bla07, Bol07, Coo07, Cox12, Cro09b, Cro11e, Cro11d, Cro11a, Csi08, Dek10, Dra08b, Dra09d, Dra11b, Dra12c, Dra13d, Fil11, Gay11, Gho09b, Gho10b, Gho11a, Gho12a, Gho12d, Häg12a, Han07h, Han08f, Han08c, Han08g, Han09a, Han10g, Han10h, Han12d, Han13g, Kau12, Kos10b, Kos12a, Kos12b, Kos12c, Lis10c, Liu07, Liu10b, Mai11b, Mar07a, Mer09a, O’B09b, O’B10d, Pol11a, Pol12a, Pol12b, Pun11a, Ric10b, Sen11c, Spe08a, Spe10d, Spe11b, Sta08c, Sta08b, Sta10b, Sta12b, Veh07b, Veh08d, Veh08e, Pod14a]. Editing [Gra97, PHO86, dWC05, Mus75]. Edition [Alv11, Cox07, Cro10b, Cro10a, Cro10f, Cro10g, Cro11b, Cro11c, Dra09c, Dra09a, Dra10b, Dra11d, Dra11e, Dra12e, Emb08, Eva07a, Fil09a, Gho08a, Gho09c, Gho09g, Gho09f, Gho11c, Gho12h, Gho12b, Han08e, Han08j, Han08b, Han10e, Han12q, Han12f, Han12c, Han12l, Han13h, Khu13a, Kos12c, Kos13b, Lis07a, Lis07b, Lis09a, Liu10a, Liu20b, Mai09c, Mai11a, Mai11f, Mai12b, Man07, Nor09b, Nor09c, Num09b, O’B08c, O’B11a, O’B11b, O’B12g, O’B12a, Oja08a, Oja11, Pol13d, Pun07a, Pun08b, Pun09, Pun10a, Pun11b, Pun11c, Pun12c, Pun13c, Pun13b, Ric10a, Ric11c, Ric11a, Rob11c, Rob11f, Ros09a, Ros09c, Sco09a, Sen11c, Spe10c, Sta07a, Sta08a, Szc11a, Szc11b, Tha07b, Tra07, Tra09, Unw10b, Unw11a, Unw12c, Val08, Var19, Veh07a, Veh08b, Veh09b, Veh10a, Veh10b, Veh11a]. Edition [Veh13b, Bra15, Ros15b, Esp11, Dur14b, Fon13, Han13j]. Editor [HN13a, Hee16, PA15, Alv20, Ano94d, Cha16, GKV20, HN13b, SZ09]. Editor-in-Chief [SZ09]. Editorial [Ano99e, Ano01d, AdV99, Cha16, Had10c, PA15, ST92]. Editors [Blu21, Dur21, Pin10a, Pyn14a, Ano99e, AdV99, HN12, HN13b, LSS08, NLSS08, PP98]. EditorsinChief [Pin10b]. Eds [Ros19]. Education [Ald05, Bre79b, GO16, Gan79, GG99, Haw97a, HH95, Hor16, LS04, Mac16, Men16b, OG16, Ott05, Rid16, STM95, Tam07, Utt21, VJ95, Wil05, Wil16a, Pol12b, Ros19, Gan78a]. Education-Fertility [LS04]. Educational [Han13l, Sta12f, Sta13b, Dur17]. Edward [Han09a, Kos11, Lin80b, Rob12a, Spe10e]. Edwards [Han13e, Pyn14a]. Edwin [Ios77c]. Eelko [Eva07b]. Effect [CS02, CS88, Han12w, IMG22, JRL10, O’N09, Slo23, BNC84, Hed20, Mal97]. Effective [Bon96, EMR22, Fel96, Mer24, Nor96, Til22, Wal96]. Effects [AW95, Con09, Cro10i, Dar81, Doc22, DSC03, GP03, HMW21, Lav14, Mai14c, NSP21, O’B15b, Pre89, Rao93, SH96, TO17, ZMSL19, Don86, Sco12a, TTB86, VLM12, Imb07, Num23]. Effet [Mal97]. Efficiency [BCvZ81, GO81, KX09, MR03, XK15, Iac84]. Efficient [Han12s, KS22, TTV05, WGKL22, Mon87]. Effizienz [vdL80b]. Efforts [SDGD^{+}11b, Wyc03]. Efron [Spe10d, Sun19, gXZ20, CS20, Cox20, FHT20, Nar20a, Spe11a]. Eggermont [Gho11d]. Eight [Ano74h]. eighteenth [Lan86]. eighth [Lan74b]. Eileen [Veh10c]. Eilers [Pod22a]. einiger [vdL80b]. Ejaz [Kha24]. Ekstrøm [Sko12, Häg12b]. El-Mehdi [Kha77]. El-Nasr [Liu22b]. Elaboration [ADE^{+}22]. Eldén [Han07e]. Electric [HHDG12]. Electricity [CG12, MCH13]. Électrique [Fou80b]. Electronic [MBW^{+}23, Sun17]. Elegant [Han11c]. Elementary [GP73, Ric13c, Der77b, Gho08c, Som78b, Som79d, Som79c, Sta07c, Tip77]. Elements [Nor09c, Sär72b, Coh75, Der77c]. Elena [Han08h, O’B07b]. Eleusis [Mer09d]. eleventh [Lan78]. Elias [Pod21b]. Elisabeth [Aur10]. Elizabeth [Khu13a, Ric17a, Zha19b]. Elke [Dur13c]. Ella [O’B09b]. Elliot [Han12m]. Elliott [Sen14d]. Ellipsoids [Le 90b]. elliptic [Mit88]. Elliptically [BCF^{+}14, Chm81]. Ellis [Sco07]. Elusive [Meh15, Ver13]. Embrechts [GN20]. Emerged [Rob12b]. Emerging [EIRS18, EB92]. Emigration [Red01]. Eminent [TTV05]. Emmanuel [Fis22, Mai09b, Pod17b, Tso18]. Emphasis [Pos78e, Smi80, WC75]. Empire [Bel02, Cox02, Sin02c, Sin02d, Spi02]. Empirical [BK20, Bic14, BMM08, DK01, DO73, FM95, HL90, Hum77, KP04, LNA75, LCL^{+}23, MNS77, NAL76, NT72, Num09a, Smi99, Spe11a, TWRB20, Wel92, WP99a, WP99b, ZW19, EMR88, MS87, Han12f]. Empirico [Fom88]. Empirico-Statistical [Fom88]. empiriques [EMR88]. Employees [CN98]. Employment [GMZ03]. Emrouznejad [O’B13b]. Encompassing [ZHW23]. Encounter [Ber77]. Encyclopedia [Veh08h, Han09a]. end [MMC13, Kha20]. End-to-End [Kha20]. Endogenous [Frö08]. Energy [Had10a, MFAG10, Ney75, Had10b, HN12]. Enforcement [Unw10a]. Engel [Ano75e, Per75]. Engin [Dur21]. Engineering [Cre20, Cro10g, Dur14a, O’B13f, TP73, Var14, dO79a, Gho12f, Gho13c, Ng21, Oja08b]. Engineers [Law07, Lis09a, Liu19, Sta10c, Tip77, Alb19, Cro07b, Els18]. England [And77b, Sti88]. English [Ald03, Ric11c]. Enhancement [CBPAK22, God97]. Enhancing [GO16, Hor16, Mac16, Men16b, OG16, Utt21, Wil16a]. Enigma [Rob12b]. Enough [Mic19]. Enquiries [Ben75]. Enquiry [Smi99, WP99a, WP99b]. Enrique [Bol07]. Ensemble [Nor13a]. Enslein [Nel80]. Enterprise [BP92, Dur14a, Sta07f]. Entre [Lec76]. entries [HL74]. Entropy [CQOR^{+}22, Ken73c, NC21]. Entropy-Complexity [CQOR^{+}22]. Envelope [Car05, LS20]. Environment [Box75, Ole03, Rad99]. Environmental [And73b, Cro10j, ESTTD05, Gut03, Han12n, O’B12e, O’B14c, RYG^{+}21, Wik03, vBS73, Mai11b]. Environments [ST07, Wil07]. Epidemiologic [Cro08e, WGKL22]. Epidemiological [BRST76, DR93, FWC^{+}21a, FWC^{+}21b, RYG^{+}21, Sun17, Sun24]. Epidemiology [Doc22, SP11, Cro10j, Gho09a, Mai13c, Tha07a, Zha14]. Epochs [Ios76f]. Epsilon [Dra09d]. Equal [AS73]. Equation [Fon13, OH13, Ros15b, San21, Sun16c, Veh11a, Veh14, Sri80]. Equations [AET04, BS85, BCO82, Dem18, Luo09, PS13, TP73, Han11f, Pin75]. Equilibria [Pos76d]. Equivalence [Han12q, WLC82, Ric17a]. Equivariant [BF91]. Era [RR21, Sne99a, Wat98, Ket12]. Erhardt [Lid80]. Eric [Cro11g, Fis18, Han10f, Han13g, Mai11b, Mer17, Unw13b, O’B13e]. Erich [Sco10b, Spe08b]. Erik [Kul78]. Erkki [Liu07]. Erlang [AK95]. Ernest [Mai11d]. Ernst [Mic77, Wol76]. Errata [Ano72c, Ano74d, Ano75f, Ano76f, Ano78g, BT76]. Erratum [Ano21, FWC^{+}21a]. erroneous [BH80]. Error [AES18, Bai95, Bie95, BS85, Blu21, CDM19, CW06, Cro10h, DO73, Dra13f, FMR^{+}20, Ful95, Han11h, KT03, NC10, PBH19, SH00, TC07, Veh08f, Veh10a]. Error-Free [Dra13f]. Errors [AJS81, CD94, Hai72, KP04, Kir82, LvR74, ŞA09, Smi21, VL96, Dal77a, Dal77b, Dal77c, NW82, ZTK13, Sun18]. Erwin [Eag74]. ESA [Ano78k, Lis09b]. Escobar [Var19]. Esposito [Han12g]. Essay [Dem77]. Essays [Kin75c, Tin80b]. Essentials [Cha14, Som80a]. Established [CFL^{+}21]. Establishment [Mag09]. Estimate [CG12, LvR74]. Estimated [GT17, KL09]. Estimates [Ano80e, BCM19, BGV18, BF91, ES06, Har90, Hou88, Lit86, PK80, Rao05, Tam82, MS87]. Estimating [Ano79e, CHR18, Chu74, DW84, EKIKW77, Gue08, Har99, Kuk22, Red01, Seb92, She06, SYSK02, Smi21, Tho98, WHT24, Zha96, ZMSL19, KDTP20]. Estimation [AS72, AET04, And76, BCM19, Bai95, BSS18, BT80, BE17, Bha91, BFST18, Bie95, BK16, BMM08, Buo06, CS20, CFP22, Cau80, CDM19, CL19, Che24, CM12, CKS17, Col81, Cox75, Cro07b, DH19, DNC06, DC19, DO73, DFG^{+}94, Dor18, Efr20a, Eid07, FS94, FS08, FHT20, Ful95, FS10, GW23, Gho08d, Gho11d, GS19, God80, GH87, HM90, HH93, Han08m, Han13g, HL01, HL74, Khu00, KP10, Koe79, KBM21, LdBP90, Lia20, Liu12b, Mer24, MFAG10, MSS20, Nai93, NC10, NT72, Nor72, Nor73, O’B16b, O’B20, Pfe02, PBH19, Pro11, Rob93, dMRAMR18, Sär72c, Sch78, SB17, SJ89, Sim98, SHGK01, Som78e, Som79a, Sør93, Spe11a, Spr75b, Sri73, Sri80, SST20, Tho00, TTB86, Til98, Tra13, TC72, UT10, VT19, Van16, WOP20, WS79, WS21, gXZ20]. Estimation [YAA20, dO79a, GT86, Hig77, Jac84, JOT85, RLS23, SS87, Sco12a, Sør83, Sør86, HL74, Num23, Rob11b, Sen11a, Tin80a, Tra07, Cox16, Han10b]. Estimator [Cox13, Efr13, Fra13, KBM21, Par13, Rob13, SH13, SH23, gXS13, gX13]. Estimators [BE17, BDS04, BCF^{+}14, CNS85, CT11, DIV17, Dra10b, GNM21, GT17, HM04, LSD^{+}11, O’B16b, Sen76, Smi72, SH21, SS07, WQS19, YK20, YYY20, ZZLZ15, Bin83, La 16, Roy86]. eText [Ana21]. Ethan [Dek22]. Ethical [Sta10b]. Ethically [Han13b]. Ethics [Ano86h, Ole03, Utt21]. Ethnic [Hau01]. Étienne [O’B13e]. étude [Lec76]. EU [Lon15, Esp15b]. EU-SILC [Lon15, Esp15b]. Eubank [Sco13c]. Eugene [Gay11, Man07, Ros09c]. Eun [Lis14]. Eun-Young [Lis14]. EUR55.99 [Pod18]. EUR85.39 [Num20]. Europe [Esp15b, Lon15]. European [Ios79b, Pre78b, HV15, Hor91]. Eva [Gho11a]. evaluate [VLM12]. Evaluating [Lan18, SBB^{+}12, Han08e]. Evaluation [Bec73, BS92, Bre79b, Dor18, Fel73, Gho09b, Han13m, Mak78, MP21, Mur74, Nel74, SJFM95, Swe95, TSR96, Tho78a, Veh08c, BKV13, SB75, Ano78g]. Evans [Zha17b]. Event [FST19, Gho09h, Liu17, Zhu14]. Events [Ano79e, EKIKW77, dM84, Cro11c, Han08n]. Everitt [Dra11e, Rob11c, Unw12a, Eva07a, Han09a, Pun10a, Ric13a, Tha07b, Veh10d]. Everybody [Wil05]. Evidence [Blu05, Cur03, FCE^{+}03, HP22, Kaf15, KMS14, Mai22, Sun19, Tsi03, Bic13, MDS^{+}19, Han07g, Han08h, Mai09a]. Evidential [TRRL23]. Evolution [BJPT99, Con91, Fis09, GMRS99, Hor72, Kok07, Pan89, BKV13, O’B14a, Gho09g]. Évolutions [Mal97]. Evolving [PAF18]. Ewout [Nee09]. Exact [Dim01, Gho08g, LS06, LLP08, OH19, RG23]. Examination [MNS77]. Examine [Unw21]. Example [BD94, EM93, Ela73, EPP08, Har73, Mai11c, Nov73, Sta07a, Sta07d, Unw11a, Han12v, Khu13b, Pun13c]. Example-based [Sta07a, Unw11a]. Examples [And73c, Liu12b, Nor98, Rui17, Tho78b, BKV13, You86, Buc77d, Han13f, Luo09, Mai10c, Pol13d, Spr77b, Amu73, Mer14]. Excavating [Mai12a]. Excel [Han07a, O’B12c, Sco10b, Sta13b]. Excess [HPHdAM22]. Exchange [Sch78]. Exchangeability [Mod08, O’N09, Ham77b, dF79]. Excluding [KM79b]. Exclusion [Hop93]. Exercise [Sav80]. Exercises [Aoy72, Pod16, MM77, Num13b, Pod14b]. Existing [PZ21]. Exits [FMS12b]. Exotic [Jau14]. Expanding [Cle01]. Expansion [Ham72b, wB09]. Expansions [Ano89a, Sko86b, Ios77a, GP12]. Expectancies [Con09]. Expectancy [GMZ03]. Expectation [Ros87]. Expected [Kei87, SB17]. Experience [GP03, HHDG12, PHO86, TH98, WB97]. Experiences [Che85a, Fel73, FB73]. Experiential [Cob07, Wil07]. Experiment [Blu18, Pre89, Fed80, Fed81a, Fed81b]. Experimental [AP81, Ano88a, Bil77a, DPS07, Dra09b, FT87, FT88, Fil09a, Fin77c, Ham79, Luo07, Man72, Puk07, SVCV88, Tan07, Atk88, DM12, Puk87]. Experimentation [FT96, Pre77a]. Experiments [AM93a, Ald07, Ana21, AW95, Atk82, Cox09, DPS07, Dra10c, Dra11a, Dra12b, Dra12d, Dra12c, Fil09a, FMP20, Fon11, HH11, Mai19, Mas73, Nor98, Pau19, SS88, Wim75, Dra09a, Dra14b, Nel72b, Pat73, Pre78a, Spr80]. Expert [Gho08g]. Explain [Unw21]. Explanatory [San91]. Explicit [LS04]. Exploratory [Gel03, UT10, Alv11, Mai11c, Unw21, Zha21]. Explore [Unw21]. Exploring [Ken76b, Sta10b]. Exponential [Bil80b, CR93, Han09b, Jør92, KS89b, Væt85, McC85, Sør86]. exponential-family [McC85]. Exponentielles [Cas91]. Exposition [EIRS18]. Expository [Ani08, UT10]. Exposure [Wan16, Sun16a]. Exposure-Response [Wan16, Sun16a]. Expression [Bla07, CC08]. Extending [Hoo05]. Extension [HG94, HHMR18, Joh83]. Extensions [Fre83, Goo86b, Jør92, TGM09, And87, GS87, Goo86a]. Extensive [CFL^{+}21]. Extent [Sta10b]. External [BC20, Dor18]. Extinction [Alb95, AK95]. Exton [Mor80a, RR79b]. Extraction [DCC98, KTM01, ZYW23]. Extrapolation [DCC98, Wan19]. Extrapolation-based [Wan19]. Extreme [CP96, DY16, GG15, Har78, LR91, dO79b, Dur22, Zha17a]. Extreme-Value [Har78]. Extremes [Cro11c, GG15, STY90, Mat73]. Eye [Dur21, Lis14, Dra13e]. Ezra [Mer09a].

F [Bar72c, Bat08b, Bil77a, Bre79c, Buc77c, Cor75a, Cox75, Cro11e, Dei83, Dra14d, Eng83, Fin77b, Fou80b, Gal80a, Gan78c, Gho09d, Gra83, Han08n, Han12o, Hoe83, Kin75a, Klo83, LS06, Lis08a, Mai08, Mar73, Miz83, MB19, New83, O’B07b, Pos76c, Pre78b, Pre80b, Pun08a, Pun13b, Rao80c, Roh78, Ros09c, Sco77b, Smi94b, Sta11a, Sta11b, Szc11b, Tip77, Veh08f]. F. [Dei83, Eng83, Gra83, Klo83, Miz83, New83]. Faber [Dor73, Var14]. Fabrication [Fri93]. Fabrizio [Han12t, Kos12a, Tso17]. Face [Dag76]. Facilitate [SAA22]. Facteurs [Dam74]. Factor [Bar80b, DS99, Han13h, MCH13, OH13, Ros15b, UT10, Veh10b, Veh08d]. Factorial [AW95, CSW93, Fou80a, SB87, Tal93]. Factorielle [Dag75]. Factorisation [HGKJ20]. Factors [HM19, HBK05, Mai22, Pro11, Sär11, Dam74, Hoi13]. Facts [Mar21]. Fahrmeir [Lis10c]. Fail [Dra13g]. Failure [DS22]. Fairclough [Cro10f]. Fairness [ZZN^{+}22]. Falcon [O’B09a]. Falissard [O’B12b]. Faliva [Liu10a]. Falk [Cro11c]. Fall [Dra13e]. Fallacy [GTF18, SH96, Han12s]. Families [Abe15, Alb95, AK95, AG15, Bil80b, DR90, Jon15a, Jon15b, KS89b, LP15, Ma15, Pew15, SR15, Væt85, Sør86, Cro08b, Dur14c, Huz73, Kha74]. Familles [Cas91]. Family [Alv20, BL24, BJS92, GKV19, GKV20, NT72, Som78f, McC85]. Fan [Han07f]. Farcomeni [Pyn14b]. Fast [RDM^{+}19]. Fatalities [LH73, CP80]. Faul [Liu20a]. Favorable [DR90]. Feasible [ZYW23]. Feature [ZYW23].

Features [OA91]. Federer [Fin75]. Feichtinger [Kro80c]. Felicitations [Ano74h]. Felix [Ano73e, Sch05]. Ferreira [Gho08f]. Fertility [AdB04, Ins73, KP04, Kró80b, LS04, PHO86, Yim90]. Feryaal [Kha24]. Festival [Ano83a, PF82]. Festschrift [Fil09b, Lis10c, Spe08a, Csi08, Liu07]. Few [Bic14]. Fiducial [Ano96d, Bar95, FB96, Ped78]. Fieguth [Gho12e]. Field [BS89, Cle01, EIRS18, Gan79, Gri15, Rou07, SVCV88, Unw13a, WD84, Unw13a]. fields [Ish81]. Fienberg [Spe11b, Stu80]. Fifteenth [Lan83]. Fifth [Gan80c, Khu13a, Pol12b, Pun13c, Ros09a, Lan72, Mar96]. Fifty [Ahn14, Cia14, Loh14a, Loh14b, RZ14, SZ14, Str14, Spe08a]. Fildes [Pol13c]. File [Chi22]. Filling [Ast94]. Filtering [Piz10, Mar07b]. Finance [Car05, Kos12c, Pun12b, JT13, Feh07, Kos13c, Kos13a, PS13, Sta12b]. Financial [Emb08, FS10, Gho12c, Gho12f, Gho12h, Han07h, Han10j, Han12s, Kos09b, Kos10b, Kos11, Lep14, Liu11, Pod11, Ryd00, Sco13b, TTV05, Han08f, Pod18]. Find [Der75b, Mor74]. Finding [DW15, FMP20, Wil16a]. Findings [DO73]. Finetti [RR79c]. Fingerprint [Das13]. Fingerprint-Based [Das13]. Fingerprints [Kay03]. Finie [Lef79]. Finish [Mai11e]. Finite [Ano86a, GS19, KT21, LdBP90, Lis10b, O’B20, dBPR83, RBR85, Sen11b, SS07, ZL22, Lef79, dBPR86, Han07d]. Finkelstein [Cro10d]. Finn [Gho08a]. Finney [Kha73]. Fiona [Han08b]. Fioredistella [Rod21]. First [BCvZ81, Bre79a, Dra13b, Ios76a, Ios80a, Kos10a, Mai14a, Mar73, Mer09c, Mis10, Moo05, Nor09a, Qiw20, Ric13b, Sco13b, Sen14b, Sta10a, Wak94]. Fish [Bar75b]. Fishburn [Pos76b]. Fisher [Els20, Sco07, Ald00, Ald07, Bar87, FT96, Fin80c, Jaz72, Jef74, LV14, MPR12, Men73, Osb92, OH19, Pea74, Qui21, SHLW22, WR05]. Fisher-von [LV14]. Fisheries [O’B12g]. Fit [Feu16, RB90, Gel03]. Fitting [Dou95, LD24, Rob12a]. Fitzmaurice [Cro09b]. Five [Kis94, Ric13b]. Fix [SJ89]. Fixed [Dar81, TO17]. Fixed-Effects [Dar81]. Flaws [Bel02, Cox02, Sin02c, Sin02d, Spi02]. Fleiss [Pla75]. Flexible [OZW23, SLWH22]. Florian [O’B09a]. Flour [Van16]. Flow [FMS12b, Gal80b]. flowgraph [CH13]. FM [DS20]. fMRI [LEGW01]. Fokoué [Mai11d]. Foltz [BPV21]. Forbrig [EV73]. Force [CB00, Som78g]. Forecast [AL86, LLT04]. Forecasting [And73a, FM95, FMS12b, Han12v, LQ04, LG04, Mak78, McL93, MCH13, SSOL04, Smi92, Tin78e, ZG12, PMM87, Fer80b, Pol13c, Han09b, Ord80b]. Forecasts [AdB04, AES18, Gol04, KP04, LO04, LH04, SM93, TLL04]. Foreign [Red01]. Foreman [Ric10c]. Forensic [Cur03, DS78, Kaf15]. Forensics [Han09d]. Forest [Has08c, O’B12f]. Forests [Mai21]. Foreword [Ano22, HR20, HRR20, HR21]. Form [Ast94, Ios79a, MW76]. Form-Filling [Ast94]. Formal [MGMRPA90]. Formation [Ano74m, Ano74n, Ano75o, Ano76k, Ano76l, BHBP72, Bra75, FOK73, Hor73, Nai73, Nel72a, NPT^{+}74, Nel75, Zar76]. Formula [RAC83, Tan92]. Formulae [Ham78, HJ83]. Formulas [Liu14b, Vos75, Liu14b]. Formulation [Gel03]. fortuitous [dM84]. Fortuna [Hau18]. Fortunato [Ros10b]. Forward [Haw97a]. Fostering [FMS95]. Foulkes [Fis10a]. Foundation [Sta11a, Sta11b]. Foundational [Bow04a, SB04, Tam07]. Foundations [Bel02, Cox02, Gho12c, Han12r, JR08, KG10, Nor13a, Rob11g, Sin02c, Sin02d, Spi02, dF74, Bre80, Cox19, Pos79b, Veh10b]. Founding [Syl14]. Fourier [Gan78b]. Fourteen [Ber03, Cox03, KSS^{+}03b, KSS^{+}03a, Sch03]. fourteenth [Lan82]. Fourth [Dra11e, Gan72a, Han12l, Oja08a, Pun11c, Ric10a, Ros09c, Wak96]. Fowler [Dek22]. Fox [Spr77b, GP11b]. Fractals [Ios79a]. Fraction [AT15]. Fractional [CSW93, DS99]. Fragmented [LA96]. Frailty [Här08]. Frame [Cob07, dMRAMR18, SHH94, Tam82]. Framework [BD20, HGKJ20, KMT03, SAA22, SCC14, Liu21]. Francesca [Liu08]. Francesco [Pyn14b, BBS07]. Francisco [Ban14, Rob11b]. Francois [Mai11c, Unw13b, O’B12h]. Francq [Liu11]. Frank [Cro11e, Nad74, Ric11b, Sär72b, Str76, Hea95]. Franke [Emb08, Gho12h]. Franklin [Hau72, Mai18]. Fraser [Ios77d, Spe08a, Cop77]. Fraud [EIRS18, RCA18]. Fréchet [Dug74b]. Frederic [Han13f]. Frederick [Bec73, Spe11b, Hau72, Ho05]. Frederik [Dur13c]. Free [CKN^{+}24, Dra13f, Kuk20, Sen03, HM19, NN20b, NN20a]. Freedman [Spe10c, Spe10b]. Frees [Kos11]. French [AM93b, Col81, Dag76, Dam74, Dam77, EMR88, Fér72, Hor72, HL74, LN80, Lec76, Lef79, Mal97, RR90, Sau76]. Frenkel [Ng21]. Frequencies [FL83, NK14]. Frequency [Kha74, Shl07]. Frequentist [Bas16, Cox13, Efr13, Fra13, GC95, Liu23, Par13, PB17, Rob11b, Rob13, SH13, gXS13, gX13]. Fricker [Esp14b]. Friday [Sär72b]. Fried [Fin80a]. Friedman [Nor09c]. Friedrich [Han08m]. Frisch [Rei76]. Frisén [Han08f]. Fritz [Has08a, Sta08b]. Front [Ano72g, Ano72h, Ano72i, Ano73f, Ano73g, Ano73h, Ano74e, Ano74f, Ano74g, Ano75g, Ano75h, Ano75i, Ano76g, Ano76h, Ano76i, Ano77g, Ano77h, Ano77i, Ano78h, Ano78i, Ano78j, Ano79f, Ano79g, Ano79h, Ano80f, Ano80g, Ano80h, Ano81d, Ano81e, Ano81f, Ano82d, Ano82e, Ano82f, Ano83e, Ano83f, Ano83g, Ano84e, Ano84f, Ano84g, Ano85d, Ano85e, Ano85f, Ano86e, Ano86f, Ano86g, Ano87e, Ano87f, Ano87g, Ano88e, Ano88f, Ano88g, Ano89e, Ano89f, Ano89g, Ano90d, Ano90e, Ano90f, Ano91d, Ano91e, Ano91f, Ano92d, Ano92e, Ano92f, Ano93e, Ano93f, Ano93g, Ano94e, Ano94f, Ano94g, Ano95d, Ano95e, Ano95f, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97f, Ano98d, Ano98e, Ano98f, Ano99f, Ano99g, Ano99h, Ano00e, Ano00f, Ano00g, Ano01e]. Front [Ano01f, Ano01g, Ano02d, Ano02e, Ano02f, Ano03d, Ano03e, Ano03f, Ano04d, Ano04e, Ano04f, Ano05e, Ano05g, Ano05f, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano08f, Ano08g, Ano08h]. Frontier [SW15]. Frontiers [Dek10, Gho17, Pol11a]. Frühwirth [Han07d]. Fujikoshi [Pun10d]. Fujiwara [Han10a]. Fullerton [Mai16b]. fully [OC13]. Fulvia [Pyn14b]. Function [CR93, DW84, wB09, Lia20, OS73, RGSO17, Sri73, MS87]. Functional [AP94, CG12, FBGGM17, Gho09i, GHK20, You14, O’B14b]. Functionals [IOS12]. Functions [DPCM21, FFS11, Fis72, Háj72, Ios79b, Ken75, Kuk20, Kuk22, RR79b, vdL03, AL86, Cro11e]. Fundamental [ET16, FT96, Ole03, Zha17b]. Fundamentalness [ABC11]. Fundamentals [Cro11h, Han13k, Cam07, Gho22, Som76a]. Fung [Han09d]. Funke [Knü80]. Furberg [Han08e, Han08e]. Furno [Dek14]. Furrow [MW81]. Further [Hus72]. fused [SLY23]. Future [ANK03, Dur23a, ESTTD05, Hau08a, Lin14a, Mai15, Mas73, Nic01, Sta10b, Veh08d, PSKL^{+}23]. Fuzzy [BGV18, KX09, O’B13g].

G [Bar77a, Bar80a, Bar76, Ber78, Coc73, Cox75, Cra73, Cro10c, Cro10g, Cro11a, Dem83, Der77b, EV73, Eng78, Fou80b, Gan80c, Gho08g, Gho09d, Gho12b, Goo73, Gow79, GP12, Háj72, Han08h, Han10e, Han11h, Han12c, Han12l, Han13c, Han13l, Ios76a, Ios76d, Ios77c, Ken80, Ken73a, Khu13a, Kin75a, Kos10b, Mai12c, Mar07b, Mer12a, Mer20a, Mor73, Nor14b, O’B09i, O’B14b, Pos76c, Pos78a, Pos79c, Pos80b, Pre78b, Pre80b, Pun08b, Ric11a, Sco77a, Smi75, Som78f, Tha07a, Tra13, Veh11a, Wol77, Yan73, vdL80c, Dal77a]. G. [Spe85]. Gabasov [Rao80c]. Gabriel [Kos12b]. Gabriele [Sun22]. Gaetan [Nan10b]. Galambos [dO79b]. Galbraith [Han08b]. Galecki [Mai14c]. Galileo [Hal86]. Galin [Pol12a]. Galit [Han07a]. Gallhofer [Veh08c]. Gallian [Dra11b]. Galore [Ric13b]. Gama [Han12h]. ‘Gambler [Edw83]. Game [Liu22b]. games [dM84, Mer09c]. Gamma [FS08]. Gan [Han08d]. Ganapati [O’B12e]. Gang [Pun10b]. Gani [Buc77a]. Gänssler [Roh80]. Gap [KMT03, Zhu14]. GARCH [PWKO24, Liu11]. Gardner [Dra13b, Mer09c, Mer09d, Mer09c, Mer09d]. Gareth [Nor14a]. Garfield [Ros19]. Garrett [Cro09b]. Garrity [Spe10e]. Gart [Ano72c]. Garvey [Dur14a]. Gas [FMS12b, Rao93]. Gatelli [Liu08]. gathering [HGA76]. Gauss [Ald98, Ald99]. Gaussian [Dur18, Cas20, DPCM21, Gho21, LS86]. Gautier [Han13g]. Gave [Dra09d]. Gayen [Mit79]. Gearheart [Kha20]. Geary [MM77]. Gebhard [Sai08]. Geer [Cox16, Gho12g]. Geert [Cro09b]. Gelman [Pol12a]. gen [MCF12]. Gender [Ayh01, MCF12]. Gene [And07, BR00, Bla07, CC08, Som79d, Tho98, Tho00]. General [BR09, Cox86, GP73, Gri15, Hoo05, KFL76, Kos13d, Liu12a, RBR85, Sär72c, Sen76, Tal93, Val08, Gre87, SS87, HHDG12]. Generalisations [NC21]. Generalised [CO14, SH23, TT20, YAA20, DTGG24]. Generalizations [Pun07a, Yas93, O’B16c]. Generalized [BP12, BCO82, HK77, Lis09b, Liu12a, Liu14a, Lov00, Mai19, Meh75, Pre89, Pun11d, SM14, Tri09, Vid15, Cro10g, Dra08a, Lis13, Num15, Pun09, Roh76a, Tra73, Zha16b]. Generating [FRS08, Kag72a]. Generation [EH95, Goo73, MCH13, Rip83, Sow72, Sow78]. Generator [Dod96]. Genes [Cla00b]. Genetic [CD00, Ken76a, She00]. Genetics [Cro08a, Cro11d, Fis10a, San17, Spe08a, Cro11h]. Geng [Num13a]. Genome [Jon00, Tho00]. Genomic [Bra14, Sch16a]. Genomics [Fis22, Pod21a, Cro08c, Cro11d]. Genshiro [Lis08b]. Gentle [Kos13c, Pun07b]. Gentleman [O’B09a]. Genuer [Mai21]. Geof [Kos13b]. Geoff [Han12w, Ric13a]. Geoffrey [Cro10g, Häg07, Khu13a]. Geographic [O’B09g, Pod24]. Geographical [WB97]. Geometric [Gho11a, O’B10a, Sch16b]. Geometrical [GPdMLMGGSG20, Wat74, You86]. Geometry [iA21, Sto90, Bad82, BNCR86]. Georg [Elf72, Nor13b]. George [Coc73, Cox07, Cro08a, Dra09a, Dra12f, Dra13a, Dur23a, Han11e, Han12s, Han12m, Kau12, Lis08a, Liu07, Liu14b, Pol13a, Pun08d, Sco10a, Spe08a, Spe10a, Sta13a, Veh09a, Zha19b, Kem74, San93]. Geostatistical [GDSN18]. Geostatistics [AA78, All79a]. Gerald [Han10i, Kró80b, O’B09b, Var19]. Gerard [Feh07]. Gerd [Zha21]. Gerda [Lis09c]. Gerhard [Lis10c, Liu17, Sta07b, Unw12d]. Gerring [O’B07c]. Gertrude [Coc79]. Gestaltung [Dri78]. Gesundheitswesen [Lin75b]. GGobi [Mai08]. Gheorghe [Oni83]. Ghosh [Bar72b]. Gian [Han12t]. Gian-Carlo [Han12t]. Gibbs [MM01]. Gibilisco [Gho11a]. Gihman [Pin75]. Gilbert [Mar07a, Sta13a, Fin77d]. Gilberto [Pat74]. Gilchrist [Fer80b]. Gill [Aal82, CV82, Duf82, Har82, LT82, Ano87a, Gho09c]. Gilles [Han13g, Kos12b, Unw20]. Gillet [O’B12h]. Gillian [Der75b, Dra08a]. Gilthorpe [Mai13c]. Gimenez [Gho11b]. Gini [For82, GG17, OY92, Rad10]. Giorgio [Bil80a]. Giovagnoli [Fil11]. Giovanni [Cam07, O’B10c]. Giri [Roh76b, Roh76b]. Girolami [Gho13b, Sen14b]. Girshick [BRST76]. GIS [Arb93]. Giuffre [Tin78b]. Giuseppe [Hau18, Mai14a]. Given [BS89]. Givens [Kos13b]. Gjessing [Gho09h]. Glass [Cur03, All79b]. GLIM [Ait18, Jør84]. Glitters [Han08e]. Global [NSZH22, OA91, SAA22, Liu08]. Globalisation [NP97]. Globalization [Car05, Mal05]. Globally [PAF18]. Glossary [Ken74]. Glymour [Tso20]. GMM [BMM08, HM04]. Gnanadesikan [Nel72b, Loy80]. Goals [Blu07]. Göbel [Gan78c]. Gobet [Pod17b, Tso18]. Gokhale [Csi80]. Golan [Cox19]. Golbeck [Ric17a]. Gold [Han08e]. Golden [Liu21]. Goldman [O’B10a]. Goldstein [Sun17, Cro11d]. Goll [EV73]. Gonzales [Dra09b]. González [Han12m]. Good [Ano74h, Dra13f, Han13a]. Goodman [And86, And87, Cau86, Clo86, EJ86, GS86, GS87, Ken86]. Goodness [Feu16, Gel03, RB90]. Goodness-of-fit [Gel03]. Goon [Som76a]. Goos [Dra12d, Pos79c]. Gordon [Han11b]. Gore [O’B12e]. Götz [Fil09b]. Gould [Lan18]. Governance [HRR07]. Government [ANK03, CS00, FMS95]. Gower [Cox15]. Grabiner [Han12b]. Graduate [Rus10]. Graduation [DMS06]. Graham [Han12x, Mai12a, O’B07b, Kal05]. Gram [Hal00, Hoe83]. Grand [HL74]. Grands [Pre79b]. Granger [Ord80a]. Graph [BN07, Cas20, Fra80, Gal80b]. Graphic [Veh10c]. Graphical [And73d, Che24, EM19, Fis83, GP11a, Han12x, MYA95, Mod20, Pol13b, Pou02, SGLS02, SRN17, LOD^{+}85, Han13e, Sun16b]. Graphics [EM93, Esp15b, Mai08, O’B10a, Sta07a, Unw11a, Lon15, Pol11b, Sco10b, Unw12c, Unw11b]. Graphische [Dri78]. Graphs [Gho08a, Nad74, Smi80]. Grätzer [Dra12f]. Graver [Smi80]. Gray [Kos13d, Meh75, Sen14d, Sta12a]. Grazia [Liu10a]. Great [Pos78d]. Green [Ham79, O’B13d, Rad99]. Greenacre [Sta08a, Veh08b, Veh11b]. Greenfield [Dra08b]. GREG [KBM21]. Gregory [Ric10c]. Grenoble [Pre78b]. Greville [Roh76a].

Grimmett [Häg07]. Grondlagen [Mui73]. Groningen [KSS^{+}03a]. Ground [Nem73]. Group [DC19, Dur13b, HG94, Lec76]. Groupe [Cas91, Lec76]. Grouped [JK78, PS85, Spr75b]. Groups [Han13d, Har96, Hau01, Roh16, RR79a]. Growth [Blu07, Hoe74, Som78e, Som79a, O’B15a]. Gruber [Dra10b, Dra14a]. Gruijter [Veh08g]. Grünwald [Csi08]. Gruppierung [Lin75a]. Guarascio [All79a]. Guerra [Cro11d]. Guessing [RR79c]. GUI [O’B12c]. Guide [Dek22, Dim01, Dra12a, Gho08b, Han08h, Han11h, Han13b, Han13a, Kok07, Muk14a, Nel74, Pun10a, Ric14, Rus10, Sco07, Sta12d, Sta13b, Tha07a, Tha07b, Var19, Veh10a, Veh11a, Kal84, Kha20, Sta07e, Sta07d, Veh10c]. Guided [DGS18, Pod14b, SW15]. Guidelines [MM96]. Guide(R) [Sta07f]. Gumbel [HS12]. Gunnar [Bje72]. Guo [Ken10, Lin14b]. Guo-Liang [Ken10, Lin14b]. Guojun [Han08d]. Gupta [Gho12c, Ken73a, Pos80c, Som76a]. Gusak [Han10j]. Gustafson [Haa12, Pun12b]. Gustav [Kro80c]. Gut [Cro10a]. Güte [vdL80b]. Guttman [Pal72]. Guttorp [Cox12]. Guyader [Unw13b]. Guyon [Nan10b]. Guzman [MG05].

H [Bil77a, Buc76a, Buc76b, Cac74, Cha80, Cox80, Cro09a, Cro09c, Cro10g, Dra10b, Dra14a, Fin80b, Gan78a, Gho08f, Gho13b, Goo78, Ham72a, Han13m, Ho05, Hoe83, Huz80, Ios76c, Ken73b, Ken73a, Knü80, Kos13b, Lin75b, Liu07, Liu14b, Mai11f, Meh75, Mer09b, Moo78, Mui73, Nel80, Num19, Oja08b, Ord72, Par05, Pat75b, Pod22a, Pol13a, Pol13d, Pos78c, Pos79b, Ros09c, Rus74, Sar73, Sco77b, Sco07, Spr80, Sta08c, Sta12e, TW92, Tin80b, Tip75b, Tip75a, Tra13, Veh09b, Vin73, Wil75, YK20, Dal77b]. h-Likelihood [YK20]. H. [Dav82, Hal73]. Håkon [Gho09h]. Haan [Mat73]. Haddon [O’B12g]. Hadi [Pun13c]. Hadlock [Dra13e]. Haenszel [Dar81]. Hafner [Emb08, Gho12h]. Hagan [Ric18]. Hagenaars [Kos10c]. Hahn [Han10i, Var19]. Hahne [O’B09a]. Hahs [Han12c, Han12l]. Hahs-Vaughn [Han12c, Han12l]. Haipeng [Kos09b]. Haitovsky [Spr75b]. Hájek [Žal75]. Hak [Dur17]. Half [Jon07, SGLS02]. Half-Sib [SGLS02]. Hall [Dur23a, Fis22, Fis24, Gay11, Gho22, Kha24, Liu18, Liu21, Liu23, Mai18, Mai22, O’B21, Pod17b, Pod17c, Ric17b, Slo23, Sun16a, Sun18, Sun24, Tso18, Unw21, Unw24, Zha17a, Zha17b, Zha18, Zha19b, BPV21, Pod21a, Pod22b]. Hall/CRC [Dur23a, Fis22, Fis24, Gho22, Kha24, Liu18, Liu21, Liu23, Mai18, Mai22, O’B21, Pod17b, Pod17c, Ric17b, Slo23, Sun16a, Sun18, Sun24, Tso18, Unw21, Unw24, Zha17a, Zha17b, Zha18, Zha19b, BPV21, Pod21a, Pod22b]. Hallin [Gen24]. Halloran [Aur10]. Halmos [Spe10a]. Hamada [Gho09d]. Hamaker [Moo78, Goo78]. Hamann [O’B12f]. Hammett [Els18]. Handbook [Cro11e, Eva07a, Han10h, Joh15, Lis08a, Rob11c, Ros19, Sen14c, Som78e, Som79a, Tip80, Bre79c, O’B11a, BPV21, Dur18, Gho13b, Gho13a, Han12g, Hee16, Kos10b, Kos13c, Mor80a, Pol12a, Pun13a, Rob12a, Veh08e]. Handle [Ize03]. Handling [Esp14a]. Hands [Kha20]. Hands-On [Kha20]. Hanna [Rou07]. Hannah [Ros09b]. Hanoi [Mer09c]. Hans [Dri78, Lin75a, Ric14, Spr77a, ERS77, Für73a, Mor80b]. Hans-Michael [Ric14]. Hansen [Ano94d, Gro95, Mar07b, Mar96, Wak94, Wak95, Wak96, WG97, Pos76d, TW92]. Hansheng [O’B08c]. Hanson [Han11a]. Hao [Mai11d]. happened [GT12]. Harald [Kin75c, Lea88]. Hard [Sti14]. hardback [Dek22, Fis22, Fis24, Kha24, Liu22a, Liu22b, Mai22, Pod17b, Pod17c, Ric17b, Sun16a, You20, Zha17a, Zha17b]. hardcover [Alb19, BPV21, Blu18, Blu20, Blu21, Cap19, Cox19, Dur17, Dur21, Dur23a, Dur23b, Els18, Els20, Fis18, Gho17, Gho20, Gho22, Hee16, Kha17, Lan18, Liu17, Liu18, Liu19, Liu20b, Liu21, Liu23, Mai16b, Mai17, Mai18, Mai19, Mar21, McE17, Mer18, Mer20c, Mer20a, Mer20b, Ng21, O’B20, O’B21, Pod18, Pod21a, Pod21b, Pod22a, Pod22b, Qiw20, Ric18, Rob19, Ros19, Rui17, San17, Sen23, Sun18, Sun19, Sun22, Sun24, Tso17, Tso18, Unw20, Unw21, Unw24, Var19, Var20, Zha16b, Zha18, Zha21]. Harding [Dek10, Pos76c]. Härdle [Emb08, Gho12h, Kos12c, Kos13c, Veh08e, Kos12a]. Hardness [Pos73]. Harezlak [Num20]. harm [O’K12, Rei12, Ski12a, Ski12b, Wil12]. Harmonization [Des00]. Harold [Cor77b, Mor80a, RR79b, Ald05, Hoe76]. Harring [Han13d]. Harry [Gho09d, Kle76, Lin80c]. Harter [Buc76a, Buc76b]. Hartley [Dav82, Huz80, Ken73b]. Haruo [Pun11d]. Harvard [Blu18]. Hase [Han11h]. Hassan [Sar72a]. Hassanien [O’B09b]. Hasselby [Sch80a]. Hastie [Nor09c, Nor14a, Sun19]. Hau [Dur17]. Haughton [Han13i, Han13i]. Hauschke [Mat07]. Havbro [Var14]. Hayes [Cro09a, Háj72]. Hays [Han13l]. Hazard [Kuk20, Kuk22]. Hazards [Bre75, Cro08d, Dur22, Kuk20]. hbk [Liu23, Sun24]. Health [Ano76m, CB00, Cap19, Doc22, DSC03, FFBK76, Gho10a, Gre06, Hee16, HK77, MBW^{+}23, Ple15, Sun17, SAA22, VL96, Dam74, Joh15, JF14, Ney75, Gan80a]. Healthcare [Yoo21]. Heaven [Haa12, KBH10]. Heavy [Tsi03]. Heck [Veh09b]. Hector [Cam77]. Hedayat [Han13o]. Hedges [Pol13b, Ros09b]. Heeringa [Cro10c]. Heiberger [Sco10b]. Heinrich [Für76]. Heinz [Liu20b]. Held [Sch80a]. Helen [Mai11d]. Heleno [RBB^{+}21]. Helga [Han08l]. Helge [Pun07a]. Helland [Cox10]. Helle [Sko12]. Heller [Dra08a]. Hendrix [Pol13b]. Hendry [Dei83, Eng83, Gra83, Klo83, Miz83, New83]. Henle [Dra13b]. Henri [Dug74a]. Henrik [Liu12a]. Henry [Gho11a, Gho13a, Mer18, Sen14d, Bil77b, Leh78]. Henschke [Fin77b]. Henseler [Han12g]. Henson [Han11d]. Heping [Mai11f]. Herbert [Gho08f, Gho09b, Pos76d, Duf78]. Hermann [Lin75a, Shi78]. Hermans [Cha80]. Herrera [Esp14a]. Hertz [Gan72a]. Herzog [Has08a]. Hesketh [Eva07a]. Hesterberg [Sco12b]. Heston [Kha80]. Heterogeneity [DC19, TO17]. Heteroscedastic [BS85]. Heteroscedasticity [Buo06]. Hettmansperger [Oja11]. Heumann [Pun07a]. Hexaflexagons [Mer09c]. Heyde [Gay11]. Hidden [Cro09d, Mar11, MP21, Cro11b]. Hideaki [Han10a]. Hierarchical [CDM19, Gho09a, NSP21, O’B13e, Tut21, Wik03]. Higgins [Ros09b, Sch78]. High [CD18, GOZV94, Gho10b, Gho12g, GM19, Han13g, Kha15, Kha24, Koc14, LD24, Liu10b, Mod20, Nov73, Oja07, Pun10d, You15, ZYW23, Gir15, ZT13, You20]. High-Dimension [YAA20].

High-Dimensional [Gho12g, Han13g, Kha15, Koc14, LD24, Mod20, Oja07, Pun10d, Gho10b, You15, Gir15, ZT13, You20]. High-Risk [GOZV94]. Higher [HG94, HH95, LL87]. Higher-order [LL87]. Highly [Veh13b]. Hilbe [Han11g, O’B08b, O’B11b]. Hildebrand [Pos78c]. Hill [DIV17, Gho23, Kró80a, RR79a]. Hillier [Bec73]. Hills [Cor77c]. Hine [Pos78a]. Hinkelmann [Fil09a, Dra12c]. Hinkley [Pos76e]. Hiroaki [O’B11c]. Hiroshi [Han10a]. Hirotsu [Hir05]. Hiroyuki [Han13n]. Hirsa [Kos13a]. Histograms [Joi75]. Histoire [Hor79]. Historian [Han12b]. Historical [Hor72, IKS^{+}81, Mis08, Ros10a, Sti14, Veh08d]. Histories [Sun15, Wil14b]. Historique [Hor72, San77]. Historisch [San77]. Historisch-Demographische [San77]. History [Ait18, Edw74, Gho09h, GR01, Hal00, Ken76c, Mis10, GT12, KM80, Dra09d, Spe07]. HIV [BD94, SJFM95]. Hjort [Cro11a, Lis09c]. Hoaglin [Spe11b, Sen11c]. Hobson [Sta11a, Sta11b]. Hoc [MDG^{+}92]. Hodges [SJ89]. Hoel [Der77b]. Hoeting [Kos13b]. Hoff [Kos10a]. Hogg [Lis09a]. Höglund [Kos09a]. Höijer [Wol76]. Hoijtink [Gho09b]. Højsgaard [Han13e]. Hold’em [Han13f]. Holland [Han08g]. Hollander [Kró80d]. Hollberg [Lin75b]. Holmes [Cro11a]. Homer [Els20, Fis09]. Homogeneity [KS22, SM19, HMT22, WLC82]. Hongyu [Gho13a]. Honor [Csi08, Huz80, Pol11a]. Honorary [Ano72j, Ano74h]. Honour [Buc77a, Fil09b, Ios77c, Ken75, Kin75c, Lis10c, Tin80b, Ada80, HR21]. Hooker [Kru89]. Hopkins [Dra13b]. Horgan [Mai10b]. Horng [Gho13a]. Horng-Shing [Gho13a]. Horses [Tan97]. Horseshoe [BDLP20]. Horton [Liu18, Pol11b]. Horváth [You14]. Hot [AL10]. Hotelling [Hoe76, KS89a]. Hothorn [Ric11b, Ric17b, Rob11c, Unw12a]. houh [Veh08e]. House [CS00]. Household [CS02, Han13i, HRR07, KBM21, LO04, Uni78, ZTK13]. Housing [Ano78k, TG94]. Hout [Han12d]. Howard [Han13b, Han13m, Ios72, WG97]. Hox [Szc11a]. Hu [Kha17, Han09d]. Hua [Nor13a, Num14, Zha18]. Huber [O’B09a]. Hubert [Han13b]. Huck [Ric09]. Hugen [Pod18]. Hughes [Han11h]. Hugo [Sco13b]. Huh [Huh05]. Hui [Pol11a]. Huijbregts [All79a]. Huiwen [Han12g]. Huizingh [Eva07b]. Human [Cla00b, Goo78, Has08d, Hoe74, Spe08a, Sta08b, WRH05]. Humanity [Gho23]. Hunted [Rob12b]. Huntington [Slo23]. Huntington-Klein [Slo23]. Hüsler [Cro11c]. Hussein [Sar72a]. Husson [Mai11c, Unw13b]. Huy [Liu22b]. Huygens [Hol84]. Huzurbazar [Tin78c]. Hybrid [PWKO24, ZZLZ15]. Hyndman [Ano21, Han09b]. Hyper [KS22]. Hyper-sphere [KS22]. Hypercubic [Lov00]. Hypergeometric [Mor80a, RR79b]. Hypothèse [Lef79]. Hypotheses [Han12q, OS73, Pér93, SC05, Hoi13, LN78, Bar72b, Gho09b]. Hypothesis [BS89, BM91, BR02, Cau80, KX09, SB17, Bic13, Lef79].

I. [BJS92, Lin76, Sen98]. Iacus [Luo09]. Iain [Cro11b, Gho10b]. Ibern [Dag75, Rio76]. IBM [Sta12d, Sta12a]. ICS [IOS12]. ICT [Pow03, JS03, Kit03]. Ideal [Jaz72]. Ideas [BD76, Cor77a, GO16, Han11a, Som77a, Som79b]. Identifiability [Cro94, O’B21]. Identifiable [SLWH22]. Identification [Ald02, KY91, NZY18, Smi20]. Identify [AT15, Nar74]. Identities [Hop93, Pan07]. Identity [HS90, HS12]. Idiot [HY01]. Ieno [O’B07b]. Iezzi [Rod21]. Ifan [Han11h]. Ignorable [CMH19, ZL22, Hed20, MP23, Sik18]. Ignorance [PW91]. Igual [Pod17a]. II [Bar76, Eng78, Gho11d, Kim14, Pos80c, Roh76b, Tip75a, Val08, Wil73, Ano76m, Dal77b, Fed81a, FFBK76, Har74b, IKS^{+}82, KM79b, LSS80a, MM96, New84, Seb92, Sen72, Tho78b]. III [Sen11c, Dal77c, Fed81b, Har75b, KM79c, New88]. Ikeda [Han10a]. Ikuho [O’B09g]. Ilia [Ng21]. Illian [Han08l]. Illicit [Ize03]. Illustrations [Pod24]. ils [Mal97]. im [Lin75b]. Image [Gho12e, SFVP^{+}24]. Images [Mar07b, AM93b]. Imaging [LEGW01, O’B09b]. Impact [FM95, GMZ03, Hor91, JS03, Nay09, Veh13b, Pea74]. Impacts [Gan78a]. Imperfect [Chi22]. Implementation [Abr05, Bra15, Gre06, O’B08a, Wan16, Cha08, LOD^{+}85, Han12p, Sun16a]. Implementations [HWS20]. Implementing [Gal02b]. Implication [HMA^{+}18]. Implications [Gol97, PPMS11, Rid16, Rum02]. implied [AL86]. Import [BS85, BCO82]. Importance [Kru89]. Important [Cra81, SJ89]. Improbable [Veh13b]. Improper [Nie03a, Nie03b]. Improve [AG99, DFV99a, DFV99b, Mac97, TC99]. Improved [CPB94]. Improvement [Box93, HH95, San93, Sne99b, Wil95]. Improving [Bat02, CS00, Hou88, Hug06, MS23, MCH13, SBB^{+}12, Tre05b, Sun17]. Imputation [AL10, AT15, BH14, HCR18, HB07, MR03, Muk15, Nie03a, Nie03b, Nor14b, Nor98, Rei09, Rub03, Zha03]. Imputed [BE17]. Including [GO16, Hor16, Mac16, Men16b, OG16, Wil16a]. Inclusion [Hop93, Til98]. Inclusion-and-Exclusion [Hop93]. Income [DTGG24, Esp15b, VF00, Lon15, Pod72]. Incomplete [Kag72a, Kag72b, LSD^{+}11, Pat88, PPMS11, Zha10, Lin14b]. Incomplete-Block [Pat88]. Inconsistency [Par80]. incorrect [BH80]. Increasing [SFVP^{+}24, Puk87]. Increment [FS08]. Independence [DC83, DGS18, TSR96, WC98]. Index [Ano07e, Ano08b, Ano09b, Ano10b, Ano11b, Ano12b, Bal95, Bar72c, CIS06, Der77a, Fou80a, Jaz72, Nie04, Rad10, SDGD^{+}11b, DTGG24, Cra80]. indexed [Nau79]. Indexes [Har76, Rad10, Whi99]. Indexp [Ano13b]. India [DM12, GMRS99, LS04, Nay09]. Indian [Som78f]. Indicateur [Dam74]. Indicators [CP07, CP08, MSS20, NWE16, Rad99, SBB^{+}12, VT19, dV01, SS17]. Indices [Ani08, Dal98, Fer94, Ham72b, Smi21]. Indici [Bil80a]. Indirect [Has08b]. Individual [GTF18, Fis72]. Indrayan [Num09b]. Induction [RR79c]. Industrial [Box93, DR93, Dra14c, Dur22, FPD05, Pre77a, San93]. Industrie [EV73]. industriellen [Kri74]. Industries [Tip75b]. Industry [Abr05, Ast94, Don12, FPD05, HHDG12, STNW93, All79a, Bar75b, Cra73, Dra08b, Han10i]. Inequalities [BCG07, Hop93, O’B10e, Pun11b]. Inequality [BJO^{+}98, DTGG24, Esp15b, Rad10, Sen98, VF00, Hoi13, Lon15, Sha88]. Inevitability [Lin82b, Lin82a]. Infeasible [ZYW23]. Infection [KMM^{+}17]. Infectious [Kin77b, TZW^{+}20]. Inference [Alv20, Ano76f, Ano93d, BH77, Bar76, Ber03, Bic92, Bic14, Bla07, Blu16, Bow04a, BD20, BCD^{+}94, Bre99b, BBvdB18, CO14, Cox03, Cra81, Daw02a, Daw02b, Fil09b, Fin79, For82, Fra80, FPR^{+}16, GP73, Gho20, Gho22, GKV19, GKV20, GC95, Han11k, Han11j, Han13n, Hel04, Hjo92, HK16, HMW21, KSS^{+}03b, KSS^{+}03a, KT21, Koe16, KLTG21, Liu11, Liu12b, Liu19, Luo09, Mai09a, MM01, MR03, Men16c, Nad74, NC21, Nor09c, Oli14, Ped78, PB17, PCA^{+}19, Pos79b, Pur72, Rao80b, RYG^{+}21, RC15, RG23, Rob19, Sch78, Sch03, SB04, SAG^{+}16, Smi75, Smi16, Sør04, Spe10b, Spe11a, Sun19, Tri09, Tso20, Vin73, WS16, WR05, WPRP17, WS21, YWH16a, YWH16b, ZL22, Zha19a, ZHW23, ZZLZ15, BT75, BT76, Don86, Hig77, MPR12, RLS23, WGKL22, Bar77b, Blu18, Cac74, Dra13c]. Inference [Gho11c, Hau08a, Høy72, Ios76c, Num09a, Rob11h, Sil73, Wet75, You19, Cop77, Esp15c, Fra72, You14]. Inferences [Bar75d, BH80]. Inferential [Ber06, GR90, PPMS11]. Inferring [Tho98]. inflated [HMT22]. Inflation [Tin80b]. Influence [Bel09, Bor09, Daw02a, Daw02b, Gho08b, Hol84]. Influences [Cam97]. influenza [NSZH22]. Info [Cox19]. Info-Metrics [Cox19]. Informal [Pun10a, Kal84]. Informatics [Pos79c]. Information [Ald07, iA21, AT15, Ano73q, Ano74q, Ano75r, Ano75s, Ano76p, Ano77n, Ano78p, Ano79m, Ano80k, Ano81g, Ano82g, Ano83h, Ano83i, Ano84h, Ano84i, Ano85g, Ano85h, Ano86i, Ano86j, Ano87h, Ano87i, Ano88h, Ano88i, Ano89h, Ano89i, Ano90g, Ano90h, Ano91g, Ano91h, Ano92g, Ano93h, Ano94h, Ano95g, Ano96h, Ano97g, Ano98g, Ano99i, Ano00h, Ano01h, Ano02g, Ano03g, Ano04g, Ano05i, Ano06h, AES18, ADE^{+}22, BK20, BCF^{+}14, BJS92, Cro11d, Csi80, ESS10, ES06, FWS14, GP03, Han08j, KMNS05, Ken73c, NP97, Ole03, Puk87, Rad99, Ram03, RCA18, WB97, Wyc03, ZL22, JOT85, Pow03, Bil80b, Ios79b, Lis07c, Lis08b]. Informative [Gho09b]. Inge [Cox10]. Ingram [Cro08b, Pun11b]. Ingrassia [Dur13b]. Initialization [MP21]. Initiatives [TC15]. Input [Bil77b]. Input-Output [Bil77b]. Inputs [Har99]. Inquiries [BD76]. Inquiry [Han13l]. Insights [GTF18]. Inspection [CR93, CW73]. Inspirations [Pol12b]. Inspired [Til22]. insr.12432 [Ano21]. Institute [DD80, Pun11a, Ric13b, RBB^{+}21, Sun19, Ano72j, Ano72b, Ano86h, Pat75a]. Institutions [Han07h]. Instruction [Sch07]. Insurance [Dra08a, Gri15, Hoe83, Kos13d, Kos12c]. Intangible [Gol97]. Integral [SM93, TP73]. Integrals [CW91, Mor80a, NK08]. Integrate [Rei09]. Integrated [Fri93, Han08j, Han11k, HW06, Keu97, KMT03, NK14, O’B10a, Sco09a, AL86]. Integration [Col99, DD80, KT21]. Integrity [Sch80a]. Intellectual [HV15]. Intelligence [Han07a, O’B09b, Yoo21, ZZN^{+}22, Mai11a]. intensity [Jac84]. Intensive [Coo07]. Interactifs [RR90]. Interacting [Gri87]. Interaction [Cox84a, Cox84b, DFG^{+}94, SKK18]. Interactions [FMS95, LG08].

Interactive [NL07, RR90, Mai08, Pre80a]. Intercept [AS72, SCC14]. Interesting [Rou07]. Interface [Sco10b]. Interim [Ano78k]. Intermediate [Cro10a, O’B16a, Pos79d]. Internal [Teu05b]. Internat [And87, And77a, BT76, GS87, Goo78, Hoy76, Moo78, Ott76, dBPR86, Pos78b, Sch78]. International [Ano72a, Ano21, BJPT99, Car93, Car05, Deb78, Fer94, Fis05, Gan72a, Had10b, Hof97, Hor72, Hug06, Liu17, Mar21, McE17, OS73, Pol12b, RBB^{+}21, ST92, Som78g, Sti00, Sti05, Tsi03, Tso16, Mal97, Daw02a, Ano72j, Ano72b, Ano86h, Cha16, DD80, Had10a, HN13b, Kha80, PA15, Pin10a, Pin10b, Ros19, SZ09]. Internationale [Hor72, Mal97]. Internationalization [Sen98]. Internet [Ber17, TRRL23]. Interobserver [Han13j]. Interpoint [GM19, SM19]. Interpolation [DCC98]. Interpretable [DTP^{+}20, SLWH22]. Interpretation [Cur03, GPdMLMGGSG20, GW14a, GW14b, Hub14, Lis14, RM93, Sco77b, Wil14a, Fin75]. Interpreting [BM91, FCE^{+}03]. Interrelated [Bec73]. Intertwining [Lan01]. Interval [DS22, Jon07, Sco12a, SB17]. Interval-censored [DS22]. Intervals [CMH19, Han12w, KL09, LK16, Lis09b, NN20a, NN20b, PW91, Var19, Roy86, Sco12a]. Intervention [Har96, SFG10]. Interview [AMR23, Ano76m, Can95, Cit16, Cox15, DS20, Fis10b, FvZ15, FFBK76, GN14, HS19, He17, Nar20a, Red74, SV22, Sch95, Sin95, Sta17, Wil16b, RS23]. Interviewing [WD84]. Interviews [Lan76b, Pun11a]. Intra [EUC09]. Intra-Cluster [EUC09]. intraclass [Don86, TTB86]. Introducing [Nem73, SAA22, Sco10a, Veh10c]. Introduction [AP94, BG03, Blu18, Cro11b, Dem73, DR93, Fil09a, Fis24, Fon11, Gir15, Gut95, Han11a, Han12c, Han13c, Hoy73, Ios77d, Jau14, Kha73, Le 90a, Lee01, Lep14, Lis07a, Lis07b, Liu20a, LG04, Mai10b, MPJS03, Nor14a, O’B10a, O’B13a, O’B13c, Ord80a, Ple15, Pos76a, Pre79a, STNW93, Slo23, Spr77a, Tre00a, Unw12a, Unw24, Val08, Veh09a, Veh09b, You19, Zha19b, Bad82, Hoy76, Ish81, JF14, Roh76b, Emb08, Gan78b, Gho12h, Kró80d, Mai13b, O’B12f, Sta10c, Sta12c, Var20, Blu16, Cac74, Cor75a, Cor77b, Esp14b, Fin77b, Fin77a, Fin80a, Gho08e, Gho08d, Gho12b, Han10c, Han11e, Han13f, Ios76c, Ios79c, Khu13a, Lin80b, Liu12a, Man08, Mer20a, Num09a, O’B09c, O’B09d, O’B13e, Oja07, Pod17a, Pos80a, Roh76b, Roh78, Ros09b, Sai08, Sen11a, Sen15, Sko12, Tra07]. Introduction [You15]. Introductory [Alb19, Eag74, Fin80b, Mai10a, Nor09b, Rus74, Som78c, vdL80c]. Invariance [Hey81]. Invariant [IOS12, NA16]. Invariantes [Cas91]. Invention [Dur23a]. Inventories [Has08c]. Inverse [Ald00, Dem18, Jef74, Pun11d, Tra73, LS86, Han13g]. Inverses [Liu14a, Roh76a]. Inversion [CKN^{+}24]. Inversion-Free [CKN^{+}24]. Inversive [EH92]. Investigate [Pau19]. Investigating [Gar93]. Investigation [Bic00, Hoe74]. Investment [Som78f]. Investment. [Rui72]. Investments [Bec73]. Invited [BBG^{+}76, Fel04]. Involvement [Fis09]. Involving [Gan78c]. Ioan [Csi08]. Iosifescu [Cor75b, Gan80c]. Ira [Aur10]. Irene [Gho09b]. Irini [Han08b, Han13h]. Iris [Mat07]. Irit [Sta12f].

Irmtraud [Veh08c]. Irving [Kha80]. Irwin [Bar84]. Isaac [Ken74]. Isaacson [Haj77]. ISBN [Alb19, Ana21, BPV21, Blu18, Blu20, Blu21, Cap19, Cox16, Cox19, Dek22, Dur17, Dur18, Dur21, Dur22, Dur23a, Dur23b, Els18, Els20, Fis18, Fis19, Fis22, Fis24, Gho17, Gho20, Gho22, Gho23, Hau18, Hee16, Kha17, Kha20, Kha24, Lan18, Liu17, Liu18, Liu19, Liu20a, Liu20b, Liu21, Liu22a, Liu22b, Liu23, Mai16b, Mai17, Mai18, Mai19, Mai21, Mai22, Mar21, McC22, McE17, Mer17, Mer18, Mer20c, Mer20a, Mer20b, Ng21, Num20, Num23, Num19, O’B20, O’B21, Pod17a, Pod17b, Pod17c, Pod18, Pod21a, Pod21b, Pod22a, Pod22b, Pod24, Qiw20, Ric17a, Ric17b, Ric18, Rob19, Rod21, Ros19, Rui17, San17, Sen23, Slo23, Slo24, Sun16a, Sun17, Sun18, Sun19, Sun22, Sun24, Tso16, Tso17, Tso18, Tso20, Unw20, Unw21, Unw24, Var19, Var20]. ISBN [You19, You20, Zha16b, Zha17a, Zha17b, Zha18, Zha19b, Zha21]. ISBN-10 [Liu22b]. ISBN-13 [Liu22b]. Ishwar [Gay11]. ISI [Ano05a, FL05, Ano76a, Fel04, Gan79, Ole03, Teu05b]. Ismael [Mai11b]. Isotalo [Liu07, Liu14b, Pol13a]. Isotonic [Vin73]. Israel [Der75a, Roh76a]. Issue [BR00, Gho10b, Had10a, Had10b, SS09a, STNW93, STM95, HRR20, HR21]. Issued [Ano72q, Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76o, Ano76n, Ano77l, Ano77m, Ano78m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j]. Issues [Bai85a, Ber03, Bic00, CB00, Cox03, DK87, EB92, Fis01, GMZ03, Han12e, Han12j, HL01, Ize03, Kaf15, KSS^{+}03b, KSS^{+}03a, LEGW01, Mar74, ME93, Rao93, Sch03, SJFM95, Smi01, Zha18]. Italian [Dur13b, Bor09]. Italy [BCG07]. Item [CH19, GP11b, Tut21]. IV [Buc77b, Pos78b, HM04, Har75c, Har75a, KM80]. Iverson [Pol12b]. Ivo [Cap19]. ix [Alb19]. Iyetomi [Han10a]. Izenman [Pun08c].

J [Ano21, Aoy72, Arm85a, Aur10, Ban14, BCL^{+}85, Bar80a, Bat85b, Ber78, Bin07, Bre80, BJS92, Buc77a, Cac74, Cha80, Che85b, vC75, Cor75a, Cor77b, Cro08c, Cro09a, Cro10g, Cro11e, Dag75, Dal72, Dra10b, Dra14a, Dur14b, Eng78, Fin77c, Fin80b, Fis18, Gan72a, Gan80a, Gho08g, Gho10b, Gho11b, Gho13b, Gho17, Háj72, Ham79, Han08n, Han08i, Han08m, Han08b, Han08g, Han09b, Han10i, Han13h, Has08a, Hoe74, Hoe83, Hor79, Ios76c, Ios76d, Ios77c, Ios79b, Ken76b, Kha73, Kha74, Kin77b, Kos13d, Kró80b, Lin80b, Liu14a, Liu18, Lou85, Luo07, MM77, Mar21, Mer07, Mus75, Nel72b, Nor09a, O’B11a, Ord80a, Pet85, Pod14a, Pod18, Pol11b, Pol12b, Pos76a, Pos78a, Pos78c, Pos79b, Pre78b, Ric18, Rob11b, Rob12a, Ros09c, Rui72, San77, Sch78, Sen98, Sen03, Som78d]. J [Som78e, Som78c, Som80b, Spr75a, Spr77b, Spr80, Sta07g, Sta08b, Sta10c, Sta12e, Sta13b, Sto75, Stu72, Szc11a, Tin78d, Tin80b, Tip77, Unw11b, Val08, Var19, Veh08g, Vin73, Wet80, Whi85, You19, You20, dO79a]. J. [Bar84, CWCT19, Dei83, Eng83, Gra83, Klo83, Miz83, New83, RAC83, SJ89]. J.-F. [Dei83, Eng83, Gra83, Klo83, Miz83, New83]. Jack [Boo79, Smi80, Wil16b]. Jackknife [BGV18, Meh75, PS80]. Jackman [Nan10a]. Jacob [Oja07, Kle78, Sen14a, Sti14, Syl14, Wei84]. Jacobs [Für76, Pos80a]. Jacquard [Ken76a]. Jacques [Kos10c, Mar07a, Mal75]. Jae [FvZ15, Lee05]. Jagers [Kin75c, Rad77]. Jahn [Bje72]. Jake [Ric10b]. Jakob [Dur13c]. James [Han12f, Kha20, Kos13c, Mai12c, Mar07b, Nor14b, Nor14a, O’B09d, O’B12d, Pol11a, Pun07b, Ric13b, Ros09a, Sta07f, dO79a, GN14]. Jamie [Dur18]. Jan [Dur13c, Liu20b]. Jan-Frederik [Dur13c]. Jana [Sta08b]. Jane [Blu07, Han08b, Mai10b]. Janine [Han08l]. Janos [dO79b]. Jansen [Pod22b]. Janssen [Här08]. Japan [Fis09, Kit03]. Japanese [Els20]. Jarkko [Liu07, Liu14b, Pol13a]. Jarmo [Liu07]. Jaroslav [Žal75]. Jaroslaw [Num20]. Jasjeet [Spe10c]. Jason [O’B13d]. Jaworski [Kos12a]. Jayant [Des05]. Jean [Cha14, GP11b, Liu11, Mai21, Mar07a, Ric10a, Sap15, Ves73]. Jean-Jacques [Mar07a]. Jean-Louis [Sap15]. Jean-Michel [Cha14, Liu11, Mai21]. Jean-Paul [GP11b]. Jef [Fis10b, Teu05a]. Jeff [Gho09c, O’B12f, Ros10a]. Jeffrey [Alv11, Han13d, Pod14a, Pun13a]. Jeffreys [Ald05]. Jégou [Unw13b]. Jen [Dur17, Gho09f]. Jen-pei [Gho09f]. Jenkins [Tin78e, NSW72]. Jennifer [Kos13b]. Jens [Kos10b, O’B09f]. Jens-Peter [Kos10b]. Jensen [Gho08a]. Jérôme [Mai11c]. Jerald [Dra14d, Han08n]. Jeremiah [Mai10c]. Jeremy [Unw13a]. Jerome [Nor09c, Sun22, Szc11b]. Jerry [Wil16b]. Jerzy [Ano74h]. Jessen [Som80b]. Jessica [Liu08]. Jewell [Tso20]. Jialiang [San17]. Jiang [O’B09b]. Jianguo [He14]. Jianhong [Han08d]. Jianmin [O’B09b]. Jim [Bau14, Rob11d, Sta08c]. Jin [Kos13c]. Jin-Chuan [Kos13c]. Jixian [Sun16a]. JMP [Ric13c]. JMP(R) [Sta07c]. Joachim [Für76]. Joan [Ros19]. João [Han12h]. Joaquim [Veh07a]. Joaquín [Rio76]. Job [CB00]. Job-related [CB00]. Johannes [Lis09a, Luo07]. Johansen [Wed80]. John [Alb19, Cro08d, Cro10h, Dra11a, Dra12b, Els18, Els20, Gho13c, Han07c, Hee16, Hol08, Kró80b, Lan18, Liu20b, Mer17, Mer20a, Nor09a, Num19, O’B07c, Pod24, Pun08a, Pun09, Pun12a, Ric18, Sen11c, Som78g, Som79e, Spr80, Sta07a, Sun22, Tso20, Unw11a, Var19, Var20, You19, Bel89, Cox15, Hav14, SV22, Bel16, Mer15]. Johnson [Buc74, Han11a, Hee16, Ios78, Som77b, Dra09a, Pos80b, Som77b]. Johnstone [Gho10b]. Joiner [Bar72c, Bre79c]. joins [Cha16]. Joint [FST19, DFM^{+}22, MMC13]. Jonathan [Han11d, Han13i, Mai09c, Mai14b]. Jones [O’B09c, Pol12a, Rob19, Dra12d, Val08]. Jong [Dra08a]. Joop [Szc11a]. Jörg [Han12p, Han12g]. Jorma [Csi08, Lis07c]. Jose [Liu22a, Han12m, Jer78]. Joseph [Blu21, Coo07, Dra11b, Gho13c, Han11g, Kró80b, Lis12, O’B08b, O’B11b, O’B14c, Oja11, Pla75]. Joshua [Ros14]. Josse [Unw13b]. Journals [Bar72c]. Journey [Han11c]. Joyce [Lid80]. Joys [Pod22a]. Jr [Aur10]. Jr. [Szc11b]. Juan [Han12m]. Judea [Gho11c, Tso20, Sin02b]. Judith [Han12b, Spe11b]. Juditsky [Gho20]. Julia [Fis19]. Julian [Pun08c, Ros09b]. Julie [Unw13b, Auk78]. Julien [Mai09b]. Julio [Lis10b]. Julious [Ros10c]. Julius [Ano79i]. July [BBS07]. Jump [WS01]. Jumps [PS13]. Jun [Liu23, Mai16b, O’B08c, Zha17a]. Jung [Kim14]. Junghenn [Sco13b]. Jürg [Cro11c]. Jürgen [Emb08, Gho12h, Sai08, vdL80b]. Justification [Bow04a, SB04]. Justify [GPdMLMGGSG20].

Kadane [Lis12]. Kagaku [Els20]. Kai [Ken10]. Kailath [Tin80a]. Kalbfleisch [Wil16b]. Kali [Der77a]. Kalman [Piz10]. Kalof [Han10c]. Kaltenbach [Ric14]. Kalton [Kal05]. Kam [Han09d]. Kamal [Pod14a]. Kamp [Veh08g]. Kang [Mai13c]. Kaplan [Liu18]. Kapteyn [SS09b]. Karen [Mai12c]. Karian [Rob12a]. Karin [Ohl77]. Karl [Gho08e, Gho12h, Haa12, Kos12c, Kos13c, Num13a, Pun12b, Ano75j, Bel09, Bly94, Bor09, FZ09, GL09, Mag09, Nay09, Pla83, Sen09, SS09a, Dav09]. Karl-Rudolf [Gho08e]. Kass [Han12d]. Kass-Hout [Han12d]. Kath [Sco07]. Katherine [Cit16]. Katie [Ros19]. Katsafados [Pod21b]. Katz [Dra09d]. Kazakos [vdL80d, vdL80d]. Keen [Unw11b]. Keener [Han12u]. Kei [Pun11d]. Keiding [Aal82, CV82, Duf82, Har82, LT82]. Keith [Dra12a, Han09b, Pol13c]. Kellerer [ERS77]. Kelley [O’B09d]. Kelly [Han13m]. Kempthorne [Fil09a]. Ken [Pol11b]. Kendall [And78, Bar76, Coc73, Pos76c, Bar77a, Cra73, Mor73]. Kendig [Dra12a]. Kenneth [Mai19, Rob11f, dO79a]. Kenward [Nor14b]. Kenya [SMW79]. Kept [Spe08b]. Kernel [GM91, KSS14, RDM^{+}19]. Kesar [LgX13]. Kevin [Mai11a, Tha07a, Unw11b]. Keying [Pol11a]. Keynes [Rob11a]. Khuri [Dra10a]. Kiang [Pun11a]. Kim [Bla07, Dur21]. Kim-Anh [Bla07]. King [Gho11b, Mis10]. Kinnear [Sta12a]. Kirchgässner [Sai08]. Kirillova [Rao80c]. Kirk [Tra07]. Kirkland [Liu14a]. Kitagawa [Lis08b]. Kjaerulff [Gho08b]. Kjeld [Wed80]. Klassifikation [Lin75a]. Klaus [Dra12c, Fil09a, Han08m, McE17, Veh13a, Tru05]. Klaus-J [Han08m]. Kleiber [Sco09b]. Kleijnen [Boo79]. Klein [Slo23, Ric11a]. Kleinbaum [Ric11a, Tha07a]. kleinen [vdL80b]. Kleinman [Pol11b]. Klemens [Unw09]. Klerx [Els18]. Klevmarken [Sch80a]. Kline [Fon13]. Kloareg [Unw13b]. Klugkist [Gho09b]. Kneib [Lis10c]. Knott [Han13h]. Knowledge [Mai12a, Ram03, Han12h, Unw10a]. Known [BR07, Hoe83, Zha96]. Knox [Var20]. Koch [Gho08e]. Kock [Ohl77]. Koehler [Han09b]. Koenker [He17]. Kogan [Oja07]. Kohlas [vC75]. Kokko [Rou07]. Kokoszka [You14]. Kolaczyk [Han10f]. Kolchin [Gal80a]. Kolmogorov [HzB91, Sch84b]. Kong [Kha17]. Koning [Sta08c]. Konishi [Lis08b]. Konstadinos [Mer20a]. Koopmans [Ald99]. Korb [Mai11a]. Korea [Ano05a, FL05, OMGT05, SLY23]. Korn [Dur13c]. Körner [Sta09]. Kosorok [Num09a]. Kossack [Fin77b, Ano72c]. Kotz [Ber78, Eng78, Ano74d, Buc74, Ios78]. Koushik [Mer20c, Mer20b]. Koutras [Mer20a]. Kovacsics [San77]. Kozesnik [Ios79b]. Kragh [Dra11c]. Krämer [Fil09b]. Krantz [Mer12a]. Kravis [Kha80]. Kreiß [Kos10b]. Kreyszig [Eag74]. Krider [Sco13a]. Krishnaiah [Cor80, Pos78b]. Kroese [Ros14, Tra09]. Kroonenberg [Veh08a]. Krotki [Som78e]. Krusemeyer [Sta13a]. Kruskal [Bar75c]. Kshirsagar [Str73, Tin78c]. Ku [Ham72a]. Kubinger [Veh13a]. Kubokava [Num23]. Kuebler [Lin80c]. Kuhnert [O’B09e]. Kukush [Han10j]. Kulahci [Han12v]. Kulik [Han10j]. Kulinskaya [Han08h]. Kulkarni [Gho12b, Han10e]. Kullback [Csi80]. Kumar [Dra14b, Mit79, Pun10c, O’B10d, Som78d]. Kung [Mai09c]. Kung-Sik [Mai09c]. Kunio [Muk14b]. Kupper [Num13b]. Kupresanin [Sco13c]. Kurt [Mün76]. Kurtosis [FZ09]. Kvam [Oja08b].

L [Alv11, Amu73, Bar72c, Bar77c, Bil77a, Bre79c, Buc74, Cha80, Coo07, Cox80, Cro10f, Cro11b, Dem73, DL13, Dur18, Fed81b, Gho08g, Gho08b, Gho10b, Goo73, Haj77, Han09a, Han12c, Han12l, Han13c, Hoe83, Høy72, Ios78, Ken76b, Kin75c, Lid80, Mat73, Meh75, Oga73, Pla75, Pod18, Pod72, Pol13b, Pos78d, Pos79c, Pos80a, Pre79a, Pre79b, Ric17a, Ric17b, Ric18, Rob19, Sco13c, Sen14d, SJ89, Som77b, Spe08b, Sta10a, Szc11b, Teu05a, Tin80b, Tip77, Tra13, Veh09b, Wim75]. L. [And86, And87, Big05, Cau86, Clo86, EJ86, GS86, GS87, Ken86, Sen03]. Lê [Mai11c]. Laan [Cro08c]. Labor [Som78g, Mal97]. Laboratories [Lee79]. Labour [CB00, Hof97, LA96, Ver13, Meh15]. Labs [Ket12]. LaBudde [O’B13a]. Lachenbruch [Smi77]. Ladislaus [HV15]. Lady [HG21]. Lag [Smi73]. Lai [Cro09c, Kos09b, Lin14b]. Laing [Pos78c]. Laird [Cro11h]. Lajos [You14]. Lakshmivarahan [Han07c]. Lalonde [Dur18]. Lancaster [Ord72]. Land [Som78d]. Lang [Cro10i, Kos12b, Pol12b]. Lange [Cro11h, Rob11f]. Langelütke [Für73a]. Langevin [LV14]. Language [Ald03, BF97, Hau01]. Lanza [Gho10a]. Laplace [HK16, Koe16, Men16c, Smi16, WS16, YWH16a, YWH16b]. Lapse [Ben75]. Large [BK88, CK19, HM90, HK77, Kin75a, Oja07, PHO86, Pun10d, RDM^{+}19, Wil95, HL74, Mon87, Spe11a]. Large-Sample [BK88, HM90, PHO86, Pun10d, Mon87]. Large-Scale [Spe11a]. LaRiccia [Gho11d]. Larner [Rus10]. Larose [Han07b]. Larry [Far13, O’B13f, Ros09b]. Lars [Han07e]. Larsen [Spr77b, Sun22]. Larson [Cac74, Cor77b, Ios76c, Sta13a]. Lasso [O’B16c, SLY23, FGFBGM22]. Last [NC21]. Latent [Gho10a, MSS20, Pyn14a, Gho10a, Han13h, Pyn14b]. Latin [HRR07]. Lattice [Nor08]. Laubscher [Bar72c]. Laugier [Dug74a]. Laura [Mai14a, Pod17a, Sco12b]. Laurent [Unw13b]. Lauritzen [Gho08g, Han13e]. Lavallée [Has08b]. Law [Ano75e, Per75, QS87, Unw10a, Cro10d, Ken80]. Lawless [Dra14d, Han08n, Wil16b]. Lawrence [Cro09a, Han13o, Han13b, Hau08a, Lan18, Num13b, Dur13a]. Laws [Cro11c]. Lawson [Dra11a, Gho09a]. Layouts [SKK18]. Lazar [Gho09i]. LCCN [You19]. Leach [Spe08a]. Leading [Tan97]. Leak [Cam77]. Learn [Can95, Gar95, GBZ07, Sch95, Sin95]. Learners [GO10]. Learning [And73c, BDLP20, Bie97, Bie99, Cle05, Cob07, For22, FPR^{+}16, GBZ07, Han10d, Han13k, HKZ07, HCV^{+}21, Kha24, Liu21, Mai12c, Men16b, Nic01, Nor09c, Nor14a, O’B09h, O’B16c, Pod14a, Ric11a, ST07, Var20, Wil07, ZZN^{+}22, Liu20a, Mai11d, Pun08c, Sen14b, Muk14b, Sta07d]. Least [Ald98, DR90, FBGGM17, GP73, Han07c, Han12g, Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, Hum77, LNA75, Pun07a, Tin80a]. Least-squares [FBGGM17]. leave [CB00]. Lebart [Pre79b]. Leclerc [Mar75]. Lecture [Fel04, Knü80, Gro95, Mar96, Wak94, Wak95, Wak96]. Lectures [Cha80, Ano94d]. Ledolter [Lis09a, Luo07]. Lee [Hol08, FvZ15, Gho08f, Lee05, LLT04, LQ04]. Legacy [Hav14, LZ22, Bel16, Mer15]. Legendre [O’B12h]. Legislation [Mar74]. Lehmann [Spe08b]. Lejeune [Mar07a, Esp11]. Lemuel [Gho08c]. Lene [Dra11c]. Length [Koe79, Lee01]. Lennart [Gal80b, Gan76]. Lens [Han13i, HMT22]. Leo [Veh08g]. Leon [Buc76a, Buc76b]. Leoncio [Bar75b]. Leong [Pun11a]. LeSage [O’B09d]. Less [O’B10e]. Lessons [HRR07]. Let [HHD97]. Letter [Alv20, GKV20, LSS08, NLSS08, PP98]. Letters [Kin75b]. Leukaemia [Gar93]. Leung [Cro09c, Kos09b]. Level [CSW93, GW23, Nic01, Dam74, KBM21, Gan76]. Levels [GOZV94, HG94, Han09c, RW09, She06, SYSK02]. Levi [Pod24]. Levinson [Mor84]. Levy [Bar72c]. Lewis [Han07c]. Li [Kau12, Mai14b, Pol12a, San17]. Liaison [Dam77]. Liang [Ken10, Lin14b]. Liber [Spe08a]. Liberati [PS13]. Library [All72]. Lid [Cro11a, Lis09c]. Lié [Dam74]. Lies [Sun18]. Liese [Han08m]. Life [Bar80b, Bel16, BJPT99, Buc77d, Con09, Cro10f, Dav09, Dra13a, Fer94, Han10a, Hav14, Hoe83, Kei05b, Koe79, Law07, Mai09a, Mer15, Rao19, Sko12, Spe08a, Sta10b, Str76, Sun15, TC72, TP73, Wil14b, Cro08b]. Life-Length [Koe79]. Life-Testing [TC72]. Light [BM16, Ole03]. Like [Liu14b]. Likelihood [Ald02, AG08, BNS94, BK20, Bow04a, BCD^{+}94, CPB94, Edw74, FS94, Gho11d, GH87, HL90, Han11k, HK16, HH11, Koe16, Le 90a, Liu12b, Mac14, MM01, Men16c, Nor72, Nor73, Oak81a, Oak81b, PS20, PB17, Ros87, Sch78, SB04, Smi16, Tra13, WS16, YWH16a, YWH16b, You19, YK20, ZL22, Zha10, ZW19, BNC84, Gre87, Hig77, Jac84, JT13, LK16, Roy86, Sko86a, Sør83, Sør86, O’B09e]. Likelihood-Based [ZL22]. Likelihood-Frequentist [PB17]. Likelihoods [HH14]. Likert [Tut21]. Lima [Lis10b]. Limit [AP94, Cro09c, Mer18]. Limitation [CKK^{+}11b]. Limitations [GR90]. Limited [LLT04, Liu23, Num15, SM14]. Limits [DR90]. Limnios [Cro09d]. Lin [Han13o]. Linda [Gho10a, Han10c, Unw11c]. Lindberg [Lut76]. Lindgren [Cor77a, Fin78, Nor13b, Som79b]. Lindley [Bar82, Ber82, Din82, Sha82, Sil82, Sin02a, Smi82, Zad82]. Line [Gho13c, Jon07, Mac97, Tri09, Vov01]. Linear [AS72, Ald99, ABGK82a, ABGK82b, Ano84a, Ano86a, Bat07, CFP22, Cau80, CDM19, CO14, Cro10g, CP07, DNC06, Dra08a, Dra11c, Dra14a, Fer80a, Fon11, FGDS18, Goo86b, GNM21, HH07, HH14, HMW21, IKS^{+}81, IKS^{+}82, JK78, Khu13a, Kle76, KPT05, Lia20, Lis13, LLP08, Liu12a, Liu14b, Mai19, NC10, NK14, Num15, O’B10c, dBPR83, Pod17b, Pol13a, Pos73, Pre89, Pun08a, Pun09, Pun11c, Pun12c, RSG06, San91, Sen72, SRN17, SM14, Spr77a, SS07, Tia07, Tri09, vdL03, And87, GS87, Gob16, Goo86a, McC85, dBPR86, PMM87, RLS23, Sri80, Tam87, Pod17b, Bar75d, Dra10a, Han11f, Kle76, Mai14c, Pre80b, Pun07a, Pun08b, Pun10b, Pun12c, Tin80a, Lis14].

Linearisation [VT19]. Lininger [Som76b]. Link [LS04]. Linkage [Bic00, CHR18, Chi22, Cla00b, Cro08a, Has08a, MS23]. Linked [Dam74, HL19]. Linking [Han08g]. links [Dam77]. L’Institut [Ano72a]. Liouville [GR01]. Lipidomics [Gho17]. Liptser [Rév80]. Lirong [Ng21]. Liski [Liu07]. Lisniansku [Ng21]. List [Ano75k, Lan83, Lan85, Bad82, Lan72, Lan73, Lan74a, Lan74b, Lan75, Lan76a, Lan78, Lan79, Lan81, Lan82, Lan84, Lan86, Lan87]. Literacy [Ahl02, Bat02, Blu07, Gal02a, Gal02b, Mac16, Rum02, Tre05b, Wat02, Wel02]. Literature [EFP19, KM79b, KM79c, KM79a, KS85, KZ76]. Little [Hoe83, RAC83]. Little-Known [Hoe83]. Liu [Han13m, Veh07b, Gho09f, Han11j, Liu10b]. Lives [Law07, Pun11a]. Living [Der77a, Joi75, Fou80a, Han13i, Som79e]. Lloyd [Fin80c]. Load [CG12, CS00]. Løber [Unw13b]. Local [Ano80e, HH14, LV14, MB19, PK80, Tam82]. Localised [BCM19]. Locating [Wil16a]. Location [AW95, Ell91, Tho00, TB22]. Location-Shift [TB22]. Loci [SGLS02]. Locus [Fin80d, Tho00]. Log [CFP22, Goo86b, IKS^{+}81, IKS^{+}82, And87, GS87, Goo86a, Sko86a, Lis14]. Log-Linear [CFP22, Goo86b, And87, GS87, Goo86a, Lis14]. Logarithms [Bel16, Hav14, Mer15]. Logic [Can95, Han10c, Sch95, SS08, Sin95]. Logiciels [Pre79b]. Logistic [IKS^{+}81, IKS^{+}82, RG23, RJ91, Han11g, Ric11a]. Loglikelihood [CR89]. Lognormal [Sch16b]. Lohnes [Gow72]. Loïc [Pod14b]. Lois [Mar75]. Lomax [Han12c, Han12l, Veh11a]. Lombardo [Fis18]. Lonardi [Ric10b]. Long [OA91, RM93, SYSK02, Han13d, Bha07]. Long-Memory [Bha07]. Long-Term [OA91, RM93, SYSK02]. Longford [Has08d]. Longini [Aur10]. Longitudinal [Blu21, Coo07, FST19, GNM21, KRR^{+}11, Kos10c, Mar11, Nee15, Pyn14b, WQS19, WY02, AHZK15, Lai22, NSZH22, Cro09b]. Longitudinally [FRS08]. Look [HG21, Tad80]. Looks [Han12b]. Loon [Veh10c]. Looney [Ric18]. Lorch [Szc11b]. Loren [Sta13a]. Lorenzo [Pod17c]. Loss [BCF^{+}14, BW12a, BCZ12, BW12b, Dat12]. Lothar [Dei76]. Louis [Sap15]. Louise [Ano21]. Low [YAA20]. Low/High [YAA20]. Low/High-Dimension [YAA20]. Lower [NK14]. Loyo [Pat74]. Lozier [Cro11e]. Ltd [Dur17]. Lu [Gho13a]. Luc [Här08]. Lucila [Han13m]. Ludwig [Lis10c, Kin77a]. Luigi [Hau18, Pos79a, Ros10b, Fér78]. Luis [Var19, RS23]. Luisa [Roy14]. Lukacs [Sch72b]. Lund [Rui72]. Lütjohann [Kle76]. Lutz [Mui73]. Lwanga [Gri77]. Lyle [O’B10b, Zha14]. Lynch [Gho08d].

M [All79a, Ano21, Aoy72, Aur10, Bar77a, Bar76, BFST18, Bil77a, Blu07, Blu20, Bol07, Bra14, Bre79a, Buc77d, Buc77c, Coc73, Cor75b, Cox80, Cra73, Cro11h, Dav09, Dor73, Dur23b, Els20, Emb08, Fis72, Gan80c, Gan80d, Gho08d, Gho10b, Gho10a, Gir73, Han07h, Han07c, Han09c, Han11g, Han13j, Høy72, Ken80, Kos13d, Lis10b, Liu21, Luo09, Mor73, Num13b, Num14, O’B07b, O’B08b, O’B11b, Oga73, Par73, Pod14a, Pos79b, Pos79a, Pun09, SV22, Sco10b, Sco12b, Smi94b, Som76a, Spe11b, Spr80, Sta10b, Sta11a, Sta11b, Str73, Tha07a, Tin78c, Tin80b, Tip77, Tra07, Tra73, Veh08g, Veh08a, Veh14, Vin73, Wet80]. M-quantile [BFST18]. Ma [Cro08a, Han08d, San17, Zha16b]. Maarten [Pod22b]. MacCallum [Pyn14a, Veh08d]. MacDonald [Cro11b]. Machado [Han13m]. Machine [BDLP20, Han13k, HCV^{+}21, Kha24, Liu20a, Liu21, Mai11d, Pod14a, PHO86, Sen14b, ZZN^{+}22, Pos78d, Han10d, Var20]. Macro [LA96, Smi72]. Macro-Models [Smi72]. Madansky [Bre80]. Made [Buc77d, O’B12c, Sta12a, Lau81]. Madelyn [Tso20]. Madsen [Liu12a, Gho08b, Haj77, Lin13]. Maela [Unw13b]. Magallanes [Liu22a]. Magnello [Veh10c]. Magnus [Dra14c, Liu20b]. Magy [Liu22b]. Mahalanobis [RBB^{+}21, Ano72j, Rao73]. Mai [Dur13c]. Maillardet [O’B09c]. Maindonald [Sta07a, Unw11a]. Mairesse [Hor79]. Major [HT92]. Majorization [Pun11b]. Majumdar [Han08k]. Makar [Ros19]. Makes [Riv02]. Making [Buc79, Gol04, Hor16, OS93, Pat88, Pol11a, Sco13a, Sta09, Sun96, Elf72, Yan73]. Makridakis [And77a, Ord80b]. Malcolm [O’B12g]. Malik [Pun10c]. Maller [Gay11]. Malley [Mai12c]. Mallick [Gho12a]. Mallows [DL13]. Malozemov [Roh78]. Man [Cro09c, Lin14b]. Man-Lai [Lin14b]. Management [Col99, Fis09, Pod15a, Pol11b, Sne99b, Sun17, Tam07, Kos09a, Ord80b]. Managing [Sco07]. Mandal [Dra14b]. Mandallaz [Has08c]. Mandelbrot [Ios79a]. Mandl [Sch80b]. Manfred [Veh10a]. Manifold [Pun08c]. manifolds [Mit88]. Manipulating [HP22]. Manipulation [Bat08a]. Manisha [Dra14b]. Manjunatha [Liu14a]. Mann [Nem73, ZMSL19]. Mantel [Dar81]. Manual [Moo75, Gho09e, Sta11b]. Manuel [Liu22a]. Manufacturing [EM93]. Manuscript [Bel89, Bel00]. Manuscripts [Had10c]. Many [Dra13g, SS08, Tuk91]. Many-valued [SS08]. Mapping [BR00, Cla00b, Gho09a, GDSN18, Jon00, SGLS02, WOP20, WS01, And07]. Maps [Cro08a]. Marc [Pod14b, Gen24]. Marcel [Kos10c]. Marcelo [Csi08]. Marchés [Mal97]. Marchi [Bau14]. Marco [Gho08f, Liu08]. Marcoulides [Han11e, Veh09a]. Mardia [Huz73]. Margaret [Dur17]. Margerison [Ham79]. Marginal [Dar81, Ham72b, WLC82, Kos10c]. Marginalised [SCC14]. Margins [Veh08f]. Maria [Bat08b, Gho11a, Liu10a]. Marian [Han12n]. Marianna [Slo24]. Marianne [Han08f]. Marie [Cro09b, Sun24]. Marilena [Dek14, Sta12b]. Marin [Cha14]. Marina [Bla07, O’B11a]. Mario [Liu10a, Bar72a]. Marius [AK95]. Mark [Cro08c, Don12, Dra13c, Gho13b, Gho13c, Han07h, Han12j, Han12m, Mai13c, O’B07a, O’B08a, O’B12a, Pod18, Pol12b, Ric13b, Sen14b, Smi80, Sta13a]. Markers [SJFM95]. Market [Han12s, Hor16, MCH13]. Marketing [Knü80, Luo07]. Markets [Emb08, Gho12h, TTV05, Mal97, Kos09b]. Markos [Mer20a]. Markov [Cro09d, Cro11b, Han07d, Han07b, Ios76d, Mai14a, Pol12a, Pyn14b, BM89, GT12, Ish81, KLTG21, Mar11, MP21, PWKO24, RR21, Sen96, She00, Haj77, Cro09d]. Markov-switching [PWKO24]. Marks [DMMG22, EM19, Som78e]. Marlow [Dra09d]. Maros [Tin78b]. Marques [Veh07a]. Marriott [Sco77b]. Marschak [Kle78]. Marshall [Cro08b, Gho23, Pun11b, Duf78]. Marsland [Han10d]. Martin [Aoy72, Dra14a, Han13h, Mer09c, Mer09d, You20, MG05, Dra13b]. Martin-Guzman [MG05]. Martinez [Alv11]. Martingales [RR90, Shi81]. Martz [Gho09d]. Marvin [Dra10b]. Marx [Pod22a]. Mary [Coc79, Han08g]. Masking [KMNS05]. Maslow [RS79]. Mass [Gho17]. ’Massively [Mur94]. Matched [Imb07]. Matching [BW12a, LN77, MP23, Smi20, BCZ12, BW12b, Dat12]. Math [Pod11]. Mathai [Ken80]. Mathematical [Ban14, Buc76a, Buc76b, Buc77b, Buc78, Cra81, Cro11e, Dra09d, Dra12e, Eag74, GL19, Hoe74, Ios76a, Ios77d, Kin77b, Knü80, Mag09, MM77, Mer09c, Mer09d, Mer14, O’B11c, Pun11a, Sta13a, Sun19, Syl14, Bar75c, Cra73, Knü80, Kok07, Kos09a, Sco12b, Sta12b, You19]. Mathematician [Dra13e, Mer12a, Spe10a, Far13, Haa12]. Mathematics [Han10j, Han12t, Liu10b, O’B13d, Pol12b, Spe10e, Sta09, Sta11a, Sta11b, Dur13a, Gho12c, Han11c, Hoe83, Mer09b, Ric13b, Sco13b, Dra11b, Gan80a, Kin75c, Mer07, Ric13b, Ros10a, Spe07]. Mathématique [Rao80a]. Mathématiques [Pos76b]. Mathematische [Kri74]. MATLAB [O’B13h, Veh07a, Alv11]. Matrices [Mar07b, Pou02, Pun11d, Sch16b, Tra73, KDTP20, Liu14a, Mer20c, Mer20b]. Matrix [Fil09b, Gus06, Han08o, HGKJ20, Lis08a, Liu10a, Liu14a, Liu14b, Rao06, Tsa24, Dra14a, Han07e, Liu20b, Pol13a, Pun07b, Pun10c]. Matrix-Variate [Tsa24]. Matt [Num20]. Matter [Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano73a, Ano73b, Ano73c, Ano73f, Ano73g, Ano73h, Ano74a, Ano74b, Ano74c, Ano74e, Ano74f, Ano74g, Ano75b, Ano75c, Ano75d, Ano75g, Ano75h, Ano75i, Ano76c, Ano76d, Ano76e, Ano76g, Ano76h, Ano76i, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78d, Ano78e, Ano78f, Ano78h, Ano78i, Ano78j, Ano79b, Ano79c, Ano79d, Ano79f, Ano79g, Ano79h, Ano80b, Ano80c, Ano80d, Ano80f, Ano80g, Ano80h, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86b, Ano86c, Ano86d, Ano86e]. Matter [Ano86f, Ano86g, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c, Ano93e, Ano93f, Ano93g, Ano94a, Ano94b, Ano94c, Ano94e, Ano94f, Ano94g, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96a, Ano96b, Ano96c, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99b, Ano99c, Ano99d, Ano99f, Ano99g, Ano99h, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano01a, Ano01b]. Matter [Ano01c, Ano01e, Ano01f, Ano01g, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05b, Ano05d, Ano05c, Ano05e, Ano05g, Ano05f, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h]. Matthias [Dur13c, Gho12h, Liu17]. Matthiopoulos [O’B13d]. Matzner [Unw13b]. Matzner-Løber [Unw13b]. Maurice [And78, Tip75b, Dug74b, Wil75]. Maurizio [Dur13b]. Mavromatidis [Pod21b]. Max [Bau14, Fon11]. Maxima [HH14]. Maximal [GP73]. Maximisation [Mac14]. Maximum [BCD^{+}94, FS94, Jac84, Le 90a, Nor72, Nor73, Sch78, Hig77, Roy86, Sør83, Sør86, Gho11d, Liu12b, Tra13]. Mayaffre [Rod21]. Mayes [vdL80c, vdL80c]. Mayo [Spe08a]. MC [Num02]. McCormick [Gow79]. McCulloch [Pun09]. McCullough [Ana21]. McElreath [Kok07, Rui17]. McElroy [Qiw20]. McGrayne [Rob12b]. McKean [Mer18, Oja11]. McLean [Wim75]. MCMC [SGLS02, Tho00, WS01]. MCMC’ists [Num02]. McNeil [Pre80a]. McNemar [Hoo05]. McNicholas [Fis19]. Meal [Bar75b]. Mean [AAF18, BR07, CM12, KS22, PBH19, Sär72c, SG77, SH21, TS06, ZL22, Zha96, BNKS82, VLM12]. Meaningful [dV01]. Meanings [Ahl02, Gal02a, Wel02]. Means [LNA75, Lit86, VT19, Hum77]. Measure [Ber17, Ber06, KPT05, LS06, LHSD15, Mon01, Pod14b, Yim90, Ios79c]. Measurement [Bai85a, Bai95, BJO^{+}98, Bie95, Car93, DO73, Des00, Dur17, FMR^{+}20, Ful95, GT99, Hal84a, Ham72a, Han11d, Har99, JRL10, KT03, LvR74, RW09, Sne99b, TC07, Whi99, Wyc03, Dam74, dM84, Dam77, Som78e, Veh08f, Blu21, Bre79b, Cro10h, Som79a]. Measurements [Hai72, KASG80, Han11h]. Measures [AES18, Ayh01, BGV18, CS88, ESS10, FMP20, Gou94, O’B10b, SRN17, SG77, dV01, OC13, TTB86, Han13j]. Measuring [AW95, Blu05, GP03, Han13o, JS03, NZY18, Ram03, Rao05, dV99, vdBZT20, Num19]. Meat [Nem73]. Mechanics [Pun12b]. Mechanisms [Pea02a, Sin02a, SS88]. Meddis [Tip80]. Media [PZ21, Gri87]. Median [CG12, SG77]. Medians [Sma90]. Medical [Arm85a, BCL^{+}85, Bat85b, Cha20b, Che85b, Cro08e, EIRS18, Gri77, Kró80a, Lan18, Lou85, MM96, Num13c, O’B09b, Pet85, Ric13a, SAG^{+}16, Tha07b, Whi85, Bat85a, Num09b, Sen11c, Sen11b, Tha07b]. Medicine [Cam07, Pos79c, Pre77b, Yoo21, Bra14, Cor75b, Han07f, San17, Sch16a]. Medieval [Bel00]. Mediocre [Sch07]. Meeker [Var19]. Meenakshi [Pol12b]. Meet [Car05, SHLW22]. Meeting [Ano88a, Ios79b, Pol12b, Pre78b, FT87, FT88]. Mehdi [Kha77]. Mehran [Han10g]. Mehrens [Bre79b]. Meister [Cro10e]. Melese [Gan72a]. Melnick [Han09a]. Members [Ano75a, Ano76b, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano80a]. Membership [Ano76a]. Membres [Ano76b, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79a, Ano80a]. Memoir [Far13]. Memoirs [Haa12]. Memories [Dra13a, Pea74]. Memory [GG17, LgX13, Tin80b, Bha07]. Memoryless [KBH10]. Mendenhall [Fin77a]. Meng [Pol12a]. Mensura [Hal84a]. Mentors [Rus10]. Mertens [Gho17]. Mesquita [Dek22]. Messy [Dra09a]. Mesure [Dam74, Dam77]. Meta [Mai13a, MBW^{+}23, Han12w, Ros09b, Cro11d, Han08h, O’B09e, Sun16c]. Meta-Analysis [Mai13a, Han12w, Ros09b, Sun16c, Cro11d, O’B09e]. Metabolomics [Gho17]. Metadata [Col99]. Metcalfe [Mai10a]. Method [Ano76f, CKN^{+}24, Chu74, DO73, GL14, Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, Kei87, LvR74, LLT04, MB19, O’B12e, Osb92, SKK18, Tal93, TS06, WvdH06, Zha03, BT75, BT76, Lef79, VLM12, Par73, Tra09]. Méthode [Lef79]. Methoden [Lin75a, Dei76]. Méthodes [Dag75, Pre79b]. Methodological [Ano79e, EKIKW77, Hor91, IKS^{+}81, MCF12]. Methodologists [Bin00, Cla00a, Dil00, Lyb00, Tre00b]. Methodology [BCO82, Con09, Der75a, HL90, Han08i, KFL76, TG94, Zha16a, lRG05, Dra10a, Sun22]. Methods [Alv14, Ano75f, Aus17, AG08, BR00, Bat07, Ber14, Bic14, Bin07, BO15, BBvdB18, CFL^{+}21, CW06, CKL14, CKS17, CP96, Cro10j, Cro10h, Cro11f, DH19, DS99, DMS06, Dek94, DR90, Dur21, Dur23b, DG17, EH95, Esp14b, Esp15a, FRS08, Feh07, Fom88, FGDS18, FMR19, Fri03, Gho08g, Gho09c, Gho11a, Gho12g, Gil15, GDSN18, GSW17, GM91, GPW05, Han07e, Han09d, Han10f, Han11d, Han12a, Han12d, Han12j, Han12n, Han12g, Han13d, HHK03, Har99, HK18, HP17a, HP17b, Hum77, Ios76d, JLS20, Jon00, KSS14, Kos09b, Kos13a, Lan18, Lar15, Lav14, LW13, LD24, LNA75, Lis10b, Lis13, Liu10a, Mai11b, MM96, Mat07, Mur94, NN20a, NN20b, NAL76, NSW72, NPT08, Nel80, Nor13a, Nor98, Num19, O’B07a, O’B09i, O’B10b, O’B12g, O’B12a, O’B12i, O’B13a, O’B13g, O’B15b]. Methods [Oga73, OCRC17, Ord80b, PJK17, PGN^{+}96, Pla75, Pod15a, Pod17b, Pos80e, RYG^{+}21, RGSO17, Ros15a, Sco77a, Sco10a, She00, Shl07, SHGK01, Spe11a, Sta08b, Sta12b, Sun16a, Sun17, Tip77, Tip75b, Tso18, Unw10b, VF00, Wan16, WD84, Wet80, WL16, Yas93, ZM17, Zha14, ZW19, dGAGR85, dWvDS20, vdL80d, Atk88, DY16, Fis83, Gob16, Joh15, LN78, OC13, SSJ77, WGKL22, Bha07, Cox10, Dra09c, Hee16, Ken74, Kos10a, Mor73, Oja08a, Oja11, Pos78e, Pos80b, Bas16, Loy80, Pun13b, Sta13c]. Metodo [Tin78b]. Metrics [FMR19, GS02, Cox19]. Metrology [GL14]. Michael [Cor77c, Cro10d, Cro11c, Dra13b, Dur14b, Dur23b, Gho09d, Gho10b, Gho13b, Mai07, Muk14a, Nor14b, Num09a, O’B13a, Pod14a, Pun07a, Pyn14a, Ric14, Rob11g, Ros09b, Som76c, Sta07g, Sta08a, Sun22, Var14, Veh08b, Veh11b]. Michaela [Liu08]. Michel [Cha14, Esp11, Hol08, Liu11, Mai09b, Mai21, Mar07a]. Michelangelo [Rod21]. Mickey [Blu20]. Microbiology [Gho09e]. Microdata [Ich09, KRR^{+}11, MDS94, NZY18, Rei09, WvdH06, Cro10b]. Micromaps [Unw11c]. Microsimulation [AET04, Dek10, KCB00]. Microsoft [Han07a]. Microstrata [NPT08]. Miescke [Han08m]. Migration [Hof97, Hug06]. Mihoc [Oni83]. Mike [Häg12a, Han10h, Rob11h, Sen14c]. Mikhail [O’B11a]. Mikosch [Kos10b]. Mikrozensus [Mui73]. Miles [Unw13a]. Militino [Bat08b]. Millar [Liu12b]. Millennium [JS03, Sti00]. Miller [Pos79d].

Milliken [Dra09a, Zha19b]. Mills [Han12r]. Miloslav [vdL80a]. Milton [Dal72, Sch73]. Minds [Pun11a]. Ming [Ken10, Pol11a]. Ming-Hui [Pol11a]. Mini [HRR20, HR21]. Mini-issue-Statistical [HRR20]. Minimax [Bic13, Roh78]. Minimax-optimal [Bic13]. Minimizing [Gow79]. Minimum [Azo01, Han13n, Lee01, NW82]. Mining [All79a, Han07a, Han07b, Han07e, Han08o, Han10g, Han13a, Mai11d, Mai12a, Mai13b, Nor09c, Pod14a, Sco13a, Han11l, Kau12, O’B10d, Ric10b]. Minitab [Bre79c]. Minoru [Ued74]. Misclassification [vdHvdH02]. Misconceptions [Ric09]. Mises [LV14, Sch84b]. Mises- [Sch84b]. Mises-Langevin [LV14]. Mishura [Han10j]. Missing [AT15, BH14, CMH19, DGS18, Esp14a, GNM21, Ken10, Muk14a, Muk15, ZHW23, ZZLZ15, Mai16a]. Misspecification [HMW21, HMT22]. Misuraca [Rod21]. Misuses [LLP01]. Mitchel [Ric11a]. Mitra [Pun10c]. Mitsa [Han11l]. Mitsunori [Kau12]. Mitteilungen [San77]. Mitteilungen/Communications [San77]. Mixed [Dar81, FMR19, GNM21, HH14, HMW21, Hus72, KX05, Lav14, LSS05, LS06, Lis13, Mai14c, Mar11, MDS^{+}19, NLSS08, Pun09, SCW08, SRN17, Zha19b, RLS23, VLM12, Cro10i, Num23, O’B15b]. Mixed-device [MDS^{+}19]. Mixed-Effects [Mai14c, Num23]. mixed-mode [VLM12]. Mixed-Models [NLSS08]. Mixed-type [FMR19]. MixedTS [HHMR18]. Mixing [BF91, FWS14]. Mixture [CP19, Dra14b, Han07d, RW09, Sen11b, SM19, YYY20, AT15]. Mixtures [Cox07, TO17, BNKS82, Han10b, Dra12b]. Modal [Cha20a, Men16a]. Mode [JRL10, VLM12]. Model [AS72, Ald99, BD94, Blu05, BGS24, Bre75, Cas20, Cau80, CDM19, CW06, CM12, Cra73, DT92, DS99, Dra10a, Eng78, Eri73, EGS72, Fil11, FGDS18, FBL18, GW23, Han12f, Han13c, HH07, Hoe87, Hus72, IKS^{+}81, IKS^{+}82, Joh83, Kab09, Kuk20, Lan01, LO04, Lis09c, LLP08, Liu10a, LQ04, McL93, MNS77, Men16c, MR10, MW10, OZW23, Pau19, Pre89, RW09, Red01, Rob11e, Roy86, San91, SGLS02, SRN17, SH21, SAA22, Tia07, TC72, Unw21, WQS19, BH80, DFM^{+}22, Don86, HMT22, Jac84, RLS23, SGM21, Tam87, Häg07, Här08, Lis09c, Mai09a, Unw20]. Model-Assisted [CM12, WQS19]. Model-Based [Han13c, Unw20]. Model-Oriented [Fil11]. Modèles [Dag75]. Modeling [AN01, Ban14, Dur13b, Dur18, Dur21, Fon13, Gho08f, Gho09a, Gho12b, Gho13d, Gho21, GP11b, Han11f, Hau18, Kos11, Lis14, O’B11c, O’B13e, OH13, Pet07, Pfe93, PAF18, Pou02, Rob11h, Ros14, SW94, Sta12f, Sun16c, Veh09b, Veh10d, Veh14, Wan16, Ber16, DY16, Gho12e, Han13a, Lis07c, Lis08b, Nan10b, O’B10a, Rob11e, Veh11a, Han10e, Han12i, Unw09]. Modelled [God97]. Modelling [CP96, DC19, Daw02a, Daw02b, EM19, FMS12b, FS10, Gho12a, GTF18, Han08l, HPHdAM22, KMM^{+}17, Kos13d, KMT03, Lis10c, MBW^{+}23, MSS01, MV22, O’B15a, OZW23, Pan07, PMM87, Ryd00, SP11, Sen23, SR92, SD11, SL13, Sun16a, TT20, TO17, Zha17a, JT13, MMC13, Dek10, Häg12a, Pod17c, Ros15b, Liu17, O’B12g, Rou07]. Models [ABC11, AT15, Ano96d, ADE^{+}22, Aus17, Bar95, Bat02, Ber78, BD94, Buo06, CFP22, Che24, CO14, CPB94, Cox10, Cro09d, Cro10h, Cro10g, Cro10i, Cro11b, CP07, DQS92, Dar81, DG90, Dra08a, DW15, DW84, Dur13c, DQ97, Dur15b, FZ15, FS08, FMR99, FPR^{+}16, FB96, Fro06, FGC03, FWS14, FST19, Gan78c, Gan80d, Gho09g, Gho11c, Gho12c, Goo86b, GNM21, GP11a, GK01, HM90, Han12f, Han13e, Han13h, HM04, HH14, Hjo92, Hon15, HWS20, HS15, HMW21, Ios78, Jas15, JK78, Jør92, Kal83, Knü80, Kok07, KY91, Kos09b, Kos10c, KPT05, Lav14, LS20, LH73, LSS05, LS06, LD24, Lia20, LR91, Lis13, Liu11, Lov00, LSD^{+}11, MYA95, ME89, Mai14c, Mai16b, Mai18, Man72, MN98, Mar11, MP21, MSW76, Mer20a, MFAG10, MS23, MCH13, NSS23, NSW72]. Models [Nee09, NLSS08, Ng21, NSP21, Num13c, Num15, Num23, Num19, O’B10c, O’B12d, O’B15c, O’B15b, Pan07, PMLL15a, PMLL15b, dBP77, PGV22, PS85, PSI19, Pol13a, Pun07a, Pun08b, Pyn14b, RSG06, RJ91, Rob93, RvA72, San15, SLWH22, Sen76, SCW08, SW15, Smi72, Smi73, Smi92, SM14, SM19, Sør93, SB87, Spe10b, Spe10c, Sto75, Sto98, SAA22, Swe95, SCC14, TZW^{+}20, TC07, TS06, Tju84, TBSW84, Tri09, Tso16, Tut21, TB22, War76, Wik03, WY02, YAA20, Zha19b, AL86, And87, CP80, FX16, GS87, Goo86a, Gre87, HMT22, McC85, PWKO24, Sri80, ZT13, Dra14a, Fon11, Han07d, Han10f, Han11g, Lis09b, Liu12a, Mai19, Ord80a, Pun08a, Pun09, Pun11c, Pun12c, Pyn14a, Rao80b, Sen11b, Str76, Unw21, Han10b, Coo07, Ric16]. models-Which [PWKO24]. Models1 [BDLP20]. moderated [Spe10e]. Modern [AM14, BR00, CHPR20, Cro11h, Gou10, Han11h, Han12r, Hun14, Ios78, LCZ14a, Lis13, Ole03, PB17, Rob14, RDMP16, Sai08, SHLW22, BKV13, Ano21, Dra09c, Han12j, Liu18, Liu23, Pun08c, Sta12c, Tan07, Zha16c]. modest [Hed20]. Modica [Mai14a]. Modified [DNC06, GNM21, KDTP20]. Moeschberger [Cox80]. Mohamed [Han13j, Sar72a]. Moivre [BD09, HJ83, Hal84a]. Molenberghs [Cro09b]. Moment [Hua89]. Moments [Gou94, TS06, HS12, SB75]. Monahan [Pun08a]. Monetary [MA05, Pan07]. Monitoring [Bol07, SBB^{+}12, SAP97]. Monographs [Sun19]. monopoly [Ket12]. Monotone [HCR18]. Monotonic [Kuk22]. Monte [KLTG21, Pol12a, RR21, Sco10a, Tra09, BO15, Gob16, JLS20, She00, vC75, Don12, Pod17b, Tso18]. Monte-Carlo [Gob16]. Montgomery [Cro10g, Khu13a, Dra08b, Pre78a]. Moon [Huh05]. Moore [Spe07, Pos80c]. Morbidité [Dam77]. Morbidity [Lid80, Dam77]. Morgan [Gho11b]. Morgenthaler [Han08h]. Morimoto [O’B11c]. Morineau [Pre79b]. Morlat [Fou80b]. Morphological [AM93b]. Morphologiques [AM93b]. Morphometrics [Mai09b]. Morris [Gro95, Mar96, Spe10d, Wak94, Wak95, Wak96, Ano94d, Fon11, TW92, WG97]. Morris-Howard [WG97]. Morrison [Sta10c]. Mortalité [Dam77]. Mortality [DMS06, LLT04, LHSD15, LQ04, RM93, Dam77, Lid80]. Most [HHD97, Mer24, Für73b]. Mosteller [Spe11b, Spe11b]. motion [IK86]. Motivation [GR90]. Moulin [Liu19]. Moulton [Cro09a]. Moustaki [Han08b, Han13h]. Mouzakitis [Dur13a]. Moving [Bha91, KY91, AL86, dGAGR85]. Moyé [Gho08c]. Mplus [Veh14]. Mr [Buc77a]. MR0731145 [dBPR86]. MR2990344 [BW12b]. MR2990345 [BW12b]. MR3006147 [O’K12, Rei12, Ski12a, Wil12]. MR3006148 [Ski12a]. MR3006149 [Ski12a]. MR3047496 [Cox13, Efr13, Fra13, Par13, Rob13, SH13]. MR3047497 [gX13]. MR3047499 [gX13]. MR3047500 [gX13]. MR3047501 [gX13]. MR3047502 [gX13]. MRI [Gho09i]. Much [LSS05]. Muenchen [Mai12b, Voe09]. Muhsam [Gan78a]. Mukerji [Pol12b]. Mukul [Han08k]. Mulaik [Han11f, Veh10b]. Müller [Bla07, Cro11a, Pol11a]. Multi [BCM19, CH13, LV14, Mur74, Ovu80, Sch07, TZW^{+}20, Tho00, dWvDS20]. Multi-Channel [Ovu80]. Multi-Compartment [TZW^{+}20]. Multi-dimensional [BCM19]. Multi-locus [Tho00]. Multi-Mediocre [Sch07]. Multi-sample [LV14]. Multi-source [dWvDS20]. Multi-state [CH13]. Multi-Subject [Mur74]. Multicollinearity [CP07, CP08, KPT05]. Multidimensional [BGV18, Gho12e, MSS20, Sma90, EMR88, Slo24]. multidimensionnels [EMR88]. Multifactorial [CD00]. Multilevel [Aus17, BCG07, Coh05, GTF18, JLS20, SD11, Veh09b, Szc11a]. multinomial [AAF18]. Multiple [Ano78g, Ano79e, BK20, BH14, DQS92, EKIKW77, GT17, HH14, MR03, Nie03a, Nie03b, Rei09, Rub03, dMRAMR18, Sco77b, SC05, SCW08, Sha23, SKK18, SHH94, SAA22, Tho78a, Tho78b, Zha03, Zhu14, SL13, Cro08c, Muk15, Nor14b, Ric11b, RR79b]. Multiple-Record [Ano79e, EKIKW77]. Multiplication [GL14]. Multiplicative [Hoe87, Jac84]. Multipoint [Tho00]. Multipopulation [Kis94]. Multiresponse [Nel72b]. Multiscale [Gho08f]. Multistate [SS08, Wil14b, Sun15]. Multivariable [Veh10d]. Multivariate [Ano76m, Ano89a, Bil77a, Buc74, DC83, DIV17, EM19, FFBK76, Han08b, HHMR18, wB09, Ken73a, Kha15, Koc14, Lee79, Loy80, Mai11c, MM04, MSS20, NK06, NK08, Nor08, Oga73, Pos78b, Pun08c, Pun13b, Ros09a, SHH94, Sko86b, STY90, SM19, SAP97, Unw12a, Veh09a, Veh10d, lRG05, Har79, KFF75, Sax78, SGM21, Sha88, Gir73, Gow72, Pod16, Pol13b, Pos78b, Pun10d, Str73]. Multiway [Veh08a]. Mun [Lis14]. Munich [Hen78]. Murat [Han12v]. Murdoch [Nor09a]. Murphy [Unw20]. Murray [Han11k, Sta12f]. Murrell [Unw12c]. Music [Kau12]. Musielak [Dra12e]. Must [Sne99b]. Mustafa [Rao80a]. Mutual [Gan78a]. My [Ber77, Rao19]. Myers [Cro10g, Szc11b]. Myles [Kró80d]. Myllymäki [Csi08]. Myslivec [Žal79]. Mythical [Moo78].

N [Aal82, Bar72c, Bil77a, Buc74, CWCT19, Coh75, Cra80, Cro07a, CV82, Duf82, Els20, Esp14a, Gan80a, Gho11d, Hal81, Har82, Has08a, Hau08a, Hoe83, Ios77a, LT82, Lau81, Lin75b, Mar73, Mer20a, Nel72b, Num19, O’B07b, Pod14a, Pos78e, Pre79b, Qiw20, Rév80, Roh76a, Roh76b, Roh78, Spe10d, Sta13c, Veh08g, Veh08c]. Nagaraja [Num19]. Nagy [Mar07b]. Nagylaki [Fin80d]. Naitee [Zha17b]. Naive [Sta09]. Nalini [Cha16]. Namboodiri [Bil77a]. Nan [Cro11h]. Nancy [Bje73, Tha07a, Sta17]. Napier [Bel16, Mer15, Hav14]. Narrative [Fom88]. Nash [Mai18]. Nason [O’B09i]. Nasr [Liu22b]. Nassim [Veh13b]. Nate [Dra13g]. Nathan [O’B14b]. National [AH17, Ano76m, Bin00, Cla00a, Dil00, FMS95, FL05, FFBK76, Has72, LvR74, Lyb00, MA05, OMGT05, OS73, PP98, Pul97, Pun11a, Sta12f, Tre00b, Uni78, dV99]. Nations [Ano72q, Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76o, Ano76n, Ano77l, Ano77m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j, SHL05, Har73]. Natural [Dod96, Dur22, GPdMLMGGSG20, Hop93]. Nature [Ber03, Cox03, Dur17, Han07g, KSS^{+}03b, KSS^{+}03a, Num23, Sch03]. Naturelles [Cas91]. Naus [Mus75]. Neal [Sun17]. Nearest [BM89, TRRL23]. Nearest-Neighbour [BM89]. Necessity [WWY23]. Necip [Han10i]. Nécrologique [Büh75]. Need [Hoe99, Sch80a]. Neelon [Num13b]. negative [HGKJ20, O’B08b, O’B11b]. Negativwachstum [Kro80c]. Neglected [Mac14]. Neighbour [BM89, TRRL23]. Neil [Han08g, Som79c]. Nekrasov [Sen03]. Nelder [Dal72]. Nelsen [Han11c, O’B10e]. Nemirovski [Gho20]. Neo [Ber77]. Neo-Bayesianism [Ber77]. Nested [CDM19, Pre89, SH21]. Net [Ast94]. Netherlands [HW06, LA96, SS09b, dVKW96]. Network [CNS85, Han10f, KC14b, LC20, Nor15, RDM^{+}19]. Networks [FMR^{+}20, FMS12b, Gho08g, Gho08a, Gho08b, Han08a, Kha17, Mur94, O’B14a, Sen23, Gho08g, Han10a, Han12i]. Neudecker [Liu20b]. Neuhaus [Pun09]. Neural [Han08a, Mur94]. Neuroscience [Gho13d]. Neusser [McE17]. Neutral [DER03]. Neuwirth [Sco10b]. Newell [Kin75a]. Newman [Goo73]. Newton [CKN^{+}24]. Next [Fis01]. Ney [Ios76b, Rad75]. Neyman [Ano74h, FT96]. Ng [Zha18, Ken10]. Nguyen [Liu22b]. Niche [Pet07]. Nicholas [Has08d, Liu18, Pol11b, Tso20, Veh13b, Hal84b, Sho85]. Nicholson [Ken76b, Mai11a]. nicht [vdL80b]. nicht-parametrischer [vdL80b]. Nick [Slo23]. Nicola [PS13]. Nicolas [Dur22, Unw13b]. Nicole [Gho09i]. Nielsen [Bil80b, Gho08a]. Niemelä [Liu07]. Nien [Han07f]. Nigel [Oja08a]. Nikolaos [Cro09d]. Nillas [Gho20]. Nillsen [Han11i]. Nils [Cro11a, Lis09c]. nineteenth [Lan87]. Ninth [Ano74h, Lan75]. NIST [Cro11e]. Nitin [Han07a]. Niveau [Dam74]. Niyati [Sen23]. No [Gow79, San77, And87, And77a, BT76, GS87, Hoy76, Ott76, dBPR86, Sch78]. Noel [Mai13a]. Noise [CQOR^{+}22, Dra13g]. Nolan [McC22]. Nombre [HL74]. nominal [Sco12a]. Non [ABC11, AC21, AL10, BSS18, BBvdB18, CH19, Col81, Dur18, Fer80a, FMR99, HCR18, HGKJ20, JSW11, KM79a, Kuk20, LN74, Lia20, Lin14b, Mai16b, MP23, NWE16, PCA^{+}19, Ric10c, Sär11, ŞA09, Sik18, SW15, SD11, Tsi03, VT19, You20, ZL22, Zhu14, Zur91, BBS77, Dal77a, Dal77b, Dal77c, FX16, Gob16, Hed20, MPR12, NL75, ZTK13, Tip80, Lin13, Pod17b]. Non-asymptotic [You20]. Non-Central [LN74, NL75]. Non-compact [AC21]. Non-Contact [SD11]. Non-Fundamentalness [ABC11]. Non-Gaussian [Dur18]. Non-Ignorable [ZL22, MP23, Sik18, Hed20]. Non-Linear [Fer80a, Gob16, Pod17b]. Non-negative [HGKJ20]. Non-normal [ŞA09]. non-normality [BBS77]. Non-Parallel [Mai16b, FX16]. Non-parametric [FMR99, Lia20, SW15, Zhu14]. Non-paramétrique [Col81]. Non-probability [BBvdB18].

Non-Proportional [Kuk20]. Non-Randomized [Lin14b]. Non-randomly [PCA^{+}19]. Non-Response [JSW11, Sär11, ZL22, AL10, CH19, HCR18, MP23, NWE16, Sik18, VT19]. non-sampling [Dal77a, Dal77b, Dal77c, ZTK13]. Non-Scientific [KM79a]. non-standard [MPR12]. Non-Stationarity [Tsi03, BSS18]. Non-Stationary [Zur91]. Non-Statisticians [Ric10c, Tip80, Lin13]. noncentral [NL75]. Nondestructive [Swe95]. Nonignorable [WQS19]. Noninferiority [Han12q, Zha18]. Noniterative [Ken10]. Nonlinear [DW15, EH95, Hou88, KS89a, LL87, Hau18, Hol08, O’B09f]. Nonlinearity [HzB91]. Nonparametric [ABGK82a, ABGK82b, Ano84a, BT80, BF91, Cor14, Cro08b, Cro10e, DMS06, DK01, DR90, DMMG22, Frö08, GPW05, Hal89, HLC97, Oja08a, Oja11, Pos78e, Pur72, Rui16, Sen11a, SJ89, Sim98, Stu72, WY02, Fis83, KASG80, Col81, FH89, Høy72, O’B16b, Oga73, Oja08b, Ric10c, Ric15, vdL80d]. Nonparametrics [Cro11a]. Nonresponse [Lit86, Hed20, SS87]. Nonsampling [Kir82]. Nordic [KP04]. Norma [Fin77d]. Normal [BNKS82, BR07, CG23, CW91, Dur14c, FGDS18, FWS14, Ken80, TC07, TS06, Bin83, DFM^{+}22, ŞA09, Ios77a]. normal-binary [DFM^{+}22]. Normality [GHK20, MM04, BBS77, JB87]. Normalized [Cro09c]. Norman [Cox07, Ios78, Kin77b]. North [NGC^{+}78]. Norway [LTS98]. Norwegian [TH98]. Nosal [vdL80a]. Note [Ano79e, Ano94d, Ano99e, AdV99, CP08, Edi77, Eid07, HN13a, HN13b, Ham75, Kag72a, LN77, Mis08, BH77, HN12, Sko86a, WLC82]. Notes [Knü80]. Nothing [LSS05]. Notice [Dav82, Büh75, Ios86]. Novel [O’B12e, Dra12e]. Nozer [Gho08h, Pea02a]. Nripes [Dra14b]. NSOs [OMGT05]. Nuclear [Gar93]. Nudging [CBPAK22]. Null [BS89, Kro80c]. Null- [Kro80c]. Nullity [NW06]. Number [Ano79e, Dod96, EH95, EKIKW77, HP22, Ios76f, Kei87, Mer09a, MV22, Sow72, Sow78, Sri73, HL74, Dur23a]. Numbers [CIS06, CK19, Cro11c, Der77a, Dur23a, EH92, QS87]. Numerical [FGC03, Mac14, Pun12b, Háj72, Ios76d, O’B12h, Pod21b, PS13, Rob11f]. Nunnari [Hau18]. Nyberg [Alb19]. Nyhoff [O’B13f].

O [Bar84, Cro10d, Dav82, GN14, Gho09h, Huz80, Ken73b, Ord72, Pol11a, Som77b, Tin78e]. O.R. [Tip75a]. Oakes [AG81, BC81, Cox81, Joh81, Kei81]. Oberste [Dur13a]. Oberste-Vorth [Dur13a]. Obituary [Büh75, Dav82, Ios86]. Objective [AAF18, Hoi13, Ham77b]. Objectivity [Goo78]. Obradović [Mac76]. O’Brien [Num13b]. Observation [Blu18]. Observational [Gre06, Lis10a, Mai22, O’B12e, SS88]. Observations [FHTV93, GM19, Ken97, PCA^{+}19, Sri73, Hal86, JB87, Loy80, Sco77b, Spr75b]. Observed [NK14, Sør04, Han09c]. Obtained [vdBZT20]. Occasion [Csi08]. occupancy [Hol86]. Occupational [HK77]. Oceanography [BEW03]. O’Connor [Bil77b]. Octav [Ios86]. October [WG97]. Odd [Gho09h]. Odds [vdL03, ZT13]. Odeh [Fin80c, Buc77b, Buc78]. Odell [Goo73]. Odile [Han10b]. OECD [Wyc03]. Oesterreich [Kro80c]. Office [Han07a, PP98, Ano72q, Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76o, Ano76n, Ano77l, Ano77m, Ano78m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j, Pul97]. Offices [CN98]. Official [Fel04, JS03, Kit03, LA96, LH04, Mal05, Ole03, Ram03, TC15, Mar21]. Offspring [Ste95]. Ofungwu [O’B14c]. Ogawa [Fin77c]. Ogihara [Kau12]. O’Hagan [Han10h]. Ohno [Han13m]. Ohno-Machado [Han13m]. Oil [Pos80e]. O’Kelly [Muk14a]. Olav [Alb19]. Old [SHLW22]. Ole [Bil80b]. O’Leary [Mar07b]. Olive [Ric10a, Spr75a]. Oliver [Spe08a]. Olivier [Gho11b]. Olkin [Pun11b, Cro08b]. Olsen [Ric13c]. OLSEs [Tia07]. Olver [Cro11e]. Oman [CBPAK22]. Omission [IMG22]. Omofilia [Bil80a]. O’Muircheartaigh [Pos80d]. On-Line [Mac97, Tri09, Vov01]. One [AS72, Cob07, Fin80d, KL09, LV14, Mic19, YB20, Don86]. One- [Fin80d]. One-sample [LV14]. One-Sided [KL09]. one-way [Don86]. Onefold [SH21]. Onghena [Man07]. Onicescu [Ios86]. Online [CD18, MDS^{+}19, TRRL23]. Only [Liu14b, Hed20]. Opening [Gro95, Mar96, MBW^{+}23, Wak94, Wak95, Wak96]. Operating [CMH19]. Operation [TTV05]. Operational [Feu16, Gan72a, IK86, O’B13b, Tip75a]. opérationnelle [Ios77b]. Operations [Aoy72, vC75, Pre77b, ZG12, Cam07, Luo07]. Operative [CS88]. Operators [Pun10c]. opinion [JOT85]. Opinions [FMR^{+}20]. Oppermann [Hoe83]. ’Optimal [RN92, BMM08, Dem18, DW15, GH87, LV14, LCL^{+}23, Loy76, MCH13, Rao80c, RDMP16, Sen76, War76, Bic13, FMS12a, Dra12d, Dra14b]. Optimality [Arm85a, Arm85b, BCL^{+}85, Bat85b, Che85b, DNC06, Lou85, Pet85, Sch78, Whi85, Hig77, Fri03]. Optimisation [Els18, HK18]. Optimisation-based [HK18]. Optimization [AM14, FD02, Gho21, Hun14, LCZ14a, Rob14, Bol07, Gho20, Pun12b, Feh07, Man75]. Optimizing [GM91]. Optimum [ACV84b, ACV84a, Rao77, Atk88, Puk07]. Option [AET04, Jau14, Sco13b]. O’Quigley [Cro08d]. Orchard [Sta13a]. Ord [Han09b, Pol13c, Ber78, Eng78, Kha74]. Order [Bha91, BCvZ81, Hua89, SCW08, Tan92, Vin73, dO79b, Har79, LL87, Sch83, dGAGR85, Sar73]. Order-Restricted [SCW08]. Ordered [AN01, FX16, Mai16b]. Ordering [Fom88, Ros10b]. orderings [Puk87]. Orders [Pun10c]. Ordinal [CQOR^{+}22, GT09, GR90, dMRAMR18, TB22, WC98]. Ordinary [Dem18, GG17]. ordination [KDTP20]. Organisation [Ano72l, Ano72m, Ano73i, Ano73j, Ano73k, Ano74j, Ano74k, Ano74i, Ano75m, Ano75n, Ano75l, Ano76j]. Organisational [AG99, DFV99a, DFV99b, TC99]. Organization [Ano72k, Ano72l, Ano72m]. Organizing [Sco07]. Oriented [Fil11]. Origami [Pol12b, Mer09d, Pol12b]. Origin [FZ09, Han12s, Sta10b]. original [Els20]. Origins [Mer12b]. Orkin [Som76c]. Ørnulf [Gho09h]. Orthogonal [Nar79, Tju84, TBSW84]. Oscar [Fil09a]. Österreichischen [Mui73]. Other [CB00, Dra09d, Rou07, AL86, Hol86]. Ottestad [Man72]. Otto [Für73b]. Our [Gho23, HHD97, Hoe99, Pol13a, Hau08b]. Ourselves [WPRP17].

Out-of-Sight [Far13]. Outcomes [BD20, Nee15, SAA22, SCC14, AHZK15]. outdated [TTB86]. Outliers [MGMRPA90]. Outlook [Dra14d]. Output [Dix73, KLTG21, Bil77b]. Ove [Nad74]. Oven [RM93]. Over-Parametrization [PPMS11]. Overdispersion [MSS01]. Overview [Che24, FBGGM17, Gri15, Har99, HCV^{+}21, IKS^{+}81, Mar11, MW08, Mur94, RB90, TC07, YYY20, Zhu14, EMR88, FMS12a, Gri87]. Owen [Buc77b, Buc78, Fin80c, O’B09c, Ken76c, Pos80e, Buc76a, Buc76b]. Owen-Jones [O’B09c]. Oxford [Blu20, Blu21, Cox19, Han10h, Liu22b, McC22]. Ozaki [Gho13d]. Ozone [SYSK02].

P [Arm85a, BCL^{+}85, Bat85b, Ber78, Boo79, Bre80, Buc77d, Cha14, Cha80, Che85b, Cor80, Cor75b, Cox07, Cro10h, Dra09c, Dra13a, Eng78, Gal80a, Gan78b, Gho11d, Gho11b, Gho11a, Gho13b, Goo73, Gow72, Hal73, Han11h, Hee16, Hoe83, Ios76d, Ken73a, Kin75c, Liu07, Liu14b, Lou85, Mai10a, Man72, Mar07b, Num20, Num19, O’B09i, O’B12e, O’B12f, Oga73, Oja11, Ord80a, Pet85, Pol13a, Pos78b, Pos78e, Pos79c, Pun10c, Rad75, Rad77, Roh80, Ros09a, Ros09c, Ros09b, Ros14, RS79, Sco10a, Smi75, Som76b, Sta11a, Sta11b, Sun22, Tan07, Tip80, Tra09, Tso20, Veh07a, Whi85, vdL80d, Pod22a]. P-splines [Pod22a]. P. [Aal82, CV82, Duf82, Har82, LT82, Sen03]. Pace [O’B09d]. Package [Ait18]. Packages [MW76, Nel74]. Paenson [Ken74]. Pagendam [Ano21]. Pagenkopf [vdL80b]. pages [Alb19, Ana21, BPV21, Blu18, Blu20, Blu21, Cap19, Cox16, Cox19, Dek22, Dur17, Dur18, Dur21, Dur22, Dur23a, Dur23b, Els18, Els20, Fis18, Fis19, Fis22, Fis24, Gho17, Gho20, Gho22, Gho23, Hau18, Hee16, Kha17, Kha20, Kha24, Lan18, Liu17, Liu18, Liu19, Liu20a, Liu20b, Liu21, Liu22a, Liu22b, Liu23, Mai16b, Mai17, Mai18, Mai19, Mai22, Mar21, McC22, McE17, Mer17, Mer18, Mer20c, Mer20a, Mer20b, Ng21, Num20, Num23, Num19, O’B20, O’B21, Pod17a, Pod17b, Pod17c, Pod18, Pod21a, Pod21b, Pod22a, Pod22b, Pod24, Qiw20, Ric17a, Ric17b, Ric18, Rob19, Rod21, Ros19, Rui17, San17, Sen23, Slo23, Slo24, Sun16a, Sun17, Sun18, Sun19, Sun22, Sun24, Tso16, Tso17, Tso18, Tso20, Unw20, Unw21, Unw24, Var19, Var20, You19]. pages [You20, Zha16b, Zha17a, Zha17b, Zha18, Zha19b, Zha21, Mai11c]. Paik [Han13f]. Pairwise [DFG^{+}94]. Pajevic [Mai12c]. Pal [Dra14b]. Palaeomagnetism [SHLW22]. Pallavi [O’B12d]. Palm [CMW17]. Palma [Bha07]. Palombi [Han12t]. Pandemic [LC20, NSZH22]. Panel [CHXY19, HM04, He14, Sti05]. Panorama [Mer17]. Paolo [BBS07, Gho11a]. Papatoni [vdL80d]. Papatoni-Kazakos [vdL80d]. Paper [Aal82, AG81, ABGK82a, And86, Ano84a, Atk82, Atk84, Bai84, Bar82, Bar83, Ber82, Box84, BC81, CS20, Cau86, CW73, Clo86, Cox81, Cox20, CV82, Dar81, Dar84, Dei83, Din82, Duf82, Eng83, EJ86, GS86, GH87, Gra83, Har82, Hey81, Joh81, JK82, Kei81, Kei87, Ken86, Klo83, Koe79, LT82, Miz83, New81, New83, New84, New88, PS85, Rip84, Sen84, Sha82, Sil82, Smi82, SS88, Spe84, Spe85, Spj84, Tam84, Tho84, Wah84, Wyn84, Zad82, And87, GS87]. paperback [Fis19, Gho23, Liu20a, Liu22b, Mai21, McC22, Num23, Pod24, Sen23, Slo23, Slo24, Tso20]. Papers [BBG^{+}76, Buc77a, Coc73, Had10a, Had10b, Huz80, Ios77c, Pod72, Spe10d]. paradata [ZTK13]. Paradigm [Hoe99, KMNS05]. Paradigms [CKK^{+}11b]. Paradis [Fis22, Mai09b]. Paradoxes [Mer09c, Mer17, Dra13c]. Parallel [Mai16b, Mur94, FX16]. Parallels [Ano88a, FT87, FT88]. Parameter [Blu05, Cox13, DFG^{+}94, Efr13, FS08, Fra13, Hou88, KS89a, MSW76, Par13, Rob13, SH13, UT10, Joh77, Joh80, gXS13, gX13, Cro07b, O’B21, dO79a]. Parameterization [vdL03]. Parameters [BMM08, DQS92, GT86, HH93, Jon15a, Sen76, RLS23, Abe15, AG15, Jon15b, LP15, Ma15, Pew15, SR15]. Parametric [BGS24, CFP22, Cro08b, Frö08, HM90, Hjo92, Sør04, TB22, FMR99, Gre87, KASG80, Lia20, SW15, Zhu14]. paramétrique [Col81]. parametrischer [vdL80b]. Parametrization [PPMS11]. Parent [O’B13e]. Pareto [SDT^{+}23]. Park [Han13n, Par05]. Parsimony [McL93]. Parsons [Buc79]. Part [Ano78g, HT92, Lee01, Roh76b, STM95, Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, IKS^{+}81, IKS^{+}82, LSS80b, Nor73, Tho78a, Tho78b]. Parthasarathy [Ios79c]. Partial [FBGGM17, Han12g, Mai16b, Oak81a, Oak81b, Pun10c, San21, SBB^{+}12, SS17, FX16, LHK78, LN78, Rém84]. Partially [Lia20, Pat75b]. particle [Gri87]. Particular [DS78]. Particularly [Red74]. Partitioned [Tia07]. Partitioning [Mai11f]. Pascal [Edw82, Edw83]. Past [Kal19, Khu00, Lin14a, Mai15]. Patel [Han07a]. Path [SLY23]. patients [MMC13]. Patil [Ber78, Eng78, O’B12e]. Patricia [Cro10c, Han08c]. Patrick [Els18, Man07]. Patrizia [O’B10c]. Patsy [Pol12b]. Pattern [AT15, Han07e, Han08a, MW81]. Pattern-mixture [AT15]. Patterns [AC21, BCG07, CQOR^{+}22, DC83, DMMG22, Han07b, MM01, MV22, NSS73, RDM^{+}19, Unw11c, NSZH22, Han08l]. Patterson [Gho13c]. Paul [Blu18, Dek10, Der77b, Dur14a, Fis19, Gho09b, Gho11d, Gho12e, GP11b, Han08g, Här08, Ken75, Kos12b, Lis10a, Mai10a, Mai22, Oja08b, Pod22a, Spe10a, Sta12a, Unw12c, GN20, RS79, Sau80, Van05]. Paulson [Gho09e]. Pavel [Kos12c]. Payne [Pos80d]. pbk [Liu23, Pod24]. Peace [Num13a]. Pearl [Tso20, Gho11c, Sin02b]. Pearson [Dav09, Ken73b, Bel09, Bly94, Bor09, CR89, FZ09, GL09, Mag09, McC85, Nay09, Pla83, Sen09, SS09a, SS09b]. Peck [Khu13a]. Pedagogy [Gar97, Gho97, Haw97b, Ket97b, Moo97a, Moo97b, Sch97]. Pederzoli [Ken80]. Pedigrees [She00]. Pedroso [Lis10b]. pei [Gho09f]. Peltzer [Mic77]. Peña [Cro09c]. Penalisation [LD24]. Penalized [GT09, Gho11d, Gre87, YL17]. Pennoni [Pyn14b]. Penttinen [Han08l]. People [Mar21, Rou07]. Percentage [Gue08, NK06]. Performance [AG99, DFV99a, DFV99b, FMP20, Han12a, Han13m, KCB00, MNS77, TC99, dV99]. Pericchi [RS23]. Period [O’B15c, San15]. Periodic [CC08, McL93, SM89]. Periodogram [Har90]. Periodograms [NK14]. Perkins [Mer12b]. Permutation [Ber14, HG21, HS90, Ros15a, Ros10b]. Permuted [Bar72c]. Perna [Sta12b]. Perplexed [Han13b, Kok07]. Personal [Cox97a, Cox97b, Cra81, Dem72, Gut03, Haw97a, Kal19, Pol13a, Teu97, Sch80a]. Personality [GG17]. Personelle [Oni79, Oni80]. Perspective [Dra13e, ESS10, Fro06, HPHdAM22, Kal19, Ken97, SGLS02, ZYW23, CH13, PWKO24, Dra14c, Gho08f, Gho08h, Han10d, O’B09h, Pod22b]. Perspectives [Ald98, Dur13b, Mer17, Roy14, Sta10b, Buc77a]. Perturbed [Har90]. Peruco [WvdH06]. Pesarin [Ros10b]. Peston [Tin80b]. Peter [BPV21, Bla07, Cox12, Cro11a, Csi08, Dra12d, Fis19, Gay11, Gho10b, Gho12g, Gho13c, Han07a, Hau08b, Ios76b, Jau14, Kos10b, Kos10a, Man75, Mer17, Nor09b, O’B09g, Oja08a, Pod11, Pol11a, Pos76b, Ric11b, Sen11b, Smi77, Veh07b]. Petersburg [Sen14a]. Petocz [Mer17]. Petri [Csi08]. Petris [O’B10c]. Petrone [O’B10c]. Petros [Pod21b]. Pfaff [Sco09a]. Pharmaceutical [Don12, Sta07e]. Phase [Cra81, Dem77, Kim14, MM96, SS87, Sha80]. phases [SS87]. Phelps [Sco07]. Phenomenology [Han12t]. Phil [Bat08a, O’B12c]. Philip [Bin07, Gho08g, Häg12a, Han13a, Mer07, Roy14, Rui72, Spe10c]. Phillip [Dra13f]. Phillips [Tin80b, Tin80b]. Philosophical [Sta10b, Gho10b]. Philosophy [Gre06]. Photograph [Ano72j, Ano74h]. Phylogenetic [O’B14b]. Phylogenetics [CKL14, Lar15]. Physical [Lis09a, Rob19, Sta11a, Sta11b]. Physiology [Gho13c]. Pickle [Unw11c]. Pictorial [lRG05]. Piera [Gho11a]. Pierre [Han13g, Has08b, Ios80b, Lin80a, Liu19, O’B12h, Unw13b]. Pierre-André [Unw13b]. Piet [Dra08a]. Pieter [Veh08a]. Pietro [Dur22]. Pigeot [Mat07]. Pilar [MG05]. Pilipenko [Han10j]. Pindaros [Val73]. Ping [Dur17]. Pinto [Dur14a]. Piotr [Kos12a, You14]. Pitman [Ios77c]. Pitts [Kos13d]. Pivotal [Ano96d, Bar95, FB96]. Plackett [Bar77c, Dem73]. Plain [Ric11c]. Plan [Cle01]. Plane [CQOR^{+}22, Dra12a, Pun11c]. Planes [EGS72]. Planning [SPC73, Som78f]. Plans [Dru11]. Plant [Gar93]. Plate [Cha93]. Platen [PS13]. Plausible [Yim90]. play [HJ83]. Players [Dav72]. Pleasures [Spe11b]. Pleuss [Lin75b]. Plots [EH91, LN74, NL75]. plusieurs [Lin80a]. Pocket [Sta12e, Tha07a, Fin80c]. Poggi [Mai21]. Poggiolini [Mai14a]. Point [BM89, BS89, CMW17, DM73, DC83, DFG^{+}94, DMMG22, EM19, Gal80b, Gho09h, Han08l, MM01, Mic19, MV22, NT72, RDM^{+}19, Sie88, Val08, GT12, Ham77b, Ish81, Kal84, Sha80, dF79]. Points [DMMG22, Edw82, Han08c, Has72, MV22, NK06, Sma88, Sør04, Han10b]. Poirier [Tin78d]. Poisson [BP12, BD09, BDS04, Dem07, EGS72, HH93, KX05, SGM21]. Polansky [Han09c]. Polaschewski [EV73]. Policy [Ano99e, AdV99, KFL76, Liu22a, MA05, ST92, Smi01, WB97, Wyc03]. Political [All72]. Politis [Mer20a, Qiw20]. Pollard [Rus74]. polls [JOT85]. Pollutants [DR93]. Pollution [Doc22, Ney75, DSC03, vBS73]. Pólya [Ma17]. Pólya-Type [Ma17]. Polynomial [MB19, Nar79]. Pommerich [Han08g]. Pons [Han10b]. Pooling [XK15]. Popescu [Gan80c]. Population [AdB04, And73a, Ano78k, CT11, Che85a, Chu74, CM12, Cro11g, Dek94, EB92, Fis10a, Fra80, Gan78a, Gho11b, GS19, Gol04, HH93, JSW11, KT21, Kin77a, Lav14, LdBP90, LH04, LG04, LS04, Mis10, NKH91, O’B15b, O’N04, Red74, SSOL04, Sär72c, Som79a, SS07, Sun17, TG94, Tho19, ZL22, ZBA76, Zha96, GT86, WGKL22, Fis22, Som78d, Som78e]. Population-Averaged [NKH91]. Population-Based [JSW11, Fis10a]. Populations [Ano86a, Cla00b, Has08d, HK18, Hoe74, Hor06, LLT04, Meh15, dBPR83, PS85, RBR85, SAA22, Ver13, MPR12, dBPR86, Ken76a, O’B20]. Port [Ios76b]. Porter [Dav09]. Posed [Wyc03]. Position [Cox97a, Cox97b, Teu97]. Positive [Sch16b]. Positively [MR10]. Possible [Cob07]. Post [GC95, JOT85, Jag86, Mon87, NA16, NZY18, ZL22, Kha24]. Post-Data [GC95]. Post-randomisation [NZY18]. Post-randomization [NA16]. Post-sampling [Mon87]. Post-Shrinkage [Kha24]. Post-Stratification [Jag86, ZL22, JOT85]. Postcensal [Ano80e, PK80, Tam82]. Posterior [Gou94, HH14, HK16, Koe16, Men16c, Smi16, WS16, YWH16a, YWH16b]. Postgraduates [Sco07]. Potential [BD20, O’K12, Rei12, Ski12a, Ski12b, Wil12]. Poul [Liu12a]. £45.99 [Sun17]. £49.99/$66.99 [Mai17]. £49.99/$79.95 [Mai16b]. £67.99 [Rui17]. £78.99 [San17]. £82.00 [Dur17]. pour [AM93b, Pre79b, Hor79]. Pourraient [Mal97]. Pourraient-ils [Mal97]. Poverty [BJO^{+}98, BGV18, DH19, Esp15b, HRR07, VF00, Lon15]. Power [Hoo05, Jon07, Rya13, ZG12, Kha80, Kim15]. Powers [LV14, Unw10b]. pp [Mai21]. Praag [Fis72]. Practical [CHXY19, FHTV93, GNM21, Han11h, HMA^{+}18, Nee09, Nor98, Pre80a, Sco07, Sta07e, Sta12c, Sta13b, Sun16a, Wan16, Zha16c, ZW19, Pod22a, Rob19, Stu72]. Practice [Bre99a, CWCT19, Der77a, Dra08b, Fin79, Fis18, Fon13, Kal83, NL07, Rum02, Sar08, Utt21, BPV21, Dur22, Han12m, Kos13d, Loy77, Rob11d, Som76b, Spe10c, Sta08a, Veh08b, Veh11b, Mai16a]. Practices [O’B07c]. Practitioner [Han13a]. Practitioners [Var19, Muk14a]. Prado [Rob11h]. Pragmatic [Pod15a]. Prague [Ios79b]. Prakasa [Sta10a]. praktische [Lin75a]. Pranab [Lis10b]. Prasad [Liu14a]. Prasanta [Ano72j, Rao73]. Pre [GMZ03, TTV05]. Pre-Eminent [TTV05]. Pre-Employment [GMZ03]. Precise [BM91]. Precision [CHR18, Der77a, SB17, Yoo21, Ham72a]. Precursor [MB19]. Predicting [GOZV94]. Prediction [Ano86a, Bre99b, CS20, Efr20a, FHT20, Hau08a, Law07, MS23, NC10, Nee09, dBPR83, Pod21b, RBR85, gXZ20, Mon87, dBPR86, Nor09c, Slo24, Spe11a, Pos78c]. Prediction-Based [Bre99b]. Predictions [Dra13g, Pet07]. Predictive [Bra14, Cap19, KLTG21, LK16, LG08, MNS77, Sch16a, Unw21, PWKO24]. Predictors [GT09, Dra11c]. Preface [SW07, SS09a]. Preference [KFL76, NSS73]. Preliminary [AS72]. Premadhis [Dra14b]. Preparation [Sta07b]. Presence [Bai95, Bie95, CMH19, Frö08, Ful95]. Present [Khu00, Lin14a, Mai15]. President [Ano72j, Ano74h, Fel04]. Presidential [Spe10e]. Press [Blu18, Blu20, Blu21, Cox19, Dek22, Dur18, Dur23b, Fis19, Gho20, Kha20, Kha24, Liu19, Liu20a, Liu22a, Liu22b, Mai16b, Mai19, McC22, Mer18, Mer20c, Mer20b, Ng21, Pod18, Pod22a, Pod24, Qiw20, Ric17a, Rob19, Rui17, San17, Sen23, Sun17, Sun19, Sun24, Tso17, Unw20, You20]. Pressures [Fel04]. Prevalence [GDSN18, HG94, WOP20]. Previously [AK95]. Pribram [Ano75j]. Price [Bal95, Cra80, Dal98, MCH13, Smi21, Whi99]. Prices [MFAG10]. Pricing [TS06, JT13, Jau14]. Prime [AET04]. Primer [Dra12b, Mai09a, Pun08a, Ric10a, Rob19, Häg12b, Liu08, Pre80a, Tso20]. Primers [HRR20]. Princeton [Dek22, Gho20]. Principal [CD18, FBGGM17, Vid15, Zha16b, Kha17]. Principle [GM91, Lee01]. Principles [Han12m, Hey81, Kos13d, O’B07c, Ole03, RC15, Rob11d, Fon13, Han12k, Lis12, Mai11d, Pol13c, Tso17]. Pringle [Tra73]. printing [Sun19]. Prior [Ano83a, HM19, PW91, PF82, RCA18, Rao77]. Prior-free [HM19]. Priorities [Edi77]. Priors [BS89, BW12a, DER03, FMR99, For22, Gou94, BCZ12, BW12b, Dat12]. Private [FWC^{+}21a, FWC^{+}21b]. Probabilistic [Bin07, CHR18, Gol04, GK01, Ios72, LO04, LS04, MS23, O’N04, SSOL04, Els18, Gan80b, Gho08g]. Probabilité [Mar75]. probabilités [Van16, Rao80a]. Probabilities [AS73, Til98, Ham77b, HB80]. Probability [Ald00, Bel88, Bel00, Buc77a, Cac74, Cra81, Cro10d, Dan87, Fin77a, Fis18, FG72, Gan72b, Gan80c, Gho17, GS02, GL14, Han13f, Hau08a, Hol84, Ios76c, Ios76b, Ios77d, Ios77c, Ios78, Ios79c, Jef74, Ken73c, Kos12b, LH73, LN74, LSS80b, Lin82b, Lin82a, Mai14a, Mer09c, Mer20a, Mon01, NK08, NL75, NT72, Pod14b, Pun10a, Roh76b, Sär72b, SB17, SM93, Sha85a, Sha85b, Sri73, Sta10a, Ste95, Sti14, Str76, Syl14, TC84, Var14, Veh07b, Wat98, dF79, vdL80a, BBvdB18, CP80, LSS80a, Mor74, Sun86, dM84, Sti88, Cro10a, Ios80a, Ken76c, Kin75b, Lin80b, Oja09, Roh76b, Bat08b, Dow72, Gho09g, Gho12c, Han08j, Han13k, Ios77d, Mai07, Mai10b, Mer18, Pod11, Rob11a, RR79c, Sen15]. Probable [Fér72]. Probably [Hau08a]. probit [DFM^{+}22]. Problem [AK95, Bre80, DS78, Edw82, Edw83, Ell91, GK16, Hua89, Liu10b, dBV82, MPR12, Sha80, Sun86, Fre83]. Problem-Solving [Bre80]. Problèmes [Dag76, Ios80b]. Problems [Cro10e, Gho12c, GPW05, Hal89, Han13g, Ken10, KTM01, Mer14, Nar74, Pat73, PGV22, Sen72, Sie88, Smi73, Sør93, Sta13a, Yim90, Dag76, Hol86, Sta13b, Fer80a, Pod15a]. Procedure [Ber17, Hoo05, SCW08, SH23]. Procedures [AT15, CW73, Cro08c, HTW01, KY91, Loy76, NPT08, Sen72, Tho78b, AL86, Don86, Ham72a]. Proceedings [Cha80, Dur13b, Ken75, Pos78b, Pre78b, Sch80a, Gan72a, Gan80c]. Process [AH17, AAF18, Ani08, Bha91, Bol07, Dur15b, EM93, FHTV93, Gho09h, Gho21, Had10c, Mac97, SAP97, Tip75b, Tso16, YB20, GT12, dGAGR85]. Processes [ABGK82a, ABGK82b, AP94, Ano84a, BM89, BNS94, BF91, CMW17, Cor75b, Cro09c, Cro11g, DM73, DFG^{+}94, EM19, EPP08, Gal80b, God97, Han10j, Ios79b, KS89b, Ma17, MM01, Mic19, NK14, Nor13b, Num09a, Pod14b, Pod17b, Pos79a, Rad77, Sør04, Val08, Wel92, CH13, EMR88, Gob16, GT12, Ish81, Kal84, Sør83, Sør86, Coh75, Han12o, Ios76e, Rad75, Rao80c, Rév80, Sto75, Tso18]. Processing [Dek94, Gho12e, TG94]. Processus [EMR88, Ios77b]. Producing [WvdH06]. Product [EPP08, Kha80, Lan18, Law98]. Productifs [Mal97]. Production [CR93, SPC73]. productive [Mal97]. Productivity [Har99]. Products [BJPT99, MFAG10]. Prof [HR21]. Profession [BW20, Car00, Ham77a, Wei05, Goo78, Moo78]. Professional [Ano86h, Dem72, Fis05, Kei05a, Sti05, Wat98]. Professionalism [Fis05, Sti05]. Professor [Ken75, DS20, LgX13, gXZ20]. Profile [O’B09e, Zha10]. Program [Ano05h, Pat79]. Programme [Azo01, Ins73]. Programmed [Pos78a]. Programmer [Sta07d]. Programmieren [Dor73]. Programming [DW15, Kha20, Nor09a, O’B09c, SPC73, Fer80a, Fin77b, Veh14, O’B13f]. Programming-Based [DW15]. Programs [MW76, Mor80a]. Progress [Hor16, Ruk95, SDGD^{+}11b, Uni78, dV01, Sta12f]. Progressing [Sti00]. Progression [SJFM95]. Projection [Zha19a, ZYW23, Pun11d]. Projection-Based [Zha19a]. Projections [LS04, O’N04]. Promotion [ANK03]. Proof [BJS92, Kay03]. Proofs [Han11c, Rao06]. Propensity [ZZLZ15, WGKL22]. propensity-score-based [WGKL22]. Proper [Nie03a, Nie03b]. Properties [Alv20, CT11, FH89, GKV19, GKV20, HHMR18, Kab09, Smi72, SS07, YK20, RR90]. property [GT12]. Proportion [KL09, NY19, WHT24, And76]. Proportional [Bre75, Kuk20, ZT13, Cro08d]. Proportions [Ano75f, Pla75]. Proposals [Azo01]. Propriétés [RR90]. Proschan [Str76]. prospects [MCF12]. Prosperity [FL05, OMGT05]. protected [CO14]. Protection [Lan76b, vBS73, OC13]. Proteomics [Gho17, Bla07]. Protocol [Dor18]. Providence [Sho85]. Providing [Fou80a]. Provinces [She06]. Proxy [AT15]. Przemyslaw [Unw21]. Pseudo [BKBP21, CBPAK22, WS21]. Pseudo-Bayesian [WS21]. Pseudo-ranks-Surprising [BKBP21]. Pseudo-Science [CBPAK22]. Pseudorandom [EH92, EH95]. Psychological [Sta13b]. Psychology [Veh13a, Bre79b]. Psychometric [Han11e]. Pte [Dur17]. Public [BP92, CP19, Fis05, KRR^{+}11, Smi01, Sti05, WB97]. Publication [Ano72q, Ano76n]. Publications [Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76o, Ano77l, Ano77m, Ano78m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j]. Publishing [Liu17, Mar21, McE17, Tso16]. Pukkila [Liu07]. Pulling [WPRP17]. Pullinger [PP98]. Puntanen [Liu07, Liu14b, Pol13a, Liu14a]. Purchasing [Kha80]. Puri [Oga73, Høy72]. Puritan [Bel88]. Purposes [Ber06]. Pursuit [Var14, ZYW23]. Pushing [Car05]. Putler [Sco13a]. Puzzles [Mer09c, Mer17, Dra12f]. Python [Fis24, Liu22a, Pod24].

Q [Rob11d, Var19, Dal77b, Men16b]. Q-q [Men16b]. Qi [Cro09c]. Qi-Man [Cro09c]. Qing [Han09d]. QR [Mon87]. QR-prediction [Mon87]. QTL [WS01, Cro08a]. QTL-mapping [WS01]. Quadratic [BK88, KT03]. Qualitative [GO10, GO16, Hor16, Lec76, Mac16, Men16b, OG16, PR10, Rao80c, Wil16a, Han13l, Mai11e]. Qualitatives [Lec76]. Qualities [Koe79].

’Quality [San93, Bai85a, Box93, CR93, Che85a, CS00, Cro10f, Des00, Els20, Fer94, Fis09, Fri93, Han07f, Has08a, HBK05, HH95, Mac97, Moo05, Mus75, Rao05]. Quantile [Alv20, BGS24, CL19, GKV19, GKV20, HK16, Koe16, Men16c, Smi16, WS16, YWH16a, YWH16b, YL17, BFST18, Dek14]. Quantile-based [Alv20, GKV19, GKV20]. Quantiles [DC19, NN20b, NN20a]. Quantitative [Cro08a, Dek22, DW84, Han09a, Nel72b, O’B12g, O’B13d, Sen72, SGLS02, Hau08b, Sta12d]. Quantum [Pun12b]. Quasi [GH87]. Quasi-Likelihood [GH87]. quelques [EMR88, Ios80b]. Quenouille [GT12, Spr80]. Quensel [Kul78]. Questioning [Kay03]. Questionnaire [O’B12b]. Questionnaires [Can95, Sch95, SVCV88, Sin95, Veh08c]. Questions [Lin14b, Pun11c, Sti14, dV99, HGA76]. Quetelet [Sen03]. Queueing [Gan78c]. Queues [Ovu80, Ios76d]. Quirk [Sta13b].

R [Aal82, Alv11, Amu73, Aoy72, Atk84, Bar75d, Bar80b, Bar77c, Ben75, Bil77b, Blu18, Box84, Bre80, Buc77b, Buc78, Cor80, Cor75a, Cox03, Cox07, Cro07a, Cro11d, CV82, Dal72, Dar84, Dem73, Dra13e, Duf82, Dur14b, Dur14a, Esp15b, Fer73, Fin80b, Fis10a, Fis24, Gho08f, Gow72, Häg12b, Haj77, Ham79, Han07a, Han12f, Han12k, Han13d, Har82, HR21, Ios72, Ios76e, Ios77a, Ios79c, Kau13, LT82, Lin80c, Lis10a, Liu12b, Liu20b, Liu22a, Liu23, Loy80, Luo09, Mai08, Mai09a, Mai10c, Mai10b, Mai12a, Mai22, MM77, Meh75, Moo75, Nel72b, Nor14b, Num09a, O’B12f, O’B13a, O’B13d, O’B13c, Pet07, Pol11b, Pol13d, Pos76e, Pos78b, Pos79d, Pos80b, Pre78b, Pre80a, Pun09, Pun12c, Rao80b, Rao80c, Rév80, Ros09b, Rui17, Som78c, Som79d, Som79c, Spj84]. R [Sta07a, Sta07g, Sta12a, Tin80b, Tip77, Tip80, Tra73, Unw11a, Veh13a, Vin73, Wah84, Wim75, Cha08, Cha15, Dal77c, Hot16, KC14b, Lon15, Per14, PNR13, Wil14b, Ana21, Bat08a, Bat08b, Bau14, Bra15, Cap19, Cha14, Cro11b, Fis22, Gho22, Han13e, Han13d, Lep14, Liu18, Mai09c, Mai09b, Mai10a, Mai11c, Mai14b, Mai14c, Mai21, Nor08, Nor09b, Nor09a, Nor14a, Num13a, Num20, O’B08a, O’B09i, O’B09c, O’B09f, O’B10c, O’B12h, O’B12b, O’B13a, O’B14b, Pod18, Pod21a, Ric11b, Ric17b, Rob11c, Rob12a, Ros10b, Sco09b, Sco09a, Sco10a, Sco12b, Sco13c, Sco13a, Sen14c, Unw11b, Unw12a, Unw13a, Unw24, Veh07a, Dur15a, Liu15, Liu16, Mai12b, Nor15, Pod15b, Sco10b, Sun15, Sun16b, Unw12c, Unw13b, Voe09]. R. [Bar87]. Råde [Gan76, Kin75c]. Rabe [Eva07a]. Rabe-Hesketh [Eva07a]. Rabi [Han08k]. Rade [Gal80b]. Radermacher [Mar21]. Radhakrishna [KK21, Pun07a]. Radiation [Hau18]. Rafał [Kos12c]. Raffaele [Bar75a]. Raftery [Unw20, AMR23]. Raghavarao [O’B12d, Pat73]. Ragnar [Rei76]. Rainer [Cro10b]. Rainfall [CBPAK22]. Ralf [Dur13c]. Ralph [Dur13a, Han09b, Sta07e]. Ralston [Nel80]. Ramsay [GN14]. Rand [Sta12c]. Randall [Puk07, Sco13c, Veh11a]. Random [BM89, BS92, Bre99b, Con09, DQS92, Dod96, Goo73, Häg07, HMW21, Ios79b, Ize03, KL09, MNS77, Mon01, Pod14b, Pre89, Rév80, Rip83, Sow72, Sow78, Sri73, Sto98, SCC14, Til98, TLL04, TO17, Don86, Gri87, Ish81, BN07, Gal80a, Han08k, Mai21, Mer20c, Mer20b]. Random-Cluster [Häg07]. Random-Intercept [SCC14]. Randomisation [HG21, PBH19, NZY18]. Randomised [BD20, Cro09a, God80, IMG22, NSS23, Zha19a]. Randomization [Bai85b, Cox09, Han13d, NA16, Man07]. Randomized [ACV84b, ACV84a, BD76, DW84, EUC09, Eri73, HGA76, Lan76b, Lin14b, Loy76, Sen76, War76, Zha16c, vdHvdH02, And76, Sco12a, Kim14]. randomly [PCA^{+}19, Sch84b, Sør83]. Randomness [Han11i]. Ranga [Ios77a]. Ranges [Gou94]. Rank [BKBP21, LN77, Mor73]. Ranked [Esp14a, KFL76]. Ranking [Alv14, AES18, Dav72, Esp15a]. Ranks [BKBP21]. Rao [Pun07a, Bar75d, CWCT19, HR21, Ios77a, KK21, RR21, Sta10a]. Raphael [Der75a]. Raquel [Rob11h]. Rare [Cro11c]. Rasch [Veh13a]. Rate [Pau19, SH00, Hed20]. Rates [Ano75f, BF91, GW23, Gar93, Hal89, Hoe87, Pla75, KT93]. Rating [PSI19]. Ratings [Nem73]. Ratio [CPB94, CR89, CS88, Kuk20, Kuk22, Ros87, vdL03, JOT85]. Ratitch [Muk14a]. Rattanasiri [O’B09e]. Rattle [Mai12a]. Ratto [Liu08]. Raúl [RS23]. Ravindra [Liu14a]. Ravishanker [Cha16]. Raykov [Han11e, Veh09a]. Raymond [Cro10g, Der75a, Han13c, Som80b, Tra13, Cha16]. Rayner [Tra73]. Re [Smi20, MCF12]. Re-identification [Smi20]. Read [BL24]. Reading [Rob11a, Bad82]. reading-list [Bad82]. Real [Buc77d, Jon07, Kha80, Pod15a, Spr77b]. Realisation [Mic19]. Realistic [Ryd00, Wim75]. Reality [CD00, Han12f]. Really [FPD05]. Realm [BD76]. Reappraisal [BPP17]. Reasoning [Dek22, Gho11c, Han12i, Sun18, Han10c]. Recall [CHR18, CM06, Ben75]. Recapture [HH11, LdBP90]. Received [Ano75k]. Receiver [CMH19]. Recently [Ano72q, Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76o, Ano76n, Ano77l, Ano77m, Ano78m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j]. récents [EMR88]. Recherche [Ano74m, Ano74n, Ano75o, Ano76k, Ano76l, BHBP72, Bra75, FOK73, Hor73, Nai73, Nel72a, NPT^{+}74, Nel75, Zar76, Ios77b]. Rechnungsführung [EV73]. Reciprocal [WHT24]. Recognition [Das13, Han07e, Han08a]. Recollections [Cra81]. Recommendations [Ano78k]. Reconciliation [Ale94, Roy94, Smi94a, Smi94b, Tam94]. Reconsidering [FT96]. Reconstructions [Cha93]. Record [Ano79e, CHR18, EKIKW77, Has08a, MS23, TH98]. Recorded [Mis10]. Records [Hof97, MBW^{+}23, Sun17, Sun22]. Recurrence [Han11i]. Recurrent [Han08n]. recursion [HJ83]. Recursive [Bha91, Mai11f]. Reddy [Mai17]. Reduce [Ayh01]. Reducing [CS00, LTS98, dVKW96]. Reduction [AM93a, HRR07, MZ13, Slo24, vdL03, BH80, Hol08]. Reductions [LHSD15]. Redundancy [O’B21]. Reese [Gho09d]. Refereeing [Ano99a, Ano00a, Ano06a, Ano07a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a, Ano08a]. Reference [BR02, DS78, Sen72]. Referenced [GDSN18]. Referendum [CM06]. Reflections [Ber03, Cox03, Dem83, IKS^{+}81, IKS^{+}82, KSS^{+}03b, KSS^{+}03a, Sch03]. Reflowable [Ana21]. Reformulated [Jaz72]. Refusal [SD11]. Regime [YAA20]. Regimes [LCL^{+}23]. Region [DMS06]. Regional [BCG07, Chu74, Deb78, OA91]. Regions [EGS72, Kab09, Pal72]. Registers [HS88, LTS98, NP97]. Regressing [Sti00]. Regression [Ahn14, AS72, AG76, Ano78g, Bas16, BFST18, Blu05, BF91, BGS24, BS92, CDM19, Cha93, CHXY19, Chu74, Cia14, CHPR20, CP07, DQS92, DNC06, DT92, Dek14, Dem07, DG90, Dra09c, EGS72, FZ15, FBGGM17, FWC^{+}21a, FWC^{+}21b, FGDS18, Frö08, GPdMLMGGSG20, GSGdMLM16, GT09, Gil08, Hai72, Han11g, HK16, Hon15, Hou88, HS15, Hus72, IKS^{+}81, IKS^{+}82, JR08, JR10, Joh83, KG10, Khu13a, KS89a, Koe16, KSS14, KPT05, KT03, LR91, Lis09b, Lis10c, LLP08, Liu14b, Loh14a, Loh14b, Mai16b, Men16c, Mer24, Muk14b, MB19, Nar79, NC10, Num20, O’B09h, O’B09f, OY92, PMLL15a, PMLL15b, PS85, PCA^{+}19, Pos73, Pre80b, Pun08c, Pun10b, Pun13a, RGSO17, RCA18, Ric11a, RG23, RJ91, RZ14, SCLR16, San21, STI06, Sen72, Sim98, Smi16, SZ14, SH21, SH23, Str14, Tho78a]. Regression [Tho78b, Tut21, TB22, VBL13, Vin73, WS16, WQS19, WWY23, YWH16a, YWH16b, YL17, YYY20, YAA20, dBV82, Col81, Col81, FX16, Gre87, JB87, Joh77, Joh80, LS86, NW82, Sha80, Ano21, Cro08d, Gho11d, Han11j, Kle76, O’B08b, O’B11b, Spr75a, Dra10b, Dra11c, Kos11, Num13c, Pun13c, Spr75b, Unw12d]. Regressions [Liu23, Wer12]. Regular [Mat73]. Regularisation [BDLP20, BC20, San21]. Regularity [Sen03]. regulated [Ket12]. Reha [Feh07]. Reichertz [Pos79c]. Reid [Cro07a, Gan80b, Sta17]. Reigns [KBH10]. Reilly [Sun24]. Reinhard [O’B13g]. Reiss [Cro11c]. Reiter [Sun22]. Rejoinder [Bow04b, BW12b, Cox03, CKK11a, Efr20b, GO16, GW14b, Ho05, HP17a, JR10, Jon15b, LCZ14b, Loh14b, Lum11, Nie03b, PMLL15a, Pea02c, SDGD11a, Sin02b, Sin02d, Ski12a, gX13, YWH16b]. Related [AG08, DM05, Dur14c, Gan80b, Ios76b, Liu14b, Mur94, Pos79a, Pos80c, SHGK01, Sør93, Tad80, AL86, Atk88, CB00]. Relation [FB96, Lec76]. relationship [Lec76, Bar76]. relationships [GT86, Som78d]. Relative [Kru89, KX09, XK15]. Release [NZY18]. Relevance [Moo05]. Relevant [Buc77d, FPD05]. Relever [Mal97]. Reliability [AES18, Bel02, Cox02, Gho12c, Han13j, ME89, Nai93, Ng21, Pos78e, SPC73, SW94, Sin02c, Sin02d, SAG^{+}16, Spi02, Str76, TC72, Veh08f, Gho09d, Gho08h, Law07, Mai18]. Reliable [PCA^{+}19, SCLR16]. Religious [Hau01]. Reluctant [TT20]. Remarks [FB96, Gro95, Hus72, Mar96, Wak94, Wak95, Wak96]. Remedial [SRN17]. Remedies [Cha20b]. Reminiscences [Spe08b]. Remote [CG05, OC13]. renal [MMC13]. Renata [Sco07]. Rencher [Pun08b, Pun13b]. Rene [Els18, Han07h, Zha16b, Dug78]. Renewal [Gal80b, Sun77]. Rense [O’B14a]. Renshaw [Cro11g]. Rényi [Kin75b, Sch72a]. Repeated [BE17, HS89, vdBZT20, KASG80, SSJ77, SL13, Tam87]. Replacement [AS73, HB80, Sun86]. Replicability [PS20]. Replication [Pau19, Mai22].

Reply [ACV84a, ABGK82b, Arm85a, Bat85b, Cox84b, Goo86b, HR83a, Lin82a, Oak81b, Sha85b, Smi94b, TBSW84, Goo78, Moo78]. Report [Ano78k, Car00, DD80, RvA72, Uni78, WD84, Wat74, MDG^{+}92]. Representation [Liu10a, vdV91]. Representations [Han12x, RR79a]. Representative [KM79b, KM79c, KM79a, KM80]. Representativeness [Ber17, Ber06, SBB^{+}12]. Reprinted [LSS80b, LSS80a]. Reproducibility [Pau19]. Republic [Ano05a, FL05]. Requirements [Rad99]. Resampling [BO15, DR90, Han13a, Sco12b]. Rescaling [Sar08]. Research [Ano74m, Ano74n, Ano75o, Ano76k, Ano76l, BHBP72, Bra75, Cam07, CWCT19, FWC^{+}21a, FWC^{+}21b, FOK73, GO10, GO16, GBZ07, HHDG12, Han08e, Hor73, Hor16, Ins73, Kal19, KSRS72, Lee79, Liu22a, Mac16, Mai11b, Mai11e, Mai14b, Mas73, ME93, Men16b, Nai73, Nar74, Nel72a, NPT^{+}74, Nel75, NGC^{+}78, O’B07c, O’B13b, OG16, Ott05, Pre77b, Ros19, Rum02, SH00, Sco07, Slo23, Tip75a, Wil16a, Zar76, AHZK15, DM12, Ket12, Aoy72, vC75, Cro08e, Dra13f, Gan72a, Han08c, Mai13a, Num13c, O’B08c, Veh08c, Zha21, Nee15, Szc11b]. Researchers [Cha20b, MM96, JF14, Dur17, Var19, Ple15]. residual [LL87]. Residuals [HH07, JB87]. Resolvable [Kag72b]. Respect [BS89]. Respiratory [KMM^{+}17]. Respondents [Can95, MDS^{+}19, Sch95, Sin95]. Responsabilités [Sau76]. Response [ACV84b, ACV84a, AW95, CS00, Cox97b, Dar81, DFV99b, DW84, Eri73, Gal02b, GKV20, God80, GP11b, JSW11, KSS^{+}03a, LSS08, Lin14b, Liu23, Loy76, Moo97b, NSS23, NLSS08, PGN^{+}96, Pul98, Sär11, SBB^{+}12, Sen76, Sun77, Tut21, Wan16, War76, WP99b, ZL22, dVKW96, vdHvdH02, And76, AL10, CH19, FMS12a, HCR18, HGA76, MP23, NWE16, Sik18, SL13, VT19, Cox07, Sun16a]. response-adaptive [FMS12a]. Responses [CP19]. Responsibilities [Ahl02, Gal02a, Wel02, Sau76]. Responsibility [Wil05]. Restricted [Bai85b, HH14, San91, SCW08]. Restrictions [Vin73]. résultats [EMR88]. Resulting [Hop93]. Results [BKBP21, CM06, Fel73, Piz10, RJ91, EMR88, Fed80, Fed81a, Fed81b, IK86, Shi81]. Résumé [Ano76m, Ano75e, Ano77j]. Resumes [Ano77k, Ano78l, Ano79j, Ano79k]. Rethinking [EMR22, Rui17]. Reticent [Hoe83]. Returns [FS10]. Reuven [Tra09]. Rev [And77a, BT76, Goo78, Hoy76, Moo78, Ott76, dBPR86, Sch78]. Rev. [And87, GS87]. Reversibility [Law91]. Reversible [WS01]. Review [Alb19, ABC11, All79a, Alv11, Amu73, Ana21, And78, And07, AL10, Ani08, Ano75k, Ano78g, Ano79e, Ano21, Aoy72, Aur10, Bal76a, Bal76b, Ban14, Bar77a, BNS94, BN07, Bar75c, Bar75b, Bar72b, Bar75d, Bar80b, Bar80a, Bar72c, Bar76, Bar77c, Bar77b, Bas16, Bat07, Bat08a, Bat08b, Bau14, Bec73, Beh04, BPV21, Bel07, Bel16, BPP17, Ben75, Ber78, Bha07, Bil77a, Bil77b, Bil80a, Bil80b, Bin07, Bje73, Bla07, Blu07, Blu16, Blu18, Blu20, Blu21, Bol07, Boo79, Bra14, Bra15, Bre79a, Bre79b, Bre79c, Buc79, Buc74, Buc76a, Buc76b, Buc77d, Buc77a, Buc77b, Buc77c, Buc78, Cac74, Cam07, Cap19, Cau80, Cha14, CH19, CHXY19, CW73, CIS06, Coc73, Coh75, CP96, vC75, Coo07, Cop77, Cor80, Cor75a, Cor75b, Cor77c]. Review [Cor77b, Cor77a, Cox75, Cox80, Cox07, Cox10, Cox12, Cox13, Cox16, Cox19, Cra73, Cra80, CR89, Cro07a, Cro07b, Cro08a, Cro08b, Cro08c, Cro08e, Cro08d, Cro09b, Cro09a, Cro09e, Cro09d, Cro09c, Cro10j, Cro10e, Cro10b, Cro10d, Cro10a, Cro10h, Cro10f, Cro10c, Cro10g, Cro10i, Cro11e, Cro11b, Cro11d, Cro11f, Cro11a, Cro11g, Cro11h, Cro11c, Csi80, Csi08, Dag75, Dal72, Dav09, Dei76, Dek10, Dek14, Dek16, Dek22, Dem73, Der75b, Der75a, Der77c, Der77a, Der77b, DSC03, Don12, Dor73, Dow72, Dra08a, Dra08b, Dra09b, Dra09c, Dra09d, Dra09a, Dra10c, Dra10a, Dra10b, Dra11b, Dra11a, Dra11d, Dra11e, Dra11c, Dra12f, Dra12a, Dra12b, Dra12d, Dra12e, Dra12c, Dra13d, Dra13f, Dra13g, Dra13a, Dra13e, Dra13c, Dra13b, Dra14a, Dra14b, Dra14c, Dra14d, Dri78, Dru11]. Review [Dur13c, Dur13a, Dur13b, Dur14b, Dur14a, Dur14c, Dur15a, Dur17, Dur18, Dur21, Dur22, Dur23a, Dur23b, Eag74, EFP19, Edi77, Efr13, EV73, EKIKW77, EUC09, Elf72, Els18, Els20, Emb08, Eng78, Esp11, Esp14a, Esp14b, Esp15a, Esp15b, Esp15c, Eva07b, Eva07a, FW04, FZ15, Far13, Feh07, Fer73, Fer80b, Fer80a, Fil09b, Fil09a, Fil11, Fin75, Fin77b, Fin77a, Fin77c, Fin77d, Fin78, Fin80a, Fin80d, Fin80c, Fin80b, Fis10a, Fis18, Fis19, Fis22, Fis24, Fis72, Fon11, Fon13, FCE^{+}03, For22, Fou80b, Fou80a, FD02, FBL18, Fra72, Fra13, Fre83, FGFBGM22, FST19, Gal80a, Gal80b, Gan72a, Gan76, Gan78b, Gan78c, Gan78a, Gan80c, Gan80a, Gan80d, Gan80b, GBZ07, Gay11, Gho08f, Gho08e, Gho08g, Gho08a, Gho08c, Gho08d, Gho08h, Gho08b, Gho09c, Gho09g, Gho09b]. Review [Gho09e, Gho09i, Gho09a, Gho09f, Gho09h, Gho09d, Gho10b, Gho10a, Gho11d, Gho11b, Gho11a, Gho11c, Gho12a, Gho12c, Gho12f, Gho12h, Gho12d, Gho12g, Gho12b, Gho12e, Gho13b, Gho13c, Gho13a, Gho13d, Gho17, Gho20, Gho22, Gho23, Gil15, Gir73, GG15, Goo73, GNM21, Gow72, Gow79, Gri77, Gun15, GP11b, GP12, Haa12, Had10b, Häg07, Häg12b, Häg12a, Háj72, Haj77, HN13b, Ham72a, Ham79, Han07g, Han07d, Han07a, Han07h, Han07b, Han07c, Han07f, Han07e, Han08a, Han08d, Han08o, Han08n, Han08k, Han08e, Han08f, Han08i, Han08c, Han08j, Han08m, Han08b, Han08g, Han08l, Han08h, Han09a, Han09d, Han09b, Han09c, Han10g, Han10f, Han10i, Han10d, Han10c, Han10a, Han10j, Han10e, Han10h, Han10b, Han11k, Han11c, Han11d, Han11h, Han11a, Han11g, Han11j, Han11e, Han11f, Han11i, Han11l, Han11b, Han12b, Han12h, Han12u, Han12q]. Review [Han12i, Han12f, Han12d, Han12o, Han12e, Han12j, Han12n, Han12k, Han12r, Han12w, Han12s, Han12t, Han12m, Han12v, Han12x, Han12p, Han12g, Han12c, Han12l, Han13k, Han13j, Han13o, Han13n, Han13c, Han13g, Han13l, Han13b, Han13e, Han13a, Han13d, Han13i, Han13h, Han13f, Han13m, HLC97, Här08, Has08b, Has08d, Has08a, Has08c, Hau08b, Hau08a, Hau18, He14, Hee16, Hil15, Hoe74, Hoe83, Hoe87, Hol08, Hon15, Hor06, Hor79, Høy72, Hua89, HS15, HH11, Huz73, Huz80, Iac82, Imb07, Ios72, Ios76e, Ios76f, Ios76a, Ios76c, Ios76b, Ios76d, Ios77a, Ios77b, Ios77c, Ios78, Ios79b, Ios79c, Ios79a, Ios80a, Ios80b, IMG22, Jau14, Jon00, Jør92, Kau12, Kau13, Ken80, Ken73b, Ken73a, Ken74, Ken75, Ken76a, Ken76b, Ken76c, Ken10, Kha73, Kha80, Kha74, Kha15, Kha17, Kha20]. Review [Kha24, Khu13b, Khu13a, Kim14, Kim15, Kin75a, Kin75b, Kin75c, Kin77b, Kin77a, Kir82, Kle76, Knü80, KSS14, Kok07, Kos09a, Kos09b, Kos10b, Kos10a, Kos10c, Kos11, Kos12a, Kos12b, Kos12c, Kos13d, Kos13c, Kos13a, Kos13b, Kri74, Kró80d, Kró80a, Kro80c, Kró80b, Lan18, Lar15, Law07, LS20, Lep14, LD24, LCL^{+}23, Lid80, Lin13, Lin14b, Lin75a, Lin75b, Lin80a, Lin80b, Lin80c, Lis07c, Lis07a, Lis07b, Lis08b, Lis08a, Lis09c, Lis09b, Lis09a, Lis10b, Lis10c, Lis10a, Lis12, Lis13, Lis14, Liu07, Liu08, Liu10a, Liu10b, Liu11, Liu12a, Liu12b, Liu14a, Liu14b, Liu15, Liu16, Liu17, Liu18, Liu19, Liu20a, Liu20b, Liu21, Liu22a, Liu23, LC20, Loy77, Loy80, Luo07, Luo09, MZ13, Mai07, Mai08, Mai09c, Mai09b, Mai09a, Mai10c, Mai10b, Mai10a, Mai11d, Mai11b, Mai11a]. Review [Mai11c, Mai11e, Mai11f, Mai12a, Mai12b, Mai12c, Mai13c, Mai13a, Mai13b, Mai14b, Mai14a, Mai14c, Mai15, Mai16a, Mai16b, Mai17, Mai18, Mai19, Mai21, Mai22, MM77, Man72, Man75, Man07, Man08, Mar07a, MM96, Mar73, Mar07b, Mar75, Mar21, Mat07, Mat73, McC22, McE17, MR22, ME93, Meh75, Meh15, Men16a, Men73, Mer07, Mer09c, Mer09d, Mer09b, Mer09a, Mer12a, Mer12b, Mer14, Mer15, Mer17, Mer18, Mer20c, Mer20a, Mer20b, Moo75, Mor73, Mor80a, Mui73, Muk14a, Muk14b, Muk15, Mus75, Nad74, Nan10b, Nan10a, Nee09, Nee15, Nel72b, Nel80, Ng21, Nor08, Nor09b, Nor09a, Nor09c, Nor13b, Nor13a, Nor14b, Nor14a, Nor15, NWS23, Num09a, Num09b, Num13a, Num13b, Num13c, Num14, Num15, Num20, Num23]. Review [Num19, O’B07a, O’B07b, O’B07c, O’B08a, O’B08d, O’B08c, O’B08b, O’B09h, O’B09e, O’B09a, O’B09i, O’B09c, O’B09d, O’B09f, O’B09g, O’B09b, O’B10e, O’B10b, O’B10a, O’B10c, O’B10d, O’B11a, O’B11c, O’B11b, O’B12h, O’B12g, O’B12e, O’B12d, O’B12f, O’B12a, O’B12b, O’B12i, O’B12c, O’B13a, O’B13d, O’B13g, O’B13e, O’B13h, O’B13c, O’B13f, O’B13b, O’B14a, O’B14b, O’B14c, O’B15a, O’B15b, O’B16a, O’B16b, O’B16c, O’B20, O’B21, Oga73, Oja07, Oja08b, Oja08a, Oja09, Oja11, Oni79, Oni80, Ord72, Ord80a, Ord80b, Osb91, OZW23, Pal72, Par73, PMLL15a, PMLL15b, Par13, Pat73, Pat75b, Pet07, Pin75, Pla75, Ple15, Pod11, Pod14a, Pod14b, Pod15a, Pod15b, Pod16, Pod17a, Pod17b, Pod17c, Pod18, Pod21a, Pod21b, Pod22a, Pod22b, Pod24, Pod72, Pol11b, Pol11a, Pol12a, Pol12b]. Review [Pol13a, PS13, Pol13d, Pol13c, Pol13b, Pos76e, Pos76d, Pos76a, Pos76b, Pos76c, Pos78d, Pos78a, Pos78c, Pos78b, Pos78e, Pos79b, Pos79c, Pos79a, Pos79d, Pos80a, Pos80d, Pos80e, Pos80b, Pos80c, Pre77b, Pre77a, Pre78a, Pre78b, Pre79a, Pre79b, Pre80a, Pre80b, Puk07, Pun07b, Pun07a, Pun08d, Pun08b, Pun08a, Pun08c, Pun09, Pun10b, Pun10a, Pun10d, Pun10c, Pun11a, Pun11b, Pun11d, Pun11c, Pun12a, Pun12b, Pun12c, Pun13c, Pun13b, Pun13a, Pyn14b, Pyn14a, Qiw20, Rad75, Rad77, Rao80a, Rao80b, Rao80c, RYG^{+}21, Rév73, Rév80, Ric09, Ric10b, Ric10c, Ric10a, Ric11c, Ric11a, Ric11b, Ric13c, Ric13a, Ric13b, Ric14, Ric15, Ric16, Ric17a, Ric17b, Ric18, Rob11b, Rob11e, Rob11d, Rob11c, Rob11h, Rob11g, Rob11f]. Review [Rob12a, Rob12b, Rob13, Rob19, Rod21, Roh76a, Roh78, Roh80, Ros09a, Ros09c, Ros09b, Ros10b, Ros10c, Ros10a, Ros14, Ros15a, Ros15b, Ros19, RR79b, RR79c, RR79a, Rou07, Roy14, Rui72, Rui16, Rui17, Rus10, Rus74, RDMP16, Sai08, San15, San17, San77, Sar73, Sär72b, Sat15, Sch80a, Sch80b, Sch16a, Sch72b, Sch72c, SH13, Sco77b, Sco77a, Sco07, Sco09b, Sco09a, Sco10b, Sco10a, Sco12b, Sco13c, Sco13b, Sco13a, Seb92, Sen11c, Sen11a, Sen11b, Sen14d, Sen14c, Sen14b, Sen15, Sen23, Sik18, ST92, Sil73, SDT^{+}23, Sko12, Slo23, Slo24, Smi73, Smi75, Smi77, Smi80, Som76c, Som76a, Som76b, Som77a, Som78g, Som78d, Som78e, Som78f, Som78a, Som78b, Som78c, Som79e, Som79a, Som79b, Som79d, Som79c, Som80a, Som80c, Som80b, Spe07, Spe08b, Spe08a, Spe10e, Spe10a, Spe10b]. Review [Spe10c, Spe10d, Spe11b, Spe11a, Spr75a, Spr75b, Spr77b, Spr77a, Spr80, Sta07a, Sta07g, Sta07f, Sta07b, Sta07e, Sta07c, Sta07d, Sta08a, Sta08c, Sta08b, Sta09, Sta10a, Sta10c, Sta10b, Sta11a, Sta11b, Sta12a, Sta12d, Sta12e, Sta12f, Sta12c, Sta12b, Sta13a, Sta13b, Sta13c, Sto75, Str76, Str73, Stu72, Stu80, Sun15, Sun16a, Sun16b, Sun16c, Sun17, Sun18, Sun19, Sun22, Sun24, SAA22, Szc11a, Szc11b, Tan07, TZW^{+}20, Tha07a, Tha07b, Tho78a, Tin78e, Tin78b, Tin78c, Tin78d, Tin78a, Tin80a, Tin80b, Tip75b, Tip75a, Tra07, Tra09, Tra13, Tra73, Tsa24, Tso16, Tso17, Tso18, Tso20, UT10, Unw09, Unw10a, Unw10b, Unw11b, Unw11a, Unw11c, Unw12c, Unw12a, Unw12b, Unw12d, Unw13b, Unw13a, Unw20, Unw21, Unw24, Val08, Var14, Var19, Var20, Veh07a]. Review [Veh07b, Veh08b, Veh08g, Veh08c, Veh08f, Veh08d, Veh08a, Veh08e, Veh08h, Veh09a, Veh09b, Veh10a, Veh10d, Veh10b, Veh10c, Veh11b, Veh11a, Veh13b, Veh13a, Veh14, Vin73, Voe09, Vos75, Wan19, Wel92, Wet75, Wet80, Wil73, Wim75, Yan73, You14, You15, You19, You20, Zha14, Zha16a, Zha16b, Zha16c, Zha17a, Zha17b, Zha18, Zha19b, Zha21, dO79a, dO79b, vdHvdH02, vdL80c, vdL80a, vdL80b, vdL80d, Chm81, Col81, Don86, Fis83, HB80, LHK78, MMC13, Rao77, WC75, gXS13, gX13, Daw02a, Cha16, Had10a, HN13b, PA15, Pin10a, Pin10b, SZ09]. Reviews [Bre80, Cha80, Ios77d, Liu22b, Roh76b, Som77b, Tip77, Tip80]. Revised [Spe10c]. Revisited [Ano75e, BS85, GBZ07, Gil08, LSS05, Per75, Qui21, SH96]. Revolution [Box93, Fri01, Ken76b, Rid16, San93]. Revue [Col81]. Rex [Fon13]. Rey [Pod24]. Reyes [Liu22a]. Ricciardi [Pos79a]. Riccomagno [Gho11a]. Richard [Bje73, Bra14, Cro09a, Dra09b, Gho09g, Han08n, Han12o, Han12n, Han12c, Han12l, Kok07, Kos10b, Liu21, Mai18, O’B09h, Rui17, Sco10b, Som76c, Spr77b, Str76, Tso16, Veh11a, You19, Dei83, Eng83, Gra83, Klo83, Miz83, New83, Sti88]. Richardson [Ric05]. Richter [Mor80b]. Rick [BN07, Han08i]. Ridge [AG76, Cox07, GPdMLMGGSG20, GSGdMLM16, JR08, JR10, KG10, MW81, YAA20]. Riedwyl [Dri78]. Riggs [Dur18]. Right [DER03, Ole03]. Rights [Sta08b]. Riley [Sta11a, Sta11b]. Rios [Tin78a]. Risk [CFL^{+}21, CFP22, CKK^{+}11b, Dur14a, Feu16, GOZV94, Gho08h, Han09a, Han10j, NZY18, NWE16, O’B14c, Pun10a, Zha17a, DY16, O’K12, Rei12, Ski12a, Ski12b, Wil12, Kos13d]. Risk-Utility [CKK^{+}11b]. Risks [Cro94, Han07h, JK78, O’B16a, Cox80, Kró80b]. Risky [Bec73]. Rissanen [Csi08, Lan01, Lis07c]. Rittersbacher [Lin75b]. Ritz [O’B09f]. Rivot [O’B13e]. Road [Els20]. Rob [Ano21, Han09b]. Robbins [GK16, PS20]. Robert [Buc79, Dra11d, Fin80a, Fin80c, Gho08g, Han08a, Han08h, Han11d, Han12u, Han13c, Kha80, Kok07, Lis09a, Mai11e, Mai12b, Nor09c, Nor14a, O’B09a, O’B09c, O’B09d, O’B13a, Pol12b, Pol13d, Pol13c, Pyn14a, Sco10a, Sco13a, Spe08a, Spe10d, Szc11b, Veh08d, Voe09, Cha14]. Roberto [Dur13b]. Robin [Dav72, Dra09d, Mai21]. Robinson [O’B12f, Cro10g, O’B09c]. Robust [Ano86a, AG08, BSS18, BS89, BF91, CDM19, CKS17, DS99, Dra14c, FS94, dBPR83, dBPR86, PW91, RCA18, SCLR16, Sen72, VF00, YYY20, Roy86, WGKL22, Oja11]. Robustness [Els18, Pér93, Iac84]. ROC [Han13m]. Rocci [Dur13b]. Rockette [Han13m]. Rodney [Han11i]. Rodrigues [Sch73]. Roel [Rus10]. Rogantin [Gho11a]. Roger [Han11c, Kos13d, O’B10e, He17]. Rogers [Sen14b]. Rogerson [O’B09g]. Rohatgi [Ios77d]. Rohlf [Cor75a]. Role [Bat02, Bin00, Car93, CBPAK22, Cla00a, Dil00, FL05, Fri01, Gan79, Han10i, Lyb00, MDG^{+}92, OMGT05, Pfe93, Tre00b, BNCR86, dF74]. Roles [Kei05b, Red74]. Rolf [Cro11c, Men76]. Rolf-Dieter [Cro11c]. Romania [Gan80c]. Romier [Pre78b]. Ron [Khu13b, Sta07d]. Ronald [Cro11e, Esp14b, Han11a, O’B10a, Pun11c, Som78b, Veh09b]. Rongling [Cro08a]. Ronny [O’B09e]. Root [Liu10a]. roots [MCF12]. Rosander [Som78a]. Rosaria [Fis18]. Rosenbaum [Blu18, Mai22, Lis10a]. Rosenblatt [Gan80d]. Rosenkranz [Zha21]. Rosenthal [Pos78c]. Ross [Gay11, Ios80a]. Rossi [You19]. Rota [Han12t]. Rotating [BE17]. Rotation [Sun77]. Rothamsted [Gow15]. Rothstein [Ros09b]. Round [Dav72]. Roussas [Ios76a]. Route [Smi01]. Routine [Dor18]. Routledge [Slo24]. Rouvière [Unw13b]. Roy [Lin80c, Nel72b, Dug78]. Royal [Gho10b]. Rubin [Imb07]. Rubinstein [Tra09]. Rudolf [Gho08e]. Rudy [Cro11d]. Ruey [Lep14]. Ruggles [Bje73, Bje73]. Ruin’ [Edw83]. Rule [Rob12b, TGM09]. Rules [Hau01, Lin82b, Lin82a, Pun10a, BNC84]. Ruling [SB17]. Runs [CSW93]. Rupp [BPV21, Han11d]. Ruppert [Num20, Gho12f]. Russell [Liu12b, Mai19, Zha19b]. Russian [Rob12b, Sen09]. Ruth [Blu20, Gho11b]. Ruud [Sta08c]. Ryan [Ano21, Bre79c, Bre79c, Dra09c, Tan07]. Rychlik [Kos12a]. Ryoichi [Pun10d, Sai79].

S [And77a, Bal76a, Bal76b, Bar75b, Bar72c, Bar77b, Bec73, Ber78, Bre80, Buc74, Buc77a, Cop77, Cox10, Cra80, Cro09e, Der77a, Dra11e, Eng78, Eva07a, Fer80a, Fis10a, Fis72, Fou80a, Gho09e, Gho09d, Gri77, Häg12a, Han07c, Han09a, Han13o, Han13d, Ios77d, Ken73b, Ken73a, Ken75, Lep14, Liu18, Mai10a, Mai13c, Mai16b, Man07, Mer09b, Nel72b, Nel80, Ord80b, Pol13d, Pos80c, Pre77b, Pre79a, Pun12a, Pun13c, Pun13a, Qiw20, Rév73, Rév80, Ric13a, Ric18, Rob11c, Sco13a, Sil73, Som77b, Som78e, Spe10c, Sta10a, Sta10c, Sun18, Tha07b, Tin78c, Tip77, Veh10d, Zha16b]. S1 [Ano21]. S179 [Ano21]. S204 [Ano21]. Sá [Veh07a]. Sachs [Dei76]. Sadanori [Lis08b]. ’safe [Hed20, MDS94, WvdH06]. Safety [Lan18]. Saha [Mer20c, Mer20b]. Sahami [Han10g]. Saharan [HRR07]. Said [Kha77]. Saint [Van16]. Saint-Flour [Van16]. Saisana [Liu08]. Sakshaug [Blu21].

Salazar [Han12m]. Salmaso [Ros10b]. Salsburg [Sun18]. Saltelli [Liu08]. Salvatore [Dur13b]. Sam [Law07]. Samaniego [Ban14, Rob11b]. Samelson [Spr77a]. Samiran [Gho12a]. Sample [AHZK15, Ale94, And73a, BK88, BP12, Ber06, BCD^{+}94, Coh05, Cro07a, EMR22, GO81, HM90, HH93, Hoo05, KT21, KL09, LdBP90, Lis10b, Mat73, Mur74, PHO86, Pun10d, Roy94, Rya13, Sär72c, Smi72, Smi94a, Smi94b, Som76b, Som78a, Sun77, Tam94, Tra13, vdBZT20, GNM21, LV14, Mon87, Rao77, Sch84b, Kim15, Nee15, O’B08c, Ros10c]. Sampled [DC83]. Samplers [Jab96, Kis96, Ver96]. Samples [BBvdB18, Kis79, NPT08, Tam82, TC84]. Sampling [ACV84b, ACV84a, AC21, Ano88a, AS73, BP12, BPP17, Ber72, BD20, Bre99b, Buc77c, CR93, CWCT19, CS02, CS88, Dem77, Der77c, Dur13c, FT96, Fra80, GS19, Han13c, HB80, HL01, Huz80, Iac82, Iac84, Imb07, Ize03, Jag86, Kal83, KP10, KW19, KM79b, KM79c, KM79a, MP23, Mer24, Moo75, NPT08, O’B12e, Pfe93, RN92, RSG06, RG23, dMRAMR18, SHH94, Smi73, SS88, Smi21, Tam82, TSR96, Til98, Til22, VT19, VL96, Ver13, WS21, ZHW23, BNC84, Dal77a, Dal77b, Dal77c, FT87, FT88, JOT85, KM80, Mon87, SS87, SB75, Sun86, WC75, ZTK13, Esp14a, Has08b, Pos79b, Sär72b, Som77a, Has08c, Meh15, O’B20]. Sampling-based [BD20]. Samprit [Pun13c, Pun13a]. Samuel [Ios78]. Sand [Sør93]. Sandra [Sta07c]. Sandrine [Cro08c, Dra13d]. Sanitaires [Dam74]. Sanitation [Dam74]. Santé [Dam74]. Santi [Pod17a]. Sanz [Jer78]. Saporta [Mar07a]. Sara [Cox16, Gho12g, McC22]. Sarasohn [Els20, Fis09]. Sard [Rév73]. Saris [Veh08c]. Sarkar [Nor08]. Särndal [Pos79b]. Saroj [Pun10c]. SAS [Bra15, Kha20, Liu12b, Mai12b, O’B08a, Ros15b, Voe09, BH14, Cha08, Dra11a, O’B13b, Puk07, Ric13a, Muk15, Zha19b, Khu13b]. Sasivimol [O’B09e]. SAS(R) [Sta07d, OH13, Sta07b, Sta07f, Sta07e]. Sastry [Zha16b]. Satellite [Pat79, RvA72, Wat74]. Saunders [Law07]. Sava [Mac76]. Savvy [Dur23a]. Saxer [Büh75]. Scalable [Sha23]. Scalar [RGSO17]. Scalar-on-Function [RGSO17]. Scale [DG17, GSW17, HP17a, HP17b, OCRC17, PJK17, ZM17, Bic14, Spe11a]. Scales [Tut21, Han08g]. Scaling [Sch83]. Scandinavian [GL09]. Scarf [Pos76d]. Scargle [Pod14a]. Scarpa [Mai13b]. Scenario [TLL04]. Scenarios [Gol04]. Schaalje [Pun08b]. Schaefer [O’B09b]. Schafer [Coo07]. Schedules [SVCV88]. Scheffé [Leh78]. Schelling [Shi78]. Scheme [Fal72]. Schemes [RSG06, Yas93, WC75]. Scherer [Dur13c]. Scheuer [Buc77d]. Scheuren [Has08a, Sta08b]. Schier [Gho13c]. Schipp [Fil09b]. Schlattmann [Sen11b]. Schlotzhauer [Sta07c]. Schlove [Pre79a]. Schmid [Liu17]. Schmidt [Bre80]. Schnatter [Han07d]. Schoenberg [Han13f]. ScholarOne [Had10c]. Schölkopf [Gho13a]. Schomaker [Pun07a]. School [Ald03, Blu07, Gan76, GL09, Han13g, Liu10b, Nov73]. Schools [Tre05b]. schrumpfende [Kro80c]. Schucany [Meh75]. Schumacker [Veh11a]. Schuyler [Ric09]. Schwarz [Lan01]. Sciadas [Dur23a]. Sciarra [Kró80b]. Science [Cap19, CBPAK22, Cha15, Cle01, Cle05, Cre20, DS78, Dur15a, Fis19, FL05, Han08b, KK21, Kha73, Kha20, Lin14a, Liu18, Mai10b, Mai13a, Oja08b, Pat88, Pod24, Pol12b, Spe10d, Sun18, Utt21, Wik03, YB20, All72, Bin07, Liu22b, Mai17, McC22, O’B13f, Pod17a, Sun19, Unw20, dO79a, Mai15]. Sciences [Bar75c, Blu16, Gho09c, Gho17, Gho21, Gre06, Kei05b, Mai09a, Mai11b, O’B13h, Pol13b, Pun11a, Sch80a, Sta11b, Sta12c, Sta12b, TP73, Bar75c, Bar80b, Cro10g, Gho10a, Han12n, Nan10a, Ric10a, Ros09a, Sko12, Som78c, Spe10b, Sta11a, Sta13c, Veh08g, Veh10d, Dek16]. Scientific [Ano05h, Bar80b, Ber78, Cox10, Cra81, Dav09, Dra13c, Eng78, KM79b, KM79a, LLP01, O’B09c, Pat75a, Som77a, Unw09, Par73]. Scientist [Rus10]. Scientists [Cro07b, Han11a, Tip77, Dur23b, Gho08d, Law07, Lis09a, Liu19, Mai07, Rob19, Sta12d]. score [WGKL22]. Scores [Han08g, Tho00, WR05, ZZLZ15]. Scoring [BPV21, Lin82b, Lin82a, Osb92]. Scott [Gho08d, Ric17a, Veh09b, Zha17b, Han12n]. Screening [Dem77, LHSD15, TC84]. Sean [Num13b]. Search [Arm85a, Arm85b, BCL^{+}85, Bat85b, Che85b, Lou85, Pet85, Pos80e, Whi85]. Searle [Pun09]. Seasonal [MR22, NSZH22]. Seasonality [Der75a]. Sébastien [Mai11c]. Seber [Cox75, Lis08a, Pre80b]. Second [Alv11, BCvZ81, Bra15, Cro10b, Cro10a, Cro10f, Cro10g, Cro11b, Dra09c, Dra09a, Dra10b, Dra11d, Dra12e, Dur14b, Fil09a, Gho09c, Gho09g, Gho11c, Gho12b, Han08j, Han08b, Han10e, Han12q, Han12f, Han12l, Han13j, Kos12c, Kos13b, Liu10a, Mai09c, Mai11a, Mai11f, Mai12b, Nor09b, Nor09c, Num09b, O’B08c, O’B11a, O’B11b, O’B12g, O’B12a, Oja11, Pun08b, Pun09, Pun10a, Pun11b, Rob11c, Rob11f, Ros15b, Sco09a, Sow78, Sta08a, Szc11a, Tan92, Unw10b, Unw12c, Var19, Veh08b, Veh09b, Veh10a, Veh10b, Veh13b, Ott76, Wak95]. Second-Order [Tan92]. Secretariat [Ano78k]. Secretary [Fre83]. Section [PS85]. Sectional [HK77, Pro11]. Security [PGV22]. Seeing [Lin02, Pea02b, Sin02a]. Seguí [Pod17a]. Seif [Liu22b]. Seismic [OA91]. Seismology [KTM01]. Sekhon [Spe10c]. Selecta [Cop77]. Selected [Bar72c, Che24, Han12b, Smi72, Spe10d, Tin78c, dWC05, KASG80, Buc76a, Buc76b, Buc77b, Buc78, Cox12, Dra13d, Gay11, Gho12d]. Selecting [SS17, Yim90]. Selection [Bet10, BS92, CR93, Cas20, DS22, FGDS18, FGFBGM22, HK77, Kab09, KSS14, Lan01, Lis09c, MW10, NC10, OS93, Rob11e, Tho78a, Tho78b, SS87, Ano78g, Eng78, Han08m, Fin80d]. Selective [YYY20, KS85]. Selectors [Wan19]. Self [MR03, Ric11a, WD84, Cro09c]. Self-efficiency [MR03]. Self-Learning [Ric11a]. Self-Normalized [Cro09c]. Self-Report [WD84]. Sellafield [Gar93]. Sellers [Som79d]. Seltzer [Som78e]. Selvin [Cro08e]. Semi [CFP22, Cro09d, FMR99, MP21, Gre87, Cro09d]. Semi- [FMR99]. Semi-Markov [Cro09d, MP21, Cro09d]. Semi-Parametric [CFP22, Gre87]. Semidefinite [DW15]. Semimartingale [KS89b]. Semiparametric [CHXY19, Cro08b, LSD^{+}11, Num09a, Num20]. Sempi [Tso17]. Sen [Lis10b, Oga73]. Seneta [Gay11]. Senn [Häg12a]. Sense [Dra11d, Mer07]. Sensing [CG05]. Sensitive [HG94, Lin14b, NSS23, NY19, HGA76]. Sensitivity [BD94, Liu08, MSW76]. Sentencing [Ize03]. Separate [dBP77]. separating [O’K12, Rei12, Ski12a, Ski12b, Wil12]. September [Dur13b, Gan80c, Han13g]. Sequence [Gho09g, Lef79]. Sequences [PS20]. Sequential [Arm85a, BCL^{+}85, Bat85b, BEW03, BRST76, Che85b, LdBP90, Lou85, Pet85, SM93, Whi85, Bat85a, Sør86, Bar72b, Sco77a]. Sergio [Pod24]. Series [Abr82, And77a, Ano83a, Ano21, Bha07, CD94, CHPR20, Cro11b, DCC98, Der75a, EFP19, FM95, Fis18, Gan78b, Gho13d, Gow79, Gue90, Hal00, Han12r, Han12v, HLC97, HHK03, HR83b, HR83a, HzB91, Jas15, KP04, Law91, Loy77, Mai09c, Mai10a, Mak76, Mak78, McE17, McL93, New81, New84, New88, Ord80a, Per14, PF82, Pod15b, Pol13d, Qiw20, Rob11h, RC93, Sai08, SP11, Sco09a, Sen14d, SR92, Slo24, Sun18, Sun24, SM89, Tin78e, TRRL23, Tsa24, Zha17b, Zur91, vdBZT20, Lau81, LL87, Mor84, Neu16, PMM87, SSJ77, Bar77a, Hau18, Kos10b]. Server [Kin75a]. Service [BP92, GO10, Kin75a, Law07, Luo07, Rus74]. Services [BJPT99]. ses [Esp11]. SESAME [Keu97]. Session [Ano72a, Ano72b]. Set [Esp14a, Hau01, O’B11a]. Seth [O’B09a]. Sets [BM89, Kag72a, Sie88, Sma88, Sto98]. Setting [Edi77, Roh16]. Settings [MDS94, Han13l]. Sevast’yanov [Gal80a]. Seven [Kis94, NC21]. Seventeenth [Bel88, Lan85, Sti88]. seventh [Lan74a]. Several [MP21, NAL76, WLC82]. Severe [HMW21]. Sha [Els20]. Shafer [Spe85]. Shahbaba [O’B13c]. Shalabh [Pun07a]. Shall [Moo99]. Shane [Gho09d]. Shankar [Mai11b, Zha16b]. Shannon [NC21]. Shao [O’B08c, Cro09c]. Shape [Abe15, AG15, Jon15a, Jon15b, LP15, Ma15, O’B16b, Pew15, Sma88, SR15]. Shapes [DS78]. Sharad [O’B12e]. Shared [Kis94]. Sharing [Tho00]. Sharma [Veh07b]. Sharon [Rob12b]. sharp [GT12]. Shayle [Pun09]. Shein [Gho09f, Han12e, O’B07a, O’B08d, O’B08c, O’B12a]. Shein-Chung [Gho09f, Han12e, O’B07a, O’B08d, O’B08c, O’B12a]. Sheldon [Ios80a]. Shelemyahu [Mer14]. Sheppard [MB19]. Shift [TB22]. Shifts [LQ04]. Shimi [Pos78e]. Shimizu [Pun10d]. Shing [Gho13a]. Shinko [Els20]. Shioya [Han13n]. Shirley [Dra08b]. Shiryayev [Rév80]. Shiskin [Ano79i]. Shmueli [Han07a]. shocks [SGM21]. Shonosuke [Num23]. Short [Alv11, And07, Aur10, BN07, Bat07, Bat08a, Bat08b, Bel07, Bha07, Bin07, Bla07, Blu07, Bol07, Cam07, Coo07, Cox07, Cox10, Cox12, Cro07a, Cro07b, Cro08a, Cro08b, Cro08c, Cro08e, Cro08d, Cro09b, Cro09a, Cro09e, Cro09d, Cro09c, Cro10j, Cro10e, Cro10b, Cro10d, Cro10a, Cro10h, Cro10f, Cro10c, Cro10g, Cro10i, Cro11e, Cro11b, Cro11d, Cro11f, Cro11a, Cro11g, Cro11h, Cro11c, Csi08, Dek10, Don12, Dra08a, Dra08b, Dra09b, Dra09c, Dra09d, Dra09a, Dra10c, Dra10a, Dra10b, Dra11b, Dra11a, Dra11d, Dra11e, Dra11c, Dra12f, Dra12a, Dra12b, Dra12d, Dra12e, Dra12c, Dra13d, Dra13f, Dra13g, Dra13a, Dra13e, Dra13c, Dra13b, Dru11, Dur13c, Dur13a, Dur13b, Emb08, Esp11, Eva07b, Eva07a, Far13, Feh07, Fil09b, Fil09a, Fil11, Fis10a, Fon11, Fon13]. Short [Gay11, Gho08f, Gho08e, Gho08g, Gho08a, Gho08c, Gho08d, Gho08h, Gho08b, Gho09c, Gho09g, Gho09b, Gho09e, Gho09i, Gho09a, Gho09f, Gho09h, Gho09d, Gho10b, Gho10a, Gho11d, Gho11b, Gho11a, Gho11c, Gho12a, Gho12c, Gho12f, Gho12h, Gho12d, Gho12g, Gho12b, Gho12e, Gho13b, Gho13c, Gho13a, Gho13d, GP11b, GP12, Haa12, Häg07, Häg12b, Häg12a, Han07g, Han07d, Han07a, Han07h, Han07b, Han07c, Han07f, Han07e, Han08a, Han08d, Han08o, Han08n, Han08k, Han08e, Han08f, Han08i, Han08c, Han08j, Han08m, Han08b, Han08g, Han08l, Han08h, Han09a, Han09d, Han09b, Han09c, Han10g, Han10f, Han10i, Han10d, Han10c, Han10a, Han10j, Han10e, Han10h, Han10b, Han11k, Han11c, Han11d, Han11h, Han11a, Han11g, Han11j, Han11e, Han11f, Han11i, Han11l, Han11b, Han12b, Han12h, Han12u, Han12q, Han12i, Han12f, Han12d, Han12o, Han12e, Han12j, Han12n]. Short [Han12k, Han12r, Han12w, Han12s, Han12t, Han12m, Han12v, Han12x, Han12p, Han12g, Han12c, Han12l, Han13k, Han13j, Han13o, Han13n, Han13c, Han13g, Han13l, Han13b, Han13e, Han13a, Han13d, Han13i, Han13h, Han13f, Han13m, Här08, Has08b, Has08d, Has08a, Has08c, Hau08b, Hau08a, Hol08, Imb07, Kau12, Kau13, Ken10, Khu13b, Khu13a, Kok07, Kos09a, Kos09b, Kos10b, Kos10a, Kos10c, Kos11, Kos12a, Kos12b, Kos12c, Kos13d, Kos13c, Kos13a, Kos13b, Kró80a, Law07, Lin13, Lis07c, Lis07a, Lis07b, Lis08b, Lis08a, Lis09c, Lis09b, Lis09a, Lis10b, Lis10c, Lis10a, Lis12, Lis13, Liu07, Liu08, Liu10a, Liu10b, Liu11, Liu12a, Liu12b, Luo07, Luo09, Mai07, Mai08, Mai09c, Mai09b, Mai09a, Mai10c, Mai10b, Mai10a, Mai11d, Mai11b, Mai11a, Mai11c, Mai11e, Mai11f, Mai12a, Mai12b]. Short [Mai12c, Mai13c, Mai13a, Mai13b, Man07, Man08, Mar07a, Mar07b, Mat07, Mer07, Mer09c, Mer09d, Mer09b, Mer09a, Mer12a, Mer12b, Mur94, Nan10b, Nan10a, Nee09, Nor08, Nor09b, Nor09a, Nor09c, Nor13b, Nor13a, Num09a, Num09b, Num13a, Num13b, Num13c, O’B07a, O’B07b, O’B07c, O’B08a, O’B08d, O’B08c, O’B08b, O’B09h, O’B09e, O’B09a, O’B09i, O’B09c, O’B09d, O’B09f, O’B09g, O’B09b, O’B10e, O’B10b, O’B10a, O’B10c, O’B10d, O’B11a, O’B11c, O’B11b, O’B12h, O’B12g, O’B12e, O’B12d, O’B12f, O’B12a, O’B12b, O’B12i, O’B12c, O’B13a, O’B13d, O’B13g, O’B13e, O’B13h, O’B13c, O’B13f, O’B13b, Oja07, Oja08b, Oja08a, Oja09, Oja11, Pea02c, Pet07, Pod11, Pol11b, Pol11a, Pol12a, Pol12b, Pol13a, PS13, Pol13d, Pol13c, Pol13b, Puk07, Pun07b, Pun07a, Pun08d]. Short [Pun08b, Pun08a, Pun08c, Pun09, Pun10b, Pun10a, Pun10d, Pun10c, Pun11a, Pun11b, Pun11d, Pun11c, Pun12a, Pun12b, Pun12c, Pun13c, Pun13b, Pun13a, Ric09, Ric10b, Ric10c, Ric10a, Ric11c, Ric11a, Ric11b, Ric13c, Ric13a, Ric13b, Rob11b, Rob11e, Rob11d, Rob11c, Rob11h, Rob11g, Rob11f, Rob12a, Rob12b, Ros09a, Ros09c, Ros09b, Ros10b, Ros10c, Ros10a, Rou07, Rus10, Sai08, Sco07, Sco09b, Sco09a, Sco10b, Sco10a, Sco12b, Sco13c, Sco13b, Sco13a, Sen11c, Sen11a, Sen11b, Sko12, Spe07, Spe08b, Spe08a, Spe10e, Spe10a, Spe10b, Spe10c, Spe10d, Spe11b, Spe11a, Sta07a, Sta07g, Sta07f, Sta07b, Sta07e, Sta07c, Sta07d, Sta08a, Sta08c, Sta08b, Sta09, Sta10a, Sta10c, Sta10b, Sta11a, Sta11b, Sta12a, Sta12d, Sta12e, Sta12f, Sta12c]. Short [Sta12b, Sta13a, Sta13b, Sta13c, Szc11a, Szc11b, Tan07, Tha07b, Tra07, Tra09, Tra13, Unw09, Unw10a, Unw10b, Unw11b, Unw11a, Unw11c, Unw12c, Unw12a, Unw12b, Unw12d, Unw13b, Unw13a, Val08, Veh07a, Veh07b, Veh08b, Veh08g, Veh08c, Veh08f, Veh08d, Veh08a, Veh08e, Veh08h, Veh09a, Veh09b, Veh10a, Veh10d, Veh10b, Veh10c, Veh11b, Veh11a, Veh13b, Veh13a, Voe09, ZG12]. Short-Term [ZG12]. Shorted [Pun10c]. Shortlist [OS93]. Shoukri [Han13j]. Should [Men16c, MV22]. Shravan [Rob11g]. Shrinkage [BDS04, YAA20, Kha24]. Shuangge [San17]. Shuangzhe [Veh07b]. Shumway [Pol13d]. Sib [SGLS02]. Sibillo [Sta12b]. Sick [CB00]. Sick-leave [CB00]. Side [Far13, Fis24]. Side-by-Side [Fis24]. Sided [KL09]. Sides [Sim98]. Sidney [Ios76b]. Siege [Gho23]. Siegmund [And07]. Sight [Far13]. Sigma [MW08]. Sign [Hoo05, MR80]. Signal [DCC98, Dra13g, KTM01, Sär11, SLY23]. Signals [God97]. Significance [LLP01]. Significant [CC08]. Sik [Mai09c]. SILC [Lon15, Esp15b]. Silva [Sch73]. Silver [Dra13g]. Silverstone [Tin80b]. Silvey [Bar77b]. Silvia [Hau18]. Simmons [Som78f]. Simo [Liu07, Liu14a, Liu14b, Pol13a]. Simon [Nan10a, Sen14b, Bra14]. Simonoff [Pun13a]. Simple [Beh04, CW91, KL09, LLP08, O’B12c, Pos73, SC05, Til98, KT93, Sta12a]. Simpler [Gol04]. Simplicity [Häg12a]. Simplified [BGV18]. Simulating [Dur13c]. Simulation [BDS04, vC75, DK01, Dim01, Don12, FS08, IMG22, MS87, Muk14b, NWE16, Nor98, O’B09c, Rob11g, Boo79, Luo09, Tra09]. Simulation-Based [Rob11g]. simulations [LOD^{+}85, You86, Cro11g]. Simulationsmodelle [Sch80b]. Simulationstechnik [Sch80b]. Simultaneous [EGS72, GP73, LLP08, RLS23, UT10, Sri80, Han11j]. simultaneous-equation [Sri80]. Sin [Kim14]. Sin-Ho [Kim14]. Singapore [Dur17, Pun11a]. Singer [Lis10b, Mai11f]. Singh [Han13l, LgX13]. Single [Bic14, Sri80]. single-equation [Sri80]. Singpurwalla [Gho08h, Pea02a]. Singular [Pun11d]. Sinha [Dra14b]. Sinisa [Mai12c]. Sint [Cha80]. Situations [dWvDS20]. Six [MDS^{+}19, MW08, Dra13e, Pod72]. Sixteenth [Bel88, Lan84]. sixth [Lan73]. Sixto [Tin78a]. Size [CT11, Coh05, EMR22, Hoo05, Kim15, LdBP90, Nee15, O’B08c, Pod72, Rya13, AHZK15, Rao77, TTB86]. Sizes [Han12w, Ros10c]. Skarobis [Cha80]. Skew [AG08, Dur14c]. Skew- [AG08]. Skew-Normal [Dur14c]. Skewed [HK18, Hor06, TS06]. Skewness [NW06]. Skillicorn [Han08o, Unw10a]. Skinner [HS19]. Skorohod [Pin75]. Skovgaard [Dra11c]. Skrondal [Dra11e]. Skysurveys [BM16]. Slepian [LR91]. Slott [Sch80a]. Small [BCM19, BSS18, BFST18, Bic14, CDM19, CSW93, CL19, Con09, Cro07a, Cro11c, DC19, Dor18, Eid07, GW23, GNM21, GO81, MSS20, NY19, Num23, Pfe02, PBH19, QS87, Smi72, SH21, SST20, WOP20, RLS23, GP12]. Small-Sample [Cro07a, GO81, GNM21]. Small-scale [Bic14]. Smeeton [Oja08a]. Smirnov [HzB91, Sch84b]. Smirnov-type [Sch84b]. Smith [Gow79, Mai11b, Buc77c, DS20, Lin80c, O’B07b, Pol13b, Rob11d, Smi94b]. Smooth [RB90]. Smoothers [BP92]. Smoothing [GW23, Gre06, Han09b, MB19, Pod22a, RDM^{+}19, Sim98, SST20, dAY92, Tit85, O’B12i]. Smoothness [Gue08]. Snedecor [Kem74]. Snieder [Rus10]. Snooping [HP22]. Snyder [Han09b]. Soccer [Unw24]. Social [Bar75c, Bar83, Blu16, Des00, GT99, Gho08d, Gho09c, Gho10a, Han08b, Han10c, Keu97, Kok07, Liu22a, Mai11b, Mai13a, Nan10a, O’B14a, PZ21, Pol13b, Ros09a, Sch80a, SMW79, Spe10b, Sta09, Sta12d, Sta12c, Sta13c, Sto75, Sen23]. Société [Sau76]. Sociétés [Ano72p, Ano72o, Ano73l, Ano73m, Ano73n, Ano74l]. Societies [Ano72n, Ano72o, Ano72p, Ano73l, Ano73m, Ano73n, Ano74l]. Society [Dur13b, Fin75, GP03, Gho10b, NP97, Ott02, Sun18, Wyc03, Sau76]. Soft [Sti14]. softback [Cox16]. softcopy [Ana21]. softcover [Dur18, Hau18, Mer17, Pod17a, Ric17a, Sun17, Tso16]. Software [Bie97, SW94]. Sokal [Cor75a]. Solar [Hau18]. Solari [Fér78, Ros10b]. Solka [Alv11]. Solomon [Csi80]. Soltow [Pod72]. Solution [AK95, PS13, Sta11b]. Solutions [Kag72b, Num13b, Pod15a, Sun86, Til22, Sta13a, Pod16]. Solving [Bre80, Sta13b]. Som [Ben75, Moo75]. Soma [Mer09d]. Somalia [KMM^{+}17]. Some [And87, And77b, BNS94, Bar83, Bel02, BF91, BD76, CFL^{+}21, Cox86, Cox02, Cra81, DPCM21, DO73, DER03, DC83, Dra13g, FHTV93, Fed80, Fed81a, Fed81b, Fel73, FB73, FB96, Fri93, Gan72b, GS19, GS87, Goo86a, Goo86b, Gou94, Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, Har79, Hir93, Hoe83, Hus72, IKS^{+}81, IKS^{+}82, KG10, Kir82, KASG80, LH73, NSS73, New81, New84, New88, NW06, OA91, Pat88, Pos80e, Pur72, Rao06, RC15, RJ91, SRN17, Sin02c, Sin02d, Smi92, Spi02, TC15, Tia07, Til22, Tsa24, ZgZ89, ZW19, dVKW96, vBS73, CP80, EMR88, IK86, LN78, You86, GT12]. son [Hor72]. sondages [Ios80b]. Sonia [O’B10c]. Sons [Alb19, Els18, Els20, Hee16, KBH10, Lan18, Liu20b, Mer17, Mer20a, Num19, Ric18, Sun22, Tso20, Var19, Var20, You19]. Sophia [Eva07a]. Sophie [Dra12e]. Søren [Han12v, Han13e]. Sørensen [Sko12]. Sortis’ [Hal84a]. Souma [Han10a]. Source [HBK05, Tam82, dWvDS20]. Sources [Ber17, Dra13e, Kir82, SFVP^{+}24, Tho19]. Sowey [Mer17]. Sozial [Sch80b]. Sozial- [Sch80b]. Space [DQ97, DG17, GSW17, Han09b, HP17a, HP17b, OCRC17, Pan07, PJK17, Per14, Pod15b, SR92, ZM17]. Space-Time [Pod15b]. Spain [DMS06]. Spanish [BJO^{+}98]. Sparse [Bat07]. Sparsely [DC83]. Sparsely-Sampled [DC83]. Sparser [TB22]. Sparsity [O’B16c, Van16, Cox16]. Spatial [AC21, Arb93, BSS18, CM12, CMW17, DC83, DMMG22, EM19, GW23, Gho09a, Han08l, MM01, MV22, NSP21, O’B09d, OZW23, Per14, Pod15b, RYG^{+}21, Rip84, San21, WB97, Ish81, RR90, Nan10b]. spatiale [Mar07a]. Spatially [BPP17, PAF18]. Spatiaux [RR90]. Spatio [GDSN18]. Spatio-temporally [GDSN18]. Special [Ano05a, BR00, BP12, Dra12c, FL05, Had10a, Had10b, Riv02, Sen72, STNW93, STM95, MCF12]. Spécialisation [Mal97]. specialization [Mal97]. Specific [NKH91]. Specification [dBV82, BH77, BH80]. Specified [Gou94]. Spector [Bat08a]. Spectra [Mar07b]. Spectral [MW81, ZgZ89, Zur91]. Spectrometry [Gho17]. Speech [God97]. Speed [Hau18, ZG12, Dra13d]. Speidel [Kró80b]. Spencer [MM77]. sphere [KS22]. Spherical [Cha93]. Spiegelhalter [Gho08g]. Spirit [Goo78]. Splines [O’B12i, Lia20, Rév73, Pod22a]. Splitting [HM19]. Sports [Dra11b, Sta08c]. Sposito [Gow79]. spread [SLY23].

Spreadsheet [Sco10b]. Sprent [Oja08a, Tip80]. Springer [Cap19, Cox16, Dur17, Dur22, Gho17, Hau18, Kha17, Liu17, Mai17, Mai21, Mar21, McE17, Num20, Num23, Pod17a, Rod21, Ros19, Tso16, Zha16b]. SPSS [Mai12b, Sta12a, Sta12d, Veh07a, Voe09, Eva07b, Veh13a]. Spyrou [Pod21b]. Squandering [HHD97]. Square [HK77, PBH19]. Squared [Ord72, Pla83]. Squares [Ald98, GP73, Han07c, Han12g, Har74a, Har74b, Har75b, Har75c, Har75d, Har75a, Har76, Hum77, LNA75, Pun07a, Tin80a, FBGGM17]. Srivastava [Han10g, Nel72b]. Srivstava [Pod14a]. ST [Ano78k, Ric13b]. St. [Sen14a]. ST/ESA/STAT/91 [Ano78k]. Stability [BC20, O’B08d, Tin80b]. Stabilize [Ela73]. Stable [DTP^{+}20, La 16]. Stage [Ber72, Buo06, HTW01, MMC13]. Stan [Liu23, Rui17]. Standard [Le 90b, MPR12]. Standardise [SAA22]. Standardization [GSGdMLM16, Hoe87]. Standardized [MW76]. Standards [BJPT99, Car93, Han13i]. Stange [Mün76]. Stanley [Han11f, Veh10b]. Star [Han12t]. Stark [Spe10c]. Stars [BM91]. Start [Mai11e]. Starter [Qiw20]. STAT [Ano78k]. STATA [Ple15, JF14, Tho14, Eva07a, Dek16, Hil15]. State [DQ97, Han09b, Nie04, Ole03, Ott02, Pan07, SR92, CH13, Iac84, NW82]. Statement [Ano99e, AdV99, ST92]. States [Ano76m, FFBK76, Spe10e, Bel09, Lid80, O’B16a]. Stationaere [Kro80c]. Stationarity [Tsi03, BSS18]. Stationary [FS08, MN98, Slo24, SM89, Zur91, Nor13b]. Stationary-Increment [FS08]. Statist [And87, And77a, BT76, GS87, Goo78, Hoy76, Moo78, Ott76, dBPR86, Sch78]. Statistic [CR89, HzB91, BBS77, McC85, Meh75]. STATISTICA [Veh07a, Cop77]. Statistical [Ahl02, ANK03, Ait18, Ald05, AG99, Alv14, And77b, Ano72j, Ano72b, Ano72k, Ano72l, Ano72m, Ano72n, Ano72o, Ano72p, Ano72q, Ano72r, Ano73i, Ano73j, Ano73k, Ano73l, Ano73m, Ano73n, Ano73o, Ano73p, Ano74j, Ano74k, Ano74i, Ano74l, Ano74m, Ano74n, Ano74o, Ano74p, Ano75f, Ano75m, Ano75n, Ano75l, Ano75o, Ano75p, Ano75q, Ano76f, Ano76j, Ano76k, Ano76l, Ano76o, Ano76n, Ano77l, Ano77m, Ano78m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j, Ano86h, Ano21, Arb93, Ast94, BM16, Bal76a, Bal76b, BR00, Bar72b, Bar75d, Bar72c, Bat02, BHBP72, Ber03, Ber14, BW20, Bil80b, Bin00, Blu05, Bon96, BJPT99, Bra75, Bre99a, Cac74, Car93, Car05, Cha20b, Cha16, CN98, Cla00a, Col99, CS00, CM06, Cox03, CW04, CKK^{+}11b]. Statistical [Cra81, Cro09c, Cro11h, Cur03, Dav09, Daw02a, Der75a, Dil00, DSC03, DFV99a, DFV99b, DD80, Dur13b, Dur21, Dur22, DG17, Edi77, EIRS18, Esp14b, Eva07a, FW04, FHTV93, Fel96, Fer73, Fin79, Fin80c, Fis10a, FL05, Fom88, FOK73, FPR^{+}16, Fri03, FGC03, Gal02a, Gal02b, Gan79, GMZ03, Gho09i, Gho10b, Gho12c, Gho13b, Gho13a, Gho17, GSW17, Gow15, GPW05, Hab82, Had10a, Had10b, HN13b, Ham72a, Ham77a, Ham79, HL19, Han07g, Han08a, Han08n, Han08h, Han10c, Han10a, Han12q, Han12e, Han12x, Han12p, Han12c, Han13b, Hau01, Hau08a, Haw97a, HW06, Hel04, Hoe87, HRR20, HP17a, HP17b, HWS20, Hor73, Høy72, Ios76c, Ios79b, Jon00, Kaf15, KK21, KSS^{+}03b, KSS^{+}03a, Ken73a, Ken74, Kha73, Kha24, Kir05, Koe79, KC14b, Kos10a, KSRS72, KM79c, Kru89, KMT03, KMS14]. Statistical [Lai22, Lan83, Lan85, Law98, LEGW01, LW13, Lin14a, Lis07c, Lis08b, Liu11, Lon15, LC20, Loy80, Lyb00, Mac16, Mai15, MP23, MM96, Mar74, Mit88, MS23, Mul76, MW76, Nai73, NC21, Nay09, NA16, Nel72a, NPT^{+}74, Nel74, Nel75, NP97, NL07, Nor09a, Nor09c, Nor14a, Nor96, NGC^{+}78, O’K12, OA91, Oja08a, Oja11, Ole03, Oli14, OCRC17, Osb91, OZW23, Par73, PJK17, Pat75a, Pat79, PZ21, Pau19, PA15, Pet07, Pin10a, Pin10b, Pod22b, Pol11b, Pol11a, Pol13a, Pos76a, Pos80e, Pre79a, Pun08c, Rao93, Rao06, RC15, Rei12, Rob11e, Rob11c, Rob12a, Ros15a, RvA72, Roy14, RBB^{+}21, Rum02, Ryd00, SDGD^{+}11b, Sär72c, Sav80, SLWH22, ST07, Sch07, Sch03, SJFM95, Sen72, Sen03, SZ09, Shl07, ST92, STM95, Sil73, SW15, Ski12a, Ski12b, Sko12]. Statistical [Smi75, Smi99, Smi01, Sne99a, Som78b, Sta12b, Sun18, Sun19, Sun96, Swe95, Szc11b, TC99, Tip77, Tre00b, Tre05b, Utt21, VJ95, Wal96, Wat02, Wei05, Wel02, WS79, Wet75, WP99a, WP99b, WPRP17, Wil12, Yas93, Zar76, ZM17, Zha16c, dV99, vBS73, vdHvdH02, BH80, BNCR86, BT75, BT76, Bic13, CH13, Goo78, Hal86, Kot74, Lan72, Lan73, Lan74a, Lan74b, Lan75, Lan76a, Lan78, Lan79, Lan81, Lan82, Lan84, Lan86, Lan87, LOD^{+}85, Moo78, Mor74, Sax78, Tip77, Tip80, Bar77b, Ber78, Bin07, Blu07, Boo79, Buc79, Buc77c, Coc73, Cro08a, Cro10j, Cro11f, Dal72, Dra08b, Dra10c, Dra14c, Dur23b, Eng78, Esp15a, Esp15b, Esp15c, Fer80b, Fil09b, Fin77c, Fin78, Gan80d, Gho12e, Gho17, Gho20, Gil15, Han07f, Han08m, Han08l, Han09d, Han10f]. Statistical [Han11k, Han12n, Han12m, Han12l, Han13o, Han13n, Han13m, He14, Kos09b, Kos12c, Lan18, Lis10c, Liu19, Liu21, Mai12c, Mai13c, Man72, Nad74, Nel80, Nor15, O’B08d, O’B09h, O’B09g, O’B13g, O’B14c, O’B16c, Pal72, Pla75, Pos80b, Pos80c, Pun08d, Rao80b, Ric09, Ros14, Rui17, Sco13c, Som80b, Spe10b, Spe10c, Sta08c, Sta08b, Sta12e, Sta12f, Str76, Tip75b, Unw10b, Var19, Veh08g, Vin73, Vos75]. Statistically [Bie99, CC08]. Statistician [All72, Dra13a, HV15, Hea95, Kei05a, LN80, PR10, Rao19, Spe08b, Dag76, Sau76]. Statisticians [AM14, FMS95, FPD05, Gol97, Hun14, Ios79b, KSS^{+}03a, Ken73b, LCZ14a, Mac97, MDG^{+}92, Pre78b, Ric10c, Rob14, Sen05, Sne99b, Mal97, Tip80, Han11a, Lin13, Lis08a, Rob11f]. Statisticien [Dag76, LN80, Sau76]. Statisticiens [Mal97]. Statistico [Tin78b]. Statistics [Abr05, Alb19, AG99, And78, And73c, And07, Ano73d, Ayh01, Azo01, Bar83, Bar76, BK88, BS03, Bat08b, BL24, Bie97, Blu16, Blu20, BO15, Bor09, Box75, Buc74, Buc76a, Buc76b, Buc77a, Buc77b, Buc78, Cam97, Cam07, Car93, CG05, Car00, Cha20a, CBPAK22, Cha93, Cle01, Cle05, Cor14, Cox86, Cox97a, Cox97b, Cro10d, Deb78, Des00, DFV99a, DFV99b, ESTTD05, Elf72, Eva07b, Fel04, Fin77b, Fis18, Fis01, FPD05, FB73, FG72, Fri01, GO10, GO16, Gan72b, Gan80a, Gar95, Gar97, GG99, GBZ07, Gho97, GMRS99, Gho08d, Gol97, GG15, Gou10, Gow79, Gri15, Gus06, Gut03, GL09, GL19, Had10a, HHDG12, Hal81, Han10i, Han10c, Han12u, Han12w, Han13k, HKZ07, Haw97b, HW06, Hen78, HMA^{+}18, Hey81, HBK05, Hof97, Hor16, Hor72, Hor91, Hot16]. Statistics [Hua89, Hug06, Huz80, Ios77d, Ios77c, JS03, JM89, Kei05b, Ken75, Ken76c, Ket97b, Ket12, Khu13b, Kin75c, Kit03, KM79b, LvR74, LA96, LSS80b, Lin80b, Lis09a, Liu15, Liu20b, Mac16, Mag09, Mai07, Mal05, Mar21, MA05, Men16b, Mer17, Moo97a, Moo97b, Moo05, Mur94, NSS73, Nan10b, Nar74, Nic01, Nor09b, O’B09i, O’B13c, OG16, Oja08b, Oja09, Ole03, Ott02, Ott05, Par80, PNR13, Pod11, Pod16, Pos79d, Pow03, Pre77a, Pre78b, PP98, Pun10d, Pun12b, Rad99, Ram03, Ric10c, Ric10a, Ric13c, Ric13a, Ric15, Rid16, Rip84, Riv02, Rob11g, Roh76b, Ros09a, Ros19, RR79c, Rus74, SHLW22, Sch97, Sco10b, Sco12b, STNW93, SMW79, STY90, Som78e, Som78c, Som79a, Som79c, Spe11b, Spr75a, SS09b, Sta07e, Sta07c, Sta12a, Sta12e]. Statistics [Sta12c, Sta13b, Sta13c, Sti00, SHL05, Sto98, Sun19, TC15, Tam07, Tan92, Teu97, Tha07b, TC99, Tin78c, Tip77, TTV05, Tso20, Unw11b, Unw13a, Veh07a, Veh07b, Veh10c, Vov01, Wat74, Wat98, WRH05, Wil95, Wil16a, You19, You20, Zar90, dWvDS20, AM93b, Bak73, BNC84, BG83, Fis83, Gir15, HN12, Har79, KM80, LSS80a, MCF12, Tit85, You86, dF74, Cha80, Deb78, FMS95, Had10b, Ios77d, LN80, Pow03, Pul97, Roh76b, SHGK01, Som77b, TC15, Amu73, Ban14, Cha80, Cor80, Cor77b, Cor77a, Cro07a, Cro10e, Dem73, Der75b, Der77b, Dow72, Dra11e, Eag74, Fin77a, Fin80a, Fin80b, Gho08e, Gho11a, GP12, Han12k, Has08d, Ios76a, Ken76b, Kos12b, Kos13b, Kró80a, Lis10b, Mar73, Mer14, Mer17, O’B13d, Pos76e, Pos78a, Pos80a, Pre77b]. Statistics [Pun07b, Pun08b, Pun10a, Ric14, Sar73, Sco77a, Sen11c, Som76c, Som76a, Som79e, Som79b, Som79d, Som80a, Som80c, Sta10a, Stu72, Tra73, Unw13b, Veh08h, dO79b, vdL80c, Bre80, Buc77d, Cor77c, Dra11d, Dra14d, DD80, Emb08, Fin75, Fin77d, Gan76, Gho12f, Gho12h, Gho12g, Kró80d, Lin13, Lin80c, Mai10c, O’B13h, Rév80, Ric11c, Ric17b, Ros09c, Som77b, Som80c, Spr77b, Sta07f, Sta10c, Tip80, Var14, Veh13a, Wil73, You15]. Statistics-The [GL19]. Statistik [EV73, Kri74, Dei76]. Statistique [Ano72o, Ano72p, Ano73l, Ano74l, Ano75l, BHBP72, Bra75, Hor72, Hor79, LN80, Nai73, Pre79b, Rao80a, Lin80a, Mar07a, Ano72a]. Statistique. [Esp11]. Statistiques [Ano72l, Ano72m, Ano73i, Ano73j, Ano73k, Ano73m, Ano73n, Ano74j, Ano74k, Ano74i, Ano74m, Ano74n, Ano75m, Ano75n, Ano75o, Ano76j, Ano76k, Ano76l, AM93b, FOK73, Hor73, Mar75, Nel72a, NPT^{+}74, Nel75, Zar76, Dru11]. Statistische [Dei76, Oni80, Oni79]. Stats [Han13g]. Status [Car00]. Staudte [Han08h]. Stayton [Mai09b]. Steel [Tra13]. Steele [Han08b, Mai17]. Stefan [Cro09d, Cro10b, Han12q]. Stefano [Liu08, Luo09, Ric10b]. Steffen [Gho08g, Han13e]. Stegmüller [Oni79, Oni80]. Stein [Sta07e]. Steinijans [Mat07]. Sten [Wol77]. Step [Ber17, Blu20, Cor14, Liu22a, OH13, Ric10c, Ric15, Ros15b]. Step-by-Step [Blu20, Cor14, Liu22a, OH13, Ric10c, Ric15, Ros15b]. Stephan [Han08h, Hau72]. Stephanie [Ano21, Gho10a]. Stephen [Cro11a, Gho11b, Häg12a, Han10d, Som80a, Spe11b, Stu80]. Steps [Cox10].

Stepwise [SC05]. Stereology [JBGS85, Sto90]. Steve [Cro08e, Liu14a, Pol12a]. Steven [Cro10c, Mer12a, Ros10c, Var20]. Stevens [Ros09a]. Stewardson [Dra08b]. Stewart [Unw12b]. Steyerberg [Nee09]. Stichproben [vdL80b]. Stöber [Dur13c]. Stochastic [AdB04, AET04, AP94, AES18, Bad82, BNS94, Ber16, CIS06, Coh75, Dur13c, FD02, Gho12b, Han10j, Han10e, Joh77, Joh80, Kin75a, KS89b, LH04, Luo09, Ma17, MFAG10, Nor13b, Pod17b, PS13, Ros10b, STI06, SS08, Sør93, Sto90, Tso18, TP73, TLL04, CH13, Gob16, GT12, Sør86, Sri80, Ban14, Cor75b, Cro09e, Cro11g, Han12o, Ios76e, Kin77a, Ng21, O’B11c, Pin75, Pod17c, Pos76c, Sch72b, Sto75]. Stochastisches [Dor73]. Stockholm [Sch80a]. Stockwell [Pet07]. Stoffer [Pol13d]. Stoltz [Han13g]. Stop [HHD97]. stopped [Sør83]. Storia [Tin78b]. Story [Kru89]. Stoudt [McC22]. Stoyan [Han08l, Han08l]. Strata [SS17]. Strategic [HHD97]. Strategies [ACV84b, ACV84a, Ano76m, Ano79e, Ano21, CHPR20, EKIKW77, FFBK76, Kha24, KFF75, MP21, YAA20, Iac84, SL13, Fis18]. Stratification [BR09, HK18, Hor06, Jag86, SHH94, ZL22, JOT85]. Stratified [Bre99b, KL09, NPT08]. Streams [Han12h]. Streibig [O’B09f]. Strength [Pos73, Bic13]. Stringer [Ano93d, Bic92]. Strode [Sti88]. Strong [BF91]. Stroup [Zha19b, Lis13, Spr77b]. Struchiner [Aur10]. Structural [ABC11, Ano76f, Fon13, FB96, Han11f, ME89, OH13, Ros15b, Sun16c, TC07, Tin78d, Veh11a, BT75, BT76, Veh14]. Structure [Cro08b, CS88, DMMG22, Fra72, Han07b, Hoe74, Ken76a, Liu11, Ma17, Tit85, ZT13, Val08]. Structures [Ano88a, Cas20, Lis10c, FT87, FT88, LL87, PWKO24, Ber78]. Strukturierung [Lin75a]. Stuart [Coc73, O’B09c, And78, Bar76, Som77a]. Student [Bre79c, Gou10, Kau13, Sta11b]. Students [Gar95, GBZ07, Rus10, Gri77, Tha07b]. Studies [BO15, Cro07a, Cro10f, Dal98, DR93, Doc22, DTP^{+}20, Esp15b, FM95, GNM21, JSW11, Mat07, Meh15, RM93, Smi73, SS88, Som78a, Ver13, WRH05, KC14a, Lai22, Lon15, Aur10, Cor77c, Don12, Fis10a, Gho09f, Han12d, Lis10a, Mai18, Mai22, O’B08d, O’B09a, O’B12e, Sun24, Ios77c, Sat15]. Study [Bar75b, BDS04, Bic00, Dra12d, FBGGM17, IMG22, NWE16, O’B07c, OMGT05, Pat75a, PPMS11, Rui72, SVCV88, Veh08f, Wan19, Wil95, Lec76, NSZH22, Dra10b, Gal80b]. Studying [O’B14a, Has08d]. Stulz [Han07h]. Stumpf [Gho13b]. Stunting [KMM^{+}17]. Stupid [HY01]. Stute [Roh80]. Styan [Liu14b, Pol13a, Ken73a, Liu07]. Sub [HL01, HRR07]. Sub-annual [HL01]. Sub-Saharan [HRR07]. Subclass [NAL76]. Subgroup [Zha21]. Subgroups [DTP^{+}20, Rad10]. Subject [HP22, Mur74, Har76]. Subjective [Ham77b, dF79]. Submarines [Rob12b]. Submission [Had10c]. Submodel [BS92]. Subsampling [MPR12]. Subsigma [BS89]. Success [Ios76f]. Sudarshan [Han07c]. sufficiency [Rém84]. Sufficient [HMA^{+}18, Tin78c]. Sugasawa [Num23]. Suite [Lef79]. Sujit [Pun10c]. Sukhatme [Kem80]. Sullivan [Tha07a]. Sum [Gow79, NW82]. summary [OC13]. Summer [Han13g]. Summers [Kha80]. sums [HS12]. Sun [He14, Pol11a]. Sung [Par05]. Sungur [Dur21]. Sunny [Far13]. Suojin [Tra13]. superpopulation [GT86]. Supervised [HCV^{+}21]. supplement [Kot74]. Support [Car93, O’B13h, Var14]. Surfaces [Cox07]. Surgailis [Kos12b]. Surname [FL83]. Surnames [Con91, Pan89]. Surprising [BKBP21, RJ91]. Surrogates [Gho21]. Surveillance [Fri03, O’B09g, Han08f]. Survey [AN01, And77a, AL10, AM14, BCM19, Bal95, Bin00, BCD^{+}94, Can95, CW06, CWCT19, Chi22, CD94, Cla00a, CS00, Cro10c, Dil00, ES06, Fel73, FWC^{+}21a, FWC^{+}21b, FMR99, Han13c, Han13i, Hee16, HK77, HBK05, Hun14, Huz80, Kag72b, Kal83, Kal19, KP10, KT21, Kir82, Kis94, LCZ14a, Lit86, LSD^{+}11, Lyb00, Mak76, Mer24, MSS01, Mur74, Nor72, Nor73, Ovu80, PGN^{+}96, Pfe93, Pos79b, Pos80d, PHO86, RAC83, Rao05, RG23, Rob14, Ruk95, Sär72c, SBB^{+}12, SS17, Sch95, Sin95, Sma90, Som76b, Som80b, Sør04, SD11, Sun22, Til22, Tre00b, Uni78, Veh08c, Veh08f, VBL13, WWY23, WD84, ZHW23, GT86, Jac84, JOT85, Joh15, KS85, NW82, dGAGR85, Ano76m, FFBK76, Ins73]. Surveys [Ale94, Ano76m, Arb93, Ayh01, BP12, BE17, BK20, Bet10, CB00, CG12, CH19, CP19, CKS17, Coh05, Dem77, DK87, FFBK76, GDSN18, GT17, HB07, HS88, HRR07, HL01, Hof97, HS89, KBM21, MDS^{+}19, PS85, PHO86, Red74, Roy94, dMRAMR18, Ski19, SS88, Smi94a, Smi94b, Sun77, Tam94, Tho19, TH98, Til98, VL96, WL16, ZW19, dWC05, vdBZT20, Bin83, Dal77a, Dal77b, Dal77c, KFF75, Kot74, Mon87, SSJ77, Tam87, VLM12, ZTK13, Lin14b, Tra13]. Survival [ABGK82a, ABGK82b, Ano84a, Aus17, Bre75, DG90, Gho09e, Gil15, Hjo92, JK78, LW13, MBW^{+}23, Oak81a, Oak81b, SAG^{+}16, MMC13, Cro08e, Gho09h, San17]. Susan [Kos13d]. Susceptibility [Cla00b]. Susmita [Gho17]. Süssmilch [Hor91]. Sustainable [Rad99, Hau08b]. Suzanne [Pol13b]. Suzuki [Ros10a]. SV [PWKO24]. SV-GARCH [PWKO24]. Svein [Alb19]. Svolba [Sta07b]. Swan [Veh13b]. Swarna [Mai17]. Swayne [Mai08]. Swedish [GL19, LQ04]. Swenson [O’B14b]. Swersey [Luo07]. Switching [Han07d, PWKO24]. Syed [Kha24]. Sylvia [Han07d, Ric05]. Symmetric [ADE^{+}22, BCF^{+}14, Sch16b, Chm81, Ken75]. Symmetry [Dar81, Hel04, Lov00]. Symposium [Deb78, Ken75, Pos78b, RvA72, Sch80a, Wat74]. Synergy [Doc22]. Synthetic [KRR^{+}11, Han12p]. System [AH17, Bon96, CNS85, CT11, Fel96, HW06, IOS12, Keu97, Nor96, Ott75a, SCZZ99, Wal96, ZG12, NGC^{+}78, EV73]. Systematic [FBL18, Ham78, Iac82, Nar74, Ken74]. Systèmes [Mal97]. Systems [Ano79e, Cro09e, EKIKW77, Gan80a, Gho08g, Gho13b, Han08k, Kin75a, Knü80, MW76, Mur74, Nel74, NWS23, SS08, SHGK01, Sne99b, TH98, WB97, dV99, Gri87, Mal97, OC13, Bat07, Dur14a, Fin80d, Gho12b, Han10e, Han11i, Ios72]. Szabados [Slo24].

T [Bat08b, Bre79c, Buc77c, Cro10c, Fin75, Fin80b, Gan80a, Gan80b, Gho10a, Gho11b, Goo73, Han07b, Han12m, Has08d, Hoe83, Ios76d, Ken73a, Ken10, Kin77b, Liu18, Mer09a, Pre79a, Ros09b, Som77b, Sta09, Sta13a, Tip77, Unw20]. T. [Bai84, Hal81, Lau81, Smi94b, Spe84, Wyn84]. Tabard [Pre79b].

Table [Kag72a, KZ76]. Tableaux [HL74, Pre79b]. Tables [Buc76a, Buc76b, Buc77b, Buc78, GP73, Goo86b, Ken73b, Lov00, Mor80a, Shl07, TSR96, Vos75, And87, GS87, Goo86a, HL74, LHK78, Csi80, Fin80c, Pat75b]. Tabular [Smi73]. Tabulated [Bar80a]. Tabus [Csi08]. Tachi [Ued74]. Taha [Han12d]. Tail [Dan87]. Taillie [O’B12e]. Tails [Jon07, Tsi03]. Tait [Fis19]. Takane [Pun11d]. Takeuchi [Pun11d, Kag72a]. Takezawa [Muk14b]. Taking [Red01]. Takuya [Veh13a]. Taleb [Veh13b]. Tales [Dra09d]. Tam [Dur17]. Tamás [Slo24]. Tamhane [Dra10c]. Tan [Ken10]. Tanaka [Els20]. Tang [Num14, Lin14b]. Tango [Cro11f]. Tanton [Ric13b]. Tanur [Spe11b]. Tao [Kau12]. Tarantola [Liu08]. Tarmo [Liu07]. Tasks [Lav14, O’B15b]. Tasting [HG21]. Tatsuoka [Gir73]. Tatsuya [Num23]. Tautu [Cor75b]. Taxonomy [DGS18, GK01]. Taylor [Fis24, wB09]. Tea [HG21]. Teach [Moo99, Sne02]. Teachers [Wat98]. Teaching [Ano73d, DPS07, FB73, GBZ07, Ham78, Nic01, NL07, Rum02, Wil95, Bak73, Ric13c]. Technical [Cle01, Mal97]. Technique [BCM19]. Techniques [BEW03, Boo79, Fri93, Han07a, Han08c, Has08c, Hir93, Høy72, HCV^{+}21, KW19, Lin14b, Mal97, Mas73, O’B09b, Pun08c, dMRAMR18, Sma88, Szc11a, Unw09, WD84, ZHW23, SB75, Tit85, Has08a, Moo75, Pos80e, Som80b, Veh09b, Pre79b]. Techno [Sch07]. Techno-Trash [Sch07]. Technology [EPP08, GP03, Pow03, Som80c, Wat98]. Tectonic [Cha93]. Teletypewriter [And73d]. Tell [Sun18]. Template [OZW23]. Templeton [Ios76d]. Templin [Han11d]. Temporal [Abr82, Gue90, Pro11, SAA22, Han11l]. Temporally [PAF18, GDSN18]. Ten [O’B10d, SH00]. Tenko [Han11e, Veh09a]. Tensile [Pos73]. Tension [Mer07]. tenth [Lan76a]. Tercentenary [Syl14]. Terence [Han12r]. Terje [Pos76d]. Term [OA91, RM93, ZG12, SYSK02]. Terminals [And73d]. Terminology [Ken74]. Terry [Dra13d, Liu08]. Tertiary [Nic01]. Test [AS72, BRST76, CQOR^{+}22, Dar81, Feu93, HzB91, KS22, OH19, Pla83, SM19, Tan92, Væt85, Veh08g, JB87, WLC82]. Testing [BS89, BM91, BR02, BFST18, Cox16, Cro08c, DMMG22, Gel03, Hab82, HM90, HH93, HG94, Han08m, HTW01, KS89a, KX09, Law07, LN78, ME93, NW06, Pér93, SKK18, Sow72, Sow78, Spe11a, Str76, TWRB20, TC72, Zha18, HMT22, Sha88, Van16, Cau80, Han12q, Luo07]. Tests [ABGK82a, ABGK82b, Ano84a, Bar72b, BR07, BKBP21, CPB94, Dar81, DC83, GMZ03, GHK20, HG21, Hoo05, LLP01, LS06, LV14, MR80, MM04, Mod08, OS73, RB90, Ros10b, SC05, SKK18, TSR96, WWY23, vdL80b, Han10b, LHK78, Sch84b, Man07]. Teugels [Fis10b, Teu05a]. Texas [Han13f]. Text [Pos78a, Els20, Bre80, Ric11a, Han10g, Rod21]. Textbook [Kró80a]. Texts [Fom88]. Teyssière [Kos12b]. Th [Veh08g]. Theil [Dra11c]. Their [DPS07, Ios78, Man72, GT86, GP11a, Han11h, Rus10]. Them [Liu14b]. Theme [Gho10b]. Themes [Lan01]. Theodore [Coo07, Dav09]. Theological [Sta10b]. Theophano [Han11l]. Theorem [BJS92, Hal84b, vdV91, Sen14a]. Theorems [AP94, Rao06, Liu10a, Mer18]. Theoretical [ZW19, Han12u, Pos76e]. Theoretische [Lin75a, Mui73]. théorie [Esp11]. Theories [Lan01, Kin77a]. Theory [And78, Bal95, BNS94, Bar76, BPV21, Bha07, Bra15, Cac74, CWCT19, Cor77b, Cox80, Cox86, Cro09c, DM73, Dek14, Dem73, Dra10c, Dur22, Esp15c, Fin77c, Fis18, For82, Gan80c, GT99, Gho12g, GG15, GP11b, Haj77, Han08d, Han08k, Han08m, Han09d, Han09c, Han11d, Han12p, Har78, Ios76f, Ios76c, Ios77d, Ios79b, JM89, Kin77b, Kos12a, Lee01, LSS80b, LR91, Liu14a, Liu14b, Loy77, Mai11d, Mor80a, NSS23, Nor13b, O’B08a, O’B11a, O’B15a, Oja09, Oli14, Pod14b, Pos78e, Pos80c, Pun07b, Pun10b, Pun11b, Pun11c, Pyn14a, Rao80c, Rob12b, RBR85, Roh76a, Ros10b, Sär72c, Sen96, Sil73, Smi80, Som76b, Spe10c, Str76, TC07, Til22, Val08, Var14, Veh08g, Vin73, Zha03, dO79b, BNCR86, Cha08, JT13, Kal84, LSS80a, Mor74, Puk87, Sha88]. Theory [And78, Bil80b, Der77c, Der77a, Fin78, Han10j, Han11e, Ios78, Mer09a, Num13b, Pun12a, Tso17, Han10j]. there [Hed20]. Therese [Blu20]. Thiele [Hoe83, Hal81, Lau81]. Things [Dra13e, TRRL23]. Think [Bie99, CKK^{+}11b]. Thinking [AG99, Bre99a, Cob07, DFV99a, DFV99b, Mac16, Mai13c, ST07, Smi99, Sne99a, Sta08c, TC99, WP99a, WP99b, Dek22]. Thinning [Gal80b]. Thionet [Ios80b]. Third [Cro11c, Fon13, Gho09f, Gho12h, Han12c, Han13h, Lis09a, Pol13d, Pun12c, Pun13b, Ric11c, Ric11a, Sen11c, Szc11b, Unw11a, Veh11a, Gro95]. thirteenth [Lan81]. Thomas [Dra09c, Fin80d, Gho08a, Han10c, Han11h, Has08a, Kos09a, Kos10b, Lis10c, Oja11, Roh76a, Spe10e, Sta13b, Tan07, Tin80a, Cro10j, Pos76a, Sti88, Veh09b]. Thompson [Gho13c, Han07g, Han12f]. Thorn [Häg12b, Sko12]. Threat [Goo78, Ham77a, Moo78]. Three [CSW93, HH07, LKC^{+}04, Sär11, Sim98, TGM09, LHK78]. Three-Level [CSW93]. three-way [LHK78]. Thyregod [Liu12a]. Tian [Ken10, Lin14b]. Tiao [Smi75]. Tibshirani [Nor09c, Nor14a, Spe10d]. Tie [Zha18]. Tie-Hua [Zha18]. Tillé [O’B20]. Tim [Sco12b]. Time [Abr82, And77a, Ano83a, Ano21, Bha07, CD94, CHPR20, Cro11b, DCC98, DS22, DK87, EFP19, FM95, FST19, Gan78b, Gue90, Han12r, Han12x, HLC97, HHK03, Hau18, HR83b, HR83a, HzB91, Jas15, KP04, Kos10b, Kuk22, Lau81, Law91, Liu17, Loy77, Mai10a, Mak76, Mak78, McL93, Neu16, New81, New84, New88, Ord80a, Per14, PF82, Pod15b, Rao80b, RC93, Sai08, SP11, Sco09a, Sen14d, SR92, Slo24, Sør04, SM89, TRRL23, Tsa24, Zur91, vdBZT20, KDTP20, LL87, Mor84, PMM87, SSJ77, Bar77a, Gho13d, Han12v, Mai09c, McE17, Pol13d, Qiw20, Rob11h, Tin78e]. Time-Series [Mak78]. Time-to-Event [FST19, Liu17]. time-varying [KDTP20]. Times [JK78, Oak81a, Oak81b, Zhu14]. Timothy [Bat07, Cro10g, Dur23b, Han11a, Hee16, Pol13b, Ric11c]. Ting [Zha17b]. Tinsson [Dru11]. Tiny [Sen14c]. Tippett [Dur23b]. Title [Bar72c]. Titterington [Gho10b]. Tjur [Bai84, Spe84, Wyn84]. Tobias [Puk07]. Tohru [Gho13d]. Toine [Ros09c]. Tolerance [Pal72]. Tomasz [Kos12a, Mai14c, Unw21]. Tome [Hor79]. Tomohiro [Rob11e]. Tomoyasu [Sai79]. Ton [Ros09c, Sta12e]. Tony [Dra08b]. Tool [And73d, Num13c, OZW23, Ott02, Pat88]. Tools [FGDS18, Han13o, Kos12c, Lav14, SRN17, Unw09, O’B15b]. Top [O’B10d, Pol13a]. Topics [Che24, Han12u, Han13m, Ios76b, Pos79a, Pos80c, Pyn14a, Gan80b, Han13k]. Topological [Ell91]. Topology [Bre79a]. Torben [Kos10b]. Tore [Sch80a]. Torrens [Rio76, Dag75]. Torrens-Ibern [Dag75]. Torsney [Fil11]. Torsten [Ric11b, Rob11c, Unw12a]. Toshiro [Cro11f]. Total [Ano79e, EKIKW77, SS07]. Totals [GT17]. Tour [Pod14b, SW15]. Tournament [Dav72]. Toutenburg [Pun07a]. Tower [Mer09c]. Toxic [Rao93]. Toxicological [Pre89]. Toxicology [Ric17b, Hot16]. Traceable [Wer12]. Tracking [SDGD^{+}11b]. Tracy [Ken75]. Trade [Par80]. Trading [DQS92]. Trading-Day [DQS92]. Traditional [Tho19, WD84]. Traffic [LH73, CP80]. Traffickers [Ize03]. Trailblazers [BW20]. Train [Dra12f]. Training [ANK03, Ano74m, Ano74n, Ano75o, Ano76k, Ano76l, BHBP72, Bra75, CN98, FOK73, Hen78, Hor73, Kir05, KSRS72, Nai73, Nel72a, NPT^{+}74, Nel75, NGC^{+}78, Zar76]. Trait [SGLS02]. Traits [Cro08a]. Tramontano [Man08]. Transactions [Gho10b, Ios79b]. Transform [SM93]. Transformation [Ela73, Rob93]. Transformations [BS85, Hoy73, Hoy76]. Transforms [BMM08]. Transition [Gho10a, MA05]. Translated [Els20]. Translation [SVCV88]. Transmission [BD94, FMS12b]. Transport [Sør93]. Trash [Sch07]. Travail [Mal97]. Treasury [LKC^{+}04]. Treatise [Rob11a]. Treatment [Ham79, LCL^{+}23, Roh16, FMS12a, Sco12a]. Treatments [Arm85a, BCL^{+}85, Bat85b, Che85b, Lou85, Pet85, Whi85, Bat85a]. Trees [Ahn14, Cia14, Loh14a, Loh14b, RZ14, SZ14, Str14]. Trend [Han12n, LQ04]. Trends [Der75a, Gue08, JBGS85, SYSK02, Som78d, Zha16a]. Trenkler [Fil09b]. Trent [Dur18]. Trevor [Nor09c, Nor14a, Sun19]. Trial [Arm85a, BCL^{+}85, Bat85b, CBPAK22, Che85b, Koe79, Lou85, Num13a, Pet85, Whi85]. Trialists [ET16, Zha17b]. Trials [ACV84b, ACV84a, Arm85b, Bra14, EUC09, God80, Hir93, Ios76f, IMG22, MM96, Muk14a, Sch16a, Til22, Tuk91, Zha19a, Zha16c, Zha18, Bat85a, Sco12a, Cro09a, Cro10f, Han12e, Kim14, O’B07a, O’B12a, Ros09c, Ros10c]. Tribute [RBB^{+}21, Sch84a]. Tricks [Pol13a]. Trigonometry [Gus06]. Trinomial [Ham75]. Trio [Hoe83]. Tristan [Mai09b]. Triumphant [Rob12b]. Trivariate [CW91]. Tropospheric [SYSK02]. True [Fou80a, Bar87]. Truong [Liu22b]. Truong-Huy [Liu22b]. Trust [SFVP^{+}24]. Truth [Par73]. Truthful [CP19]. Truthful-Untruthful [CP19]. Trutzel [Tru05]. Tsay [Lep14]. Tsokos [Pos78e]. Tsung [Dur17]. Tsung-Hau [Dur17]. Tsybakov [Sen11a]. Tu [Mai13c, Num14]. Tucker [Qiw20]. Tuning [HT92]. Turbulence [RvA72, Gan80d]. Turner [Han08i]. Turning [Han08c]. Tutorial [Aus17, CMW17, EH92, HMA^{+}18, Rip83, WS01]. Tutuncu [Feh07]. Tutz [Liu17, Lis10c, Unw12d]. twelfth [Lan79]. Twelve [OS73]. Twenty [Dra13b, Pol13a]. Twenty-First [Dra13b]. Two [AJS81, Ber17, Ber72, Buo06, CSW93, Dem77, Fin80d, Fis01, HTW01, Koe79, KBM21, LKC^{+}04, Lee01, NSS23, NAL76, PZ21, Pat75b, RDM^{+}19, Rob12b, Sch84b, SKK18, Sør93, SM89, TSR96, CKS17, SS87, Sha80]. Two-Associate [Pat75b]. Two-dimensional [RDM^{+}19]. Two-Level [CSW93, KBM21]. Two-Locus [Fin80d]. Two-Part [Lee01]. Two-Phase [Dem77, SS87, Sha80]. Two-sample [Sch84b]. Two-Stage [Ber72, Buo06]. Two-Step [Ber17]. Two-Way [AJS81, NAL76, SKK18, SM89, TSR96]. Type [HzB91, HS90, Ma17, FMR19, Sch84b, Sør83, ZMSL19]. Types [HH07, Kis94]. Tzanetakis [Kau12]. Tze [Cro09c, Kos09b].

U [Coh75, Knü80, Tip75a]. U-Boat [Tip75a]. U.S. [Har99, KT93]. Udny [Coc73]. Uffe [Gho08b]. Ugarte [Bat08b]. Uggè [Col72]. UK [MDS^{+}19, Red74]. Ulmann [Som79c]. Ulyanov [Pun10d]. Unavoidable [LLP01]. Unbalanced [BKBP21]. Uncertainties [Han11h, Veh10a]. Uncertainty [BD94, LG04, WS21, Lis12]. Uncovering [Han07b]. Under-Development [Kir05]. Undercount [Red01]. Understanding [Bat02, BE99, Bor09, GO10, Han08o, Han12w]. Unequal [Ham72b, Tal93, HB80, Sun86]. Unified [Cox10, DCC98, DGS18, Han13h, JM89, Liu21, RN92, Sha88]. Uniform [BF91]. Uniformisation [Sha23]. Uniformity [Rad10]. Unifying [Cle05, GM91, HGKJ20, KS89b, SCC14, dF74]. Uniqueness [Kay03]. Unit [CS02, Liu10a]. Unit-Root [Liu10a]. United [Lid80, Spe10e, Ano72q, Ano72r, Ano73o, Ano73p, Ano74o, Ano74p, Ano75p, Ano75q, Ano76m, Ano76o, Ano76n, Ano77l, Ano77m, Ano78n, Ano78o, Ano79l, Ano80i, Ano80j, Bel09, FFBK76, Har73]. Univariate [BR09, GG15, Mit88]. University [Blu18, Blu20, Blu21, Cox19, Dek22, Dur18, Dur23b, Gho20, Liu19, Liu22a, Liu22b, McC22, Mer18, Pod18, Pod22a, Pun11a, Rob19, Sun19, Unw20, You20, NGC^{+}78]. Unlucky [HP22]. Unobservables [LK16]. Unpublished [AK95]. Unraveling [CD00]. Unreplicated [SKK18]. Unrestricted [KRR^{+}11]. Unstructured [Pou02]. Untruthful [CP19]. Unweighted [Hum77, LNA75]. Unwin [Veh08e]. Update [Mak78]. Updated [EFP19]. Updating [Nee09]. upon [dM84]. Upton [Bar80a]. Urban [Deb78]. Urbanization [She06, KT93]. Urdan [Ric11c]. urn [Hol86, Ios78]. Usage [Han07b]. usages [JOT85]. Use [Arb93, Bar80b, BRST76, Chu74, DPS07, Ela73, FMR^{+}20, Fri93, GMZ03, HMA^{+}18, HTW01, HH11, KRR^{+}11, LTS98, MA05, Mer24, RCA18, Sne99a, Tuk91, Væt85, ZTK13]. Used [NP97]. Useful [Goo86b, Liu14b, And87, GS87, Goo86a]. User [Gue08, Nel74]. Users [GO10, Mai12b, Voe09]. Uses [Des00, LLP01, Sen11c, Bal76a, Bal76b]. Using [AG99, AT15, BCM19, BL24, BH14, BO15, Bra15, Cap19, CG05, CG12, CO14, CS02, Cro11b, DC19, DFV99a, DFV99b, Esp15b, Eva07a, FB73, Gho17, GDSN18, Han13d, Hau08a, LLT04, Liu22a, Mac97, Mai11c, Mai14c, MBW^{+}23, MS23, Muk15, Mul76, Num13a, O’B08a, O’B09e, O’B09c, OH13, OS73, Pre89, Rei09, Ric11b, Ric13a, Ric17b, RG23, Rob11c, Ros15b, Sco13a, SAP97, Sta07a, Sta07f, Sta07b, Sta07e, Sta07c, Sun17, Tam82, TC99, Til98, Unw11a, Unw13a, Unw24, Veh07a, Veh13a, Cha08, GP11a, Hot16, Lon15, MS87, MDS^{+}19, RDM^{+}19, Roy86, Tam87, TTB86, WGKL22, Pol11b]. Usual [Goo86b, And87, GS87, Goo86a]. utilisés [Ios77b]. Utility [CKK^{+}11b, HKZ07]. Uusipaikka [Lis09b].

V [Bar82, Ber82, Buc78, Csi80, Der75a, Der77c, Din82, Gan78a, Gho08a, Gho12b, Gow79, Han12b, Huz73, Ios77d, Kem80, Lis09a, Liu19, Mai10a, Mer20a, Pin75, Pos76e, Pun10d, Roh78, Ros09b, Sha82, Sil82, Smi82, Tin78c, Wet75, Wim75, Zad82, dO79a, Har75d]. Vaccaro [Mai11b]. Vaccine [Aur10]. Vach [Num13c]. Václav [Žal79]. Vague [SS08]. Vajda [Fer80a]. Valencia [DMS06]. Valentin [Gal80a]. Valeriu [Oni78]. Validation [Nee09]. Validity [BC20, GMZ03]. Valter [Lut76]. Valuation [Sco13b]. Value [CFL^{+}21, CP96, Dur22, GG15, Har78, KX09, LR91, Pun11d, Tam07, Zha17a, DY16]. valued [SS08]. Values [Kuk20, OH19, Tho06, CK19]. Vannucci [Bla07]. Variability [BO15]. Variable [AS72, CG12, DS22, FGDS18, FGFBGM22, Han13h, HK77, Lec76, NC10, Pyn14a, STI06, KDTP20, VLM12]. Variables [AT15, Ano78g, Frö08, GSGdMLM16, HS88, Lec76, Liu23, Mer24, MR10, Rip83, San91, Sär11, Smi20, SM14, Tho78a, Tho78b, Sch83, Lin80a, Num15]. Variance [AJS81, BE17, BGV18, BS85, Ela73, FS08, GW23, Har90, Hir93, KS85, Khu00, Sah79, Spr75a, SH23, SM89, TS06, Tju84, TBSW84, Tra07, VT19, Zha96, BNKS82, VLM12, Pos79d]. Variance-Gamma [FS08]. Variance-Mean [TS06, BNKS82]. Variance-Smoothing [GW23]. Variances [Ela73, Tal93, Bin83]. Variants [GS19, TGM09]. Variate [Tsa24]. Variates [Goo73]. Variation [AM93a, BE99, BR07, Bie99, Mat73, OA91]. Variational [GM91]. Variations [DQS92, EH91]. Varying [FZ15, Hon15, HS15, MSW76, PMLL15a, PMLL15b, WY02, KDTP20]. Varying-Coefficient [WY02]. Vasishth [Rob11g]. Vaughn [Han12c, Han12l]. VEC [PWKO24]. Vector [LD24, Ma17, RC93]. Veeravalli [Liu19]. Vehicle [CS00]. Veldman [Som78c]. Venezuelan [CM06]. Venugopal [Liu19]. Verbatim [SVCV88]. Verbeke [Cro09b]. Vere [Val08]. Vere-Jones [Val08]. Vergottini [Bar72a]. Verified [Mar21]. Verleysen [Hol08]. Versus [TLL04, Bai85b, HK18, MDS94, SVCV88]. Vervaat [Ios76f]. Very [RCA18]. Vetula [Bel00]. VI [Har76]. Via [BJPT99, AT15, DTP^{+}20, Gho20, Pod14b, Sha23]. Viable [GL14]. Vichi [Dur13b]. Vicious [Kir05]. Victor [Cro09c, Dra09d, RR79a]. Vidakovic [O’B13h, Oja08b]. Vidal [Zha16b]. Vidyadhar [Han10e]. Viertl [O’B13g]. View [Cox97a, Cox97b, CW04, DCC98, Dem72, Gra97, Gut03, Has72, Haw97a, JM89, NP97, PR10, Teu97, WS01, Ham77b, SS87, dF79, Gho09h]. Viewpoint [Cro09e, You20]. views [KASG80]. vii [McC22, Tso16]. viii [Blu20, Num23]. Vijayan [Sen84, Tam84, Tho84]. Vincent [Gho11d, Sau80]. Vincenzo [Han12g]. Vincze [Kri74]. Vine [HCR18]. Vingt [Fou80b]. Vining [Cro10g, Khu13a]. Vinzi [Han12g]. Vipin [O’B10d]. Virginia [Ric10a, Spr75a]. Virtual [DPS07, Wil07]. Vistocco [Dek14]. Visualisation [GDSN18]. Visualization [Liu22a, Nor08, Veh08e]. Visualizing [O’B10e, Han12x, Unw11c]. Vital [FL05, Hoe87, Som78e, Som79a, Som76c].

VitalSource [Ana21]. Vivek [Cro09e]. Vlad [Cro09d]. Vladimir [Gal80a, Pun10d, Ser78]. Vol [Bar76, Buc76a, Buc76b, Buc77b, Cop77, Ios76b, Ken73b, Knü80]. Volatility [MFAG10, Pod17c, Ber16]. Volker [Mat07]. Volume [And78, Ano73q, Ano74q, Ano75r, Ano75s, Ano76p, Ano77n, Ano78p, Ano79m, Ano80k, Ano81g, Ano82g, Ano83h, Ano83i, Ano84h, Ano84i, Ano85g, Ano85h, Ano86i, Ano86j, Ano87h, Ano87i, Ano88h, Ano88i, Ano89h, Ano89i, Ano90g, Ano90h, Ano91g, Ano91h, Ano92g, Ano93h, Ano94h, Ano95g, Ano96h, Ano97g, Ano98g, Ano99i, Ano00h, Ano01h, Ano02g, Ano03g, Ano04g, Ano05i, Ano06h, Ano13b, Ber78, Buc78, Dra09a, Dra12c, Eng78, Fil09a, Gan80b, Gho11d, Ios79b, O’B11a, Tin80b, Val08, Wil73]. Vorth [Dur13a]. Votes [Har73]. Vranić [Ser78]. vs [Bre99b, Has72, Sär72c].

W [BPV21, Bil77b, Blu21, Cor77a, Cro08b, Cro11e, Far13, Fin77a, Fin78, Fin80c, Gho13c, Gow72, Gow79, Haj77, Han08g, Han12u, Han12g, Ho05, Ios76f, Kem74, Ken73a, Ken76b, Kos11, Kró80d, Lis13, Meh75, Nee09, Nor09a, Oja11, Oni79, Oni80, Ord80a, Pat75b, Pod24, Pol13b, Pos79d, Pre79a, Pun11b, Ric09, Ric10c, Ric18, Roh80, Sch72c, Som78e, Som79b, Sta09, Stu72, Tin80b, Var20, Wol77, Zha19b]. W. [Dem83, MB19]. Waddington [Tip75a]. Wagenführ [Men76]. Wahlund [Wol77]. Wahrscheinlichkeit [Oni79, Oni80]. Wahrscheinlichkeitstheorie [Roh80]. Wainer [Han13b]. Wainwright [You20]. Wald [Le 90b, Væt85]. Walker [Cro11a]. Wall [Mer09b]. Wallace [Lan01]. Wallman [Cit16]. Walls [Coo07]. Walpole [Tip77, Som78b]. Walter [Cro11b, Dru11, Fil09b, Fin75, Lis13, Mar21, Men73, Zha19b, Büh75]. Wan [Num14]. Wand [Num20]. Wang [Ken10, Pol12b, Sun16a, Tra13, Han12g, O’B08c, O’B12i]. Wang-Iverson [Pol12b]. Want [Spe10a]. War [Tip75a]. Warranty [Law98]. Warren [Fer80b]. Warwick [Som76b]. Wataru [Han10a]. Water [Tan97]. Watkins [Smi80]. Watson [Blu07]. Wavelet [BDS04, O’B09i]. Wavelets [Pun12b, Pod22b]. Way [Alb19, AJS81, GPdMLMGGSG20, NAL76, Pod14a, SKK18, SM89, TSR96, Don86, LHK78]. Wayne [Sen14d]. Ways [Nar74]. Weak [Mat73]. Wealth [Pod72, SHL05]. Weapon [HHD97]. Weather [Pod21b]. Web [Han07b, Bet10, Had10c]. Web-Based [Had10c]. Wei [Gho12c, Han11j]. Wei-Bin [Gho12c]. Weibull [EPP08, SDT^{+}23, TC72]. Weighing [Pre77b]. Weight [WWY23]. Weighted [Bow04a, FS94, SB04]. Weighting [Chi22, WL16]. Weights [Pfe93]. Weinberger [Csi08]. Weintraub [Rév73]. Welfare [Fis72, WRH05]. Well [CFL^{+}21, MSS20, Sne99b, Szc11b]. Well-being [MSS20]. Well-Established [CFL^{+}21]. Wellbeing [AH17]. Wellek [Han12q]. Welsh [Tra13]. Wendy [Alv11]. Went [Ald02]. Wenting [Han13o]. Werner [Num13c]. Weron [Kos12c]. Wesley [Han11a]. West [Cro10c, Han10h, Rob11h]. Westfall [Ric11b]. Wetherill [Pos78a, Sco77a]. Wheelwright [Ord80b]. where [Hed20]. Which [CD18, God97, PWKO24]. White [CQOR^{+}22]. Whitney [Nem73, ZMSL19]. Whitney-type [ZMSL19]. Whittle [Man75]. Who [Dra09d]. Whole [Han12t]. Wicher [Kos10c]. Wichern [Pos79d]. Wiedermann [Dur21]. Wilcox [Sta12c]. Wiley [Alb19, Ana21, Dur21, Els18, Els20, Fis18, Gho23, Hee16, Lan18, Liu20b, Mer17, Mer20a, Num19, O’B12d, O’B20, Pod21b, Ric18, Sun22, Tso20, Var19, Var20, You19, Zha21]. Wilf [Nel80]. Wilfredo [Bha07]. Wilhelm [Ada80, Sch84a]. Will [Sen03]. Willem [Veh08c]. William [Bar75c, Bre79b, Has08a, O’B13b, Pun13b, Var19]. Williams [Unw11c, Ios77c, Mai12a]. Williamson [Dek10]. Wills [Han12x]. Wilson [Gho09d, Dra09d]. WinBUGS [O’B13h, O’B12c]. Wind [Hau18, Sør93, ZG12]. Wing [Han09d]. Winkel [Han07f]. Winkelmann [Cro10b]. Winkler [Ada80, Has08a, Sch84a, Sch72c]. Wirtschaftswissenschaften [Sch80b]. Wishes [Ano74h]. within [Teu05b]. without [AS73, CW06, HB80, Sun86]. Witten [Nor14a]. Wolf [Pod24]. Wolfgang [Dur21, Emb08, Gho12h, Kos12a, Kos12c, Kos13c, O’B09a, Veh08e]. Wolfinger [Zha19b]. Wolfowitz [Wei84]. Wolter [Tra07]. Wolters [Sai08]. Women [BW20, Car00]. Wonnacott [Fin80b, Fin80b]. Won’t [LLP01]. Woo [Han11b]. Wood [Gho13c]. Woodward [Sen14d, O’B12c]. Work [Ano99a, Ano00a, Ano06a, Ano07a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a, Ber78, Eng78, Gol04, Hea95, Hol84, Kir82, Men16c, Nay09, Pat88, Pat75a, Spe08a, Sti88, Ano08a, Pea74]. Workbook [Som77b]. Workers [RM93]. Working [Ano83a, Blu05, Men16c, PF82]. Works [Cox12, Dra13d, Gay11, Gho12d]. World [Mal05, Par80, PHO86, Som78d, Spr77b, Tip75a, Wil16a, Ins73, Sta09]. Worth [Dra12f]. Would [Rob12b]. Wretman [Pos79b]. Writing [McC22]. Writings [Han12b]. Written [Bel89]. Wu [Cro08a, Dur17, O’B10d, Cro10i, Gho12c, Han08d, Han13o]. Wynn [Gho11a]. Wynne [Han12g].

X [Liu22b, You19, Liu17, Mai21, Pod21b, Rob19]. x2007 [Ano07e]. Xavier [Nan10b]. xi [Gho22, Num20, Rod21]. Xiangyu [Kha17]. Xiao [Pol12a, Pun10b]. Xiao-Li [Pol12a]. Xiaohui [Han12d]. Xie [Unw10b]. xii [Blu21, Mar21, Mer17, Pod22a]. xiii [Cox16, Kha20, Mer20a, Pod17b]. Xin [Num14, Pun10b]. Xindong [O’B10d]. Xing [Cro08a, Kos09b]. Xingqiu [He14]. xiv [Dur17, Gho20, Mai19, Pod17a, Pod24, Slo23, Slo24, Unw21]. xix [Mer20c, Mer20b, Pod22b, Qiw20, Ros19, Sun17, Sun19]. XLMiner [Han07a]. Xu [Liu23, Mai16b]. xv [Hau18, Liu23, Num19, Sun16a, You19]. xvi [Liu22b, Mai16b, Pod17c, Unw24]. xvii [Sen23, Unw20, You20]. xviii [Mai22]. xx [Cox19, O’B21]. xxi [Ana21, Mer18]. xxii [Dur22, Hee16, Kha17, Pod21a]. xxiii [Mai17]. xxiv [McE17]. xxv [Gho23]. xxvi [Ric17a]. xxviii [O’B20]. xxxii [Sun24]. xxxiv [Cap19]. xxxv [Var19].

Yadolah [Veh08h]. Yakir [And07]. Yamada [O’B09g]. Yamauti [Vos75].

Yan [BPV21, Pun10b, Zha17a]. Yanagida [Veh13a]. Yanai [Pun11d]. Yanhong [Gho12c]. Yasunori [Pun10d]. Yates [Hea95]. Ye [Pol11a]. Year [Dra12f]. Years [Ahn14, Cia14, Kal19, KMM^{+}17, Loh14a, Loh14b, Ric13b, RZ14, SH00, SZ14, Spe08a, Str14]. Yi [Zha16b]. Yield [MW81]. Yim [Pol12b]. Yin [Mai11e]. Yor [Pod14b]. Yosef [Mai10c]. Yoshi [Han10a]. Yoshio [Pun11d]. Young [Lis14, Som78c, Sun22]. Yu [Mai13c, Pun11a, Unw10b, Csi08]. Yu-Kang [Mai13c]. Yue [Han09d]. Yue-Qing [Han09d]. Yuedong [O’B12i]. Yuelin [Mai14b]. Yugoslavia [Zar90]. Yuichi [Han10a]. Yule [Coc73, Ald98, Kru89, Coc73]. Yuliya [Han10j]. Yüzbasi [Kha24]. Yves [O’B20].

Z [Dra08a, Dra13f, Fin80c, Tha07b, Dal77c, Fed81b]. Zacks [Mer14, Sil73]. Zahlenmaterial [Dri78]. Zaidi [Dek10]. Zakoian [Liu11]. Zarkovich [Bal76a, Bal76b]. Zatuchni [Kró80b]. Zaven [Rob12a]. Zdravko [Han07b]. Zeileis [Sco09b]. Zeng [Gho12c]. Zenga [Rad10]. zero [HMT22]. zero-inflated [HMT22]. Zeros [HPHdAM22]. Zhang [Mai11f, Han07f, Han12d, Mai11d]. Zhansheng [Kha17]. Zhao [Gho13a, He14]. Zhi [Nor13a]. Zhi-Hua [Nor13a]. Zhou [Nor13a]. Zieffler [Han13d]. Ziro [Vos75]. Zoë [Unw13a]. Zoia [Liu10a]. Zou [Han13m]. Zucchini [Cro11b]. zur [Lin75a]. Zuur [O’B07b]. Zvi [Ros19]. Zwinderman [Ros09c, Sta12e].

References

Alldredge:1978:GB

[AA78] J. Richard Alldredge and Noreen Gilb Alldredge. Geostatistics: A bibliography. International Statistical Review = Revue Internationale de Statistique, 46(1):77–88, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

                  5823 (electronic). URL http://www.jstor.org/stable/
                  1402511.

Alvares:2018:WDO

[AAF18] Danilo Alvares, Carmen Armero, and Anabel Forte. What does objective mean in a Dirichlet-multinomial process? International Statistical Review = Revue Internationale de Statistique, 86(1):106–118, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Aalen:1982:DPP

[Aal82] Odd O. Aalen. Discussion of paper by P. K. Andersen, Ø. Borgan, R. Gill and N. Keiding. International Statistical Review = Revue Internationale de Statistique, 50(3):244– 246, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402490.

Alessi:2011:NFS

[ABC11] Lucia Alessi, Matteo Barigozzi, and Marco Capasso. Nonfundamentalness in structural econometric models: A review. International Statistical Review = Revue Internationale de Statistique, 79(1):16–47, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41306163.

Abe:2015:DFD

[Abe15] Toshihiro Abe. Discussion on On Families of Distributions with Shape Parameters: Discussions. International Statistical Review = Revue Internationale de Statistique, 83(2): 193–197, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Andersen:1982:LNTa

[ABGK82a] Per Kragh Andersen, Ørnulf Borgan, Richard Gill, and Niels Keiding. Linear nonparametric tests for comparison of counting processes, with applications to censored survival data, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 50(3):219– 244, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor. org/stable/1402489. With discussion, and with a reply by the authors.

Andersen:1982:LNTb

[ABGK82b] Per Kragh Andersen, Ørnulf Borgan, Richard Gill, and Niels Keiding. [linear nonparametric tests for comparison of counting processes, with applications to censored survival data]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 50(3):256– 258, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402495.

Abraham:1982:TAT

[Abr82] Bovas Abraham. Temporal aggregation and time series. International Statistical Review = Revue Internationale de Statistique, 50(3):285–291, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402498.

Abraham:2005:SBI

[Abr05] Bovas Abraham. Statistics in business and industry: Implementation. International Statistical Review = Revue Internationale de Statistique, 73(2):173–176, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472651.

Altieri:2021:SSN

[AC21] Linda Altieri and Daniela Cocchi. Spatial sampling for non-compact patterns. International Statistical Review = Revue Internationale de Statistique, 89(3):532–549, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Adhikari:1984:OSSb

[ACV84a] A. K. Adhikari, A. Chaudhuri, and K. Vijayan. [optimum sampling strategies for randomized response trials]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 52(2):124–125, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403094.

Adhikari:1984:OSSa

[ACV84b] Arun Kumar Adhikari, Arijit Chaudhuri, and K. Vijayan. Optimum sampling strategies for randomized response trials. International Statistical Review = Revue Internationale de Statistique, 52(2):115–120, August 1984. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403090.

With discussion by P. K. Sen, S. M. Tam and M. E. Thompson, and a reply by the authors.

Adam:1980:WWH

[Ada80] Adolf Adam. Wilhelm Winkler—in honour of his 95th birthday. International Statistical Review = Revue Internationale de Statistique, 48(1):1–2, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402399.

Alders:2004:AFS

[AdB04] Maarten Alders and Joop de Beer. Assumptions on fertility in stochastic population forecasts. International Statistical Review = Revue Internationale de Statistique, 72 (1):65–79, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403843.

Asadi:2022:EMS

[ADE^{+}22] Majid Asadi, Karthik Devarajan, Nader Ebrahimi, Ehsan Soofi, and Lauren Spirko-Burns. Elaboration models with symmetric information divergence. International Statistical Review = Revue Internationale de Statistique, 90(3): 499–524, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Arjas:1999:ENS

[AdV99] Elja Arjas and Willem de Vries. Editors’ note: Statement of editorial policy. International Statistical Review = Revue Internationale de Statistique, 67(3):221, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403698.

Ardakani:2018:RFS

[AES18] Omid M. Ardakani, Nader Ebrahimi, and Ehsan S. Soofi. Ranking forecasts by stochastic error distance, information and reliability measures. International Statistical Review = Revue Internationale de Statistique, 86(3):442– 468, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Allende:2004:EOP

[AET04] Héctor Allende, Carlos Elías, and Soledad Torres. Estimation of the option prime: Microsimulation of backward stochastic differential equations. International Statistical Review = Revue Internationale de Statistique, 72(1): 107–121, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403846.

Alldredge:1976:RRA

[AG76] J. R. Alldredge and N. S. Gilb. Ridge regression: An annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 44(3):355–360, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402352.

Andersen:1981:DPD

[AG81] Per Kragh Andersen and Richard Gill. Discussion of paper by D. Oakes. International Statistical Review = Revue Internationale de Statistique, 49(3):253–255, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402607.

Allende:1999:USS

[AG99] Héctor Allende and Jorge Galbiati. [using statistics and statistical thinking to improve organisational performance]: Discussion. International Statistical Review = Revue Internationale de Statistique, 67(2):122–125, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403390.

Azzalini:2008:RLM

[AG08] Adelchi Azzalini and Marc G. Genton. Robust likelihood methods based on the skew-t and related distributions. International Statistical Review = Revue Internationale de Statistique, 76(1):106–129, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919586.

Azzalini:2015:DFD

[AG15] Adelchi Azzalini and Marc G. Genton. Discussion on On Families of Distributions with Shape Parameters: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(2):198–202, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Allin:2017:SNA

[AH17] Paul Allin and David J. Hand. From a system of national accounts to a process of national wellbeing accounting. International Statistical Review = Revue Internationale de Statistique, 85(2):355–370, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ahlgren:2002:ASL

[Ahl02] Andrew Ahlgren. [adults’ statistical literacy: Meanings, components, responsibilities]: Discussion. International Statistical Review = Revue Internationale de Statistique, 70(1):41–43, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403718.

Ahn:2014:DFY

[Ahn14] Hongshik Ahn. Discussions on fifty years of classification and regression trees: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(3):357– 359, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ahn:2015:SSC

[AHZK15] Chul Ahn, Moonseoung Heo, Song Zhang, and Mimi (Mimi Y.) Kim. Sample size calculations for clustered and longitudinal outcomes in clinical research. Chapman and Hall/CRC biostatistics series. CRC Press, Taylor and Francis, Boca Raton, FL, USA, 2015. ISBN 1-4665-5626-9 (hardcover). xv + 244 pp. LCCN R853.S7 A325 2015.

Aitkin:2018:HGS

[Ait18] Murray Aitkin. A history of the GLIM statistical package. International Statistical Review = Revue Internationale de Statistique, 86(2):275–299, August 2018. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Andersen:1981:TWA

[AJS81] Anders Holst Andersen, Eva Bjørn Jensen, and Geert Schou. Two-way analysis of variance with correlated errors. International Statistical Review = Revue Internationale de Statistique, 49(2):153–167, August 1981. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403069.

Albertsen:1995:PUS

[AK95] K. Albertsen and E. Kristensen. A previously unpublished solution to Erlang’s “Problem of Extinction of Families” by Carl Marius Christensen. International Statistical Review = Revue Internationale de Statistique, 63(2):242– 245, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403619.

Abraham:1986:FFI

[AL86] Bovas Abraham and Johannes Ledolter. Forecast functions implied by autoregressive integrated moving average models and other related forecast procedures. International Statistical Review = Revue Internationale de Statistique, 54(1):51–66, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403258.

Andridge:2010:RHD

[AL10] Rebecca R. Andridge and Roderick J. A. Little. A review of hot deck imputation for survey non-response. International Statistical Review = Revue Internationale de Statistique, 78(1):40–64, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919794.

Albertsen:1995:EF

[Alb95] K. Albertsen. The extinction of families. International Statistical Review = Revue Internationale de Statistique, 63(2):234–239, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403617.

Albert:2019:BRB

[Alb19] Jim Albert. Book review: The Bayesian Way: Introductory Statistics for Economists and Engineers, Svein Olav Nyberg, John Wiley and Sons, 2018, ix + 512 pages, £87.50, hardcover ISBN: 978-1-119-24687-9. International Statistical Review = Revue Internationale de Statistique, 87(1):179–181, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Aldrich:1998:DLS

[Ald98] John Aldrich. Doing least squares: Perspectives from Gauss and Yule. International Statistical Review = Revue Internationale de Statistique, 66(1):61–81, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403657.

Aldrich:1999:DLM

[Ald99] John Aldrich. Determinacy in the linear model: Gauss to Bose and Koopmans. International Statistical Review = Revue Internationale de Statistique, 67(2):211–219, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403399.

Aldrich:2000:FIP

[Ald00] John Aldrich. Fisher’s “inverse probability” of 1930. International Statistical Review = Revue Internationale de Statistique, 68(2):155–172, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403666.

Aldrich:2002:HLI

[Ald02] John Aldrich. How likelihood and identification went Bayesian. International Statistical Review = Revue Internationale de Statistique, 70(1):79–98, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403727.

Aldrich:2003:LEB

[Ald03] John Aldrich. The language of the English biometric school. International Statistical Review = Revue Internationale de Statistique, 71(1):109–129, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403877.

Aldrich:2005:SEH

[Ald05] John Aldrich. The statistical education of Harold Jeffreys. International Statistical Review = Revue Internationale de Statistique, 73(3):289–307, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472677.

Aldrich:2007:IEF

[Ald07] John Aldrich. Information and economics in Fisher’s design of experiments. International Statistical Review = Revue Internationale de Statistique, 75(2):131–149, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41508457.

Alexander:1994:SSA

[Ale94] Charles H. Alexander. [sample surveys 1975–1990; an age of reconciliation?]: Discussion. International Statistical Review = Revue Internationale de Statistique, 62(1): 21–28, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403541.

Allen:1972:SBL

[All72] R. G. D. Allen. The statistician and the British Library of Political and Economic Science. International Statistical Review = Revue Internationale de Statistique, 40(3):373– 375, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402473.

Alldredge:1979:BRB

[All79a] J. R. Alldredge. Book review: Advanced Geostatistics in the Mining Industry by M. Guarascio; M. David; C. Huijbregts. International Statistical Review = Revue Internationale de Statistique, 47(1):89, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403218.

Allen:1979:DG

[All79b] R. G. D. Allen. David Glass, 1911–1978. International Statistical Review = Revue Internationale de Statistique, 47(1):95–96, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403228.

AlvarezP:2011:SBR

[Alv11] Adolfo Alvarez P. Short book review: Exploratory Data Analysis with MATLAB, Second Edition, by Wendy L. Martinez, Angel R. Martinez, Jeffrey L. Solka. International Statistical Review = Revue Internationale de Statistique, 79(3):492, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305073.

Alvo:2014:SMR

[Alv14] Mayer Alvo. Statistical Methods for Ranking Data. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4939-1470-7. xi + 273 pp. LCCN QA278.75 .A48 2014. URL http://www.loc.gov/

catdir/enhancements/fy1411/2014945645-d.html;
http://www.loc.gov/catdir/enhancements/fy1411/2014945645-
t.html.

Alvarez:2020:LEQ

[Alv20] Francisco J. Rubio Alvarez. Letter to the Editor: ‘on quantile-based asymmetric family of distributions: Properties and inference’. International Statistical Review = Revue Internationale de Statistique, 88(3):793–796, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See response [GKV20].

Abraham:1993:VRD

[AM93a] Bovas Abraham and Jock MacKay. Variation reduction and designed experiments. International Statistical Review = Revue Internationale de Statistique, 61(1):121– 129, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403598.

Archambault:1993:SMP

[AM93b] Sylvain Archambault and Marc Moore. Statistiques morphologiques pour l’ajustement d’images. (French) [Morphological statistics for adjusting images]. International Statistical Review = Revue Internationale de Statistique, 61(2):283–297, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403629.

Atchade:2014:DBS

[AM14] Yves Atchade and George Michailidis. Discussions on “A Brief Survey of Modern Optimization for Statisticians”: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(1):71–75, April 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [LCZ14a].

Alkema:2023:IAR

[AMR23] Leontine Alkema, Thomas Brendan Murphy, and Adrian E. Raftery. Interview with Adrian Raftery. International Statistical Review = Revue Internationale de Statistique, 91(3):349–367, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Amundsen:1973:BRB

[Amu73] H. T. Amundsen. Book review: Examples in Statistics by R. L. Bolt. International Statistical Review = Revue Internationale de Statistique, 41(2):291, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402852.

Agresti:2001:MCO

[AN01] Alan Agresti and Ranjini Natarajan. Modeling clustered ordered categorical data: A survey. International Statistical Review = Revue Internationale de Statistique, 69(3): 345–371, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403450.

Anand:2021:BRB

[Ana21] Sathyanarayan Anand. Book review: Business Experiments with R, B. D. McCullough Wiley, 2021, xxi + 361 pages, $104, Reflowable softcopy on VitalSource eText ISBN: 978-1-119-68988-1, 1-119-68988-0. International Statistical Review = Revue Internationale de Statistique, 89(3):663–666, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anderson:1973:PSE

[And73a] Henry Anderson. Population and sample in economic forecasting. International Statistical Review = Revue Internationale de Statistique, 41(2):149–153, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402831.

Andrews:1973:CAE

[And73b] D. F. Andrews. Car accidents: Environmental aspects. International Statistical Review = Revue Internationale de Statistique, 41(2):235–239, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402839.

Andrews:1973:DEL

[And73c] D. F. Andrews. Developing examples for learning statistics; data and computing. International Statistical Review = Revue Internationale de Statistique, 41(2):225– 228, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402837.

Andrews:1973:GTT

[And73d] D. F. Andrews. A graphical tool for teletypewriter terminals. International Statistical Review = Revue Internationale de Statistique, 41(2):272–275, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402846.

Anderson:1976:EPT

[And76] Harald Anderson. Estimation of a proportion through randomized response. International Statistical Review = Revue Internationale de Statistique, 44(2):213–217, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1403279. International Association of Survey Statisticians Meeting on Randomized Response (Warsaw, 1975).

Anderson:1977:CST

[And77a] O. D. Anderson. A commentary on: “A Survey of Time Series” (Internat. Statist. Rev. 44 (1976), no. 1, 29–70) by S. Makridakis. International Statistical Review = Revue Internationale de Statistique, 45(3):273–297, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402541.

Anderson:1977:SSC

[And77b] O. D. Anderson. Some statistical conferences at Cambridge, England. International Statistical Review = Revue Internationale de Statistique, 45(1):91–93, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403006.

Anderson:1978:BRB

[And78] O. D. Anderson. Book review: The Advanced Theory of Statistics: Volume 1, Distribution Theory by Maurice Kendall; A. Stuart. International Statistical Review = Revue Internationale de Statistique, 46(3):320, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402383.

Andersen:1986:DPG

[And86] Erling B. Andersen. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):271–272, December 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403054.

Andersen:1987:ADP

[And87] Erling B. Andersen. Amendment to discussion of paper by L. A. Goodman: “Some useful extensions of the usual correspondence analysis approach and the usual log-linear models approach in the analysis of contingency tables” [Internat. Statist. Rev. 54 (1986), no. 3, 243–309]. International Statistical Review = Revue Internationale de Statistique, 55(2):229, 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Andrews:2007:SBR

[And07] David F. Andrews. Short book review: The Statistics of Gene Mapping by David Siegmund, Benjamin Yakir. International Statistical Review = Revue Internationale de Statistique, 75(2):275–276, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508490.

Anis:2008:BPC

[Ani08] Mohammed Z. Anis. Basic process capability indices: An expository review. International Statistical Review = Revue Internationale de Statistique, 76(3):347–367, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919651.

Ahn:2003:FDS

[ANK03] Yunkee Ahn, Buhn Nam, and Chul Eung Kim. The future direction of statistical training for the promotion of cooperation between government and academy. International Statistical Review = Revue Internationale de Statistique, 71(2):461–467, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403898.

Anonymous:1972:SBI

[Ano72a] Anonymous. 39ième session de L’Institut International de Statistique. International Statistical Review = Revue Internationale de Statistique, 40(1):97, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402713.

Anonymous:1972:SIS

[Ano72b] Anonymous. 39th session of the International Statistical Institute. International Statistical Review = Revue Internationale de Statistique, 40(1):96, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402712.

Anonymous:1972:AEG

[Ano72c] Anonymous. Addenda and errata for Gart, Kossack and Baxi. International Statistical Review = Revue Internationale de Statistique, 40(2):221–222, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402764.

Anonymous:1972:BMa 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402725.

Anonymous:1972:BMb

[Ano72e] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 40(2):??, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402781.

Anonymous:1972:BMc

[Ano72f] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 40(3):??, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402486.

Anonymous:1972:FMa

[Ano72g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 40(1):82–100, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402699.

Anonymous:1972:FMb

[Ano72h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 40(2):247– 248, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402753.

Anonymous:1972:FMc

[Ano72i] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 40(3):274– 392, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402463.

Anonymous:1972:PPC

[Ano72j] Anonymous. [photograph]: Prasanta Chandra Mahalanobis, 1893–1972. Honorary President of the International Statistical Institute. International Statistical Review = Revue Internationale de Statistique, 40(3):??, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402464.

Anonymous:1972:SOAa

[Ano72k] Anonymous. Statistical organization and administration. International Statistical Review = Revue Internationale de Statistique, 40(1):83–89, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402709.

Anonymous:1972:SOAb

[Ano72l] Anonymous. Statistical organization and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 40(2):223–235, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402765.

Anonymous:1972:SOAc

[Ano72m] Anonymous. Statistical organization and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 40(3):377–383, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402475.

Anonymous:1972:SS

[Ano72n] Anonymous. Statistical societies. International Statistical Review = Revue Internationale de Statistique, 40 (1):95, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402711.

Anonymous:1972:SSSa

[Ano72o] Anonymous. Statistical societies / Sociétés de statistique. International Statistical Review = Revue Internationale de Statistique, 40(2):245–246, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402767.

Anonymous:1972:SSSb

[Ano72p] Anonymous. Statistical societies / sociétés de statistique. International Statistical Review = Revue Internationale de Statistique, 40(3):389–390, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402477.

Anonymous:1972:UNSa

[Ano72q] Anonymous. United Nations Statistical Office publication recently issued. International Statistical Review = Revue Internationale de Statistique, 40(1):107–108, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402721.

Anonymous:1972:UNSb

[Ano72r] Anonymous. United Nations Statistical Office publications recently issued. International Statistical Review = Revue Internationale de Statistique, 40(2):256–258, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402777.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 41(1):148, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402807.

Anonymous:1973:BMb

[Ano73b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 41(2):??, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402867.

Anonymous:1973:BMc

[Ano73c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 41(3):??, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402643.

Anonymous:1973:BTS

[Ano73d] Anonymous. Bibliography on the teaching of statistics. International Statistical Review = Revue Internationale de Statistique, 41(3):381–386, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402626.

Anonymous:1973:FB

[Ano73e] Anonymous. Felix Burkhardt, 1888–1973. International Statistical Review = Revue Internationale de Statistique, 41(3):401–402, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402640.

Anonymous:1973:FMa

[Ano73f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 41(1):137– 138, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402783.

Anonymous:1973:FMb

[Ano73g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 41(2):287– 288, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402830.

Anonymous:1973:FMc

[Ano73h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 41(3):393– 394, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402618.

Anonymous:1973:SOAa

[Ano73i] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 41(1):129–130, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402793.

Anonymous:1973:SOAb

[Ano73j] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 41(2):277–282, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402847.

Anonymous:1973:SOAc

[Ano73k] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 41(3):387–388, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402627.

Anonymous:1973:SSSa

[Ano73l] Anonymous. Statistical societies / Sociétés de statistique. International Statistical Review = Revue Internationale de Statistique, 41(1):135, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402795.

Anonymous:1973:SSSb

[Ano73m] Anonymous. Statistical societies / Sociétés de statistiques. International Statistical Review = Revue Internationale de Statistique, 41(2):285–286, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402849.

Anonymous:1973:SSSc

[Ano73n] Anonymous. Statistical societies / Sociétés de statistiques. International Statistical Review = Revue Internationale de Statistique, 41(3):391–392, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402629.

Anonymous:1973:UNSa

[Ano73o] Anonymous. United Nations Statistical Office publications recently issued. International Statistical Review = Revue Internationale de Statistique, 41(1):145–146, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402804.

Anonymous:1973:UNSb

[Ano73p] Anonymous. United Nations Statistical Office publications recently issued. International Statistical Review = Revue Internationale de Statistique, 41(2):298–299, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402864.

Anonymous:1973:VI

[Ano73q] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 41 (1):??, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402782.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 42(1):115, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402698.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 42(2):??, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403088.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 42(3):??, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402996.

Anonymous:1974:EK

[Ano74d] Anonymous. Errata for Kotz. International Statistical Review = Revue Internationale de Statistique, 42(3):

314, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402992.

Anonymous:1974:FMa

[Ano74e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 42(1):103– 105, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402677.

Anonymous:1974:FMb

[Ano74f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 42(2):228– 230, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403074.

Anonymous:1974:FMc

[Ano74g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 42(3):315– 316, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402982.

Anonymous:1974:PJN

[Ano74h] Anonymous. [photograph]: Jerzy Neyman — Honorary President. Felicitations on eight decades — good wishes for the ninth. International Statistical Review = Revue Internationale de Statistique, 42(1):??, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402678.

Anonymous:1974:SOAc

[Ano74i] Anonymous. Statistical organisation and administration / organisation et administratien statistiques. International Statistical Review = Revue Internationale de Statistique, 42(3):313, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402990.

Anonymous:1974:SOAa

[Ano74j] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 42(1):71–78, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402686.

Anonymous:1974:SOAb

[Ano74k] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 42(2):213–223, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403082.

Anonymous:1974:SSS

[Ano74l] Anonymous. Statistical societies / Sociétés de statistique. International Statistical Review = Revue Internationale de Statistique, 42(1):89–102, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402688.

Anonymous:1974:STRa

[Ano74m] Anonymous. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 42(2):225– 227, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403083.

Anonymous:1974:STRb

[Ano74n] Anonymous. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 42(3):314, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402991.

Anonymous:1974:UNSa

[Ano74o] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 42(1):110–112, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402695.

Anonymous:1974:UNSb

[Ano74p] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 42(2):232–233, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403086.

Anonymous:1974:VI

[Ano74q] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 42 (1):??, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402676.

Anonymous:1975:AMA

[Ano75a] Anonymous. Announcement to members / avis aux members. International Statistical Review = Revue Internationale de Statistique, 43(2):237–238, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402905.

Anonymous:1975:BMa

[Ano75b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 43(1):124, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402675.

Anonymous:1975:BMb

[Ano75c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 43(2):??, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402928.

Anonymous:1975:BMc cember 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403136.

Anonymous:1975:ELR

[Ano75e] Anonymous. Engel’s law revisited: Résumé. International Statistical Review = Revue Internationale de Statistique, 43(2):210, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402899.

Anonymous:1975:ESM

[Ano75f] Anonymous. Errata: Statistical methods for rates and proportions. International Statistical Review = Revue Internationale de Statistique, 43(3):375, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403135.

Anonymous:1975:FMa

[Ano75g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 43(1):114– 115, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402657.

Anonymous:1975:FMb

[Ano75h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 43(2):233– 236, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402896.

Anonymous:1975:FMc

[Ano75i] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 43(3):365– 366, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403108.

Anonymous:1975:KP

[Ano75j] Anonymous. Karl Pribram, 1877–1973. International Statistical Review = Revue Internationale de Statistique, 43 (3):374–375, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403134.

Anonymous:1975:RBR

[Ano75k] Anonymous. Review books received (continuation list). International Statistical Review = Revue Internationale de Statistique, 43(3):367, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403122.

Anonymous:1975:SOAc

[Ano75l] Anonymous. Statistical organisation and administration / organisation et administration statistique. International Statistical Review = Revue Internationale de Statistique, 43(3):351–353, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403120.

Anonymous:1975:SOAa

[Ano75m] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 43(1):109–111, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402663.

Anonymous:1975:SOAb

[Ano75n] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 43(2):221–225, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402902.

Anonymous:1975:STR

[Ano75o] Anonymous. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 43(1):112– 113, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402664.

Anonymous:1975:UNSa

[Ano75p] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 43(1):121–122, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402673.

Anonymous:1975:UNSb

[Ano75q] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 43(2):250–252, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402927.

Anonymous:1975:VIa

[Ano75r] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 43 (1):??, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402656.

Anonymous:1975:VIb

[Ano75s] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 43(3): ??, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403107.

Anonymous:1976:AIM

[Ano76a] Anonymous. 50th anniversary of ISI membership. International Statistical Review = Revue Internationale de Statistique, 44(3):388, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402369.

Anonymous:1976:AMA

[Ano76b] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 44(1):166–167, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402969.

Anonymous:1976:BMa

[Ano76c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 44(1):??, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402981.

Anonymous:1976:BMb

[Ano76d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 44(2):??, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403308.

Anonymous:1976:BMc

[Ano76e] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 44(3):??, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402370.

Anonymous:1976:EAS

[Ano76f] Anonymous. Errata: Applications of the structural method in statistical inference. International Statistical Review = Revue Internationale de Statistique, 44(2): 300, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403290.

Anonymous:1976:FMa

[Ano76g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 44(1):164– 165, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402962.

Anonymous:1976:FMb

[Ano76h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 44(2):298– 299, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403275.

Anonymous:1976:FMc

[Ano76i] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 44(3):374– 376, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402345.

Anonymous:1976:SOA

[Ano76j] Anonymous. Statistical organisation and administration / organisation et administration statistiques. International Statistical Review = Revue Internationale de Statistique, 44(1):160–161, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402967.

Anonymous:1976:STRa

[Ano76k] Anonymous. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 44(1):161– 163, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402968.

Anonymous:1976:STRb

[Ano76l] Anonymous. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 44(3):369– 373, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402355.

Anonymous:1976:SMA

[Ano76m] Anonymous. Strategies in the multivariate analysis of dates from complete surveys II: An application to the United States National Health Interview Survey: Résumé. International Statistical Review = Revue Internationale de Statistique, 44(3):348, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402350.

Anonymous:1976:UNSb

[Ano76n] Anonymous. United Nations Statistical Office: Publication recently issued. International Statistical Review = Revue Internationale de Statistique, 44(2):309, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403306.

Anonymous:1976:UNSa

[Ano76o] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 44(1):173–177, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402975.

Anonymous:1976:VI

[Ano76p] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 44(3):389– 390, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402344.

Anonymous:1977:AMAa

[Ano77a] Anonymous. Announcement to members / avis aux members. International Statistical Review = Revue Internationale de Statistique, 45(1):94–95, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403007.

Anonymous:1977:AMAb

[Ano77b] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 45(2):201, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403319.

Anonymous:1977:AMAc

[Ano77c] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 45(3):318–321, December 1977. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402544.

Anonymous:1977:BMa 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403032.

Anonymous:1977:BMb

[Ano77e] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 45(2):??, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403331.

Anonymous:1977:BMc

[Ano77f] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 45(3):??, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402561.

Anonymous:1977:FMa

[Ano77g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 45(1):??, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402997.

Anonymous:1977:FMb

[Ano77h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 45(2):??, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403309.

Anonymous:1977:FMc

[Ano77i] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 45(3):??, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402534.

Anonymous:1977:Rb

[Ano77j] Anonymous. Résumé. International Statistical Review = Revue Internationale de Statistique, 45(3):256, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402539.

Anonymous:1977:Ra

[Ano77k] Anonymous. Resumes. International Statistical Review = Revue Internationale de Statistique, 45(2):200, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403318.

Anonymous:1977:UNSa

[Ano77l] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 45(1):106, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403027.

Anonymous:1977:UNSb

[Ano77m] Anonymous. United Nations Statistical Office; publications recently issued. International Statistical Review = Revue Internationale de Statistique, 45(2):205–208, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403329.

Anonymous:1977:VI

[Ano77n] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 45(3):329– 330, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402533.

Anonymous:1978:AMAa

[Ano78a] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 46(1):115–116, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402514.

Anonymous:1978:AMAb

[Ano78b] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 46(2):227–228, August 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402818.

Anonymous:1978:AMAc

[Ano78c] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 46(3):317–318, December 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402378.

Anonymous:1978:BMa

[Ano78d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 46(1):??, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402532.

Anonymous:1978:BMb

[Ano78e] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 46(2):??, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402829.

Anonymous:1978:BMc

[Ano78f] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 46(3):??, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402397.

Anonymous:1978:ESV

[Ano78g] Anonymous. Errata: Selection of Variables in Multiple Regression: Part I. A Review and Evaluation. International Statistical Review = Revue Internationale de Statistique, 46(2):146, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402810. See [Tho78a].

Anonymous:1978:FMa

[Ano78h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 46(1):??, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402504.

Anonymous:1978:FMb

[Ano78i] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 46(2):??, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402808.

Anonymous:1978:FMc

[Ano78j] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 46(3):??, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402372.

Anonymous:1978:RPH

[Ano78k] Anonymous. Recommendations for 1980 population and housing censuses: An interim report by the UN Secretariat (ST/ESA/STAT/91). International Statistical Review = Revue Internationale de Statistique, 46(1):103–114, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402513.

Anonymous:1978:R

[Ano78l] Anonymous. Resumes. International Statistical Review = Revue Internationale de Statistique, 46(2):236, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402828.

Anonymous:1978:SOP

[Ano78m] Anonymous. UN Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 46(3):323–324, December 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402392.

Anonymous:1978:UNSa

[Ano78n] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 46(1):123–124, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402527.

Anonymous:1978:UNSb

[Ano78o] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 46(2):230–232, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402824.

Anonymous:1978:VI

[Ano78p] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 46(3):329– 330, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402371.

Anonymous:1979:AMA

[Ano79a] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 47(2):229–230, August 1979. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402570.

Anonymous:1979:BMa

[Ano79b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 47(1):??, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403232.

Anonymous:1979:BMb

[Ano79c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 47(2):??, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402582.

Anonymous:1979:BMc cember 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402655.

Anonymous:1979:CNE

[Ano79e] Anonymous. Corrigendum note: Estimating the total number of events with data from multiple-record systems: A review of methodological strategies. International Statistical Review = Revue Internationale de Statistique, 47 (1):36, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403205.

Anonymous:1979:FMa

[Ano79f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 47(1):??, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403200.

Anonymous:1979:FMb

[Ano79g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 47(2):??, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402562.

Anonymous:1979:FMc

[Ano79h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 47(3):??, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402645.

Anonymous:1979:JS

[Ano79i] Anonymous. Julius Shiskin, 1912–1978. International Statistical Review = Revue Internationale de Statistique, 47 (3):301, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402654.

Anonymous:1979:Ra

[Ano79j] Anonymous. Résumés. International Statistical Review = Revue Internationale de Statistique, 47(1):46, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403207.

Anonymous:1979:Rb

[Ano79k] Anonymous. Résumés. International Statistical Review = Revue Internationale de Statistique, 47(1):84, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403210.

Anonymous:1979:UNS

[Ano79l] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 47(1):92–94, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403226.

Anonymous:1979:VI

[Ano79m] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 47(3):303– 304, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402644.

Anonymous:1980:AMA

[Ano80a] Anonymous. Announcement to members / avis aux membres. International Statistical Review = Revue Internationale de Statistique, 48(1):117, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402411.

Anonymous:1980:BMa

[Ano80b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 48(1):??, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402446.

Anonymous:1980:BMb

[Ano80c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 48(2):??, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403188.

Anonymous:1980:BMc

[Ano80d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 48(3):??, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402961.

Anonymous:1980:CPE

[Ano80e] Anonymous. Corrections: Postcensal estimates for local areas (or domains). International Statistical Review = Revue Internationale de Statistique, 48(2):196, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403152.

Anonymous:1980:FMa

[Ano80f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 48(1):??, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402398.

Anonymous:1980:FMb

[Ano80g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 48(2):??, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403149.

Anonymous:1980:FMc

[Ano80h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 48(3):??, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402939.

Anonymous:1980:UNSa

[Ano80i] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 48(1):133–134, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402440.

Anonymous:1980:UNSb

[Ano80j] Anonymous. United Nations Statistical Office: Publications recently issued. International Statistical Review = Revue Internationale de Statistique, 48(2):246–248, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403187.

Anonymous:1980:VI

[Ano80k] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 48(3):375– 377, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402938.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 49(1):??, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403041.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 49(2):??, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403073.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 49(3):??, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402617.

Anonymous:1981:FMa

[Ano81d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 49(1):??, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403033.

Anonymous:1981:FMb

[Ano81e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 49(2):??, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403065.

Anonymous:1981:FMc

[Ano81f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 49(3):??, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402604.

Anonymous:1981:VI

[Ano81g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 49(3): ??, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402603.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 50(1):??, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402462.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 50(2):??, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402602.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 50(3):??, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402503.

Anonymous:1982:FMa

[Ano82d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 50(1):??, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402447.

Anonymous:1982:FMb

[Ano82e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 50(2):??, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402594.

Anonymous:1982:FMc

[Ano82f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 50(3):??, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402488.

Anonymous:1982:VI

[Ano82g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 50(3): 331, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402487.

Anonymous:1983:AFW

[Ano83a] Anonymous. Amendment: Festival and working days prior adjustments in economic time series. International Statistical Review = Revue Internationale de Statistique, 51(2): 223, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402751.

Anonymous:1983:BMa

[Ano83b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 51(1):??, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402736.

Anonymous:1983:BMb

[Ano83c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 51(2):??, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402752.

Anonymous:1983:BMc

[Ano83d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 51(3):??, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402593.

Anonymous:1983:FMa

[Ano83e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 51(1):??, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402727.

Anonymous:1983:FMb

[Ano83f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 51(2):??, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402737.

Anonymous:1983:FMc

[Ano83g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 51(3):??, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402584.

Anonymous:1983:VIa

[Ano83h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 51 (1):??, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402726.

Anonymous:1983:VIb

[Ano83i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 51(3): ??, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402583.

Anonymous:1984:ALN

[Ano84a] Anonymous. Amendment: Linear nonparametric tests for comparison of counting processes, with applications to censored survival data, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 52(2):??, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403105.

Anonymous:1984:BMa

[Ano84b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 52(1):??, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403251.

Anonymous:1984:BMb

[Ano84c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 52(2):??, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403106.

Anonymous:1984:BMc

[Ano84d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 52(3):??, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403050.

Anonymous:1984:FMa

[Ano84e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 52(1):??, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403234.

Anonymous:1984:FMb

[Ano84f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 52(2):??, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403089.

Anonymous:1984:FMc

[Ano84g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 52(3):??, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403043.

Anonymous:1984:VIa

[Ano84h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 52 (1):??, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403233.

Anonymous:1984:VIb

[Ano84i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 52(3):327– 328, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403042.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 53(1):??, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402884.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 53(2):??, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402937.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 53(3):??, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402895.

Anonymous:1985:FMa

[Ano85d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 53(1):??, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402869.

Anonymous:1985:FMb

[Ano85e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 53(2):??, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402929.

Anonymous:1985:FMc

[Ano85f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 53(3):??, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402886.

Anonymous:1985:VIa

[Ano85g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 53 (1):??, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402868.

Anonymous:1985:VIb

[Ano85h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 53(3): ??, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402885.

Anonymous:1986:ARL

[Ano86a] Anonymous. Amendment: Robust linear prediction in finite populations. International Statistical Review = Revue Internationale de Statistique, 54(1):115, April 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403263.

Anonymous:1986:BMa

[Ano86b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 54(1):??, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403264.

Anonymous:1986:BMb

[Ano86c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 54(2):??, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403148.

Anonymous:1986:BMc

[Ano86d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 54(3):??, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403064.

Anonymous:1986:FMa

[Ano86e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 54(1):??, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403253.

Anonymous:1986:FMb

[Ano86f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 54(2):??, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403137.

Anonymous:1986:FMc

[Ano86g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 54(3):??, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403052.

Anonymous:1986:ISI

[Ano86h] Anonymous. International Statistical Institute declaration on professional ethics. International Statistical Review = Revue Internationale de Statistique, 54(2):227– 242, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403147.

Anonymous:1986:VIa

[Ano86i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 54 (1):??, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403252.

Anonymous:1986:VIb

[Ano86j] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 54(3): ??, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403051.

Anonymous:1987:AAG

[Ano87a] Anonymous. Amendments for Anderson and Gill. International Statistical Review = Revue Internationale de Statistique, 55(2):229, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403198.

Anonymous:1987:BMa

[Ano87b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 55(1):??, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403274.

Anonymous:1987:BMb

[Ano87c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 55(2):??, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403199.

Anonymous:1987:BMc

[Ano87d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 55(3):??, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403409.

Anonymous:1987:FMa

[Ano87e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 55(1):??, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403266.

Anonymous:1987:FMb

[Ano87f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 55(2):??, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403189.

Anonymous:1987:FMc

[Ano87g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 55(3):??, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403402.

Anonymous:1987:VIa

[Ano87h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 55 (1):??, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403265.

Anonymous:1987:VIb

[Ano87i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 55(3): 331, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403401.

Anonymous:1988:AES

[Ano88a] Anonymous. Amendment: Experimental and sampling structures: Parallels diverging and meeting. International Statistical Review = Revue Internationale de Statistique, 56(2):197, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403642.

Anonymous:1988:BMa

[Ano88b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 56(1):??, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403365.

Anonymous:1988:BMb

[Ano88c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 56(2):??, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403643.

Anonymous:1988:BMc

[Ano88d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 56(3):??, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403355.

Anonymous:1988:FMa

[Ano88e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 56(1):??, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403357.

Anonymous:1988:FMb

[Ano88f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 56(2):??, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403634.

Anonymous:1988:FMc

[Ano88g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 56(3):??, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403348.

Anonymous:1988:VIa

[Ano88h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 56 (1):??, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403356.

Anonymous:1988:VIb

[Ano88i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 56(3): ??, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403347.

Anonymous:1989:AME

[Ano89a] Anonymous. Amendment: On multivariate Edgeworth expansions. International Statistical Review = Revue Internationale de Statistique, 57(2):183, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403386.

Anonymous:1989:BMa

[Ano89b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 57(1):??, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403587.

Anonymous:1989:BMb

[Ano89c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 57(2):??, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403387.

Anonymous:1989:BMc

[Ano89d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 57(3):??, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403800.

Anonymous:1989:FMa

[Ano89e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 57(1):??, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403580.

Anonymous:1989:FMb

[Ano89f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 57(2):??, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403380.

Anonymous:1989:FMc

[Ano89g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 57(3):??, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403792.

Anonymous:1989:VIa

[Ano89h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 57 (1):??, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403579.

Anonymous:1989:VIb

[Ano89i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 57(3): 295, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403791.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 58(1):??, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403476.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 58(2):??, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403466.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 58(3):??, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403810.

Anonymous:1990:FMa

[Ano90d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 58(1):??, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403468.

Anonymous:1990:FMb

[Ano90e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 58(2):??, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403460.

Anonymous:1990:FMc

[Ano90f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 58(3):??, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403802.

Anonymous:1990:VIa

[Ano90g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 58 (1):??, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403467.

Anonymous:1990:VIb

[Ano90h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 58(3): ??, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403801.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 59(1):??, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403578.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 59(2):??, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403447.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 59(3):??, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403695.

Anonymous:1991:FMa

[Ano91d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 59(1):??, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403570.

Anonymous:1991:FMb

[Ano91e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 59(2):??, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403438.

Anonymous:1991:FMc

[Ano91f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 59(3):??, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403686.

Anonymous:1991:VIa

[Ano91g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 59 (1):??, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403569.

Anonymous:1991:VIb

[Ano91h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 59(3): ??, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403685.

Anonymous:1992:BMa

[Ano92a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 60(1):??, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403505.

Anonymous:1992:BMb

[Ano92b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 60(2):??, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403652.

Anonymous:1992:BMc

[Ano92c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 60(3):??, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403684.

Anonymous:1992:FMa

[Ano92d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 60(1):??, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403495.

Anonymous:1992:FMb

[Ano92e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 60(2):??, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403644.

Anonymous:1992:FMc

[Ano92f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 60(3):??, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403676.

Anonymous:1992:VI

[Ano92g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 60 (1):??, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403494.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 61(1):??, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403603.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 61(2):??, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403633.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 61(3):??, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403759.

Anonymous:1993:CIA

[Ano93d] Anonymous. Correction: “Inference and Auditing: The Stringer Bound”. International Statistical Review = Revue Internationale de Statistique, 61(3):487, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1403758. See [Bic92].

Anonymous:1993:FMa

[Ano93e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 61(1):??, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403588.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 61(2):??, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403622.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 61(3):??, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403747.

Anonymous:1993:VI

[Ano93h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 61(3): ??, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403746.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 62(1):??, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403551.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 62(2):??, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403513.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 62(3):??, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403771.

Anonymous:1994:ENM

[Ano94d] Anonymous. Editor’s note: The Morris Hansen Lectures. International Statistical Review = Revue Internationale de Statistique, 62(1):1, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403537.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 62(1):??, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403536.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 62(2):??, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403506.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 62(3):??, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403761.

Anonymous:1994:VI

[Ano94h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 62(3): ??, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403760.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 63(1):??, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403781.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 63(2):??, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403621.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 63(3):??, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403485.

Anonymous:1995:FMa

[Ano95d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 63(1):??, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403772.

Anonymous:1995:FMb

[Ano95e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 63(2):??, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403604.

Anonymous:1995:FMc

[Ano95f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 63(3):??, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403478.

Anonymous:1995:VI

[Ano95g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 63(3): ??, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403477.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 64(1):??, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403428.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 64(2):??, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403745.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 64(3):??, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403790.

Anonymous:1996:CPM

[Ano96d] Anonymous. Corrigenda: Pivotal models and the fiducial argument. International Statistical Review = Revue Internationale de Statistique, 64(1):??, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403427.

Anonymous:1996:FMa

[Ano96e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 64(1):??, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403420.

Anonymous:1996:FMb

[Ano96f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 64(2):139, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403732.

Anonymous:1996:FMc

[Ano96g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 64(3):??, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403783.

Anonymous:1996:VI

[Ano96h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 64(3): ??, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403782.

Anonymous:1997:BMa

[Ano97a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 65(1):??, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403437.

Anonymous:1997:BMb

[Ano97b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 65(2):??, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403346.

Anonymous:1997:BMc

[Ano97c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 65(3):??, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403379.

Anonymous:1997:FMa

[Ano97d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 65(1):??, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403429.

Anonymous:1997:FMb

[Ano97e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 65(2):??, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403332.

Anonymous:1997:FMc

[Ano97f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 65(3):??, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403367.

Anonymous:1997:VI

[Ano97g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 65(3): ??, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403366.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 66(1):??, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403661.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 66(2):??, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403493.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 66(3):??, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403521.

Anonymous:1998:FMa

[Ano98d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 66(1):??, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403653.

Anonymous:1998:FMb

[Ano98e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 66(2):??, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403486.

Anonymous:1998:FMc

[Ano98f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 66(3):??, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403515.

Anonymous:1998:VI

[Ano98g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 66(3): ??, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403514.

Anonymous:1999:ARW

[Ano99a] Anonymous. Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 67(3):355, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:1999:BMa

[Ano99b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 67(1):??, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403568.

Anonymous:1999:BMb

[Ano99c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 67(2):??, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403400.

Anonymous:1999:BMc

[Ano99d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 67(3):??, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403711.

Anonymous:1999:ENS

[Ano99e] Anonymous. Editors’ note: Statement of editorial policy. International Statistical Review = Revue Internationale de Statistique, 67(3):221, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:1999:FMa

[Ano99f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 67(1):??, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403561.

Anonymous:1999:FMb

[Ano99g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 67(2):??, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403388.

Anonymous:1999:FMc

[Ano99h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 67(3):??, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403697.

Anonymous:1999:VI

[Ano99i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 67(3): ??, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403696.

Anonymous:2000:ARW

[Ano00a] Anonymous. Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 68(3):i, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2000:BMa

[Ano00b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 68(1):??, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403560.

Anonymous:2000:BMb

[Ano00c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 68(2):??, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403675.

Anonymous:2000:BMc

[Ano00d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 68(3):??, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403419.

Anonymous:2000:FMa

[Ano00e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 68(1):??, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403552.

Anonymous:2000:FMb

[Ano00f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 68(2):??, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403662.

Anonymous:2000:FMc

[Ano00g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 68(3):??, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403411.

Anonymous:2000:VI

[Ano00h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 68(3): ??, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403410.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 69(1):??, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403535.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 69(2):??, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403820.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 69(3):??, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403459.

Anonymous:2001:E

[Ano01d] Anonymous. Editorial. International Statistical Review = Revue Internationale de Statistique, 69(1):1–2, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2001:FMa

[Ano01e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 69(1):1–2, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403522.

Anonymous:2001:FMb

[Ano01f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 69(2):??, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403811.

Anonymous:2001:FMc

[Ano01g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 69(3):??, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403449.

Anonymous:2001:VI

[Ano01h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 69(3): ??, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403448.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 70(1):??, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403731.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 70(2):??, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403912.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 70(3):??, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403867.

Anonymous:2002:FMa

[Ano02d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 70(1):??, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403712.

Anonymous:2002:FMb

[Ano02e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 70(2):??, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403900.

Anonymous:2002:FMc

[Ano02f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 70(3):??, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403860.

Anonymous:2002:VI

[Ano02g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 70(3): ??, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403859.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 71(1):??, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403880.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 71(2):??, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403899.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 71(3):??, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403836.

Anonymous:2003:FMa

[Ano03d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 71(1):??, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403868.

Anonymous:2003:FMb

[Ano03e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 71(2):??, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403881.

Anonymous:2003:FMc

[Ano03f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 71(3):??, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403822.

Anonymous:2003:VI

[Ano03g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 71(3):667– 669, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403821.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 72(1):??, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403849.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 72(2):??, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403858.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 72(3):??, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472632.

Anonymous:2004:FMa

[Ano04d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 72(1):??, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403837.

Anonymous:2004:FMb

[Ano04e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 72(2):??, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403850.

Anonymous:2004:FMc

[Ano04f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 72(3):??, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472623.

Anonymous:2004:VI

[Ano04g] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 72(3): ??, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472622.

Anonymous:2005:ISC

[Ano05a] Anonymous. The 2004 ISI Special Conference: August 29–31, 2004, Daejon, Republic of Korea. International Statistical Review = Revue Internationale de Statistique, 73(2):151, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2005:BMa

[Ano05b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 73(1):??, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472645.

Anonymous:2005:BMd

[Ano05c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 73(2):??, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472672.

Anonymous:2005:BMb

[Ano05d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 73(3):??, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472684.

Anonymous:2005:FMa

[Ano05e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 73(1):??, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472634.

Anonymous:2005:FMd

[Ano05f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 73(2):??, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472647.

Anonymous:2005:FMb

[Ano05g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 73(3):??, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472675.

Anonymous:2005:SP

[Ano05h] Anonymous. Scientific program. International Statistical Review = Revue Internationale de Statistique, 73(2):154– 156, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2005:VI

[Ano05i] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 73(3): ??, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472674.

Anonymous:2006:ARW

[Ano06a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 74(3): 421, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2006:BMa

[Ano06b] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 74(1):??, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472695.

Anonymous:2006:BMb

[Ano06c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 74(2):??, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472707.

Anonymous:2006:BMc

[Ano06d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 74(3):??, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472718.

Anonymous:2006:FMa

[Ano06e] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 74(1):??, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472686.

Anonymous:2006:FMb

[Ano06f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 74(2):??, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472697.

Anonymous:2006:FMc

[Ano06g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 74(3):??, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472710.

Anonymous:2006:VI

[Ano06h] Anonymous. Volume information. International Statistical Review = Revue Internationale de Statistique, 74(3): ??, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472709.

Anonymous:2007:ARW

[Ano07a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 75(3): 408, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2007:FMa

[Ano07b] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 75(1):??, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508445.

Anonymous:2007:FMb

[Ano07c] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 75(2):??, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41508455.

Anonymous:2007:FMc

[Ano07d] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 75(3):??, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41509869.

Anonymous:2007:IAI

[Ano07e] Anonymous. Index: Annual index 75/x2007. International Statistical Review = Revue Internationale de Statistique, 75(3):406–407, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2008:ARW

[Ano08a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 76(3): 435, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2008:AIA

[Ano08b] Anonymous. Annual index 76/2008. International Statistical Review = Revue Internationale de Statistique, 76 (3):433–434, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2008:BMa

[Ano08c] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 76(1):??, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919606.

Anonymous:2008:BMb

[Ano08d] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 76(2):??, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919647.

Anonymous:2008:BMc

[Ano08e] Anonymous. Back matter. International Statistical Review = Revue Internationale de Statistique, 76(3):??, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919683.

Anonymous:2008:FMa

[Ano08f] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 76(1):??, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919580.

Anonymous:2008:FMb

[Ano08g] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 76(2):??, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919608.

Anonymous:2008:FMc

[Ano08h] Anonymous. Front matter. International Statistical Review = Revue Internationale de Statistique, 76(3):??, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919649.

Anonymous:2009:ARW

[Ano09a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 77(3): 462, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2009:AIA

[Ano09b] Anonymous. Annual index 77/2009. International Statistical Review = Revue Internationale de Statistique, 77 (3):460–461, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2010:ARW

[Ano10a] Anonymous. Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 78(3):442, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2010:AIA

[Ano10b] Anonymous. Annual index 78/2010. International Statistical Review = Revue Internationale de Statistique, 78 (3):443–444, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2011:ARW

[Ano11a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 79(3): 477, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2011:AIA

[Ano11b] Anonymous. Annual index 79/2011. International Statistical Review = Revue Internationale de Statistique, 79 (3):478–479, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2012:AAR

[Ano12a] Anonymous. Acknowledgement: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 80(3):468, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2012:AIA

[Ano12b] Anonymous. Annual index 80/2012. International Statistical Review = Revue Internationale de Statistique, 80 (3):469–470, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2013:ARW

[Ano13a] Anonymous. Acknowledgement of refereeing work: Acknowledgement of refereeing work. International Statistical Review = Revue Internationale de Statistique, 81(3): 456, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2013:VI

[Ano13b] Anonymous. Volume indexp. International Statistical Review = Revue Internationale de Statistique, 81(3):457– 458, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anonymous:2021:EMS

[Ano21] Anonymous. Erratum: [Modern Strategies for Time Series Regression. Stephanie Clark, Rob J. Hyndman, Dan Pagendam, Louise M. Ryan. International Statistical Review (2020), 88, S1, S179–S204 doi:10.1111/insr.12432]. International Statistical Review = Revue Internationale de Statistique, 89(1):213, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [CHPR20].

Anonymous:2022:F

[Ano22] Anonymous. Foreword. International Statistical Review = Revue Internationale de Statistique, 90:S1, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Aoyama:1972:BRB

[Aoy72] H. Aoyama. Book review: Case Exercises in Operations Research by M. J. C. Martin; R. A. Denison. International Statistical Review = Revue Internationale de Statistique, 40(3):396–397, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402484.

Abdelbasit:1981:EDC

[AP81] K. M. Abdelbasit and R. L. Plackett. Experimental design for categorized data. International Statistical Review = Revue Internationale de Statistique, 49(2):111– 126, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403066.

Andrews:1994:IFC

[AP94] Donald W. K. Andrews and David Pollard. An introduction to functional central limit theorems for dependent stochastic processes. International Statistical Review = Revue Internationale de Statistique, 62(1):119–132, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403549.

Arbia:1993:UGS

[Arb93] G. Arbia. The use of GIS in spatial statistical surveys. International Statistical Review = Revue Internationale de Statistique, 61(2):339–359, August 1993. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403632.

Armitage:1985:RDA

[Arm85a] P. Armitage. Reply to discussion [‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage]. International Statistical Review = Revue Internationale de Statistique, 53(1):36, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402878.

Armitage:1985:SOC

[Arm85b] P. Armitage. The search for optimality in clinical trials. International Statistical Review = Revue Internationale de Statistique, 53(1):15–24, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402871.

Ahsanullah:1972:EIL

[AS72] M. Ahsanullah and A. K. Md. Ehsanes Saleh. Estimation of intercept in a linear regression model with one dependent variable after a preliminary test on the regression coefficient. International Statistical Review = Revue Internationale de Statistique, 40(2):139–145, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402756.

Avadhani:1973:CSE

[AS73] M. S. Avadhani and B. V. Sukhatme. Controlled sampling with equal probabilities and without replacement. International Statistical Review = Revue Internationale de Statistique, 41(2):175–182, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402834.

Astin:1994:SFF

[Ast94] J. A. Astin. Statistical form-filling by industry: Net burden or net benefit? International Statistical Review = Revue Internationale de Statistique, 62(1):87–97, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403547.

Andridge:2015:UFM

[AT15] Rebecca Andridge and Katherine Jenny Thompson. Using the fraction of missing information to identify auxiliary variables for imputation procedures via proxy patternmixture models. International Statistical Review = Revue Internationale de Statistique, 83(3):472–492, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Atkinson:1982:DDE

[Atk82] A. C. Atkinson. Developments in the design of experiments, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 50(2):161– 177, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402599.

Atkinson:1984:DPD

[Atk84] A. C. Atkinson. Discussion of paper by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 52(1):24–26, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403236.

Atkinson:1988:RDM

[Atk88] A. C. Atkinson. Recent developments in the methods of optimum and related experimental designs. International Statistical Review = Revue Internationale de Statistique, 56(2):99–115, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403635.

Aukrust:1978:JB

[Auk78] Odd Aukrust. Julie E. Backer, 1890–1977. International Statistical Review = Revue Internationale de Statistique, 46(3):325, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402393.

Auranen:2010:SBR

[Aur10] Kari Auranen. Short book review: Design and Analysis of Vaccine Studies by M. Elisabeth Halloran, Ira M. Longini, Jr., Claudio J. Struchiner. International Statistical Review = Revue Internationale de Statistique, 78(3): 462, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919878.

Austin:2017:TMS

[Aus17] Peter C. Austin. A tutorial on multilevel survival analysis: Methods, models and applications. International Statistical Review = Revue Internationale de Statistique, 85(2):

185–203, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Anderson:1995:MLE

[AW95] Christine M. Anderson and C. F. Jeff Wu. Measuring location effects from factorial experiments with a directional response. International Statistical Review = Revue Internationale de Statistique, 63(3):345–363, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403484.

Ayhan:2001:SGM

[Ayh01] H. Öztaş Ayhan. Statistics by gender: Measures to reduce gender bias in agricultural surveys. International Statistical Review = Revue Internationale de Statistique, 69(3):447–460, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403456.

Azouvi:2001:PMP

[Azo01] Alain Azouvi. Proposals for a minimum programme for statistics in developing countries. International Statistical Review = Revue Internationale de Statistique, 69(2):333– 343, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403819.

Baddeley:1982:SGI

[Bad82] Adrian Baddeley. Stochastic geometry: an introduction and reading-list. International Statistical Review = Revue Internationale de Statistique, 50(2):179–193, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402600.

Bailey:1984:DPT

[Bai84] R. A. Bailey. Discussion of paper by T. Tjur. International Statistical Review = Revue Internationale de Statistique, 52(1):65–77, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403243.

Bailar:1985:QIM

[Bai85a] Barbara A. Bailar. Quality issues in measurement. International Statistical Review = Revue Internationale de Statistique, 53(2):123–139, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402931.

Bailey:1985:RRV

[Bai85b] R. A. Bailey. Restricted randomization versus blocking. International Statistical Review = Revue Internationale de Statistique, 53(2):171–182, August 1985. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402933.

Bailar:1995:EPM

[Bai95] Barbara A. Bailar. [estimation in the presence of measurement error]: Discussion. International Statistical Review = Revue Internationale de Statistique, 63(2):142– 143, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403607.

Bakker:1973:BTS

[Bak73] T. Bakker. Bibliography on the teaching of statistics. International Statistical Review = Revue Internationale de Statistique, 41:381–386, 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Balaban:1976:BRBa

[Bal76a] V. Balaban. Book review: Uses of Statistical Data by S. S. Zarkovich. International Statistical Review = Revue Internationale de Statistique, 44(2):306–307, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403299.

Balaban:1976:BRBb

[Bal76b] V. Balaban. Book review: Uses of Statistical Data by S. S. Zarkovich. International Statistical Review = Revue Internationale de Statistique, 44(3):381, December 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402363.

Balk:1995:API

[Bal95] Bert M. Balk. Axiomatic price index theory: A survey. International Statistical Review = Revue Internationale de Statistique, 63(1):69–93, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403778.

Banerjee:2014:BRS

[Ban14] Moulinath Banerjee. Book review: Stochastic Modeling and Mathematical Statistics by Francisco J. Samaniego. International Statistical Review = Revue Internationale de Statistique, 82(3):481–483, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Barberi:1972:MV

[Bar72a] Benedetto Barberi. Mario de Vergottini, 1901–1971. International Statistical Review = Revue Internationale de Statistique, 40(1):109–110, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402723.

Bartholomew:1972:BRB

[Bar72b] D. J. Bartholomew. Book review: Sequential Tests of Statistical Hypotheses by B. K. Ghosh. International Statistical Review = Revue Internationale de Statistique, 40(2): 250, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402769.

Barton:1972:BRB

[Bar72c] D. E. Barton. Book review: An Author and Permuted Title Index to Selected Statistical Journals by B. L. Joiner; N. F. Laubscher; E. S. Brown; B. Levy. International Statistical Review = Revue Internationale de Statistique, 40(2):252–253, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402772.

Barberi:1975:RA

[Bar75a] B. Barberi. Raffaele D’Addario, 1899–1974. International Statistical Review = Revue Internationale de Statistique, 43(3):372–373, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403132.

Bartels:1975:BRBb

[Bar75b] Cornelis P. A. Bartels. Book review: An Econometric Study of the Fish Meal Industry by E. S. Leoncio. International Statistical Review = Revue Internationale de Statistique, 43(1):118–119, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402669.

Bartels:1975:BRBa

[Bar75c] Cornelis P. A. Bartels. Book review: Mathematical Sciences and Social Sciences by William Kruskal. International Statistical Review = Revue Internationale de Statistique, 43(1):117–118, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402668.

Bartholomew:1975:BRB

[Bar75d] D. J. Bartholomew. Book review: Linear Statistical Inferences and Its Applications by C. R. Rao. International Statistical Review = Revue Internationale de Statistique, 43(1):119–120, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402670.

Barton:1976:BRB

[Bar76] D. E. Barton. Book review: The Advanced Theory of Statistics; Vol. II, Inference and Relationship by M. G. Kendall; A. Stuart. International Statistical Review = Revue Internationale de Statistique, 44(1):167–169, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402971.

BarOn:1977:BRB

[Bar77a] R. R. V. BarOn. Book review: Time Series by M. G. Kendall. International Statistical Review = Revue Internationale de Statistique, 45(2):203–204, August 1977. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403324.

Barton:1977:BRBb

[Bar77b] D. E. Barton. Book review: Statistical Inference by S. D. Silvey. International Statistical Review = Revue Internationale de Statistique, 45(1):103, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403024.

Barton:1977:BRBa

[Bar77c] D. E. Barton. Book review: The Analysis of Categorical Data by R. L. Plackett. International Statistical Review = Revue Internationale de Statistique, 45(1):102, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403022.

Bartholomew:1980:BRBb

[Bar80a] D. J. Bartholomew. Book review: The Analysis of Cross-Tabulated Data by G. J. G. Upton. International Statistical Review = Revue Internationale de Statistique, 48(2): 244, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403184.

Bartholomew:1980:BRBa

[Bar80b] D. J. Bartholomew. Book review: The Scientific Use of Factor Analysis in Behavioural and Life Sciences by R. B. Cattell. International Statistical Review = Revue Internationale de Statistique, 48(2):232, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403162.

Barnard:1982:DPD

[Bar82] G. A. Barnard. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):11–14, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402449.

Bartholomew:1983:SRD

[Bar83] D. J. Bartholomew. Some recent developments in social statistics, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 51 (1):1–9, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402728.

Bartlett:1984:JI

[Bar84] M. S. Bartlett. J. O. Irwin, 1898–1982. International Statistical Review = Revue Internationale de Statistique, 52(1):109–114, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403250.

Barnard:1987:RFT

[Bar87] G. A. Barnard. R. A. Fisher — a true Bayesian? International Statistical Review = Revue Internationale de Statistique, 55(2):183–189, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403194.

Barnard:1995:PMF

[Bar95] George A. Barnard. Pivotal models and the fiducial argument. International Statistical Review = Revue Internationale de Statistique, 63(3):309–323, December 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403482.

Basu:2016:BRB

[Bas16] Sanjib Basu. Book review: Bayesian and Frequentist Regression Methods. International Statistical Review = Revue Internationale de Statistique, 84(1):159–161, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bather:1985:ATS

[Bat85a] J. A. Bather. On the allocation of treatments in sequential medical trials. International Statistical Review = Revue Internationale de Statistique, 53(1):1–13, April 1985. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402870.

Bather:1985:RDA

[Bat85b] J. A. Bather. Reply to discussion [‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage]. International Statistical Review = Revue Internationale de Statistique, 53(1):35–36, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402877.

Batanero:2002:DRM

[Bat02] Carmen Batanero. Discussion: The role of models in understanding and improving statistical literacy. International Statistical Review = Revue Internationale de Statistique, 70(1):37–40, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403717.

Bates:2007:SBR

[Bat07] Douglas M. Bates. Short book review: Direct Methods for Sparse Linear Systems by Timothy A. Davis. International Statistical Review = Revue Internationale de Statistique, 75(2):260–261, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508477.

Bates:2008:SBRa

[Bat08a] Douglas M. Bates. Short book review: Data Manipulation with R by Phil Spector. International Statistical Review = Revue Internationale de Statistique, 76(2):324– 325, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919643.

Bates:2008:SBRb

[Bat08b] Douglas M. Bates. Short book review: Probability and Statistics with R by Maria Dolores Ugarte, Ana F. Militino, Alan T. Arnholt. International Statistical Review = Revue Internationale de Statistique, 76(3):447–448, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919667.

Baumer:2014:BRA

[Bau14] Ben S. Baumer. Book review: Analyzing Baseball Data with R by Max Marchi, Jim Albert. International Statistical Review = Revue Internationale de Statistique, 82(2):

313–315, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Barnard:1976:DIC

[BBG^{+}76] G. A. Barnard, D. J. Bartholomew, V. P. Godambe, M. W. Hansen, L. Kish, R. K. C. Li, J. J. A. Moors, C. Scott, A. K. Shahani, D. Horvitz, J. Lanke, and L. Warner. Discussion on the invited and contributed papers. International Statistical Review = Revue Internationale de Statistique, 44(2):226–230, August 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403283.

Bowman:1977:DSU

[BBS77] K. O. Bowman, J. J. Beauchamp, and L. R. Shenton. The distribution of the t-statistic under non-normality. International Statistical Review = Revue Internationale de Statistique, 45(3):233–242, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402537.

Benzi:2007:FPC

[BBS07] Margherita Benzi, Michele Benzi, and Eugene Seneta. Francesco Paolo Cantelli: b. 20 December 1875 d. 21 July 1966. International Statistical Review = Revue Internationale de Statistique, 75(2):127–130, August 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41508456.

Buelens:2018:CIM

[BBvdB18] Bart Buelens, Joep Burger, and Jan A. van den Brakel. Comparing inference methods for non-probability samples. International Statistical Review = Revue Internationale de Statistique, 86(2):322–343, August 2018. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Breslow:1981:DPD

[BC81] Norman Breslow and John Crowley. Discussion of paper by D. Oakes. International Statistical Review = Revue Internationale de Statistique, 49(3):255–257, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402608.

Buhlmann:2020:DCR

[BC20] Peter Bühlmann and Domagoj ’Cevid. Deconfounding and causal regularisation for stability and external validity. International Statistical Review = Revue Internationale de Statistique, 88(S1):S114–S134, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Breckling:1994:MLI

[BCD^{+}94] J. U. Breckling, R. L. Chambers, A. H. Dorfman, S. M. Tam, and A. H. Welsh. Maximum likelihood inference from sample survey data. International Statistical Review = Revue Internationale de Statistique, 62(3):349– 363, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403766.

Boisbunon:2014:AIC

[BCF^{+}14] Aurélie Boisbunon, Stéphane Canu, Dominique Fourdrinier, William Strawderman, and Martin T. Wells. Akaike’s information criterion, C_{p} and estimators of loss for elliptically symmetric distributions. International Statistical Review = Revue Internationale de Statistique, 82 (3):422–439, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bono:2007:RIC

[BCG07] Filippa Bono, Miranda Cuffaro, and Rosa Giaimo. Regional inequalities in consumption patterns: A multilevel approach to the case of Italy. International Statistical Review = Revue Internationale de Statistique, 75(1):44–57, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41508448.

Barnard:1985:DAT

[BCL^{+}85] G. A. Barnard, Herman Chernoff, Thomas A. Louis, Richard Peto, and John Whitehead. Discussion of ‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage. International Statistical Review = Revue Internationale de Statistique, 53(1):25–26, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1402872. With replies by Bather and Armitage.

Baffour:2019:LED

[BCM19] Bernard Baffour, Hukum Chandra, and Arturo Martinez. Localised estimates of dynamics of multi-dimensional disadvantage: An application of the small area estimation technique using Australian survey and census data. International Statistical Review = Revue Internationale de Statistique, 87(1):1–23, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Boylan:1982:IDE

[BCO82] T. A. Boylan, M. P. Cuddy, and I. G. O’Muircheartaigh. Import demand equations: An application of a generalized Box–Cox methodology. International Statistical Review = Revue Internationale de Statistique, 50(1):103–112, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402461.

Bickel:1981:EFS

[BCvZ81] P. J. Bickel, D. M. Chibisov, and W. R. van Zwet. On efficiency of first and second order. International Statistical Review = Revue Internationale de Statistique, 49(2): 169–175, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403070.

Bouchard-Cote:2012:DBP

[BCZ12] Alexandre Bouchard-Côté and James V. Zidek. Discussion: “Bayesian priors from loss matching” [2990343]. International Statistical Review = Revue Internationale de Statistique, 80(1):83–86, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bourke:1976:SNI

[BD76] P. D. Bourke and T. Dalenius. Some new ideas in the realm of randomized inquiries. International Statistical Review = Revue Internationale de Statistique, 44(2):219– 221, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403280.

Blower:1994:SUA

[BD94] S. M. Blower and H. Dowlatabadi. Sensitivity and uncertainty analysis of complex models of disease transmission: An HIV model, as an example. International Statistical Review = Revue Internationale de Statistique, 62(2):229– 243, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403510.

Bellhouse:2009:MPA

[BD09] David R. Bellhouse and Matt Davison. De Moivre’s Poisson approximation to the binomial. International Statistical Review = Revue Internationale de Statistique, 77(3): 451–459, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919768.

Branson:2020:SBR

[BD20] Zach Branson and Tirthankar Dasgupta. Sampling-based randomised designs for causal inference under the potential outcomes framework. International Statistical Review = Revue Internationale de Statistique, 88(1):101– 121, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bhadra:2020:HRM

[BDLP20] Anindya Bhadra, Jyotishka Datta, Yunfan Li, and Nicholas Polson. Horseshoe regularisation for machine learning in complex and deep models1. International Statistical Review = Revue Internationale de Statistique, 88 (2):302–320, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Besbeas:2004:CSS

[BDS04] Panagiotis Besbeas, Italia De Feis, and Theofanis Sapatinas. A comparative simulation study of wavelet shrinkage estimators for Poisson counts. International Statistical Review = Revue Internationale de Statistique, 72(2): 209–237, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403855.

Bergman:1999:DUV

[BE99] Bo Bergman and Frederik Ekdahl. Discussion: Understanding variation. International Statistical Review = Revue Internationale de Statistique, 67(2):126–131, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403391.

Berger:2017:VEI

[BE17] Yves G. Berger and Emilio L. Escobar. Variance estimation of imputed estimators of change for repeated rotating surveys. International Statistical Review = Revue Internationale de Statistique, 85(3):421–438, December 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Beckmann:1973:BRB

[Bec73] M. Beckmann. Book review: The Evaluation of Risky Interrelated Investments by Frederick S. Hillier. International Statistical Review = Revue Internationale de Statistique, 41(3):399, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402636.

Beh:2004:SCA

[Beh04] Eric J. Beh. Simple correspondence analysis: A bibliographic review. International Statistical Review = Revue Internationale de Statistique, 72(2):257–284, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403857.

Bellhouse:1988:PSS

[Bel88] D. R. Bellhouse. Probability in the sixteenth and seventeenth centuries: An analysis of Puritan casuistry. International Statistical Review = Revue Internationale de Statistique, 56(1):63–74, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403362.

Bellhouse:1989:MCW

[Bel89] D. R. Bellhouse. A manuscript on chance written by John Arbuthnot. International Statistical Review = Revue Internationale de Statistique, 57(3):249–259, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403798.

Bellhouse:2000:BVM

[Bel00] D. R. Bellhouse. De Vetula: A medieval manuscript containing probability calculations. International Statistical Review = Revue Internationale de Statistique, 68(2):123– 136, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403664.

Belyaev:2002:SCE

[Bel02] Yuri Belyaev. [some cracks in the empire of chance (flaws in the foundations of reliability)]: Comment. International Statistical Review = Revue Internationale de Statistique, 70(1):65–67, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403723.

Bellhouse:2007:SBR

[Bel07] David Bellhouse. Short book review. International Statistical Review = Revue Internationale de Statistique, 75(1): 124–125, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bellhouse:2009:KPI

[Bel09] David R. Bellhouse. Karl Pearson’s influence in the United States. International Statistical Review = Revue Internationale de Statistique, 77(1):51–63, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919690.

Bellhouse:2016:BRJ

[Bel16] David Bellhouse. Book review: John Napier: Life, Logarithms and Legacy. International Statistical Review = Revue Internationale de Statistique, 84(1):161–162, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Benjamin:1975:BRB

[Ben75] B. Benjamin. Book review: Recall Lapse in Demographic Enquiries by R. K. Som. International Statistical Review = Revue Internationale de Statistique, 43(3):368– 369, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403125.

Berg:1972:DSC

[Ber72] Sven Berg. Double sampling for cluster- or two-stage sampling. International Statistical Review = Revue Internationale de Statistique, 40(1):13–22, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402701.

Berkson:1977:MEN

[Ber77] Joseph Berkson. My encounter with neo-Bayesianism. International Statistical Review = Revue Internationale de Statistique, 45(1):1–8, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402998.

Berg:1978:BRB

[Ber78] Sven Berg. Book review: Statistical Distributions in Scientific Work. Volume I: Models and Structures by G. P. Patil; S. Kotz; J. K. Ord. International Statistical Review = Revue Internationale de Statistique, 46(1):121– 122, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402525.

Bernardo:1982:DPD

[Ber82] Jose M. Bernardo. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):14–15, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402450.

Bernardo:2003:RFC

[Ber03] José M. Bernardo. [reflections on fourteen cryptic issues concerning the nature of statistical inference]: Discussion. International Statistical Review = Revue Internationale de Statistique, 71(2):307–314, August 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403889.

Bertino:2006:MRS

[Ber06] Salvatore Bertino. A measure of representativeness of a sample for inferential purposes. International Statistical Review = Revue Internationale de Statistique, 74(2):149– 159, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472699.

Berry:2014:CPS

[Ber14] Kenneth Berry. A Chronicle of Permutation Statistical Methods: 1920–2000, and Beyond. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-02743-3, 3-319-02744-1. xix + 1 + 517 pp. LCCN QA165 .B47 2014;

QA165 .C57 2014. URL http://www.loc.gov/catdir/
enhancements/fy1407/2014935885-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy1407/2014935885-
d.html;   http://www.loc.gov/catdir/enhancements/
fy1407/2014935885-t.html.

Bergomi:2016:SVM

[Ber16] Lorenzo Bergomi. Stochastic volatility modeling. Chapman and Hall/CRC financial mathematics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2016. ISBN 1-4822-4406-3. xvi + 506 pp. LCCN HG106 .B39 2016.

Beresewicz:2017:TSP

[Ber17] Maciej Beresewicz. A two-step procedure to measure representativeness of Internet data sources. International Statistical Review = Revue Internationale de Statistique, 85 (3):473–493, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bethlehem:2010:SBW

[Bet10] Jelke Bethlehem. Selection bias in Web surveys. International Statistical Review = Revue Internationale de Statistique, 78(2):161–188, August 2010. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919830.

Bertino:2003:SDA

[BEW03] Laurent Bertino, Geir Evensen, and Hans Wackernagel. Sequential data assimilation techniques in oceanography. International Statistical Review = Revue Internationale de Statistique, 71(2):223–241, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403885.

Boente:1991:SUC

[BF91] Graciela Boente and Ricardo Fraiman. Strong uniform convergence rates for some robust equivariant nonparametric regression estimates for mixing processes. International Statistical Review = Revue Internationale de Statistique, 59(3):355–372, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403692.

Bellhouse:1997:LC

[BF97] D. R. Bellhouse and J. Franklin. The language of chance. International Statistical Review = Revue Internationale de Statistique, 65(1):73–85, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403433.

Bianchi:2018:ETM

[BFST18] Annamaria Bianchi, Enrico Fabrizi, Nicola Salvati, and Nikos Tzavidis. Estimation and testing in m-quantile regression with applications to small area estimation. International Statistical Review = Revue Internationale de Statistique, 86(3):541–570, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Becker:1983:CCD

[BG83] Niels Becker and Ian Gordon. On Cox’s criterion for discriminating between alternative ancillary statistics. International Statistical Review = Revue Internationale de Statistique, 51(1):89–92, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402734.

Balding:2003:I

[BG03] David J. Balding and Joseph L. Gastwirth. Introduction. International Statistical Review = Revue Internationale de Statistique, 71(3):469–471, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403823.

Bourguignon:2024:PQB

[BGS24] Marcelo Bourguignon, Diego I. Gallardo, and Helton Saulo. Parametric quantile beta regression model. International Statistical Review = Revue Internationale de Statistique, 92(1):106–129, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Betti:2018:SJV

[BGV18] Gianni Betti, Francesca Gagliardi, and Vijay Verma. Simplified jackknife variance estimates for fuzzy measures of multidimensional poverty. International Statistical Review = Revue Internationale de Statistique, 86(1):68–86, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bancroft:1977:IBC

[BH77] T. A. Bancroft and Chien Pai Han. Inference based on conditional specification: a note and a bibliography. International Statistical Review = Revue Internationale de Statistique, 45(2):117–127, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403311.

Bancroft:1980:RES

[BH80] T. A. Bancroft and Chien Pai Han. On the reduction of erroneous statistical inferences due to incorrect specification of the model under analysis. International Statistical Review = Revue Internationale de Statistique, 48(3):309– 316, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402943.

Berglund:2014:MIM

[BH14] Patricia Berglund and Steven Heeringa. Multiple Imputation of Missing Data Using SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2014. ISBN 1-61290-452-1, 1-62959-203-X, 1-62959-205-6 (e-book). xii + 150 pp. LCCN HA31.2 .B37 2014.

Bhansali:1991:CRE

[Bha91] R. J. Bhansali. Consistent recursive estimation of the order of an autoregressive moving average process. International Statistical Review = Revue Internationale de Statistique, 59(1):81–96, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403576.

Bhansali:2007:SBR

[Bha07] R. J. Bhansali. Short book review: Long-Memory Time Series: Theory and Methods by Wilfredo Palma. International Statistical Review = Revue Internationale de Statistique, 75(2):271–272, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508486.

Berent:1972:SRT

[BHBP72] J. Berent, E. Hofsten, G. Beyer, and G. P. Patil. Statistical research and training / formation et recherche statistique. International Statistical Review = Revue Internationale de Statistique, 40(2):237–243, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402766.

Bickel:1992:IAS

[Bic92] Peter J. Bickel. Inference and auditing: The Stringer bound. International Statistical Review = Revue Internationale de Statistique, 60(2):197–209, August 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403650.

See correction [Ano93d].

Bickeboller:2000:ILA

[Bic00] Heike Bickeböller. Investigation of linkage and association: Issues on study design. International Statistical Review = Revue Internationale de Statistique, 68(1):75–81, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403558.

Bickel:2013:MOS

[Bic13] David R. Bickel. Minimax-optimal strength of statistical evidence for a composite alternative hypothesis. International Statistical Review = Revue Internationale de Statistique, 81(2):188–206, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bickel:2014:SSI

[Bic14] David R. Bickel. Small-scale inference: Empirical Bayes and confidence methods for as few as a single comparison. International Statistical Review = Revue Internationale de Statistique, 82(3):457–476, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Biemer:1995:EPM

[Bie95] Paul Biemer. [estimation in the presence of measurement error]: Discussion. International Statistical Review = Revue Internationale de Statistique, 63(2):143–147, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403608.

Biehler:1997:SLD

[Bie97] Rolf Biehler. Software for learning and for doing statistics. International Statistical Review = Revue Internationale de Statistique, 65(2):167–189, August 1997. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403342.

Biehler:1999:DLT

[Bie99] Rolf Biehler. Discussion: Learning to think statistically and to cope with variation. International Statistical Review = Revue Internationale de Statistique, 67(3):259– 262, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403704.

Biggeri:2005:DB

[Big05] L. Biggeri. Discussant: L. Biggeri. International Statistical Review = Revue Internationale de Statistique, 73(2): 249–250, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Billeter:1977:BRBa

[Bil77a] Ernst P. Billeter. Book review: Applied Multivariate Analysis and Experimental Design by N. K. Namboodiri; L. F. Carter; H. M. Blalock. International Statistical Review = Revue Internationale de Statistique, 45(1):100–101, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403019.

Billeter:1977:BRBb

[Bil77b] Ernst P. Billeter. Book review: Input-Output Analysis and Its Applications by R. O’Connor; E. W. Henry. International Statistical Review = Revue Internationale de Statistique, 45(1):101, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403020.

Billeter:1980:BRBa

[Bil80a] Ernst P. Billeter. Book review: Gli Indici di Omofilia by Giorgio Angelini. International Statistical Review = Revue Internationale de Statistique, 48(1):118, April 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402413.

Billeter:1980:BRBb

[Bil80b] Ernst P. Billeter. Book review: Information and Exponential Families in Statistical Theory by Ole Barndorff-Nielsen. International Statistical Review = Revue Internationale de Statistique, 48(2):229, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403156.

Binder:1983:VAN

[Bin83] David A. Binder. On the variances of asymptotically normal estimators from complex surveys. International Statistical Review = Revue Internationale de Statistique, 51 (3):279–292, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402588.

Binder:2000:RBS

[Bin00] David A. Binder. [the role of behavioral survey methodologists in national statistical agencies]: Discussion. International Statistical Review = Revue Internationale de Statistique, 68(2):213–215, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403671.

Bingham:2007:SBR

[Bin07] N. H. Bingham. Short book review: Statistical and Probabilistic Methods in Actuarial Science by Philip J. Boland. International Statistical Review = Revue Internationale de Statistique, 75(2):253, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508468.

Bjerve:1972:GJ

[Bje72] Petter Jakob Bjerve. Gunnar Jahn, 1883–1971. International Statistical Review = Revue Internationale de Statistique, 40(1):109, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402722.

Bjerke:1973:BRB

[Bje73] Kjeld Bjerke. Book review: The Design of Economic Accounts by Nancy Ruggles; Richard Ruggles. International Statistical Review = Revue Internationale de Statistique, 41(2):295–297, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402861.

Bellido:1998:MAP

[BJO^{+}98] N. P. Bellido, M. D. Jano, F. J. López Ortega, M. P. Martín-Guzmán, and M. I. Toledo. The measurement and analysis of poverty and inequality: An application to Spanish conurbations. International Statistical Review = Revue Internationale de Statistique, 66(1):115–131, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403659.

Boulanger:1999:EIS

[BJPT99] Michèle Boulanger, Mark Johnson, Christophe Perruchet, and Poul Thyregod. Evolution of international statistical standards via life cycle of products and services. International Statistical Review = Revue Internationale de Statistique, 67(2):151–171, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403396.

Bru:1992:JBF

[BJS92] Bernard Bru, François Jongmans, and Eugene Seneta. I. J. Bienaymé: Family information and proof of the criticality theorem. International Statistical Review = Revue Internationale de Statistique, 60(2):177–183, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403648.

Basawa:1988:LSS

[BK88] I. V. Basawa and H. L. Koul. Large-sample statistics based on quadratic dispersion. International Statistical Review = Revue Internationale de Statistique, 56(3):199– 219, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403349.

Bott:2016:AEC

[BK16] Ann-Kathrin Bott and Michael Kohler. Adaptive estimation of a conditional density. International Statistical Review = Revue Internationale de Statistique, 84(2):291– 316, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Berger:2020:ELA

[BK20] Yves G. Berger and Ewa Kabzińska. Empirical likelihood approach for aligning information from multiple surveys. International Statistical Review = Revue Internationale de Statistique, 88(1):54–74, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Brunner:2021:RPR

[BKBP21] Edgar Brunner, Frank Konietschke, Arne C. Bathke, and Markus Pauly. Ranks and pseudo-ranks-surprising results of certain rank tests in unbalanced designs. International Statistical Review = Revue Internationale de Statistique, 89(2):349–366, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Baffour:2013:MCE

[BKV13] Bernard Baffour, Thomas King, and Paolo Valente. The modern census: evolution, examples, and evaluation. International Statistical Review = Revue Internationale de Statistique, 81(3):407–425, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Beh:2024:CAU

[BL24] Eric J. Beh and Rosaria Lombardo. Correspondence analysis using the Cressie–Read family of divergence statistics. International Statistical Review = Revue Internationale de Statistique, 92(1):17–42, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Black:2007:SBR

[Bla07] Mik Black. Short book review: Bayesian Inference for Gene Expression and Proteomics edited by Kim-Anh Do, Peter Müller, Marina Vannucci. International Statistical Review = Revue Internationale de Statistique, 75(3):433– 434, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509905.

Blume:2005:HCW

[Blu05] Jeffrey D. Blume. How to choose a working model for measuring the statistical evidence about a regression parameter. International Statistical Review = Revue Internationale de Statistique, 73(3):351–363, December 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472680.

Blumberg:2007:SBR

[Blu07] Carol Joyce Blumberg. Short book review: Statistical Literacy at School: Growth and Goals by Jane M. Watson. International Statistical Review = Revue Internationale de Statistique, 75(2):259, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508475.

Blumberg:2016:BRC

[Blu16] Carol Joyce Blumberg. Book review: Causal Inference for Statistics, Social, and Biomedical Sciences: An Introduction. International Statistical Review = Revue Internationale de Statistique, 84(1):159, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Blumberg:2018:BRO

[Blu18] Carol Joyce Blumberg. Book review: Observation and Experiment: An Introduction to Causal Inference, Paul R. Rosenbaum, Harvard University Press, 2017, 374 pages, £25.95, hardcover ISBN: 978-0-674-97557-6. International Statistical Review = Revue Internationale de Statistique, 86(1):165–166, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Blumberg:2020:BRB

[Blu20] Carol Joyce Blumberg. Book review: Bayesian Statistics for Beginners: A Step-by-Step Approach. Therese M. Donovan and Ruth M. Mickey Oxford University Press, 2019, viii + 419 pages, $100, hardcover ISBN: 978-0-19-884129-6. International Statistical Review = Revue Internationale de Statistique, 88(2):514–515, August 2020. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Blumberg:2021:BRM

[Blu21] Carol Joyce Blumberg. Book review: Measurement Error in Longitudinal Data, Alexandru Cernat and Joseph W. Sakshaug, Editors Oxford University Press, 2021, xii + 448 pages, $115.00, hardcover ISBN: 978-0-19-885998-7. International Statistical Review = Revue Internationale de Statistique, 89(3):658–659, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Blyth:1994:KPC

[Bly94] Stephen Blyth. Karl Pearson and the correlation curve. International Statistical Review = Revue Internationale de Statistique, 62(3):393–403, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403769.

Baddeley:1989:NNM

[BM89] Adrian Baddeley and Jesper Møller. Nearest-neighbour Markov point processes and random sets. International Statistical Review = Revue Internationale de Statistique, 57(2):89–121, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403381.

Berger:1991:ISP

[BM91] James O. Berger and Julia Mortera. Interpreting the stars in precise hypothesis testing. International Statistical Review = Revue Internationale de Statistique, 59(3): 337–353, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403691.

Babu:2016:SLC

[BM16] G. Jogesh Babu and Ashish Mahabal. Skysurveys, light curves and statistical challenges. International Statistical Review = Revue Internationale de Statistique, 84(3):506– 527, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Braun:2008:ODA

[BMM08] Maria P. Braun, Simos G. Meintanis, and Viatcheslav B. Melas. Optimal design approach to GMM estimation of parameters based on empirical transforms. International Statistical Review = Revue Internationale de Statistique, 76(3):387–400, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919653.

Barndorff-Nielsen:2007:SBR

[BN07] Ole E. Barndorff-Nielsen. Short book review: Random Graph Dynamics by Rick Durrett. International Statistical Review = Revue Internationale de Statistique, 75(3): 428, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509897.

Barndorff-Nielsen:1984:ESR

[BNC84] O. E. Barndorff-Nielsen and D. R. Cox. The effect of sampling rules on likelihood statistics. International Statistical Review = Revue Internationale de Statistique, 52(3):309– 326, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403049.

Barndorff-Nielsen:1986:RDG

[BNCR86] O. E. Barndorff-Nielsen, D. R. Cox, and N. Reid. The role of differential geometry in statistical theory. International Statistical Review = Revue Internationale de Statistique, 54(1):83–96, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403260.

Barndorff-Nielsen:1982:NVM

[BNKS82] O. Barndorff-Nielsen, J. Kent, and M. Sørensen. Normal variance-mean mixtures and z distributions. International Statistical Review = Revue Internationale de Statistique, 50(2):145–159, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402598.

Barndorff-Nielsen:1994:RSA

[BNS94] O. E. Barndorff-Nielsen and M. Sørensen. A review of some aspects of asymptotic likelihood theory for stochastic processes. International Statistical Review = Revue Internationale de Statistique, 62(1):133–165, April 1994. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403550.

Boos:2015:AVC

[BO15] Dennis D. Boos and Jason A. Osborne. Assessing variability of complex descriptive statistics in Monte Carlo studies using resampling methods. International Statistical Review = Revue Internationale de Statistique, 83(2): 228–238, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bolstad:2007:SBR

[Bol07] Bill Bolstad. Short book review: Bayesian Process Monitoring, Control and Optimization edited by Enrique del Castillo, Bianca M. Colosimo. International Statistical Review = Revue Internationale de Statistique, 75(3):412, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41509883.

Bonnen:1996:CES

[Bon96] James T. Bonnen. [characteristics of an effective statistical system]: Discussion. International Statistical Review = Revue Internationale de Statistique, 64(2):193– 197, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403741.

Boomsma:1979:BRB

[Boo79] A. Boomsma. Book review: Statistical Techniques in Simulation by Jack P. C. Kleijnen. International Statistical Review = Revue Internationale de Statistique, 47 (1):90, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403220.

Borroni:2009:UKP

[Bor09] Claudio Giovanni Borroni. Understanding Karl Pearson’s influence on Italian statistics in the early 20th century. International Statistical Review = Revue Internationale de Statistique, 77(1):81–95, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919693.

Bowater:2004:FJW

[Bow04a] Russell J. Bowater. A foundational justification for a weighted likelihood approach to inference. International Statistical Review = Revue Internationale de Statistique, 72(3):307–330, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/25472625.

Bowater:2004:R

[Bow04b] Russell J. Bowater. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 72(3):335– 336, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Box:1975:SE

[Box75] George E. P. Box. Statistics and the environment. International Statistical Review = Revue Internationale de Statistique, 43(3):279–287, December 1975. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403112.

Box:1984:DPD

[Box84] George E. P. Box. Discussion of paper by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 52(1):26, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403237.

Box:1993:QIN

[Box93] George Box. Quality improvement: The new industrial revolution. International Statistical Review = Revue Internationale de Statistique, 61(1):3–19, April 1993. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403590.

Breiman:1992:CAS

[BP92] Leo Breiman and Steve Peters. Comparing automatic smoothers (a public service enterprise). International Statistical Review = Revue Internationale de Statistique, 60 (3):271–290, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403679.

Beaumont:2012:GBS

[BP12] Jean-François Beaumont and Zdenek Patak. On the generalized bootstrap for sample surveys with special attention to Poisson sampling. International Statistical Review = Revue Internationale de Statistique, 80(1):127–148, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257170.

Benedetti:2017:SBS

[BPP17] Roberto Benedetti, Federica Piersimoni, and Paolo Postiglione. Spatially balanced sampling: A review and a reappraisal. International Statistical Review = Revue Internationale de Statistique, 85(3):439–454, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Beiting-Parrish:2021:BRH

[BPV21] Magdalen Beiting-Parrish and Jay Verkuilen. Book review: Handbook of Automated Scoring: Theory into Practice, Duanli Yan, André A. Rupp and Peter W. Foltz Chapman and Hall/CRC, 2020, 580 pages, £120, hardcover ISBN: 978-1-138-57827-2. International Statistical Review = Revue Internationale de Statistique, 89(2):423– 426, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Balding:2000:SIM

[BR00] David Balding and Sylvia Richardson. Special issue: Modern statistical methods for disease gene mapping. International Statistical Review = Revue Internationale de Statistique, 68(1):1–3, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403553.

Bernardo:2002:BHT

[BR02] José M. Bernardo and Raúl Rueda. Bayesian hypothesis testing: A reference approach. International Statistical Review = Revue Internationale de Statistique, 70(3):351– 372, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403862.

Bhat:2007:TNM

[BR07] Kavitha Bhat and K. Aruna Rao. On tests for a normal mean with known coefficient of variation. International Statistical Review = Revue Internationale de Statistique, 75(2):170–182, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41508459.

Baillargeon:2009:GAU

[BR09] Sophie Baillargeon and Louis-Paul Rivest. A general algorithm for univariate stratification. International Statistical Review = Revue Internationale de Statistique, 77(3):331– 344, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919760.

Bradley:1975:STR

[Bra75] Ralph A. Bradley. Statistical training and research / formation et recherche statistique. International Statistical Review = Revue Internationale de Statistique, 43(3):355– 363, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403121.

Braun:2014:BRG

[Bra14] Thomas M. Braun. Book review: Genomic Clinical Trials and Predictive Medicine by Richard M. Simon. International Statistical Review = Revue Internationale de Statistique, 82(3):478–479, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Braun:2015:BRA

[Bra15] Thomas M. Braun. Book review: Adaptive Design Theory and Implementation Using SAS and R, Second Edition. International Statistical Review = Revue Internationale de Statistique, 83(3):512–513, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Breslow:1975:ASD

[Bre75] N. E. Breslow. Analysis of survival data under the proportional hazards model. International Statistical Review = Revue Internationale de Statistique, 43(1):45–57, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402659.

Breny:1979:BRBa

[Bre79a] H. Breny. Book review: Algebraic Topology: A First Course by M. K. Agoston. International Statistical Review = Revue Internationale de Statistique, 47(1):85, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403211.

Breny:1979:BRBb

[Bre79b] H. Breny. Book review: Measurement and Evaluation in Education and Psychology by William A. Mehrens. International Statistical Review = Revue Internationale de Statistique, 47(1):91, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403221.

Breny:1979:BRBc

[Bre79c] H. Breny. Book review: Minitab Student Handbook by T. A. Ryan; B. L. Joiner; B. F. Ryan. International Statistical Review = Revue Internationale de Statistique, 47 (1):92, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403224.

Breny:1980:BRB

[Bre80] H. Breny. Book reviews: Statistics and Econometrics. A Problem-Solving Text by B. R. Chiswick; S. J. Chiswick; Foundations of Econometrics by A. Madansky; Econometrics by P. Schmidt. International Statistical Review = Revue Internationale de Statistique, 48(3):373, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402958.

Breslow:1999:DST

[Bre99a] N. E. Breslow. Discussion: Statistical thinking in practice. International Statistical Review = Revue Internationale de Statistique, 67(3):252–255, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403702.

Brewer:1999:DBP

[Bre99b] K. R. W. Brewer. Design-based or prediction-based inference? stratified random vs stratified balanced sampling. International Statistical Review = Revue Internationale de Statistique, 67(1):35–47, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403564.

Bharucha-Reid:1976:UGS

[BRST76] A. T. Bharucha-Reid, T. Smith, and C. P. Tsokos. On the use of Girshick’s sequential test for epidemiological data. International Statistical Review = Revue Internationale de Statistique, 44(3):331–340, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402348.

Blaylock:1985:BCT

[BS85] James R. Blaylock and David M. Smallwood. Box–Cox transformations and a heteroscedastic error variance: Import demand equations revisited. International Statistical Review = Revue Internationale de Statistique, 53(1): 91–97, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402882.

Bas:1989:TPN

[BS89] E. Moreno Bas and J. A. Cano Sanchez. Testing a point null hypothesis: Asymptotic robust Bayesian analysis with respect to the priors given on a subsigma field. International Statistical Review = Revue Internationale de Statistique, 57(3):221–232, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403795.

Breiman:1992:SSE

[BS92] Leo Breiman and Philip Spector. Submodel selection and evaluation in regression. The X-random case. International Statistical Review = Revue Internationale de Statistique, 60(3):291–319, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403680.

Basmann:2003:CDS

[BS03] Robert L. Basmann and Daniel J. Slottje. Copyright damages and statistics. International Statistical Review = Revue Internationale de Statistique, 71(3):557–564, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403828.

Baldermann:2018:RSA

[BSS18] Claudia Baldermann, Nicola Salvati, and Timo Schmid. Robust small area estimation under spatial non-stationarity. International Statistical Review = Revue Internationale de Statistique, 86(1):136–159, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bernholtz:1975:ASM

[BT75] B. Bernholtz and P. Tan. Applications of the structural method in statistical inference. International Statistical Review = Revue Internationale de Statistique, 43(3):289– 300, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403113.

Bernholtz:1976:EAS

[BT76] B. Bernholtz and P. Tan. Errata: “Applications of the structural method in statistical inference” (Internat. Statist. Rev. 43 (1975), no. 3, 289–300). International Statistical Review = Revue Internationale de Statistique, 44 (2):300, 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bean:1980:DND

[BT80] Steven J. Bean and Chris P. Tsokos. Developments in nonparametric density estimation. International Statistical Review = Revue Internationale de Statistique, 48(3): 267–287, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402941.

Buckland:1974:BRB

[Buc74] W. R. Buckland. Book review: Distributions in Statistics: Continuous Multivariate Distributions by N. L. Johnson; S. Kotz. International Statistical Review = Revue Internationale de Statistique, 42(3):317, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402993.

Buckland:1976:BRBa

[Buc76a] W. R. Buckland. Book review: Selected Tables in Mathematical Statistics. Vol. 1 by H. Leon Harter; D. B. Owen. International Statistical Review = Revue Internationale de Statistique, 44(1):169–170, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402972.

Buckland:1976:BRBb

[Buc76b] W. R. Buckland. Book review: Selected Tables in Mathematical Statistics. Vol. 3 by H. Leon Harter; D. B. Owen.

International Statistical Review = Revue Internationale de Statistique, 44(3):378–379, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402359.

Buckland:1977:BRBb

[Buc77a] W. R. Buckland. Book review: Perspectives in Probability and Statistics. Papers in Honour of Mr S. Bartlett by J. Gani. International Statistical Review = Revue Internationale de Statistique, 45(3):323–324, December 1977. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402550.

Buckland:1977:BRBc

[Buc77b] W. R. Buckland. Book review: Selected Tables in Mathematical Statistics. Vol. IV by D. B. Owen; R. E. Odeh. International Statistical Review = Revue Internationale de Statistique, 45(3):326, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402555.

Buckland:1977:BRBd

[Buc77c] W. R. Buckland. Book review: Statistical Sampling for Accountants by T. M. F. Smith. International Statistical Review = Revue Internationale de Statistique, 45(3):326– 327, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402556.

Buckland:1977:BRBa

[Buc77d] W. R. Buckland. Book review: Statistics Made Relevant: A Casebook of Real Life Examples by P. Baum; E. M. Scheuer. International Statistical Review = Revue Internationale de Statistique, 45(3):321, December 1977. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402545.

Buckland:1978:BRB

[Buc78] W. R. Buckland. Book review: Selected Tables in Mathematical Statistics, Volume V by D. B. Owen; R. E. Odeh. International Statistical Review = Revue Internationale de Statistique, 46(2):229, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402822.

Buck:1979:BRB

[Buc79] Dominic Buck. Book review: Statistical Analysis: A Decision-Making Approach by Robert Parsons. International Statistical Review = Revue Internationale de Statistique, 47(3):299, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402651.

Buhlmann:1975:ONN

[Büh75] Hans Bühlmann. Obituary notice: Notice nécrologique: Walter Saxer, 1896–1974. International Statistical Review = Revue Internationale de Statistique, 43(1):123, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402674.

Buonaccorsi:2006:ETS

[Buo06] John Buonaccorsi. Estimation in two-stage models with heteroscedasticity. International Statistical Review = Revue Internationale de Statistique, 74(3):403–418, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472717.

Brown:2012:BPL

[BW12a] Philip J. Brown and Stephen G. Walker. Bayesian priors from loss matching. International Statistical Review = Revue Internationale de Statistique, 80(1):60–82, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257167.

Brown:2012:RBP

[BW12b] Philip J. Brown and Stephen G. Walker. Rejoinder: “Bayesian priors from loss matching” [2990343; MR2990344; MR2990345]. International Statistical Review = Revue Internationale de Statistique, 80(1):89–92, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Billard:2020:WTS

[BW20] Lynne Billard and Katherine K. Wallman. Women trailblazers in the statistical profession. International Statistical Review = Revue Internationale de Statistique, 88(2): 280–301, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cacoullos:1974:BRB

[Cac74] T. Cacoullos. Book review: Introduction to Probability Theory and Statistical Inference by H. J. Larson. International Statistical Review = Revue Internationale de Statistique, 42(1):108–109, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402691.

Campion:1977:HL

[Cam77] H. Campion. Hector Leak, 1887–1976. International Statistical Review = Revue Internationale de Statistique, 45 (3):328, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402559.

Cameron:1997:DCI

[Cam97] Murray Cameron. Discussion: Current influences of computing on statistics. International Statistical Review = Revue Internationale de Statistique, 65(3):277–280, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403369.

Campbell:2007:SBR

[Cam07] M. J. Campbell. Short book review: Fundamentals of Clinical Research: Bridging Medicine, Statistics and Operations by Antonella Bacchieri, Giovanni Della Cioppa. International Statistical Review = Revue Internationale de Statistique, 75(3):411–412, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509882.

Cantor:1995:WRL

[Can95] David Cantor. [what respondents learn from questionnaires: The survey interview and the logic of conversation]: Discussion. International Statistical Review = Revue Internationale de Statistique, 63(2):173–177, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403612.

Capaldi:2019:BRD

[Cap19] Mindy Capaldi. Book review: Data Science and Predictive Analytics: Biomedical and Health Applications Using R, Ivo D. Dinov, Springer, 2018, xxxiv + 832 pages, £89.99, hardcover ISBN: 978-3-319-72346-4. International Statistical Review = Revue Internationale de Statistique, 87(1): 181–182, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Carey:1993:MAR

[Car93] Michèle Boulanger Carey. Measurement assurance: Role of statistics and support from international statistical standards. International Statistical Review = Revue Internationale de Statistique, 61(1):27–40, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403592.

Carlson:2000:WSP

[Car00] Beverley A. Carlson. Women in the statistics profession: A status report. International Statistical Review = Revue Internationale de Statistique, 68(3):339–352, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403418.

Carson:2005:IFP

[Car05] Carol S. Carson. International finance: Pushing the statistical envelope to meet the challenges of globalization. International Statistical Review = Revue Internationale de Statistique, 73(2):157–160, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472649.

Casalis:1991:FEN

[Cas91] Muriel Casalis. Familles exponentielles naturelles sur R^{d} invariantes par un groupe. International Statistical Review = Revue Internationale de Statistique, 59(2):241– 262, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403445.

Castelletti:2020:BMS

[Cas20] Federico Castelletti. Bayesian model selection of Gaussian directed acyclic graph structures. International Statistical Review = Revue Internationale de Statistique, 88(3):752– 775, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Caussinus:1980:BRB

[Cau80] H. Caussinus. Book review: Disturbances in the Linear Model, Estimation and Hypothesis Testing by C. Dubbelman. International Statistical Review = Revue Internationale de Statistique, 48(2):235, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403168.

Caussinus:1986:DPG

[Cau86] Henri Caussinus. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):274–278, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403056.

Campostrini:2000:SLO

[CB00] Stefano Campostrini and Pierantonio Bellini. Sick-leave and other job-related health issues from labour force surveys data. International Statistical Review = Revue Internationale de Statistique, 68(2):189–198, August 2000. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403668.

Chambers:2022:NPS

[CBPAK22] Ray Chambers, Stephen Beare, Scott Peak, and Mohammed Al-Kalbani. Nudging a pseudo-science towards a science — the role of statistics in a rainfall enhancement trial in Oman. International Statistical Review = Revue Internationale de Statistique, 90(2):346–373, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chen:2008:DSS

[CC08] Jie Chen and Kuang-Chao Chang. Discovering statistically significant periodic gene expression. International Statistical Review = Revue Internationale de Statistique, 76(2):228–246, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919614.

Cholette:1994:BTS

[CD94] Pierre A. Cholette and Estela Bee Dagum. Benchmarking time series with autocorrelated survey errors. International Statistical Review = Revue Internationale de Statistique, 62(3):365–377, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403767.

Clerget-Darpoux:2000:UGC

[CD00] F. Clerget-Darpoux. Unraveling the genetic component of multifactorial diseases: Dream or reality. International Statistical Review = Revue Internationale de Statistique, 68(1):45–51, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403556.

Cardot:2018:OPC

[CD18] Hervé Cardot and David Degras. Online principal component analysis in high dimension: Which algorithm to choose? International Statistical Review = Revue Internationale de Statistique, 86(1):29–50, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chakraborty:2019:RHB

[CDM19] Adrijo Chakraborty, Gauri Sankar Datta, and Abhyuday Mandal. Robust hierarchical Bayes small area estimation for the nested error linear regression model. International Statistical Review = Revue Internationale de Statistique, 87(S1):S158–S176, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Calmon:2021:ECS

[CFL^{+}21] Wilson Calmon, Eduardo Ferioli, Davi Lettieri, Johann Soares, and Adrian Pizzinga. An extensive comparison of some well-established value at risk methods. International Statistical Review = Revue Internationale de Statistique, 89(1):148–166, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Carota:2022:ABS

[CFP22] Cinzia Carota, Maurizio Filippone, and Silvia Polettini. Assessing Bayesian semi-parametric log-linear models: an application to disclosure risk estimation. International Statistical Review = Revue Internationale de Statistique, 90(1):165–183, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Carfagna:2005:URS

[CG05] Elisabetta Carfagna and F. Javier Gallego. Using remote sensing for agricultural statistics. International Statistical Review = Revue Internationale de Statistique, 73(3):389– 404, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472682.

Chaouch:2012:UCS

[CG12] Mohamed Chaouch and Camelia Goga. Using complex surveys to estimate the L_{1}-median of a functional variable: Application to electricity load curves. International Statistical Review = Revue Internationale de Statistique, 80(1):40–59, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257166.

Chen:2023:YAN

[CG23] Wanfang Chen and Marc G. Genton. Are you all normal? It depends! International Statistical Review = Revue Internationale de Statistique, 91(1):114–139, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Collins:2013:MSS

[CH13] David H. Collins and Aparna V. Huzurbazar. Multi-state stochastic processes: a statistical flowgraph perspective. International Statistical Review = Revue Internationale de Statistique, 81(1):78–106, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chen:2019:RDD

[CH19] Sixia Chen and David Haziza. Recent developments in dealing with item non-response in surveys: a critical review. International Statistical Review = Revue Internationale de Statistique, 87(S1):S192–S218, May 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chambers:1980:BRB

[Cha80] John M. Chambers. Book reviews: Compstat 78: Proceedings in Computational Statistics by L. C. A. Corsten; J. Hermans; Compstat Lectures I: Lectures in Computational Statistics by H. Skarobis; P. P. Sint. International Statistical Review = Revue Internationale de Statistique, 48(2): 233–234, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403165.

Chang:1993:SRS

[Cha93] Ted Chang. Spherical regression and the statistics of tectonic plate reconstructions. International Statistical Review = Revue Internationale de Statistique, 61(2):299– 316, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403630.

Chang:2008:ADT

[Cha08] Mark Chang. Adaptive design theory and implementation using SAS and R. Chapman and Hall/CRC biostatistics series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-962-4. xxii + 418 pp. LCCN

R853.C55 C42 2008. URL http://www.loc.gov/catdir/
enhancements/fy0801/2007011412-d.html;   http:/
/www.loc.gov/catdir/toc/ecip0713/2007011412.html.

Chakraborty:2014:BRB

[Cha14] Sounak Chakraborty. Book review: Bayesian Essentials with R by Jean-Michel Marin and Christian P. Robert. International Statistical Review = Revue Internationale de Statistique, 82(3):488–490, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Charpentier:2015:CAS

[Cha15] Arthur Charpentier, editor. Computational Actuarial Science with R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-9259-1 (hardcover). xxxi + 618 pp. LCCN HG8781 .C637 2014.

Chambers:2016:ENR

[Cha16] Raymond Chambers. Editorial: Nalini Ravishanker joins Raymond Chambers as co-Editor-in-Chief of the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 84(1):1, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chacon:2020:MAS

[Cha20a] José E. Chacón. The modal age of statistics. International Statistical Review = Revue Internationale de Statistique, 88(1):122–141, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chakraborty:2020:SRM

[Cha20b] Bibhas Chakraborty. Statistical remedies for medical researchers. International Statistical Review = Revue Internationale de Statistique, 88(3):802–804, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chengrui:1985:CEQ

[Che85a] Li Chengrui. China’s experiences in the quality control of its 1982 population census. International Statistical Review = Revue Internationale de Statistique, 53(2):109– 121, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402930.

Chernoff:1985:DAT

[Che85b] Herman Chernoff. Discussion of ‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage. International Statistical Review = Revue Internationale de Statistique, 53(1):26–28, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402873.

Chen:2024:EGM

[Che24] Li-Pang Chen. Estimation of graphical models: An overview of selected topics. International Statistical Review = Revue Internationale de Statistique, 92(2):194– 245, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chipperfield:2022:SWA

[Chi22] James Chipperfield. Survey weighting after imperfect linkage to an administrative file. International Statistical Review = Revue Internationale de Statistique, 90(3):419– 436, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chmielewski:1981:ESD

[Chm81] M. A. Chmielewski. Elliptically symmetric distributions: a review and bibliography. International Statistical Review = Revue Internationale de Statistique, 49(1):67–74, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403038.

Clark:2020:MST

[CHPR20] Stephanie Clark, Rob J. Hyndman, Dan Pagendam, and Louise M. Ryan. Modern strategies for time series regression. International Statistical Review = Revue Internationale de Statistique, 88(S1):S179–S204, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See erratum [Ano21].

Chipperfield:2018:EPR

[CHR18] James Chipperfield, Noel Hansen, and Peter Rossiter. Estimating precision and recall for deterministic and probabilistic record linkage. International Statistical Review = Revue Internationale de Statistique, 86(2):219–236, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chu:1974:URM

[Chu74] S. F. Chu. On the use of regression method in estimating regional population. International Statistical Review = Revue Internationale de Statistique, 42(1):17–28, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402682.

Chiou:2019:SRA

[CHXY19] Sy Han Chiou, Chiung-Yu Huang, Gongjun Xu, and Jun Yan. Semiparametric regression analysis of panel count data: a practical review. International Statistical Review = Revue Internationale de Statistique, 87(1):24–43, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ciampi:2014:DFY

[Cia14] Antonio Ciampi. Discussions on fifty years of classification and regression trees: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(3):352– 357, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Clements:2006:SIN

[CIS06] Kenneth W. Clements, Izan H. Y. Izan, and E. Antony Selvanathan. Stochastic index numbers: A review. International Statistical Review = Revue Internationale de Statistique, 74(2):235–270, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472705.

Citro:2016:ICK

[Cit16] Constance F. Citro. Interview: a conversation with Katherine K. Wallman. International Statistical Review = Revue Internationale de Statistique, 84(2):210–223, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2019:ALN

[CK19] David R. Cox and Christiana Kartsonaki. On the analysis of large numbers of p-values. International Statistical Review = Revue Internationale de Statistique, 87(3):505– 513, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2011:R

[CKK11a] Lawrence H. Cox, Alan F. Karr, and Satkartar K. Kinney. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 79(2):197–199, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2011:RUP

[CKK^{+}11b] Lawrence H. Cox, Alan F. Karr, Satkartar K. Kinney, Josep Domingo-Ferrer, George T. Duncan, Christine M. O’Keefe, and Natalie Shlomo. Risk-utility paradigms for statistical disclosure limitation: How to think, but not how to act [with discussions]. International Statistical Review = Revue Internationale de Statistique, 79(2):160– 183, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305021.

Chen:2014:BPM

[CKL14] Ming-Hui Chen, Lynn Kuo, and Paul O. Lewis, editors. Bayesian Phylogenetics: Methods, Algorithms, and Applications. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-0079-4 (hardcover). xxx + 365 pp. LCCN QH367.5 .B39 2014.

Chau:2024:IFN

[CKN^{+}24] Huy N. Chau, J. Lars Kirkby, Dang H. Nguyen, Duy Nguyen, Nhu N. Nguyen, and Thai Nguyen. On the inversion-free Newton’s method and its applications. International Statistical Review = Revue Internationale de Statistique, 92(2):284–321, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Clark:2017:CTR

[CKS17] Robert Graham Clark, Philip Kokic, and Paul A. Smith. A comparison of two robust estimation methods for business surveys. International Statistical Review = Revue Internationale de Statistique, 85(2):270–289, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chen:2019:SAQ

[CL19] Jiahua Chen and Yukun Liu. Small area quantile estimation. International Statistical Review = Revue Internationale de Statistique, 87(S1):S219–S238, May 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Clark:2000:RBS

[Cla00a] Cynthia Z. F. Clark. [the role of behavioral survey methodologists in national statistical agencies]: Discussion. International Statistical Review = Revue Internationale de Statistique, 68(2):217–220, August 2000. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403673.

Clayton:2000:LDM

[Cla00b] David Clayton. Linkage disequilibrium mapping of disease susceptibility genes in human populations. International Statistical Review = Revue Internationale de Statistique, 68(1):23–43, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403555.

Cleveland:2001:DSA

[Cle01] William S. Cleveland. Data science: An action plan for expanding the technical areas of the field of statistics. International Statistical Review = Revue Internationale de Statistique, 69(1):21–26, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403527.

Cleveland:2005:LDU

[Cle05] William S. Cleveland. Learning from data: Unifying statistics and computer science. International Statistical Review = Revue Internationale de Statistique, 73(2):217– 221, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472660.

Clogg:1986:DPG

[Clo86] Clifford C. Clogg. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):284–288, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403058.

Cordero:2006:SAA

[CM06] Maria M. Febres Cordero and Bernardo Márquez. A statistical approach to assess referendum results: the Venezuelan Recall Referendum 2004. International Statistical Review = Revue Internationale de Statistique, 74(3):379– 389, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472715.

Cicchitelli:2012:MAE

[CM12] Giuseppe Cicchitelli and Giorgio E. Montanari. Modelassisted estimation of a spatial population mean. International Statistical Review = Revue Internationale de Statistique, 80(1):111–126, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/23257169.

Cho:2019:CIA

[CMH19] Hunyong Cho, Gregory J. Matthews, and Ofer Harel. Confidence intervals for the area under the receiver operating characteristic curve in the presence of ignorable missing data. International Statistical Review = Revue Internationale de Statistique, 87(1):152–177, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Coeurjolly:2017:TPD

[CMW17] Jean-François Coeurjolly, Jesper Møller, and Rasmus Waagepetersen. A tutorial on palm distributions for spatial point processes. International Statistical Review = Revue Internationale de Statistique, 85(3):404–420, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chinganda:1998:TEA

[CN98] Enock F. Ching’anda and James P. M. Ntozi. Training employees for African statistical offices. International Statistical Review = Revue Internationale de Statistique, 66(2): 235–244, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403492.

Casady:1985:ADS

[CNS85] Robert J. Casady, Gad Nathan, and Monroe G. Sirken. Alternative dual system network estimators. International Statistical Review = Revue Internationale de Statistique, 53(2):183–197, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402934.

Chipperfield:2014:DPI

[CO14] James O. Chipperfield and Christine M. O’Keefe. Disclosure-protected inference using generalised linear models. International Statistical Review = Revue Internationale de Statistique, 82(3):371–391, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cobb:2007:OPF

[Cob07] George W. Cobb. One possible frame for thinking about experiential learning. International Statistical Review = Revue Internationale de Statistique, 75(3):336–347, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41509874.

Cochran:1973:BRB

[Coc73] W. G. Cochran. Book review: Statistical Papers of George Udny Yule by A. Stuart; M. G. Kendall; George Udny Yule. International Statistical Review = Revue Internationale de Statistique, 41(1):143–144, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402803.

Cochran:1979:GMC

[Coc79] William G. Cochran. Gertrude Mary Cox, 1900–1978. International Statistical Review = Revue Internationale de Statistique, 47(1):97–98, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403231.

Cohen:1975:BRB

[Coh75] J. W. Cohen. Book review: Elements of Applied Stochastic Processes by U. N. Bhat. International Statistical Review = Revue Internationale de Statistique, 43(1):116, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402665.

Cohen:2005:SSC

[Coh05] Michael P. Cohen. Sample size considerations for multilevel surveys. International Statistical Review = Revue Internationale de Statistique, 73(3):279–287, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472676.

Colombo:1972:AU

[Col72] Bernardo Colombo. Albino Uggè, 1899–1971. International Statistical Review = Revue Internationale de Statistique, 40(3):398, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402485.

Collomb:1981:ENP

[Col81] Gérard Collomb. Estimation non-paramétrique de la régression: Revue bibliographique. (French) [Nonparametric estimation of regression: bibliographical review]. International Statistical Review = Revue Internationale de Statistique, 49(1):75–93, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403039.

Colledge:1999:SIT

[Col99] Michael J. Colledge. Statistical integration through metadata management. International Statistical Review = Revue Internationale de Statistique, 67(1):79–98, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403567.

Consul:1991:ES

[Con91] P. C. Consul. Evolution of surnames. International Statistical Review = Revue Internationale de Statistique, 59 (3):271–278, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403687.

Congdon:2009:LES

[Con09] Peter Congdon. Life expectancies for small areas: A Bayesian random effects methodology. International Statistical Review = Revue Internationale de Statistique, 77 (2):222–240, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919724.

Cook:2007:SBR

[Coo07] Richard J. Cook. Short book review: Models for Intensive Longitudinal Data edited by Theodore A. Walls, Joseph L. Schafer. International Statistical Review = Revue Internationale de Statistique, 75(2):274–275, August 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41508489.

Copas:1977:BRB

[Cop77] J. B. Copas. Book review: Inference and Decision, Vol. 1, Selecta Statistica Canadiana (1973) by D. A. S. Fraser. International Statistical Review = Revue Internationale de Statistique, 45(2):202, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403321.

Cormack:1975:BRBa

[Cor75a] R. M. Cormack. Book review: Introduction to Biostatistics by R. R. Sokal; F. J. Rohlf. International Statistical Review = Revue Internationale de Statistique, 43(2): 250, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402925.

Cormack:1975:BRBb

[Cor75b] R. M. Cormack. Book review: Stochastic Processes and Applications in Biology and Medicine by M. Iosifescu; P. Tautu. International Statistical Review = Revue Internationale de Statistique, 43(3):368, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403124.

Cormack:1977:BRBc

[Cor77a] R. M. Cormack. Book review: Basic Ideas of Statistics by Bernard W. Lindgren. International Statistical Review = Revue Internationale de Statistique, 45(2):204– 205, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403326.

Cormack:1977:BRBb

[Cor77b] R. M. Cormack. Book review: Introduction to the Theory of Statistics by Harold J. Larson. International Statistical Review = Revue Internationale de Statistique, 45(2): 204, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403325.

Cormack:1977:BRBa

[Cor77c] R. M. Cormack. Book review: Statistics for Comparative Studies by Michael Hills. International Statistical Review = Revue Internationale de Statistique, 45(2):202– 203, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403322.

Cor:1980:BRB

[Cor80] D. R. Cor. Book review: Applications of Statistics by P. R. Krishnaiah. International Statistical Review = Revue Internationale de Statistique, 48(1):126, April 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402429.

Corder:2014:NSS

[Cor14] Gregory W. Corder. Nonparametric Statistics: A Step-by-Step Approach. John Wiley, New York, NY, USA, 2014. ISBN 1-118-84031-3 (hardcover), 1-118-84040-2 (e-book). LCCN QA278.8.

Cox:1975:BRB

[Cox75] D. R. Cox. Book review: The Estimation of Animal Abundance by G. A. F. Seber. International Statistical Review = Revue Internationale de Statistique, 43(2):249, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402924.

Cox:1980:BRB

[Cox80] D. R. Cox. Book review: The Theory of Competing Risks by H. A. David; M. L. Moeschberger. International Statistical Review = Revue Internationale de Statistique, 48(2): 234–235, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403167.

Cox:1981:DPD

[Cox81] D. R. Cox. Discussion of paper by D. Oakes. International Statistical Review = Revue Internationale de Statistique, 49(3):258, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402609.

Cox:1984:I

[Cox84a] D. R. Cox. Interaction. International Statistical Review = Revue Internationale de Statistique, 52(1):1–24, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1403235. With discussion and a reply by the author.

Cox:1984:IRD

[Cox84b] D. R. Cox. [interaction]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 52(1):30–31, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403241.

Cox:1986:SGA

[Cox86] D. R. Cox. Some general aspects of the theory of statistics. International Statistical Review = Revue Internationale de Statistique, 54(2):117–126, August 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403138.

Cox:1997:CPSa

[Cox97a] D. R. Cox. The current position of statistics: A personal view. International Statistical Review = Revue Internationale de Statistique, 65(3):261–276, December 1997. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403368.

Cox:1997:CPSb

[Cox97b] David Cox. [the current position of statistics: A personal view]: Response. International Statistical Review = Revue Internationale de Statistique, 65(3):289–290, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403372.

Cox:2002:SCE

[Cox02] D. R. Cox. [some cracks in the empire of chance (flaws in the foundations of reliability)]: Comment. International Statistical Review = Revue Internationale de Statistique, 70(1):67–68, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403724.

Cox:2003:RFC

[Cox03] D. R. Cox. [reflections on fourteen cryptic issues concerning the nature of statistical inference]: Rejoinder by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 71(2):318, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403891.

Cox:2007:SBR

[Cox07] D. R. Cox. Short book review: Response Surfaces, Mixtures, and Ridge Analyses, 2nd Edition by George E. P. Box, Norman R. Draper. International Statistical Review = Revue Internationale de Statistique, 75(2):265– 266, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508480.

Cox:2009:RDE

[Cox09] D. R. Cox. Randomization in the design of experiments. International Statistical Review = Revue Internationale de Statistique, 77(3):415–429, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919766.

Cox:2010:SBR

[Cox10] David Cox. Short book review: Steps Towards a Unified Basis for Scientific Models and Methods by Inge S. Helland. International Statistical Review = Revue Internationale de Statistique, 78(2):319, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919844.

Cox:2012:SBR

[Cox12] David Cox. Short book review: Selected Works of David Brillinger edited by Peter Guttorp and David Brillinger. International Statistical Review = Revue Internationale de Statistique, 80(2):326–327, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2013:DCD

[Cox13] David R. Cox. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1): 40–41, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2015:ICJ

[Cox15] David R. Cox. Interview: a conversation with John C. Gower. International Statistical Review = Revue Internationale de Statistique, 83(3):339–356, December 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2016:BRE

[Cox16] D. R. Cox. Book review: Estimation and Testing under Sparsity, Sara van de Geer, Springer, 2016, xiii + 274 pages, $59.99, softback ISBN: 978-3-319-32773-0. International Statistical Review = Revue Internationale de Statistique, 84(3):550, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2019:BRF

[Cox19] D. R. Cox. Book review: Foundations of Info-Metrics, Amos Golan, Oxford University Press, 2018, xx + 465 pages, $99.00, hardcover, ISBN: 978-0-19-934953-1. International Statistical Review = Revue Internationale de Statistique, 87(2):443, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cox:2020:DPB

[Cox20] D. R. Cox. Discussion of paper by Brad Efron. International Statistical Review = Revue Internationale de Statistique, 88(S1):S64, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Cacoullos:1980:SBP

[CP80] T. Cacoullos and H. Papageorgiou. On some bivariate probability models applicable to traffic accidents and fatalities. International Statistical Review = Revue Internationale de Statistique, 48(3):345–356, December 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402946.

Coles:1996:BME

[CP96] Stuart G. Coles and Elwyn A. Powell. Bayesian methods in extreme value modelling: A review and new developments. International Statistical Review = Revue Internationale de Statistique, 64(1):119–136, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403426.

Curto:2007:NMI

[CP07] José Dias Curto and José Castro Pinto. New multicollinearity indicators in linear regression models. International Statistical Review = Revue Internationale de Statistique, 75(1):114–121, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41508452.

Curto:2008:CNN

[CP08] José Dias Curto and José Castro Pinto. Correction note on new multicollinearity indicators. International Statistical Review = Revue Internationale de Statistique, 76(2):298– 299, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919618.

Christofides:2019:MTU

[CP19] Tasos C. Christofides and Pier Francesco Perri. Mixture of truthful-untruthful responses in public surveys. International Statistical Review = Revue Internationale de Statistique, 87(3):557–579, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cordeiro:1994:ILR

[CPB94] Gauss M. Cordeiro, Gilberto A. Paula, and Denise A. Botter. Improved likelihood ratio tests for dispersion models. International Statistical Review = Revue Internationale de Statistique, 62(2):257–274, August 1994. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403512.

Chagas:2022:WNT

[CQOR^{+}22] Eduarda T. C. Chagas, Marcelo Queiroz-Oliveira, Osvaldo A. Rosso, Heitor S. Ramos, Cristopher G. S. Freitas, and Alejandro C. Frery. White noise test from ordinal patterns in the entropy-complexity plane. International Statistical Review = Revue Internationale de Statistique, 90 (2):374–396, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Cressie:1989:PLR

[CR89] Noel Cressie and Timothy R. C. Read. Pearson’s χ^{2} and the loglikelihood ratio statistic G^{2}: A comparative review. International Statistical Review = Revue Internationale de Statistique, 57(1):19–43, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403582.

Carlsson:1993:QSU

[CR93] Olle Carlsson and Sture Rydin. Quality selection under sampling inspection with an exponential production cost function. International Statistical Review = Revue Internationale de Statistique, 61(1):109–119, April 1993. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403597.

Cramer:1973:BRB

[Cra73] J. S. Cramer. Book review: Mathematical Model Building in Economics and Industry by M. G. Kendall. International Statistical Review = Revue Internationale de Statistique, 41(2):293, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402856.

Cramer:1980:BRB

[Cra80] J. S. Cramer. Book review: The Price Index by S. N. Afriat. International Statistical Review = Revue Internationale de Statistique, 48(1):117–118, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402412.

Cramer:1981:MPS

[Cra81] Harald Cramér. Mathematical probability and statistical inference: Some personal recollections from an important phase of scientific development. International Statistical Review = Revue Internationale de Statistique, 49(3):309– 317, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor. org/stable/1402615. Some personal recollections from an important phase of scientific development.

Cressie:2020:CWI

[Cre20] Noel Cressie. Comment: When is it data science and when is it data engineering? International Statistical Review = Revue Internationale de Statistique, 88(S1):S65–S69, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Crowder:1994:ICC

[Cro94] Martin Crowder. Identifiability crises in competing risks. International Statistical Review = Revue Internationale de Statistique, 62(3):379–391, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403768.

Crowder:2007:SBRa

[Cro07a] M. J. Crowder. Short book review: Applied Asymptotics: Case Studies in Small-Sample Statistics by A. R. Brazzale, A. C. Davison, N. Reid. International Statistical Review = Revue Internationale de Statistique, 75(3):424– 425, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509894.

Crowder:2007:SBRb

[Cro07b] M. J. Crowder. Short book review: Parameter Estimation for Scientists and Engineers by Adriaan van den Bos. International Statistical Review = Revue Internationale de Statistique, 75(3):436–437, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509908.

Crowder:2008:SBRa

[Cro08a] M. J. Crowder. Short book review: Statistical Genetics of Quantitative Traits: Linkage, Maps and QTL by Rongling Wu, Chang-Xing Ma, George Casella. International Statistical Review = Revue Internationale de Statistique, 76 (1):145–146, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919595.

Crowder:2008:SBRb

[Cro08b] Martin Crowder. Short book review: Life Distributions: Structure of Nonparametric, Semiparametric, and Parametric Families by Albert W. Marshall, Ingram Olkin. International Statistical Review = Revue Internationale de Statistique, 76(2):303–304, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919623.

Crowder:2008:SBRc

[Cro08c] Martin Crowder. Short book review: Multiple Testing Procedures with Applications to Genomics by Sandrine Dudoit, Mark J. van der Laan. International Statistical Review = Revue Internationale de Statistique, 76(2):309– 310, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919629.

Crowder:2008:SBRe

[Cro08d] Martin Crowder. Short book review: Proportional Hazards Regression by John O’Quigley. International Statistical Review = Revue Internationale de Statistique, 76(3): 451, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919671.

Crowder:2008:SBRd

[Cro08e] Martin Crowder. Short book review: Survival Analysis for Epidemiologic and Medical Research by Steve Selvin. International Statistical Review = Revue Internationale de Statistique, 76(3):444, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919664.

Crowder:2009:SBRb

[Cro09a] Martin Crowder. Short book review: Cluster Randomised Trials by Richard J. Hayes, Lawrence H. Moulton. International Statistical Review = Revue Internationale de Statistique, 77(2):305, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919734.

Crowder:2009:SBRa

[Cro09b] Martin Crowder. Short book review: Longitudinal Data Analysis edited by Garrett Fitzmaurice, Marie Davidian, Geert Verbeke, Geert Molenberghs. International Statistical Review = Revue Internationale de Statistique, 77(1): 148–149, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919697.

Crowder:2009:SBRe

[Cro09c] Martin Crowder. Short book review: Self-Normalized Processes: Limit Theory and Statistical Applications by Victor H. de la Peña, Tze Leung Lai, Qi-Man Shao. International Statistical Review = Revue Internationale de Statistique, 77(3):467–468, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919772.

Crowder:2009:SBRd

[Cro09d] Martin Crowder. Short book review: Semi-Markov Chains and Hidden Semi-Markov Models Toward Applications by Vlad Stefan Barbu, Nikolaos Limnios. International Statistical Review = Revue Internationale de Statistique, 77 (2):307, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919736.

Crowder:2009:SBRc

[Cro09e] Martin Crowder. Short book review: Stochastic Approximation: A Dynamical Systems Viewpoint by Vivek S. Borkar. International Statistical Review = Revue Internationale de Statistique, 77(2):306, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919735.

Crowder:2010:SBRe

[Cro10a] Martin Crowder. Short book review: An Intermediate Course in Probability, Second Edition, by Allan Gut. International Statistical Review = Revue Internationale de Statistique, 78(1):154, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919820.

Crowder:2010:SBRc

[Cro10b] Martin Crowder. Short book review: Analysis of Microdata, Second Edition, by Rainer Winkelmann, Stefan Boes. International Statistical Review = Revue Internationale de Statistique, 78(1):143, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919808.

Crowder:2010:SBRh

[Cro10c] Martin Crowder. Short book review: Applied Survey Data Analysis by Steven G. Heeringa, Brady T. West, Patricia A. Berglund. International Statistical Review = Revue Internationale de Statistique, 78(3):463, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919879.

Crowder:2010:SBRd

[Cro10d] Martin Crowder. Short book review: Basic Concepts of Probability and Statistics in the Law by Michael O. Finkelstein. International Statistical Review = Revue Internationale de Statistique, 78(1):153–154, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919819.

Crowder:2010:SBRb

[Cro10e] Martin Crowder. Short book review: Deconvolution Problems in Nonparametric Statistics by Alexander Meister. International Statistical Review = Revue Internationale de Statistique, 78(1):142–143, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919807.

Crowder:2010:SBRg

[Cro10f] Martin Crowder. Short book review: Design and Analysis of Quality of Life Studies in Clinical Trials, Second Edition, by Diane L. Fairclough. International Statistical Review = Revue Internationale de Statistique, 78(3):454– 455, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919871.

Crowder:2010:SBRi

[Cro10g] Martin Crowder. Short book review: Generalized Linear Models: with Applications in Engineering and the Sciences, Second Edition, by Raymond H. Myers, Douglas C. Montgomery, G. Geoffrey Vining, Timothy J. Robinson. International Statistical Review = Revue Internationale de Statistique, 78(3):472–473, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919888.

Crowder:2010:SBRf

[Cro10h] Martin Crowder. Short book review: Measurement Error: Models, Methods, and Applications by John P. Buonaccorsi. International Statistical Review = Revue Internationale de Statistique, 78(3):447, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919865.

Crowder:2010:SBRj

[Cro10i] Martin Crowder. Short book review: Mixed Effects Models for Complex Data by Lang Wu. International Statistical Review = Revue Internationale de Statistique, 78(3): 480, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919895.

Crowder:2010:SBRa

[Cro10j] Martin Crowder. Short book review: Statistical Methods in Environmental Epidemiology by Duncan C. Thomas. International Statistical Review = Revue Internationale de Statistique, 78(1):137–138, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919803.

Crowder:2011:SBRe

[Cro11a] Martin Crowder. Short book review: Bayesian Nonparametrics edited by Nils Lid Hjort, Chris Holmes, Peter Müller, Stephen G. Walker. International Statistical Review = Revue Internationale de Statistique, 79(2):275– 276, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305028.

Crowder:2011:SBRb

[Cro11b] Martin Crowder. Short book review: Hidden Markov Models for Time Series: An Introduction Using R, Second Edition, by Walter Zucchini, Iain L. MacDonald. International Statistical Review = Revue Internationale de Statistique, 79(1):132–133, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41306185.

Crowder:2011:SBRh

[Cro11c] Martin Crowder. Short book review: Laws of Small Numbers: Extremes and Rare Events, Third Edition, by Michael Falk, Jürg Hüsler, Rolf-Dieter Reiss. International Statistical Review = Revue Internationale de Statistique, 79(3):489, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305070.

Crowder:2011:SBRc

[Cro11d] Martin Crowder. Short book review: Meta-analysis and Combining Information in Genetics and Genomics edited by Rudy Guerra, Darlene R. Goldstein. International Statistical Review = Revue Internationale de Statistique, 79 (1):134–135, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41306186.

Crowder:2011:SBRa

[Cro11e] Martin Crowder. Short book review: NIST Handbook of Mathematical Functions edited by Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, Charles W. Clark. International Statistical Review = Revue Internationale de Statistique, 79(1):131–132, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41306184.

Crowder:2011:SBRd

[Cro11f] Martin Crowder. Short book review: Statistical Methods for Disease Clustering by Toshiro Tango. International Statistical Review = Revue Internationale de Statistique, 79(2):274–275, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305027.

Crowder:2011:SBRf

[Cro11g] Martin Crowder. Short book review: Stochastic Population Processes: Analysis, Approximations, Simulations by Eric Renshaw. International Statistical Review = Revue Internationale de Statistique, 79(3):487–488, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305068.

Crowder:2011:SBRg

[Cro11h] Martin Crowder. Short book review: The Fundamentals of Modern Statistical Genetics by Nan M. Laird, Christoph Lange. International Statistical Review = Revue Internationale de Statistique, 79(3):488–489, December 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305069.

Cuevas:1988:ESO

[CS88] E. A. Molina Cuevas and T. M. F. Smith. The effect of sampling on operative measures of association with a ratio structure. International Statistical Review = Revue Internationale de Statistique, 56(3):235–242, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403351.

Colledge:2000:ACG

[CS00] Michael J. Colledge and Rosslyn Starick. Australian Commonwealth government statistical clearing house: Vehicle for reducing response load and improving survey quality. International Statistical Review = Revue Internationale de Statistique, 68(2):221–231, August 2000. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403674.

Clark:2002:EUH

[CS02] Robert G. Clark and David G. Steel. The effect of using household as a sampling unit. International Statistical Review = Revue Internationale de Statistique, 70(2):289– 314, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403911.

Candes:2020:DPP

[CS20] Emmanuel Candès and Chiara Sabatti. Discussion of the paper “Prediction, Estimation, and Attribution” by B. Efron. International Statistical Review = Revue Internationale de Statistique, 88(S1):S60–S63, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Csiszar:1980:BRB

[Csi80] I. Csiszár. Book review: The Information in Contingency Tables by D. V. Gokhale; Solomon Kullback. International Statistical Review = Revue Internationale de Statistique, 48(2):237–238, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403172.

Csiszar:2008:SBR

[Csi08] Imre Csiszár. Short book review: Festschrift in Honor of Jorma Rissanen on the Occasion of his 75th Birthday edited by Peter Grünwald, Petri Myllymäki, Ioan Tabus, Marcelo Weinberger, Bin Yu. International Statistical Review = Revue Internationale de Statistique, 76(3):436– 437, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919656.

Chen:1993:CTL

[CSW93] Jiahua Chen, D. X. Sun, and C. F. J. Wu. A catalogue of two-level and three-level fractional factorial designs with small runs. International Statistical Review = Revue Internationale de Statistique, 61(1):131–145, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403599.

Chen:2011:PCD

[CT11] Song Xi Chen and Cheng Yong Tang. Properties of census dual system population size estimators. International Statistical Review = Revue Internationale de Statistique, 79(3):336–361, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305055.

Curran:2003:SIF

[Cur03] James M. Curran. The statistical interpretation of forensic glass evidence. International Statistical Review = Revue Internationale de Statistique, 71(3):497–520, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403825.

Crowley:1982:DPP

[CV82] John Crowley and Joseph Voelkel. Discussion of paper by P. K. Andersen, Ø. Borgan, R. Gill and N. Keiding. International Statistical Review = Revue Internationale de Statistique, 50(3):247–248, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402491.

Chiu:1973:EDC

[CW73] W. K. Chiu and G. B. Wetherill. The economic design of continuous inspection procedures: A review paper. International Statistical Review = Revue Internationale de Statistique, 41(3):357–373, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402624.

Cox:1991:SAB

[CW91] D. R. Cox and Nanny Wermuth. A simple approximation for bivariate and trivariate normal integrals. International Statistical Review = Revue Internationale de Statistique, 59(2):263–269, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403446.

Cox:2004:CSV

[CW04] D. R. Cox and Nanny Wermuth. Causality: A statistical view. International Statistical Review = Revue Internationale de Statistique, 72(3):285–305, December 2004. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472624.

Chen:2006:CBM

[CW06] Zhao-Guo Chen and Ka Ho Wu. Comparison of benchmarking methods with and without a survey error model. International Statistical Review = Revue Internationale de Statistique, 74(3):285–304, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472711.

Chen:2019:CTP

[CWCT19] Jiahua Chen, Changbao Wu, Song Cai, and Mahmoud Torabi. Contemporary theory and practice of survey sampling: a celebration of research contributions of J. N. K. Rao. International Statistical Review = Revue Internationale de Statistique, 87(S1):S1–S2, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dagnelie:1975:BRB

[Dag75] P. Dagnélie. Book review: Modèles et Méthodes de l’Analyse Factorielle by J. Torrens-Ibern. International Statistical Review = Revue Internationale de Statistique, 43(1):120, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402671.

Dagnelie:1976:SFP

[Dag76] Pierre Dagnelie. Le statisticien face aux problèmes de documentation. (French) [The statistician confronting problems of documentation]. International Statistical Review = Revue Internationale de Statistique, 44(2):231–239, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403284.

Dalenius:1972:BRB

[Dal72] Tore Dalenius. Book review: Statistical Computation by R. C. Milton; J. A. Nelder. International Statistical Review = Revue Internationale de Statistique, 40(2):254– 256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402775.

Dalenius:1977:BNSa

[Dal77a] Tore Dalenius. Bibliography on non-sampling errors in surveys. I. (A to G). International Statistical Review = Revue Internationale de Statistique, 45(1):71–89, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403005.

Dalenius:1977:BNSb

[Dal77b] Tore Dalenius. Bibliography on non-sampling errors in surveys. II. H to Q. International Statistical Review = Revue Internationale de Statistique, 45(2):181–197, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403316.

Dalenius:1977:BNSc

[Dal77c] Tore Dalenius. Bibliography on non-sampling errors in surveys. III. R to Z. International Statistical Review = Revue Internationale de Statistique, 45(3):303–317, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402543.

Dalen:1998:SCC

[Dal98] Jörgen Dalén. Studies on the comparability of consumer price indices. International Statistical Review = Revue Internationale de Statistique, 66(1):83–113, April 1998. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403658.

Damiani:1974:CMD

[Dam74] Paul Damiani. Contribution à la mesure du niveau de santé: Définition d’un indicateur lié aux facteurs sanitaires. (French) [Contribution to the measurement of the level of health: Definition of factors linked to sanitation]. International Statistical Review = Revue Internationale de Statistique, 42(1):29–37, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402683.

Damiani:1977:MML

[Dam77] Paul Damiani. Mesure de la morbidité: Liaison avec la mortalité. (French) [Measurement of morbidity: links with mortality]. International Statistical Review = Revue Internationale de Statistique, 45(1):39–50, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403002.

Daniels:1987:TPA

[Dan87] H. E. Daniels. Tail probability approximations. International Statistical Review = Revue Internationale de Statistique, 55(1):37–48, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403269.

Darroch:1981:MHT

[Dar81] J. N. Darroch. The Mantel–Haenszel test and tests of marginal symmetry; fixed-effects and mixed models for a categorical response, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 49(3):285–307, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402614.

Darroch:1984:DPD

[Dar84] J. N. Darroch. Discussion of paper by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 52(1):26–28, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403238.

Dass:2013:FBR

[Das13] Sarat C. Dass. Fingerprint-based recognition. International Statistical Review = Revue Internationale de Statistique, 81(2):175–187, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Datta:2012:DBP

[Dat12] Gauri Sankar Datta. Discussion: “Bayesian priors from loss matching” [2990343]. International Statistical Review = Revue Internationale de Statistique, 80(1):87–89, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

David:1972:ARP

[Dav72] H. A. David. Addendum: Ranking the players in a round robin tournament. International Statistical Review = Revue Internationale de Statistique, 40(3):376, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402474.

David:1982:ONH

[Dav82] H. A. David. Obituary notice: H. O. Hartley, 1912–1980. International Statistical Review = Revue Internationale de Statistique, 50(3):327–330, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402502.

David:2009:BPS

[Dav09] Herbert A. David. Karl Pearson — The Scientific Life in a Statistical Age by Theodore M. Porter: A review. International Statistical Review = Revue Internationale de Statistique, 77(1):30–39, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919688.

Davison:2020:D

[Dav20] A. C. Davison. Discussion. International Statistical Review = Revue Internationale de Statistique, 88(S1):S70– S72, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Dawid:2002:CID

[Daw02a] A. P. Dawid. Corrigenda: “Influence Diagrams for Causal Modelling and Inference”, International Statistical Review (2002) 70(2), 161–189. International Statistical Review = Revue Internationale de Statistique, 70(3):437, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403866.

Dawid:2002:IDC

[Daw02b] A. P. Dawid. Influence diagrams for causal modelling and inference. International Statistical Review = Revue Internationale de Statistique, 70(2):161–189, August 2002. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403901.

Dawson:2005:DDD

[Daw05] Donald A. Dawson. Discussant: Donald A. Dawson. International Statistical Review = Revue Internationale de Statistique, 73(2):239–240, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deAngelis:1992:SB

[dAY92] Daniela de Angelis and G. Alastair Young. Smoothing the bootstrap. International Statistical Review = Revue Internationale de Statistique, 60(1):45–56, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403500.

Pereira:1977:DAS

[dBP77] Basilio de B. Pereira. Discriminating among separate models: A bibliography. International Statistical Review = Revue Internationale de Statistique, 45(2):163–172, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403314.

Pereira:1983:RLP

[dBPR83] Carlos Alberto de Bragança Pereira and Josemar Rodrigues. Robust linear prediction in finite populations. International Statistical Review = Revue Internationale de Statistique, 51(3):293–300, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402589.

Pereira:1986:ARL

[dBPR86] C. A. de B. Pereira and J. Rodrigues. Amendment: “Robust linear prediction in finite populations” [Internat. Statist. Rev. 51 (1983), no. 3, 293–300; MR0731145 (85k:62019)]. International Statistical Review = Revue Internationale de Statistique, 54(1):115, 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

denButter:1982:SPR

[dBV82] F. A. G. den Butter and H. A. A. Verbon. The specification problem in regression analysis. International Statistical Review = Revue Internationale de Statistique, 50 (3):267–283, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402497.

Diggle:1983:SDB

[DC83] Peter J. Diggle and Trevor F. Cox. Some distance-based tests of independence for sparsely-sampled multivariate spatial point patterns. International Statistical Review = Revue Internationale de Statistique, 51(1):11–23, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402729.

Dawber:2019:MGH

[DC19] James Dawber and Raymond Chambers. Modelling group heterogeneity for small area estimation using M-quantiles. International Statistical Review = Revue Internationale de Statistique, 87(S1):S50–S63, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dagum:1998:UVS

[DCC98] Estela Bee Dagum, Pierre A. Cholette, and Zhao-Guo Chen. A unified view of signal extraction, benchmarking, interpolation and extrapolation of time series. International Statistical Review = Revue Internationale de Statistique, 66(3):245–269, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403516.

Duncan:1980:RIS

[DD80] J. W. Duncan and J. Durbin. Report of the International Statistical Institute Committee on the Integration of Statistics. International Statistical Review = Revue Internationale de Statistique, 48(2):139–168, August 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403150.

Debonneuil:1978:CIA

[Deb78] Xavier Debonneuil. The 11th Conference of International Association for Regional and Urban Statistics: An international symposium on regional and urban statistics. International Statistical Review = Revue Internationale de Statistique, 46(3):313–316, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402377.

Deininger:1976:BRB

[Dei76] R. Deininger. Book review: Angewandte Statistik by Lothar Sachs; Statistische Methoden. International Statistical Review = Revue Internationale de Statistique, 44 (2):304–305, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403296.

Deistler:1983:DPD

[Dei83] Manfred Deistler. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):148–149, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402739.

Dekker:1994:CMP

[Dek94] Arij L. Dekker. Computer methods in population census data processing. International Statistical Review = Revue Internationale de Statistique, 62(1):55–70, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403545.

Dekkers:2010:SBR

[Dek10] Gijs Dekkers. Short book review: New Frontiers in Microsimulation Modelling edited by Asghar Zaidi, Ann Harding, Paul Williamson. International Statistical Review = Revue Internationale de Statistique, 78(3):481– 482, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919896.

Dekkers:2014:BRQ

[Dek14] Gijs Dekkers. Book review: Quantile Regression: Theory and Applications by Cristina Davino, Marilena Furno, Domenico Vistocco. International Statistical Review = Revue Internationale de Statistique, 82(2):320–322, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dekkers:2016:BRS

[Dek16] Gijs Dekkers. Book review: STATA for the Behavioral Sciences. International Statistical Review = Revue Internationale de Statistique, 84(2):323–324, August 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dekkers:2022:BRT

[Dek22] Gijs Dekkers. Book review: Thinking Clearly with Data: a Guide to Quantitative Reasoning and Analysis. Ethan Bueno de Mesquita and Anthony Fowler. Princeton University Press, 2021, 400 pages, $95.00/£74.00, hardback ISBN: 978-0-691-21436-8. International Statistical Review = Revue Internationale de Statistique, 90(3):622–625, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Deming:1972:CPC

[Dem72] W. Edwards Deming. Code of professional conduct: A personal view. International Statistical Review = Revue Internationale de Statistique, 40(2):215–219, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402763.

Dempster:1973:BRB

[Dem73] A. P. Dempster. Book review: An Introduction to the Theory of Statistics by R. L. Plackett. International Statistical Review = Revue Internationale de Statistique, 41 (2):294, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402858.

Deming:1977:EST

[Dem77] W. Edwards Deming. An essay on screening, or on twophase sampling, applied to surveys of a community. International Statistical Review = Revue Internationale de Statistique, 45(1):29–37, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403001.

Dempster:1983:RWG

[Dem83] A. P. Dempster. Reflections on W. G. Cochran, 1909– 1980. International Statistical Review = Revue Internationale de Statistique, 51(3):321–322, December 1983. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402591.

Demidenko:2007:PRC

[Dem07] Eugene Demidenko. Poisson regression for clustered data. International Statistical Review = Revue Internationale de Statistique, 75(1):96–113, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508451.

Demidenko:2018:OAD

[Dem18] Eugene Demidenko. Optimal adaptive designs with inverse ordinary differential equations. International Statistical Review = Revue Internationale de Statistique, 86(2): 169–188, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Derksen:1975:BRBb

[Der75a] J. B. D. Derksen. Book review: Analysis of Seasonality and Trends in Statistical Series: Methodology and Applications in Israel by Raphael Raymond V. BarOn. International Statistical Review = Revue Internationale de Statistique, 43(2):240–241, August 1975. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402909.

Derksen:1975:BRBa

[Der75b] J. B. D. Derksen. Book review: How to Find out about Statistics by Gillian A. Burrington. International Statistical Review = Revue Internationale de Statistique, 43(1): 116–117, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402666.

Derksen:1977:BRBb

[Der77a] J. B. D. Derksen. Book review: Cost of Living Index Numbers; Practice, Precision and Theory by Kali S. Banerjee. International Statistical Review = Revue Internationale de Statistique, 45(2):201–202, August 1977. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403320.

Derksen:1977:BRBc

[Der77b] J. B. D. Derksen. Book review: Elementary Statistics by Paul G. Hoel. International Statistical Review = Revue Internationale de Statistique, 45(3):324–325, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402552.

Derksen:1977:BRBa

[Der77c] J. B. D. Derksen. Book review: Elements of Sampling Theory by V. Barnett. International Statistical Review = Revue Internationale de Statistique, 45(1):96–97, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403011.

Dey:2003:SAN

[DER03] Jyotirmoy Dey, R. V. Erickson, and R. V. Ramamoorthi. Some aspects of neutral to right priors. International Statistical Review = Revue Internationale de Statistique, 71 (2):383–401, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403895.

Desrosieres:2000:MUH

[Des00] Alain Desrosières. Measurement and its uses: Harmonization and quality in social statistics. International Statistical Review = Revue Internationale de Statistique, 68(2): 173–187, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403667.

Deshpande:2005:DJD

[Des05] Jayant Deshpande. Discussant: Jayant Deshpande. International Statistical Review = Revue Internationale de Statistique, 73(2):183–185, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deFinetti:1974:BUR

[dF74] Bruno de Finetti. Bayesianism: its unifying role for both the foundations and applications of statistics. International Statistical Review = Revue Internationale de Statistique, 42(2):117–130, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403075.

deFinetti:1979:PES

[dF79] Bruno de Finetti. Probability and exchangeability from a subjective point of view. International Statistical Review = Revue Internationale de Statistique, 47(2):129– 135, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402565.

Domingo-Ferrer:2011:D

[DF11] Josep Domingo-Ferrer. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2): 184–186, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Diggle:1994:PEP

[DFG^{+}94] Peter J. Diggle, Thomas Fiksel, Pavel Grabarnik, Yosihiko Ogata, Dietrich Stoyan, and Masaharu Tanemura. On parameter estimation for pairwise interaction point processes. International Statistical Review = Revue Internationale de Statistique, 62(1):99–117, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403548.

Delporte:2022:JNB

[DFM^{+}22] Margaux Delporte, Steffen Fieuws, Geert Molenberghs, Geert Verbeke, Simeon Situma Wanyama, Elpis Hatziagorou, and Christiane De Boeck. A joint normal-binary (probit) model. International Statistical Review = Revue Internationale de Statistique, 90:S37–S51, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dransfield:1999:USSa

[DFV99a] S. B. Dransfield, N. I. Fisher, and N. J. Vogel. Using statistics and statistical thinking to improve organisational performance. International Statistical Review = Revue Internationale de Statistique, 67(2):99–122, August 1999. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403389.

Dransfield:1999:USSb

[DFV99b] S. B. Dransfield, N. I. Fisher, and N. J. Vogel. [using statistics and statistical thinking to improve organisational performance]: Response. International Statistical Review = Revue Internationale de Statistique, 67(2):146– 150, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403395.

Doksum:1990:CBM

[DG90] Kjell A. Doksum and Miriam Gasko. On a correspondence between models in binary regression analysis and in survival analysis. International Statistical Review = Revue Internationale de Statistique, 58(3):243–252, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403807.

Dutta:2017:DSS

[DG17] Subhajit Dutta and Anil K. Ghosh. Discussions on statistical scale space methods: Discussion. International Statistical Review = Revue Internationale de Statistique, 85 (1):40–43, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deGooijer:1985:MDO

[dGAGR85] Jan G. de Gooijer, Bovas Abraham, Ann Gould, and Lecily Robinson. Methods for determining the order of an autoregressive-moving average process: a survey. International Statistical Review = Revue Internationale de Statistique, 53(3):301–329, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402894.

Doretti:2018:MDU

[DGS18] Marco Doretti, Sara Geneletti, and Elena Stanghellini. Missing data: A unified taxonomy guided by conditional independence. International Statistical Review = Revue Internationale de Statistique, 86(2):189–204, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Das:2019:CMP

[DH19] Sumonkanti Das and Stephen Haslett. A comparison of methods for poverty estimation in developing countries. International Statistical Review = Revue Internationale de Statistique, 87(2):368–392, August 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dillman:2000:RBS

[Dil00] Don A. Dillman. The role of behavioral survey methodologists in national statistical agencies. International Statistical Review = Revue Internationale de Statistique, 68(2): 200–213, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403670.

Dimakos:2001:GES

[Dim01] Xeni K. Dimakos. A guide to exact simulation. International Statistical Review = Revue Internationale de Statistique, 69(1):27–48, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403528.

Dinges:1982:DPD

[Din82] H. Dinges. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):15–17, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402451.

Dominicy:2017:MHE

[DIV17] Yves Dominicy, Pauliina Ilmonen, and David Veredas. Multivariate hill estimators. International Statistical Review = Revue Internationale de Statistique, 85(1):108– 142, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dixon:1973:AO

[Dix73] W. J. Dixon. Annotated output. International Statistical Review = Revue Internationale de Statistique, 41(2): 259, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402843.

Duncan:1987:IDA

[DK87] Greg J. Duncan and Graham Kalton. Issues of design and analysis of surveys across time. International Statistical Review = Revue Internationale de Statistique, 55(1): 97–117, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403273.

Dellaportas:2001:SAN

[DK01] Petros Dellaportas and Dimitris Karlis. A simulation approach to nonparametric empirical Bayes analysis. International Statistical Review = Revue Internationale de Statistique, 69(1):63–79, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403530.

Denby:2013:CCM

[DL13] Lorraine Denby and Jim Landwehr. A conversation with Colin L. Mallows. International Statistical Review = Revue Internationale de Statistique, 81(3):338–360, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Daley:1973:TPP

[DM73] D. J. Daley and R. K. Milne. The theory of point processes: A bibliography. International Statistical Review = Revue Internationale de Statistique, 41(2):183–201, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402835.

deMoivre:1984:MCP

[dM84] Abraham de Moivre. On the measurement of chance, or, on the probability of events in games depending upon fortuitous chance. International Statistical Review = Revue Internationale de Statistique, 52(3):237–262, 1984. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). Translated from the Latin by Bruce McClintock.

Demarta:2005:CRC

[DM05] Stefano Demarta and Alexander J. McNeil. The t copula and related copulas. International Statistical Review = Revue Internationale de Statistique, 73(1):111–129, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472643.

Dey:2012:DRE

[DM12] Aloke Dey and Rahul Mukerjee. Development of research in experimental design in India. International Statistical Review = Revue Internationale de Statistique, 80(2):231– 252, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dvorak:2022:NTD

[DMMG22] Jirí Dvorák, Tomás Mrkvicka, Jorge Mateu, and Jonatan A. González. Nonparametric testing of the dependence structure among points–marks–covariates in spatial point patterns. International Statistical Review = Revue Internationale de Statistique, 90(3):592–621, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rueda:2018:ETO

[dMRAMR18] Maria del Mar Rueda, Antonio Arcos, David Molina, and Maria Giovanna Ranalli. Estimation techniques for ordinal data in multiple frame surveys with complex sampling designs. International Statistical Review = Revue Internationale de Statistique, 86(1):51–67, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Debon:2006:CNM

[DMS06] Ana Debón, Francisco Montes, and Ramón Sala. A comparison of nonparametric methods in the graduation of mortality: Application to data from the Valencia Region (Spain). International Statistical Review = Revue Internationale de Statistique, 74(2):215–233, August 2006. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472704.

Davies:2006:EOC

[DNC06] Simon L. Davies, Andrew A. Neath, and Joseph E. Cavanaugh. Estimation optimality of corrected AIC and modified Cp in linear regression. International Statistical Review = Revue Internationale de Statistique, 74(2): 161–168, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472700.

Denton:1973:MEC

[DO73] Frank T. Denton and Ernest H. Oksanen. Measurement error and choice of econometric estimation method: Some empirical findings. International Statistical Review = Revue Internationale de Statistique, 41(3):343–349, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402622.

deOliveira:1979:BRBa

[dO79a] J. Tiago de Oliveira. Book review: Parameter Estimation in Engineering and Science by James V. Beck; Kenneth J. Arnold. International Statistical Review = Revue Internationale de Statistique, 47(1):85, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403212.

deOliveira:1979:BRBb

[dO79b] J. Tiago de Oliveira. Book review: The Asymptotic Theory of Extreme Order Statistics by Janos Galambos.

International Statistical Review = Revue Internationale de Statistique, 47(2):230–231, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402572.

deOliveira:2013:BD

[dO13] A. Gouveia de Oliveira. Biostatistics Decoded. John Wiley, New York, NY, USA, 2013. ISBN 1-119-95337-5 (paperback), 1-118-67077-9. x + 335 pp. LCCN QH323.5 .O5613 2013.

Dockery:2022:SBE

[Doc22] Douglas W. Dockery. Synergy of biostatistics and epidemiology in air pollution health effects studies. International Statistical Review = Revue Internationale de Statistique, 90:S67–S81, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dodge:1996:NRN

[Dod96] Yadolah Dodge. A natural random number generator. International Statistical Review = Revue Internationale de Statistique, 64(3):329–344, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403789.

Donner:1986:RIP

[Don86] Allan Donner. A review of inference procedures for the intraclass correlation coefficient in the one-way random effects model. International Statistical Review = Revue Internationale de Statistique, 54(1):67–82, April 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403259.

Donnet:2012:SBR

[Don12] Sophie Donnet. Short book review: Monte Carlo Simulation for the Pharmaceutical Industry: Concepts, Algorithms, and Case Studies by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 80(1):186–187, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257181.

Dorhout:1973:BRB

[Dor73] B. Dorhout. Book review: Stochastisches Programmieren by M. M. Faber. International Statistical Review = Revue Internationale de Statistique, 41(1):139, April 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402796.

Dorfman:2018:TRE

[Dor18] Alan H. Dorfman. Towards a routine external evaluation protocol for small area estimation. International Statistical Review = Revue Internationale de Statistique, 86(2): 259–274, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Douglas:1995:FBD

[Dou95] J. B. Douglas. Fitting bivariate discrete distributions. International Statistical Review = Revue Internationale de Statistique, 63(1):117–118, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403780.

Downton:1972:BRB

[Dow72] F. Downton. Book review: Probability and Statistics by D. E. Bailey. International Statistical Review = Revue Internationale de Statistique, 40(2):249–250, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402768.

Dunson:2020:D

[DP20] David Dunson and Theodore Papamarkou. Discussions. International Statistical Review = Revue Internationale de Statistique, 88(2):321–324, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

DeIaco:2021:SCG

[DPCM21] Sandra De Iaco, Donato Posa, Claudia Cappello, and Sabrina Maggio. On some characteristics of Gaussian covariance functions. International Statistical Review = Revue Internationale de Statistique, 89(1):36–53, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Darius:2007:VET

[DPS07] Paul L. Darius, Kenneth M. Portier, and Eddie Schrevens. Virtual experiments and their use in teaching experimental design. International Statistical Review = Revue Internationale de Statistique, 75(3):281–294, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41509871.

Durbin:1997:BSS

[DQ97] J. Durbin and B. Quenneville. Benchmarking by state space models. International Statistical Review = Revue Internationale de Statistique, 65(1):23–48, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403431.

Dagum:1992:TDV

[DQS92] Estela Bee Dagum, Benoît Quenneville, and Brajendra Sutradhar. Trading-day variations multiple regression models with random parameters. International Statistical Review = Revue Internationale de Statistique, 60(1): 57–73, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403501.

DiCiccio:1990:NCL

[DR90] Thomas J. DiCiccio and Joseph P. Romano. Nonparametric confidence limits by resampling methods and least favorable families. International Statistical Review = Revue Internationale de Statistique, 58(1):59–76, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403474.

Diggle:1993:ESI

[DR93] Peter J. Diggle and Sylvia Richardson. Epidemiological studies of industrial pollutants: An introduction. International Statistical Review = Revue Internationale de Statistique, 61(2):203–206, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403623.

Draper:2008:SBRa

[Dra08a] Norman R. Draper. Short book review: Generalized Linear Models for Insurance Data by Piet de Jong, Gillian Z. Heller. International Statistical Review = Revue Internationale de Statistique, 76(2):315, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919635.

Draper:2008:SBRb

[Dra08b] Norman R. Draper. Short book review: Statistical Practice in Business and Industry edited by Shirley Coleman, Tony Greenfield, Dave Stewardson, Douglas C. Montgomery. International Statistical Review = Revue Internationale de Statistique, 76(3):453, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919673.

Draper:2009:SBRd

[Dra09a] Norman R. Draper. Short book review: Analysis of Messy Data Volume 1: Designed Experiments, Second Edition, by George A. Milliken, Dallas E. Johnson. International Statistical Review = Revue Internationale de Statistique, 77(2):321–322, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919751.

Draper:2009:SBRa

[Dra09b] Norman R. Draper. Short book review: Data Analysis for Experimental Design by Richard Gonzales. International Statistical Review = Revue Internationale de Statistique, 77(1):147, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919696.

Draper:2009:SBRb

[Dra09c] Norman R. Draper. Short book review: Modern Regression Methods, Second Edition, by Thomas P. Ryan. International Statistical Review = Revue Internationale de Statistique, 77(2):303, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919732.

Draper:2009:SBRc

[Dra09d] Norman R. Draper. Short book review: Who Gave You the Epsilon? & Other Tales of Mathematical History edited by Marlow Anderson, Victor Katz, Robin Wilson. International Statistical Review = Revue Internationale de Statistique, 77(2):309–310, August 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919738.

Draper:2010:SBRb

[Dra10a] Norman R. Draper. Short book review: Linear Model Methodology by André I. Khuri. International Statistical Review = Revue Internationale de Statistique, 78(2): 316, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919840.

Draper:2010:SBRc

[Dra10b] Norman R. Draper. Short book review: Regression Estimators: A Comparative Study, Second Edition, by Marvin H. J. Gruber. International Statistical Review = Revue Internationale de Statistique, 78(3):460, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919876.

Draper:2010:SBRa

[Dra10c] Norman R. Draper. Short book review: Statistical Analysis of Designed Experiments: Theory and Applications by Ajit C. Tamhane. International Statistical Review = Revue Internationale de Statistique, 78(1):136, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919801.

Draper:2011:SBRb

[Dra11a] Norman R. Draper. Short book review: Design and Analysis of Experiments with SAS by John Lawson. International Statistical Review = Revue Internationale de Statistique, 79(1):118, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41306171.

Draper:2011:SBRa

[Dra11b] Norman R. Draper. Short book review: Mathematics and Sports edited by Joseph A. Gallian. International Statistical Review = Revue Internationale de Statistique, 79 (1):116, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306169.

Draper:2011:SBRe

[Dra11c] Norman R. Draper. Short book review: Regression with Linear Predictors by Per Kragh Andersen, Lene Theil Skovgaard. International Statistical Review = Revue Internationale de Statistique, 79(2):279, August 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305031.

Draper:2011:SBRc

[Dra11d] Norman R. Draper. Short book review: Statistics for Archaeologists: A Common Sense Approach, Second Edition, by Robert D. Drennan. International Statistical Review = Revue Internationale de Statistique, 79(1):128–129, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41306181.

Draper:2011:SBRd

[Dra11e] Norman R. Draper. Short book review: The Cambridge Dictionary of Statistics, Fourth Edition, by B. S. Everitt, A. Skrondal. International Statistical Review = Revue Internationale de Statistique, 79(2):273–274, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41305026.

Draper:2012:SBRb

[Dra12a] Norman R. Draper. Short book review: A Guide to Plane Algebraic Curves by Keith Kendig. International Statistical Review = Revue Internationale de Statistique, 80 (1):183, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/23257178.

Draper:2012:SBRc

[Dra12b] Norman R. Draper. Short book review: A Primer on Experiments with Mixtures by John A. Cornell. International Statistical Review = Revue Internationale de Statistique, 80(2):338, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2012:SBRf

[Dra12c] Norman R. Draper. Short book review: Design and Analysis of Experiments, Volume 3, Special Designs and Applications edited by Klaus Hinkelmann. International Statistical Review = Revue Internationale de Statistique, 80(3): 484, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2012:SBRd

[Dra12d] Norman R. Draper. Short book review: Optimal Design of Experiments: A Case Study Approach by Peter Goos and Bradley Jones. International Statistical Review = Revue Internationale de Statistique, 80(2):341–342, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2012:SBRe

[Dra12e] Norman R. Draper. Short book review: Sophie’s Diary: A Mathematical Novel, Second Edition, by Dora Musielak. International Statistical Review = Revue Internationale de Statistique, 80(3):481–482, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2012:SBRa

[Dra12f] Norman R. Draper. Short book review: Train Your Brain: A Year’s Worth of Puzzles by George Grätzer. International Statistical Review = Revue Internationale de Statistique, 80(1):182, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/23257177.

Draper:2013:SBRd

[Dra13a] Norman R. Draper. Short book review: An Accidental Statistician: The Life and Memories of George E. P. Box by George E. P. Box. International Statistical Review = Revue Internationale de Statistique, 81(3):465–466, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRg

[Dra13b] Norman R. Draper. Short book review: Martin Gardner in the Twenty-First Century by Michael Henle, Brian Hopkins. International Statistical Review = Revue Internationale de Statistique, 81(3):484–485, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRf

[Dra13c] Norman R. Draper. Short book review: Paradoxes in Scientific Inference by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 81(3):483– 484, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRa

[Dra13d] Norman R. Draper. Short book review: Selected Works of Terry Speed edited by Sandrine Dudoit. International Statistical Review = Revue Internationale de Statistique, 81(1):170, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRe

[Dra13e] Norman R. Draper. Short book review: Six Sources of Collapse: A Mathematician’s Perspective on How Things Can Fall Apart in the Blink of an Eye by Charles R. Hadlock. International Statistical Review = Revue Internationale de Statistique, 81(3):474–475, December 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRb

[Dra13f] Norman R. Draper. Short book review: The A–Z of Error-Free Research by Phillip I. Good. International Statistical Review = Revue Internationale de Statistique, 81(1): 171–172, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2013:SBRc

[Dra13g] Norman R. Draper. Short book review: The Signal and the Noise: Why So Many Predictions Fail – But Some Don’t by Nate Silver. International Statistical Review = Revue Internationale de Statistique, 81(2):332–333, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2014:BRM

[Dra14a] Norman R. Draper. Book review: Matrix Algebra for Linear Models by Martin H. J. Gruber. International Statistical Review = Revue Internationale de Statistique, 82 (2):309, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2014:BRO

[Dra14b] Norman R. Draper. Book review: Optimal Mixture Experiments by Bikas Kumar Sinha, Nripes Kumar Mandal, Manisha Pal, Premadhis Das. International Statistical Review = Revue Internationale de Statistique, 82(3):478, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2014:BRSa

[Dra14c] Norman R. Draper. Book review: Statistical Robust Design: An Industrial Perspective by Magnus Arner. International Statistical Review = Revue Internationale de Statistique, 82(2):317–318, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Draper:2014:BRSb

[Dra14d] Norman R. Draper. Book review: Statistics in Action: A Canadian Outlook by Jerald F. Lawless. International Statistical Review = Revue Internationale de Statistique, 82(3):479–480, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Drion:1978:BRB

[Dri78] E. F. Drion. Book review: Graphische Gestaltung von Zahlenmaterial by Hans Riedwyl. International Statistical Review = Revue Internationale de Statistique, 46(3):321– 322, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402386.

Druilhet:2011:SBR

[Dru11] Pierre Druilhet. Short book review: Plans d’expérience: constructions et analyses statistiques by Walter Tinsson. International Statistical Review = Revue Internationale de Statistique, 79(1):129–130, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41306182.

Dudley:1978:CDS

[DS78] R. J. Dudley and K. W. Smalldon. The comparison of distributional shapes with particular reference to a problem in forensic science. International Statistical Review = Revue Internationale de Statistique, 46(1):53–63, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402507.

DeSantis:1999:MDR

[DS99] Fulvio De Santis and Fulvio Spezzaferri. Methods for default and robust Bayesian model comparison: The fractional Bayes factor approach. International Statistical Review = Revue Internationale de Statistique, 67(3):267– 286, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403706.

Dellaportas:2020:IPA

[DS20] Petros Dellaportas and David A. Stephens. Interview with Professor Adrian FM Smith. International Statistical Review = Revue Internationale de Statistique, 88(2): 265–279, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Du:2022:VSI

[DS22] Mingyue Du and Jianguo Sun. Variable selection for interval-censored failure time data. International Statistical Review = Revue Internationale de Statistique, 90(2): 193–215, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dominici:2003:HEA

[DSC03] Francesca Dominici, Lianne Sheppard, and Merlise Clyde. Health effects of air Pollution: A statistical review. International Statistical Review = Revue Internationale de Statistique, 71(2):243–276, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403886.

Davison:1992:RMD

[DT92] A. C. Davison and C.-L. Tsai. Regression model diagnostics. International Statistical Review = Revue Internationale de Statistique, 60(3):337–353, December 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403682.

Dong:2024:GII

[DTGG24] Ziqing Dong, Yves Tille, Giovanni Maria Giorgi, and Alessio Guandalini. Generalised Income Inequality Index. International Statistical Review = Revue Internationale de Statistique, 92(1):87–105, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dwivedi:2020:SDI

[DTP^{+}20] Raaz Dwivedi, Yan Shuo Tan, Briton Park, Mian Wei, Kevin Horgan, David Madigan, and Bin Yu. Stable discovery of interpretable subgroups via calibration in causal studies. International Statistical Review = Revue Internationale de Statistique, 88(S1):S135–S178, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Duffett:1978:HM

[Duf78] W. Duffett. Herbert Marshall, 1887–1977. International Statistical Review = Revue Internationale de Statistique, 46(1):125, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402528.

Duflo:1982:DPP

[Duf82] Marie Duflo. Discussion of paper by P. K. Andersen, Ø. Borgan, R. Gill and N. Keiding. International Statistical Review = Revue Internationale de Statistique, 50(3):248– 250, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402492.

Dugue:1974:HL

[Dug74a] Daniel Dugué. Henri Laugier, 1888–1973. International Statistical Review = Revue Internationale de Statistique, 42(1):113, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402696.

Dugue:1974:MF

[Dug74b] Daniel Dugué. Maurice Fréchet, 1878–1973. International Statistical Review = Revue Internationale de Statistique, 42(1):113–114, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402697.

Dugue:1978:RR

[Dug78] Daniel Dugué. René Roy, 1894–1977. International Statistical Review = Revue Internationale de Statistique, 46(1): 125–126, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402529.

Duncan:2011:D

[Dun11] George T. Duncan. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):186– 191, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2013:SBRb

[Dur13a] Fabrizio Durante. Short book review: Bridge to Abstract Mathematics by Ralph Oberste-Vorth, Aristides Mouzakitis, Bonita A. Lawrence. International Statistical Review = Revue Internationale de Statistique, 81(2):325, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2013:SBRc

[Dur13b] Fabrizio Durante. Short book review: New Perspectives in Statistical Modeling and Data Analysis: Proceedings of the 7th Conference of the Classification and Data Analysis Group of the Italian Statistical Society, Catania, September 9–11, 2009 by Salvatore Ingrassia, Roberto Rocci, Maurizio Vichi. International Statistical Review = Revue Internationale de Statistique, 81(3):460–461, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2013:SBRa

[Dur13c] Fabrizio Durante. Short book review: Simulating Copulas: Stochastic Models, Sampling Algorithms and Applications by Jan-Frederik Mai, Matthias Scherer, with contributions by Claudia Czado, Elke Korn, Ralf Korn, Jakob Stöber. International Statistical Review = Revue Internationale de Statistique, 81(2):307, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2014:BRBb

[Dur14a] Fabrizio Durante. Book review: Advanced Risk Analysis in Engineering Enterprise Systems by Cesar Ariel Pinto, Paul R. Garvey. International Statistical Review = Revue Internationale de Statistique, 82(1):146–147, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2014:BRBa

[Dur14b] Fabrizio Durante. Book review: The R Book, Second Edition by Michael J. Crawley. International Statistical Review = Revue Internationale de Statistique, 82(1): 145–146, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2014:BRS

[Dur14c] Fabrizio Durante. Book review: The Skew-Normal and Related Families by Adelchi Azzalini. International Statistical Review = Revue Internationale de Statistique, 82 (3):483, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2015:BRC

[Dur15a] Fabrizio Durante. Book review: Computational Actuarial Science with R. International Statistical Review = Revue Internationale de Statistique, 83(3):511, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durrett:2015:BPM

[Dur15b] Richard Durrett. Branching Process Models of Cancer, volume 1 of Mathematical Biosciences Institute lecture series. Stochastics in biological systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-16065-6 (e-book), 3-319-16064-8 (print). vii + 63 pp. LCCN RC267. URL http:/

/public.eblib.com/choice/publicfullrecord.aspx?
p=2096172;   http://www.myilibrary.com?id=802056.

Durante:2017:BRE

[Dur17] Fabrizio Durante. Book review: Educational Measurement for Applied Researchers, Margaret Wu Hak Ping Tam and Jen Tsung-Hau Springer Nature Singapore Pte Ltd., 2016, xiv + 306 pages, £82.00, hardcover ISBN: 978-981-10-3300-1. International Statistical Review = Revue Internationale de Statistique, 85(2):372–373, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2018:BRH

[Dur18] Fabrizio Durante. Book review: Handbook for Applied Modeling: Non-Gaussian and Correlated Data, Jamie D. Riggs and Trent L. Lalonde, Cambridge University Press, 2017, 228 pages, £29.99, softcover ISBN: 978-1-316-60105-1. International Statistical Review = Revue Internationale de Statistique, 86(3):599–600, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2021:BRD

[Dur21] Fabrizio Durante. Book review: Direction Dependence in Statistical Modeling: Methods of Analysis, Wolfgang Wiedermann, Daeyoung Kim, Engin A. Sungur and Alexander von Eye, Editors Wiley, 2021, 432 pages, £102, hardcover ISBN: 978-1-119-52307-9. International Statistical Review = Revue Internationale de Statistique, 89(3):659–660, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2022:BRE

[Dur22] Fabrizio Durante. Book review: Extreme Value Theory with Applications to Natural Hazards: From Statistical Theory to Industrial Practice. Edited by Nicolas Bousquet and Pietro Bernardara. Springer Cham, 2021, xxii + 481 pages, $199.99. ISBN: 978-3-030-74941-5. International Statistical Review = Revue Internationale de Statistique, 90(2):411–412, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2023:BRN

[Dur23a] Fabrizio Durante. Book review: Number Savvy: From the Invention of Numbers to the Future of Data, George Sciadas, Chapman & Hall/CRC, 2022, 312 pages, £56.99/ $74.95, hardcover, ISBN 978-1-032-36215-1. International Statistical Review = Revue Internationale de Statistique, 91(2):348, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2023:BRS

[Dur23b] Fabrizio Durante. Book review: Statistical Methods for Climate Scientists Timothy M. DelSole and Michael K. Tippett Cambridge University Press, 2022, 542 pages, £54.99, hardcover. ISBN: 978-1-108-47241-8. International Statistical Review = Revue Internationale de Statistique, 91(3):535, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deVries:1999:WML

[dV99] Willem de Vries. Are we measuring up … ? questions on the performance of national statistical systems. International Statistical Review = Revue Internationale de Statistique, 67(1):63–77, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403566.

deVries:2001:MMI

[dV01] Willem F. M. de Vries. Meaningful measures: Indicators on progress, progress on indicators. International Statistical Review = Revue Internationale de Statistique, 69(2): 313–331, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403818.

deVries:1996:RRB

[dVKW96] Willem de Vries, Wouter Keller, and Ad Willeboordse. Reducing the response burden: Some developments in the Netherlands. International Statistical Review = Revue Internationale de Statistique, 64(2):199–213, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403742.

Duffy:1984:RRM

[DW84] John C. Duffy and Jennifer J. Waterton. Randomized response models for estimating the distribution function of a quantitative character. International Statistical Review = Revue Internationale de Statistique, 52(2):165– 171, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403099.

Duarte:2015:FBO

[DW15] Belmiro P. M. Duarte and Weng Kee Wong. Finding Bayesian optimal designs for nonlinear models: A semidefinite programming-based approach. International Statistical Review = Revue Internationale de Statistique, 83(2): 239–262, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deWaal:2005:AEB

[dWC05] Ton de Waal and Wieger Coutinho. Automatic editing for business surveys: An assessment of selected algorithms. International Statistical Review = Revue Internationale de Statistique, 73(1):73–102, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472641.

deWaal:2020:MSS

[dWvDS20] Ton de Waal, Arnout van Delden, and Sander Scholtus. Multi-source statistics: Basic situations and methods. International Statistical Review = Revue Internationale de Statistique, 88(1):203–228, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Dey:2016:EVM

[DY16] Dipak Dey and Jun Yan, editors. Extreme value modeling and risk analysis: methods and applications. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4987-0129-9. xx + 520 pp. LCCN QA273.6 .E9685 2016.

Eagleson:1974:BRB

[Eag74] G. K. Eagleson. Book review: Introductory Mathematical Statistics by Erwin Kreyszig. International Statistical Review = Revue Internationale de Statistique, 42 (1):108, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402690.

El-Badry:1992:EPI

[EB92] M. A. El-Badry. Emerging population issues in Africa. International Statistical Review = Revue Internationale de Statistique, 60(2):119–127, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403645.

Editor:1977:SSP

[Edi77] The Executive Editor. “setting statistical priorities”: A review note. International Statistical Review = Revue Internationale de Statistique, 45(2):198–199, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403317.

Edwards:1974:HL

[Edw74] A. W. F. Edwards. The history of likelihood. International Statistical Review = Revue Internationale de Statistique, 42(1):9–15, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402681.

Edwards:1982:PPP

[Edw82] A. W. F. Edwards. Pascal and the problem of points. International Statistical Review = Revue Internationale de Statistique, 50(3):259–266, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402496.

Edwards:1983:PPG

[Edw83] A. W. F. Edwards. Pascal’s problem: The ‘Gambler’s Ruin’. International Statistical Review = Revue Internationale de Statistique, 51(1):73–79, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402732.

Edelmann:2019:ULR

[EFP19] Dominic Edelmann, Konstantinos Fokianos, and Maria Pitsillou. An updated literature review of distance correlation and its applications to time series. International Statistical Review = Revue Internationale de Statistique, 87(2):237–262, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Efron:2013:DCD

[Efr13] Bradley Efron. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1): 41–42, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Efron:2020:PEA

[Efr20a] Bradley Efron. Prediction, estimation, and attribution. International Statistical Review = Revue Internationale de Statistique, 88(S1):S28–S59, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [CS20, Cox20, Cox20, FHT20, gXZ20, Cre20, Dav20, YB20] and rejoinder [Efr20b].

Efron:2020:R

[Efr20b] Bradley Efron. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 88(S1):S87– S90, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Erlander:1972:ASC

[EGS72] S. Erlander, J. Gustavsson, and Å. Svensson. On asymptotic simultaneous confidence regions for regression planes in a Poisson model. International Statistical Review = Revue Internationale de Statistique, 40(2):111–122, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402754.

Embrechts:1991:VAP

[EH91] Paul Embrechts and Agnes M. Herzberg. Variations of Andrews’ plots. International Statistical Review = Revue Internationale de Statistique, 59(2):175–194, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403442.

Eichenauer-Herrmann:1992:ICP

[EH92] Jürgen Eichenauer-Herrmann. Inversive congruential pseudorandom numbers: a tutorial. International Statistical Review = Revue Internationale de Statistique, 60(2): 167–176, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403647.

Eichenauer-Herrmann:1995:PNG

[EH95] Jürgen Eichenauer-Herrmann. Pseudorandom number generation by nonlinear methods. International Statistical Review = Revue Internationale de Statistique, 63(2): 247–255, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403620.

Eideh:2007:CNS

[Eid07] Abdulhakeem A. H. Eideh. A correction note on small area estimation. International Statistical Review = Revue Internationale de Statistique, 75(1):122–123, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508453.

Ekin:2018:SMF

[EIRS18] Tahir Ekin, Francesca Ieva, Fabrizio Ruggeri, and Refik Soyer. Statistical medical fraud assessment: Exposition to an emerging field. International Statistical Review = Revue Internationale de Statistique, 86(3):379–402, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Escoufier:1986:DPG

[EJ86] Y. Escoufier and S. Junca. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):279–283, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403057.

El-Khorazaty:1977:ETN

[EKIKW77] M. Nabil El-Khorazaty, Peter B. Imrey, Gary G. Koch, and H. Bradley Wells. Estimating the total number of events with data from multiple-record systems: A review of methodological strategies. International Statistical Review = Revue Internationale de Statistique, 45(2):129– 157, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403312.

Elashoff:1973:UTS

[Ela73] Janet D. Elashoff. The use of transformation to stabilize variances in an analysis of variance example. International Statistical Review = Revue Internationale de Statistique, 41(2):228–234, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402838.

Elfving:1972:BRB

[Elf72] G. Elfving. Book review: Applied Statistics in Decision-Making by Georg K. Chacko. International Statistical Review = Revue Internationale de Statistique, 40(3):393, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402478.

Ellis:1991:TAL

[Ell91] Steven P. Ellis. Topological aspects of the location problem for directional and axial data. International Statistical Review = Revue Internationale de Statistique, 59(3):389– 394, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403694.

Elsayed:2018:BRP

[Els18] E. A. Elsayed. Book review: Probabilistic Design for Optimisation and Robustness for Engineers, Bryan Dodson, Patrick C. Hammett and Rene Klerx, John Wiley and Sons, 2014, 272 pages, £65.50, hardcover ISBN: 978-1-118-79619-1. International Statistical Review = Revue Internationale de Statistique, 86(3):601–602, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Elsayed:2020:BRR

[Els20] Elsayed Elsayed. Book review: The Road to Quality Control, Homer M. Sarasohn, Translated by N. I. Fisher and Y. Tanaka from the original Japanese text by Kagaku Shinko Sha, John Wiley and Sons, 2019, 160 pages, £89.95, hardcover, ISBN: 978-1-119-51493-0. International Statistical Review = Revue Internationale de Statistique, 88(1):254–255, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Eddy:1993:EDG

[EM93] William F. Eddy and Audris Mockus. An example of dynamic graphics for manufacturing process data. International Statistical Review = Revue Internationale de Statistique, 61(1):81–95, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403595.

Eckardt:2019:AMS

[EM19] Matthias Eckardt and Jorge Mateu. Analysing multivariate spatial point processes with continuous marks: a graphical modelling approach. International Statistical Review = Revue Internationale de Statistique, 87(1):44– 67, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Embrechts:2008:SBR

[Emb08] Paul Embrechts. Short book review: Statistics of Financial Markets: An Introduction, 2nd Edition, by Jürgen Franke, Wolfgang K. Härdle, Christian M. Hafner. International Statistical Review = Revue Internationale de Statistique, 76(2):313–314, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919633.

Einmahl:1988:PEM

[EMR88] J. H. J. Einmahl, D. M. Mason, and F. H. Ruymgaart. Processus empiriques multidimensionnels: aperçu de quelques résultats récents. (French) [Multidimensional empirical processes: overview of some recent results]. International Statistical Review = Revue Internationale de Statistique, 56(2):139–151, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403638.

Elvira:2022:RES

[EMR22] Víctor Elvira, Luca Martino, and Christian P. Robert. Rethinking the effective sample size. International Statistical Review = Revue Internationale de Statistique, 90(3):525– 550, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Engen:1978:BRB

[Eng78] Steinar Engen. Book review: Statistical Distributions in Scientific Work: Volume II: Model Building and Model Selection by G. P. Patil; S. Kotz; J. K. Ord. International Statistical Review = Revue Internationale de Statistique, 46(3):321, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402385.

Engle:1983:DPD

[Eng83] Robert F. Engle. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):149–150, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402740.

Erto:2008:EDT

[EPP08] Pasquale Erto, Giuliana Pallotta, and Sung H. Park. An example of data technology product: A control chart for Weibull processes. International Statistical Review = Revue Internationale de Statistique, 76(2):157–166, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919609.

Eriksson:1973:NMR

[Eri73] Sven A. Eriksson. A new model for randomized response. International Statistical Review = Revue Internationale de Statistique, 41(1):101–113, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402791.

Esenwein-Rothe:1977:HK

[ERS77] I. Esenwein-Rothe and H. Strecker. Hans Kellerer, 1902– 1976. International Statistical Review = Revue Internationale de Statistique, 45(1):108–109, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403029.

Estevao:2006:SEC

[ES06] Victor M. Estevao and Carl-Erik Särndal. Survey estimates by calibration on complex auxiliary information. International Statistical Review = Revue Internationale de Statistique, 74(2):127–147, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472698.

Espejo:2011:SBR

[Esp11] Mariano Ruiz Espejo. Short book review: Statistique. La théorie et ses applications, Deuxième édition, by Michel Lejeune. International Statistical Review = Revue Internationale de Statistique, 79(3):501, December 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305082.

Espejo:2014:BRB

[Esp14a] Mariano Ruiz Espejo. Book review: Handling Missing Data in Ranked Set Sampling by Carlos N. Bouza-Herrera. International Statistical Review = Revue Internationale de Statistique, 82(1):157–158, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Espejo:2014:BRI

[Esp14b] Mariano Ruiz Espejo. Book review: Introduction to Statistical Methods for Biosurveillance by Ronald D. Fricker. International Statistical Review = Revue Internationale de Statistique, 82(2):312, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Espejo:2015:BRSa

[Esp15a] Mariano Ruiz Espejo. Book review: Statistical Methods for Ranking Data. International Statistical Review = Revue Internationale de Statistique, 83(1):172–173, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Espejo:2015:BRSb

[Esp15b] Mariano Ruiz Espejo. Book review: Statistical Studies of Income, Poverty and Inequality in Europe: Computing and Graphics in R Using EU-SILC. International Statistical Review = Revue Internationale de Statistique, 83(1): 173–174, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Espejo:2015:BRSc

[Esp15c] Mariano Ruiz Espejo. Book review: Statistical Theory and Inference. International Statistical Review = Revue Internationale de Statistique, 83(1):174, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ebrahimi:2010:IMP

[ESS10] Nader Ebrahimi, Ehsan S. Soofi, and Refik Soyer. Information measures in perspective. International Statistical Review = Revue Internationale de Statistique, 78(3):383– 412, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919860.

El-Shaarawi:2005:ESC

[ESTTD05] Abdel H. El-Shaarawi, Jef Teugels, Dennis Trewin, and Donald A. Dawson. Environmental statistics: Current and future. International Statistical Review = Revue Internationale de Statistique, 73(2):233–236, August 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472662.

Evans:2016:FCN

[ET16] Scott R. Evans and Naitee Ting. Fundamental Concepts for New Clinical Trialists. Chapman and Hall/CRC biostatistics series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4200-9087-9 (hardcover). xix + 348 pp. LCCN R853.C55 E936 2016.

Eldridge:2009:ICC

[EUC09] Sandra M. Eldridge, Obioha C. Ukoumunne, and John B. Carlin. The intra-cluster correlation coefficient in cluster randomized trials: A review of definitions. International Statistical Review = Revue Internationale de Statistique, 77(3):378–394, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919763.

Egermayer:1973:BRB

[EV73] F. Egermayer and J. Vysušil. Book review: System von Rechnungsführung und Statistik in der Industrie by G. Forbrig; G. Goll; E. Polaschewski. International Statistical Review = Revue Internationale de Statistique, 41(1): 142–143, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402801.

Evans:2007:SBRb

[Eva07a] Michael Evans. Short book review: A Handbook of Statistical Analyses Using Stata, 4th Edition, by Sophia Rabe-Hesketh, Brian S. Everitt. International Statistical Review = Revue Internationale de Statistique, 75(2):251, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508466.

Evans:2007:SBRa

[Eva07b] Michael Evans. Short book review: Applied Statistics with SPSS by Eelko Huizingh. International Statistical Review = Revue Internationale de Statistique, 75(2):250, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508465.

Faleschini:1972:ABD

[Fal72] Luigi Faleschini. The applicability of Bayes decisions Scheme. International Statistical Review = Revue Internationale de Statistique, 40(1):41–45, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402704.

Farebrother:2013:SBR

[Far13] Richard William Farebrother. Short book review: In the Dark on the Sunny Side: A Memoir of an Out-of-Sight Mathematician by Larry W. Baggett. International Statistical Review = Revue Internationale de Statistique, 81 (2):322–323, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Forsythe:1973:SEU

[FB73] Alan B. Forsythe and Donald Bleich. Some experiences using computers to assist in the teaching of basic statistics. International Statistical Review = Revue Internationale de Statistique, 41(2):245–252, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402841.

Fraser:1996:SRP

[FB96] D. A. S. Fraser and George A. Barnard. Some remarks on pivotal models and the fiducial argument in relation to structural models. International Statistical Review = Revue Internationale de Statistique, 64(2):231–236, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403744.

Febrero-Bande:2017:FPC

[FBGGM17] Manuel Febrero-Bande, Pedro Galeano, and Wenceslao González-Manteiga. Functional principal component regression and functional partial least-squares regression: An overview and a comparative study. International Statistical Review = Revue Internationale de Statistique, 85 (1):61–83, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fragoso:2018:BMA

[FBL18] Tiago M. Fragoso, Wesley Bertoli, and Francisco Louzada. Bayesian model averaging: A systematic review and conceptual classification. International Statistical Review = Revue Internationale de Statistique, 86(1):1–28, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Foreman:2003:IDE

[FCE^{+}03] L. A. Foreman, C. Champod, I. W. Evett, J. A. Lambert, and S. Pope. Interpreting DNA evidence: A review. International Statistical Review = Revue Internationale de Statistique, 71(3):473–495, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403824.

Fouskakis:2002:SOR

[FD02] Dimitris Fouskakis and David Draper. Stochastic optimization: A review. International Statistical Review = Revue Internationale de Statistique, 70(3):315–349, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403861.

Federer:1980:SRR

[Fed80] Walter T. Federer. Some recent results in experiment design with a bibliography. I. International Statistical Review = Revue Internationale de Statistique, 48(3):357– 368, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402947.

Federer:1981:SRRa

[Fed81a] Walter T. Federer. Some recent results in experiment design with a bibliography. II. Bibliography A–K. International Statistical Review = Revue Internationale de Statistique, 49(1):95–109, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403040.

Federer:1981:SRRb

[Fed81b] Walter T. Federer. Some recent results in experiment design with a bibliography. III. Bibliography L–Z. International Statistical Review = Revue Internationale de Statistique, 49(2):185–197, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403072.

Fehr:2007:SBR

[Feh07] Max Fehr. Short book review: Optimization Methods in Finance by Gerard Cornuejols, Reha Tutuncu. International Statistical Review = Revue Internationale de Statistique, 75(3):420, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41509891.

Fellegi:1973:EAS

[Fel73] Ivan P. Fellegi. The evaluation of the accuracy of survey results: Some Canadian experiences. International Statistical Review = Revue Internationale de Statistique, 41 (1):1–14, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402784.

Fellegi:1996:CES

[Fel96] Ivan P. Fellegi. Characteristics of an effective statistical system. International Statistical Review = Revue Internationale de Statistique, 64(2):165–187, August 1996. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403738.

Fellegi:2004:OSP

[Fel04] Ivan P. Fellegi. Official statistics — pressures and challenges ISI President’s invited lecture, 2003. International Statistical Review = Revue Internationale de Statistique, 72(1):139–155, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403848.

Feraud:1972:DDP

[Fér72] Lucien Féraud. Dialectique du probable. (French) [Dialectic of the probable]. International Statistical Review = Revue Internationale de Statistique, 40(2):173–200, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402760.

Ferschl:1973:BRB

[Fer73] F. Ferschl. Book review: The Statistical Argument by R. Cormack. International Statistical Review = Revue Internationale de Statistique, 41(2):291–292, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402853.

Feraud:1978:LS

[Fér78] Lucien Féraud. Luigi Solari, 1932–1977. International Statistical Review = Revue Internationale de Statistique, 46(3):326–327, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402396.

Ferschl:1980:BRBb

[Fer80a] F. Ferschl. Book review: Problems in Linear and Non-Linear Programming by S. Vajda. International Statistical Review = Revue Internationale de Statistique, 48(3): 374, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402960.

Ferschl:1980:BRBa

[Fer80b] Franz Ferschl. Book review: Statistical Forecasting by Warren Gilchrist. International Statistical Review = Revue Internationale de Statistique, 48(3):370–371, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402952.

Fergany:1994:QLI

[Fer94] Nader Fergany. Quality of life indices for Arab countries in an international context. International Statistical Review = Revue Internationale de Statistique, 62(2): 187–202, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403508.

Feuerverger:1993:CTB

[Feu93] Andrey Feuerverger. A consistent test for bivariate dependence. International Statistical Review = Revue Internationale de Statistique, 61(3):419–433, December 1993. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403753.

Feuerverger:2016:GFO

[Feu16] Andrey Feuerverger. On goodness of fit for operational risk. International Statistical Review = Revue Internationale de Statistique, 84(3):434–455, December 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Freeman:1976:SMA

[FFBK76] Daniel H. Freeman, Jr., Jean L. Freeman, Dwight B. Brock, and Gary G. Koch. Strategies in the multivariate analysis of data from complex surveys II: An application to the United States National Health Interview Survey. International Statistical Review = Revue Internationale de Statistique, 44(3):317–330, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402347.

Finlay:2011:AF

[FFS11] Richard Finlay, Thomas Fung, and Eugene Seneta. Autocorrelation functions. International Statistical Review = Revue Internationale de Statistique, 79(2):255–271, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41305024.

Freiberger:1972:CPS

[FG72] Walter Freiberger and Ulf Grenander. Computational probability and statistics. International Statistical Review = Revue Internationale de Statistique, 40(1):67–72, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402707.

Fuentes:2003:SAN

[FGC03] Montserrat Fuentes, Peter Guttorp, and Peter Challenor. Statistical assessment of numerical models. International Statistical Review = Revue Internationale de Statistique, 71(2):201–221, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403884.

Forte:2018:MTB

[FGDS18] Anabel Forte, Gonzalo Garcia-Donato, and Mark Steel. Methods and tools for Bayesian variable selection and model averaging in normal linear regression. International Statistical Review = Revue Internationale de Statistique, 86(2):237–258, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Freijeiro-Gonzalez:2022:CRL

[FGFBGM22] Laura Freijeiro-González, Manuel Febrero-Bande, and Wenceslao González-Manteiga. A critical review of LASSO and its derivatives for variable selection under dependence among covariates. International Statistical Review = Revue Internationale de Statistique, 90(1):118– 145, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fix:1989:DAN

[FH89] Evelyn Fix and J. L. Hodges, Jr. Discriminatory analysis. Nonparametric discrimination: Consistency properties. International Statistical Review = Revue Internationale de Statistique, 57(3):238–247, December 1989. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403797.

Friedman:2020:DPE

[FHT20] Jerome Friedman, Trevor Hastie, and Robert Tibshirani. Discussion of “Prediction, Estimation, and Attribution” by Bradley Efron. International Statistical Review = Revue Internationale de Statistique, 88(S1):S73–S74, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Faltin:1993:ASP

[FHTV93] Frederick W. Faltin, Gerald J. Hahn, William T. Tucker, and Scott A. Vander Wiel. Algorithmic statistical process control: Some practical observations. International Statistical Review = Revue Internationale de Statistique, 61 (1):67–80, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403594.

Filipiak:2009:SBRb

[Fil09a] Katarzyna Filipiak. Short book review: Design and Analysis of Experiments, Volume 1: Introduction to Experimental Design, Second Edition, by Klaus Hinkelmann, Oscar Kempthorne. International Statistical Review = Revue Internationale de Statistique, 77(3):470, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919775.

Filipiak:2009:SBRa

[Fil09b] Katarzyna Filipiak. Short book review: Statistical Inference, Econometric Analysis and Matrix Algebra: Festschrift in Honour of Götz Trenkler by Bernhard Schipp, Walter Krämer. International Statistical Review = Revue Internationale de Statistique, 77(2):324–325, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919753.

Filipiak:2011:SBR

[Fil11] Katarzyna Filipiak. Short book review: Advances in Model-Oriented Design and Analysis edited by Alessandra Giovagnoli, Anthony C. Atkinson, Bernard Torsney. International Statistical Review = Revue Internationale de Statistique, 79(3):481–482, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305062.

Finney:1975:BRB

[Fin75] D. J. Finney. Book review: Statistics and Society: Data Collection and Interpretation by Walter T. Federer. International Statistical Review = Revue Internationale de Statistique, 43(2):244, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402914.

Finney:1977:BRBb

[Fin77a] D. J. Finney. Book review: Introduction to Probability and Statistics by W. Mendenhall. International Statistical Review = Revue Internationale de Statistique, 45 (1):100, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403018.

Finney:1977:BRBa

[Fin77b] D. J. Finney. Book review: Introduction to Statistics and Computer Programming by C. F. Kossack; C. I. Henschke. International Statistical Review = Revue Internationale de Statistique, 45(1):98, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403014.

Finney:1977:BRBc

[Fin77c] D. J. Finney. Book review: Statistical Theory of the Analysis of Experimental Designs by J. Ogawa. International Statistical Review = Revue Internationale de Statistique, 45(1):101–102, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403021.

Finney:1977:BRBd

[Fin77d] D. J. Finney. Book review: Statistics by Norma Gilbert. International Statistical Review = Revue Internationale de Statistique, 45(3):324, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402551.

Finney:1978:BRB

[Fin78] D. J. Finney. Book review: Statistical Theory by Bernard W. Lindgren. International Statistical Review = Revue Internationale de Statistique, 46(1):119, April 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402521.

Finney:1979:BPS

[Fin79] D. J. Finney. Bioassay and the practice of statistical inference. International Statistical Review = Revue Internationale de Statistique, 47(1):1–12, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403201.

Finney:1980:BRBa

[Fin80a] D. J. Finney. Book review: Introduction to Statistics by Robert Fried. International Statistical Review = Revue Internationale de Statistique, 48(1):122, April 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402422.

Finney:1980:BRBd

[Fin80b] D. J. Finney. Book review: Introductory Statistics by T. H. Wonnacott; R. J. Wonnacott. International Statistical Review = Revue Internationale de Statistique, 48(2): 244, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403185.

Finney:1980:BRBc

[Fin80c] D. J. Finney. Book review: Pocket Book of Statistical Tables by Robert E. Odeh; Donald B. Owen; Z. W. Birnbaum; Lloyd Fisher. International Statistical Review = Revue Internationale de Statistique, 48(1):131, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402435.

Finney:1980:BRBb

[Fin80d] D. J. Finney. Book review: Selection in One- and Two-Locus Systems by Thomas Nagylaki. International Statistical Review = Revue Internationale de Statistique, 48 (1):131, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402434.

Fisk:1972:BRB

[Fis72] P. R. Fisk. Book review: Individual Welfare Functions and Consumer Behaviour by B. M. S. van Praag. International Statistical Review = Revue Internationale de Statistique, 40(1):102, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402716.

Fisher:1983:GMN

[Fis83] Nicholas I. Fisher. Graphical methods in nonparametric statistics: a review and annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 51(1):25–58, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402730.

Fisher:2001:CIS

[Fis01] Nicholas I. Fisher. Crucial issues for statistics in the next two decades. International Statistical Review = Revue Internationale de Statistique, 69(1):3–4, April 2001. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403523.

Fisher:2005:PAP

[Fis05] Nicholas I. Fisher. Public awareness, professionalism & professional accreditation; international collaboration. International Statistical Review = Revue Internationale de Statistique, 73(2):269–270, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472667.

Fisher:2009:HSA

[Fis09] Nicholas I. Fisher. Homer Sarasohn and American involvement in the evolution of quality management in Japan, 1945–1950. International Statistical Review = Revue Internationale de Statistique, 77(2):276–299, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919728.

Fischer:2010:SBR

[Fis10a] Daniel Fischer. Short book review: Applied Statistical Genetics with R: For Population-based Association Studies by Andrea S. Foulkes. International Statistical Review = Revue Internationale de Statistique, 78(1):151–152, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919817.

Fisher:2010:ICJ

[Fis10b] Nicholas I. Fisher. Interview: a conversation with Jef Teugels. International Statistical Review = Revue Internationale de Statistique, 78(3):419–441, December 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919862.

Fischer:2018:BRC

[Fis18] Daniel Fischer. Book review: Correspondence Analysis-Theory, Practice and New Strategies, Eric J. Beh and Rosaria Lombardo, Wiley Series in Probability and Statistics, 2014, 592 pages, 82.20EUR, hardcover ISBN: 978-1-119-95324-1. International Statistical Review = Revue Internationale de Statistique, 86(3):603–604, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fischer:2019:BRD

[Fis19] Daniel Fischer. Book review: Data Science with Julia, Paul D. McNicholas and Peter A. Tait, CRC Press, 2019, 220 pages, £37.59, paperback, ISBN: 978-1-138-49998-0. International Statistical Review = Revue Internationale de Statistique, 87(2):445–446, August 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fischer:2022:BRP

[Fis22] Daniel Fischer. Book review: Population Genomics with R, Emmanuel Paradis, Chapman and Hall/CRC, 2020, 394 pages, $120, hardback. ISBN: 978-1-1386-0818-4. International Statistical Review = Revue Internationale de Statistique, 90(1):184–185, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fischer:2024:BRI

[Fis24] Daniel Fischer. Book review: An Introduction to R and Python for Data Analysis: a Side-by-Side Approach Taylor R. Brown Chapman and Hall/CRC, 2023, 246 pages (hardback $99.95, ebook $74.96) ISBN 978-10322032-56. International Statistical Review = Revue Internationale de Statistique, 92(1):132–134, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fox:1983:DSF

[FL83] Wendy R. Fox and Gabriel W. Lasker. The distribution of surname frequencies. International Statistical Review = Revue Internationale de Statistique, 51(1):81–87, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402733.

Fisher:2005:ISC

[FL05] N. I. Fisher and J. J. Lee. The 2004 ISI Special Conference: The Vital Role of Statistical Science in Assuring National Prosperity: August 29–31, 2004 Daejon, Republic of Korea. International Statistical Review = Revue Internationale de Statistique, 73(2):153, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472648.

Fildes:1995:IEA

[FM95] Robert Fildes and Spyros Makridakis. The impact of empirical accuracy studies on time series analysis and forecasting. International Statistical Review = Revue Internationale de Statistique, 63(3):289–308, December 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403481.

Flournoy:2020:PMD

[FMP20] Nancy Flournoy, José Moler, and Fernando Plo. Performance measures in dose-finding experiments. International Statistical Review = Revue Internationale de Statistique, 88(3):728–751, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Florens:1999:SNP

[FMR99] J.-P. Florens, M. Mouchart, and J.-M. Rolin. Semi- and non-parametric Bayesian analysis of duration models with Dirichlet priors: A survey. International Statistical Review = Revue Internationale de Statistique, 67(2):187– 210, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403398.

Foss:2019:DMC

[FMR19] Alexander H. Foss, Marianthi Markatou, and Bonnie Ray. Distance metrics and clustering methods for mixed-type data. International Statistical Review = Revue Internationale de Statistique, 87(1):80–109, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Farr:2020:COU

[FMR^{+}20] A. Charisse Farr, Kerrie Mengersen, Fabrizio Ruggeri, Daniel Simpson, Paul Wu, and Prasad Yarlagadda. Combining opinions for use in Bayesian networks: A measurement error approach. International Statistical Review = Revue Internationale de Statistique, 88(2):335–353, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fienberg:1995:CNS

[FMS95] Stephen E. Fienberg, Margaret E. Martin, and Miron L. Straf. The Committee on National Statistics: Fostering interactions between statisticians in academia and government. International Statistical Review = Revue Internationale de Statistique, 63(3):257–269, December 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403479.

Flournoy:2012:AOR

[FMS12a] Nancy Flournoy, Caterina May, and Piercesare Secchi. Asymptotically optimal response-adaptive designs for allocating the best treatment: an overview. International Statistical Review = Revue Internationale de Statistique, 80(2):293–305, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Friedl:2012:MFG

[FMS12b] Herwig Friedl, Radoslava Mirkov, and Ansgar Steinkamp. Modelling and forecasting gas flow on exits of gas transmission networks. International Statistical Review = Revue Internationale de Statistique, 80(1):24–39, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257165.

Food:1973:STR

[FOK73] Food, Agriculture Organization, and C. C. Kisiel. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 41(2):283–284, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402848.

Fomenko:1988:ESM

[Fom88] A. T. Fomenko. Empirico-statistical methods in ordering narrative texts. International Statistical Review = Revue Internationale de Statistique, 56(3):279–301, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403354.

Fonseca:2011:SBR

[Fon11] Miguel Fonseca. Short book review: Design of Experiments: An Introduction Based on Linear Models by Max D. Morris. International Statistical Review = Revue Internationale de Statistique, 79(1):139–140, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41306191.

Fonseca:2013:SBR

[Fon13] Miguel Fonseca. Short book review: Principles and Practice of Structural Equation Modeling, Third Edition by Rex B. Kline. International Statistical Review = Revue Internationale de Statistique, 81(1):172–173, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Forcina:1982:GCT

[For82] Antonio Forcina. Gini’s contributions to the theory of inference. International Statistical Review = Revue Internationale de Statistique, 50(1):65–70, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402459.

Fortuin:2022:PBD

[For22] Vincent Fortuin. Priors in Bayesian deep learning: a review. International Statistical Review = Revue Internationale de Statistique, 90(3):563–591, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fourastie:1980:BRBb

[Fou80a] Jacqueline Fourastié. Book review: On the Factorial Approach Providing the True Index of Cost of Living by K. S.

Banerjee. International Statistical Review = Revue Internationale de Statistique, 48(3):369, December 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402948.

Fourastie:1980:BRBa

[Fou80b] Jacqueline Fourastié. Book review: Vingt cinq ans d’Économie Électrique by G. Morlat; F. Bessiere. International Statistical Review = Revue Internationale de Statistique, 48(2):240–241, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403177.

Fisher:2005:ISD

[FPD05] Nicholas I. Fisher, Sung H. Park, and Jayant Deshpande. The industrial statistics divide: Are statisticians really relevant to business and industry? International Statistical Review = Revue Internationale de Statistique, 73(2): 177–180, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472652.

Franke:2016:SIL

[FPR^{+}16] Beate Franke, Jean-François Plante, Ribana Roscher, En shiun Annie Lee, Cathal Smyth, Armin Hatefi, Fuqi Chen, Einat Gil, Alexander Schwing, Alessandro Selvitella, Michael M. Hoffman, Roger Grosse, Dieter Hendricks, and Nancy Reid. Statistical inference, learning and models in big data. International Statistical Review = Revue Internationale de Statistique, 84(3):371–389, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fraser:1972:BRB

[Fra72] D. A. S. Fraser. Book review: The Structure of Inference. International Statistical Review = Revue Internationale de Statistique, 40(2):259–260, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402778.

Frank:1980:SIP

[Fra80] Ove Frank. Sampling and inference in a population graph. International Statistical Review = Revue Internationale de Statistique, 48(1):33–41, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402403.

Fraser:2013:DCD

[Fra13] Donald A. S. Fraser. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1):42–48, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Freeman:1983:SPE

[Fre83] P. R. Freeman. The Secretary Problem and its extensions: A review. International Statistical Review = Revue Internationale de Statistique, 51(2):189–206, August 1983. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402748.

Friedman:1993:SCU

[Fri93] D. J. Friedman. Some considerations in the use of quality control techniques in integrated circuit fabrication. International Statistical Review = Revue Internationale de Statistique, 61(1):97–107, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403596.

Friedman:2001:RSD

[Fri01] Jerome H. Friedman. The role of statistics in the data revolution? International Statistical Review = Revue Internationale de Statistique, 69(1):5–10, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403524.

Frisen:2003:SSO

[Fri03] Marianne Frisén. Statistical surveillance. Optimality and methods. International Statistical Review = Revue Internationale de Statistique, 71(2):403–434, August 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403896.

Frosini:2006:CCM

[Fro06] Benito V. Frosini. Causality and causal models: A conceptual perspective. International Statistical Review = Revue Internationale de Statistique, 74(3):305–334, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472712.

Frolich:2008:PNR

[Frö08] Markus Frölich. Parametric and nonparametric regression in the presence of endogenous control variables. International Statistical Review = Revue Internationale de Statistique, 76(2):214–227, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919613.

Farrell:2008:MGL

[FRS08] Patrick J. Farrell and Katrina Rogers-Stewart. Methods for generating longitudinally correlated binary data. International Statistical Review = Revue Internationale de Statistique, 76(1):28–38, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919582.

Field:1994:REW

[FS94] C. Field and B. Smith. Robust estimation: A weighted maximum likelihood approach. International Statistical Review = Revue Internationale de Statistique, 62(3):405– 424, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403770.

Finlay:2008:SIV

[FS08] Richard Finlay and Eugene Seneta. Stationary-increment variance-gamma and t models: Simulation and parameter estimation. International Statistical Review = Revue Internationale de Statistique, 76(2):167–186, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919610.

Fung:2010:MEB

[FS10] Thomas Fung and Eugene Seneta. Modelling and estimation for bivariate financial returns. International Statistical Review = Revue Internationale de Statistique, 78(1): 117–133, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919798.

Furgal:2019:RCC

[FST19] Allison K. C. Furgal, Ananda Sen, and Jeremy M. G. Taylor. Review and comparison of computational approaches for joint longitudinal and time-to-event models. International Statistical Review = Revue Internationale de Statistique, 87(2):393–418, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fienberg:1987:ESS

[FT87] Stephen E. Fienberg and Judith M. Tanur. Experimental and sampling structures: parallels diverging and meeting. International Statistical Review = Revue Internationale de Statistique, 55(1):75–96, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403272.

Fienberg:1988:AES

[FT88] S. E. Fienberg and J. M. Tanur. An amendment to: “Experimental and sampling structures: parallels diverging and meeting”. International Statistical Review = Revue Internationale de Statistique, 56(2):197, 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fienberg:1996:RFC

[FT96] Stephen E. Fienberg and Judith M. Tanur. Reconsidering the fundamental contributions of Fisher and Neyman on experimentation and sampling. International Statistical Review = Revue Internationale de Statistique, 64(3):237– 253, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403784.

Fuller:1995:EPM

[Ful95] Wayne A. Fuller. Estimation in the presence of measurement error. International Statistical Review = Revue Internationale de Statistique, 63(2):121–141, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403606.

Furst:1973:HL

[Für73a] Gerhard Fürst. Hans Langelütke, 1892–1972. International Statistical Review = Revue Internationale de Statistique, 41(1):147, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402805.

Furst:1973:OM

[Für73b] Gerhard Fürst. Otto Most, 1881–1971. International Statistical Review = Revue Internationale de Statistique, 41 (1):147, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402806.

Furst:1976:ACJ

[Für76] Gerhard Fürst. Alfred Carl Joachim Heinrich Jacobs, 1897–1975. International Statistical Review = Revue Internationale de Statistique, 44(1):179–180, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402979.

Fisher:2015:IIJ

[FvZ15] Nicholas I. Fisher and Willem R. van Zwet. Interview: An interview with Jae C. Lee. International Statistical Review = Revue Internationale de Statistique, 83(3):371– 404, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Falta:2004:RDS

[FW04] Michael Falta and Rodney Wolff. Recent developments of statistical approaches in aspects of accounting: A review. International Statistical Review = Revue Internationale de Statistique, 72(3):377–396, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472629.

Ficek:2021:ESD

[FWC^{+}21a] Joseph Ficek, Wei Wang, Henian Chen, Getachew Dagne, and Ellen Daley. Erratum: A survey of differentially private regression for clinical and epidemiological research. International Statistical Review = Revue Internationale de Statistique, 89(2):433, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [FWC^{+}21b].

Ficek:2021:SDP

[FWC^{+}21b] Joseph Ficek, Wei Wang, Henian Chen, Getachew Dagne, and Ellen Daley. A survey of differentially private regression for clinical and epidemiological research. International Statistical Review = Revue Internationale de Statistique, 89(1):132–147, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See erratum [FWC^{+}21a].

Fung:2014:DIC

[FWS14] Thomas Fung, Joanna J. J. Wang, and Eugene Seneta. The deviance information criterion in comparison of normal mixing models. International Statistical Review = Revue Internationale de Statistique, 82(3):411–421, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Fullerton:2016:ORM

[FX16] Andrew S. Fullerton and Jun Xu. Ordered regression models: parallel, partial, and non-parallel alternatives. Chapman and Hall/CRC statistics in the social and behavioral sciences. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4665-6973-5. xv + 171 pp. LCCN QA278.2 .F85 2016.

Fiori:2009:KPO

[FZ09] Anna M. Fiori and Michele Zenga. Karl Pearson and the origin of kurtosis. International Statistical Review = Revue Internationale de Statistique, 77(1):40–50, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919689.

Fan:2015:DVC

[FZ15] Jianqing Fan and Wenyang Zhang. Discussions on varying coefficient regression models: A review and new developments: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(1):65–68, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Galambos:1980:BRBa

[Gal80a] J. Galambos. Book review: Random Allocations by Valentin F. Kolchin; Boris A. Sevast’yanov; Vladimir P. Chistyakov. International Statistical Review = Revue Internationale de Statistique, 48(2):238–239, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403175.

Galambos:1980:BRBb

[Gal80b] J. Galambos. Book review: Thinning of Renewal Point Processes. A Flow Graph Study by Lennart Rade. International Statistical Review = Revue Internationale de Statistique, 48(2):243, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403181.

Gal:2002:ASL

[Gal02a] Iddo Gal. Adults’ statistical literacy: Meanings, components, responsibilities. International Statistical Review = Revue Internationale de Statistique, 70(1):1–25, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403713.

Gal:2002:RDS

[Gal02b] Iddo Gal. Response: Developing statistical literacy: Towards implementing change. International Statistical Review = Revue Internationale de Statistique, 70(1):46–51, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403721.

Gani:1972:BRB

[Gan72a] J. Gani. Book review: Proceedings of the Fourth International Conference on Operational Research by D. B. Hertz; J. Melese. International Statistical Review = Revue Internationale de Statistique, 40(1):102–104, April 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402717.

Gani:1972:SCB

[Gan72b] J. Gani. Some comments on the bibliography of probability and statistics. International Statistical Review = Revue Internationale de Statistique, 40(2):201–207, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402761.

Gani:1976:BRB

[Gan76] J. Gani. Book review: Statistics at the School Level by Lennart Råde. International Statistical Review = Revue Internationale de Statistique, 44(1):170, April 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402973.

Gani:1978:BRBc

[Gan78a] J. Gani. Book review: Education and Population; Mutual Impacts by H. V. Muhsam. International Statistical Review = Revue Internationale de Statistique, 46(1): 121, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402524.

Gani:1978:BRBa

[Gan78b] J. Gani. Book review: Fourier Analysis of Time Series: An Introduction by P. Bloomfield. International Statistical Review = Revue Internationale de Statistique, 46(1): 116–117, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402516.

Gani:1978:BRBb

[Gan78c] J. Gani. Book review: Queueing Models Involving Buffers by F. Göbel. International Statistical Review = Revue Internationale de Statistique, 46(1):118, April 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402518.

Gani:1979:RIF

[Gan79] J. Gani. The role of ISI in the field of statistical education: 1949–1979. International Statistical Review = Revue Internationale de Statistique, 47(3):237–244, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402646.

Gani:1980:BRBb

[Gan80a] J. Gani. Book review: Mathematics, Statistics and Systems for Health by N. T. J. Bailey. International Statistical Review = Revue Internationale de Statistique, 48(1): 118–119, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402415.

Gani:1980:BRBd

[Gan80b] J. Gani. Book review: Probabilistic Analysis and Related Topics, Volume 1 by A. T. Bharucha-Reid. International Statistical Review = Revue Internationale de Statistique, 48(2):231, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403160.

Gani:1980:BRBa

[Gan80c] J. Gani. Book review: Proceedings of the Fifth Conference on Probability Theory. 1–6 September 1974, Brasov, Romania by B. Bereanu; M. Iosifescu; G. Popescu. International Statistical Review = Revue Internationale de Statistique, 48(1):118, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402414.

Gani:1980:BRBc

[Gan80d] J. Gani. Book review: Statistical Models and Turbulence by M. Rosenblatt; C. van Atta. International Statistical Review = Revue Internationale de Statistique, 48(1): 131–132, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402436.

Gardner:1993:ICL

[Gar93] Martin J. Gardner. Investigating childhood leukaemia rates around the Sellafield nuclear plant. International Statistical Review = Revue Internationale de Statistique, 61(2):231–244, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403626.

Garfield:1995:HSL

[Gar95] Joan Garfield. How students learn statistics. International Statistical Review = Revue Internationale de Statistique, 63(1):25–34, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403775.

Garfield:1997:NPN

[Gar97] Joan Garfield. [new pedagogy and new content: The case of statistics]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2):137– 141, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403334.

Gay:2011:SBR

[Gay11] Roger Gay. Short book review: Selected Works of C. C. Heyde edited by Ross Maller, Ishwar Basawa, Peter Hall, Eugene Seneta. International Statistical Review = Revue Internationale de Statistique, 79(2):283, August 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305034.

Garfield:2007:HSL

[GBZ07] Joan Garfield and Dani Ben-Zvi. How students learn statistics revisited: A current review of research on teaching and learning statistics. International Statistical Review = Revue Internationale de Statistique, 75(3):372– 396, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509878.

Goutis:1995:FPD

[GC95] Constantinos Goutis and George Casella. Frequentist post-data inference. International Statistical Review = Revue Internationale de Statistique, 63(3):325–344, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403483.

Giorgi:2018:GMD

[GDSN18] Emanuele Giorgi, Peter J. Diggle, Robert W. Snow, and Abdisalan M. Noor. Geostatistical methods for disease mapping and visualisation using data from spatiotemporally referenced prevalence surveys. International Statistical Review = Revue Internationale de Statistique, 86(3):571–597, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2020:D

[GDV20] Soumya Ghosh and Finale Doshi-Velez. Discussion. International Statistical Review = Revue Internationale de Statistique, 88(2):324–326, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gelman:2003:BFE

[Gel03] Andrew Gelman. A Bayesian formulation of exploratory data analysis and goodness-of-fit testing. International Statistical Review = Revue Internationale de Statistique, 71(2):369–382, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403894.

Genest:2024:CMH

[Gen24] Christian Genest. A conversation with Marc Hallin. International Statistical Review = Revue Internationale de Statistique, 92(2):137–159, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Garfield:1999:ASE

[GG99] Joan B. Garfield and Iddo Gal. Assessment and statistics education: Current challenges and directions. International Statistical Review = Revue Internationale de Statistique, 67(1):1–12, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403562.

Gomes:2015:EVT

[GG15] M. Ivette Gomes and Armelle Guillou. Extreme value theory and statistics of univariate extremes: A review. International Statistical Review = Revue Internationale de Statistique, 83(2):263–292, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Giorgi:2017:CMC

[GG17] Giovanni M. Giorgi and Stefania Gubbiotti. Celebrating the memory of Corrado Gini: a personality out of the ordinary. International Statistical Review = Revue Internationale de Statistique, 85(2):325–339, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Godambe:1987:QLO

[GH87] V. P. Godambe and C. C. Heyde. Quasi-likelihood and optimal estimation, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 55(3):231–244, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403403.

Gorecki:2020:TNF

[GHK20] Tomasz Górecki, Lajos Horváth, and Piotr Kokoszka. Tests of normality of functional data. International Statistical Review = Revue Internationale de Statistique, 88 (3):677–697, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:1997:NPN

[Gho97] Jayanta K. Ghosh. [new pedagogy and new content: The case of statistics]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2):154– 155, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403338.

Ghosh:2008:SBRd

[Gho08a] Jayanta K. Ghosh. Short book review: Bayesian Networks and Decision Graphs, 2nd Edition, by Finn V. Jensen, Thomas D. Nielsen. International Statistical Review = Revue Internationale de Statistique, 76(2):307–308, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919627.

Ghosh:2008:SBRh

[Gho08b] Jayanta K. Ghosh. Short book review: Bayesian Networks and Influence Diagrams: A Guide to Construction and Analysis by Uffe B. Kjaerulff, Anders L. Madsen. International Statistical Review = Revue Internationale de Statistique, 76(3):461–462, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919681.

Ghosh:2008:SBRe

[Gho08c] Jayanta K. Ghosh. Short book review: Elementary Bayesian Biostatistics by Lemuel A. Moyé. International Statistical Review = Revue Internationale de Statistique, 76(2):310–311, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919630.

Ghosh:2008:SBRf

[Gho08d] Jayanta K. Ghosh. Short book review: Introduction to Applied Bayesian Statistics and Estimation for Social Scientists by Scott M. Lynch. International Statistical Review = Revue Internationale de Statistique, 76(2):311– 312, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919631.

Ghosh:2008:SBRb

[Gho08e] Jayanta K. Ghosh. Short book review: Introduction to Bayesian Statistics by Karl-Rudolf Koch. International Statistical Review = Revue Internationale de Statistique, 76(1):147–148, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919597.

Ghosh:2008:SBRa

[Gho08f] Jayanta K. Ghosh. Short book review: Multiscale Modeling: A Bayesian Perspective by Marco A. R. Ferreira, Herbert K. H. Lee. International Statistical Review = Revue Internationale de Statistique, 76(1):146–147, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919596.

Ghosh:2008:SBRc

[Gho08g] Jayanta K. Ghosh. Short book review: Probabilistic Networks and Expert Systems: Exact Computational Methods for Bayesian Networks by Robert G. Cowell, A. Philip Dawid, Steffen L. Lauritzen, David J. Spiegelhalter. International Statistical Review = Revue Internationale de Statistique, 76(2):306–307, August 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919626.

Ghosh:2008:SBRg

[Gho08h] Jayanta K. Ghosh. Short book review: Reliability and Risk: A Bayesian Perspective by Nozer D. Singpurwalla. International Statistical Review = Revue Internationale de Statistique, 76(2):316–317, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919637.

Ghosh:2009:SBRf

[Gho09a] Jayanta K. Ghosh. Short book review: Bayesian Disease Mapping: Hierarchical Modeling in Spatial Epidemiology by Andrew B. Lawson. International Statistical Review = Revue Internationale de Statistique, 77(2):325– 326, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919754.

Ghosh:2009:SBRc

[Gho09b] Jayanta K. Ghosh. Short book review: Bayesian Evaluation of Informative Hypotheses edited by Herbert Hoijtink, Irene Klugkist, Paul A. Boelen. International Statistical Review = Revue Internationale de Statistique, 77(2):308– 309, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919737.

Ghosh:2009:SBRa

[Gho09c] Jayanta K. Ghosh. Short book review: Bayesian Methods: A Social and Behavioral Sciences Approach, Second Edition, by Jeff Gill. International Statistical Review = Revue Internationale de Statistique, 77(2):301–302, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919730.

Ghosh:2009:SBRi

[Gho09d] Jayanta K. Ghosh. Short book review: Bayesian Reliability by Michael S. Hamada, Alyson G. Wilson, Shane C. Reese, Harry F. Martz. International Statistical Review = Revue Internationale de Statistique, 77(3):469– 470, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919774.

Ghosh:2009:SBRd

[Gho09e] Jayanta K. Ghosh. Short book review: Biostatistics and Microbiology: A Survival Manual by Daryl S. Paulson. International Statistical Review = Revue Internationale de Statistique, 77(2):312, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919741.

Ghosh:2009:SBRg

[Gho09f] Jayanta K. Ghosh. Short book review: Design and Analysis of Bioavailability and Bioequivalence Studies, Third Edition, by Shein-Chung Chow, Jen-pei Liu. International Statistical Review = Revue Internationale de Statistique, 77(2):326–328, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919755.

Ghosh:2009:SBRb

[Gho09g] Jayanta K. Ghosh. Short book review: Probability Models for DNA Sequence Evolution, Second Edition, by Richard Durrett. International Statistical Review = Revue Internationale de Statistique, 77(2):304, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919733.

Ghosh:2009:SBRh

[Gho09h] Jayanta K. Ghosh. Short book review: Survival and Event History Analysis: A Process Point of View by Odd O. Aalen, Ørnulf Borgan, Håkon K. Gjessing. International Statistical Review = Revue Internationale de Statistique, 77(3):463–464, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919769.

Ghosh:2009:SBRe

[Gho09i] Jayanta K. Ghosh. Short book review: The Statistical Analysis of Functional MRI Data by Nicole A. Lazar. International Statistical Review = Revue Internationale de Statistique, 77(2):312–313, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919742.

Ghosh:2010:SBRb

[Gho10a] Jayanta K. Ghosh. Short book review: Latent Class and Latent Transition Analysis: With Applications in the Social, Behavioral, and Health Sciences by Linda M. Collins, Stephanie T. Lanza. International Statistical Review = Revue Internationale de Statistique, 78(3):449–450, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919867.

Ghosh:2010:SBRa

[Gho10b] Jayanta K. Ghosh. Short book review: Philosophical Transactions of The Royal Society A, 367 (1906) Theme Issue ‘Statistical Challenges of High-dimensional Data’ edited by David L. Banks, Peter J. Bickel, Iain M. Johnstone, D. Michael Titterington. International Statistical Review = Revue Internationale de Statistique, 78(2):322– 323, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919848.

Ghosh:2011:SBRc

[Gho11a] Jayanta K. Ghosh. Short book review: Algebraic and Geometric Methods in Statistics edited by Paolo Gibilisco, Eva Riccomagno, Maria Piera Rogantin, Henry P. Wynn. International Statistical Review = Revue Internationale de Statistique, 79(2):281–282, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305033.

Ghosh:2011:SBRb

[Gho11b] Jayanta K. Ghosh. Short book review: Bayesian Analysis for Population Ecology by Ruth King, Byron J. T. Morgan, Olivier Gimenez, Stephen P. Brooks. International Statistical Review = Revue Internationale de Statistique, 79(1):130–131, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41306183.

Ghosh:2011:SBRd

[Gho11c] Jayanta K. Ghosh. Short book review: Causality: Models, Reasoning and Inference, Second Edition, by Judea Pearl. International Statistical Review = Revue Internationale de Statistique, 79(2):289–290, August 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305040.

Ghosh:2011:SBRa

[Gho11d] Jayanta K. Ghosh. Short book review: Maximum Penalized Likelihood Estimation, Volume II: Regression by Paul P. Eggermont, Vincent N. LaRiccia. International Statistical Review = Revue Internationale de Statistique, 79(1):123, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306176.

Ghosh:2012:SBRa

[Gho12a] Jayanta K. Ghosh. Short book review: Bayesian Modelling in Bioinformatics edited by Dipak K. Dey, Samiran Ghosh and Bani K. Mallick. International Statistical Review = Revue Internationale de Statistique, 80(2):332– 333, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRg

[Gho12b] Jayanta K. Ghosh. Short book review: Introduction to Modeling and Analysis of Stochastic Systems, Second Edition, by V. G. Kulkarni. International Statistical Review = Revue Internationale de Statistique, 80(3):487, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRb

[Gho12c] Jayanta K. Ghosh. Short book review: Probability and Statistical Models: Foundations for Problems in Reliability and Financial Mathematics by Arjun K. Gupta, Wei-Bin Zeng and Yanhong Wu. International Statistical Review = Revue Internationale de Statistique, 80(2):345–346, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRe

[Gho12d] Jayanta K. Ghosh. Short book review: Selected Works of Debabrata Basu edited by Anirban DasGupta. International Statistical Review = Revue Internationale de Statistique, 80(3):484–486, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRh

[Gho12e] Jayanta K. Ghosh. Short book review: Statistical Image Processing and Multidimensional Modeling by Paul Fieguth. International Statistical Review = Revue Internationale de Statistique, 80(3):492–493, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRc

[Gho12f] Jayanta K. Ghosh. Short book review: Statistics and Data Analysis for Financial Engineering by David Ruppert. International Statistical Review = Revue Internationale de Statistique, 80(2):347–348, August 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRf

[Gho12g] Jayanta K. Ghosh. Short book review: Statistics for High-Dimensional Data: Methods, Theory and Applications by Peter Bühlmann, Sara van de Geer. International Statistical Review = Revue Internationale de Statistique, 80(3): 486–487, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2012:SBRd

[Gho12h] Jayanta K. Ghosh. Short book review: Statistics of Financial Markets: An Introduction, Third Edition by Jürgen Franke, Wolfgang Karl Härdle, Christian Matthias Hafner. International Statistical Review = Revue Internationale de Statistique, 80(3):471–472, December 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2013:SBRc

[Gho13a] Jayanta K. Ghosh. Short book review: Handbook of Statistical Bioinformatics by Henry Horng-Shing Lu, Bernhard Schölkopf, Hongyu Zhao. International Statistical Review = Revue Internationale de Statistique, 81(3):467–469, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2013:SBRa

[Gho13b] Jayanta K. Ghosh. Short book review: Handbook of Statistical Systems Biology by Michael P. H. Stumpf, David J. Balding, Mark Girolami. International Statistical Review = Revue Internationale de Statistique, 81(2):313–314, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2013:SBRb

[Gho13c] Jayanta K. Ghosh. Short book review: Physiology, Biophysics, and Biomedical Engineering by Andrew W. Wood with contributions by Anthony Bartel, Peter Cadusch, Joseph Ciorciari, David Crewther, Per Line, John Patterson, Mark Schier, and Bruce Thompson. International Statistical Review = Revue Internationale de Statistique, 81(3):466–467, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2013:SBRd

[Gho13d] Jayanta K. Ghosh. Short book review: Time Series Modeling of Neuroscience Data by Tohru Ozaki. International Statistical Review = Revue Internationale de Statistique, 81(3):475–476, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2017:BRS

[Gho17] Debashis Ghosh. Book review: Statistical Analysis of Proteomics, Metabolomics, and Lipidomics Data Using Mass Spectrometry, Susmita Datta and Bart J. A. Mertens Springer Frontiers in Probability and the Statistical Sciences, 2017, 295 pages, £82.00, hardcover ISBN: 978-3-319-45807-6. International Statistical Review = Revue Internationale de Statistique, 85(3):544–545, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2020:BRS

[Gho20] Debashis Ghosh. Book review: Statistical Inference via Convex Optimization. Nillas, Alice Anatoli Juditsky and Arkadi Nemirovski. Princeton University Press, 2020, xiv + 656 pages, £70/$85, hardcover ISBN: 978-0-691-19729-6. International Statistical Review = Revue Internationale de Statistique, 88(3):806–808, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2021:SGP

[Gho21] Debashis Ghosh. Surrogates: Gaussian process modeling, design, and optimization for the applied sciences. International Statistical Review = Revue Internationale de Statistique, 89(1):207–209, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2022:BRF

[Gho22] Debashis Ghosh. Book review: Fundamentals of Causal Inference with R, Babette A. Brumback, Chapman and Hall/CRC, 2021, xi + 236 pages, $69.95, hardcover. ISBN: 978-0-3677-0505-3. International Statistical Review = Revue Internationale de Statistique, 90(1):185–186, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:2023:BRD

[Gho23] Debashis Ghosh. Book review: Data Conscience: Algorithmic Siege on Our Humanity, Brandeis Hill Marshall, Wiley, 2022, xxv + 326 pages, paperback £30.99, ISBN: 978-1-119-82118-2. International Statistical Review = Revue Internationale de Statistique, 91(2):347–348, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gilchrist:2008:RR

[Gil08] Warren Gilchrist. Regression revisited. International Statistical Review = Revue Internationale de Statistique, 76 (3):401–418, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919654.

Gillespie:2015:BRS

[Gil15] Brenda W. Gillespie. Book review: Statistical Methods for Survival Data Analysis. International Statistical Review = Revue Internationale de Statistique, 83(1):167–168, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Giri:1973:BRB

[Gir73] N. Giri. Book review: Multivariate Analysis by M. M. Tatsuoka. International Statistical Review = Revue Internationale de Statistique, 41(2):297–298, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402862.

Giraud:2015:IHD

[Gir15] Christophe Giraud. Introduction to high-dimensional statistics, volume 139 of Monographs on statistics and applied probability. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4822-3794-6 (hardcover). xv + 255 pp. LCCN QC20.7.D55 G57 2015.

Gyllenberg:2001:PMB

[GK01] M. Gyllenberg and T. Koski. Probabilistic models for bacterial taxonomy. International Statistical Review = Revue Internationale de Statistique, 69(2):249–276, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403815.

Gu:2016:PR

[GK16] Jiaying Gu and Roger Koenker. On a problem of Robbins. International Statistical Review = Revue Internationale de Statistique, 84(2):224–244, August 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gijbels:2019:QBA

[GKV19] Irène Gijbels, Rezaul Karim, and Anneleen Verhasselt. On quantile-based asymmetric family of distributions: Properties and inference. International Statistical Review = Revue Internationale de Statistique, 87(3):471–504, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gijbels:2020:RLE

[GKV20] Irène Gijbels, Rezaul Karim, and Anneleen Verhasselt. Response to the letter to the Editor on ‘on quantile-based asymmetric family of distributions: Properties and inference’. International Statistical Review = Revue Internationale de Statistique, 88(3):797–801, December 2020. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Alv20].

Guttorp:2009:KPS

[GL09] Peter Guttorp and Georg Lindgren. Karl Pearson and the Scandinavian school of statistics. International Statistical Review = Revue Internationale de Statistique, 77(1): 64–71, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919691.

Grientschnig:2014:CPD

[GL14] Dieter Grientschnig and Ignacio Lira. Combining probability distributions by multiplication in metrology: A viable method? International Statistical Review = Revue Internationale de Statistique, 82(3):392–410, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Guttorp:2019:WDB

[GL19] Peter Guttorp and Georg Lindgren. Why distinguish between statistics and mathematical statistics-the case of Swedish academia. International Statistical Review = Revue Internationale de Statistique, 87(1):110–126, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Granovsky:1991:OKM

[GM91] Boris L. Granovsky and Hans-Georg Müller. Optimizing kernel methods: A unifying variational principle. International Statistical Review = Revue Internationale de Statistique, 59(3):373–388, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403693.

Guo:2019:IDC

[GM19] Lingzhe Guo and Reza Modarres. Interpoint distance classification of high dimensional discrete observations. International Statistical Review = Revue Internationale de Statistique, 87(2):191–206, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ghosh:1999:ESI

[GMRS99] J. K. Ghosh, P. Maiti, T. J. Rao, and B. K. Sinha. Evolution of statistics in India. International Statistical Review = Revue Internationale de Statistique, 67(1):13–34, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403563.

Gastwirth:2003:SIA

[GMZ03] Joseph L. Gastwirth, Weiwen Miao, and Gang Zheng. Statistical issues arising in disparate impact cases and the use of the expectancy curve in assessing the validity of pre-employment tests. International Statistical Review = Revue Internationale de Statistique, 71(3):565– 580, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403829.

Genest:2014:ICJ

[GN14] Christian Genest and Johanna G. Nešlehová. Interview: a conversation with James O. Ramsay. International Statistical Review = Revue Internationale de Statistique, 82(2): 161–183, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Genest:2020:CPE

[GN20] Christian Genest and Johanna G. Neslehová. A conversation with Paul Embrechts. International Statistical Review = Revue Internationale de Statistique, 88(3):521–547, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gosho:2021:PRC

[GNM21] Masahiko Gosho, Hisashi Noma, and Kazushi Maruo. Practical review and comparison of modified covariance estimators for linear mixed models in small-sample longitudinal studies with missing data. International Statistical Review = Revue Internationale de Statistique, 89(3):550– 572, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Groeneboom:1981:BES

[GO81] P. Groeneboom and J. Oosterhoff. Bahadur efficiency and small-sample efficiency. International Statistical Review = Revue Internationale de Statistique, 49(2):127– 141, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403067.

Gal:2010:QRS

[GO10] Iddo Gal and Irena Ograjenšek. Qualitative research in the service of understanding learners and users of statistics. International Statistical Review = Revue Internationale de Statistique, 78(2):287–296, August 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919838.

Gal:2016:RES

[GO16] Iddo Gal and Irena Ograjenšek. Rejoinder to enhancing statistics education by including qualitative research:

Rejoinder: More on enhancing statistics education with qualitative ideas. International Statistical Review = Revue Internationale de Statistique, 84(2):202–209, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gobet:2016:MCMg

[Gob16] Emmanuel Gobet. Monte-Carlo methods and stochastic processes: from linear to non-linear. Chapman and Hall CRC, Boca Raton, FL, USA, 2016. ISBN 1-4987-4622-5, 1-4987-4625-X (e-book), 1-4987-4623-3 (PDF), 1-4987-4624-1 (VitalBook), 1-315-36875-7 (e-book). xxv + 309 pp. LCCN QA274 .G63 2016.

Godambe:1980:ERR

[God80] V. P. Godambe. Estimation in randomised response trials. International Statistical Review = Revue Internationale de Statistique, 48(1):29–32, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402402.

Godsill:1997:BES

[God97] Simon J. Godsill. Bayesian enhancement of speech and audio signals which can be modelled as ARMA processes. International Statistical Review = Revue Internationale de Statistique, 65(1):1–21, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403430.

Goldfinger:1997:IEI

[Gol97] Charles Goldfinger. Intangible economy and its implications for statistics and statisticians. International Statistical Review = Revue Internationale de Statistique, 65(2): 191–220, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403343.

Goldstein:2004:SPP

[Gol04] Joshua R. Goldstein. Simpler probabilistic population forecasts: Making scenarios work. International Statistical Review = Revue Internationale de Statistique, 72(1): 93–106, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403845.

Good:1973:BRB

[Goo73] I. J. Good. Book review: The Generation of Random Variates by T. G. Newman; P. L. Odell. International Statistical Review = Revue Internationale de Statistique, 41(1):139–140, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402798.

Good:1978:AOT

[Goo78] I. J. Good. Alleged objectivity; a threat to the human spirit? A reply to “Bayesianism; a threat to the statistical profession?” (Internat. Statist. Rev. 45 (1977), 111– 115) by H. C. Hamaker. International Statistical Review = Revue Internationale de Statistique, 46(1):65–66, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402508.

Goodman:1986:SUEa

[Goo86a] Leo A. Goodman. Some useful extensions of the usual correspondence analysis approach and the usual log-linear models approach in the analysis of contingency tables. International Statistical Review = Revue Internationale de Statistique, 54(3):243–270, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403053. With a

discussion and reply by the author.

Goodman:1986:SUEb

[Goo86b] Leo A. Goodman. [some useful extensions of the usual correspondence analysis approach and the usual log-linear models approach in the analysis of contingency tables]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 54(3):291–309, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403060.

Goutis:1994:RPM

[Gou94] Constantinos Goutis. Ranges of posterior measures for some classes of priors with specified moments. International Statistical Review = Revue Internationale de Statistique, 62(2):245–256, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403511.

Gould:2010:SMS

[Gou10] Robert Gould. Statistics and the modern student. International Statistical Review = Revue Internationale de Statistique, 78(2):297–315, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919839.

Gower:1972:BRB

[Gow72] J. C. Gower. Book review: Multivariate Data Analysis by W. W. Cooley; P. R. Lohnes. International Statistical Review = Revue Internationale de Statistique, 40(3): 394, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402480.

Gower:1979:BRB

[Gow79] J. C. Gower. Book review: Minimizing the Sum of Absolute Deviations (No. 12 in the Series Applied Statistics and Econometrics) by V. Sposito; W. Smith; G. McCormick. International Statistical Review = Revue Internationale de Statistique, 47(3):299–300, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402652.

Gower:2015:DSC

[Gow15] John C. Gower. The development of statistical computing at Rothamsted. International Statistical Review = Revue Internationale de Statistique, 83(3):357–370, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Garber:1994:PHR

[GOZV94] Alan M. Garber, Richard A. Olshen, Heping Zhang, and E. S. Venkatraman. Predicting high-risk cholesterol levels. International Statistical Review = Revue Internationale de Statistique, 62(2):203–228, August 1994. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403509.

Gabriel:1973:LSM

[GP73] K. R. Gabriel and Eric Peritz. Least squares and maximal contrasts: An elementary general approach and its application to simultaneous inference in contingency tables. International Statistical Review = Revue Internationale de Statistique, 41(2):155–164, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402832.

Gault:2003:MDI

[GP03] Fred Gault and Greg Peterson. Measuring the diffusion of information and communication technology in society and its effects: Canadian experience. International Statistical Review = Revue Internationale de Statistique, 71 (1):49–57, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403873.

Greenland:2011:ATC

[GP11a] Sander Greenland and Judea Pearl. Adjustments and their consequences — collapsibility analysis using graphical models. International Statistical Review = Revue Internationale de Statistique, 79(3):401–426, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305058.

Gutierrez-Pena:2011:SBR

[GP11b] Eduardo Gutiérrez-Peña. Short book review: Bayesian Item Response Modeling: Theory and Applications by Jean-Paul Fox. International Statistical Review = Revue Internationale de Statistique, 79(3):480, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305061.

Gutierrez-Pena:2012:SBR

[GP12] Eduardo Gutiérrez-Peña. Short book review: Expansions and Asymptotics for Statistics by Christopher G. Small. International Statistical Review = Revue Internationale de Statistique, 80(2):327, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Garcia-Perez:2020:GIC

[GPdMLMGGSG20] José García-Pérez, María del Mar López-Martín, Catalina García-García, and Román Salmerón-Gómez. A geometrical interpretation of collinearity: A natural way to justify ridge regression and its anomalies. International Statistical Review = Revue Internationale de Statistique, 88(3): 776–792, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gutierrez-Pena:2005:SDP

[GPW05] Eduardo Gutiérrez-Peña and Stephen G. Walker. Statistical decision problems and Bayesian nonparametric methods. International Statistical Review = Revue Internationale de Statistique, 73(3):309–330, December 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472678.

Gilula:1990:IOC

[GR90] Zvi Gilula and Yaakov Ritov. Inferential ordinal correspondence analysis: Motivation, derivation and limitations. International Statistical Review = Revue Internationale de Statistique, 58(2):99–108, August 1990. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403461.

Gupta:2001:HDL

[GR01] Rameshwar D. Gupta and Donald St. P. Richards. The history of the Dirichlet and Liouville distributions. International Statistical Review = Revue Internationale de Statistique, 69(3):433–446, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403455.

Granger:1983:DPD

[Gra83] C. W. J. Granger. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):151–153, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402741.

Granquist:1997:NVE

[Gra97] Leopold Granquist. The new view on editing. International Statistical Review = Revue Internationale de Statistique, 65(3):381–387, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403378.

Gramacy:2020:D

[Gra20] Robert B. Gramacy. Discussion. International Statistical Review = Revue Internationale de Statistique, 88(2):326– 329, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Greenberg:1980:CB

[Gre80] B. G. Greenberg. Chester I. Bliss, 1899–1979. International Statistical Review = Revue Internationale de Statistique, 48(1):135–136, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402441.

Green:1987:PLG

[Gre87] Peter J. Green. Penalized likelihood for general semiparametric regression models. International Statistical Review = Revue Internationale de Statistique, 55(3):245– 259, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403404.

Greenland:2006:SOD

[Gre06] Sander Greenland. Smoothing observational data: A philosophy and implementation for the health sciences. International Statistical Review = Revue Internationale de Statistique, 74(1):31–46, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/25472689.

Grimshaw:1977:BRB

[Gri77] J. J. Grimshaw. Book review: Biostatics for Medical Students by S. K. Lwanga. International Statistical Review = Revue Internationale de Statistique, 45(3):325– 326, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402554.

Grimmett:1987:IPS

[Gri87] Geoffrey Grimmett. Interacting particle systems and random media: an overview. International Statistical Review = Revue Internationale de Statistique, 55(1):49–62, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403270.

Grize:2015:ASF

[Gri15] Yves L. Grize. Applications of statistics in the field of general insurance: An overview. International Statistical Review = Revue Internationale de Statistique, 83(1): 135–159, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Groves:1995:ORT

[Gro95] Robert M. Groves. Opening remarks: Third Morris Hansen Lecture. International Statistical Review = Revue Internationale de Statistique, 63(2):149–151, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403609.

Gill:1986:DPG

[GS86] R. D. Gill and B. F. Schriever. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):289–291, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403059.

Gill:1987:ADP

[GS87] R. D. Gill and B. F. Schriever. Amendment to discussion of paper by L. A. Goodman: “Some useful extensions of the usual correspondence analysis approach and the usual log-linear models approach in the analysis of contingency tables” [Internat. Statist. Rev. 54 (1986), no. 3, 243–309]. International Statistical Review = Revue Internationale de Statistique, 55(2):229, 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gibbs:2002:CBP

[GS02] Alison L. Gibbs and Francis Edward Su. On choosing and bounding probability metrics. International Statistical Review = Revue Internationale de Statistique, 70(3): 419–435, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403865.

Ghosh:2019:SVC

[GS19] Malay Ghosh and Rebecca C. Steorts. Some variants of constrained estimation in finite population sampling. International Statistical Review = Revue Internationale de Statistique, 87(S1):S90–S103, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Garcia:2016:SVC

[GSGdMLM16] José García, Román Salmerón, Catalina García, and María del Mar López Martín. Standardization of variables and collinearity diagnostic in ridge regression. International Statistical Review = Revue Internationale de Statistique, 84(2):245–266, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Godtliebsen:2017:DSS

[GSW17] Fred Godtliebsen, Stein O. Skrøvseth, and Susan Wei. Discussions on statistical scale space methods: Discussion. International Statistical Review = Revue Internationale de Statistique, 85(1):36–37, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Godambe:1986:PSS

[GT86] V. P. Godambe and M. E. Thompson. Parameters of superpopulation and survey population: their relationships and estimation. International Statistical Review = Revue Internationale de Statistique, 54(2):127–138, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403139.

Garonna:1999:SCM

[GT99] Paolo Garonna and Umberto Triacca. Social change: Measurement and theory. International Statistical Review = Revue Internationale de Statistique, 67(1):49–62, April 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403565.

Gertheiss:2009:PRO

[GT09] Jan Gertheiss and Gerhard Tutz. Penalized regression with ordinal predictors. International Statistical Review = Revue Internationale de Statistique, 77(3):345–365, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919761.

Guttorp:2012:WHD

[GT12] Peter Guttorp and Thordis L. Thorarinsdottir. What happened to discrete chaos, the Quenouille process, and the sharp Markov property? Some history of stochastic point processes. International Statistical Review = Revue Internationale de Statistique, 80(2):253–268, August 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Guandalini:2017:DBE

[GT17] Alessio Guandalini and Yves Tillé. Design-based estimators calibrated on estimated totals from multiple surveys. International Statistical Review = Revue Internationale de Statistique, 85(2):250–269, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gnaldi:2018:EFC

[GTF18] Michela Gnaldi, Venera Tomaselli, and Antonio Forcina. Ecological fallacy and covariates: New insights based on multilevel modelling of individual data. International Statistical Review = Revue Internationale de Statistique, 86 (1):119–135, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Guarini:1977:BB

[Gua77] R. Guarini. Benedetto Barberi, 1901–1976. International Statistical Review = Revue Internationale de Statistique, 45(1):107, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403028.

Guerrero:1990:TDT

[Gue90] Victor M. Guerrero. Temporal disaggregation of time series: An ARIMA-based approach. International Statistical Review = Revue Internationale de Statistique, 58(1): 29–46, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403472.

Guerrero:2008:ETP

[Gue08] Victor M. Guerrero. Estimating trends with percentage of smoothness chosen by the user. International Statistical Review = Revue Internationale de Statistique, 76(2):187– 202, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919611.

Gunderson:2015:BRB

[Gun15] Brenda Gunderson. Book review: Biostatistics Decoded. International Statistical Review = Revue Internationale de Statistique, 83(1):161–163, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Gustafson:2006:TMS

[Gus06] Karl Gustafson. The trigonometry of matrix statistics. International Statistical Review = Revue Internationale de Statistique, 74(2):187–202, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472702.

Guttorp:1995:I

[Gut95] Peter Guttorp. [introduction]. International Statistical Review = Revue Internationale de Statistique, 63(2): 233, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403616.

Guttorp:2003:ESP

[Gut03] Peter Guttorp. Environmental statistics — a personal view. International Statistical Review = Revue Internationale de Statistique, 71(2):169–179, August 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403882.

Grolemund:2014:CID

[GW14a] Garrett Grolemund and Hadley Wickham. A cognitive interpretation of data analysis. International Statistical Review = Revue Internationale de Statistique, 82 (2):184–204, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See discussion [Hub14, Wil14a] and rejoinder [GW14b].

Grolemund:2014:RCI

[GW14b] Garrett Grolemund and Hadley Wickham. Rejoinder to a cognitive interpretation of data analysis: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 82(2):210–213, August 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [GW14a].

Gao:2023:SVS

[GW23] Peter A. Gao and Jonathan Wakefield. A spatial variancesmoothing area level model for small area estimation of demographic rates. International Statistical Review = Revue Internationale de Statistique, 91(3):493–510, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Xie:2013:RCD

[gX13] Min ge Xie. Rejoinder: “Confidence distribution, the frequentist distribution estimator of a parameter: a review” [MR3047497; 3047498; MR3047499; MR3047500; MR3047501; MR3047502; 3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1): 68–77, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Xie:2013:CDF

[gXS13] Min ge Xie and Kesar Singh. Confidence distribution, the frequentist distribution estimator of a parameter: a review. International Statistical Review = Revue Internationale de Statistique, 81(1):3–39, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Xie:2020:DPB

[gXZ20] Min ge Xie and Zheshi Zheng. Discussion of Professor Bradley Efron’s Article on “Prediction, Estimation, and Attribution”. International Statistical Review = Revue Internationale de Statistique, 88(S1):S75–S82, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Haataja:2012:SBR

[Haa12] Juha Haataja. Short book review: The Crossing of Heaven: Memoirs of a Mathematician by Karl Gustafson. International Statistical Review = Revue Internationale de Statistique, 80(2):343–344, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Haber:1982:CCS

[Hab82] Michael Haber. The continuity correction and statistical testing. International Statistical Review = Revue Internationale de Statistique, 50(2):135–144, August 1982. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402597.

Hadi:2010:ACP

[Had10a] Ali S. Hadi. Announcements: Call for papers for a special issue of the International Statistical Review on energy statistics. International Statistical Review = Revue Internationale de Statistique, 78(1):2, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hadi:2010:CPS

[Had10b] Ali S. Hadi. Call for papers for a special issue of the international statistical review on Energy Statistics. International Statistical Review = Revue Internationale de Statistique, 78(1):2, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919791.

Hadi:2010:EAW

[Had10c] Ali S. Hadi. Editorial: Announcing Web-based submission process: ScholarOne manuscripts. International Statistical Review = Revue Internationale de Statistique, 78 (1):1, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919790.

Haggstrom:2007:SBR

[Häg07] Olle Häggström. Short book review: The Random-Cluster Model by Geoffrey Grimmett. International Statistical Review = Revue Internationale de Statistique, 75(2):273, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41508488.

Haggstrom:2012:SBRb

[Häg12a] Olle Häggström. Short book review: Simplicity, Complexity and Modelling edited by Mike Christie, Andrew Cliffe, Philip Dawid, Stephen S. Senn. International Statistical Review = Revue Internationale de Statistique, 80(3):472– 473, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Haggstrom:2012:SBRa

[Häg12b] Olle Häggström. Short book review: The R Primer by Claus Thorn Ekstrøm. International Statistical Review = Revue Internationale de Statistique, 80(1):197, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257192.

Haitovsky:1972:EMR

[Hai72] Yoel Haitovsky. On errors of measurements in regression analysis in economics. International Statistical Review = Revue Internationale de Statistique, 40(1):23–35, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402702.

Hajek:1972:BRB

[Háj72] J. Hájek. Book review: Numerical Approximation to Functions and Data by J. G. Hayes. International Statistical Review = Revue Internationale de Statistique, 40(2): 250–251, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402770.

Hajnal:1977:BRB

[Haj77] J. Hajnal. Book review: Markov Chains: Theory and Applications by D. L. Isaacson; R. W. Madsen. International Statistical Review = Revue Internationale de Statistique, 45(1):97–98, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403013.

Hall:1973:HPB

[Hal73] A. R. Hall. H. P. Brown, 1916–1971. International Statistical Review = Revue Internationale de Statistique, 41(3): 401, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402639.

Hald:1981:TCS

[Hal81] A. Hald. T. N. Thiele’s contributions to statistics. International Statistical Review = Revue Internationale de Statistique, 49(1):1–20 (one plate), April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403034.

Hald:1984:CMS

[Hal84a] A. Hald. A. de Moivre: ‘De Mensura Sortis’ or ‘On the Measurement of Chance’. International Statistical Review = Revue Internationale de Statistique, 52(3):229– 236, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403045.

Hald:1984:NBT

[Hal84b] A. Hald. Nicholas Bernoulli’s theorem. International Statistical Review = Revue Internationale de Statistique, 52 (1):93–99, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403248.

Hald:1986:GSA

[Hal86] A. Hald. Galileo’s statistical analysis of astronomical observations. International Statistical Review = Revue Internationale de Statistique, 54(2):211–220, August 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403145.

Hall:1989:CRN

[Hal89] Peter Hall. On convergence rates in nonparametric problems. International Statistical Review = Revue Internationale de Statistique, 57(1):45–58, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403583.

Hald:2000:EHC

[Hal00] Anders Hald. The early history of the cumulants and the Gram–Charlier series. International Statistical Review = Revue Internationale de Statistique, 68(2):137– 153, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403665.

Hamaker:1972:BRB

[Ham72a] H. C. Hamaker. Book review: Precision Measurement and Calibration: Statistical Concepts and Procedures by H. H. Ku. International Statistical Review = Revue Internationale de Statistique, 40(2):253, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402773.

Hamdan:1972:CEB

[Ham72b] M. A. Hamdan. Canonical expansion of the bivariate binomial distribution with unequal marginal indices. International Statistical Review = Revue Internationale de Statistique, 40(3):277–280, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402468.

Hamdan:1975:NTD

[Ham75] M. A. Hamdan. A note on the trinomial distribution. International Statistical Review = Revue Internationale de Statistique, 43(2):219–220, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402901.

Hamaker:1977:BTS

[Ham77a] Hugo C. Hamaker. Bayesianism; a threat to the statistical profession? International Statistical Review = Revue Internationale de Statistique, 45(2):111–115, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403310.

Hamaker:1977:SPE

[Ham77b] Hugo C. Hamaker. Subjective probabilities and exchangeability from an objective point of view. International Statistical Review = Revue Internationale de Statistique, 45 (3):223–231, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402536.

Hamdan:1978:SAT

[Ham78] M. A. Hamdan. A systematic approach to teaching counting formulae. International Statistical Review = Revue Internationale de Statistique, 46(2):219–220, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402815.

Hamaker:1979:BRB

[Ham79] H. C. Hamaker. Book review: The Statistical Treatment of Experimental Data by J. R. Green; D. Margerison. International Statistical Review = Revue Internationale de Statistique, 47(1):87–89, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403217.

Hand:2007:SBRc

[Han07a] David J. Hand. Short book review: Data Mining for Business Intelligence: Concepts, Techniques, and Applications in Microsoft Office Excel with XLMiner by Galit Shmueli, Nitin R. Patel, Peter C. Bruce. International Statistical Review = Revue Internationale de Statistique, 75(2): 256, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508472.

Hand:2007:SBRe

[Han07b] David J. Hand. Short book review: Data Mining the Web: Uncovering Patterns in Web Content, Structure, and Usage by Zdravko Markov, Daniel T. Larose. International Statistical Review = Revue Internationale de Statistique, 75(3):409, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41509880.

Hand:2007:SBRf

[Han07c] David J. Hand. Short book review: Dynamic Data Assimilation: A Least Squares Approach by John M. Lewis, S. Lakshmivarahan, Sudarshan Dhall. International Statistical Review = Revue Internationale de Statistique, 75(3): 410, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509881.

Hand:2007:SBRb

[Han07d] David J. Hand. Short book review: Finite Mixture and Markov Switching Models by Sylvia Frühwirth–Schnatter. International Statistical Review = Revue Internationale de Statistique, 75(2):255, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508471.

Hand:2007:SBRh

[Han07e] David J. Hand. Short book review: Matrix Methods in Data Mining and Pattern Recognition by Lars Eldén. International Statistical Review = Revue Internationale de Statistique, 75(3):418, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509889.

Hand:2007:SBRg

[Han07f] David J. Hand. Short book review: Statistical Development of Quality in Medicine by Per Winkel, Nien Fan Zhang. International Statistical Review = Revue Internationale de Statistique, 75(3):417, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509888.

Hand:2007:SBRa

[Han07g] David J. Hand. Short book review: The Nature of Statistical Evidence by Bill Thompson. International Statistical Review = Revue Internationale de Statistique, 75(2):254– 255, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508470.

Hand:2007:SBRd

[Han07h] David J. Hand. Short book review: The Risks of Financial Institutions edited by Mark Carey, Rene M. Stulz. International Statistical Review = Revue Internationale de Statistique, 75(2):266–267, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508481.

Hand:2008:SBRa

[Han08a] David J. Hand. Short book review: A Statistical Approach to Neural Networks for Pattern Recognition by Robert A. Dunne. International Statistical Review = Revue Internationale de Statistique, 76(1):140, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919589.

Hand:2008:SBRl

[Han08b] David J. Hand. Short book review: Analysis of Multivariate Social Science Data, Second Edition, by David J. Bartholomew, Fiona Steele, Irini Moustaki, Jane Galbraith. International Statistical Review = Revue Internationale de Statistique, 76(3):456, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919676.

Hand:2008:SBRi

[Han08c] David J. Hand. Short book review: Applied Data Analytic Techniques for Turning Points Research edited by Patricia Cohen. International Statistical Review = Revue Internationale de Statistique, 76(3):448–449, December 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919668.

Hand:2008:SBRb

[Han08d] David J. Hand. Short book review: Data Clustering: Theory, Algorithms, and Applications by Guojun Gan, Chaoqun Ma, Jianhong Wu. International Statistical Review = Revue Internationale de Statistique, 76(1):141, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919590.

Hand:2008:SBRf

[Han08e] David J. Hand. Short book review: Evaluating Clinical Research: All That Glitters is not Gold, 2nd Edition, by Bengt D. Furberg, Curt D. Furberg. International Statistical Review = Revue Internationale de Statistique, 76(1): 144–145, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919594.

Hand:2008:SBRg

[Han08f] David J. Hand. Short book review: Financial Surveillance edited by Marianne Frisén. International Statistical Review = Revue Internationale de Statistique, 76(2):304– 305, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919624.

Hand:2008:SBRm

[Han08g] David J. Hand. Short book review: Linking and Aligning Scores and Scales edited by Neil J. Dorans, Mary Pommerich, Paul W. Holland. International Statistical Review = Revue Internationale de Statistique, 76(3):457– 458, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919677.

Hand:2008:SBRo

[Han08h] David J. Hand. Short book review: Meta Analysis: A Guide to Calibrating and Combining Statistical Evidence by Elena Kulinskaya, Stephan Morgenthaler, Robert G. Staudte. International Statistical Review = Revue Internationale de Statistique, 76(3):459, December 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919679.

Hand:2008:SBRh

[Han08i] David J. Hand. Short book review: New Drug Development: Design, Methodology, and Analysis by J. Rick Turner. International Statistical Review = Revue Internationale de Statistique, 76(2):305–306, August 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919625.

Hand:2008:SBRj

[Han08j] David J. Hand. Short book review: Probability and Information: An Integrated Approach, Second Edition, by David Applebaum. International Statistical Review = Revue Internationale de Statistique, 76(3):449–450, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919669.

Hand:2008:SBRe

[Han08k] David J. Hand. Short book review: Random Dynamical Systems: Theory and Applications by Rabi Bhattacharya, Mukul Majumdar. International Statistical Review = Revue Internationale de Statistique, 76(1):143–144, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919593.

Hand:2008:SBRn

[Han08l] David J. Hand. Short book review: Statistical Analysis and Modelling of Spatial Point Patterns by Janine Illian, Antti Penttinen, Helga Stoyan, Dietrich Stoyan. International Statistical Review = Revue Internationale de Statistique, 76(3):458, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919678.

Hand:2008:SBRk

[Han08m] David J. Hand. Short book review: Statistical Decision Theory: Estimation, Testing, and Selection by Friedrich Liese, Klaus-J. Miescke. International Statistical Review = Revue Internationale de Statistique, 76(3):450, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919670.

Hand:2008:SBRd

[Han08n] David J. Hand. Short book review: The Statistical Analysis of Recurrent Events by Richard J. Cook, Jerald F.

Lawless. International Statistical Review = Revue Internationale de Statistique, 76(1):142–143, April 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919592.

Hand:2008:SBRc

[Han08o] David J. Hand. Short book review: Understanding Complex Datasets: Data Mining with Matrix Decompositions by David B. Skillicorn. International Statistical Review = Revue Internationale de Statistique, 76(1):142, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919591.

Hand:2009:SBRa

[Han09a] David J. Hand. Short book review: Encyclopedia of Quantitative Risk Analysis and Assessment edited by Edward L. Melnick, Brian S. Everitt. International Statistical Review = Revue Internationale de Statistique, 77(2):313– 314, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919743.

Hand:2009:SBRc

[Han09b] David J. Hand. Short book review: Forecasting with Exponential Smoothing: The State Space Approach by Rob J. Hyndman, Anne B. Koehler, J. Keith Ord, Ralph D. Snyder. International Statistical Review = Revue Internationale de Statistique, 77(2):315–316, August 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919745.

Hand:2009:SBRd

[Han09c] David J. Hand. Short book review: Observed Confidence Levels: Theory and Application by Alan M. Polansky. International Statistical Review = Revue Internationale de Statistique, 77(2):316, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919746.

Hand:2009:SBRb

[Han09d] David J. Hand. Short book review: Statistical DNA Forensics: Theory, Methods and Computation by Wing Kam Fung, Yue-Qing Hu. International Statistical Review = Revue Internationale de Statistique, 77(2):314– 315, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919744.

Hand:2010:SBRf

[Han10a] David J. Hand. Short book review: Econophysics and Companies: Statistical Life and Death in Complex Business Networks by Hideaki Aoyama, Yoshi Fujiwara, Yuichi Ikeda, Hiroshi Iyetomi, Wataru Souma. International Statistical Review = Revue Internationale de Statistique, 78 (3):445, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919863.

Hand:2010:SBRj

[Han10b] David J. Hand. Short book review: Estimation and tests in distribution mixtures and change-points models by Odile Pons. International Statistical Review = Revue Internationale de Statistique, 78(3):475–476, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919890.

Hand:2010:SBRe

[Han10c] David J. Hand. Short book review: Introduction to Social Statistics: The Logic of Statistical Reasoning by Thomas Dietz, Linda Kalof. International Statistical Review = Revue Internationale de Statistique, 78(2):326–327, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919851.

Hand:2010:SBRd

[Han10d] David J. Hand. Short book review: Machine Learning: An Algorithmic Perspective by Stephen Marsland. International Statistical Review = Revue Internationale de Statistique, 78(2):325, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919850.

Hand:2010:SBRh

[Han10e] David J. Hand. Short book review: Modeling and Analysis of Stochastic Systems, Second Edition, by Vidyadhar G. Kulkarni. International Statistical Review = Revue Internationale de Statistique, 78(3):467, December 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919883.

Hand:2010:SBRb

[Han10f] David J. Hand. Short book review: Statistical Analysis of Network Data: Methods and Models by Eric D. Kolaczyk. International Statistical Review = Revue Internationale de Statistique, 78(1):135, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919800.

Hand:2010:SBRa

[Han10g] David J. Hand. Short book review: Text Mining: Classification, Clustering, and Applications edited by Ashok Srivastava, Mehran Sahami. International Statistical Review = Revue Internationale de Statistique, 78(1):134– 135, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919799.

Hand:2010:SBRi

[Han10h] David J. Hand. Short book review: The Oxford Handbook of Applied Bayesian Analysis edited by Anthony O’Hagan, Mike West. International Statistical Review = Revue Internationale de Statistique, 78(3):474–475, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919889.

Hand:2010:SBRc

[Han10i] David J. Hand. Short book review: The Role of Statistics in Business and Industry by Gerald J. Hahn, Necip Doganaksoy. International Statistical Review = Revue Internationale de Statistique, 78(1):152–153, April 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919818.

Hand:2010:SBRg

[Han10j] David J. Hand. Short book review: Theory of Stochastic Processes: With Applications to Financial Mathematics and Risk Theory by Dmytro Gusak, Alexander Kukush, Alexey Kulik, Yuliya Mishura, Andrey Pilipenko. International Statistical Review = Revue Internationale de Statistique, 78(3):461, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919877.

Hand:2011:SBRe

[Han11a] David J. Hand. Short book review: Bayesian Ideas and Data Analysis: An Introduction for Scientists and Statisticians by Ronald Christensen, Wesley Johnson, Adam Branscum, Timothy E. Hanson. International Statistical Review = Revue Internationale de Statistique, 79(2):285– 286, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305036.

Hand:2011:SBRl

[Han11b] David J. Hand. Short book review: Calculating Catastrophe by Gordon Woo. International Statistical Review = Revue Internationale de Statistique, 79(3):494–495, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41305076.

Hand:2011:SBRb

[Han11c] David J. Hand. Short book review: Charming Proofs: A Journey into Elegant Mathematics by Claudi Alsina, Roger B. Nelsen. International Statistical Review = Revue Internationale de Statistique, 79(1):122, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41306175.

Hand:2011:SBRc

[Han11d] David J. Hand. Short book review: Diagnostic Measurement: Theory, Methods, and Applications by André A. Rupp, Jonathan Templin, Robert A. Henson. International Statistical Review = Revue Internationale de Statistique, 79(1):135–136, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41306187.

Hand:2011:SBRh

[Han11e] David J. Hand. Short book review: Introduction to Psychometric Theory by Tenko Raykov, George A. Marcoulides. International Statistical Review = Revue Internationale de Statistique, 79(2):298–299, August 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305048.

Hand:2011:SBRi

[Han11f] David J. Hand. Short book review: Linear Causal Modeling with Structural Equations by Stanley A. Mulaik. International Statistical Review = Revue Internationale de Statistique, 79(2):299–300, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305049.

Hand:2011:SBRf

[Han11g] David J. Hand. Short book review: Logistic Regression Models by Joseph M. Hilbe. International Statistical Review = Revue Internationale de Statistique, 79(2):287– 288, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305038.

Hand:2011:SBRd

[Han11h] David J. Hand. Short book review: Measurements and their Uncertainties: A Practical Guide to Modern Error Analysis by Ifan G. Hughes, Thomas P. A. Hase. International Statistical Review = Revue Internationale de Statistique, 79(2):280, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305032.

Hand:2011:SBRj

[Han11i] David J. Hand. Short book review: Randomness and Recurrence in Dynamical Systems by Rodney Nillsen. International Statistical Review = Revue Internationale de Statistique, 79(3):484, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41305065.

Hand:2011:SBRg

[Han11j] David J. Hand. Short book review: Simultaneous Inference in Regression by Wei Liu. International Statistical Review = Revue Internationale de Statistique, 79(2):288– 289, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305039.

Hand:2011:SBRa

[Han11k] David J. Hand. Short book review: Statistical Inference: An Integrated Bayesian/Likelihood Approach by Murray Aitkin. International Statistical Review = Revue Internationale de Statistique, 79(1):121, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41306174.

Hand:2011:SBRk

[Han11l] David J. Hand. Short book review: Temporal Data Mining by Theophano Mitsa. International Statistical Review = Revue Internationale de Statistique, 79(3):485, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305066.

Handcock:2011:D

[Han11m] Mark S. Handcock. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):223– 225, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:APC

[Han12a] David J. Hand. Assessing the performance of classification methods. International Statistical Review = Revue Internationale de Statistique, 80(3):400–414, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRa

[Han12b] David J. Hand. Short book review: A Historian Looks Back: The Calculus as Algebra and Selected Writings by Judith V. Grabiner. International Statistical Review = Revue Internationale de Statistique, 80(1):180–181, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257175.

Hand:2012:SBRv

[Han12c] David J. Hand. Short book review: An Introduction to Statistical Concepts, Third Edition, by Richard G. Lomax, Debbie L. Hahs-Vaughn. International Statistical Review = Revue Internationale de Statistique, 80(3):489–491, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRg

[Han12d] David J. Hand. Short book review: Biosurveillance: Methods and Case Studies edited by Taha Kass-Hout, Xiaohui Zhang. International Statistical Review = Revue Internationale de Statistique, 80(1):193–194, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257188.

Hand:2012:SBRi

[Han12e] David J. Hand. Short book review: Controversial Statistical Issues in Clinical Trials by Shein-Chung Chow. International Statistical Review = Revue Internationale de Statistique, 80(1):195–196, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/23257190.

Hand:2012:SBRf

[Han12f] David J. Hand. Short book review: Empirical Model Building: Data, Models, and Reality, Second Edition, by James R. Thompson. International Statistical Review = Revue Internationale de Statistique, 80(1):192, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257187.

Hand:2012:SBRu

[Han12g] David J. Hand. Short book review: Handbook of Partial Least Squares: Concepts, Methods and Applications by Vincenzo Esposito Vinzi, Wynne W. Chin, Jörg Henseler, Huiwen Wang. International Statistical Review = Revue Internationale de Statistique, 80(3):488–489, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRb

[Han12h] David J. Hand. Short book review: Knowledge Discovery from Data Streams by João Gama. International Statistical Review = Revue Internationale de Statistique, 80(1): 181–182, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/23257176.

Hand:2012:SBRe

[Han12i] David J. Hand. Short book review: Modeling and Reasoning with Bayesian Networks by Adnan Darwiche. International Statistical Review = Revue Internationale de Statistique, 80(1):191, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/23257186.

Hand:2012:SBRj

[Han12j] David J. Hand. Short book review: Modern Issues and Methods in Biostatistics by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 80(1):196–197, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257191.

Hand:2012:SBRl

[Han12k] David J. Hand. Short book review: Principles of Applied Statistics by D. R. Cox and Christl A. Donnelly. International Statistical Review = Revue Internationale de Statistique, 80(2):336–338, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRw

[Han12l] David J. Hand. Short book review: Statistical Concepts: A Second Course, Fourth Edition, by Richard G. Lomax, Debbie L. Hahs-Vaughn. International Statistical Review = Revue Internationale de Statistique, 80(3):491, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRq

[Han12m] David J. Hand. Short book review: Statistical Confidentiality: Principles and Practice by George T. Duncan, Mark Elliot, Juan–José Salazar–González. International Statistical Review = Revue Internationale de Statistique, 80(3):479–480, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRk

[Han12n] David J. Hand. Short book review: Statistical Methods for Trend Detection and Analysis in the Environmental Sciences by Richard Chandler and Marian Scott. International Statistical Review = Revue Internationale de Statistique, 80(2):324–325, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRh

[Han12o] David J. Hand. Short book review: Stochastic Processes by Richard F. Bass. International Statistical Review = Revue Internationale de Statistique, 80(1):194–195, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257189.

Hand:2012:SBRt

[Han12p] David J. Hand. Short book review: Synthetic Datasets for Statistical Disclosure Control: Theory and Implementation by Jörg Drechsler. International Statistical Review = Revue Internationale de Statistique, 80(3):483, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRd

[Han12q] David J. Hand. Short book review: Testing Statistical Hypotheses of Equivalence and Noninferiority, Second Edition, by Stefan Wellek. International Statistical Review = Revue Internationale de Statistique, 80(1):188–189, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257183.

Hand:2012:SBRm

[Han12r] David J. Hand. Short book review: The Foundations of Modern Time Series Analysis by Terence C. Mills. International Statistical Review = Revue Internationale de Statistique, 80(2):340–341, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRo

[Han12s] David J. Hand. Short book review: The Origin of Financial Crises: Central Banks, Credit Bubbles and the Efficient Market Fallacy by George Cooper. International Statistical Review = Revue Internationale de Statistique, 80(3):478–479, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRp

[Han12t] David J. Hand. Short book review: The Star and the Whole: Gian-Carlo Rota on Mathematics and Phenomenology by Fabrizio Palombi. International Statistical Review = Revue Internationale de Statistique, 80(3): 479, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRc

[Han12u] David J. Hand. Short book review: Theoretical Statistics: Topics for a Core Course by Robert W. Keener. International Statistical Review = Revue Internationale de Statistique, 80(1):187–188, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/23257182.

Hand:2012:SBRr

[Han12v] David J. Hand. Short book review: Time Series Analysis and Forecasting by Example by Søren Bisgaard, Murat Kulahci. International Statistical Review = Revue Internationale de Statistique, 80(3):480–481, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRn

[Han12w] David J. Hand. Short book review: Understanding The New Statistics: Effect Sizes, Confidence Intervals, and Meta-Analysis by Geoff Cumming. International Statistical Review = Revue Internationale de Statistique, 80(2): 344–345, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2012:SBRs

[Han12x] David J. Hand. Short book review: Visualizing Time: Designing Graphical Representations for Statistical Data by Graham Wills. International Statistical Review = Revue Internationale de Statistique, 80(3):482–483, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRj

[Han13a] David J. Hand. Short book review: A Practitioner’s Guide to Resampling for Data Analysis, Data Mining, and Modeling by Philip I. Good. International Statistical Review = Revue Internationale de Statistique, 81(2):326, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRh

[Han13b] David J. Hand. Short book review: A Statistical Guide for the Ethically Perplexed by Lawrence Hubert, Howard Wainer. International Statistical Review = Revue Internationale de Statistique, 81(2):314–316, August 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRe

[Han13c] David J. Hand. Short book review: An Introduction to Model-Based Survey Sampling with Applications by Raymond L. Chambers, Robert G. Clark. International Statistical Review = Revue Internationale de Statistique, 81 (1):163, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRk

[Han13d] David J. Hand. Short book review: Comparing Groups: Randomization and Bootstrap Methods Using R by Andrew S. Zieffler, Jeffrey R. Harring, Jeffrey D. Long. International Statistical Review = Revue Internationale de Statistique, 81(2):326–328, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRi

[Han13e] David J. Hand. Short book review: Graphical Models with R by Søren Højsgaard, David Edwards, Steffen Lauritzen. International Statistical Review = Revue Internationale de Statistique, 81(2):316, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRn

[Han13f] David J. Hand. Short book review: Introduction to Probability with Texas Hold’em Examples by Frederic Paik Schoenberg. International Statistical Review = Revue Internationale de Statistique, 81(2):334, August 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRf

[Han13g] David J. Hand. Short book review: Inverse Problems and High-Dimensional Estimation: Stats in the Château Summer School, August 31–September 4, 2009 edited by Pierre Alquier, Eric Gautier, Gilles Stoltz. International Statistical Review = Revue Internationale de Statistique, 81 (1):166, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRm

[Han13h] David J. Hand. Short book review: Latent Variable Models and Factor Analysis: A Unified Approach, Third Edition, by David J. Bartholomew, Martin Knott, Irini Moustaki. International Statistical Review = Revue Internationale de Statistique, 81(2):333–334, August 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRl

[Han13i] David J. Hand. Short book review: Living Standards Analytics: Development through the Lens of Household Survey Data by Dominique Haughton, Jonathan Haughton. International Statistical Review = Revue Internationale de Statistique, 81(2):331–332, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRb

[Han13j] David J. Hand. Short book review: Measures of Interobserver Agreement and Reliability, Second Edition by Mohamed M. Shoukri. International Statistical Review = Revue Internationale de Statistique, 81(1):155–156, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRa

[Han13k] David J. Hand. Short book review: Probability for Statistics and Machine Learning: Fundamentals and Advanced Topics by Anirban DasGupta. International Statistical Review = Revue Internationale de Statistique, 81(1):155, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRg

[Han13l] David J. Hand. Short book review: Qualitative Inquiry in Clinical and Educational Settings by Danica G. Hays, Anneliese A. Singh. International Statistical Review = Revue Internationale de Statistique, 81(1):167–168, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRo

[Han13m] David J. Hand. Short book review: Statistical Evaluation of Diagnostic Performance: Topics in ROC Analysis by Kelly H. Zou, Aiyi Liu, Andriy I. Bandos, Lucila Ohno-Machado, Howard E. Rockette. International Statistical Review = Revue Internationale de Statistique, 81(2):335, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRd

[Han13n] David J. Hand. Short book review: Statistical Inference: The Minimum Distance Approach by Ayanendranath Basu, Hiroyuki Shioya, Chanseok Park. International Statistical Review = Revue Internationale de Statistique, 81(1):162, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRc

[Han13o] David J. Hand. Short book review: Statistical Tools for Measuring Agreement by Lawrence Lin, A. S. Hedayat, Wenting Wu. International Statistical Review = Revue Internationale de Statistique, 81(1):156–157, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hartigan:1973:UNV

[Har73] J. A. Hartigan. United Nations votes: A clustering example. International Statistical Review = Revue Internationale de Statistique, 41(2):259–263, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402844.

Harter:1974:MLSa

[Har74a] H. Leon Harter. The method of least squares and some alternatives: Part I. International Statistical Review = Revue Internationale de Statistique, 42(2):147–174, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403077.

Harter:1974:MLSb

[Har74b] H. Leon Harter. The method of least squares and some alternatives: Part II. International Statistical Review = Revue Internationale de Statistique, 42(3):235–264+282, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402983.

Harter:1975:MLSd

[Har75a] H. Leon Harter. The method of least squares and some alternatives. Addendum to Part IV. International Statistical Review = Revue Internationale de Statistique, 43(3):273– 278, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403111.

Harter:1975:MLSa

[Har75b] H. Leon Harter. The method of least squares and some alternatives: Part III. International Statistical Review = Revue Internationale de Statistique, 43(1):1–44, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402658.

Harter:1975:MLSb

[Har75c] H. Leon Harter. The method of least squares and some alternatives: Part IV. International Statistical Review = Revue Internationale de Statistique, 43(2):125–190, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402897.

Harter:1975:MLSc

[Har75d] H. Leon Harter. The method of least squares and some alternatives: Part V. International Statistical Review = Revue Internationale de Statistique, 43(3):269–272, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403110.

Harter:1976:MLS

[Har76] H. Leon Harter. The method of least squares and some alternatives: Part VI. Subject and author indexes. International Statistical Review = Revue Internationale de Statistique, 44(1):113–159, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402966.

Harter:1978:BEV

[Har78] H. Leon Harter. A bibliography of extreme-value theory. International Statistical Review = Revue Internationale de Statistique, 46(3):279–306, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402375.

Harter:1979:SEA

[Har79] H. Leon Harter. Some early applications of order statistics to multivariate analysis. International Statistical Review = Revue Internationale de Statistique, 47(3):267– 282, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402648.

Harrington:1982:DPP

[Har82] David P. Harrington. Discussion of paper by P. K. Andersen, Ø. Borgan, R. Gill and N. Keiding. International Statistical Review = Revue Internationale de Statistique, 50(3):250–254, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402493.

Hartigan:1990:PPE

[Har90] J. A. Hartigan. Perturbed periodogram estimates of variance. International Statistical Review = Revue Internationale de Statistique, 58(1):1–7, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403469.

Harvey:1996:IAC

[Har96] Andrew Harvey. Intervention analysis with control groups. International Statistical Review = Revue Internationale de Statistique, 64(3):313–328, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403788.

Harper:1999:ECI

[Har99] Michael J. Harper. Estimating capital inputs for productivity measurement: An overview of U.S. concepts and methods. International Statistical Review = Revue Internationale de Statistique, 67(3):327–337, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403709.

Harkanen:2008:SBR

[Här08] Tommi Härkänen. Short book review: The Frailty Model by Luc Duchateau, Paul Janssen. International Statistical Review = Revue Internationale de Statistique, 76(3):438– 439, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919658.

Hashim:1972:CCN

[Has72] J. M. Hashim. Capital consumption: National vs. business accounting points of view. International Statistical Review = Revue Internationale de Statistique, 40(1):53–65, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402706.

Haslett:2008:SBRc

[Has08a] Stephen Haslett. Short book review: Data Quality and Record Linkage Techniques by Thomas N. Herzog, Fritz J. Scheuren, William E. Winkler. International Statistical Review = Revue Internationale de Statistique, 76(2):325– 326, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919644.

Haslett:2008:SBRa

[Has08b] Stephen Haslett. Short book review: Indirect Sampling by Pierre Lavallée. International Statistical Review = Revue Internationale de Statistique, 76(2):320–321, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919640.

Haslett:2008:SBRd

[Has08c] Stephen Haslett. Short book review: Sampling Techniques for Forest Inventories by Daniel Mandallaz. International Statistical Review = Revue Internationale de Statistique, 76(2):326–327, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919645.

Haslett:2008:SBRb

[Has08d] Stephen Haslett. Short book review: Studying Human Populations: An Advanced Course in Statistics by Nicholas T. Longford. International Statistical Review = Revue Internationale de Statistique, 76(2):322–323, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919641.

Hauser:1972:FFS

[Hau72] Philip M. Hauser. Frederick Franklin Stephan, 1903– 1971. International Statistical Review = Revue Internationale de Statistique, 40(2):261–262, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402780.

Haug:2001:ERL

[Hau01] Werner Haug. Ethnic, religious and language groups: Towards a set of rules for data collection and statistical analysis. International Statistical Review = Revue Internationale de Statistique, 69(2):303–311, August 2001. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403817.

Hauke:2008:SBRb

[Hau08a] Jan Hauke. Short book review: Probably Not: Future Prediction Using Probability and Statistical Inference by Lawrence N. Dworsky. International Statistical Review = Revue Internationale de Statistique, 76(3):455, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919675.

Hauke:2008:SBRa

[Hau08b] Jan Hauke. Short book review: Quantitative Economics: How sustainable are our economies? by Peter Bartelmus. International Statistical Review = Revue Internationale de Statistique, 76(3):454, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919674.

Haupt:2018:BRN

[Hau18] Sue Ellen Haupt. Book review: Nonlinear Modeling of Solar Radiation and Wind Speed Time Series, Luigi Fortuna, Giuseppe Nunnari and Silvia Nunnari, Springer, 2016, xv + 98 pages, $54.99, softcover ISBN: 978-3-319-38764-2. International Statistical Review = Revue Internationale de Statistique, 86(1):161, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Havil:2014:JNL

[Hav14] Julian Havil. John Napier: Life, Logarithms, and Legacy. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15570-4, 0-691-14342-0. xv + 279 pp. LCCN QA247.5 .H38 2012.

Hawkins:1997:DFB

[Haw97a] Anne Hawkins. Discussion: Forward to basics! A personal view of developments in statistical education. International Statistical Review = Revue Internationale de Statistique, 65(3):280–287, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403370.

Hawkins:1997:NPN

[Haw97b] Anne Hawkins. [new pedagogy and new content: The case of statistics]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2):141– 146, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403335.

Hanif:1980:SUP

[HB80] Muhammad Hanif and K. R. W. Brewer. Sampling with unequal probabilities without replacement: a review. International Statistical Review = Revue Internationale de Statistique, 48(3):317–335, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402944.

Haziza:2007:CIC

[HB07] David Haziza and Jean-François Beaumont. On the construction of imputation classes in surveys. International Statistical Review = Revue Internationale de Statistique, 75(1):25–43, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41508447.

Ho:2005:SSS

[HBK05] Frederick W. H. Ho, L. Biggeri, and Graham Kalton. Survey as a source of statistics and factors affecting the quality of survey statistics. International Statistical Review = Revue Internationale de Statistique, 73(2):245– 248, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472664.

Hasler:2018:VCI

[HCR18] Caren Hasler, Radu V. Craiu, and Louis-Paul Rivest. Vine copulas for imputation of monotone non-response. International Statistical Review = Revue Internationale de Statistique, 86(3):488–511, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hu:2021:SML

[HCV^{+}21] Linwei Hu, Jie Chen, Joel Vaughan, Soroush Aramideh, Hanyu Yang, Kelly Wang, Agus Sudjianto, and Vijayan N. Nair. Supervised machine learning techniques: an overview with applications to banking. International Statistical Review = Revue Internationale de Statistique, 89(3):573–604, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

He:2014:BRS

[He14] Xin He. Book review: Statistical Analysis of Panel Count Data by Jianguo Sun, Xingqiu Zhao. International Statistical Review = Revue Internationale de Statistique, 82 (2):324, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

He:2017:ICR

[He17] Xuming He. Interview: a conversation with Roger Koenker. International Statistical Review = Revue Internationale de Statistique, 85(1):46–60, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Healy:1995:FYW

[Hea95] M. J. R. Healy. Frank Yates, 1902–1994: The work of a statistician. International Statistical Review = Revue Internationale de Statistique, 63(3):271–288, December 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403480.

Heckman:2008:EC

[Hec08] James J. Heckman. Econometric causality. International Statistical Review = Revue Internationale de Statistique, 76(1):1–27, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919581.

Hedlin:2020:TSA

[Hed20] Dan Hedlin. Is there a ’safe area’ where the nonresponse rate has only a modest effect on bias despite non-ignorable nonresponse? International Statistical Review = Revue Internationale de Statistique, 88(3):642–657, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Heeringa:2016:BRH

[Hee16] Steven G. Heeringa. Book review: Handbook of Health Survey Methods, Timothy P. Johnson, Editor, John Wiley and Sons, Inc., 2015, xxii + 816 pages, £100.00/ EUR125.00, hardcover ISBN: 978-1-118-00232-2. International Statistical Review = Revue Internationale de Statistique, 84(3):552–553, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Helland:2004:SIU

[Hel04] Inge S. Helland. Statistical inference under symmetry. International Statistical Review = Revue Internationale de Statistique, 72(3):409–422, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472631.

Henderson:1978:MCA

[Hen78] K. V. Henderson. The Munich Centre for Advanced Training in Applied Statistics for Developing Countries. International Statistical Review = Revue Internationale de Statistique, 46(2):225–226, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402817.

Heyde:1981:IPS

[Hey81] C. C. Heyde. Invariance principles in statistics, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 49(2):143–152, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403068.

Hammick:1994:GTS

[HG94] Patricia A. Hammick and Joseph L. Gastwirth. Group testing for sensitive characteristics: Extension to higher prevalence levels. International Statistical Review = Revue Internationale de Statistique, 62(3):319–331, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403764.

Hemerik:2021:ALL

[HG21] Jesse Hemerik and Jelle J. Goeman. Another look at the Lady Tasting Tea and differences between permutation tests and randomisation tests. International Statistical Review = Revue Internationale de Statistique, 89(2):367– 381, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Horvitz:1976:RRD

[HGA76] D. G. Horvitz, B. G. Greenberg, and J. R. Abernathy. Randomized response: a data-gathering device for sensitive questions. International Statistical Review = Revue Internationale de Statistique, 44(2):181–196, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1403276. International Association of Survey Statisticians Meeting on Randomized Response (Warsaw, 1975).

Hobolth:2020:UFC

[HGKJ20] Asger Hobolth, Qianyun Guo, Astrid Kousholt, and Jens Ledet Jensen. A unifying framework and comparison of algorithms for non-negative matrix factorisation. International Statistical Review = Revue Internationale de Statistique, 88(1):29–53, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hamdan:1993:ETP

[HH93] M. A. Hamdan and Hisham Hilow. Estimation and testing for the parameters of the bivariate Poisson population based on a 2 × 2 sample. International Statistical Review = Revue Internationale de Statistique, 61(3):361– 367, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403748.

Hogg:1995:CQI

[HH95] Robert V. Hogg and Mary C. Hogg. Continuous quality improvement in higher education. International Statistical Review = Revue Internationale de Statistique, 63(1): 35–48, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403776.

Haslett:2007:TBT

[HH07] John Haslett and Stephen J. Haslett. The three basic types of residuals for a linear model. International Statistical Review = Revue Internationale de Statistique, 75 (1):1–24, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508446.

Huggins:2011:RUC

[HH11] Richard Huggins and Wen-Han Hwang. A review of the use of conditional likelihood in capture-recapture experiments. International Statistical Review = Revue Internationale de Statistique, 79(3):385–400, December 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305057.

Henn:2014:MLM

[HH14] Lisa Henn and James S. Hodges. Multiple local maxima in restricted likelihoods and posterior distributions for mixed linear models. International Statistical Review = Revue Internationale de Statistique, 82(1):90–105, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hoerl:1997:DLS

[HHD97] Roger Hoerl, Gerry Hahn, and Necip Doganaksoy. Discussion: Let’s stop squandering our most strategic weapon. International Statistical Review = Revue Internationale de Statistique, 65(2):147–153, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403336.

Hahn:2012:DDS

[HHDG12] Gerald J. Hahn, Roger Hoerl, Necip Doganaksoy, and Martha Gardner. Discussion: Discussion: Statistics research in business and industry — the General Electric experience. International Statistical Review = Revue Internationale de Statistique, 80(2):219–230, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hardle:2003:BMT

[HHK03] Wolfgang Härdle, Joel Horowitz, and Jens-Peter Kreiss. Bootstrap methods for time series. International Statistical Review = Revue Internationale de Statistique, 71(2): 435–459, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403897.

Hitaj:2018:PMD

[HHMR18] Asmerilda Hitaj, Friedrich Hubalek, Lorenzo Mercuri, and Edit Rroji. On properties of the MixedTS distribution and its multivariate extension. International Statistical Review = Revue Internationale de Statistique, 86(3):512– 540, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Higgins:1977:BIO

[Hig77] J. J. Higgins. Bayesian inference and the optimality of maximum likelihood estimation. International Statistical Review = Revue Internationale de Statistique, 45(1): 9–11, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor. org/stable/1402999. See comment [Sch78].

Hilbe:2015:BRB

[Hil15] Joseph M. Hilbe. Book review: Bayesian Analysis with Stata. International Statistical Review = Revue Internationale de Statistique, 83(2):331–332, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hirotsu:1993:BAV

[Hir93] Chihiro Hirotsu. Beyond analysis of variance techniques: Some applications in clinical trials. International Statistical Review = Revue Internationale de Statistique, 61(1): 183–201, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403602.

Hirotsu:2005:DCH

[Hir05] C. Hirotsu. Discussant: C. Hirotsu. International Statistical Review = Revue Internationale de Statistique, 73(2): 265–268, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hald:1983:MRF

[HJ83] A. Hald and S. Johansen. On de Moivre’s recursion formulae for the duration of play. International Statistical Review = Revue Internationale de Statistique, 51(3):239– 253, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402586.

Hjort:1992:IPS

[Hjo92] Nils Lid Hjort. On inference in parametric survival data models. International Statistical Review = Revue Internationale de Statistique, 60(3):355–387, December 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403683.

Higgins:1977:VSG

[HK77] James E. Higgins and Gary G. Koch. Variable selection and generalized chi-square analysis of categorical data applied to a large cross-sectional occupational health survey. International Statistical Review = Revue Internationale de Statistique, 45(1):38–38 + 51–62, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403003.

Hobert:2016:DPI

[HK16] James P. Hobert and Kshitij Khare. Discussions on posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Discussion. International Statistical Review = Revue Internationale de Statistique, 84 (3):349–356, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Hidiroglou:2018:SSP

[HK18] Michael A. Hidiroglou and Marcin Kozak. Stratification of skewed populations: A comparison of optimisation-based versus approximate methods. International Statistical Review = Revue Internationale de Statistique, 86(1):87–105, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hardle:2007:ULS

[HKZ07] Wolfgang Härdle, Sigbert Klinke, and Uwe Ziegenhagen. On the utility of e-learning in statistics. International Statistical Review = Revue Internationale de Statistique, 75(3):355–364, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41509876.

Huber:1974:ETC

[HL74] C. Huber and J. Lellouch. Estimation dans les tableaux de contingence à un grand nombre d’entrées. (French) [Estimation in contingency tables with a large number of entries]. International Statistical Review = Revue Internationale de Statistique, 42(2):193–203, August 1974. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403079.

Hall:1990:MAE

[HL90] Peter Hall and Barbara La Scala. Methodology and algorithms of empirical likelihood. International Statistical Review = Revue Internationale de Statistique, 58(2):109– 127, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403462.

Hidiroglou:2001:SEI

[HL01] Michael A. Hidiroglou and Normand Laniel. Sampling and estimation issues for annual and sub-annual Canadian business surveys. International Statistical Review = Revue Internationale de Statistique, 69(3):487–504, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403458.

Han:2019:SAL

[HL19] Ying Han and Partha Lahiri. Statistical analysis with linked data. International Statistical Review = Revue Internationale de Statistique, 87(S1):S139–S157, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hardle:1997:RNT

[HLC97] Wolfgang Härdle, Helmut Lütkepohl, and Rong Chen. A review of nonparametric time series analysis. International Statistical Review = Revue Internationale de Statistique, 65(1):49–72, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403432.

Hall:1990:LSE

[HM90] W. J. Hall and David J. Mathiason. On large-sample estimation and testing in parametric models. International Statistical Review = Revue Internationale de Statistique, 58(1):77–97, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403475.

Harris:2004:CAD

[HM04] Mark N. Harris and László Mátyás. A comparative analysis of different IV and GMM estimators of dynamic panel data models. International Statistical Review = Revue Internationale de Statistique, 72(3):397–408, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472630.

Hart:2019:PFB

[HM19] Jeffrey D. Hart and Matthew Malloure. Prior-free Bayes factors based on data splitting. International Statistical Review = Revue Internationale de Statistique, 87(2):419– 442, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hermans:2018:TPU

[HMA^{+}18] Lisa Hermans, Geert Molenberghs, Marc Aerts, Michael G. Kenward, and Geert Verbeke. A tutorial on the practical use and implication of complete sufficient statistics.

International Statistical Review = Revue Internationale de Statistique, 86(3):403–414, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hsu:2022:THZ

[HMT22] Wei-Wen Hsu, Nadeesha R. Mawella, and David Todem. On testing for homogeneity with zero-inflated models through the lens of model misspecification. International Statistical Review = Revue Internationale de Statistique, 90(1):62–77, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hui:2021:REM

[HMW21] Francis K. C. Hui, Samuel Müller, and Alan H. Welsh. Random effects misspecification can have severe consequences for random effects inference in linear mixed models. International Statistical Review = Revue Internationale de Statistique, 89(1):186–206, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hallin:2012:ENE

[HN12] Marc Hallin and Vijay Nair. Editors’ note [Energy statistics]. International Statistical Review = Revue Internationale de Statistique, 80(1):1, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/23257163.

Hallin:2013:ENE

[HN13a] Marc Hallin and Vijay Nair. Editors’ note. International Statistical Review = Revue Internationale de Statistique, 81(1):1, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hallin:2013:ENN

[HN13b] Marc Hallin and Vijay Nair. Editors’ note: New book review Editor for the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 81(3):337, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ho:2005:RFW

[Ho05] Frederick W. H. Ho. Rejoinder: Frederick W. H. Ho. International Statistical Review = Revue Internationale de Statistique, 73(2):253–254, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hoem:1974:BRB

[Hoe74] J. M. Hoem. Book review: The Growth and Structure of Human Populations; A Mathematical Investigation by A. J. Coale. International Statistical Review = Revue Internationale de Statistique, 42(1):107–108, April 1974. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402689.

Hoeffding:1976:HH

[Hoe76] Wassily Hoeffding. Harold Hotelling, 1895–1973. International Statistical Review = Revue Internationale de Statistique, 44(3):385–386, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402366.

Hoem:1983:BRB

[Hoe83] Jan M. Hoem. [book review:] The Reticent Trio: Some Little-Known Early Discoveries in Life Insurance Mathematics by L. H. F. Oppermann, T. N. Thiele and J. P. Gram. International Statistical Review = Revue Internationale de Statistique, 51(2):213–221, August 1983. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402750.

Hoem:1987:SAM

[Hoe87] Jan M. Hoem. Statistical analysis of a multiplicative model and its application to the standardization of vital rates: A review. International Statistical Review = Revue Internationale de Statistique, 55(2):119–152, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403190.

Hoerl:1999:DWN

[Hoe99] Roger W. Hoerl. Discussion: We need to change our paradigm. International Statistical Review = Revue Internationale de Statistique, 67(2):132–138, August 1999. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403392.

Hoffmann:1997:ARS

[Hof97] Eivind Hoffmann. Administrative records and surveys as basis for statistics on international labour migration. International Statistical Review = Revue Internationale de Statistique, 65(2):221–246, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403344.

Hoijtink:2013:OBF

[Hoi13] Herbert Hoijtink. Objective Bayes factors for inequality constrained hypotheses. International Statistical Review = Revue Internationale de Statistique, 81(2):207–229, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Holgate:1984:IHW

[Hol84] P. Holgate. The influence of Huygens’ work in dynamics on his contribution to probability. International Statistical Review = Revue Internationale de Statistique, 52(2):137– 140, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403096.

Holst:1986:BCO

[Hol86] Lars Holst. On birthday, collectors’, occupancy and other classical urn problems. International Statistical Review = Revue Internationale de Statistique, 54(1):15–27, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403255.

Holmstrom:2008:SBR

[Hol08] Lasse Holmström. Short book review: Nonlinear Dimensionality Reduction by John A. Lee, Michel Verleysen. International Statistical Review = Revue Internationale de Statistique, 76(2):308–309, August 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919628.

Honda:2015:DVC

[Hon15] Toshio Honda. Discussions on varying coefficient regression models: A review and new developments: Discussion.

International Statistical Review = Revue Internationale de Statistique, 83(1):68–70, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hoover:2005:EPS

[Hoo05] Donald R. Hoover. Extending power and sample size approaches developed for McNemar’s procedure to general sign tests. International Statistical Review = Revue Internationale de Statistique, 73(1):103–110, April 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472642.

Hoppe:1993:BIE

[Hop93] Fred M. Hoppe. Beyond inclusion-and-exclusion: Natural identities for P [Exactly t Events] and P [at Least t Events] and resulting inequalities. International Statistical Review = Revue Internationale de Statistique, 61(3):435– 446, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403754.

Horvath:1972:CSI

[Hor72] R. A. Horváth. Le concept de statistique internationale et son evolution historique. (French) [The concept of international statistics and its historical evolution]. International Statistical Review = Revue Internationale de Statistique, 40(3):281–298, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402469.

Horstmann:1973:STR

[Hor73] K. Horstmann. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 41(3):389– 390, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402628.

Horvath:1979:BRB

[Hor79] Robert A. Horváth. Book review: Pour une Histoire de la Statistique, Tome I, Contributions by J. Mairesse. International Statistical Review = Revue Internationale de Statistique, 47(2):232–233, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402576.

Horvath:1991:SMI

[Hor91] Robert A. Horváth. Süssmilch’s methodological impact on European statistics. International Statistical Review = Revue Internationale de Statistique, 59(1):59–66, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403574.

Horgan:2006:SSP

[Hor06] Jane M. Horgan. Stratification of skewed populations: A review. International Statistical Review = Revue Internationale de Statistique, 74(1):67–76, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472691.

Horton:2016:DES

[Hor16] Nicholas Jon Horton. Discussions on enhancing statistics education by including qualitative research: Discussion: Making progress in a crowded market. International Statistical Review = Revue Internationale de Statistique, 84 (2):179–181, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hothorn:2016:STU

[Hot16] L. (Ludwig) Hothorn. Statistics in toxicology using R. The R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4987-0127-2 (hardcover). xviii + 234 pp. LCCN RA1199.4.S73 H68 2016.

Hougaard:1988:ADN

[Hou88] Philip Hougaard. The asymptotic distribution of nonlinear regression parameter estimates: Improving the approximation. International Statistical Review = Revue Internationale de Statistique, 56(3):221–233, December 1988. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403350.

Hoyland:1972:BRB

[Høy72] Arnljot Høyland. Book review: Nonparametric Techniques in Statistical Inference by M. L. Puri. International Statistical Review = Revue Internationale de Statistique, 40(1): 106–107, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402720.

Hoyle:1973:TIB

[Hoy73] M. H. Hoyle. Transformations: An introduction and a bibliography. International Statistical Review = Revue Internationale de Statistique, 41(2):203–223, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402836.

Hoyle:1976:CTI

[Hoy76] M. H. Hoyle. Comment: “Transformations — an introduction and bibliography” (Internat. Statist. Rev. 41 (1973), no. 2, 203–223). International Statistical Review = Revue Internationale de Statistique, 44(3):368, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402354.

Holmstrom:2017:RSS

[HP17a] Lasse Holmström and Leena Pasanen. Rejoinder to statistical scale space methods: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 85 (1):43–45, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Holmstrom:2017:SSS

[HP17b] Lasse Holmström and Leena Pasanen. Statistical scale space methods. International Statistical Review = Revue Internationale de Statistique, 85(1):1–30, April 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hassler:2022:UNM

[HP22] Uwe Hassler and Marc-Oliver Pohle. Unlucky number 13? Manipulating evidence subject to snooping. International Statistical Review = Revue Internationale de Statistique, 90(2):397–410, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Haslett:2022:MEZ

[HPHdAM22] John Haslett, Andrew C. Parnell, John Hinde, and Rafael de Andrade Moral. Modelling excess zeros in count data: a new perspective on modelling approaches. International Statistical Review = Revue Internationale de Statistique, 90(2):216–236, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hendry:1983:EAEb

[HR83a] D. F. Hendry and J.-F. Richard. [the econometric analysis of economic time series]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 51(2):160–163, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402745.

Hendry:1983:EAEa

[HR83b] David F. Hendry and Jean-François Richard. The econometric analysis of economic time series. International Statistical Review = Revue Internationale de Statistique, 51 (2):111–148, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402738.

Holan:2020:F

[HR20] Scott H. Holan and Nalini Ravishanker. Foreword. International Statistical Review = Revue Internationale de Statistique, 88(S1):S1, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Holan:2021:FMI

[HR21] Scott H. Holan and Nalini Ravishanker. Foreword for the mini issue in honour of Prof C. R. Rao. International Statistical Review = Revue Internationale de Statistique, 89(2):215, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Herrera:2007:GDP

[HRR07] Javier Herrera, Mireille Razafindrakoto, and François Roubaud. Governance, democracy and poverty reduction: Lessons drawn from household surveys in Sub-Saharan Africa and Latin America. International Statistical Review = Revue Internationale de Statistique, 75(1):70–95, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508450.

Holan:2020:FCM

[HRR20] Scott H. Holan, Nalini Ravishanker, and J. Sunil Rao. Foreword: COVID-19 mini-issue-statistical primers. International Statistical Review = Revue Internationale de Statistique, 88(2):396–397, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Heldal:1988:CSR

[HS88] Johan Heldal and Emil Spjøtvoll. Combination of surveys and registers: A calibration approach with categorical variables. International Statistical Review = Revue Internationale de Statistique, 56(2):153–164, August 1988. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403639.

Holt:1989:CCR

[HS89] D. Holt and C. J. Skinner. Components of change in repeated surveys. International Statistical Review = Revue Internationale de Statistique, 57(1):1–18, April 1989. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403581.

Hoppe:1990:BTI

[HS90] Fred M. Hoppe and Eugene Seneta. A Bonferroni-type identity and permutation bounds. International Statistical Review = Revue Internationale de Statistique, 58(3):253– 261, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403808.

Hoppe:2012:GIB

[HS12] Fred M. Hoppe and Eugene Seneta. Gumbel’s identity, binomial moments, and Bonferroni sums. International Statistical Review = Revue Internationale de Statistique, 80(2):269–292, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Huang:2015:DVC

[HS15] Jianhua Z. Huang and Ya Su. Discussions on varying coefficient regression models: A review and new developments: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(1):70–72, April 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Haziza:2019:ICS

[HS19] David Haziza and Paul A. Smith. An interview with Chris Skinner. International Statistical Review = Revue Internationale de Statistique, 87(3):451–470, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hansen:1992:TMP

[HT92] Katherine M. Hansen and John W. Tukey. Tuning a major part of a clustering algorithm. International Statistical Review = Revue Internationale de Statistique, 60(1): 21–43, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403499.

Holst:2001:CTU

[HTW01] Erik Holst, Poul Thyregod, and Peter-Th. Wilrich. On conformity testing and the use of two stage procedures. International Statistical Review = Revue Internationale de Statistique, 69(3):419–432, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403454.

Huang:1989:MPO

[Hua89] J. S. Huang. Moment problem of order statistics: A review. International Statistical Review = Revue Internationale de Statistique, 57(1):59–66, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403584.

Huber:2014:DCI

[Hub14] Peter J. Huber. Discussions on “A Cognitive Interpretation of Data Analysis”: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(2):

205–207, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [GW14a].

Hugo:2006:ISI

[Hug06] Graeme Hugo. Improving statistics on international migration in Asia. International Statistical Review = Revue Internationale de Statistique, 74(3):335–355, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472713.

Huh:2005:DMH

[Huh05] Moon Y. Huh. Discussant: Moon Y. Huh. International Statistical Review = Revue Internationale de Statistique, 73(2):227–229, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hummel:1977:CEC

[Hum77] Thomas J. Hummel. Comment on ’An Empirical Comparison of the Methods of Least Squares and Unweighted Means… ’. International Statistical Review = Revue Internationale de Statistique, 45(3):299–301, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402542.

Hunter:2014:DBS

[Hun14] David R. Hunter. Discussions on “A Brief Survey of Modern Optimization for Statisticians”: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(1):76–79, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [LCZ14a].

Hussain:1972:SFR

[Hus72] Ashiq Hussain. Some further remarks on a mixed model for regression. International Statistical Review = Revue Internationale de Statistique, 40(1):37–40, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402703.

Huzurbazar:1973:BRB

[Huz73] V. S. Huzurbazar. Book review: Families of Bivariate Distributions by K. V. Mardia. International Statistical Review = Revue Internationale de Statistique, 41(1): 139, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402797.

Huzurbazar:1980:BRB

[Huz80] V. S. Huzurbazar. Book review: Contributions to Survey Sampling and Applied Statistics: Papers in Honor of H. O. Hartley by H. A. David. International Statistical Review = Revue Internationale de Statistique, 48(2):234, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403166.

Hardle:2015:LBS

[HV15] Wolfgang Karl Härdle and Annette B. Vogt. Ladislaus von Bortkiewicz — statistician, economist and a European intellectual. International Statistical Review = Revue Internationale de Statistique, 83(1):17–35, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Heerschap:2006:TIS

[HW06] Nico Heerschap and Leon Willenborg. Towards an integrated statistical system at statistics Netherlands. International Statistical Review = Revue Internationale de Statistique, 74(3):357–378, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472714.

Hooten:2020:SIA

[HWS20] Mevin Hooten, Christopher Wikle, and Michael Schwob. Statistical implementations of agent-based demographic models. International Statistical Review = Revue Internationale de Statistique, 88(2):441–461, August 2020. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2001:IBS

[HY01] David J. Hand and Keming Yu. Idiot’s Bayes — not so stupid after all? International Statistical Review = Revue Internationale de Statistique, 69(3):385–398, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403452.

Hong-zhi:1991:KST

[HzB91] An Hong-zhi and Cheng Bing. A Kolmogorov–Smirnov type statistic with application to test for nonlinearity in time series. International Statistical Review = Revue Internationale de Statistique, 59(3):287–307, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403689.

Amari:2021:IG

[iA21] Shun ichi Amari. Information geometry. International Statistical Review = Revue Internationale de Statistique, 89(2):250–273, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Iachan:1982:SSC

[Iac82] Ronaldo Iachan. Systematic sampling: A critical review. International Statistical Review = Revue Internationale de Statistique, 50(3):293–303, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402499.

Iachan:1984:SSR

[Iac84] Ronaldo Iachan. Sampling strategies, robustness and efficiency: the state of the art. International Statistical Review = Revue Internationale de Statistique, 52(2):209– 218, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403103.

Ichim:2009:DCB

[Ich09] Daniela Ichim. Disclosure control of business microdata: A density-based approach. International Statistical Review = Revue Internationale de Statistique, 77(2):196– 211, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919722.

Imhof:1986:ODS

[IK86] J. P. Imhof and P. Kümmerling. Operational derivation of some Brownian motion results. International Statistical Review = Revue Internationale de Statistique, 54(3):327– 341, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403062.

Imrey:1981:CDA

[IKS^{+}81] Peter B. Imrey, Gary G. Koch, Maura E. Stokes, John N. Darroch, Daniel H. Freeman, Jr., and H. Dennis Tolley. Categorical data analysis: Some reflections on the log linear model and logistic regression. Part I: Historical and methodological overview. International Statistical Review = Revue Internationale de Statistique, 49(3):265– 283, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402613.

Imrey:1982:CDA

[IKS^{+}82] Peter B. Imrey, Gary G. Koch, Maura E. Stokes, John N. Darroch, Daniel H. Freeman, Jr., and H. Dennis Tolley. Categorical data analysis: Some reflections on the log linear model and logistic regression. Part II: Data analysis. International Statistical Review = Revue Internationale de Statistique, 50(1):35–63, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402458.

Imbens:2007:SBR

[Imb07] Guido W. Imbens. Short book review: Matched Sampling for Causal Effects by Donald B. Rubin. International Statistical Review = Revue Internationale de Statistique, 75 (2):257–258, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41508473.

Ishii:2022:ECO

[IMG22] Ryota Ishii, Kazushi Maruo, and Masahiko Gosho. Effect of covariate omission in randomised controlled trials: a review and simulation study. International Statistical Review = Revue Internationale de Statistique, 90(1): 100–117, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Institute:1973:WFS

[Ins73] International Statistical Institute. The World Fertility Survey: An international programme of fertility research. International Statistical Review = Revue Internationale de Statistique, 41(3):303–314, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402619.

Iosifescu:1972:BRB

[Ios72] M. Iosifescu. Book review: Dynamic Probabilistic Systems by R. A. Howard. International Statistical Review = Revue Internationale de Statistique, 40(2):251–252, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402771.

Iosifescu:1976:BRBc

[Ios76a] M. Iosifescu. Book review: A First Course in Mathematical Statistics by G. G. Roussas. International Statistical Review = Revue Internationale de Statistique, 44(2): 308, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403302.

Iosifescu:1976:BRBe

[Ios76b] M. Iosifescu. Book review: Advances in Probability and Related Topics. Vol. 3 by Peter Ney; Sidney Port. International Statistical Review = Revue Internationale de Statistique, 44(2):308, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403304.

Iosifescu:1976:BRBd

[Ios76c] M. Iosifescu. Book review: Introduction to Probability Theory and Statistical Inference by H. J. Larson. International Statistical Review = Revue Internationale de Statistique, 44(2):308, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403303.

Iosifescu:1976:BRBf

[Ios76d] M. Iosifescu. Book review: Numerical Methods in Markov Chains and Bulk Queues by T. P. Bagchi; J. G. C. Templeton. International Statistical Review = Revue Internationale de Statistique, 44(2):309, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403305.

Iosifescu:1976:BRBa

[Ios76e] M. Iosifescu. Book review: Stochastic Processes by R. Coleman. International Statistical Review = Revue Internationale de Statistique, 44(2):307, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403300.

Iosifescu:1976:BRBb

[Ios76f] M. Iosifescu. Book review: Success Epochs in Bernoulli Trials (With Applications in Number Theory) by W. Vervaat. International Statistical Review = Revue Internationale de Statistique, 44(2):307–308, August 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403301.

Iosifescu:1977:BRBa

[Ios77a] Marius Iosifescu. Book review: Normal Approximation and Asymptotic Expansions by R. N. Bhattacharya; R. Ranga Rao. International Statistical Review = Revue Internationale de Statistique, 45(3):321–322, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402546.

Iosifescu:1977:BRBb

[Ios77b] Marius Iosifescu. Book review: Processus aléatoires utilisés en recherche opérationnelle by D. Carton. International Statistical Review = Revue Internationale de Statistique, 45(3):322, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402547.

Iosifescu:1977:BRBd

[Ios77c] Marius Iosifescu. Book review: Studies in Probability and Statistics (Papers in Honour of Edwin J. G. Pitman) by E. J. Williams. International Statistical Review = Revue Internationale de Statistique, 45(3):327, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402558.

Iosifescu:1977:BRBc

[Ios77d] Marius Iosifescu. Book reviews: Probability and Statistics: Theory and Applications by D. A. S. Fraser; An Introduction to Probability Theory and Mathematical Statistics by V. K. Rohatgi. International Statistical Review = Revue Internationale de Statistique, 45(3):323, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402549.

Iosifescu:1978:BRB

[Ios78] Marius Iosifescu. Book review: Urn Models and Their Application: An Approach to Modern Discrete Probability Theory by Norman L. Johnson; Samuel Kotz. International Statistical Review = Revue Internationale de Statistique, 46(3):319–320, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402382.

Iosifescu:1979:BRBc

[Ios79a] Marius Iosifescu. Book review: Fractals: Form, Chance and Dimension, by Benoît B. Mandelbrot. International Statistical Review = Revue Internationale de Statistique, 47(3):299, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402650.

Iosifescu:1979:BRBa

[Ios79b] Marius Iosifescu. Book review: Information Theory, Statistical Decision Functions, Random Processes. Transactions of the 7th Prague Conference and of the European Meeting of Statisticians, 1974. Volume A by J. Kožešnik. International Statistical Review = Revue Internationale de Statistique, 47(2):231, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402573.

Iosifescu:1979:BRBb

[Ios79c] Marius Iosifescu. Book review: Introduction to Probability and Measure by K. R. Parthasarathy. International Statistical Review = Revue Internationale de Statistique, 47(2):233–234, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402578.

Iosifescu:1980:BRBa

[Ios80a] Marius Iosifescu. Book review: A First Course in Probability by Sheldon Ross. International Statistical Review = Revue Internationale de Statistique, 48(2):243, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403182.

Iosifescu:1980:BRBb

[Ios80b] Marius Iosifescu. Book review: Quelques problèmes concernant les sondages by Pierre D. Thionet. International Statistical Review = Revue Internationale de Statistique, 48(2):243, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403183.

Iosifescu:1986:ONO

[Ios86] Marius Iosifescu. Obituary notice: Octav Onicescu, 1892– 1983. International Statistical Review = Revue Internationale de Statistique, 54(1):97–108, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403261.

Ilmonen:2012:ICS

[IOS12] Pauliina Ilmonen, Hannu Oja, and Robert Serfling. On invariant coordinate system (ICS) functionals. International Statistical Review = Revue Internationale de Statistique, 80(1):93–110, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257168.

Isham:1981:ISP

[Ish81] Valerie Isham. An introduction to spatial point processes and Markov random fields. International Statistical Review = Revue Internationale de Statistique, 49(1): 21–43, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403035.

Izenman:2003:SID

[Ize03] Alan J. Izenman. Sentencing illicit drug traffickers: How do the courts handle random sampling issues? International Statistical Review = Revue Internationale de Statistique, 71(3):535–556, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403827.

Jabine:1996:DSD

[Jab96] Thomas B. Jabine. [developing samplers for developing countries]: Discussion. International Statistical Review = Revue Internationale de Statistique, 64(2):153–156, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403735.

Jacobsen:1984:MLE

[Jac84] Martin Jacobsen. Maximum likelihood estimation in the multiplicative intensity model: a survey. International Statistical Review = Revue Internationale de Statistique, 52(2):193–207, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403102.

Jagers:1986:PSA

[Jag86] Peter Jagers. Post-stratification against bias in sampling. International Statistical Review = Revue Internationale de Statistique, 54(2):159–167, August 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403141.

Jasra:2015:ABC

[Jas15] Ajay Jasra. Approximate Bayesian computation for a class of time series models. International Statistical Review = Revue Internationale de Statistique, 83(3):405–435, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Jauri:2014:BRB

[Jau14] Osmo Jauri. Book review: An Introduction to Exotic Option Pricing by Peter Buchen. International Statistical Review = Revue Internationale de Statistique, 82(1):149, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Jazairi:1972:FII

[Jaz72] N. T. Jazairi. Fisher ideal index reformulated. International Statistical Review = Revue Internationale de Statistique, 40(1):47–51, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402705.

Jarque:1987:TNO

[JB87] Carlos M. Jarque and Anil K. Bera. A test for normality of observations and regression residuals. International Statistical Review = Revue Internationale de Statistique, 55(2):163–172, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403192.

Jensen:1985:RTS

[JBGS85] E. B. Jensen, A. J. Baddeley, H. J. G. Gundersen, and R. Sundberg. Recent trends in stereology. International Statistical Review = Revue Internationale de Statistique, 53(1):99–108, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402883.

Jeffreys:1974:FIP

[Jef74] Harold Jeffreys. Fisher and inverse probability. International Statistical Review = Revue Internationale de Statistique, 42(1):1–3, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402679.

Jerez:1978:JAA

[Jer78] Miguel Jerez. José Antonio de Artigas y Sanz, 1887– 1977. International Statistical Review = Revue Internationale de Statistique, 46(2):233, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402825.

Juul:2014:ISH

[JF14] Svend Juul and Morten Frydenberg. An introduction to Stata for health researchers. StataCorp LP, College Station, TX, USA, fourth edition, 2014. ISBN 1-59718-135-8 (paperback). xx + 320 pp. LCCN R853.S7 J88 2014.

Johnson:1978:LMA

[JK78] William D. Johnson and Gary G. Koch. Linear models analysis of competing risks for grouped survival times. International Statistical Review = Revue Internationale de Statistique, 46(1):21–51, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402506.

Johnson:1982:DDD

[JK82] Norman L. Johnson and Samuel Kotz. Developments in discrete distributions, 1969–1980, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 50(1):71–101, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402460.

Jasra:2020:AMM

[JLS20] Ajay Jasra, Kody Law, and Carina Suciu. Advanced multilevel Monte Carlo methods. International Statistical Review = Revue Internationale de Statistique, 88(3): 548–579, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Jupp:1989:UVT

[JM89] P. E. Jupp and K. V. Mardia. A unified view of the theory of directional statistics, 1975–1988. International Statistical Review = Revue Internationale de Statistique, 57(3): 261–294, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403799.

Johnson:1977:SPR

[Joh77] L. W. Johnson. Stochastic parameter regression: an annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 45(3):257–272, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402540.

Johnson:1980:SPR

[Joh80] L. W. Johnson. Stochastic parameter regression: an additional annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 48(1): 95–102, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402409.

Johansen:1981:DPD

[Joh81] Søren Johansen. Discussion of paper by D. Oakes. International Statistical Review = Revue Internationale de Statistique, 49(3):258–262, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402610.

Johansen:1983:ECR

[Joh83] Søren Johansen. An extension of Cox’s regression model. International Statistical Review = Revue Internationale de Statistique, 51(2):165–174, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402746.

Johnson:2015:HHS

[Joh15] Timothy P. Johnson, editor. Handbook of health survey methods. Wiley handbooks in survey methodology. John Wiley, New York, NY, USA, 2015. ISBN 1-118-00232-6 (hardcover), 1-118-59474-6 (e-pub), 1-118-59460-6 (PDF). xxii + 816 pp. LCCN RA407 .H36

2015. URL http://onlinelibrary.wiley.com/book/
10.1002/9781118594629.

Joiner:1975:LH

[Joi75] Brian L. Joiner. Living histograms. International Statistical Review = Revue Internationale de Statistique, 43 (3):339–340, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403117.

Jones:2000:RSM

[Jon00] Hywel B. Jones. A review of statistical methods for genome mapping. International Statistical Review = Revue Internationale de Statistique, 68(1):5–21, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403554.

Jones:2007:CDP

[Jon07] M. C. Jones. Connecting distributions with power tails on the real line, the half line and the interval. International Statistical Review = Revue Internationale de Statistique, 75(1):58–69, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41508449.

Jones:2015:FDS

[Jon15a] M. C. Jones. On families of distributions with shape parameters. International Statistical Review = Revue Internationale de Statistique, 83(2):175–192, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See discussion [Abe15, AG15, LP15, Ma15, Pew15, SR15] and rejoinder [Jon15b].

Jones:2015:RFD

[Jon15b] M. C. Jones. Rejoinder to On Families of Distributions with Shape Parameters: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 83(2): 223–227, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Jorgensen:1984:DAG

[Jør84] Bent Jørgensen. The delta algorithm and GLIM. International Statistical Review = Revue Internationale de Statistique, 52(3):283–300, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403047.

Jorgensen:1992:EDM

[Jør92] Bent Jørgensen. Exponential dispersion models and extensions: A review. International Statistical Review = Revue Internationale de Statistique, 60(1):5–20, April 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403498.

Jagers:1985:PSR

[JOT85] P. Jagers, A. Odén, and L. Trulsson. Post-stratification and ratio estimation: usages of auxiliary information in survey sampling and opinion polls. International Statistical Review = Revue Internationale de Statistique, 53(3): 221–238, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402887.

Jensen:2008:AFR

[JR08] Donald R. Jensen and Donald E. Ramirez. Anomalies in the foundations of ridge regression. International Statistical Review = Revue Internationale de Statistique, 76(1): 89–105, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919585.

Jensen:2010:ARR

[JR10] Donald R. Jensen and Donald E. Ramirez. Anomalies in ridge regression: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 78(2):215– 217, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919833.

Jackle:2010:AED

[JRL10] Annette Jäckle, Caroline Roberts, and Peter Lynn. Assessing the effect of data collection mode on measurement. International Statistical Review = Revue Internationale de Statistique, 78(1):3–20, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919792.

Jeskanen-Sundstrom:2003:ISN

[JS03] Heli Jeskanen-Sundström. ICT statistics at the New Millennium — developing official statistics — measuring the diffusion of ICT and its impact. International Statistical Review = Revue Internationale de Statistique, 71(1): 5–15, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403870.

Jiang:2011:ANR

[JSW11] Yannan Jiang, Alastair J. Scott, and Christopher J. Wild. Adjusting for non-response in population-based case-control studies. International Statistical Review = Revue Internationale de Statistique, 79(2):145–159, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41305020.

Janssen:2013:ALT

[JT13] Arnold Janssen and Martin Tietje. Applications of the likelihood theory in finance: modelling and pricing. International Statistical Review = Revue Internationale de Statistique, 81(1):107–133, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kabaila:2009:CPC

[Kab09] Paul Kabaila. The coverage properties of confidence regions after model selection. International Statistical Review = Revue Internationale de Statistique, 77(3):405– 414, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919765.

Kafadar:2015:SIA

[Kaf15] Karen Kafadar. Statistical issues in assessing forensic evidence. International Statistical Review = Revue Internationale de Statistique, 83(1):111–134, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kageyama:1972:NTT

[Kag72a] Sanpei Kageyama. Note on Takeuchi’s table of difference sets generating balanced incomplete block designs. International Statistical Review = Revue Internationale de Statistique, 40(3):275–276, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402467.

Kageyama:1972:SRS

[Kag72b] Sanpei Kageyama. A survey of resolvable solutions of balanced incomplete block designs. International Statistical Review = Revue Internationale de Statistique, 40(3):269– 273, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402466.

Kalton:1983:MPS

[Kal83] Graham Kalton. Models in the practice of survey sampling. International Statistical Review = Revue Internationale de Statistique, 51(2):175–188, August 1983. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402747.

Kallenberg:1984:IGT

[Kal84] Olav Kallenberg. An informal guide to the theory of conditioning in point processes. International Statistical Review = Revue Internationale de Statistique, 52(2):151– 164, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403098.

Kalton:2005:DGK

[Kal05] Graham Kalton. Discussant: Graham Kalton. International Statistical Review = Revue Internationale de Statistique, 73(2):251–252, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kalton:2019:DSR

[Kal19] Graham Kalton. Developments in survey research over the past 60 years: a personal perspective. International Statistical Review = Revue Internationale de Statistique, 87(S1):S10–S30, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koch:1980:SVP

[KASG80] Gary G. Koch, Ingrid A. Amara, Maura E. Stokes, and Dennis B. Gillings. Some views on parametric and nonparametric analysis for repeated measurements and selected bibliography. International Statistical Review = Revue Internationale de Statistique, 48(3):249–265, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402940.

Kauppinen:2012:SBR

[Kau12] Joonas Kauppinen. Short book review: Music Data Mining edited by Tao Li, Mitsunori Ogihara, George Tzanetakis. International Statistical Review = Revue Internationale de Statistique, 80(1):189–190, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/23257184.

Kauppinen:2013:SBR

[Kau13] Joonas Kauppinen. Short book review: The R Student Companion by Brian Dennis. International Statistical Review = Revue Internationale de Statistique, 81(2):309, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kaye:2003:QCP

[Kay03] David H. Kaye. Questioning a courtroom proof of the uniqueness of fingerprints. International Statistical Review = Revue Internationale de Statistique, 71(3):521– 533, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403826.

Khmaladze:2010:MRS

[KBH10] Estate V. Khmaladze, Ray Brownrigg, and John Haywood. Memoryless reigns of the “sons of heaven”. International Statistical Review = Revue Internationale de Statistique, 78(3):348–362, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919858.

Konrad:2021:TLG

[KBM21] Anne Konrad, Jan Pablo Burgard, and Ralf Münnich. A two-level GREG estimator for consistent estimation in household surveys. International Statistical Review = Revue Internationale de Statistique, 89(3):635–656, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Keogh:2014:CCS

[KC14a] Ruth H. Keogh and D. R. (David Roxbee) Cox. Casecontrol studies. Institute of mathematical statistics monographs. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-01956-7 (hardcover), 1-139-69871-0 (PDF e-book). x + 285 pp. LCCN RA652 .K46 2014.

Kolaczyk:2014:SAN

[KC14b] Eric D. Kolaczyk and Gábor Csárdi. Statistical Analysis of Network Data with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4939-0982-7 (paperback). xiii + 207 pp. LCCN QA402 .K6483 2014.

Kokic:2000:MBP

[KCB00] Philip Kokic, Ray Chambers, and Steve Beare. Microsimulation of business performance. International Statistical Review = Revue Internationale de Statistique, 68(3):259– 275, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403413.

Kang:2020:VOM

[KDTP20] Xiaoning Kang, Xinwei Deng, Kam-Wah Tsui, and Mohsen Pourahmadi. On variable ordination of modified Cholesky decomposition for estimating time-varying covariance matrices. International Statistical Review = Revue Internationale de Statistique, 88(3):616–641, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Keiding:1981:DPD

[Kei81] Niels Keiding. Discussion of paper by D. Oakes. International Statistical Review = Revue Internationale de Statistique, 49(3):262, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402611.

Keiding:1987:MEN

[Kei87] Niels Keiding. The method of expected number of deaths, 1786–1886–1986, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 55 (1):1–20, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403267.

Keiding:2005:PS

[Kei05a] Niels Keiding. The professional statistician. International Statistical Review = Revue Internationale de Statistique, 73(2):271–272, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/25472668.

Keiding:2005:RSL

[Kei05b] Niels Keiding. Roles of statistics in the life sciences. International Statistical Review = Revue Internationale de Statistique, 73(2):255–258, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472665.

Kempthorne:1974:GWS

[Kem74] Oscar Kempthorne. George W. Snedecor, 1881–1974. International Statistical Review = Revue Internationale de Statistique, 42(3):319–323, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402994.

Kempthorne:1980:BVS

[Kem80] Oscar Kempthorne. Balkrishna V. Sukhatme, 1924–1979. International Statistical Review = Revue Internationale de Statistique, 48(1):137–138, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402444.

Kendall:1973:BRBb

[Ken73a] M. G. Kendall. Book review: A Bibliography of Multivariate Statistical Analysis by T. W. Anderson; S. das Gupta; G. P. H. Styan. International Statistical Review = Revue Internationale de Statistique, 41(2):289, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402850.

Kendall:1973:BRBa

[Ken73b] M. G. Kendall. Book review: Biometrika Tables for Statisticians, Vol. 2 by E. S. Pearson; H. O. Hartley. International Statistical Review = Revue Internationale de Statistique, 41(1):143, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402802.

Kendall:1973:EPI

[Ken73c] M. G. Kendall. Entropy, probability and information. International Statistical Review = Revue Internationale de Statistique, 41(1):59–68, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402787.

Kendall:1974:BRB

[Ken74] M. G. Kendall. Book review: Systematic Glossary of the Terminology of Statistical Methods by Isaac Paenson. International Statistical Review = Revue Internationale de Statistique, 42(2):231–232, August 1974. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403085.

Kendall:1975:BRB

[Ken75] Maurice Kendall. Book review: Symmetric Functions in Statistics: Proceedings of a Symposium in Honour of Professor Paul S. Dwyer by Derrick S. Tracy. International Statistical Review = Revue Internationale de Statistique, 43(3):370, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403128.

Kendall:1976:BRBa

[Ken76a] David G. Kendall. Book review: The Genetic Structure of Populations by Albert Jacquard; D. Charlesworth; B. Charlesworth. International Statistical Review = Revue Internationale de Statistique, 44(2):302–304, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403294.

Kendall:1976:BRBb

[Ken76b] M. G. Kendall. Book review: Exploring Data Analysis: The Computer Revolution in Statistics by W. J. Dixon; W. L. Nicholson. International Statistical Review = Revue Internationale de Statistique, 44(3):377, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402356.

Kendall:1976:BRBc

[Ken76c] M. G. Kendall. Book review: On the History of Statistics and Probability by D. B. Owen. International Statistical Review = Revue Internationale de Statistique, 44(3):379– 380, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402361.

Kenda:1980:BRB

[Ken80] Maurice Kenda. Book review: Characterisations of the Normal Law by A. M. Mathai; G. Pederzoli. International Statistical Review = Revue Internationale de Statistique, 48(1):128–130, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402432.

Kent:1986:DPG

[Ken86] J. T. Kent. Discussion of paper by L. A. Goodman. International Statistical Review = Revue Internationale de Statistique, 54(3):273–274, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403055.

Kenessey:1997:PAE

[Ken97] Zoltan Kenessey. A perspective on the accuracy of economic observations. International Statistical Review = Revue Internationale de Statistique, 65(2):247–259, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403345.

Kenward:2010:SBR

[Ken10] Mike Kenward. Short book review: Bayesian Missing Data Problems: EM, Data Augmentation and Noniterative Computation by Ming T. Tan, Guo-Liang Tian, Kai Wang Ng. International Statistical Review = Revue Internationale de Statistique, 78(3):479, December 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919894.

Kettenring:1997:D

[Ket97a] Jon Kettenring. Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2):153, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kettenring:1997:NPN

[Ket97b] Jon Kettenring. [new pedagogy and new content: The case of statistics]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2):153, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403337.

Kettenring:2012:SRB

[Ket12] Jon R. Kettenring. Statistics research at Bell Labs in the regulated monopoly era. International Statistical Review = Revue Internationale de Statistique, 80(2):205–218, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Keuning:1997:SIE

[Keu97] Steven J. Keuning. SESAME: An integrated economic and social accounting system. International Statistical Review = Revue Internationale de Statistique, 65(1):111– 121, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403436.

Koch:1975:SMA

[KFF75] Gary G. Koch, Daniel H. Freeman, Jr., and Jean L. Freeman. Strategies in the multivariate analysis of data from complex surveys. International Statistical Review = Revue Internationale de Statistique, 43(1):59–78, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402660.

Koch:1976:GMA

[KFL76] Gary G. Koch, Jean L. Freeman, and Robert G. Lehnen. A general methodology for the analysis of ranked policy preference data. International Statistical Review = Revue Internationale de Statistique, 44(1):1–28, April 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402963.

Kapat:2010:AFR

[KG10] Prasenjit Kapat and Prem K. Goel. Anomalies in the foundations of ridge regression: Some clarifications. International Statistical Review = Revue Internationale de Statistique, 78(2):209–215, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919832.

Khamis:1973:BRB

[Kha73] S. H. Khamis. Book review: An Introduction to Statistical Science in Agriculture by D. J. Finney. International Statistical Review = Revue Internationale de Statistique, 41(3):398, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402634.

Khatri:1974:BRB

[Kha74] C. G. Khatri. Book review: Families of Frequency Distributions by J. K. Ord. International Statistical Review = Revue Internationale de Statistique, 42(2):231, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403084.

Khamis:1977:MS

[Kha77] Salem H. Khamis. El-Mehdi Said, 1921–1976. International Statistical Review = Revue Internationale de Statistique, 45(2):208, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403330.

Khamis:1980:BRB

[Kha80] Salem H. Khamis. Book review: International Comparisons of Real Product and Purchasing Power by Irving B. Kravis; Alan Heston; Robert Summers. International Statistical Review = Revue Internationale de Statistique, 48 (2):239–240, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403176.

Khattree:2015:BRA

[Kha15] Ravindra Khattree. Book review: Analysis of Multivariate and High-Dimensional Data. International Statistical Review = Revue Internationale de Statistique, 83(1): 169–171, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Khattree:2017:BRP

[Kha17] Ravindra Khattree. Book review: Principal Component Analysis Networks and Algorithms, Xiangyu Kong, Changhua Hu and Zhansheng Duan Springer, 2017, xxii + 323 pages, £112.00, hardcover ISBN: 978-981-10-2913-4. International Statistical Review = Revue Internationale de Statistique, 85(3):543–544, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Khattree:2020:BRE

[Kha20] Ravi Khattree. Book review: End-to-End Data Science with SAS: A Hands-On Programming Guide. James Gearheart SAS Press, 2020, xiii + 363 pages, £23.99/$29.99, e-book ISBN: 978-1-64295-806-5. International Statistical Review = Revue Internationale de Statistique, 88(3):809– 812, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Khattree:2024:BRP

[Kha24] Ravindra Khattree. Book review: Post-Shrinkage Strategies in Statistical and Machine Learning for High Dimensional Data, Syed Ejaz Ahmed, Feryaal Ahmed, Bahadir Yüzbasi. Chapman & Hall/CRC Press, 2023, 408 pages, £125/$160, hardback ISBN: 978-0-367-76344-2. International Statistical Review = Revue Internationale de Statistique, 92(2):322–323, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Khuri:2000:DVC

[Khu00] André I. Khuri. Designs for variance components estimation: Past and present. International Statistical Review = Revue Internationale de Statistique, 68(3):311– 322, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403416.

Khuri:2013:SBRb

[Khu13a] André I. Khuri. Short book review: Introduction to Linear Regression Analysis, Fifth Edition, by Douglas C. Montgomery, Elizabeth A. Peck, G. Geoffrey Vining. International Statistical Review = Revue Internationale de Statistique, 81(2):318–319, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Khuri:2013:SBRa

[Khu13b] André I. Khuri. Short book review: SAS Statistics by Example by Ron Cody. International Statistical Review = Revue Internationale de Statistique, 81(1):164–165, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kim:2014:BRR

[Kim14] KyungMann Kim. Book review: Randomized Phase II Cancer Clinical Trials by Sin-Ho Jung. International Statistical Review = Revue Internationale de Statistique, 82 (2):309–311, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kim:2015:BRS

[Kim15] Hyungjin Myra Kim. Book review: Sample Size Determination and Power. International Statistical Review = Revue Internationale de Statistique, 83(1):168–169, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kingman:1975:BRBa

[Kin75a] J. F. C. Kingman. Book review: Approximate Stochastic Behaviour of n-Server Service Systems with Large n by G. F. Newell. International Statistical Review = Revue Internationale de Statistique, 43(2):248–249, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402922.

Kingman:1975:BRBb

[Kin75b] J. F. C. Kingman. Book review: Letters on Probability by Alfréd Rényi. International Statistical Review = Revue Internationale de Statistique, 43(2):249, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402923.

Kingman:1975:BRBc

[Kin75c] J. F. C. Kingman. Book review: Mathematics and Statistics: Essays in Honour of Harald Bergström by P. Jagers; L. Råde. International Statistical Review = Revue Internationale de Statistique, 43(3):370, December 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403129.

Kingman:1977:BRBb

[Kin77a] J. F. C. Kingman. Book review: Stochastic Population Theories by Donald Ludwig. International Statistical Review = Revue Internationale de Statistique, 45(1): 99–100, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403017.

Kingman:1977:BRBa

[Kin77b] J. F. C. Kingman. Book review: The Mathematical Theory of Infectious Diseases by Norman T. J. Bailey. International Statistical Review = Revue Internationale de Statistique, 45(1):95–96, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403009.

Kiregyera:1982:SSN

[Kir82] Ben Kiregyera. Some sources of nonsampling errors in African census and survey work: A review. International Statistical Review = Revue Internationale de Statistique, 50(3):305–316, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402500.

Kiregyera:2005:STB

[Kir05] Ben Kiregyera. Statistical training to break the vicious cycle of statistical under-development. International Statistical Review = Revue Internationale de Statistique, 73 (2):215–216, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/25472659.

Kiregyera:2011:D

[Kir11] Ben Kiregyera. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(3):330– 331, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kish:1979:SC

[Kis79] Leslie Kish. Samples and censuses. International Statistical Review = Revue Internationale de Statistique, 47(2): 99–109, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402563.

Kish:1994:MSD

[Kis94] Leslie Kish. Multipopulation survey designs: Five types with seven shared aspects. International Statistical Review = Revue Internationale de Statistique, 62(2):167– 186, August 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403507.

Kish:1996:DSD

[Kis96] Leslie Kish. Developing samplers for developing countries. International Statistical Review = Revue Internationale de Statistique, 64(2):143–152, August 1996. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403734.

Kitada:2003:OSI

[Kit03] Hiroyuki Kitada. Official statistics on ICT in Japan. International Statistical Review = Revue Internationale de Statistique, 71(1):69–82, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403875.

Kannan:2021:CRR

[KK21] Nandini Kannan and Debasis Kundu. C. Radhakrishna Rao: a century in statistical science. International Statistical Review = Revue Internationale de Statistique, 89(2): 216–236, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kott:2009:OSC

[KL09] Phillip S. Kott and Yan K. Liu. One-sided coverage intervals for a proportion estimated from a stratified simple random sample. International Statistical Review = Revue Internationale de Statistique, 77(2):251–265, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919726.

Klevmarken:1976:BRB

[Kle76] Anders Klevmarken. Book review: Linear Aggregation in Linear Regression by Harry Lütjohann. International Statistical Review = Revue Internationale de Statistique, 44(1):170–173, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402974.

Klein:1978:JM

[Kle78] L. R. Klein. Jacob Marschak, 1898–1977. International Statistical Review = Revue Internationale de Statistique, 46(3):326, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402395.

Kloek:1983:DPD

[Klo83] T. Kloek. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):153–155, August 1983. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402742.

Kruger:2021:PIB

[KLTG21] Fabian Krüger, Sebastian Lerch, Thordis Thorarinsdottir, and Tilmann Gneiting. Predictive inference based on Markov Chain Monte Carlo output. International Statistical Review = Revue Internationale de Statistique, 89(2): 274–301, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kruskal:1979:RSN

[KM79a] William Kruskal and Frederick Mosteller. Representative sampling, i: Non-scientific literature. International Statistical Review = Revue Internationale de Statistique, 47 (1):13–24, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403202.

Kruskal:1979:RSIa

[KM79b] William Kruskal and Frederick Mosteller. Representative sampling, II: Scientific literature, excluding statistics. International Statistical Review = Revue Internationale de Statistique, 47(2):111–127, August 1979. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402564.

Kruskal:1979:RSIb

[KM79c] William Kruskal and Frederick Mosteller. Representative sampling, III: The current statistical literature. International Statistical Review = Revue Internationale de Statistique, 47(3):245–265, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402647.

Kruskal:1980:RSI

[KM80] William Kruskal and Frederick Mosteller. Representative sampling. IV. The history of the concept in statistics, 1895–1939. International Statistical Review = Revue Internationale de Statistique, 48(2):169–195, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403151.

Kinyoki:2017:MEC

[KMM^{+}17] Damaris K. Kinyoki, Samuel O. Manda, Grainne M. Moloney, Elijah O. Odundo, James A. Berkley, Abdisalan M. Noor, and Ngianga-Bakwin Kandala. Modelling the ecological comorbidity of acute respiratory infection, diarrhoea and stunting among children under the age of 5 years in Somalia. International Statistical Review = Revue Internationale de Statistique, 85(1):164–176, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Keller-McNulty:2005:PMC

[KMNS05] Sallie Keller-McNulty, Charles W. Nakhleh, and Nozer D. Singpurwalla. A paradigm for masking (camouflaging) information. International Statistical Review = Revue Internationale de Statistique, 73(3):331–349, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472679.

Kulinskaya:2014:CSE

[KMS14] Elena Kulinskaya, Stephan Morgenthaler, and Robert G. Staudte. Combining statistical evidence. International Statistical Review = Revue Internationale de Statistique, 82(2):214–242, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kuhnert:2003:BGB

[KMT03] Petra M. Kuhnert, Kerrie Mengersen, and Peter Tesar. Bridging the gap between different statistical approaches: An integrated framework for modelling. International Statistical Review = Revue Internationale de Statistique, 71 (2):335–368, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403893.

Knuppel:1980:BRB

[Knü80] Lothar Knüppel. Book review: Mathematical Models in Marketing. A Collection of Abstracts. Lecture Notes in Economics and Mathematical Systems. Vol. 132 by U. H. Funke. International Statistical Review = Revue Internationale de Statistique, 48(1):122–123, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402423.

Koch:2014:AMH

[Koc14] Inge Koch. Analysis of Multivariate and High-Dimensional Data. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-88793-3 (hardcover). xxv + 504 pp. LCCN QA278 .K5935 2014. URL http://assets.

cambridge.org/97805218/87939/cover/9780521887939.
jpg.

Koeze:1979:ASE

[Koe79] P. Koeze. An accurate statistical estimation of the lifelength of f100-banknotes: A circulation trial with two qualities of currency paper. International Statistical Review = Revue Internationale de Statistique, 47(3):283– 297, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402649.

Koenker:2016:DPI

[Koe16] Roger Koenker. Discussions on posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Discussion. International Statistical Review = Revue Internationale de Statistique, 84(3):345–349, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Kokko:2007:SBR

[Kok07] Hanna Kokko. Short book review: Mathematical Models of Social Evolution: A Guide for the Perplexed by Richard McElreath, Robert Boyd. International Statistical Review = Revue Internationale de Statistique, 75(3): 416, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509887.

Koskinen:2009:SBRa

[Kos09a] Lasse Koskinen. Short book review: Mathematical Asset Management by Thomas Höglund. International Statistical Review = Revue Internationale de Statistique, 77 (1):149, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919698.

Koskinen:2009:SBRb

[Kos09b] Lasse Koskinen. Short book review: Statistical Models and Methods for Financial Markets by Tze Leung Lai, Haipeng Xing. International Statistical Review = Revue Internationale de Statistique, 77(1):154–155, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919704.

Koskinen:2010:SBRb

[Kos10a] Lasse Koskinen. Short book review: A First Course in Bayesian Statistical Methods by Peter D. Hoff. International Statistical Review = Revue Internationale de Statistique, 78(1):150, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919815.

Koskinen:2010:SBRa

[Kos10b] Lasse Koskinen. Short book review: Handbook of Financial Time Series edited by Torben G. Andersen, Richard A. Davis, Jens-Peter Kreiß, Thomas Mikosch. International Statistical Review = Revue Internationale de Statistique, 78(1):148–149, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919814.

Koskinen:2010:SBRc

[Kos10c] Lasse Koskinen. Short book review: Marginal Models: For Dependent, Clustered, and Longitudinal Categorical Data by Wicher Bergsma, Marcel Croon, Jacques A. Hagenaars. International Statistical Review = Revue Internationale de Statistique, 78(1):150–151, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919816.

Koskinen:2011:SBR

[Kos11] Lasse Koskinen. Short book review: Regression Modeling with Actuarial and Financial Applications by Edward W. Frees. International Statistical Review = Revue Internationale de Statistique, 79(1):137–138, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41306189.

Koskinen:2012:SBRa

[Kos12a] Lasse Koskinen. Short book review: Copula Theory and Its Applications edited by Piotr Jaworski, Fabrizio Durante, Wolfgang Härdle and Tomasz Rychlik. International Statistical Review = Revue Internationale de Statistique, 80(2):328, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2012:SBRb

[Kos12b] Lasse Koskinen. Short book review: Dependence in Probability and Statistics edited by Paul Doukhan, Gabriel Lang, Donatas Surgailis, Gilles Teyssière. International Statistical Review = Revue Internationale de Statistique, 80(2):329, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2012:SBRc

[Kos12c] Lasse Koskinen. Short book review: Statistical Tools for Finance and Insurance, Second Edition edited by Pavel Čížek, Wolfgang Karl Härdle and Rafał Weron. International Statistical Review = Revue Internationale de Statistique, 80(2):329–330, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2013:SBRc

[Kos13a] Lasse Koskinen. Short book review: Computational Methods in Finance by Ali Hirsa. International Statistical Review = Revue Internationale de Statistique, 81(3):480– 481, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2013:SBRd

[Kos13b] Lasse Koskinen. Short book review: Computational Statistics, Second Edition, by Geof H. Givens, Jennifer A. Hoeting. International Statistical Review = Revue Internationale de Statistique, 81(3):481–482, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2013:SBRb

[Kos13c] Lasse Koskinen. Short book review: Handbook of Computational Finance by Jin-Chuan Duan, Wolfgang Karl Härdle, and James E. Gentle. International Statistical Review = Revue Internationale de Statistique, 81(3):479– 480, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Koskinen:2013:SBRa

[Kos13d] Lasse Koskinen. Short book review: Risk Modelling in General Insurance: From Principles to Practice by Roger J. Gray, Susan M. Pitts. International Statistical Review = Revue Internationale de Statistique, 81(3):478–479, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kotz:1974:CSD

[Kot74] Samuel Kotz. Characterizations of statistical distributions: a supplement to recent surveys. International Statistical Review = Revue Internationale de Statistique, 42 (1):39–65, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402684.

Keilman:2004:TSB

[KP04] Nico Keilman and Dinh Quang Pham. Time series based errors and empirical errors in fertility forecasts in the Nordic countries. International Statistical Review = Revue Internationale de Statistique, 72(1):5–18, April 2004. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403839.

Kim:2010:CES

[KP10] Jae Kwang Kim and Mingue Park. Calibration estimation in survey sampling. International Statistical Review = Revue Internationale de Statistique, 78(1):21–39, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919793.

Kovacs:2005:NMM

[KPT05] Péter Kovàcs, Tibor Petres, and László Tóth. A new measure of multicollinearity in linear regression models. International Statistical Review = Revue Internationale de Statistique, 73(3):405–412, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472683.

Krickeberg:1974:BRB

[Kri74] Klaus Krickeberg. Book review: Mathematische Statistik mit industriellen Anwendungen by I. Vincze. International Statistical Review = Revue Internationale de Statistique, 42(1):110, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402694.

Krotki:1980:BRBb

[Kró80a] Karol J. Krótki. Book review: A Short Textbook of Medical Statistics by Austin Bradford Hill. International Statistical Review = Revue Internationale de Statistique, 48(1): 124–125, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402426.

Krotki:1980:BRBd

[Kró80b] Karol J. Krótki. Book review: Risks, Benefits, and Controversies in Fertility Control by John J. Sciarra; Gerald I. Zatuchni; J. Joseph Speidel. International Statistical Review = Revue Internationale de Statistique, 48(3):373– 374, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402959.

Krotki:1980:BRBc

[Kro80c] Karol J. Krotki. Book review: Stationaere und schrumpfende Bevoelkerungen. Demographisches Nullund Negativwachstum in Oesterreich by Gustav Feichtinger. International Statistical Review = Revue Internationale de Statistique, 48(2):235–237, August 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403170.

Krotki:1980:BRBa

[Kró80d] Karol J. Krótki. Book review: Statistics: A Biomedical Introduction by Byron W. Brown,; Myles Hollander. International Statistical Review = Revue Internationale de Statistique, 48(1):119, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402416.

Kinney:2011:TUP

[KRR^{+}11] Satkartar K. Kinney, Jerome P. Reiter, Arnold P. Reznek, Javier Miranda, Ron S. Jarmin, and John M. Abowd. Towards unrestricted public use business microdata: The synthetic longitudinal business database. International Statistical Review = Revue Internationale de Statistique, 79(3):362–384, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305056.

Kruskal:1989:HYR

[Kru89] William Kruskal. Hooker and Yule on relative importance: A statistical detective story. International Statistical Review = Revue Internationale de Statistique, 57 (1):83–88, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403586.

Khuri:1985:VCA

[KS85] A. I. Khuri and Hardeo Sahai. Variance components analysis: a selective literature survey. International Statistical Review = Revue Internationale de Statistique, 53(3):279– 300, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402893.

Knowles:1989:HAT

[KS89a] Mark Knowles and David Siegmund. On Hotelling’s approach to testing for a nonlinear parameter in regression. International Statistical Review = Revue Internationale de Statistique, 57(3):205–220, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403794.

Kuchler:1989:EFS

[KS89b] Uwe Küchler and Michael Sørensen. Exponential families of stochastic processes: A unifying semimartingale approach. International Statistical Review = Revue Internationale de Statistique, 57(2):123–144, August 1989. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403382.

Kulkarni:2022:ETH

[KS22] Hemangi V. Kulkarni and Ashis SenGupta. An efficient test for homogeneity of mean directions on the hypersphere. International Statistical Review = Revue Internationale de Statistique, 90(1):41–61, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Krishnaiah:1972:SRT

[KSRS72] P. R. Krishnaiah, H. Schubnell, R. H. Randles, and A. J. Schaeffer. Statistical research and training. International Statistical Review = Revue Internationale de Statistique, 40(1):91–94, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402710.

Kardaun:2003:RFCb

[KSS^{+}03a] O. J. W. F. Kardaun, D. Salomé, W. Schaafsma, A. G. M. Steerneman, and J. C. Willems. [reflections on fourteen cryptic issues concerning the nature of statistical inference]: Response by Groningen statisticians. International Statistical Review = Revue Internationale de Statistique, 71(2):314–318, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403890.

Kardaun:2003:RFCa

[KSS^{+}03b] O. J. W. F. Kardaun, D. Salomè, W. Schaafsma, A. G. M. Steerneman, J. C. Willems, and D. R. Cox. Reflections on fourteen cryptic issues concerning the nature of statistical inference. International Statistical Review = Revue Internationale de Statistique, 71(2):277–303, August 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403887.

Kohler:2014:RCB

[KSS14] Max Köhler, Anja Schindler, and Stefan Sperlich. A review and comparison of bandwidth selection methods for kernel regression. International Statistical Review = Revue Internationale de Statistique, 82(2):243–274, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kafadar:1993:UCD

[KT93] Karen Kafadar and John W. Tukey. U.S. cancer death rates: a simple adjustment for urbanization. International Statistical Review = Revue Internationale de Statistique, 61(2):257–281, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403628.

Kuha:2003:CME

[KT03] Jouni Kuha and Jonathan Temple. Covariate measurement error in quadratic regression. International Statistical Review = Revue Internationale de Statistique, 71(1): 131–150, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403878.

Kim:2021:DIC

[KT21] Jae-Kwang Kim and Siu-Ming Tam. Data integration by combining big data and survey sample data for finite population inference. International Statistical Review = Revue Internationale de Statistique, 89(2):382–401, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kitagawa:2001:SEP

[KTM01] Genshiro Kitagawa, Tetsuo Takanami, and Norio Matsumoto. Signal extraction problems in seismology. International Statistical Review = Revue Internationale de Statistique, 69(1):129–152, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403533.

Kuk:2020:NPH

[Kuk20] Anthony Y. C. Kuk. A non-proportional hazards model with hazard ratio functions free from covariate values. International Statistical Review = Revue Internationale de Statistique, 88(3):715–727, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kuk:2022:EMH

[Kuk22] Anthony Y. C. Kuk. Estimating monotonic hazard ratio functions of time. International Statistical Review = Revue Internationale de Statistique, 90(2):285–305, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kulldorff:1978:CEQ

[Kul78] Gunnar Kulldorff. Carl-Erik Quensel, 1907–1977. International Statistical Review = Revue Internationale de Statistique, 46(2):233–234, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402826.

Kim:2019:STB

[KW19] Jae Kwang Kim and Zhonglei Wang. Sampling techniques for big data analysis. International Statistical Review = Revue Internationale de Statistique, 87(S1):S177–S191, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Karlis:2005:MPD

[KX05] Dimitris Karlis and Evdokia Xekalaki. Mixed Poisson distributions. International Statistical Review = Revue Internationale de Statistique, 73(1):35–58, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472639.

Kuk:2009:REF

[KX09] Anthony Y. C. Kuk and Jinfeng Xu. Relative efficiency of the fuzzy p-value approach to hypothesis testing. International Statistical Review = Revue Internationale de Statistique, 77(3):395–404, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919764.

Koreisha:1991:CAI

[KY91] Sergio Koreisha and Gary Yoshimoto. A comparison among identification procedures for autoregressive moving average models. International Statistical Review = Revue Internationale de Statistique, 59(1):37–57, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403573.

Killion:1976:BCT

[KZ76] Ruth A. Killion and Douglas A. Zahn. A bibliography of contingency table literature: 1900 to 1974. International Statistical Review = Revue Internationale de Statistique, 44(1):71–112, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402965.

LaVecchia:2016:SAE

[La 16] Davide La Vecchia. Stable asymptotics for M-estimators. International Statistical Review = Revue Internationale de Statistique, 84(2):267–290, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Leunis:1996:LAN

[LA96] Wim P. Leunis and Jan Willem Altena. Labour accounts in The Netherlands, 1987–1993: How to cope with fragmented macro data in official statistics. International Statistical Review = Revue Internationale de Statistique, 64 (1):1–22, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403421.

Laird:2022:SAL

[Lai22] Nan M. Laird. Statistical analysis of longitudinal studies. International Statistical Review = Revue Internationale de Statistique, 90:S2–S16, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lancaster:1972:BSB

[Lan72] H. O. Lancaster. A bibliography of statistical bibliographies: a fifth list. International Statistical Review = Revue Internationale de Statistique, 40(1):73–81, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402708.

Lancaster:1973:BSB

[Lan73] H. O. Lancaster. A bibliography of statistical bibliographies: a sixth list. International Statistical Review = Revue Internationale de Statistique, 41(3):375–379, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402625.

Lancaster:1974:BSBa

[Lan74a] H. O. Lancaster. A bibliography of statistical bibliographies: a seventh list. International Statistical Review = Revue Internationale de Statistique, 42(1):67–70, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402685.

Lancaster:1974:BSBb

[Lan74b] H. O. Lancaster. A bibliography of statistical bibliographies: an eighth list. International Statistical Review = Revue Internationale de Statistique, 42(3):307–311, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402989.

Lancaster:1975:BSB

[Lan75] H. O. Lancaster. A bibliography of statistical bibliographies: a ninth list. International Statistical Review = Revue Internationale de Statistique, 43(3):345–349, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403119.

Lancaster:1976:BSB

[Lan76a] H. O. Lancaster. A bibliography of statistical bibliographies: a tenth list. International Statistical Review = Revue Internationale de Statistique, 44(3):361–365, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402353.

Lanke:1976:DPR

[Lan76b] Jan Lanke. On the degree of protection in randomized interviews. International Statistical Review = Revue Internationale de Statistique, 44(2):197–203, August 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403277.

International Association of Survey Statisticians Meeting on Randomized Response (Warsaw, 1975).

Lancaster:1978:BSB

[Lan78] H. O. Lancaster. A bibliography of statistical bibliographies: an eleventh list. International Statistical Review = Revue Internationale de Statistique, 46(2):221–224, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402816.

Lancaster:1979:BSB

[Lan79] H. O. Lancaster. A bibliography of statistical bibliographies: a twelfth list. International Statistical Review = Revue Internationale de Statistique, 47(1):79–83, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403209.

Lancaster:1981:BSB

[Lan81] H. O. Lancaster. A bibliography of statistical bibliographies: a thirteenth list. International Statistical Review = Revue Internationale de Statistique, 49(2):177–183, August 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403071.

Lancaster:1982:BSB

[Lan82] H. O. Lancaster. A bibliography of statistical bibliographies: a fourteenth list. International Statistical Review = Revue Internationale de Statistique, 50(2):195–217, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402601.

Lancaster:1983:BSB

[Lan83] H. O. Lancaster. A bibliography of statistical bibliographies: A fifteenth list. International Statistical Review = Revue Internationale de Statistique, 51(2):207–212, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402749.

Lancaster:1984:BSB

[Lan84] H. O. Lancaster. A bibliography of statistical bibliographies: a sixteenth list. International Statistical Review = Revue Internationale de Statistique, 52(1):101–108, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403249.

Lancaster:1985:BSB

[Lan85] H. O. Lancaster. A bibliography of statistical bibliographies: A seventeenth list. International Statistical Review = Revue Internationale de Statistique, 53(2):215– 220, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402936.

Lancaster:1986:BSB

[Lan86] H. O. Lancaster. A bibliography of statistical bibliographies: an eighteenth list. International Statistical Review = Revue Internationale de Statistique, 54(1):109– 114, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403262.

Lancaster:1987:BSB

[Lan87] H. O. Lancaster. A bibliography of statistical bibliographies: a nineteenth list. International Statistical Review = Revue Internationale de Statistique, 55(2):221–227, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403197.

Lanterman:2001:SWR

[Lan01] Aaron D. Lanterman. Schwarz, Wallace, and Rissanen: Intertwining themes in theories of model selection. International Statistical Review = Revue Internationale de Statistique, 69(2):185–212, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403813.

Lan:2018:BRS

[Lan18] Ling Lan. Book review: Statistical Methods for Evaluating Safety in Medical Product Development, A. Lawrence Gould, John Wiley and Sons, 2015, 400 pages, $99.95, hardcover ISBN: 978-1-119-97966-1. International Statistical Review = Revue Internationale de Statistique, 86(3): 604–605, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Larget:2015:BRB

[Lar15] Bret Larget. Book review: Bayesian Phylogenetics: Methods, Algorithms and Applications. International Statistical Review = Revue Internationale de Statistique, 83(1): 160–161, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lauritzen:1981:TSA

[Lau81] Steffen L. Lauritzen. Time series analysis in 1880: a discussion of contributions made by T. N. Thiele. International Statistical Review = Revue Internationale de Statistique, 49(3):319–331, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402616.

Lavielle:2014:MEM

[Lav14] Marc Lavielle. Mixed Effects Models for the Population Approach: Models, Tasks, Methods and Tools. Chapman and Hall/CRC biostatistics series. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2014. ISBN 1-4822-2650-2 (hardcover). xviii + 365 pp. LCCN QA278 .L38 2014.

Lawrance:1991:DRT

[Law91] A. J. Lawrance. Directionality and reversibility in time series. International Statistical Review = Revue Internationale de Statistique, 59(1):67–79, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403575.

Lawless:1998:SAP

[Law98] J. F. Lawless. Statistical analysis of product warranty data. International Statistical Review = Revue Internationale de Statistique, 66(1):41–60, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403656.

Lawless:2007:SBR

[Law07] Jerry F. Lawless. Short book review: Reliability, Life Testing and the Prediction of Service Lives: For Engineers and Scientists by Sam C. Saunders. International Statistical Review = Revue Internationale de Statistique, 75(2): 260, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508476.

Loyal:2020:SNA

[LC20] Joshua Daniel Loyal and Yuguo Chen. Statistical network analysis: A review with applications to the coronavirus disease 2019 pandemic. International Statistical Review = Revue Internationale de Statistique, 88(2):419–440, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Li:2023:OTR

[LCL^{+}23] Zhen Li, Jie Chen, Eric Laber, Fang Liu, and Richard Baumgartner. Optimal treatment regimes: a review and empirical comparison. International Statistical Review = Revue Internationale de Statistique, 91(3):427–463, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lange:2014:BSM

[LCZ14a] Kenneth Lange, Eric C. Chi, and Hua Zhou. A brief survey of modern optimization for statisticians. International Statistical Review = Revue Internationale de Statistique, 82(1):46–70, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See comments [AM14, Hun14, Rob14] and rejoinder [LCZ14b].

Lange:2014:R

[LCZ14b] Kenneth Lange, Eric C. Chi, and Hua Zhou. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 82(1):81–89, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [LCZ14a].

Levakova:2024:PMF

[LD24] Marie Levakova and Susanne Ditlevsen. Penalisation methods in fitting high-dimensional cointegrated vector autoregressive models: A review. International Statistical Review = Revue Internationale de Statistique, 92(2): 160–193, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Leite:1990:BES

[LdBP90] José Galvão Leite and Carlos Alberto de Bragança Pereira. Bayes estimation of the size of a finite population: Capture/recapture sequential sample data. International Statistical Review = Revue Internationale de Statistique, 58 (3):201–213, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403804.

LeCam:1990:MLI

[Le 90a] L. Le Cam. Maximum likelihood: An introduction. International Statistical Review = Revue Internationale de Statistique, 58(2):153–171, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403464.

LeCam:1990:SAC

[Le 90b] L. Le Cam. On the standard asymptotic confidence ellipsoids of Wald. International Statistical Review = Revue Internationale de Statistique, 58(2):129–152, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403463.

Leadbetter:1988:HC

[Lea88] M. R. Leadbetter. Harald Cramér, 1893–1985. International Statistical Review = Revue Internationale de Statistique, 56(1):89–97, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403364.

Leclerc:1976:ERE

[Lec76] Annette Leclerc. Une étude de la relation entre une variable qualitative et un groupe de variables qualitatives. (French) [A study of the relationship between a variable and a group of qualitative variables ]. International Statistical Review = Revue Internationale de Statistique, 44(2): 241–248, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403285.

Lee:1979:CMD

[Lee79] Jack C. Lee. Computations of multivariate distributions done at the aerospace research laboratories. International Statistical Review = Revue Internationale de Statistique, 47(1):37–45, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403206.

Lee:2001:ICT

[Lee01] Thomas C. M. Lee. An introduction to coding theory and the two-part minimum description length principle. International Statistical Review = Revue Internationale de Statistique, 69(2):169–183, August 2001. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403812.

Lee:2005:DJC

[Lee05] Jae Chang Lee. Discussant: Jae Chang Lee. International Statistical Review = Revue Internationale de Statistique, 73(2):231–232, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lefort:1979:CHS

[Lef79] G. Lefort. Choix d’une Hypothèse dans une suite finie: Méthode et applications. (French) [Choosing a hypothesis in a finite sequence: method and applications]. International Statistical Review = Revue Internationale de Statistique, 47(2):137–154, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402566.

Lazar:2001:SIF

[LEGW01] Nicole A. Lazar, William F. Eddy, Christopher R. Genovese, and Joel Welling. Statistical issues in fMRI for brain imaging. International Statistical Review = Revue Internationale de Statistique, 69(1):105–127, April 2001. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403532.

Lehmann:1978:HS

[Leh78] E. L. Lehmann. Henry Scheffé, 1907–1977. International Statistical Review = Revue Internationale de Statistique, 46(1):126, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402530.

Leppisaari:2014:BRB

[Lep14] Matias Leppisaari. Book review: An Introduction to Analysis of Financial Data with R by Ruey S. Tsay. International Statistical Review = Revue Internationale de Statistique, 82(1):142–143, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lutz:2004:IHD

[LG04] Wolfgang Lutz and Joshua R. Goldstein. Introduction: How to deal with uncertainty in population forecasting? International Statistical Review = Revue Internationale de Statistique, 72(1):1–4, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403838.

Liu:2008:APC

[LG08] Juxin Liu and Paul Gustafson. On average predictive comparisons and interactions. International Statistical Review = Revue Internationale de Statistique, 76(3):419– 432, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919655.

Liu:2013:MPK

[LgX13] Regina Liu and Min ge Xie. In memory of Professor Kesar Singh. International Statistical Review = Revue Internationale de Statistique, 81(1):2, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Leiter:1973:SBP

[LH73] Robert E. Leiter and M. A. Hamdan. Some bivariate probability models applicable to traffic accidents and fatalities. International Statistical Review = Revue Internationale de Statistique, 41(1):87–100, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402790.

Long:2004:DOS

[LH04] John F. Long and Frederick W. Hollmann. Developing official stochastic population forecasts at the US Census Bureau. International Statistical Review = Revue Internationale de Statistique, 72(2):201–208, August 2004. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403854.

Landis:1978:APA

[LHK78] J. Richard Landis, Eugene R. Heyman, and Gary G. Koch. Average partial association in three-way contingency tables: a review and discussion of alternative tests.

International Statistical Review = Revue Internationale de Statistique, 46(3):237–254, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402373.

Liu:2015:CAM

[LHSD15] Zhihui (Amy) Liu, James A. Hanley, Olli Saarela, and Nandini Dendukuri. A conditional approach to measure mortality reductions due to cancer screening. International Statistical Review = Revue Internationale de Statistique, 83(3):493–510, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lian:2020:ANP

[Lia20] Heng Lian. Asymptotics of the non-parametric function for B-splines-based estimation in partially linear models. International Statistical Review = Revue Internationale de Statistique, 88(1):142–154, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liddell:1980:BRB

[Lid80] F. D. K. Liddell. Book review: Mortality and Morbidity in the United States by Carl L. Erhardt; Joyce E. Berlin. International Statistical Review = Revue Internationale de Statistique, 48(1):121, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402420.

Linder:1975:BRBa

[Lin75a] A. Linder. Book review: Automatische Klassifikation. Theoretische und praktische Methoden zur Gruppierung und Strukturierung von Daten (Cluster-Analyse) by Hans Hermann Bock. International Statistical Review = Revue Internationale de Statistique, 43(2):242, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402912.

Linder:1975:BRBb

[Lin75b] A. Linder. Book review: Computer: Aufgaben im Gesundheitswesen by N. Hollberg; B. Pleuss; H. Rittersbacher. International Statistical Review = Revue Internationale de Statistique, 43(3):371, December 1975. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403130.

Lindley:1976:B

[Lin76] D. V. Lindley. I. I. Berenblut, 1918–1975. International Statistical Review = Revue Internationale de Statistique, 44(3):384–385, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402365.

Linder:1980:BRBa

[Lin80a] A. Linder. Book review: Analyse statistique à plusieurs variables by Pierre Dagnelie. International Statistical Review = Revue Internationale de Statistique, 48(1):119– 120, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402417.

Linder:1980:BRBb

[Lin80b] A. Linder. Book review: Introduction to Statistics and Probability by Edward J. Dudewicz. International Statistical Review = Revue Internationale de Statistique, 48 (1):120, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402418.

Linder:1980:BRBc

[Lin80c] A. Linder. Book review: Statistics: A Beginning by Roy R. Kuebler; Harry Smith. International Statistical Review = Revue Internationale de Statistique, 48(1):126, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402430.

Lindley:1982:SRIb

[Lin82a] D. V. Lindley. [scoring rules and the inevitability of probability]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 50(1): 24–26, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402456.

Lindley:1982:SRIa

[Lin82b] Dennis V. Lindley. Scoring rules and the inevitability of probability. International Statistical Review = Revue Internationale de Statistique, 50(1):1–11, April 1982. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402448.

With discussion and a reply by the author.

Lindley:2002:SDC

[Lin02] Dennis V. Lindley. Seeing and doing: The concept of causation. International Statistical Review = Revue Internationale de Statistique, 70(2):191–197, August 2002. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403902.

See comment [Sin02a].

Linde:2013:SBR

[Lin13] Peter Linde. Short book review: Statistics for Non-Statisticians by Birger Madsen. International Statistical Review = Revue Internationale de Statistique, 81(1): 161–162, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lin:2014:PPF

[Lin14a] Xihong Lin, editor. Past, Present, and Future of Statistical Science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4822-0496-7 (hardcover). xxiii + 622 pp. LCCN QA276 .P276 2014.

Linde:2014:BRI

[Lin14b] Peter Linde. Book review: Incomplete Categorical Data Design: Non-Randomized Response Techniques for Sensitive Questions in Surveys by Guo-Liang Tian and Man-Lai Tang. International Statistical Review = Revue Internationale de Statistique, 82(3):485–486, December 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liski:2007:SBRb

[Lis07a] Erkki P. Liski. Short book review: An Introduction to Categorical Data Analysis, 2nd Edition, by Alan Agresti. International Statistical Review = Revue Internationale de Statistique, 75(3):414, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509885.

Liski:2007:SBRc

[Lis07b] Erkki P. Liski. Short book review: An Introduction to Categorical Data Analysis, 2nd Edition, by Alan Agresti. International Statistical Review = Revue Internationale de Statistique, 75(3):414, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liski:2007:SBRa

[Lis07c] Erkki P. Liski. Short book review: Information and Complexity in Statistical Modeling by Jorma Rissanen. International Statistical Review = Revue Internationale de Statistique, 75(2):249, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508464.

Liski:2008:SBRb

[Lis08a] Erkki P. Liski. Short book review: A Matrix Handbook for Statisticians by George A. F. Seber. International Statistical Review = Revue Internationale de Statistique, 76(3): 439, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919659.

Liski:2008:SBRa

[Lis08b] Erkki P. Liski. Short book review: Information Criteria and Statistical Modeling by Sadanori Konishi, Genshiro Kitagawa. International Statistical Review = Revue Internationale de Statistique, 76(2):313, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919632.

Liski:2009:SBRc

[Lis09a] Erkki P. Liski. Short book review: Applied Statistics for Engineers and Physical Scientists, Third Edition, by Johannes Ledolter, Robert V. Hogg. International Statistical Review = Revue Internationale de Statistique, 77(3): 481, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919785.

Liski:2009:SBRb

[Lis09b] Erkki P. Liski. Short book review: Confidence Intervals in Generalized Regression Models by Esa Uusipaikka. International Statistical Review = Revue Internationale de Statistique, 77(1):161, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919712.

Liski:2009:SBRa

[Lis09c] Erkki P. Liski. Short book review: Model Selection and Model Averaging by Gerda Claeskens, Nils Lid Hjort. International Statistical Review = Revue Internationale de Statistique, 77(1):156, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919706.

Liski:2010:SBRc

[Lis10a] Erkki P. Liski. Short book review: Design of Observational Studies by Paul R. Rosenbaum. International Statistical Review = Revue Internationale de Statistique, 78 (3):477–478, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919892.

Liski:2010:SBRa

[Lis10b] Erkki P. Liski. Short book review: From Finite Sample to Asymptotic Methods in Statistics by Pranab K. Sen, Julio M. Singer, Antonio C. Pedroso de Lima. International Statistical Review = Revue Internationale de Statistique, 78(2):318, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919843.

Liski:2010:SBRb

[Lis10c] Erkki P. Liski. Short book review: Statistical Modelling and Regression Structures: Festschrift in Honour of Ludwig Fahrmeir edited by Thomas Kneib, Gerhard Tutz. International Statistical Review = Revue Internationale de Statistique, 78(3):466, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919882.

Liski:2012:SBR

[Lis12] Erkki P. Liski. Short book review: Principles of Uncertainty by Joseph B. Kadane. International Statistical Review = Revue Internationale de Statistique, 80(1): 190, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/23257185.

Liski:2013:SBR

[Lis13] Erkki P. Liski. Short book review: Generalized Linear Mixed Models: Modern Concepts, Methods and Applications by Walter W. Stroup. International Statistical Review = Revue Internationale de Statistique, 81(3):482– 483, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liski:2014:BRB

[Lis14] Erkki P. Liski. Book review: Log-linear Modeling: Concepts, Interpretation, and Application by Alexander von Eye, Eun-Young Mun. International Statistical Review = Revue Internationale de Statistique, 82(1):144–145, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Little:1986:SNA

[Lit86] Roderick J. A. Little. Survey nonresponse adjustments for estimates of means. International Statistical Review = Revue Internationale de Statistique, 54(2):139–157, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403140.

Liu:2007:SBR

[Liu07] Shuangzhe Liu. Short book review: Festschrift for Tarmo Pukkila on his 60th Birthday edited by Erkki P. Liski, Jarkko Isotalo, Jarmo Niemelä, Simo Puntanen, George P. H. Styan. International Statistical Review = Revue Internationale de Statistique, 75(3):419–420, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41509890.

Liu:2008:SBR

[Liu08] Shuangzhe Liu. Short book review: Global Sensitivity Analysis: The Primer by Andrea Saltelli, Marco Ratto, Terry Andres, Francesca Campolongo, Jessica Cariboni, Debora Gatelli, Michaela Saisana, Stefano Tarantola. International Statistical Review = Revue Internationale de Statistique, 76(3):452, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919672.

Liu:2010:SBRa

[Liu10a] Shuangzhe Liu. Short book review: Dynamic Model Analysis: Advanced Matrix Methods and Unit-Root Econometrics Representation Theorems, Second Edition, by Mario Faliva, Maria Grazia Zoia. International Statistical Review = Revue Internationale de Statistique, 78(1):136– 137, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919802.

Liu:2010:SBRb

[Liu10b] Shuangzhe Liu. Short book review: The Alberta High School Mathematics Competitions 1957–2006: A Canadian Problem Book edited by Andy Liu. International Statistical Review = Revue Internationale de Statistique, 78(3):468, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919884.

Liu:2011:SBR

[Liu11] Shuangzhe Liu. Short book review: GARCH Models: Structure, Statistical Inference and Financial Applications by Christian Francq, Jean-Michel Zakoian. International Statistical Review = Revue Internationale de Statistique, 79(2):301, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41305051.

Liu:2012:SBRa

[Liu12a] Shuangzhe Liu. Short book review: Introduction to General and Generalized Linear Models by Henrik Madsen, Poul Thyregod. International Statistical Review = Revue Internationale de Statistique, 80(1):183–184, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257179.

Liu:2012:SBRb

[Liu12b] Shuangzhe Liu. Short book review: Maximum Likelihood Estimation and Inference: With Examples in R, SAS, and ADMB by Russell B. Millar. International Statistical Review = Revue Internationale de Statistique, 80(2):346, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2014:BRBa

[Liu14a] Shuangzhe Liu. Book review: Combinatorial Matrix Theory and Generalized Inverses of Matrices by Ravindra B. Bapat, Steve J. Kirkland, K. Manjunatha Prasad, Simo Puntanen. International Statistical Review = Revue Internationale de Statistique, 82(1):143–144, April 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2014:BRBb

[Liu14b] Shuangzhe Liu. Book review: Formulas Useful for Linear Regression Analysis and Related Matrix Theory: It’s Only Formulas But We Like Them by Simo Puntanen, George P. H. Styan, Jarkko Isotalo. International Statistical Review = Revue Internationale de Statistique, 82(1): 148–149, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2015:BRC

[Liu15] Shuangzhe Liu. Book review: Circular Statistics in R. International Statistical Review = Revue Internationale de Statistique, 83(1):164, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2016:BRA

[Liu16] Shuangzhe Liu. Book review: Analysis of Categorical Data with R. International Statistical Review = Revue Internationale de Statistique, 84(1):162–163, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2017:BRM

[Liu17] Shuangzhe Liu. Book review: Modelling Discrete Time-to-Event Data, Gerhard Tutz and Matthias, Schmid Springer International Publishing, 2016, x + 247 pages, $89.99, hardcover ISBN: 978-3-319-28156-8. International Statistical Review = Revue Internationale de Statistique, 85(3): 542, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2018:BRM

[Liu18] Shuangzhe Liu. Book review: Modern Data Science with R, Benjamin S. Baumer, Daniel T. Kaplan and Nicholas J. Horton, Chapman and Hall/CRC, 2017, 556 pages, $99.95, hardcover ISBN: 978-1-4987-2448-7. International Statistical Review = Revue Internationale de Statistique, 86(1):162, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2019:BRS

[Liu19] Shuangzhe Liu. Book review: Statistical Inference for Engineers and Data Scientists, Pierre Moulin and Venugopal V. Veeravalli, Cambridge University Press, 2018, 418 pages, £54.99, hardcover, ISBN: 978-1-107-18592-0. International Statistical Review = Revue Internationale de Statistique, 87(2):448–449, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2020:BRC

[Liu20a] Shuangzhe Liu. Book review: A Concise Introduction to Machine Learning. A. C. Faul. CRC Press, 2019, 314 pages, £46.99, paperback ISBN: 978-0-8153-8410-6. International Statistical Review = Revue Internationale de Statistique, 88(2):517–518, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2020:BRM

[Liu20b] Shuangzhe Liu. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics, 3rd Edition, Jan R. Magnus and Heinz Neudecker, John Wiley and Sons, 2019, 504 pages, $115, hardcover; $92.99 ebook, ISBN: 978-1-119-54116-5. International Statistical Review = Revue Internationale de Statistique, 88(1):

261–263, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2021:BRS

[Liu21] Shuangzhe Liu. Book review: Statistical Machine Learning: A Unified Framework. Richard M. Golden, Chapman and Hall/CRC, 2020, 524 pages, £71.99, hardcover. ISBN: 978-1-138-48469-6. International Statistical Review = Revue Internationale de Statistique, 89(1):210–212, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2022:BRD

[Liu22a] Shuangzhe Liu. Book review: Data Visualization for Social and Policy Research: a Step-by-Step Approach Using R and Python. Jose Manuel Magallanes Reyes. Cambridge University Press, 2022, 292 pages, $105, hardback. ISBN: 978-1-108-49433-5. International Statistical Review = Revue Internationale de Statistique, 90(3):626–627, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2022:BRG

[Liu22b] Shuangzhe Liu. Book reviews: Game Data Science. Magy Seif El-Nasr, Truong-Huy D. Nguyen, Alessandro Canossa, Anders Drachen, Oxford University Press, 2022, xvi + 416 pages, $105 (hardback)/$55 (paperback) ISBN-10: 0-19-289787-X, ISBN-13: 978-0-19-289787-9 (hardback), 978-0-19-289788-6 (paperback). International Statistical Review = Revue Internationale de Statistique, 90 (2):412–414, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2023:BRM

[Liu23] Shuangzhe Liu. Book review: Modern Applied Regressions: Bayesian and Frequentist Analysis of Categorical and Limited Response Variables with R and Stan, Jun Xu, Chapman & Hall/CRC, 2023, xv + 281 pages, £80.99/ $108, hardcover ISBN: 978-0-367-17387-6 (hbk); 978-1-032-37674-5 (pbk); 978-0-429-05646-8 (ebk). International Statistical Review = Revue Internationale de Statistique, 91(2):345–347, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lawless:2011:D

[LK11] Jerald F. Lawless and John D. Kalbfleisch. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):225–228, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lee:2016:LPI

[LK16] Youngjo Lee and Gwangsu Kim. H-likelihood predictive intervals for unobservables. International Statistical Review = Revue Internationale de Statistique, 84(3):487– 505, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lazar:2004:CDT

[LKC^{+}04] Nicole A. Lazar, Joseph B. Kadane, Fang Chen, William G. Cavanagh, and Clifford D. Litton. Corbelled domes in two and three dimensions: The treasury of Atreus. International Statistical Review = Revue Internationale de Statistique, 72(2):239–255, August 2004. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403856.

Lawrance:1987:HOR

[LL87] A. J. Lawrance and P. A. W. Lewis. Higher-order residual analysis for nonlinear time series with autoregressive correlation structures. International Statistical Review = Revue Internationale de Statistique, 55(1):21–35, April 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403268.

Lecoutre:2001:UAM

[LLP01] Bruno Lecoutre, Marie-Paule Lecoutre, and Jacques Poitevineau. Uses, abuses and misuses of significance tests in the scientific community: Won’t the Bayesian choice be unavoidable? International Statistical Review = Revue Internationale de Statistique, 69(3):399–417, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403453.

Liu:2008:CES

[LLP08] Wei Liu, Shan Lin, and Walter W. Piegorsch. Construction of exact simultaneous confidence bands for a simple linear regression model. International Statistical Review = Revue Internationale de Statistique, 76(1):39–57, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919583.

Li:2004:ULC

[LLT04] Nan Li, Ronald Lee, and Shripad Tuljapurkar. Using the Lee–Carter method to forecast mortality for populations with limited data. International Statistical Review = Revue Internationale de Statistique, 72(1):19–36, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403840.

Levy:1974:PDP

[LN74] Kenneth J. Levy and Subhash C. Narula. Probability density plots of the non-central t-distribution. International Statistical Review = Revue Internationale de Statistique, 42(3):305–306, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402988.

Levy:1977:NCM

[LN77] Kenneth J. Levy and Subhash C. Narula. A note concerning matching by rank. International Statistical Review = Revue Internationale de Statistique, 45(2):159– 161, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403313.

Levy:1978:THC

[LN78] Kenneth J. Levy and Subhash C. Narula. Testing hypotheses concerning partial correlations: some methods and discussion. International Statistical Review = Revue Internationale de Statistique, 46(2):215–218, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402814.

Landenna:1980:SSF

[LN80] G. Landenna and A. Naddeo. La statistique et le statisticien. (French) [Statistics and the statistician]. International Statistical Review = Revue Internationale de Statistique, 48(2):197–204, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403153.

Levy:1975:ECM

[LNA75] Kenneth J. Levy, Subhash C. Narula, and Patrick F. Abrami. An empirical comparison of the methods of least squares and unweighted means for the analysis of disproportionate cell data. International Statistical Review = Revue Internationale de Statistique, 43(3):335– 338, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403116.

Leiwen:2004:TNM

[LO04] Jiang Leiwen and Brian C. O’Neill. Toward a new model for probabilistic household forecasts. International Statistical Review = Revue Internationale de Statistique, 72 (1):51–64, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403842.

Lewis:1985:IGA

[LOD^{+}85] P. A. W. Lewis, E. J. Orav, H. W. Drueg, D. G. Linnebur, and L. Uribe. An implementation of graphical analysis in statistical simulations. International Statistical Review = Revue Internationale de Statistique, 53(1):69–90, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402881.

Loh:2014:FYC

[Loh14a] Wei-Yin Loh. Fifty years of classification and regression trees. International Statistical Review = Revue Internationale de Statistique, 82(3):329–348, December 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Loh:2014:RFY

[Loh14b] Wei-Yin Loh. Rejoinder to fifty years of classification and regression trees: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 82(3):367– 370, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Longford:2015:SSI

[Lon15] Nicholas T. Longford. Statistical studies of income, poverty and inequality in Europe: computing and graphics in R using EU-SILC. Chapman and Hall/CRC statistics in the social and behavioral sciences series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-6832-1 (hardcover). xxii + 354 pp. LCCN HC79.I5 L646 2015.

URL http://images.tandf.co.uk/common/jackets/
websmall/978146656/9781466568327.jpg.

Louis:1985:DAT

[Lou85] Thomas A. Louis. Discussion of ‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage. International Statistical Review = Revue Internationale de Statistique, 53(1):28–31, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402874.

Lovison:2000:GSM

[Lov00] Gianfranco Lovison. Generalized symmetry models for hypercubic concordance tables. International Statistical Review = Revue Internationale de Statistique, 68(3):323– 338, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403417.

Loynes:1976:AOR

[Loy76] R. M. Loynes. Asymptotically optimal randomized response procedures: An abstract. International Statistical Review = Revue Internationale de Statistique, 44(2): 225, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403282.

Loynes:1977:BRB

[Loy77] R. M. Loynes. Book review: The Analysis of Time Series: Theory and Practice by C. Chatfield. International Statistical Review = Revue Internationale de Statistique, 45(1):97, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403012.

Loynes:1980:BRB

[Loy80] R. M. Loynes. Book review: Methods for Statistical Data Analysis of Multivariate Observations by R. Gnanadesikan. International Statistical Review = Revue Internationale de Statistique, 48(2):237, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403171.

Ley:2015:DFD

[LP15] Christophe Ley and Davy Paindaveine. Discussion on On Families of Distributions with Shape Parameters: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(2):202–207, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Lundstrom:2004:MFT

[LQ04] Hans Lundström and Jan Qvist. Mortality forecasting and trend shifts: An application of the Lee–Carter model to Swedish mortality data. International Statistical Review = Revue Internationale de Statistique, 72(1):37–50, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403841.

Lindgren:1991:SMR

[LR91] Georg Lindgren and Igor Rychlik. Slepian models and regression approximations in crossing and extreme value theory. International Statistical Review = Revue Internationale de Statistique, 59(2):195–225, August 1991. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403443.

leRoux:2005:AYM

[lRG05] Niël J. le Roux and Sugnet Gardner. Analysing your multivariate data as a pictorial: A case for applying biplot methodology? International Statistical Review = Revue Internationale de Statistique, 73(3):365–387, December 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472681.

Letac:1986:CCI

[LS86] G. Letac and V. Seshadri. On a conjecture concerning inverse Gaussian regression. International Statistical Review = Revue Internationale de Statistique, 54(2):187– 190, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403143.

Lutz:2004:PPP

[LS04] Wolfgang Lutz and Sergei Scherbov. Probabilistic population projections for India with explicit consideration of the education-fertility link. International Statistical Review = Revue Internationale de Statistique, 72(1):81–92, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403844.

Lencina:2006:MME

[LS06] Viviana B. Lencina and Julio M. Singer. Measure for measure: Exact f tests and the mixed models controversy. International Statistical Review = Revue Internationale de Statistique, 74(3):391–402, December 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472716.

Lee:2020:REM

[LS20] Minji Lee and Zhihua Su. A review of envelope models. International Statistical Review = Revue Internationale de Statistique, 88(3):658–676, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lumley:2011:CBS

[LSD^{+}11] Thomas Lumley, Pamela A. Shaw, James Y. Dai, Anastasios A. Tsiatis, Marie Davidian, Mark S. Handcock, Jerald F. Lawless, John D. Kalbfleisch, Alastair J. Scott, and Christopher J. Wild. Connections between survey calibration estimators and semiparametric models for incomplete data [with discussions]. International Statistical Review = Revue Internationale de Statistique, 79(2):200– 220, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305022.

Leysieffer:1980:RASb

[LSS80a] Frederick W. Leysieffer, I. Richard Savage, and Richard J. Sylvester. Reprinted articles in statistics and probability theory: a bibliography. II. International Statistical Review = Revue Internationale de Statistique, 48(3):289– 307, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402942.

Leysieffer:1980:RASa

[LSS80b] Frederick W. Leysieffer, I. Richard Savage, and Richard J. Sylvester. Reprinted articles in statistics and probability theory: A bibliography: Part I. International Statistical Review = Revue Internationale de Statistique, 48(2):205– 228, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403154.

Lencina:2005:MAA

[LSS05] Viviana B. Lencina, Julio M. Singer, and Edward J. Stanek III. Much ado about nothing: the mixed models controversy revisited. International Statistical Review = Revue Internationale de Statistique, 73(1):9–20, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472637.

Lencina:2008:LER

[LSS08] Viviana B. Lencina, Julio M. Singer, and Edward J. Stanek III. Letter to the Editors: Response. International Statistical Review = Revue Internationale de Statistique, 76(1):135–139, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lagakos:1982:DPP

[LT82] S. W. Lagakos and A. A. Tsiatis. Discussion of paper by P. K. Andersen, Ø. Borgan, R. Gill and N. Keiding. International Statistical Review = Revue Internationale de Statistique, 50(3):254–255, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402494.

Longva:1998:RCC

[LTS98] Svein Longva, Ib Thomsen, and Paul Inge Severeide. Reducing costs of censuses in Norway through use of administrative registers. International Statistical Review = Revue Internationale de Statistique, 66(2):223–234, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403491.

Lumley:2011:R

[Lum11] Thomas Lumley. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 79(2):230– 232, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Luoma:2007:SBR

[Luo07] Arto Luoma. Short book review: Testing 1–2–3: Experimental Design with Applications in Marketing and Service Operations by Johannes Ledolter, Arthur J. Swersey. International Statistical Review = Revue Internationale de Statistique, 75(3):431, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509903.

Luoma:2009:SBR

[Luo09] Arto Luoma. Short book review: Simulation and Inference for Stochastic Differential Equations: With R Examples by Stefano M. Iacus. International Statistical Review = Revue Internationale de Statistique, 77(1):163, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919714.

Luther:1976:VL

[Lut76] Georg Luther. Valter Lindberg, 1897–1976. International Statistical Review = Revue Internationale de Statistique, 44(3):386, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402367.

Ley:2014:LPO

[LV14] Christophe Ley and Thomas Verdebout. Local powers of optimal one-sample and multi-sample tests for the concentration of Fisher-von Mises-Langevin distributions. International Statistical Review = Revue Internationale de Statistique, 82(3):440–456, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Langaskens:1974:NME

[LvR74] Y. Langaskens and W. van Rijckeghem. A new method to estimate measurement errors in national account statistics: The Belgian case. International Statistical Review = Revue Internationale de Statistique, 42(3):283–290, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402985.

Lee:2013:SMS

[LW13] Elisa T. Lee and John Wenyu Wang. Statistical Methods for Survival Data Analysis. John Wiley, New York, NY, USA, fourth edition, 2013. ISBN 1-118-09502-2 (hardcover), 1-118-59311-1. xii + 484 pp. LCCN R853.S7 L43 2013.

Lyberg:2000:RBS

[Lyb00] Lars Lyberg. [the role of behavioral survey methodologists in national statistical agencies]: Discussion. International Statistical Review = Revue Internationale de Statistique, 68(2):216–217, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403672.

Lange:2022:LEA

[LZ22] Kenneth Lange and Hua Zhou. A legacy of EM algorithms. International Statistical Review = Revue Internationale de Statistique, 90:S52–S66, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Matalik:2005:USM

[MA05] Ivan Matalík and Josef Arlt. Use of statistics in the monetary policy of the Czech National Bank: The case of a country in transition. International Statistical Review = Revue Internationale de Statistique, 73(1):59–72, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472640.

Ma:2015:DFD

[Ma15] Yanyuan Ma. Discussion on On Families of Distributions with Shape Parameters: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(2): 207–211, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Ma:2017:VSP

[Ma17] Chunsheng Ma. Vector stochastic processes with Pólyatype correlation structure. International Statistical Review = Revue Internationale de Statistique, 85(2):340–354, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Macura:1976:SO

[Mac76] Miloš Macura. Sava Obradović, 1900–1972. International Statistical Review = Revue Internationale de Statistique, 44(1):180, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402980.

MacGregor:1997:ULP

[Mac97] John F. MacGregor. Using on-line process data to improve quality: Challenges for statisticians. International Statistical Review = Revue Internationale de Statistique, 65(3):309–323, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403374.

MacDonald:2014:NML

[Mac14] Iain L. MacDonald. Numerical maximisation of likelihood: A neglected alternative to EM? International Statistical Review = Revue Internationale de Statistique, 82(2):296– 308, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

MacKay:2016:DES

[Mac16] Jock MacKay. Discussions on enhancing statistics education by including qualitative research: Discussion: Acquiring statistical literacy and thinking. International Statistical Review = Revue Internationale de Statistique, 84(2): 189–194, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Magnello:2009:KPE

[Mag09] M. Eileen Magnello. Karl Pearson and the establishment of mathematical statistics. International Statistical Review = Revue Internationale de Statistique, 77(1): 3–29, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919687.

Maindonald:2007:SBR

[Mai07] John Maindonald. Short book review: Probability and Statistics for Computer Scientists by Michael Baron. International Statistical Review = Revue Internationale de Statistique, 75(2):276, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508491.

Maindonald:2008:SBR

[Mai08] John H. Maindonald. Short book review: Interactive and Dynamic Graphics for Data Analysis — With R and GGobi by Dianne Cook; Deborah F. Swayne. International Statistical Review = Revue Internationale de Statistique, 76(3):437–438, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919657.

Maindonald:2009:SBRc

[Mai09a] John H. Maindonald. Short book review: Model Based Inference in the Life Sciences: A Primer on Evidence by David R. Anderson. International Statistical Review = Revue Internationale de Statistique, 77(3):479–480, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919784.

Maindonald:2009:SBRb

[Mai09b] John H. Maindonald. Short book review: Morphometrics with R by Julien Claude; Michel Baylac; Emmanuel Paradis; Tristan Stayton. International Statistical Review = Revue Internationale de Statistique, 77(2):302– 303, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919731.

Maindonald:2009:SBRa

[Mai09c] John H. Maindonald. Short book review: Time Series Analysis With Applications in R, Second Edition, by Jonathan D. Cryer, Kung-Sik Chan. International Statistical Review = Revue Internationale de Statistique, 77(2): 300–301, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919729.

Maindonald:2010:SBRc

[Mai10a] John H. Maindonald. Short book review: Introductory Time Series with R by Paul S. P. Cowpertwait, Andrew V. Metcalfe. International Statistical Review = Revue Internationale de Statistique, 78(3):452–453, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919869.

Maindonald:2010:SBRb

[Mai10b] John H. Maindonald. Short book review: Probability with R: An Introduction with Computer Science Applications by Jane Horgan. International Statistical Review = Revue Internationale de Statistique, 78(1):146, April 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919811.

Maindonald:2010:SBRa

[Mai10c] John H. Maindonald. Short book review: Statistics and Data with R: An Applied Approach Through Examples by Yosef Cohen, Jeremiah Y. Cohen. International Statistical Review = Revue Internationale de Statistique, 78(1): 141–142, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919806.

Maindonald:2011:SBRc

[Mai11a] John H. Maindonald. Short book review: Bayesian Artificial Intelligence, Second Edition, by Kevin B. Korb, Ann E. Nicholson. International Statistical Review = Revue Internationale de Statistique, 79(3):497, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305078.

Maindonald:2011:SBRb

[Mai11b] John H. Maindonald. Short book review: Environmental Social Sciences: Methods and Research Design edited by Ismael Vaccaro, Eric Alden Smith, Shankar Aswani. International Statistical Review = Revue Internationale de Statistique, 79(3):496, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41305077.

Maindonald:2011:SBRd

[Mai11c] John H. Maindonald. Short book review: Exploratory Multivariate Analysis by Example Using R by François Husson, Sébastien Lê, Jérôme Pagès. International Statistical Review = Revue Internationale de Statistique, 79(3): 498, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305079.

Maindonald:2011:SBRa

[Mai11d] John H. Maindonald. Short book review: Principles and Theory for Data Mining and Machine Learning by Bertrand Clarke, Ernest Fokoué, Hao Helen Zhang. International Statistical Review = Revue Internationale de Statistique, 79(1):119–120, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41306172.

Maindonald:2011:SBRe

[Mai11e] John H. Maindonald. Short book review: Qualitative Research from Start to Finish by Robert K. Yin. International Statistical Review = Revue Internationale de Statistique, 79(3):499–500, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41305080.

Maindonald:2011:SBRf

[Mai11f] John H. Maindonald. Short book review: Recursive Partitioning and Applications, Second Edition, by Heping Zhang, Burton H. Singer. International Statistical Review = Revue Internationale de Statistique, 79(3):500– 501, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305081.

Maindonald:2012:SBRa

[Mai12a] John H. Maindonald. Short book review: Data Mining with Rattle and R: The Art of Excavating Data for Knowledge Discovery by Graham Williams. International Statistical Review = Revue Internationale de Statistique, 80(1): 199–200, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/23257194.

Maindonald:2012:SBRb

[Mai12b] John H. Maindonald. Short book review: R for SAS and SPSS Users, Second Edition, by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 80(1):200–201, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/23257195.

Maindonald:2012:SBRc

[Mai12c] John H. Maindonald. Short book review: Statistical Learning for Biomedical Data by James D. Malley, Karen G. Malley, Sinisa Pajevic. International Statistical Review = Revue Internationale de Statistique, 80(3):476–478, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2013:SBRb

[Mai13a] John H. Maindonald. Short book review: Applied Meta-Analysis for Social Science Research by Noel A. Card. International Statistical Review = Revue Internationale de Statistique, 81(1):165–166, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2013:SBRc

[Mai13b] John H. Maindonald. Short book review: Data Analysis and Data Mining: An Introduction by Adelchi Azzalini, Bruno Scarpa. International Statistical Review = Revue Internationale de Statistique, 81(1):170–171, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2013:SBRa

[Mai13c] John H. Maindonald. Short book review: Statistical Thinking in Epidemiology by Yu-Kang Tu, Mark S. Gilthorpe. International Statistical Review = Revue Internationale de Statistique, 81(1):154, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2014:BRBb

[Mai14a] John H. Maindonald. Book review: A First Course in Probability and Markov Chains by Giuseppe Modica, Laura Poggiolini. International Statistical Review = Revue Internationale de Statistique, 82(1):148, April 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2014:BRBa

[Mai14b] John H. Maindonald. Book review: Behavioral Research Data Analysis with R by Yuelin Li, Jonathan Baron. International Statistical Review = Revue Internationale de Statistique, 82(1):147–148, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2014:BRL

[Mai14c] John H. Maindonald. Book review: Linear Mixed-Effects Models Using R by Andrzej Galecki, Tomasz Burzykowski. International Statistical Review = Revue Internationale de Statistique, 82(2):315–317, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2015:BRP

[Mai15] John H. Maindonald. Book review: Past, Present, and Future of Statistical Science. International Statistical Review = Revue Internationale de Statistique, 83(2):332–334, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2016:BRM

[Mai16a] John H. Maindonald. Book review: Missing Data Analysis in Practice. International Statistical Review = Revue Internationale de Statistique, 84(2):322–323, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2016:BRO

[Mai16b] John H. Maindonald. Book review: Ordered Regression Models: Parallel, Partial, and Non-Parallel Alternatives, Andrew S. Fullerton, Jun Xu, CRC Press, 2016, xvi + 172 pages, £49.99/$79.95, hardcover ISBN: 978-1-4665-6973-7. International Statistical Review = Revue Internationale de Statistique, 84(3):550–551, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2017:BRA

[Mai17] John H. Maindonald. Book review: Algorithms for Data Science, Brian Steele John Chandler and Reddy Swarna Springer, 2016, xxiii + 430 pages, £49.99/$66.99, hardcover ISBN: 978-3-319-45795-6. International Statistical Review = Revue Internationale de Statistique, 85(2):373– 374, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2018:BRR

[Mai18] John H. Maindonald. Book review: Reliability Assessments: Concepts, Models, and Case Studies, Franklin Richard Nash, Chapman and Hall/CRC, 2016, 784 pages, £65.60, hardcover ISBN: 978-1-4987-1917-9. International Statistical Review = Revue Internationale de Statistique, 86(3):602–603, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2019:BRD

[Mai19] John H. Maindonald. Book review: Design of Experiments for Generalized Linear Models, Kenneth Russell, CRC Press, 2019, xiv + 225 pages, $154.40, hardcover, ISBN: 978-1-4987-1917-9. International Statistical Review = Revue Internationale de Statistique, 87(2):447–448, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2021:BRR

[Mai21] John H. Maindonald. Book review: Random Forests with R, Robin Genuer and Jean-Michel Poggi Springer, 2019, x + 99 pp, £49.99, paperback. ISBN: 978-3-030-56484-1. International Statistical Review = Revue Internationale de Statistique, 89(2):422–423, August 2021. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2022:BRR

[Mai22] John H. Maindonald. Book review: Replication and Evidence Factors in Observational Studies, Paul R. Rosenbaum, Chapman and Hall/CRC, 2021, xviii + 254 pages, $120, hardback, ISBN: 978-036748-388-3. International Statistical Review = Revue Internationale de Statistique, 90(1):187–188, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Makridakis:1976:STS

[Mak76] Spyros Makridakis. A survey of time series. International Statistical Review = Revue Internationale de Statistique, 44(1):29–70, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402964.

Makridakis:1978:TSA

[Mak78] Spyros Makridakis. Time-series analysis and forecasting: An update and evaluation. International Statistical Review = Revue Internationale de Statistique, 46(3):255– 278, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402374.

Malinvaud:1975:JD

[Mal75] Edmond Malinvaud. Jacques Dumontier, 1914–1975. International Statistical Review = Revue Internationale de Statistique, 43(3):373–374, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403133.

Malinvaud:1997:EET

[Mal97] Edmond Malinvaud. Effet des Évolutions techniques et des changements de la spécialisation internationale sur les marchés du travail et les systèmes productifs: Comment les statisticiens pourraient-ils relever les défis?. (French) [Effect of technical developments and changes in international specialization on labor markets and productive systems: How could statisticians address challenges?]. International Statistical Review = Revue Internationale de Statistique, 65(1):97–109, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403435.

Malaguerra:2005:OSG

[Mal05] Carlo Malaguerra. Official statistics, globalization and world democracy — a challenge? International Statistical Review = Revue Internationale de Statistique, 73(2): 187–190, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472653.

Mandel:1972:BRB

[Man72] J. Mandel. Book review: Statistical Models and Their Experimental Application by P. Ottestad. International Statistical Review = Revue Internationale de Statistique, 40(1):104–105, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402718.

Mandl:1975:BRB

[Man75] Petr Mandl. Book review: Optimization under Constraints by Peter Whittle. International Statistical Review = Revue Internationale de Statistique, 43(3):369– 370, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403127.

Manly:2007:SBR

[Man07] Bryan F. J. Manly. Short book review: Randomization Tests, 4th Edition, by Eugene S. Edgington, Patrick Onghena. International Statistical Review = Revue Internationale de Statistique, 75(2):269, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508484.

Mannino:2008:SBR

[Man08] Frank Mannino. Short book review: Introduction to Bioinformatics by Anna Tramontano. International Statistical Review = Revue Internationale de Statistique, 76(2):314– 315, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919634.

Maritz:1973:BRB

[Mar73] J. S. Maritz. Book review: A First Course in Statistics by F. N. David. International Statistical Review = Revue Internationale de Statistique, 41(3):397, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402633.

Martin:1974:SLC

[Mar74] Margaret E. Martin. Statistical legislation and confidentiality issues. International Statistical Review = Revue Internationale de Statistique, 42(3):265–281, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402984.

Martin:1975:BRB

[Mar75] Léopold J. Martin. Book review: Distributions Statistiques et Lois de Probabilité by Bruno Leclerc. International Statistical Review = Revue Internationale de Statistique, 43(2):247–248, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402920.

Martin:1996:ORF

[Mar96] Margaret Martin. Opening remarks: Fifth Morris Hansen Lecture. International Statistical Review = Revue Internationale de Statistique, 64(2):163–164, August 1996. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403737.

Marcotte:2007:SBR

[Mar07a] Denis Marcotte. Short book review: Analyse statistique des données spatiale edited by Jean-Jacques Droesbeke, Michel Lejeune, Gilbert Saporta. International Statistical Review = Revue Internationale de Statistique, 75(2): 264, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508479.

Marriott:2007:SBR

[Mar07b] Paul Marriott. Short book review: Deblurring Images: Matrices, Spectra and Filtering by Per Christian Hansen, James G. Nagy, Dianne P. O’Leary. International Statistical Review = Revue Internationale de Statistique, 75(2): 253–254, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508469.

Maruotti:2011:MHM

[Mar11] Antonello Maruotti. Mixed hidden Markov models for longitudinal data: An overview. International Statistical Review = Revue Internationale de Statistique, 79(3):427– 454, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305059.

Martinez:2021:BRO

[Mar21] Wendy L. Martinez. Book review: Official Statistics 4.0: Verified Facts for People in the 21st Century, Walter J. Radermacher Springer International Publishing, 2020, xii + 158 pages, $79.99, hardcover ISBN: 978-3-030-31491-0. International Statistical Review = Revue Internationale de Statistique, 89(3):660–661, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Masuyama:1973:ACA

[Mas73] M. Masuyama. Adequacy of currently available techniques and areas of future research in biostatistics: Design of experiments. International Statistical Review = Revue Internationale de Statistique, 41(3):333–342, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402621.

Matusita:1973:BRB

[Mat73] K. Matusita. Book review: On Regular Variation and Its Application to the Weak Convergence of Sample Extremes by L. de Haan. International Statistical Review = Revue Internationale de Statistique, 41(2):293, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402855.

Matthews:1988:RDC

[Mat88] J. N. S. Matthews. Recent developments in crossover designs. International Statistical Review = Revue Internationale de Statistique, 56(2):117–127, August 1988. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403636.

Mathew:2007:SBR

[Mat07] Thomas Mathew. Short book review: Bioequivalence Studies in Drug Development: Methods and Applications by Dieter Hauschke, Volker Steinijans, Iris Pigeot. International Statistical Review = Revue Internationale de Statistique, 75(2):272–273, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508487.

Murray:2019:WFS

[MB19] Lori Murray and David Bellhouse. W. F. Sheppard’s smoothing method: a precursor to local polynomial regression. International Statistical Review = Revue Internationale de Statistique, 87(3):604–612, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

McVittie:2023:SMD

[MBW^{+}23] James H. McVittie, Ana F. Best, David B. Wolfson, David A. Stephens, Julian Wolfson, David L. Buckeridge, and Shahinaz M. Gadalla. Survival modelling for data from combined cohorts: Opening the door to meta survival analyses and survival analysis using electronic health records. International Statistical Review = Revue Internationale de Statistique, 91(1):72–87, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

McCullagh:1985:ADP

[McC85] Peter McCullagh. On the asymptotic distribution of Pearson’s statistic in linear exponential-family models. International Statistical Review = Revue Internationale de Statistique, 53(1):61–67, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402880.

McConville:2022:BRC

[McC22] Kelly McConville. Book review: Communicating with Data: The Art of Writing for Data Science, Deborah Nolan and Sara Stoudt, Oxford University Press, 2021, vii + 331 pages, $45.95, paperback, ISBN: 978-0-1988-6275-8. International Statistical Review = Revue Internationale de Statistique, 90(1):189–191, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

McElroy:2017:BRT

[McE17] Tucker S. McElroy. Book review: Time Series Econometrics, Klaus Neusser, Springer International Publishing, 2016, xxiv + 409 pages, £99.00, hardcover ISBN: 978-3-319-32861-4. International Statistical Review = Revue Internationale de Statistique, 85(1):181–183, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Mecatti:2012:SGD

[MCF12] Fulvia Mecatti, Franca Crippa, and Patrizia Farina. A special gen(d)re of statistics: roots, development and methodological prospects of gender statistics. International Statistical Review = Revue Internationale de Statistique, 80(3):452–467, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Munoz:2013:IEM

[MCH13] M. Pilar Muñoz, Cristina Corchero, and F.-Javier Heredia. Improving electricity market price forecasting with factor models for the optimal generation bid. International Statistical Review = Revue Internationale de Statistique, 81 (2):289–306, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

McLeod:1993:PMA

[McL93] A. I. McLeod. Parsimony, model adequacy and periodic correlation in time series forecasting. International Statistical Review = Revue Internationale de Statistique, 61 (3):387–393, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403750.

Moriguti:1992:RSR

[MDG^{+}92] Sigeiti Moriguti, Peter J. Diggle, John C. Gower, Katherine K. Wallman, and Wang Shou Ren. The role of statisticians. Report of the Ad Hoc Committee. International Statistical Review = Revue Internationale de Statistique, 60(3):227–246, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403677.

Marsh:1994:SDV

[MDS94] Catherine Marsh, Angela Dale, and Christopher Skinner. Safe data versus Safe settings: Access to microdata from the British census. International Statistical Review = Revue Internationale de Statistique, 62(1):35–53, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403544.

Maslovskaya:2019:WCR

[MDS^{+}19] Olga Maslovskaya, Gabriele B. Durrant, Peter W. F. Smith, Tim Hanson, and Ana Villar. What are the characteristics of respondents using different devices in mixed-device online surveys? Evidence from six UK surveys. International Statistical Review = Revue Internationale de Statistique, 87(2):326–346, August 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Madsen:1989:SRM

[ME89] Henrik O. Madsen and Thore Egeland. Structural reliability: Models and applications. International Statistical Review = Revue Internationale de Statistique, 57(3):185– 203, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403793.

Meeker:1993:RRR

[ME93] William Q. Meeker and Luis A. Escobar. A review of recent research and current issues in accelerated testing. International Statistical Review = Revue Internationale de Statistique, 61(1):147–168, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403600.

Mehran:1975:BRB

[Meh75] Farhad Mehran. Book review: The Generalized Jackknife Statistic by H. L. Gray; W. R. Schucany. International Statistical Review = Revue Internationale de Statistique, 43(2):245–246, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402917.

Mehran:2015:BRS

[Meh15] Farhad Mehran. Book review: Sampling Elusive Populations: Applications to Studies of Child Labour. International Statistical Review = Revue Internationale de Statistique, 83(2):329–331, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Menges:1973:BRB

[Men73] G. Menges. Book review: Clustering and Aggregation in Economics by Walter D. Fisher. International Statistical Review = Revue Internationale de Statistique, 41(3):398– 399, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402635.

Menges:1976:RW

[Men76] Günter Menges. Rolf Wagenführ, 1905–1975. International Statistical Review = Revue Internationale de Statistique, 44(3):386–387, December 1976. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402368.

Menardi:2016:RMC

[Men16a] Giovanna Menardi. A review on modal clustering. International Statistical Review = Revue Internationale de Statistique, 84(3):413–433, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Meng:2016:DES

[Men16b] Xiao-Li Meng. Discussions on enhancing statistics education by including qualitative research: Discussion: The Qq dynamic for deeper learning and research. International Statistical Review = Revue Internationale de Statistique, 84(2):181–189, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Meng:2016:DPI

[Men16c] Xiao-Li Meng. Discussions on posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Discussion: Should a working model actually work? International Statistical Review = Revue Internationale de Statistique, 84(3):362–367, December 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Merikoski:2007:SBR

[Mer07] Jorma Kaarlo Merikoski. Short book review: Mathematics and Common Sense: A Case of Creative Tension by Philip J. Davis. International Statistical Review = Revue Internationale de Statistique, 75(2):267–268, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508482.

Merikoski:2009:SBRd

[Mer09a] Jorma Kaarlo Merikoski. Short book review: Biscuits of Number Theory edited by Arthur T. Benjamin, Ezra Brown. International Statistical Review = Revue Internationale de Statistique, 77(3):468–469, December 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919773.

Merikoski:2009:SBRc

[Mer09b] Jorma Kaarlo Merikoski. Short book review: Creative Mathematics by H. S. Wall. International Statistical Review = Revue Internationale de Statistique, 77(3):464– 465, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919770.

Merikoski:2009:SBRa

[Mer09c] Jorma Kaarlo Merikoski. Short book review: Hexaflexagons, Probability Paradoxes, and the Tower of Hanoi: Martin Gardner’s First Book of Mathematical Puzzles and Games by Martin Gardner. International Statistical Review = Revue Internationale de Statistique, 77 (1):152, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919701.

Merikoski:2009:SBRb

[Mer09d] Jorma Kaarlo Merikoski. Short book review: Origami, Eleusis, and the Soma Cube: Martin Gardner’s Mathematical Diversions by Martin Gardner. International Statistical Review = Revue Internationale de Statistique, 77 (1):152–153, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919702.

Merikoski:2012:SBRa

[Mer12a] Jorma Kaarlo Merikoski. Short book review: A Mathematician Comes of Age by Steven G. Krantz. International Statistical Review = Revue Internationale de Statistique, 80(2):323–324, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2012:SBRb

[Mer12b] Jorma Kaarlo Merikoski. Short book review: Calculus and Its Origins by David Perkins. International Statistical Review = Revue Internationale de Statistique, 80(3):491– 492, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2014:BRE

[Mer14] Jorma K. Merikoski. Book review: Examples and Problems in Mathematical Statistics by Shelemyahu Zacks. International Statistical Review = Revue Internationale de Statistique, 82(3):483–485, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2015:BRJ

[Mer15] Jorma K. Merikoski. Book review: John Napier: Life, Logarithms, and Legacy. International Statistical Review = Revue Internationale de Statistique, 83(2):327–329, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2017:BRP

[Mer17] Jorma K. Merikoski. Book review: A Panorama of Statistics: Perspectives, Puzzles and Paradoxes in Statistics, Eric Sowey and Peter Petocz John Wiley & Sons, 2017, xii + 313 pages, softcover ISBN: 978-1-119-07582-0. International Statistical Review = Revue Internationale de Statistique, 85(2):375–376, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2018:BRP

[Mer18] Jorma K. Merikoski. Book review: Probability: The Classical Limit Theorems, Henry McKean, Cambridge University Press, 2014, xxi + 464 pages, £62.00, hardcover ISBN: 978-1-107-05321-2. International Statistical Review = Revue Internationale de Statistique, 86(3):598–599, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2020:BRI

[Mer20a] Jorma K. Merikoski. Book review: Introduction to Probability: Models and Applications, N. Balakrishnan, Markos V. Koutras and Konstadinos G. Politis, John Wiley and Sons, 2020, xiii + 608 pages, $140.00, hardcover ISBN: 978-1-118-12334-8. International Statistical Review = Revue Internationale de Statistique, 88(1):255–256, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2020:BRR

[Mer20b] Jorma K. Merikoski. Book review: Random Circulant Matrices, Arup Bose and Koushik Saha, CRC Press, 2019, xix + 192 pages, $174.95, hardcover ISBN: 978-1-138-35109-7. International Statistical Review = Revue Internationale de Statistique, 88(1):252–254, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merikoski:2020:BRB

[Mer20c] Jorma K. Merikoski. Book review: Random Circulant Matrices, Arup Bose and Koushik Saha, CRC Press, 2019, xix + 192 pages, $174.95, hardcover ISBN: 978-1-138-35109-7. International Statistical Review = Revue Internationale de Statistique, 88(1):520, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Merkouris:2024:MEU

[Mer24] Takis Merkouris. On the most effective use of continuous auxiliary variables in regression estimation in survey sampling. International Statistical Review = Revue Internationale de Statistique, 92(2):263–283, August 2024. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Montero:2010:EAS

[MFAG10] José-María Montero, Gema Fernández-Avilés, and María-Carmen García. Estimation of asymmetric stochastic volatility models: Application to daily average prices of energy products. International Statistical Review = Revue Internationale de Statistique, 78(3):330–347, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919857.

Martin-Guzman:2005:DPM

[MG05] Pilar Martin-Guzman. Discussant: Pilar Martin-Guzman. International Statistical Review = Revue Internationale de Statistique, 73(2):197–199, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Munoz-Garcia:1990:OFA

[MGMRPA90] J. Muñoz-Garcia, J. L. Moreno-Rebollo, and A. Pascual-Acosta. Outliers: A formal approach. International Statistical Review = Revue Internationale de Statistique, 58 (3):215–226, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403805.

Michalup:1977:EP

[Mic77] Erich Michalup. Ernst Peltzer, 1901–1975. International Statistical Review = Revue Internationale de Statistique, 45(1):109–110, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403031.

Micheas:2019:CPP

[Mic19] Athanasios C. Micheas. Cox point processes: Why one realisation is not enough. International Statistical Review = Revue Internationale de Statistique, 87(2):306–325, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Missiakoulis:2008:AED

[Mis08] Spyros Missiakoulis. Aristotle and earthquake data: A historical note. International Statistical Review = Revue Internationale de Statistique, 76(1):130–133, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919587.

Missiakoulis:2010:CKA

[Mis10] Spyros Missiakoulis. Cecrops, King of Athens: the first (?) recorded population census in history. International Statistical Review = Revue Internationale de Statistique, 78(3):413–418, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919861.

Mitra:1979:AKG

[Mit79] S. K. Mitra. Anil Kumar Gayen, 1919–1978. International Statistical Review = Revue Internationale de Statistique, 47(1):95, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403227.

Mitchell:1988:SMU

[Mit88] Ann F. S. Mitchell. Statistical manifolds of univariate elliptic distributions. International Statistical Review = Revue Internationale de Statistique, 56(1):1–16, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403358.

Mizon:1983:DPD

[Miz83] Grayham E. Mizon. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):155–157, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402743.

Mairesse:1977:BRB

[MM77] J. Mairesse and P. Mazodier. Book review: Exercises in Mathematical Economics and Econometrics by J. E. Spencer; R. C. Geary. International Statistical Review = Revue Internationale de Statistique, 45(1):103–105, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403025.

Mariani:1996:DAP

[MM96] L. Mariani and E. Marubini. Design and analysis of phase II cancer trials: A review of statistical methods and guidelines for medical researchers. International Statistical Review = Revue Internationale de Statistique, 64 (1):61–88, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403424.

Mateu:2001:LIG

[MM01] Jorge Mateu and Francisco Montes. Likelihood inference for Gibbs processes in the analysis of spatial point patterns. International Statistical Review = Revue Internationale de Statistique, 69(1):81–104, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403531.

Mecklin:2004:ABT

[MM04] Christopher J. Mecklin and Daniel J. Mundfrom. An appraisal and bibliography of tests for multivariate normality. International Statistical Review = Revue Internationale de Statistique, 72(1):123–138, April 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403847.

McCrink:2013:AJM

[MMC13] Lisa M. McCrink, Adele H. Marshall, and Karen J. Cairns. Advances in joint modelling: a review of recent developments with application to the survival of end stage renal disease patients. International Statistical Review = Revue Internationale de Statistique, 81(2):249–269, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Marriott:1998:BCA

[MN98] John Marriott and Paul Newbold. Bayesian comparison of ARIMA and stationary ARMA models. International Statistical Review = Revue Internationale de Statistique, 66(3):323–336, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403520.

McWhorter:1977:EEP

[MNS77] Archer McWhorter, Jr., Gorti V. L. Narasimham, and Richard R. Simonds. An empirical examination of the predictive performance of an econometric model with random coefficients. International Statistical Review = Revue Internationale de Statistique, 45(3):243–255, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402538.

Modarres:2008:TBE

[Mod08] Reza Modarres. Tests of bivariate exchangeability. International Statistical Review = Revue Internationale de Statistique, 76(2):203–213, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919612.

Modarres:2020:GCH

[Mod20] Reza Modarres. Graphical comparison of high-dimensional distributions. International Statistical Review = Revue Internationale de Statistique, 88(3):698–714, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Montanari:1987:PSE

[Mon87] G. E. Montanari. Post-sampling efficient QR-prediction in large-sample surveys. International Statistical Review = Revue Internationale de Statistique, 55(2):191–202, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403195.

Monticino:2001:HCR

[Mon01] Michael Monticino. How to construct a random probability measure. International Statistical Review = Revue Internationale de Statistique, 69(1):153–167, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403534.

Moore:1975:BRB

[Moo75] P. G. Moore. Book review: A Manual of Sampling Techniques by R. K. Som. International Statistical Review = Revue Internationale de Statistique, 43(2):250, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402926.

Moore:1978:MTB

[Moo78] P. G. Moore. The mythical threat of Bayesianism. A reply to “Bayesianism; a threat to the statistical profession?” (Internat. Statist. Rev. 45 (1977), 111–115 by H. C. Hamaker). International Statistical Review = Revue Internationale de Statistique, 46(1):67–73, April 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402509.

Moore:1997:NPNa

[Moo97a] David S. Moore. New pedagogy and new content: The case of statistics. International Statistical Review = Revue Internationale de Statistique, 65(2):123–137, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403333.

Moore:1997:NPNb

[Moo97b] David S. Moore. [new pedagogy and new content: The case of statistics]: Response. International Statistical Review = Revue Internationale de Statistique, 65(2):162– 165, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403341.

Moore:1999:DWS

[Moo99] David S. Moore. Discussion: What shall we teach beginners? International Statistical Review = Revue Internationale de Statistique, 67(3):250–252, December 1999. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403701.

Moore:2005:QRF

[Moo05] David S. Moore. Quality and relevance in the first statistics course. International Statistical Review = Revue Internationale de Statistique, 73(2):205–206, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472655.

Moran:1973:BRB

[Mor73] P. A. P. Moran. Book review: Rank Correlation Methods by M. G. Kendall. International Statistical Review = Revue Internationale de Statistique, 41(3):399–400, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402637.

Moran:1974:HFS

[Mor74] P. A. P. Moran. How to find out in statistical and probability theory. International Statistical Review = Revue Internationale de Statistique, 42(3):299–303, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402987.

Moran:1980:BRB

[Mor80a] P. A. P. Moran. Book review: Handbook of Hypergeometric Integrals. Theory, Applications, Tables, Computer Programs by Harold Exton. International Statistical Review = Revue Internationale de Statistique, 48(2):235, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403169.

Morgenstern:1980:HR

[Mor80b] Dietrich Morgenstern. Hans Richter, 1912–1978. International Statistical Review = Revue Internationale de Statistique, 48(1):137, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402443.

Morettin:1984:LAA

[Mor84] Pedro A. Morettin. The Levinson algorithm and its applications in time series analysis. International Statistical Review = Revue Internationale de Statistique, 52(1): 83–92, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403247.

Maruotti:2021:IHM

[MP21] Antonello Maruotti and Antonio Punzo. Initialization of hidden Markov and semi-Markov models: a critical evaluation of several strategies. International Statistical Review = Revue Internationale de Statistique, 89(3):447–480, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Marella:2023:ANI

[MP23] Daniela Marella and Danny Pfeffermann. Accounting for non-ignorable sampling and non-response in statistical matching. International Statistical Review = Revue Internationale de Statistique, 91(2):269–293, August 2023. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Manninen:2003:I

[MPJS03] Asta Manninen, Lea Parjo, and Heli Jeskanen-Sundström. Introduction. International Statistical Review = Revue Internationale de Statistique, 71(1):1–3, April 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403869.

McMurry:2012:SIP

[MPR12] Timothy L. McMurry, Dimitris N. Politis, and Joseph P. Romano. Subsampling inference with K populations and a non-standard Behrens–Fisher problem. International Statistical Review = Revue Internationale de Statistique, 80 (1):149–175, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257171.

Mantel:1980:DST

[MR80] Nathan Mantel and Alton J. Rahe. Differentiated sign tests. International Statistical Review = Revue Internationale de Statistique, 48(1):19–28, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402401.

Meng:2003:DES

[MR03] Xiao-Li Meng and Martin Romero. Discussion: Efficiency and self-efficiency with multiple imputation inference. International Statistical Review = Revue Internationale de Statistique, 71(3):607–618, December 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403832.

Moller:2010:MPC

[MR10] Jesper Møller and Ege Rubak. A model for positively correlated count variables. International Statistical Review = Revue Internationale de Statistique, 78(1):65–80, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919795.

McElroy:2022:RSA

[MR22] Tucker McElroy and Anindya Roy. A review of seasonal adjustment diagnostics. International Statistical Review = Revue Internationale de Statistique, 90(2):259–284, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Madan:1987:SEU

[MS87] D. B. Madan and E. Seneta. Simulation of estimates using the empirical characteristic function. International Statistical Review = Revue Internationale de Statistique, 55(2): 153–161, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403191.

Moretti:2023:IPR

[MS23] Angelo Moretti and Natalie Shlomo. Improving probabilistic record linkage using statistical prediction models. International Statistical Review = Revue Internationale de Statistique, 91(3):368–394, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Molina:2001:MOC

[MSS01] E. A. Molina, T. M. F. Smith, and R. A. Sugden. Modelling overdispersion for complex survey data. International Statistical Review = Revue Internationale de Statistique, 69(3):373–384, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403451.

Moretti:2020:MSA

[MSS20] Angelo Moretti, Natalie Shlomo, and Joseph W. Sakshaug. Multivariate small area estimation of multidimensional latent economic well-being indicators. International Statistical Review = Revue Internationale de Statistique, 88(1):1–28, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

McWhorter:1976:SAV

[MSW76] A. McWhorter, W. Allen Spivey, and W. J. Wrobleski. A sensitivity analysis of varying parameter econometric models. International Statistical Review = Revue Internationale de Statistique, 44(2):265–282, August 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403287.

Muilwijk:1973:BRB

[Mui73] J. Muilwijk. Book review: Theoretische Grondlagen des Österreichischen Mikrozensus by H. Lutz. International Statistical Review = Revue Internationale de Statistique, 41(2):293, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402857.

Mukhopadhyay:2014:BRC

[Muk14a] Purna Mukhopadhyay. Book review: Clinical Trials with Missing Data: A Guide for Practitioners by Michael O’Kelly and Bohdana Ratitch. International Statistical Review = Revue Internationale de Statistique, 82(3):480– 481, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Mukhopadhyay:2014:BRL

[Muk14b] Purna Mukhopadhyay. Book review: Learning Regression Analysis by Simulation by Kunio Takezawa. International Statistical Review = Revue Internationale de Statistique, 82(2):325, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Mukhopadhyay:2015:BRM

[Muk15] Purna Mukhopadhyay. Book review: Multiple Imputation of Missing Data Using SAS. International Statistical Review = Revue Internationale de Statistique, 83(2):326– 327, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Muller:1976:CUC

[Mul76] Mervin E. Muller. Challenges in using computers in statistical applications in developing countries. International Statistical Review = Revue Internationale de Statistique, 44(2):249–263, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403286.

Munzner:1976:KS

[Mün76] Hans Münzner. Kurt Stange, 1907–1974. International Statistical Review = Revue Internationale de Statistique, 44(2):310, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403307.

Murthy:1974:EMS

[Mur74] M. N. Murthy. Evaluation of multi-subject sample survey systems. International Statistical Review = Revue Internationale de Statistique, 42(2):175–191, August 1974. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403078.

Murtagh:1994:NNR

[Mur94] F. Murtagh. Neural networks and related ’massively parallel’ methods for statistics: A short overview. International Statistical Review = Revue Internationale de Statistique, 62(3):275–288, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403762.

Mustert:1975:BRB

[Mus75] G. R. Mustert. Book review: Data Quality Control and Editing by J. I. Naus. International Statistical Review = Revue Internationale de Statistique, 43(2):248, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402921.

Moller:2022:SWC

[MV22] Jesper Møller and Ninna Vihrs. Should we condition on the number of points when modelling spatial point patterns? International Statistical Review = Revue Internationale de Statistique, 90(3):551–562, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Muller:1976:SFD

[MW76] Mervin E. Muller and Graham N. Wilkinson. A standardized form for describing programs, packages and systems for statistical applications. International Statistical Review = Revue Internationale de Statistique, 44(3):349– 353, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402351.

McBratney:1981:DRF

[MW81] A. B. McBratney and R. Webster. Detection of ridge and furrow pattern by spectral analysis of crop yield. International Statistical Review = Revue Internationale de Statistique, 49(1):45–52, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403036.

Montgomery:2008:OSS

[MW08] Douglas C. Montgomery and William H. Woodall. An overview of six sigma. International Statistical Review = Revue Internationale de Statistique, 76(3):329–346, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919650.

Muller:2010:MSC

[MW10] Samuel Müller and Alan H. Welsh. On model selection curves. International Statistical Review = Revue Internationale de Statistique, 78(2):240–256, August 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919835.

Madigan:1995:BGM

[MYA95] David Madigan, Jeremy York, and Denis Allard. Bayesian graphical models for discrete data. International Statistical Review = Revue Internationale de Statistique, 63(2): 215–232, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403615.

Ma:2013:RDR

[MZ13] Yanyuan Ma and Liping Zhu. A review on dimension reduction. International Statistical Review = Revue Internationale de Statistique, 81(1):134–150, April 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nayak:2016:IPR

[NA16] Tapan K. Nayak and Samson A. Adeshiyan. On invariant post-randomization for statistical disclosure control. International Statistical Review = Revue Internationale de Statistique, 84(1):26–42, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Naddeo:1974:BRB

[Nad74] A. Naddeo. Book review: Statistical Inference in Graphs by Frank Ove. International Statistical Review = Revue Internationale de Statistique, 42(1):109, April 1974. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402692.

Nair:1973:SRT

[Nai73] K. R. Nair. Statistical research and training / formation et recherche statistique. International Statistical Review = Revue Internationale de Statistique, 41(1):131–134, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402794.

Nair:1993:BRE

[Nai93] Vijayan N. Nair. Bounds for reliability estimation under dependent censoring. International Statistical Review = Revue Internationale de Statistique, 61(1):169–182, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403601.

Narula:1976:ECS

[NAL76] Subhash C. Narula, Patrick F. Abrami, and Kenneth J. Levy. An empirical comparison of several methods for analysing disproportionate subclass data in a two-way classification ANOVA. International Statistical Review = Revue Internationale de Statistique, 44(3):341–347, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402349.

Nandy:2010:SBRb

[Nan10a] Karabi Nandy. Short book review: Bayesian Analysis for the Social Sciences by Simon Jackman. International Statistical Review = Revue Internationale de Statistique, 78 (3):464, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919880.

Nandy:2010:SBRa

[Nan10b] Rajesh Ranjan Nandy. Short book review: Spatial Statistics and Modeling by Carlo Gaetan, Xavier Guyon. International Statistical Review = Revue Internationale de Statistique, 78(3):459, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919875.

Narula:1974:SWI

[Nar74] Subhash C. Narula. Systematic ways to identify research problems in statistics. International Statistical Review = Revue Internationale de Statistique, 42(2):205–209, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403080.

Narula:1979:OPR

[Nar79] Sabhash C. Narula. Orthogonal polynomial regression. International Statistical Review = Revue Internationale de Statistique, 47(1):31–36, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403204.

Narasimhan:2020:IBE

[Nar20a] Balasubramanian Narasimhan. An interview with Bradley Efron. International Statistical Review = Revue Internationale de Statistique, 88(S1):S2–S27, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Narisetty:2020:D

[Nar20b] Naveen Naidu Narisetty. Discussion. International Statistical Review = Revue Internationale de Statistique, 88(2): 330–334, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Naus:1979:IBC

[Nau79] Joseph I. Naus. An indexed bibliography of clusters, clumps and coincidences. International Statistical Review = Revue Internationale de Statistique, 47(1):47–78, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403208.

Nayak:2009:IKP

[Nay09] Tapan K. Nayak. Impact of Karl Pearson’s work on statistical developments in India. International Statistical Review = Revue Internationale de Statistique, 77(1): 72–80, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919692.

Neath:2010:BEP

[NC10] Andrew A. Neath and Joseph E. Cavanaugh. Bayesian estimation of prediction error and variable selection in linear regression. International Statistical Review = Revue Internationale de Statistique, 78(2):257–270, August 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919836.

Nanda:2021:SEG

[NC21] Asok K. Nanda and Shovan Chowdhury. Shannon’s entropy and its generalisations towards statistical inference in last seven decades. International Statistical Review = Revue Internationale de Statistique, 89(1):167–185, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Neeman:2009:SBR

[Nee09] Teresa Neeman. Short book review: Clinical Prediction Models: A Practical Approach to Development, Validation, and Updating by Ewout W. Steyerberg. International Statistical Review = Revue Internationale de Statistique, 77(2):320–321, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919750.

Neeman:2015:BRS

[Nee15] Teresa Neeman. Book review: Sample Size Calculations for Clustered and Longitudinal Outcomes in Clinical Research. International Statistical Review = Revue Internationale de Statistique, 83(3):520–522, December 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nelder:1972:STR

[Nel72a] J. Nelder. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 40(3):384–388, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402476.

Nelder:1972:BRB

[Nel72b] J. A. Nelder. Book review: Analysis and Design of Certain Quantitative Multiresponse Experiments by S. N. Roy; R. Gnanadesikan; J. N. Srivastava. International Statistical Review = Revue Internationale de Statistique, 40(2): 256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402776.

Nelder:1974:UGE

[Nel74] J. A. Nelder. A User’s guide to the evaluation of statistical packages and systems. International Statistical Review = Revue Internationale de Statistique, 42(3):291– 298, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402986.

Nelder:1975:STR

[Nel75] J. A. Nelder. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 43(2):227– 232, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402904.

Nelder:1980:BRB

[Nel80] J. A. Nelder. Book review: Statistical Methods for Digital Computers by K. Enslein; A. Ralston; H. S. Wilf. International Statistical Review = Revue Internationale de Statistique, 48(1):120–121, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402419.

Nemenyi:1973:ARG

[Nem73] Peter Nemenyi. Agility ratings and ground meat: Introducing Mann–Whitney differences. International Statistical Review = Revue Internationale de Statistique, 41(2): 240–244, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402840.

Neusser:2016:TSE

[Neu16] Klaus Neusser. Time series econometrics. Springer Science+Business Media, New York, NY, 2016. ISBN 3-319-32861-1, 3-319-32862-X (e-book). ???? pp. LCCN HB141 .N48 2016.

Newbold:1981:SRD

[New81] P. Newbold. Some recent developments in time series analysis, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 49(1):53–66, April 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403037.

Newbold:1983:DPD

[New83] Paul Newbold. Discussion of paper by D. F. Hendry and J.-F. Richard. International Statistical Review = Revue Internationale de Statistique, 51(2):157–160, August 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402744.

Newbold:1984:SRD

[New84] P. Newbold. Some recent developments in time series analysis: II, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 52(2): 183–192, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403101.

Newbold:1988:SRD

[New88] P. Newbold. Some recent developments in time series analysis. III, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 56(1): 17–29, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403359.

Neyman:1975:ACE

[Ney75] Jerzy Neyman. Assessing the chain: energy crisis, pollution and health. International Statistical Review = Revue Internationale de Statistique, 43(3):253–267, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403109.

Ng:2021:BRS

[Ng21] Hon Keung Tony Ng. Book review: Stochastic Models in Reliability Engineering, Lirong Cui, Ilia Frenkel and Anatoly Lisniansku CRC Press, 2020, 482 pages, $190.00, hardcover ISBN: 978-0-367-34585-3. International Statistical Review = Revue Internationale de Statistique, 89(2): 426–429, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nourse:1978:STR

[NGC^{+}78] E. Shepley Nourse, Bernard G. Greenberg, Gertrude M. Cox, David D. Mason, James E. Grizzle, Norman L. Johnson, Lyle V. Jones, John Monroe, and Gordon D. Simons, Jr. Statistical training and research: The University of North Carolina System. International Statistical Review = Revue Internationale de Statistique, 46(2):171– 207, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402812.

Nicholls:2001:FDT

[Nic01] Des F. Nicholls. Future directions for the teaching and learning of statistics at the tertiary level. International Statistical Review = Revue Internationale de Statistique, 69(1):11–15, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403525.

Nielsen:2003:PIMa

[Nie03a] Søren Feodor Nielsen. Proper and improper multiple imputation. International Statistical Review = Revue Internationale de Statistique, 71(3):593–607, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403831.

Nielsen:2003:PIMb

[Nie03b] Søren Feodor Nielsen. [proper and improper multiple imputation]: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 71(3):625–627, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403834.

Nieto:2004:CIS

[Nie04] Fabio H. Nieto. A coincident index for the state of the economy. International Statistical Review = Revue Internationale de Statistique, 72(3):355–376, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472628.

Nadarajah:2006:PPM

[NK06] Saralees Nadarajah and Samuel Kotz. Percentage points of the multivariate t distribution. International Statistical Review = Revue Internationale de Statistique, 74(1): 15–30, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472688.

Nadarajah:2008:PIM

[NK08] Saralees Nadarajah and Samuel Kotz. Probability integrals of the multivariate t distribution. International Statistical Review = Revue Internationale de Statistique, 76 (1):58–88, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919584.

Niebuhr:2014:AAI

[NK14] Tobias Niebuhr and Jens-Peter Kreiss. Asymptotics for autocovariances and integrated periodograms for linear processes observed at lower frequencies. International Statistical Review = Revue Internationale de Statistique, 82 (1):123–140, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Neuhaus:1991:CCS

[NKH91] J. M. Neuhaus, J. D. Kalbfleisch, and W. W. Hauck. A comparison of cluster-specific and population-averaged approaches for analyzing correlated binary data. International Statistical Review = Revue Internationale de Statistique, 59(1):25–35, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403572.

Narula:1975:PDP

[NL75] Subhash C. Narula and Kenneth J. Levy. Probability density plots of the noncentral χ^{2} and non-central F distributions. International Statistical Review = Revue Internationale de Statistique, 43(1):79–82, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402661.

Nolan:2007:DID

[NL07] Deborah Nolan and Duncan Temple Lang. Dynamic, interactive documents for teaching statistical practice. International Statistical Review = Revue Internationale de Statistique, 75(3):295–321, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509872.

Nelder:2008:LEW

[NLSS08] John A. Nelder, Viviana B. Lencina, Julio M. Singer, and Edward J. Stanek III. Letter to the Editors: What is the mixed-models controversy? (with response). International Statistical Review = Revue Internationale de Statistique, 76(1):134–135, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919588.

Nagaraja:2020:CDF

[NN20a] Chaitra H. Nagaraja and Haikady N. Nagaraja. Correction to ‘Distribution-free Approximate Methods for Constructing Confidence Intervals for Quantiles’. International Statistical Review = Revue Internationale de Statistique, 88(2):519, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [NN20b].

Nagaraja:2020:DFA

[NN20b] Chaitra H. Nagaraja and Haikady N. Nagaraja. Distributionfree approximate methods for constructing confidence intervals for quantiles. International Statistical Review = Revue Internationale de Statistique, 88(1):75–100, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See correction [NN20a].

Norden:1972:SML

[Nor72] R. H. Norden. A survey of maximum likelihood estimation. International Statistical Review = Revue Internationale de Statistique, 40(3):329–354, December 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402471.

See also Part II [Nor73].

Norden:1973:SML

[Nor73] R. H. Norden. A survey of maximum likelihood estimation: Part 2. International Statistical Review = Revue Internationale de Statistique, 41(1):39–58, April 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402786.

See also Part I [Nor72].

Norwood:1996:CES

[Nor96] Janet L. Norwood. [characteristics of an effective statistical system]: Discussion. International Statistical Review = Revue Internationale de Statistique, 64(2):191– 193, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403740.

Nordholt:1998:IMS

[Nor98] Eric Schulte Nordholt. Imputation: Methods, simulation experiments and practical examples. International Statistical Review = Revue Internationale de Statistique, 66(2): 157–180, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403488.

Nordhausen:2008:SBR

[Nor08] Klaus Nordhausen. Short book review: Lattice: Multivariate Data Visualization with R by Deepayan Sarkar. International Statistical Review = Revue Internationale de Statistique, 76(3):440, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919660.

Nordhausen:2009:SBRb

[Nor09a] Klaus Nordhausen. Short book review: A First Course in Statistical Programming with R by W. John Braun, Duncan J. Murdoch. International Statistical Review = Revue Internationale de Statistique, 77(2):317, August 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919747.

Nordhausen:2009:SBRa

[Nor09b] Klaus Nordhausen. Short book review: Introductory Statistics with R, Second Edition, by Peter Dalgaard. International Statistical Review = Revue Internationale de Statistique, 77(1):155–156, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919705.

Nordhausen:2009:SBRc

[Nor09c] Klaus Nordhausen. Short book review: The Elements of Statistical Learning: Data Mining, Inference, and Prediction, Second Edition, by Trevor Hastie, Robert Tibshirani, Jerome Friedman. International Statistical Review = Revue Internationale de Statistique, 77(3):482, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919786.

Nordhausen:2013:SBRb

[Nor13a] Klaus Nordhausen. Short book review: Ensemble Methods: Foundations and Algorithms by Zhi-Hua Zhou. International Statistical Review = Revue Internationale de Statistique, 81(3):470, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nordhausen:2013:SBRa

[Nor13b] Klaus Nordhausen. Short book review: Stationary Stochastic Processes: Theory and Applications by Georg Lindgren. International Statistical Review = Revue Internationale de Statistique, 81(3):469–470, December 2013.

CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nordhausen:2014:BRBb

[Nor14a] Klaus Nordhausen. Book review: An Introduction to Statistical Learning — with Applications in R by Gareth James, Daniela Witten, Trevor Hastie & Robert Tibshirani. International Statistical Review = Revue Internationale de Statistique, 82(1):156–157, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nordhausen:2014:BRBa

[Nor14b] Klaus Nordhausen. Book review: Multiple Imputation and its Application by James R. Carpenter, Michael G. Kenward. International Statistical Review = Revue Internationale de Statistique, 82(1):151–152, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nordhausen:2015:BRS

[Nor15] Klaus Nordhausen. Book review: Statistical Analysis of Network Data with R. International Statistical Review = Revue Internationale de Statistique, 83(1):171–172, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Novick:1973:HSA

[Nov73] Melvin R. Novick. High school attainment: An example of a computer-assisted Bayesian approach to data analysis. International Statistical Review = Revue Internationale de Statistique, 41(2):264–271, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402845.

Nielsen:1997:CUS

[NP97] Peter Bøegh Nielsen and Jan Plovsing. Concepts used in statistical business registers in view of globalisation and the information society. International Statistical Review = Revue Internationale de Statistique, 65(3):351– 363, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403376.

Nelder:1974:STR

[NPT^{+}74] J. A. Nelder, G. P. Patil, J. L. Teugels, T. Dalenius, Saw Swee-Hock, and J. S. Siegel. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 42(1):79–87, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402687.

Nedyalkova:2008:SPC

[NPT08] Desislava Nedyalkova, Johan Pea, and Yves Tillé. Sampling procedures for coordinating stratified samples: Methods based on microstrata. International Statistical Review = Revue Internationale de Statistique, 76(3):368– 386, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919652.

Nobre:2021:ESC

[NSP21] Widemberg S. Nobre, Alexandra M. Schmidt, and João B. M. Pereira. On the effects of spatial confounding in hierarchical models. International Statistical Review = Revue Internationale de Statistique, 89(2):302–322, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nagi:1973:DPA

[NSS73] M. H. Nagi, E. G. Stockwell, and L. M. Snavley. Digit preference and avoidance in the age statistics of some recent African censuses: Some patterns and correlates. International Statistical Review = Revue Internationale de Statistique, 41(2):165–174, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402833.

Naatjes:2023:NRR

[NSS23] Daryan Naatjes, Stephen A. Sedory, and Sarjinder Singh. New randomised response models for two sensitive characteristics: Theory and application. International Statistical Review = Revue Internationale de Statistique, 91(3):511– 534, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Naylor:1972:BJM

[NSW72] Thomas H. Naylor, Terry G. Seaks, and D. W. Wichern. Box–Jenkins methods: An alternative to econometric models. International Statistical Review = Revue Internationale de Statistique, 40(2):123–137, August 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402755.

Naumova:2022:GSP

[NSZH22] Elena N. Naumova, Ryan B. Simpson, Bingjie Zhou, and Meghan A. Hartwick. Global seasonal and pandemic patterns in influenza: an application of longitudinal study designs. International Statistical Review = Revue Internationale de Statistique, 90:S82–S95, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nichols:1972:EBP

[NT72] W. G. Nichols and Chris P. Tsokos. Empirical Bayes point estimation in a family of probability distributions. International Statistical Review = Revue Internationale de Statistique, 40(2):147–151, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402757.

Nummelin:2002:MM

[Num02] Esa Nummelin. MC’s for MCMC’ists. International Statistical Review = Revue Internationale de Statistique, 70 (2):215–240, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403908.

Nummi:2009:SBRa

[Num09a] Tapio Nummi. Short book review: Introduction to Empirical Processes and Semiparametric Inference by Michael R. Kosorok. International Statistical Review = Revue Internationale de Statistique, 77(2):318, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919748.

Nummi:2009:SBRb

[Num09b] Tapio Nummi. Short book review: Medical Biostatistics, Second Edition, by Abhaya Indrayan. International Statistical Review = Revue Internationale de Statistique, 77 (2):319, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919749.

Nummi:2013:SBRa

[Num13a] Tapio Nummi. Short book review: Clinical Trial Data Analysis Using R by Ding-Geng (Din) Chen, Karl E. Peace. International Statistical Review = Revue Internationale de Statistique, 81(1):151, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2013:SBRb

[Num13b] Tapio Nummi. Short book review: Exercises and Solutions in Biostatistical Theory by Lawrence Kupper, Brian Neelon, Sean M. O’Brien. International Statistical Review = Revue Internationale de Statistique, 81(1):151– 152, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2013:SBRc

[Num13c] Tapio Nummi. Short book review: Regression Models as a Tool in Medical Research by Werner Vach. International Statistical Review = Revue Internationale de Statistique, 81(3):461–462, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2014:BRB

[Num14] Tapio Nummi. Book review: Applied Categorical and Count Data Analysis by Wan Tang, Hua He, Xin M. Tu. International Statistical Review = Revue Internationale de Statistique, 82(1):152–153, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2015:BRG

[Num15] Tapio Nummi. Book review: Generalized Linear Models for Categorical and Continuous Limited Dependent Variables. International Statistical Review = Revue Internationale de Statistique, 83(2):337, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Numni:2019:BRM

[Num19] Tapio Numni. Book review: Measuring Agreement-Models, Methods, and Applications, P. K. Choudhary and H. N. Nagaraja, John Wiley and Sons, 2018, xv + 360 pages, £78.99, ebook ISBN: 978-1-118-07858-7. International Statistical Review = Revue Internationale de Statistique, 87(1):188–189, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2020:BRS

[Num20] Tapio Nummi. Book review: Semiparametric Regression with R. Jaroslaw Harezlak, David Ruppert and Matt P. Wand Springer, 2018, xi + 331 pages, EUR85.39, ebook ISBN: 978-1-4939-8853-2. International Statistical Review = Revue Internationale de Statistique, 88(2):516, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2023:BRM

[Num23] Tapio Nummi. Book review: Mixed-Effects Models and Small Area Estimation, Shonosuke Sugasawa and Tatsuya Kubokava, Springer Nature, 2023, viii + 121 pages, £39.99, paperback ISBN: 978-981-19-9485-21. International Statistical Review = Revue Internationale de Statistique, 91(3):536, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Narula:1982:MSA

[NW82] Subhash C. Narula and John F. Wellington. The minimum sum of absolute errors regression: a state of the art survey. International Statistical Review = Revue Internationale de Statistique, 50(3):317–326, December 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402501.

Ngatchou-Wandji:2006:TNS

[NW06] Joseph Ngatchou-Wandji. On testing for the nullity of some skewness coefficients. International Statistical Review = Revue Internationale de Statistique, 74(1):47–65, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472690.

Nishimura:2016:AIR

[NWE16] Raphael Nishimura, James Wagner, and Michael Elliott. Alternative indicators for the risk of non-response bias: A simulation study. International Statistical Review = Revue Internationale de Statistique, 84(1):43–62, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

North:2023:RDD

[NWS23] Joshua S. North, Christopher K. Wikle, and Erin M. Schliep. A review of data-driven discovery for dynamic systems. International Statistical Review = Revue Internationale de Statistique, 91(3):464–492, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nandram:2019:BAS

[NY19] Balgobin Nandram and Yuan Yu. Bayesian analysis of a sensitive proportion for a small area. International Statistical Review = Revue Internationale de Statistique, 87 (S1):S104–S120, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nayak:2018:MIR

[NZY18] Tapan K. Nayak, Cheng Zhang, and Jiashen You. Measuring identification risk in microdata release and its control by post-randomisation. International Statistical Review = Revue Internationale de Statistique, 86(2):300–321, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ogata:1991:SSF

[OA91] Yosihiko Ogata and Katsuyuki Abe. Some statistical features of the long-term variation of the global and regional seismic activity. International Statistical Review = Revue Internationale de Statistique, 59(2):139–161, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403440.

Oakes:1981:STAa

[Oak81a] David Oakes. Survival times: Aspects of partial likelihood. International Statistical Review = Revue Internationale de Statistique, 49(3):235–252, December 1981. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402606.

With discussion and a reply by the author.

Oakes:1981:STAb

[Oak81b] David Oakes. [survival times: Aspects of partial likelihood]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 49(3):262– 264, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402612.

OBrien:2007:SBRa

[O’B07a] C. M. O’Brien. Short book review: Adaptive Design Methods in Clinical Trials by Shein-Chung Chow, Mark Chang. International Statistical Review = Revue Internationale de Statistique, 75(2):278, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508493.

OBrien:2007:SBRb

[O’B07b] C. M. O’Brien. Short book review: Analysing Ecological Data by Alain F. Zuur, Elena N. Ieno, Graham M. Smith. International Statistical Review = Revue Internationale de Statistique, 75(3):426–427, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509896.

OBrien:2007:SBRc

[O’B07c] C. M. O’Brien. Short book review: Case Study Research: Principles and Practices by John Gerring. International Statistical Review = Revue Internationale de Statistique, 75(3):432, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41509904.

OBrien:2008:SBRa

[O’B08a] C. M. O’Brien. Short book review: Adaptive Design Theory and Implementation Using SAS and R by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 76(2):300, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919619.

OBrien:2008:SBRd

[O’B08b] C. M. O’Brien. Short book review: Negative Binomial Regression by Joseph M. Hilbe. International Statistical Review = Revue Internationale de Statistique, 76(2):302– 303, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919622.

OBrien:2008:SBRc

[O’B08c] C. M. O’Brien. Short book review: Sample Size Calculations in Clinical Research, Second Edition, by Shein-Chung Chow; Jun Shao; Hansheng Wang. International Statistical Review = Revue Internationale de Statistique, 76(2):301–302, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919621.

OBrien:2008:SBRb

[O’B08d] C. M. O’Brien. Short book review: Statistical Design and Analysis of Stability Studies by Shein-Chung Chow. International Statistical Review = Revue Internationale de Statistique, 76(2):301, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919620.

OBrien:2009:SBRc

[O’B09a] C. M. O’Brien. Short book review: Bioconductor Case Studies by Florian Hahne, Wolfgang Huber, Robert Gentleman, Seth Falcon. International Statistical Review = Revue Internationale de Statistique, 77(1):158–159, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919709.

OBrien:2009:SBRi

[O’B09b] C. M. O’Brien. Short book review: Computational Intelligence in Medical Imaging: Techniques and Applications edited by Gerald Schaefer, Aboul Ella Hassanien, Jianmin Jiang. International Statistical Review = Revue Internationale de Statistique, 77(3):475, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919780.

OBrien:2009:SBRe

[O’B09c] C. M. O’Brien. Short book review: Introduction to Scientific Programming and Simulation Using R by [E. Stuart] Owen-Jones, Robert Maillardet, Andrew Robinson. International Statistical Review = Revue Internationale de Statistique, 77(3):471, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919776.

OBrien:2009:SBRf

[O’B09d] C. M. O’Brien. Short book review: Introduction to Spatial Econometrics by James LeSage, Robert Kelley Pace. International Statistical Review = Revue Internationale de Statistique, 77(3):472, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919777.

OBrien:2009:SBRb

[O’B09e] C. M. O’Brien. Short book review: Meta-analysis of Binary Data Using Profile Likelihood by Dankmar Böhning, Ronny Kuhnert, Sasivimol Rattanasiri. International Statistical Review = Revue Internationale de Statistique, 77 (1):157–158, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919708.

OBrien:2009:SBRg

[O’B09f] C. M. O’Brien. Short book review: Nonlinear Regression with R by Christian Ritz, Jens Carl Streibig. International Statistical Review = Revue Internationale de Statistique, 77(3):473, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919778.

OBrien:2009:SBRh

[O’B09g] C. M. O’Brien. Short book review: Statistical Detection and Surveillance of Geographic Clusters by Peter Rogerson, Ikuho Yamada. International Statistical Review = Revue Internationale de Statistique, 77(3):474, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919779.

OBrien:2009:SBRa

[O’B09h] C. M. O’Brien. Short book review: Statistical Learning from a Regression Perspective by Richard A. Berk. International Statistical Review = Revue Internationale de Statistique, 77(1):157, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919707.

OBrien:2009:SBRd

[O’B09i] C. M. O’Brien. Short book review: Wavelet Methods in Statistics with R by G. P. Nason. International Statistical Review = Revue Internationale de Statistique, 77 (1):159, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919710.

OBrien:2010:SBRc

[O’B10a] Carl M. O’Brien. Short book review: An Integrated Introduction to Computer Graphics and Geometric Modeling by Ronald Goldman. International Statistical Review = Revue Internationale de Statistique, 78(1):156–157, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919823.

OBrien:2010:SBRb

[O’B10b] Carl M. O’Brien. Short book review: Bayesian Methods for Measures of Agreement by Lyle D. Broemeling. International Statistical Review = Revue Internationale de Statistique, 78(1):155–156, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919822.

OBrien:2010:SBRd

[O’B10c] Carl M. O’Brien. Short book review: Dynamic Linear Models with R by Giovanni Petris, Sonia Petrone, Patrizia Campagnoli. International Statistical Review = Revue Internationale de Statistique, 78(1):157, April 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919824.

OBrien:2010:SBRe

[O’B10d] Carl M. O’Brien. Short book review: The Top Ten Algorithms in Data Mining edited by Xindong Wu, Vipin Kumar. International Statistical Review = Revue Internationale de Statistique, 78(1):158, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919825.

OBrien:2010:SBRa

[O’B10e] Carl M. O’Brien. Short book review: When Less is More: Visualizing Basic Inequalities by Claudi Alsina, Roger B. Nelsen. International Statistical Review = Revue Internationale de Statistique, 78(1):155, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919821.

OBrien:2011:SBRa

[O’B11a] Carl M. O’Brien. Short book review: Algorithms and Theory of Computation Handbook, Second Edition — 2 Volume Set by Mikhail J. Atallah, Marina Blanton. International Statistical Review = Revue Internationale de Statistique, 79(1):141–142, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41306192.

OBrien:2011:SBRc

[O’B11b] Carl M. O’Brien. Short book review: Negative Binomial Regression, Second Edition, by Joseph M. Hilbe. International Statistical Review = Revue Internationale de Statistique, 79(3):483–484, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305064.

OBrien:2011:SBRb

[O’B11c] Carl M. O’Brien. Short book review: Stochastic Control and Mathematical Modeling: Applications in Economics by Hiroaki Morimoto. International Statistical Review = Revue Internationale de Statistique, 79(1):143, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41306193.

OBrien:2012:SBRf

[O’B12a] Carl M. O’Brien. Short book review: Adaptive Design Methods in Clinical Trials, Second Edition, by Shein-Chung Chow, Mark Chang. International Statistical Review = Revue Internationale de Statistique, 80(3):474, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2012:SBRg

[O’B12b] Carl M. O’Brien. Short book review: Analysis of Questionnaire Data with R by Bruno Falissard. International Statistical Review = Revue Internationale de Statistique, 80(3):474–475, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2012:SBRi

[O’B12c] Carl M. O’Brien. Short book review: Bayesian Analysis Made Simple: An Excel GUI for WinBUGS by Phil Woodward. International Statistical Review = Revue Internationale de Statistique, 80(3):476, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2012:SBRd

[O’B12d] Carl M. O’Brien. Short book review: Choice-Based Conjoint Analysis: Models and Designs by Damaraju Raghavarao, James B. Wiley, Pallavi Chitturi. International Statistical Review = Revue Internationale de Statistique, 80(1):203, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/23257199.

OBrien:2012:SBRc

[O’B12e] Carl M. O’Brien. Short book review: Composite Sampling: A Novel Method to Accomplish Observational Economy in Environmental Studies by Ganapati P. Patil, Sharad D. Gore, Charles Taillie. International Statistical Review = Revue Internationale de Statistique, 80(1):202–203, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257198.

OBrien:2012:SBRe

[O’B12f] Carl M. O’Brien. Short book review: Forest Analytics with R: An Introduction by Andrew P. Robinson, Jeff D. Hamann. International Statistical Review = Revue Internationale de Statistique, 80(1):203–204, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/23257200.

OBrien:2012:SBRb

[O’B12g] Carl M. O’Brien. Short book review: Modelling and Quantitative Methods in Fisheries, Second Edition, by Malcolm Haddon. International Statistical Review = Revue Internationale de Statistique, 80(1):201–202, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/23257197.

OBrien:2012:SBRa

[O’B12h] Carl M. O’Brien. Short book review: Numerical Ecology with R by Daniel Borcard, François Gillet, Pierre Legendre. International Statistical Review = Revue Internationale de Statistique, 80(1):201, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/23257196.

OBrien:2012:SBRh

[O’B12i] Carl M. O’Brien. Short book review: Smoothing Splines: Methods and Applications by Yuedong Wang. International Statistical Review = Revue Internationale de Statistique, 80(3):475–476, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRa

[O’B13a] Carl M. O’Brien. Short book review: An Introduction to Bootstrap Methods with Applications to R by Michael R. Chernick, Robert A. LaBudde. International Statistical Review = Revue Internationale de Statistique, 81(1): 157–158, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRh

[O’B13b] Carl M. O’Brien. Short book review: Applied Operational Research with SAS by Ali Emrouznejad, William Ho. International Statistical Review = Revue Internationale de Statistique, 81(3):474, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRf

[O’B13c] Carl M. O’Brien. Short book review: Biostatistics with R: An Introduction to Statistics Through Biological Data by Babak Shahbaba. International Statistical Review = Revue Internationale de Statistique, 81(3):472–473, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRb

[O’B13d] Carl M. O’Brien. Short book review: How to be a Quantitative Ecologist: The ‘A to R’ of Green Mathematics and Statistics by Jason Matthiopoulos. International Statistical Review = Revue Internationale de Statistique, 81 (1):158, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRd

[O’B13e] Carl M. O’Brien. Short book review: Introduction to Hierarchical Bayesian Modeling for Ecological Data by Éric Parent, Étienne Rivot. International Statistical Review = Revue Internationale de Statistique, 81(3):471, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRg

[O’B13f] Carl M. O’Brien. Short book review: Programming in C++ for Engineering and Science by Larry Nyhoff. International Statistical Review = Revue Internationale de Statistique, 81(3):473, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRc

[O’B13g] Carl M. O’Brien. Short book review: Statistical Methods for Fuzzy Data by Reinhard Viertl. International Statistical Review = Revue Internationale de Statistique, 81(1): 158–159, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2013:SBRe

[O’B13h] Carl M. O’Brien. Short book review: Statistics for Bioengineering Sciences: With MATLAB and WinBUGS Support by Brani Vidakovic. International Statistical Review = Revue Internationale de Statistique, 81(3):471– 472, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2014:BRC

[O’B14a] Carl M. O’Brien. Book review: Computational Approaches to Studying the Co-evolution of Networks and Behavior in Social Dilemmas by Rense Corten. International Statistical Review = Revue Internationale de Statistique, 82(3):477–478, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2014:BRF

[O’B14b] Carl M. O’Brien. Book review: Functional and Phylogenetic Ecology in R by Nathan G. Swenson. International Statistical Review = Revue Internationale de Statistique, 82(3):486–487, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2014:BRS

[O’B14c] Carl M. O’Brien. Book review: Statistical Applications for Environmental Analysis and Risk Assessment by Joseph Ofungwu. International Statistical Review = Revue Internationale de Statistique, 82(3):487–488, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2015:BRG

[O’B15a] Carl M. O’Brien. Book review: Growth Curve Modelling: Theory and Applications. International Statistical Review = Revue Internationale de Statistique, 83(3):517, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2015:BRM

[O’B15b] Carl M. O’Brien. Book review: Mixed Effects Models for the Population Approach: Models, Tasks, Methods and Tools. International Statistical Review = Revue Internationale de Statistique, 83(2):334–335, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2015:APC

[O’B15c] Robert M. O’Brien. Age-Period-Cohort Models: Approaches and Analyses with Aggregate Data. Chapman and Hall/CRC statistics in the social and behavioral sciences series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-5153-4 (hardcover), 1-4665-5154-2 (e-book). xi + 204 pp. LCCN HB849.47 .O27 2015.

OBrien:2016:BRD

[O’B16a] Carl M. O’Brien. Book review: Data Analysis with Competing Risks and Intermediate States. International Statistical Review = Revue Internationale de Statistique, 84(2): 319–320, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2016:BRN

[O’B16b] Carl M. O’Brien. Book review: Nonparametric Estimation under Shape Constraints: Estimators, Algorithms and Asymptotics. International Statistical Review = Revue Internationale de Statistique, 84(2):318–319, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2016:BRS

[O’B16c] Carl M. O’Brien. Book review: Statistical Learning with Sparsity: The Lasso and Generalizations. International Statistical Review = Revue Internationale de Statistique, 84(1):156–157, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2020:BRS

[O’B20] Carl M. O’Brien. Book review: Sampling and Estimation from Finite Populations, Yves Tillé, Wiley, 2020, xxviii + 419 pages, £65/$100, hardcover ISBN: 978-0-470-68205-0. International Statistical Review = Revue Internationale de Statistique, 88(3):804–806, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2021:BRP

[O’B21] Carl M. O’Brien. Book review: Parameter Redundancy and Identifiability, Diana Cole Chapman and Hall/CRC, 2020, xx + 252 pages, £140, hardcover ISBN: 978-1-4987-2087-8. International Statistical Review = Revue Internationale de Statistique, 89(3):657–658, December 2021. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OKeefe:2013:SAM

[OC13] Christine M. O’Keefe and James O. Chipperfield. A summary of attack methods and confidentiality protection measures for fully automated remote analysis systems. International Statistical Review = Revue Internationale de Statistique, 81(3):426–455, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Oliveira:2017:DSS

[OCRC17] María Oliveira, Rosa M. Crujeiras, and Alberto Rodríguez-Casal. Discussions on statistical scale space methods: Discussion. International Statistical Review = Revue Internationale de Statistique, 85(1):31–32, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ograjensek:2016:ESE

[OG16] Irena Ograjenšek and Iddo Gal. Enhancing statistics education by including qualitative research. International Statistical Review = Revue Internationale de Statistique, 84(2):165–178, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ogawa:1973:BRB

[Oga73] J. Ogawa. Book review: Nonparametric Methods in Multivariate Analysis by M. L. Puri; P. K. Sen. International Statistical Review = Revue Internationale de Statistique, 41(2):294, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402859.

ORourke:2013:SSA

[OH13] Norm O’Rourke and Larry Hatcher. A Step-by-Step Approach to Using SAS(R) for Factor Analysis and Structural Equation Modeling. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 2013. ISBN 1-59994-230-5. xxii + 420 pp. LCCN QA276.4 H37 2013.

Ott:2019:BCV

[OH19] Manuela Ott and Leonhard Held. Bayesian calibration of p-values from Fisher’s exact test. International Statistical Review = Revue Internationale de Statistique, 87(2):285– 305, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ohlsson:1977:KK

[Ohl77] Ingvar Ohlsson. Karin Kock, 1891–1976. International Statistical Review = Revue Internationale de Statistique, 45(1):109, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403030.

Oja:2007:SBR

[Oja07] Hannu Oja. Short book review: Introduction to Clustering Large and High-Dimensional Data by Jacob Kogan. International Statistical Review = Revue Internationale de Statistique, 75(3):434–435, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41509906.

Oja:2008:SBRb

[Oja08a] Hannu Oja. Short book review: Applied Nonparametric Statistical Methods, Fourth Edition, by Peter Sprent, Nigel C. Smeeton. International Statistical Review = Revue Internationale de Statistique, 76(1):152, April 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919602.

Oja:2008:SBRa

[Oja08b] Hannu Oja. Short book review: Nonparametric Statistics with Applications to Science and Engineering by Paul H. Kvam, Brani Vidakovic. International Statistical Review = Revue Internationale de Statistique, 76(1):150– 151, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919600.

Oja:2009:SBR

[Oja09] Hannu Oja. Short book review: Asymptotic Theory of Statistics and Probability by Anirban DasGupta. International Statistical Review = Revue Internationale de Statistique, 77(1):160–161, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919711.

Oja:2011:SBR

[Oja11] Hannu Oja. Short book review: Robust Nonparametric Statistical Methods, Second Edition, by Thomas P. Hettmansperger, Joseph W. McKean. International Statistical Review = Revue Internationale de Statistique, 79 (2):300, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305050.

OKeefe:2011:D

[O’K11] Christine M. O’Keefe. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2): 191–194, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OKeefe:2012:DSD

[O’K12] Christine M. O’Keefe. Discussions: “Statistical disclosure risk: separating potential and harm” [MR3006147]. International Statistical Review = Revue Internationale de Statistique, 80(3):369–372, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Olenski:2003:CRI

[Ole03] Józef Oleński. The citizens’ right to information and the duties of a democratic state in modern IT environment: In the light of the UN fundamental principles of official statistics and the ISI declaration on statistical ethics. International Statistical Review = Revue Internationale de Statistique, 71(1):33–48, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403872.

Olive:2014:STI

[Oli14] David J. Olive. Statistical Theory and Inference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-04971-2, 3-319-04972-0 (e-book). xii + 434 pp. LCCN

QA276 .O45 2014. URL http://www.loc.gov/catdir/
enhancements/fy1407/2014934894-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy1407/2014934894-
d.html;   http://www.loc.gov/catdir/enhancements/
fy1407/2014934894-t.html.

Oh:2005:RNA

[OMGT05] Jong Nam Oh, Pilar Martin-Guzman, and Klaus Trutzel. The role of NSOs in assuring national prosperity: A case study of Korea. International Statistical Review = Revue Internationale de Statistique, 73(2):191–196, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472654.

ONeill:2004:CPP

[O’N04] Brian C. O’Neill. Conditional probabilistic population projections: An application to climate change. International Statistical Review = Revue Internationale de Statistique, 72(2):167–184, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403852.

ONeill:2009:ECB

[O’N09] Ben O’Neill. Exchangeability, correlation, and Bayes’ effect. International Statistical Review = Revue Internationale de Statistique, 77(2):241–250, August 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919725.

Onicescu:1978:VB

[Oni78] Octav Onicescu. Valeriu Bulgaru, 1898–1978. International Statistical Review = Revue Internationale de Statistique, 46(3):325, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402394.

Onicescu:1979:BRB

[Oni79] Octav Onicescu. Book review: Personelle und statistische Wahrscheinlichkeit by W. Stegmüller. International Statistical Review = Revue Internationale de Statistique, 47 (2):234, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402580.

Onicescu:1980:BRB

[Oni80] Octav Onicescu. Book review: Personelle und Statistische Wahrscheinlichkeit by W. Stegmüller. International Statistical Review = Revue Internationale de Statistique, 48(1):133, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402439.

Onicescu:1983:GM

[Oni83] O. Onicescu. Gheorghe Mihoc, 1906–1981. International Statistical Review = Revue Internationale de Statistique, 51(3):323–327, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402592.

Ord:1972:BRB

[Ord72] Keith Ord. Book review: The Chi-Squared Distribution by H. O. Lancaster. International Statistical Review = Revue Internationale de Statistique, 40(3):395–396, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402483.

Ord:1980:BRBa

[Ord80a] Keith Ord. Book review: An Introduction to Bilinear Time Series Models by C. W. J. Granger; A. P. Anderson. International Statistical Review = Revue Internationale de Statistique, 48(2):238, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403173.

Ord:1980:BRBb

[Ord80b] Keith Ord. Book review: Forecasting Methods for Management by S. C. Wheelwright; S. Makridakis. International Statistical Review = Revue Internationale de Statistique, 48(2):244–245, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403186.

Oksanen:1973:ICB

[OS73] Ernest H. Oksanen and Byron G. Spencer. International consumption behaviour: Tests of alternative consumption function hypotheses using national accounting data for twelve countries. International Statistical Review = Revue Internationale de Statistique, 41(1):69–76, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402788.

Olkin:1993:MSS

[OS93] Ingram Olkin and Michael A. Stephens. On making the shortlist for the selection of candidates. International Statistical Review = Revue Internationale de Statistique, 61 (3):477–486, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403757.

Osborne:1991:SCR

[Osb91] Christine Osborne. Statistical calibration: A review. International Statistical Review = Revue Internationale de Statistique, 59(3):309–336, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403690.

Osborne:1992:FMS

[Osb92] M. R. Osborne. Fisher’s method of scoring. International Statistical Review = Revue Internationale de Statistique, 60(1):99–117, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403504.

Ottestad:1975:CAA

[Ott75a] Per Ottestad. Component analysis: An alternative system. International Statistical Review = Revue Internationale de Statistique, 43(1):83–107, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402662.

Ottestad:1975:DA

[Ott75b] Per Ottestad. Discrimination analysis. International Statistical Review = Revue Internationale de Statistique, 43 (3):301–315, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403114.

Ottestad:1976:CCA

[Ott76] Per Ottestad. Comment: “Component analysis: a second alternative” (Internat. Statist. Rev. 43 (1975), no. 1, 83– 107). International Statistical Review = Revue Internationale de Statistique, 44(3):367–368, December 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402354.

Ottaviani:2002:DST

[Ott02] M. Gabriella Ottaviani. Discussion: Statistics, from a tool for state and society to a tool for all citizens. International Statistical Review = Revue Internationale de Statistique, 70(1):30–32, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403715.

Ottaviani:2005:RSE

[Ott05] Maria-Gabriella Ottaviani. Research into statistics education as a discipline. International Statistical Review = Revue Internationale de Statistique, 73(2):207–209, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472656.

Ovuworie:1980:MCQ

[Ovu80] Godwin C. Ovuworie. Multi-channel queues: A survey and bibliography. International Statistical Review = Revue Internationale de Statistique, 48(1):49–71, April 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402405.

Olkin:1992:GRA

[OY92] Ingram Olkin and Shlomo Yitzhaki. Gini regression analysis. International Statistical Review = Revue Internationale de Statistique, 60(2):185–196, August 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403649.

Osgood-Zimmerman:2023:SRT

[OZW23] Aaron Osgood-Zimmerman and Jon Wakefield. A statistical review of template model builder: a flexible tool for spatial modelling. International Statistical Review = Revue Internationale de Statistique, 91(2):318–342, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Perez-Abreu:2015:ENC

[PA15] Victor Perez-Abreu. Editorial: New Co-Editor-in-Chief of the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 83(1): 1, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Porcu:2018:MTE

[PAF18] Emilio Porcu, Alfredo Alegria, and Reinhard Furrer. Modeling temporally evolving and spatially globally dependent data. International Statistical Review = Revue Internationale de Statistique, 86(2):344–377, August 2018. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Palazzi:1972:BRB

[Pal72] Aurelio Palazzi. Book review: Statistical Tolerance Regions: Classical and Bayesian by I. Guttman. International Statistical Review = Revue Internationale de Statistique, 40(3):395, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402482.

Panaretos:1989:ES

[Pan89] John Panaretos. On the evolution of surnames. International Statistical Review = Revue Internationale de Statistique, 57(2):161–167, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403384.

Pandher:2007:MCM

[Pan07] Gurupdesh S. Pandher. Modelling & controlling monetary and economic identities with constrained state space models. International Statistical Review = Revue Internationale de Statistique, 75(2):150–169, August 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41508458.

Parenti:1973:BRB

[Par73] G. Parenti. Book review: Scientific Truth and Statistical Method by M. Boldrini. International Statistical Review = Revue Internationale de Statistique, 41(3):395– 396, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402631.

Parniczky:1980:IWT

[Par80] Gabor Parniczky. On the inconsistency of world trade statistics. International Statistical Review = Revue Internationale de Statistique, 48(1):43–48, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402404.

Park:2005:DSH

[Par05] Sung H. Park. Discussant: Sung H. Park. International Statistical Review = Revue Internationale de Statistique, 73(2):181–182, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Parzen:2013:DCD

[Par13] Emanuel Parzen. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1):48–52, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Patterson:1973:BRB

[Pat73] H. D. Patterson. Book review: Constructions and Combinatorial Problems in Design of Experiments by D. Raghavarao. International Statistical Review = Revue Internationale de Statistique, 41(2):294–295, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402860.

Patino:1974:GL

[Pat74] Emilio Alanĭs Patiño. Gilberto Loyo, 1901–1973. International Statistical Review = Revue Internationale de Statistique, 42(3):323–324, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402995.

Patil:1975:BSI

[Pat75a] G. P. Patil. The Advanced Study Institute on statistical distributions in scientific work. International Statistical Review = Revue Internationale de Statistique, 43(3):341– 344, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403118.

Patterson:1975:BRB

[Pat75b] H. D. Patterson. Book review: Tables of Two-Associate Class Partially Balanced Designs by W. H. Clatworthy. International Statistical Review = Revue Internationale de Statistique, 43(3):367–368, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403123.

Patil:1979:SPS

[Pat79] G. P. Patil. The satellite program in statistical ecology. International Statistical Review = Revue Internationale de Statistique, 47(2):223–228, August 1979. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402569.

Paterson:1988:SRW

[Pat88] Lindsay J. Paterson. Some recent work on making incomplete-block designs available as a tool for science. International Statistical Review = Revue Internationale de Statistique, 56(2):129–138, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403637.

Pauli:2019:SMI

[Pau19] Francesco Pauli. A statistical model to investigate the reproducibility rate based on replication experiments. International Statistical Review = Revue Internationale de Statistique, 87(1):68–79, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pierce:2017:MLF

[PB17] Donald Alan Pierce and Ruggero Bellio. Modern likelihood-frequentist inference. International Statistical Review = Revue Internationale de Statistique, 85(3):519– 541, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pfeffermann:2019:ERM

[PBH19] Danny Pfeffermann and Dano Ben-Hur. Estimation of randomisation mean square error in small area estimation. International Statistical Review = Revue Internationale de Statistique, 87(S1):S31–S49, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Plass:2019:RIC

[PCA^{+}19] Julia Plass, Marco E. G. V. Cattaneo, Thomas Augustin, Georg Schollmeyer, and Christian Heumann. Reliable inference in categorical regression analysis for non-randomly coarsened observations. International Statistical Review = Revue Internationale de Statistique, 87(3):580–603, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pearson:1974:MIF

[Pea74] E. S. Pearson. Memories of the impact of Fisher’s work in the 1920s. International Statistical Review = Revue Internationale de Statistique, 42(1):4–8, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402680.

Pearl:2002:CNS

[Pea02a] Judea Pearl. Comments on Nozer Singpurwalla’s “On Causality and Causal Mechanisms”. International Statistical Review = Revue Internationale de Statistique, 70(2): 210–212, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403905.

Pearl:2002:CSD

[Pea02b] Judea Pearl. Comments on seeing and doing. International Statistical Review = Revue Internationale de Statistique, 70(2):207–209, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403904.

Pearl:2002:SRR

[Pea02c] Judea Pearl. A short rejoinder to rejoinder. International Statistical Review = Revue Internationale de Statistique, 70(2):214, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403907.

Pedersen:1978:FI

[Ped78] J. G. Pedersen. Fiducial inference. International Statistical Review = Revue Internationale de Statistique, 46(2): 147–170, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402811.

Perthel:1975:ELR

[Per75] D. Perthel. Engel’s law revisited. International Statistical Review = Revue Internationale de Statistique, 43(2):211– 218, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402900.

Perez:1993:RHT

[Pér93] Alfonso García Pérez. On robustness for hypotheses testing. International Statistical Review = Revue Internationale de Statistique, 61(3):369–385, December 1993. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403749.

PerpinanLamigueiro:2014:DTS

[Per14] Oscar Perpiñán Lamigueiro. Displaying Time Series, Spatial, and Space–Time Data with R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-6520-9 (hardcover). vii + 200 pp. LCCN QA280 .P475 2014.

Peto:1985:DAT

[Pet85] Richard Peto. Discussion of ‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage. International Statistical Review = Revue Internationale de Statistique, 53(1):31–34, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402875.

Peterson:2007:SBR

[Pet07] A. Townsend Peterson. Short book review: Niche Modeling: Predictions from Statistical Distributions by David R. B. Stockwell. International Statistical Review = Revue Internationale de Statistique, 75(2):252, August 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41508467.

Pewsey:2015:DFD

[Pew15] Arthur Pewsey. Discussion on On Families of Distributions with Shape Parameters: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(2):211–217, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Pfeffermann:1982:FWD

[PF82] Dany Pfeffermann and Jeffrey M. Fisher. Festival and working days prior adjustments in economic time series. International Statistical Review = Revue Internationale de Statistique, 50(2):113–124, August 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402595.

Pfeffermann:1993:RSW

[Pfe93] Danny Pfeffermann. The role of sampling weights when modeling survey data. International Statistical Review = Revue Internationale de Statistique, 61(2):317–337, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403631.

Pfeffermann:2002:SAE

[Pfe02] Danny Pfeffermann. Small area estimation: New developments and directions. International Statistical Review = Revue Internationale de Statistique, 70(1):125–143, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403729.

Pendergast:1996:SMA

[PGN^{+}96] Jane F. Pendergast, Stephen J. Gange, Michael A. Newton, Mary J. Lindstrom, Mari Palta, and Marian R. Fisher. A survey of methods for analyzing clustered binary response data. International Statistical Review = Revue Internationale de Statistique, 64(1):89–118, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403425.

Perusquia:2022:BMA

[PGV22] José A. Perusquía, Jim E. Griffin, and Cristiano Villa. Bayesian models applied to cyber security anomaly detection problems. International Statistical Review = Revue Internationale de Statistique, 90(1):78–99, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pullum:1986:MEL

[PHO86] Thomas W. Pullum, Trudy Harpham, and Nuri Ozsever. The machine editing of large-sample surveys: The experience of the world fertility survey. International Statistical Review = Revue Internationale de Statistique, 54(3):311– 326, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403061.

Pinsky:1975:BRB

[Pin75] M. Pinsky. Book review: Stochastic Differential Equations by I. I. Gihman; A. V. Skorohod. International Statistical Review = Revue Internationale de Statistique, 43(2): 245, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402916.

Pinkasovych:2010:ANC

[Pin10a] Liliana Pinkasovych. Announcement: New Co-Editors-in-Chief for the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 78(3):329, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pinkasovych:2010:NCE

[Pin10b] Liliana Pinkasovych. New Co-EditorsinChief for the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 78(3):329, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919856.

Pizzinga:2010:CKF

[Piz10] Adrian Pizzinga. Constrained Kalman filtering: Additional results. International Statistical Review = Revue Internationale de Statistique, 78(2):189–208, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919831.

Park:2017:DSS

[PJK17] Cheolwoo Park, Yongho Jeon, and Kee-Hoon Kang. Discussions on statistical scale space methods: Discussion. International Statistical Review = Revue Internationale de Statistique, 85(1):33–35, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Purcell:1980:PEL

[PK80] Noel J. Purcell and Leslie Kish. Postcensal estimates for local areas (or domains). International Statistical Review = Revue Internationale de Statistique, 48(1):3–18, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402400.

Plackett:1975:BRB

[Pla75] R. L. Plackett. Book review: Statistical Methods for Rates and Proportions by Joseph L. Fleiss. International Statistical Review = Revue Internationale de Statistique, 43(2): 244, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402915.

Plackett:1983:KPC

[Pla83] R. L. Plackett. Karl Pearson and the chi-squared test. International Statistical Review = Revue Internationale de Statistique, 51(1):59–72, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402731.

Plackett:1988:DAB

[Pla88] R. L. Plackett. Data analysis before 1750. International Statistical Review = Revue Internationale de Statistique, 56(2):181–195, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403641.

Plegue:2015:BRI

[Ple15] Melissa Plegue. Book review: An Introduction to STATA for Health Researchers. International Statistical Review = Revue Internationale de Statistique, 83(3):511–512, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Park:2015:RVC

[PMLL15a] Byeong U. Park, Enno Mammen, Young K. Lee, and Eun Ryung Lee. Rejoinder to varying coefficient regression models: A review and new developments: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 83(1):72–76, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Park:2015:VCR

[PMLL15b] Byeong U. Park, Enno Mammen, Young K. Lee, and Eun Ryung Lee. Varying coefficient regression models: A review and new developments. International Statistical Review = Revue Internationale de Statistique, 83(1): 36–64, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pino:1987:MFL

[PMM87] Francisco A. Pino, Pedro A. Morettin, and Raúl P. Mentz. Modelling and forecasting linear combinations of time series. International Statistical Review = Revue Internationale de Statistique, 55(3):295–313, December 1987. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403407.

Pewsey:2013:CSR

[PNR13] Arthur Pewsey, Markus Neuhäuser, and Graeme D. Ruxton. Circular Statistics in R. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967113-3. xiv + 183 pp. LCCN

QA276 .P45 2013. URL http://www.loc.gov/catdir/
enhancements/fy1410/2013940576-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy1410/2013940576-
d.html;   http://www.loc.gov/catdir/enhancements/
fy1410/2013940576-t.html.

Podoluk:1972:BRB

[Pod72] J. R. Podoluk. Book review: Six Papers on the Size Distribution of Wealth and Income by L. Soltow. International Statistical Review = Revue Internationale de Statistique, 40(1):101, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402714.

Podgorski:2011:SBR

[Pod11] Krzysztof Podgórski. Short book review: Probability, Statistics, & Financial Math by Peter Caithamer. International Statistical Review = Revue Internationale de Statistique, 79(1):124–125, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41306177.

Podgorski:2014:BRBa

[Pod14a] Krzysztof Podgorski. Book review: Advances in Machine Learning and Data Mining for Astronomy edited by Michael J. Way, Jeffrey D. Scargle, Kamal M. Ali, and Ashok N. Srivstava. International Statistical Review = Revue Internationale de Statistique, 82(1):153–154, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2014:BRBb

[Pod14b] Krzysztof Podgorski. Book review: Exercises in Probability: A Guided Tour from Measure Theory to Random Processes,via Conditioning by Loïc Chaumont and Marc Yor. International Statistical Review = Revue Internationale de Statistique, 82(1):154–155, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2015:BRB

[Pod15a] Krzysztof Podgorski. Book review: Bayesian Methods for Management and Business-Pragmatic Solutions for Real Problems. International Statistical Review = Revue Internationale de Statistique, 83(3):519–520, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2015:BRD

[Pod15b] Krzysztof Podgorski. Book review: Displaying Time Series, Spatial and Space-Time Data with R. International Statistical Review = Revue Internationale de Statistique, 83(1):165, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2016:BRM

[Pod16] Krzysztof Podgórski. Book review: Multivariate Statistics: Exercises and Solutions. International Statistical Review = Revue Internationale de Statistique, 84(2):317– 318, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2017:BRI

[Pod17a] Krzysztof Podgorski. Book review: Introduction to Data Science, Laura Igual and Santi Seguí Springer, 2017, xiv + 218 pages, C48.14, softcover ISBN: 978-3-319-50016-4. International Statistical Review = Revue Internationale de Statistique, 85(2):371–372, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2017:BRM

[Pod17b] Krzysztof Podgórski. Book review: Monte Carlo Methods and Stochastic Processes — From Linear to Non-Linear, Emmanuel Gobet Chapman and Hall/CRC, 2016, xiii + 310 pages, £44.99, hardback ISBN: 978-1-4987-4622-9. International Statistical Review = Revue Internationale de Statistique, 85(1):179–180, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2017:BRS

[Pod17c] Krzysztof Podgórski. Book review: Stochastic Volatility Modelling, Lorenzo Bergomi, Chapman and Hall/CRC, 2016, xvi + 522 pages, £57.99, hardback ISBN: 978-1-4822-4406-9. International Statistical Review = Revue Internationale de Statistique, 85(1):180–181, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2018:BRF

[Pod18] Krzysztof Podgorski. Book review: Financial Analytics with R, Mark J. Bennet and Dirk L. Hugen, Cambridge University Press, 2016, 377 pages, EUR55.99, hardcover ISBN: 978-1-107-15075-1. International Statistical Review = Revue Internationale de Statistique, 86(1):163– 165, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2021:BRC

[Pod21a] Krzysztof Podgórski. Book review: Computational Genomics with R, Altuna Akalin Chapman and Hall/CRC, 2021, xxii + 440 pages, £99.99, hardcover ISBN: 978-1-4987-8185-5. International Statistical Review = Revue Internationale de Statistique, 89(2):420–421, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2021:BRN

[Pod21b] Krzysztof Podgórski. Book review: Numerical Weather Prediction and Data Assimilation, Petros Katsafados, Elias Mavromatidis and Christos Spyrou Wiley, 2020, x + 208 pages, $165, hardcover ISBN: 978-1-78630-141-3. International Statistical Review = Revue Internationale de Statistique, 89(2):431–432, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2022:BRP

[Pod22a] Krzysztof Podgórski. Book review: Practical Smoothing: The Joys of P-splines, Paul H. C. Eilers and Brian D. Marx, Cambridge University Press, 2021, xii + 199 pages, $59.99, hardcover, ISBN: 978-1-1084-8295-0. International Statistical Review = Revue Internationale de Statistique, 90(1):191–192, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2022:BRW

[Pod22b] Krzysztof Podgórski. Book review: Wavelets from a Statistical Perspective. Maarten Jansen. Chapman and Hall/ CRC, 2022, xix + 325 pages, $96 (hardcover). ISBN: 978-1-032-20067-5 (hardcover). International Statistical Review = Revue Internationale de Statistique, 90(2):414– 415, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2024:BRG

[Pod24] Krzysztof Podgórski. Book review: -Geographic Data Science With Python, Sergio Rey, Dani Arribas-Bel, Levi John Wolf CRC Press, 2023, xiv + 410 pages, 119 Color & 4 B/W Illustrations, £39.19/$89.95, paperback ISBN: 978-1-032-44595-3 (pbk). International Statistical Review = Revue Internationale de Statistique, 92(1):134– 135, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Polasek:2011:SBRb

[Pol11a] Wolfgang Polasek. Short book review: Frontiers of Statistical Decision Making and Bayesian Analysis: In Honor of James O. Berger edited by Ming-Hui Chen, Dipak K. Dey, Peter Müller, Dongchu Sun, Keying Ye. International Statistical Review = Revue Internationale de Statistique, 79(3):482–483, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305063.

Polasek:2011:SBRa

[Pol11b] Wolfgang Polasek. Short book review: Using R for Data Management, Statistical Analysis, and Graphics by Nicholas J. Horton, Ken Kleinman. International Statistical Review = Revue Internationale de Statistique, 79(2): 284–285, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305035.

Polasek:2012:SBRa

[Pol12a] Wolfgang Polasek. Short book review: Handbook of Markov Chain Monte Carlo edited by Steve Brooks, Andrew Gelman, Galin Jones, Xiao-Li Meng. International Statistical Review = Revue Internationale de Statistique, 80(1):184–185, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257180.

Polasek:2012:SBRb

[Pol12b] Wolfgang Polasek. Short book review: Origami 5: Fifth International Meeting of Origami Science, Mathematics, and Education edited by Patsy Wang-Iverson, Robert J. Lang, Mark Yim. Origami Inspirations by Meenakshi Mukerji. International Statistical Review = Revue Internationale de Statistique, 80(1):197–199, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/23257193.

Polasek:2013:SBRa

[Pol13a] Wolfgang Polasek. Short book review: Matrix Tricks for Linear Statistical Models: Our Personal Top Twenty by Simo Puntanen, George P. H. Styan, Jarkko Isotalo. International Statistical Review = Revue Internationale de Statistique, 81(1):152–153, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Polasek:2013:SBRe

[Pol13b] Wolfgang Polasek. Short book review: Multivariate Analysis for the Biobehavioral and Social Sciences: A Graphical Approach by Bruce L. Brown, Suzanne B. Hendrix, Dawson W. Hedges, Timothy B. Smith. International Statistical Review = Revue Internationale de Statistique, 81 (3):463–465, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Polasek:2013:SBRd

[Pol13c] Wolfgang Polasek. Short book review: Principles of Business Forecasting by Keith Ord, Robert Fildes. International Statistical Review = Revue Internationale de Statistique, 81(3):462–463, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Polasek:2013:SBRc

[Pol13d] Wolfgang Polasek. Short book review: Time Series Analysis and Its Applications: With R Examples, Third Edition by Robert H. Shumway, David S. Stoffer. International Statistical Review = Revue Internationale de Statistique, 81(2):323–325, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Postelnicu:1973:HTS

[Pos73] T. Postelnicu. Hardness and tensile strength of copper: Simple linear regression analysis. International Statistical Review = Revue Internationale de Statistique, 41(2):253– 258, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402842.

Postelnicu:1976:BRBc

[Pos76a] T. Postelnicu. Book review: An Introduction to Statistical Analysis for Economists by J. J. Thomas. International Statistical Review = Revue Internationale de Statistique, 44(2):305–306, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403298.

Postelnicu:1976:BRBd

[Pos76b] T. Postelnicu. Book review: Les Mathématiques de la Décision by Peter C. Fishburn. International Statistical Review = Revue Internationale de Statistique, 44(3): 377, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402357.

Postelnicu:1976:BRBe

[Pos76c] T. Postelnicu. Book review: Stochastic Analysis by D. G. Kendall; E. F. Harding. International Statistical Review = Revue Internationale de Statistique, 44(3):379, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402360.

Postelnicu:1976:BRBb

[Pos76d] T. Postelnicu. Book review: The Computation of Economic Equilibria by Herbert Scarf; Terje Hansen. International Statistical Review = Revue Internationale de Statistique, 44(2):305, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403297.

Postelnicu:1976:BRBa

[Pos76e] T. Postelnicu. Book review: Theoretical Statistics by D. R. Cox; D. V. Hinkley. International Statistical Review = Revue Internationale de Statistique, 44(2):301–302, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403292.

Postelnicu:1978:BRBb

[Pos78a] T. Postelnicu. Book review: A Programmed Text in Statistics by J. Hine; G. B. Wetherill. International Statistical Review = Revue Internationale de Statistique, 46 (1):118, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402519.

Postelnicu:1978:BRBd

[Pos78b] T. Postelnicu. Book review: Multivariate Analysis IV. Proceedings, 4th Internat. Symposium on Multivariate Analysis by P. R. Krishnaiah. International Statistical Review = Revue Internationale de Statistique, 46(3):320– 321, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402384.

Postelnicu:1978:BRBc

[Pos78c] T. Postelnicu. Book review: Prediction Analysis of Cross Classifications by D. K. Hildebrand; J. D. Laing; H. Rosenthal. International Statistical Review = Revue Internationale de Statistique, 46(3):319, December 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402381.

Postelnicu:1978:BRBa

[Pos78d] T. Postelnicu. Book review: The Great American Degree Machine by Douglas L. Adkins. International Statistical Review = Revue Internationale de Statistique, 46 (1):116, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402515.

Postelnicu:1978:BRBe

[Pos78e] T. Postelnicu. Book review: The Theory and Applications of Reliability with Emphasis on Bayesian and Nonparametric Methods by C. P. Tsokos; I. N. Shimi. International Statistical Review = Revue Internationale de Statistique, 46(3):322–323, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402389.

Postelnicu:1979:BRBc

[Pos79a] T. Postelnicu. Book review: Diffusion Processes and Related Topics in Biology by Luigi M. Ricciardi. International Statistical Review = Revue Internationale de Statistique, 47(1):91–92, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403223.

Postelnicu:1979:BRBa

[Pos79b] T. Postelnicu. Book review: Foundations of Inference in Survey Sampling by C. M. Cassel; C. E. Särndal; J. H. Wretman. International Statistical Review = Revue Internationale de Statistique, 47(1):86, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403214.

Postelnicu:1979:BRBb

[Pos79c] T. Postelnicu. Book review: Informatics and Medicine: An Advanced Course by P. L. Reichertz; G. Goos. International Statistical Review = Revue Internationale de Statistique, 47(1):91, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403222.

Postelnicu:1979:BRBd

[Pos79d] T. Postelnicu. Book review: Intermediate Business Statistics: Analysis of Variance by R. B. Miller; D. W. Wichern. International Statistical Review = Revue Internationale de Statistique, 47(2):233, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402577.

Postelnicu:1980:BRBa

[Pos80a] T. Postelnicu. Book review: Introduction to Statistics by D. Ary; L. C. Jacobs. International Statistical Review = Revue Internationale de Statistique, 48(2):229, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403155.

Postelnicu:1980:BRBd

[Pos80b] T. Postelnicu. Book review: Statistical Concepts and Methods by G. K. Bhattacharyya; R. A. Johnson. International Statistical Review = Revue Internationale de Statistique, 48(3):369, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402949.

Postelnicu:1980:BRBe

[Pos80c] T. Postelnicu. Book review: Statistical Decision Theory and Related Topics, II by S. S. Gupta; D. S. Moore. International Statistical Review = Revue Internationale de Statistique, 48(3):371, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402953.

Postelnicu:1980:BRBb

[Pos80d] T. Postelnicu. Book review: The Analysis of Survey Data by C. A. O’Muircheartaigh; C. Payne. International Statistical Review = Revue Internationale de Statistique, 48 (2):242, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403179.

Postelnicu:1980:BRBc

[Pos80e] T. Postelnicu. Book review: The Search for Oil. Some Statistical Methods and Techniques by D. B. Owen. International Statistical Review = Revue Internationale de Statistique, 48(2):242, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403180.

Pourahmadi:2002:GDM

[Pou02] Mohsen Pourahmadi. Graphical diagnostics for modeling unstructured covariance matrices. International Statistical Review = Revue Internationale de Statistique, 70(3):395– 417, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403864.

Power:2003:DIC

[Pow03] Tim Power. Development in Information and Communications Technology (ICT) statistics at the Australian Bureau of Statistics. International Statistical Review = Revue Internationale de Statistique, 71(1):59–67, April 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403874.

Puig:1998:LEC

[PP98] Antonio Puig and John Pullinger. Letter to the Editors: A comment on “The Creation of the Office for National Statistics” by John Pullinger. International Statistical Review = Revue Internationale de Statistique, 66(1): 133–134, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403660. See [Pul97].

Poleto:2011:IIP

[PPMS11] Frederico Z. Poleto, Carlos D. Paulino, Geert Molenberghs, and Julio M. Singer. Inferential implications of over-parametrization: A case study in incomplete categorical data. International Statistical Review = Revue Internationale de Statistique, 79(1):92–113, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41306166.

Petocz:2010:BSQ

[PR10] Peter Petocz and Anna Reid. On becoming a statistician — a qualitative view. International Statistical Review = Revue Internationale de Statistique, 78(2):271– 286, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919837.

Preece:1977:BRBb

[Pre77a] D. A. Preece. Book review: Applications of Statistics to Industrial Experimentation by C. Daniel. International Statistical Review = Revue Internationale de Statistique, 45(3):322–323, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402548.

Preece:1977:BRBa

[Pre77b] D. A. Preece. Book review: Weighing Designs for Chemistry, Medicine, Economics, Operations Research, Statistics by K. S. Banerjee. International Statistical Review = Revue Internationale de Statistique, 45(1):96, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403010.

Preece:1978:BRBa

[Pre78a] D. A. Preece. Book review: Design and Analysis of Experiments by D. C. Montgomery. International Statistical Review = Revue Internationale de Statistique, 46(1): 120–121, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402523.

Preece:1978:BRBb

[Pre78b] D. A. Preece. Book review: Recent Developments in Statistics. Proceedings, European Meeting of Statisticians, Grenoble, 1976 by J. R. Barra; F. Brodeau; G. Romier; B. van Cutsem. International Statistical Review = Revue Internationale de Statistique, 46(3):318, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402379.

Preece:1979:BRBa

[Pre79a] D. A. Preece. Book review: An Introduction to the Statistical Analysis of Data by T. W. Anderson; S. L. Schlove. International Statistical Review = Revue Internationale de Statistique, 47(2):230, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402571.

Preece:1979:BRBb

[Pre79b] D. A. Preece. Book review: Techniques de la Description Statistique: Méthodes et Logiciels pour l’Analyse des Grands Tableaux by L. Lebart; A. Morineau; N. Tabard.

International Statistical Review = Revue Internationale de Statistique, 47(2):231–232, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402574.

Preece:1980:BRBa

[Pre80a] D. A. Preece. Book review: Interactive Data Analysis: A Practical Primer by Donald R. McNeil. International Statistical Review = Revue Internationale de Statistique, 48(1):127–128, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402431.

Preece:1980:BRBb

[Pre80b] D. A. Preece. Book review: Linear Regression Analysis by G. A. F. Seber. International Statistical Review = Revue Internationale de Statistique, 48(1):132–133, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402437.

Preisler:1989:ATE

[Pre89] Haiganoush K. Preisler. Analysis of a toxicological experiment using a generalized linear model with nested random effects. International Statistical Review = Revue Internationale de Statistique, 57(2):145–159, August 1989. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403383.

Proietti:2011:ECF

[Pro11] Tommaso Proietti. Estimation of common factors under cross-sectional and temporal aggregation constraints. International Statistical Review = Revue Internationale de Statistique, 79(3):455–476, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305060.

Parr:1980:JB

[PS80] William C. Parr and William R. Schucany. The jackknife: A bibliography. International Statistical Review = Revue Internationale de Statistique, 48(1):73–78, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402406.

Pfeffermann:1985:RMG

[PS85] D. Pfeffermann and T. M. F. Smith. Regression models for grouped populations in cross-section surveys, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 53(1):37–59, April 1985. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402879.

Polasek:2013:SBRb

[PS13] Wolfgang Polasek and Evelina Shaparova. Short book review: Numerical Solution of Stochastic Differential Equations with Jumps in Finance by Eckhard Platen, Nicola Bruti-Liberati. International Statistical Review = Revue Internationale de Statistique, 81(2):311–313, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pace:2020:LRR

[PS20] Luigi Pace and Alessandra Salvan. Likelihood, replicability and Robbins’ confidence sequences. International Statistical Review = Revue Internationale de Statistique, 88(3):599–615, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Piccolo:2019:CCM

[PSI19] Domenico Piccolo, Rosaria Simone, and Maria Iannario. Cumulative and CUB models for rating data: a comparative analysis. International Statistical Review = Revue Internationale de Statistique, 87(2):207–236, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pesonen:2023:AF

[PSKL^{+}23] Henri Pesonen, Umberto Simola, Alvaro Köhn-Luque, Henri Vuollekoski, Xiaoran Lai, Arnoldo Frigessi, Samuel Kaski, David T. Frazier, Worapree Maneesoonthorn, Gael M. Martin, and Jukka Corander. ABC of the future. International Statistical Review = Revue Internationale de Statistique, 91(2):243–268, August 2023. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pukelsheim:1987:IIO

[Puk87] Friedrich Pukelsheim. Information increasing orderings in experimental design theory. International Statistical Review = Revue Internationale de Statistique, 55(2):203– 219, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403196.

Pukelsheim:2007:SBR

[Puk07] Friedrich Pukelsheim. Short book review: Optimum Experimental Designs, with SAS by Anthony Atkinson, Alexander Donev, Randall Tobias. International Statistical Review = Revue Internationale de Statistique, 75(3): 413, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509884.

Pullinger:1997:CON

[Pul97] John Pullinger. The creation of the Office for National Statistics. International Statistical Review = Revue Internationale de Statistique, 65(3):291–308, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/

1403373. See comment [PP98] and response [Pul98].

Pullinger:1998:R

[Pul98] John Pullinger. Response. International Statistical Review = Revue Internationale de Statistique, 66(1):135, April 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Pul97].

Puntanen:2007:SBRb

[Pun07a] Simo Puntanen. Short book review: Linear Models and Generalizations: Least Squares and Alternatives, 3rd Edition by C. Radhakrishna Rao, Helge Toutenburg, Shalabh, Christian Heumann (with Contributions by Michael Schomaker). International Statistical Review = Revue Internationale de Statistique, 75(3):438, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41509910.

Puntanen:2007:SBRa

[Pun07b] Simo Puntanen. Short book review: Matrix Algebra: Theory, Computations, and Applications in Statistics by James E. Gentle. International Statistical Review = Revue Internationale de Statistique, 75(3):435, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41509907.

Puntanen:2008:SBRc

[Pun08a] Simo Puntanen. Short book review: A Primer on Linear Models by John F. Monahan. International Statistical Review = Revue Internationale de Statistique, 76(3): 446, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919666.

Puntanen:2008:SBRb

[Pun08b] Simo Puntanen. Short book review: Linear Models in Statistics, Second Edition, by Alvin C. Rencher, G. Bruce Schaalje. International Statistical Review = Revue Internationale de Statistique, 76(3):445, December 2008. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919665.

Puntanen:2008:SBRd

[Pun08c] Simo Puntanen. Short book review: Modern Multivariate Statistical Techniques: Regression, Classification, and Manifold Learning by Alan Julian Izenman. International Statistical Review = Revue Internationale de Statistique, 76(3):462–463, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919682.

Puntanen:2008:SBRa

[Pun08d] Simo Puntanen. Short book review: Statistical Design by George Casella. International Statistical Review = Revue Internationale de Statistique, 76(3):443, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919663.

Puntanen:2009:SBR

[Pun09] Simo Puntanen. Short book review: Generalized, Linear, and Mixed Models, Second Edition, by Charles E. McCulloch, Shayle R. Searle, John M. Neuhaus. International Statistical Review = Revue Internationale de Statistique, 77(2):311, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919740.

Puntanen:2010:SBRb

[Pun10a] Simo Puntanen. Short book review: Chance Rules: An Informal Guide to Probability, Risk and Statistics, Second Edition, by Brian Everitt. International Statistical Review = Revue Internationale de Statistique, 78(1):145, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919810.

Puntanen:2010:SBRa

[Pun10b] Simo Puntanen. Short book review: Linear Regression Analysis: Theory and Computing by Xin Yan, Xiao Gang Su. International Statistical Review = Revue Internationale de Statistique, 78(1):144, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919809.

Puntanen:2010:SBRd

[Pun10c] Simo Puntanen. Short book review: Matrix Partial Orders, Shorted Operators and Applications by Sujit Kumar Mitra, P. Bhimasankaram, Saroj B. Malik. International Statistical Review = Revue Internationale de Statistique, 78(3):470, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919886.

Puntanen:2010:SBRc

[Pun10d] Simo Puntanen. Short book review: Multivariate Statistics: High-Dimensional and Large-Sample Approximations by Yasunori Fujikoshi, Vladimir V. Ulyanov, Ryoichi Shimizu. International Statistical Review = Revue Internationale de Statistique, 78(3):458–459, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919874.

Puntanen:2011:SBRa

[Pun11a] Simo Puntanen. Short book review: Creative Minds, Charmed Lives: Interviews at Institute for Mathematical Sciences, National University of Singapore edited by Yu Kiang Leong. International Statistical Review = Revue Internationale de Statistique, 79(1):138–139, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41306190.

Puntanen:2011:SBRb

[Pun11b] Simo Puntanen. Short book review: Inequalities: Theory of Majorization and Its Applications, Second Edition, by Albert W. Marshall, Ingram Olkin, Barry C. Arnold. International Statistical Review = Revue Internationale de Statistique, 79(2):293, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305043.

Puntanen:2011:SBRd

[Pun11c] Simo Puntanen. Short book review: Plane Answers to Complex Questions: The Theory of Linear Models, Fourth Edition, by Ronald Christensen. International Statistical Review = Revue Internationale de Statistique, 79(3):504– 505, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305085.

Puntanen:2011:SBRc

[Pun11d] Simo Puntanen. Short book review: Projection Matrices, Generalized Inverse Matrices, and Singular Value Decomposition by Haruo Yanai, Kei Takeuchi, Yoshio Takane. International Statistical Review = Revue Internationale de Statistique, 79(3):503–504, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305084.

Puntanen:2012:SBRa

[Pun12a] Simo Puntanen. Short book review: Advanced Econometric Theory by John S. Chipman. International Statistical Review = Revue Internationale de Statistique, 80(2):323, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puntanen:2012:SBRb

[Pun12b] Simo Puntanen. Short book review: Antieigenvalue Analysis: With Applications to Numerical Analysis, Wavelets, Statistics, Quantum Mechanics, Finance, and Optimization by Karl Gustafson. International Statistical Review = Revue Internationale de Statistique, 80(2):325–326, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puntanen:2012:SBRc

[Pun12c] Simo Puntanen. Short book review: Linear Algebra and Linear Models, Third Edition, by R. B. Bapat. International Statistical Review = Revue Internationale de Statistique, 80(2):347, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puntanen:2013:SBRc

[Pun13a] Simo Puntanen. Short book review: Handbook of Regression Analysis by Samprit Chatterjee, Jeffrey S. Simonoff. International Statistical Review = Revue Internationale de Statistique, 81(2):330–331, August 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puntanen:2013:SBRb

[Pun13b] Simo Puntanen. Short book review: Methods of Multivariate Analysis, Third Edition, by Alvin C. Rencher, William F. Christensen. International Statistical Review = Revue Internationale de Statistique, 81(2):328–329, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puntanen:2013:SBRa

[Pun13c] Simo Puntanen. Short book review: Regression Analysis by Example, Fifth Edition by Samprit Chatterjee, Ali S. Hadi. International Statistical Review = Revue Internationale de Statistique, 81(2):308, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Puri:1972:SAN

[Pur72] Madan L. Puri. Some aspects of nonparametric inference. International Statistical Review = Revue Internationale de Statistique, 40(3):299–327, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402470.

Pericchi:1991:RBC

[PW91] Luis Raúl Pericchi and Peter Walley. Robust Bayesian credible intervals and prior ignorance. International Statistical Review = Revue Internationale de Statistique, 59 (1):1–23, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403571.

Pajor:2024:HSG

[PWKO24] Anna Pajor, Justyna Wróblewska, Lukasz Kwiatkowski, and Jacek Osiewalski. Hybrid SV-GARCH, t-GARCH and Markov-switching covariance structures in VEC modelswhich is better from a predictive perspective? International Statistical Review = Revue Internationale de Statistique, 92(1):62–86, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pynnonen:2014:BRBb

[Pyn14a] Seppo Pynnönen. Book review: Current Topics in the Theory and Application of Latent Variable Models by Michael C. Edwards, Robert C. MacCallum (Editors). International Statistical Review = Revue Internationale de Statistique, 82(1):150–151, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Pynnonen:2014:BRBa

[Pyn14b] Seppo Pynnönen. Book review: Latent Markov Models for Longitudinal Data by Francesco Bartolucci, Alessio Farcomeni, Fulvia Pennoni. International Statistical Review = Revue Internationale de Statistique, 82(1):150, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Patone:2021:TEA

[PZ21] Martina Patone and Li-Chun Zhang. On two existing approaches to statistical analysis of social media data. International Statistical Review = Revue Internationale de Statistique, 89(1):54–71, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Qiwei:2020:BRT

[Qiw20] Yao Qiwei. Book review: Time Series: A First Course With Bootstrap Starter. Tucker S. McElroy and Dimitris N. Politis. CRC Press, 2020, xix + 566 pages, £74.99/ $99.95, hardcover ISBN 978-1-4398-7651-0. International Statistical Review = Revue Internationale de Statistique, 88(3):808–809, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Quine:1987:BDL

[QS87] M. P. Quine and E. Seneta. Bortkiewicz’s data and the law of small numbers. International Statistical Review = Revue Internationale de Statistique, 55(2):173–181, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403193.

Quinn:2021:FR

[Qui21] Barry G. Quinn. Fisher’s g revisited. International Statistical Review = Revue Internationale de Statistique, 89(2): 402–419, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ramalhoto:1983:SJL

[RAC83] M. F. Ramalhoto, J. A. Amaral, and M. Teresa Cochito. A survey of J. Little’s formula. International Statistical Review = Revue Internationale de Statistique, 51(3):255– 278, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402587.

Radcliffe:1975:BRB

[Rad75] J. Radcliffe. Book review: Branching Processes by K. B. Athreya; P. E. Ney. International Statistical Review = Revue Internationale de Statistique, 43(2):239–240, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402907.

Radcliffe:1977:BRB

[Rad77] J. Radcliffe. Book review: Branching Processes with Biological Applications by P. Jagers. International Statistical Review = Revue Internationale de Statistique, 45(2): 203, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403323.

Radermacher:1999:IGA

[Rad99] Walter Radermacher. Indicators, green accounting and environment statistics— information requirements for sustainable development. International Statistical Review = Revue Internationale de Statistique, 67(3):339–354, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403710.

Radaelli:2010:DSG

[Rad10] Paolo Radaelli. On the decomposition by subgroups of the Gini index and Zenga’s uniformity and inequality indexes. International Statistical Review = Revue Internationale de Statistique, 78(1):81–101, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919796.

Ramachandran:2003:MKD

[Ram03] R. Ramachandran. Measuring knowledge development and developing official statistics for the information age. International Statistical Review = Revue Internationale de Statistique, 71(1):83–107, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403876.

Rao:1973:PCM

[Rao73] C. R. Rao. Prasanta Chandra Mahalanobis, 1893–1972. International Statistical Review = Revue Internationale de Statistique, 41(2):301–302, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402866.

Rao:1977:OAS

[Rao77] T. J. Rao. Optimum allocation of sample size and prior distributions: a review. International Statistical Review = Revue Internationale de Statistique, 45(2):173–179, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403315.

Rao:1980:BRBa

[Rao80a] M. M. Rao. Book review: Probabilités et Statistique Mathématique by Mustafa Benyaklef. International Statistical Review = Revue Internationale de Statistique, 48(2): 229–230, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403157.

Rao:1980:BRBb

[Rao80b] M. M. Rao. Book review: Statistical Inference in Continuous Time Economic Models by A. R. Bergstrom. International Statistical Review = Revue Internationale de Statistique, 48(2):230–231, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403159.

Rao:1980:BRBc

[Rao80c] M. M. Rao. Book review: The Qualitative Theory of Optimal Processes by R. Gabasov; F. Kirillova. International Statistical Review = Revue Internationale de Statistique, 48(3):370, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402951.

Rao:1993:SIA

[Rao93] P. S. S. Sundar Rao. Statistical issues in assessing the toxic effects of Bhopal gas disaster. International Statistical Review = Revue Internationale de Statistique, 61(2):223– 229, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403625.

Rao:2005:MQS

[Rao05] J. N. K. Rao. On measuring the quality of survey estimates. International Statistical Review = Revue Internationale de Statistique, 73(2):241–244, August 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472663.

Rao:2006:SPS

[Rao06] C. Radhakrishna Rao. Statistical proofs of some matrix theorems. International Statistical Review = Revue Internationale de Statistique, 74(2):169–185, August 2006. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472701.

Rao:2019:MCL

[Rao19] J. N. K. Rao. My chancy life as a statistician. International Statistical Review = Revue Internationale de Statistique, 87(S1):S3–S9, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rayner:1990:STG

[RB90] J. C. W. Rayner and D. J. Best. Smooth tests of goodness of fit: An overview. International Statistical Review = Revue Internationale de Statistique, 58(1):9–17, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403470.

Ruggeri:2021:ISI

[RBB^{+}21] Fabrizio Ruggeri, Henrique Bolfarine, Jorge Luis Bazán, Reinaldo B. Arellano-Valle, Victor Hugo Lachos Davila, and Mário de Castro. 2021 International Statistical Institute Mahalanobis Award: a tribute to Heleno Bolfarine. International Statistical Review = Revue Internationale de Statistique, 89(3):435–446, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rodrigues:1985:GTP

[RBR85] Josemar Rodrigues, Heleno Bolfarine, and André Rogatko. A general theory of prediction in finite populations. International Statistical Review = Revue Internationale de Statistique, 53(3):239–254, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402888.

Roy:1993:VCC

[RC93] Roch Roy and Robert Cléroux. Vector cross-correlation in time series and applications. International Statistical Review = Revue Internationale de Statistique, 61(3):447– 464, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403755.

Reid:2015:SPS

[RC15] Nancy Reid and David R. Cox. On some principles of statistical inference. International Statistical Review = Revue Internationale de Statistique, 83(2):293–308, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Riani:2018:UPI

[RCA18] Marco Riani, Aldo Corbellini, and Anthony C. Atkinson. The use of prior information in very robust regression for fraud detection. International Statistical Review = Revue Internationale de Statistique, 86(2):205–218, August 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rakshit:2019:FKS

[RDM^{+}19] Suman Rakshit, Tilman Davies, M. Mehdi Moradi, Greg McSwiggan, Gopalan Nair, Jorge Mateu, and Adrian Baddeley. Fast kernel smoothing of point patterns on a large network using two-dimensional convolution. International Statistical Review = Revue Internationale de Statistique, 87(3):531–556, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ryan:2016:RMC

[RDMP16] Elizabeth G. Ryan, Christopher C. Drovandi, James M. McGree, and Anthony N. Pettitt. A review of modern computational algorithms for Bayesian optimal design. International Statistical Review = Revue Internationale de Statistique, 84(1):128–154, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Redfern:1974:DRP

[Red74] P. Redfern. The different roles of population censuses and interview surveys, particularly in the UK context. International Statistical Review = Revue Internationale de Statistique, 42(2):131–146, August 1974. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403076.

Redfern:2001:BME

[Red01] Philip Redfern. A Bayesian model for estimating census undercount, taking emigration data from foreign censuses. International Statistical Review = Revue Internationale de Statistique, 69(2):277–301, August 2001. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403816.

Reiersol:1976:RF

[Rei76] Olav Reiersøl. Ragnar Frisch, 1895–1973. International Statistical Review = Revue Internationale de Statistique, 44(1):178–179, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402976.

Reiter:2009:UMI

[Rei09] Jerome P. Reiter. Using multiple imputation to integrate and disseminate confidential microdata. International Statistical Review = Revue Internationale de Statistique, 77(2):179–195, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919721.

Reiter:2012:DSD

[Rei12] Jerome P. Reiter. Discussion: “Statistical disclosure risk: separating potential and harm” [MR3006147]. International Statistical Review = Revue Internationale de Statistique, 80(3):373–375, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Remon:1984:CPS

[Rém84] M. Rémon. On a concept of partial sufficiency: Lsufficiency. International Statistical Review = Revue Internationale de Statistique, 52(2):127–135, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403095.

Revesz:1973:BRB

[Rév73] P. Révész. Book review: A Book of Splines by A. Sard; S. Weintraub. International Statistical Review = Revue Internationale de Statistique, 41(1):141–142, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402800.

Revesz:1980:BRB

[Rév80] P. Révész. Book review: Statistics of Random Processes by R. S. Liptser; A. N. Shiryayev. International Statistical Review = Revue Internationale de Statistique, 48(3):371– 372, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402954.

Rivest:2023:USS

[RG23] Louis-Paul Rivest and Serigne Abib Gaye. Using survey sampling algorithms for exact inference in logistic regression. International Statistical Review = Revue Internationale de Statistique, 91(1):18–34, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Reiss:2017:MSF

[RGSO17] Philip T. Reiss, Jeff Goldsmith, Han Lin Shang, and R. Todd Ogden. Methods for scalar-on-function regression. International Statistical Review = Revue Internationale de Statistique, 85(2):228–249, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2005:DSR

[Ric05] Sylvia Richardson. Discussant: Sylvia Richardson. International Statistical Review = Revue Internationale de Statistique, 73(2):263–264, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2009:SBR

[Ric09] Alice Richardson. Short book review: Statistical Misconceptions by Schuyler W. Huck. International Statistical Review = Revue Internationale de Statistique, 77(2):322– 323, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919752.

Richardson:2010:SBRc

[Ric10a] Alice Richardson. Short book review: Basic Statistics: A Primer for the Biomedical Sciences, Fourth Edition, by Olive Jean Dunn, Virginia A. Clark. International Statistical Review = Revue Internationale de Statistique, 78 (3):453–454, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919870.

Richardson:2010:SBRa

[Ric10b] Alice Richardson. Short book review: Biological Data Mining edited by Jake Y. Chen, Stefano Lonardi. International Statistical Review = Revue Internationale de Statistique, 78(3):448–449, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919866.

Richardson:2010:SBRb

[Ric10c] Alice Richardson. Short book review: Nonparametric Statistics for Non-Statisticians: A Step-by-Step Approach by Gregory W. Corder, Dale I. Foreman. International Statistical Review = Revue Internationale de Statistique, 78(3):451–452, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919868.

Richardson:2011:SBRb

[Ric11a] Alice Richardson. Short book review: Logistic Regression: A Self-Learning Text, Third Edition, by David G. Kleinbaum, Mitchel Klein. International Statistical Review = Revue Internationale de Statistique, 79(2):296, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305046.

Richardson:2011:SBRc

[Ric11b] Alice Richardson. Short book review: Multiple Comparisons Using R by Frank Bretz, Torsten Hothorn, Peter Westfall. International Statistical Review = Revue Internationale de Statistique, 79(2):297, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305047.

Richardson:2011:SBRa

[Ric11c] Alice Richardson. Short book review: Statistics in Plain English, Third Edition, by Timothy C. Urdan. International Statistical Review = Revue Internationale de Statistique, 79(2):295, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305045.

Richardson:2013:SBRb

[Ric13a] Alice Richardson. Short book review: Applied Medical Statistics Using SAS by Geoff Der, Brian S. Everitt. International Statistical Review = Revue Internationale de Statistique, 81(3):476–477, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2013:SBRc

[Ric13b] Alice Richardson. Short book review: Mathematics Galore! The First Five Years of the St. Mark’s Institute of Mathematics by James Tanton. International Statistical Review = Revue Internationale de Statistique, 81(3):477– 478, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2013:SBRa

[Ric13c] Alice Richardson. Short book review: Teaching Elementary Statistics with JMP by Chris Olsen. International Statistical Review = Revue Internationale de Statistique, 81(2):329–330, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2014:BRC

[Ric14] Alice Richardson. Book review: A Concise Guide to Statistics by Hans-Michael Kaltenbach. International Statistical Review = Revue Internationale de Statistique, 82 (2):326, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2015:BRN

[Ric15] Alice M. Richardson. Book review: Nonparametric Statistics: A Step-by-Step Approach. International Statistical Review = Revue Internationale de Statistique, 83(1): 163–164, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2016:BRD

[Ric16] Alice M. Richardson. Book review: Dynamical Biostatistical Models. International Statistical Review = Revue Internationale de Statistique, 84(2):320–321, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2017:BRE

[Ric17a] Alice Richardson. Book review: Equivalence: Elizabeth L. Scott at Berkeley, A. L. Golbeck CRC Press, 2017, xxvi + 596 pages, $68, softcover ISBN: 978-1-4822-4944-6. International Statistical Review = Revue Internationale de Statistique, 85(2):374–375, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2017:BRS

[Ric17b] Alice Richardson. Book review: Statistics in Toxicology Using R. L. A. Hothorn Chapman and Hall/CRC, 2015, 234 pages, £61.99, hardback ISBN: 978-1-4987-0127-3. International Statistical Review = Revue Internationale de Statistique, 85(1):177, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Richardson:2018:BRA

[Ric18] Alice Richardson. Book review: Analysis of Biomarker Data, S. W. Looney and J. L. Hagan, John Wiley and Sons, 2015, 405 pages, $129.00, hardcover ISBN: 978-1-118-02755-4. International Statistical Review = Revue Internationale de Statistique, 86(3):600–601, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ridgway:2016:IDR

[Rid16] Jim Ridgway. Implications of the data revolution for statistics education. International Statistical Review = Revue Internationale de Statistique, 84(3):528–549, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rios:1976:JTI

[Rio76] G. Sixto Rios. Joaquín Torrens Ibern, 1909–1975. International Statistical Review = Revue Internationale de Statistique, 44(1):179, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402978.

Ripley:1983:CGR

[Rip83] B. D. Ripley. Computer generation of random variables: A tutorial. International Statistical Review = Revue Internationale de Statistique, 51(3):301–319, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402590.

Ripley:1984:SSD

[Rip84] B. D. Ripley. Spatial statistics: Developments 1980–3, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 52(2):141–150, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403097.

Riviere:2002:WMB

[Riv02] Pascal Rivière. What makes business statistics special? International Statistical Review = Revue Internationale de Statistique, 70(1):145–159, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403730.

Robinson:1991:SSR

[RJ91] Laurence D. Robinson and Nicholas P. Jewell. Some surprising results about covariate adjustment in logistic regression models. International Statistical Review = Revue Internationale de Statistique, 59(2):227–240, August 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403444.

Reluga:2023:SIL

[RLS23] Katarzyna Reluga, María-José Lombardía, and Stefan Sperlich. Simultaneous inference for linear mixed model parameters with an application to small area estimation.

International Statistical Review = Revue Internationale de Statistique, 91(2):193–217, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Redmond:1993:DAI

[RM93] Carol K. Redmond and Sati Mazumdar. Design, analysis and interpretation of long-term mortality studies of Coke oven workers. International Statistical Review = Revue Internationale de Statistique, 61(2):207–221, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403624.

Rao:1992:OCS

[RN92] J. N. K. Rao and A. K. Nigam. ’optimal’ controlled sampling: A unified approach. International Statistical Review = Revue Internationale de Statistique, 60(1):89–98, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403503.

Robinson:1993:ETM

[Rob93] P. M. Robinson. The estimation of transformation models. International Statistical Review = Revue Internationale de Statistique, 61(3):465–475, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403756.

Robert:2011:RKB

[Rob11a] Christian P. Robert. Reading Keynes’ Treatise on Probability. International Statistical Review = Revue Internationale de Statistique, 79(1):1–15, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41306162.

Robert:2011:SBRa

[Rob11b] Christian P. Robert. Short book review: A Comparison of the Bayesian and Frequentist Approaches to Estimation by Francisco J. Samaniego. International Statistical Review = Revue Internationale de Statistique, 79(1): 117–118, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306170.

Robert:2011:SBRd

[Rob11c] Christian P. Robert. Short book review: A Handbook of Statistical Analyses Using R, Second Edition, by Brian S. Everitt, Torsten Hothorn. International Statistical Review = Revue Internationale de Statistique, 79(2):276– 277, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305029.

Robert:2011:SBRc

[Rob11d] Christian P. Robert. Short book review: Bayesian Decision Analysis: Principles and Practice by Jim Q. Smith. International Statistical Review = Revue Internationale de Statistique, 79(2):272–273, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305025.

Robert:2011:SBRb

[Rob11e] Christian P. Robert. Short book review: Bayesian Model Selection and Statistical Modeling by Tomohiro Ando. International Statistical Review = Revue Internationale de Statistique, 79(1):120–121, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41306173.

Robert:2011:SBRg

[Rob11f] Christian P. Robert. Short book review: Numerical Analysis for Statisticians, Second Edition, by Kenneth Lange. International Statistical Review = Revue Internationale de Statistique, 79(3):502–503, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305083.

Robert:2011:SBRf

[Rob11g] Christian P. Robert. Short book review: The Foundations of Statistics: A Simulation-Based Approach by Shravan Vasishth, Michael Broe. International Statistical Review = Revue Internationale de Statistique, 79(3):486– 487, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305067.

Robert:2011:SBRe

[Rob11h] Christian P. Robert. Short book review: Time Series: Modeling, Computation, and Inference by Raquel Prado, Mike West. International Statistical Review = Revue Internationale de Statistique, 79(2):277–279, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41305030.

Robert:2012:SBRa

[Rob12a] Christian P. Robert. Short book review: Handbook of Fitting Statistical Distributions with R by Zaven A. Karian, Edward J. Dudewicz. International Statistical Review = Revue Internationale de Statistique, 80(1):177–178, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257173.

Robert:2012:SBRb

[Rob12b] Christian P. Robert. Short book review: The Theory That Would Not Die: How Bayes’ Rule Cracked the Enigma Code, Hunted Down Russian Submarines, and Emerged Triumphant from Two Centuries of Controversy by Sharon Bertsch McGrayne. International Statistical Review = Revue Internationale de Statistique, 80(1):178– 179, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/23257174.

Robert:2013:DCD

[Rob13] Christian P. Robert. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1):52–56, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Robert:2014:DBS

[Rob14] Christian P. Robert. Discussions on “A Brief Survey of Modern Optimization for Statisticians”: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(1):79–81, April 2014. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). See [LCZ14a].

Robert:2019:BRP

[Rob19] Christian P. Robert. Book review: Practical Bayesian Inference: a Primer for Physical Scientists, Coryn A. L. Bailer-Jones, Cambridge University Press, 2017, x + 295 pages, £69.99, hardcover ISBN 978-1-316-64221-4. International Statistical Review = Revue Internationale de Statistique, 87(1):182–184, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rodu:2021:BRT

[Rod21] Jordan Rodu. Book review: Text Analytics: Advances and Challenges, Domenica Fioredistella Iezzi, Damon Mayaffre and Michelangelo Misuraca Springer, 2020, xi + 302 pages, £95.50, eBook ISBN: 978-3-030-52680-1. International Statistical Review = Revue Internationale de Statistique, 89(2):429–431, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rohatgi:1976:BRBa

[Roh76a] V. K. Rohatgi. Book review: Generalized Inverses: Theory and Applications by Adi Ben-Israel; Thomas N. E. Greville. International Statistical Review = Revue Internationale de Statistique, 44(2):301, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403291.

Rohatgi:1976:BRBb

[Roh76b] V. K. Rohatgi. Book reviews: Introduction to Probability and Statistics. Part I: Probability by N. C. Giri; Introduction to Probability and Statistics. Part II: Statistics by N. C. Giri. International Statistical Review = Revue Internationale de Statistique, 44(2):302, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403293.

Rohatgi:1978:BRB

[Roh78] V. K. Rohatgi. Book review: Introduction to Minimax by V. F. Demyanov; V. N. Malozemov. International Statistical Review = Revue Internationale de Statistique, 46 (3):318–319, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402380.

Rohatgi:1980:BRB

[Roh80] V. K. Rohatgi. Book review: Wahrscheinlichkeitstheorie by P. Gänssler; W. Stute. International Statistical Review = Revue Internationale de Statistique, 48(1):123– 124, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402424.

Rohwer:2016:TCG

[Roh16] Götz Rohwer. Treatment and control groups in a dynamic setting. International Statistical Review = Revue Internationale de Statistique, 84(1):63–78, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ross:1987:ELR

[Ros87] W. H. Ross. The expectation of the likelihood ratio criterion. International Statistical Review = Revue Internationale de Statistique, 55(3):315–330, December 1987. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403408.

Rosenblad:2009:SBRa

[Ros09a] Andreas Rosenblad. Short book review: Applied Multivariate Statistics for the Social Sciences, Fifth Edition, by James P. Stevens. International Statistical Review = Revue Internationale de Statistique, 77(3):476, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919781.

Rosenblad:2009:SBRc

[Ros09b] Andreas Rosenblad. Short book review: Introduction to Meta-Analysis by Michael Borenstein, Larry V. Hedges, Julian P. T. Higgins, Hannah R. Rothstein. International Statistical Review = Revue Internationale de Statistique, 77(3):478–479, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919783.

Rosenblad:2009:SBRb

[Ros09c] Andreas Rosenblad. Short book review: Statistics Applied to Clinical Trials, Fourth Edition, by Ton J. Cleophas, Aeilko H. Zwinderman, Toine F. Cleophas, Eugene P. Cleophas. International Statistical Review = Revue Internationale de Statistique, 77(3):477–478, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919782.

Rosenblad:2010:SBRc

[Ros10a] Andreas Rosenblad. Short book review: Mathematics in Historical Context by Jeff Suzuki. International Statistical Review = Revue Internationale de Statistique, 78(3): 478, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919893.

Rosenblad:2010:SBRa

[Ros10b] Andreas Rosenblad. Short book review: Permutation Tests for Stochastic Ordering and ANOVA: Theory and Applications with R by Dario Basso, Fortunato Pesarin, Luigi Salmaso, Aldo Solari. International Statistical Review = Revue Internationale de Statistique, 78(3):446, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919864.

Rosenblad:2010:SBRb

[Ros10c] Andreas Rosenblad. Short book review: Sample Sizes for Clinical Trials by Steven A. Julious. International Statistical Review = Revue Internationale de Statistique, 78(3): 465, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919881.

Rosenblad:2014:BRS

[Ros14] Andreas Rosenblad. Book review: Statistical Modeling and Computation by Dirk P. Kroese, Joshua C. C. Chan. International Statistical Review = Revue Internationale de Statistique, 82(2):318–319, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rosenblad:2015:BRC

[Ros15a] Andreas Rosenblad. Book review: A Chronicle of Permutation Statistical Methods: 1920–2000, and Beyond. International Statistical Review = Revue Internationale de Statistique, 83(1):165–167, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rosenblad:2015:BRS

[Ros15b] Andreas Rosenblad. Book review: A Step-by-Step Approach to Using SAS^{(}R) for Factor Analysis and Structural Equation Modelling, Second Edition. International Statistical Review = Revue Internationale de Statistique, 83(2):325–326, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rosenblad:2019:BRI

[Ros19] Andreas Rosenblad. Book review: International Handbook of Research in Statistics Education, Dani Ben-Zvi, Katie Makar and Joan Garfield (Eds.), Springer, 2018, xix + 512 pages, £437.50/$229.00, hardcover ISBN: 978-3-319-66193-3. International Statistical Review = Revue Internationale de Statistique, 87(1):184–186, April 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Routledge:2007:SBR

[Rou07] Rick Routledge. Short book review: Modelling for Field Biologists and Other Interesting People by Hanna Kokko. International Statistical Review = Revue Internationale de Statistique, 75(3):437, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509909.

Royall:1986:MRC

[Roy86] Richard M. Royall. Model robust confidence intervals using maximum likelihood estimators. International Statistical Review = Revue Internationale de Statistique, 54(2): 221–226, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403146.

Royall:1994:SSA

[Roy94] Richard M. Royall. [sample surveys 1975–1990; an age of reconciliation?]: Discussion. International Statistical Review = Revue Internationale de Statistique, 62(1):19–21, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403540.

Roy:2014:BRC

[Roy14] Jason Roy. Book review: Causality: Statistical Perspectives and Applications by Carlo Berzuini, Philip Dawid, Luisa Bernardinelli. International Statistical Review = Revue Internationale de Statistique, 82(2):319–320, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rosenblatt-Roth:1979:BRBc

[RR79a] M. Rosenblatt-Roth. Book review: Groups, Representations and Characters by Victor E. Hill. International Statistical Review = Revue Internationale de Statistique, 47 (1):90, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403219.

Rosenblatt-Roth:1979:BRBa

[RR79b] M. Rosenblatt-Roth. Book review: Multiple Hypergeometric Functions and Applications by Harold Exton. International Statistical Review = Revue Internationale de Statistique, 47(1):86, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403215.

Rosenblatt-Roth:1979:BRBb

[RR79c] M. Rosenblatt-Roth. Book review: Probability, Induction and Statistics. The Art of Guessing by B. de Finetti. International Statistical Review = Revue Internationale de Statistique, 47(1):87, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403216.

Roelly:1990:CPM

[RR90] Sylvie Roelly and Alain Rouault. Construction et propriétés de martingales des branchements spatiaux interactifs. (French) [Construction and properties of martingales of interactive spatial connections]. International Statistical Review = Revue Internationale de Statistique, 58(2): 173–189, August 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403465.

Robert:2021:RBM

[RR21]  Christian P. Robert and Gareth Roberts.  Rao–

Blackwellisation in the Markov Chain Monte Carlo era. International Statistical Review = Revue Internationale de Statistique, 89(2):237–249, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ryabushkin:1979:PPM

[RS79] T. V. Ryabushkin and V. M. Simchera. Paul P. Maslow, 1902–1978. International Statistical Review = Revue Internationale de Statistique, 47(2):235, August 1979. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402581.

Rodriguez:2023:ILR

[RS23] Abel Rodríguez and Bruno Sansó. An interview with Luis Raúl Pericchi. International Statistical Review = Revue Internationale de Statistique, 91(1):1–17, April 2023. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Reis:2006:CSS

[RSG06] Edna A. Reis, Esther Salazar, and Dani Gamerman. Comparison of sampling schemes for dynamic linear models. International Statistical Review = Revue Internationale de Statistique, 74(2):203–214, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472703.

Rubin:2003:DMI

[Rub03] Donald B. Rubin. Discussion on multiple imputation. International Statistical Review = Revue Internationale de Statistique, 71(3):619–625, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403833.

Ruist:1972:BRB

[Rui72] Erik Ruist. Book review: Investment. The Study of an Economic Aggregate by Philip J. Lund. International Statistical Review = Revue Internationale de Statistique, 40 (2):254, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402774.

Ruiz:2016:BRB

[Rui16] Diego Andres Perez Ruiz. Book review: Bayesian Nonparametric Data Analysis. International Statistical Review = Revue Internationale de Statistique, 84(1):157– 158, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ruiz:2017:BRS

[Rui17] Diego Andrés Pérez Ruiz. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan, Richard McElreath CRC Press, 2015, 469 pages, £67.99, hardcover ISBN: 978-1-4822-5344-3. International Statistical Review = Revue Internationale de Statistique, 85(2):379–380, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rukhin:1995:ASC

[Ruk95] Andrew L. Rukhin. Admissibility: Survey of a concept in progress. International Statistical Review = Revue Internationale de Statistique, 63(1):95–115, April 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403779.

Rumsey:2002:DSL

[Rum02] Deborah J. Rumsey. Discussion: Statistical literacy: Implications for teaching, research, and practice. International Statistical Review = Revue Internationale de Statistique, 70(1):32–36, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403716.

Rustagi:1974:BRB

[Rus74] J. S. Rustagi. Book review: Introductory Statistics: A Service Course by A. H. Pollard. International Statistical Review = Revue Internationale de Statistique, 42(1):

109–110, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402693.

Russell:2010:SBR

[Rus10] Ken Russell. Short book review: The Art of Being a Scientist: A Guide for Graduate Students and their Mentors by Roel Snieder, Ken Larner. International Statistical Review = Revue Internationale de Statistique, 78(1):159, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919826.

Rosenblatt:1972:SMT

[RvA72] M. Rosenblatt and C. W. van Atta. “Statistical Models and Turbulence”: Comments on turbulence and a report on the satellite symposium. International Statistical Review = Revue Internationale de Statistique, 40(2):209– 214, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402762.

Raghavarao:2009:CMC

[RW09] Damaraju Raghavarao and James B. Wiley. Conjoint measurement with constraints on attribute levels: A mixture—amount model approach. International Statistical Review = Revue Internationale de Statistique, 77(2): 167–178, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919720.

Ryan:2013:SSD

[Rya13] Thomas P. Ryan. Sample Size Determination and Power. John Wiley, New York, NY, USA, 2013. ISBN 1-118-43760-8 (hardcover), 1-118-43920-1. xvi + 374 pp. LCCN

R853.C55 R93 2013. URL http://onlinelibrary.
wiley.com/book/10.1002/9781118439241.

Rydberg:2000:RSM

[Ryd00] Tina Hviid Rydberg. Realistic statistical modelling of financial data. International Statistical Review = Revue Internationale de Statistique, 68(3):233–258, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403412.

Reich:2021:RSC

[RYG^{+}21] Brian J. Reich, Shu Yang, Yawen Guan, Andrew B. Giffin, Matthew J. Miller, and Ana Rappold. A review of spatial causal inference methods for environmental and epidemiological applications. International Statistical Review = Revue Internationale de Statistique, 89(3):605–634, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rusch:2014:DFY

[RZ14] Thomas Rusch and Achim Zeileis. Discussions on fifty years of classification and regression trees: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(3):361–367, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Senoglu:2009:ACN

[ŞA09] Birdal Şenoğlu and Mubeccel Didem Avcıoğlu. Analysis of covariance with non-normal errors. International Statistical Review = Revue Internationale de Statistique, 77 (3):366–377, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919762.

Susmann:2022:TMD

[SAA22] Herbert Susmann, Monica Alexander, and Leontine Alkema. Temporal models for demographic and global health outcomes in multiple populations: Introducing a new framework to review and standardise documentation of model assumptions and facilitate model comparison. International Statistical Review = Revue Internationale de Statistique, 90(3):437–467, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Singpurwalla:2016:AAI

[SAG^{+}16] Nozer D. Singpurwalla, Barry C. Arnold, Joseph L. Gastwirth, Anna S. Gordon, and Hon Keung Tony Ng. Adversarial and amiable inference in medical diagnosis, reliability and survival analysis. International Statistical Review = Revue Internationale de Statistique, 84(3):390–412, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sahai:1979:BVC

[Sah79] Hardeo Sahai. A bibliography on variance components. International Statistical Review = Revue Internationale de Statistique, 47(2):177–222, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402568.

Saito:1979:RT

[Sai79] K. Saito. Ryoichi Tomoyasu, 1901–1978. International Statistical Review = Revue Internationale de Statistique, 47(1):97, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403230.

Saikkonen:2008:SBR

[Sai08] Pentti Saikkonen. Short book review: Introduction to Modern Time Series Analysis by Gebhard Kirchgässner, Jürgen Wolters. International Statistical Review = Revue Internationale de Statistique, 76(1):151–152, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919601.

Santow:1977:BRB

[San77] M. G. Santow. Book review: Historisch-Demographische Mitteilungen/Communications de Démographie Historique No. 3 by J. Kovacsics. International Statistical Review = Revue Internationale de Statistique, 45(3):325, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402553.

SantAnna:1991:LMR

[San91] Annibal Parracho Sant’Anna. The linear model with restricted explanatory variables. International Statistical Review = Revue Internationale de Statistique, 59(3):279– 285, December 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403688.

Sandland:1993:DGB

[San93] R. L. Sandland. Discussion of: George Box ’quality improvement: The new industrial revolution’. International Statistical Review = Revue Internationale de Statistique, 61(1):21–26, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403591.

Sankaran:2015:BRA

[San15] P. G. Sankaran. Book review: Age-Period-Cohort Models: Approaches and Analyses with Aggregate Data. International Statistical Review = Revue Internationale de Statistique, 83(2):335–336, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sankaran:2017:BRS

[San17] P. G. Sankaran. Book review: Survival Analysis in Genetics and Medicine, Jialiang Li and Shuangge Ma CRC Press, 2013, 381 pages, £78.99, hardcover ISBN: 978-1-4398-9311-1. International Statistical Review = Revue Internationale de Statistique, 85(2):377–379, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sangalli:2021:SRP

[San21] Laura M. Sangalli. Spatial regression with partial differential equation regularisation. International Statistical Review = Revue Internationale de Statistique, 89(3):505– 531, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sparks:1997:MPM

[SAP97] Ross Sparks, Allan Adolphson, and Aloke Phatak. Multivariate process monitoring using the dynamic biplot. International Statistical Review = Revue Internationale de Statistique, 65(3):325–349, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403375.

Saporta:2015:CJL

[Sap15] Gilbert Saporta. A conversation with Jean-Louis Bodin. International Statistical Review = Revue Internationale de Statistique, 83(1):2–16, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sarhan:1972:HMH

[Sar72a] A. E. Sarhan. Hassan Mohamed Hussein, 1910–1971. International Statistical Review = Revue Internationale de Statistique, 40(2):261, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402779.

Sarndal:1972:BRB

[Sär72b] Carl Erik Särndal. Book review: The Elements of Probability and Sampling by Frank A. Friday. International Statistical Review = Revue Internationale de Statistique, 40(3):394–395, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402481.

Sarndal:1972:SST

[Sär72c] Carl-Erik Särndal. Sample survey theory vs. general statistical theory: Estimation of the population mean. International Statistical Review = Revue Internationale de Statistique, 40(1):1–12, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402700.

Sarhan:1973:BRB

[Sar73] A. E. Sarhan. Book review: Order Statistics by H. A. David. International Statistical Review = Revue Internationale de Statistique, 41(2):292–293, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402854.

Sardy:2008:PRC

[Sar08] Sylvain Sardy. On the practice of rescaling covariates. International Statistical Review = Revue Internationale de Statistique, 76(2):285–297, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919617.

Sarndal:2011:TFS

[Sär11] Carl-Erik Särndal. Three factors to signal non-response bias with applications to categorical auxiliary variables.

International Statistical Review = Revue Internationale de Statistique, 79(2):233–254, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305023.

Satagopan:2015:BRC

[Sat15] Jaya M. Satagopan. Book review: Case-Control Studies. International Statistical Review = Revue Internationale de Statistique, 83(3):513–515, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sauvy:1976:RDS

[Sau76] Alfred Sauvy. Les responsabilités du statisticien dans la société. (French) [Responsibilities of the statistician in society]. International Statistical Review = Revue Internationale de Statistique, 44(3):311–316, December 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402346.

Sauvy:1980:PV

[Sau80] A. Sauvy. Paul Vincent, 1912–1979. International Statistical Review = Revue Internationale de Statistique, 48 (1):138, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402445.

Savage:1980:SE

[Sav80] I. Richard Savage. Statistical exercise. International Statistical Review = Revue Internationale de Statistique, 48 (1):103–116, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402410.

Saxena:1978:CMS

[Sax78] Ashok K. Saxena. Complex multivariate statistical analysis: an annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 46(2):209– 214, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402813.

Shenton:1975:DTE

[SB75] L. R. Shenton and K. O. Bowman. The development of techniques for the evaluation of sampling moments. International Statistical Review = Revue Internationale de Statistique, 43(3):317–334, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403115.

Speed:1987:FDM

[SB87] T. P. Speed and R. A. Bailey. Factorial dispersion models. International Statistical Review = Revue Internationale de Statistique, 55(3):261–277, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403405.

Shafer:2004:CFJ

[SB04] Glenn Shafer and Russell J. Bowater. Comments on “A Foundational Justification for a Weighted Likelihood Approach to Inference”. International Statistical Review = Revue Internationale de Statistique, 72(3):331–334, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/25472626.

Scosyrev:2017:EPE

[SB17] Emil Scosyrev and Fabrice Bancken. Expected precision of estimation and probability of ruling out a hypothesis based on a confidence interval. International Statistical Review = Revue Internationale de Statistique, 85(3):455– 472, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Schouten:2012:ECM

[SBB^{+}12] Barry Schouten, Jelke Bethlehem, Koen Beullens, Øyvin Kleven, Geert Loosveldt, Annemieke Luiten, Katja Rutar, Natalie Shlomo, and Chris Skinner. Evaluating, comparing, monitoring, and improving representativeness of survey response through R-indicators and partial Rindicators. International Statistical Review = Revue Internationale de Statistique, 80(3):382–399, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Seneta:2005:SST

[SC05] Eugene Seneta and John T. Chen. Simple stepwise tests of hypotheses and multiple comparisons. International Statistical Review = Revue Internationale de Statistique, 73 (1):21–34, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472638.

Swihart:2014:UFM

[SCC14] Bruce J. Swihart, Brian S. Caffo, and Ciprian M. Crainiceanu. A unifying framework for marginalised random-intercept models of correlated binary outcomes. International Statistical Review = Revue Internationale de Statistique, 82(2):275–295, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Schmetterer:1972:AR

[Sch72a] L. Schmetterer. Alfred Rényi, 1921–1970. International Statistical Review = Revue Internationale de Statistique, 40(1):110, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402724.

Schmetterer:1972:BRB

[Sch72b] L. Schmetterer. Book review: Stochastic Convergence by E. Lukacs. International Statistical Review = Revue Internationale de Statistique, 40(1):101–102, April 1972. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402715.

Schubnell:1972:BRB

[Sch72c] H. Schubnell. Book review: Demometrie by W. Winkler. International Statistical Review = Revue Internationale de Statistique, 40(1):105–106, April 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402719.

Schattan:1973:MCS

[Sch73] S. Schattan. Milton C. da Silva Rodrigues, 1904–1971. International Statistical Review = Revue Internationale de Statistique, 41(3):403, December 1973. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402642.

Schneeweiss:1978:ECC

[Sch78] H. Schneeweiss. An exchange of correspondence. Comment on “Bayesian Inference and the Optimality of Maximum Likelihood Estimation” (Internat. Statist. Rev. 45 (1977), no. 1, 9–11) by J. J. Higgins. International Statistical Review = Revue Internationale de Statistique, 46 (1):75–76, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402510. See [Hig77].

Schaich:1980:BRBa

[Sch80a] Eberhard Schaich. Book review: Personal Integrity and the Need for Data in the Social Sciences. Proceedings of a Symposium Held at Hasselby Slott, Stockholm by Tore Dalenius; Anders Klevmarken. International Statistical Review = Revue Internationale de Statistique, 48(3):369– 370, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402950.

Schaich:1980:BRBb

[Sch80b] Eberhard Schaich. Book review: Simulationstechnik und Simulationsmodelle in den Sozial- und Wirtschaftswissenschaften by Christoph E. Mandl. International Statistical Review = Revue Internationale de Statistique, 48(3): 372, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402955.

Schriever:1983:SOD

[Sch83] B. F. Schriever. Scaling of order dependent categorical variables with correspondence analysis. International Statistical Review = Revue Internationale de Statistique, 51 (3):225–238, December 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402585.

Schmetterer:1984:TWW

[Sch84a] L. Schmetterer. Tribute to Wilhelm Winkler at his 100th anniversary. International Statistical Review = Revue Internationale de Statistique, 52(3):227–228, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403044.

Schumacher:1984:TST

[Sch84b] Martin Schumacher. Two-sample tests of Cramér–von Mises- and Kolmogorov–Smirnov-type for randomly censored data. International Statistical Review = Revue Internationale de Statistique, 52(3):263–281, December 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403046.

Schwarz:1995:WRL

[Sch95] Norbert Schwarz. What respondents learn from questionnaires: The survey interview and the logic of conversation. International Statistical Review = Revue Internationale de Statistique, 63(2):153–168, August 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403610.

Scheaffer:1997:NPN

[Sch97] Richard L. Scheaffer. [new pedagogy and new content: The case of statistics]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(2): 156–158, August 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403339.

Schweder:2003:RFC

[Sch03] Tore Schweder. [reflections on fourteen cryptic issues concerning the nature of statistical inference]: Discussion. International Statistical Review = Revue Internationale de Statistique, 71(2):303–307, August 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403888.

Schappacher:2005:FB

[Sch05] Norbert Schappacher. Felix Bernstein. International Statistical Review = Revue Internationale de Statistique, 73 (1):3–7, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472636.

Schwarz:2007:CAS

[Sch07] Carl J. Schwarz. Computer-aided statistical instruction — multi-mediocre techno-trash? International Statistical Review = Revue Internationale de Statistique, 75(3):348– 354, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509875.

Schipper:2016:BRG

[Sch16a] Matthew Schipper. Book review: Genomic Clinical Trials and Predictive Medicine. International Statistical Review = Revue Internationale de Statistique, 84(1):155– 156, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Schwartzman:2016:LDG

[Sch16b] Armin Schwartzman. Lognormal distributions and geometric averages of symmetric positive definite matrices. International Statistical Review = Revue Internationale de Statistique, 84(3):456–486, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Salini:2016:RRR

[SCLR16] Silvia Salini, Andrea Cerioli, Fabrizio Laurini, and Marco Riani. Reliable robust regression diagnostics. International Statistical Review = Revue Internationale de Statistique, 84(1):99–127, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:1977:BRBb

[Sco77a] J. F. Scott. Book review: Sequential Methods in Statistics by G. Barrie Wetherill. International Statistical Review = Revue Internationale de Statistique, 45(2):205, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403328.

Scott:1977:BRBa

[Sco77b] J. F. Scott. Book review: The Interpretation of Multiple Observations by F. H. C. Marriott. International Statistical Review = Revue Internationale de Statistique, 45(2): 205, August 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403327.

Scott:2007:SBR

[Sco07] David J. Scott. Short book review: Organizing and Managing Your Research: A Practical Guide for Postgraduates by Renata Phelps, Kath Fisher, Allan H. Ellis. International Statistical Review = Revue Internationale de Statistique, 75(2):268–269, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41508483.

Scott:2009:SBRb

[Sco09a] David J. Scott. Short book review: Analysis of Integrated and Cointegrated Time Series with R, Second Edition, by Bernhard Pfaff. International Statistical Review = Revue Internationale de Statistique, 77(1):164–165, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919716.

Scott:2009:SBRa

[Sco09b] David J. Scott. Short book review: Applied Econometrics with R by Christian Kleiber, Achim Zeileis. International Statistical Review = Revue Internationale de Statistique, 77(1):164, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919715.

Scott:2010:SBRb

[Sco10a] David Scott. Short book review: Introducing Monte Carlo Methods with R by Christian P. Robert, George Casella. International Statistical Review = Revue Internationale de Statistique, 78(3):476–477, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919891.

Scott:2010:SBRa

[Sco10b] David J. Scott. Short book review: R Through Excel: A Spreadsheet Interface for Statistics, Data Analysis, and Graphics by Richard M. Heiberger, Erich Neuwirth. International Statistical Review = Revue Internationale de Statistique, 78(2):324, August 2010. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919849.

Scosyrev:2012:IET

[Sco12a] Emil Scosyrev. Interval estimation of treatment effects in randomized trials: when do confidence intervals have nominal coverage? International Statistical Review = Revue Internationale de Statistique, 80(3):439–451, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2012:SBR

[Sco12b] David Scott. Short book review: Mathematical Statistics with Resampling and R by Laura M. Chihara, Tim C. Hesterberg. International Statistical Review = Revue Internationale de Statistique, 80(3):472, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2013:SBRc

[Sco13a] David Scott. Short book review: Customer and Business Analytics: Applied Data Mining for Business Decision Making Using R by Daniel S. Putler, Robert E. Krider. International Statistical Review = Revue Internationale de Statistique, 81(2):328, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2013:SBRb

[Sco13b] David Scott. Short book review: Option Valuation: A First Course in Financial Mathematics by Hugo D. Junghenn. International Statistical Review = Revue Internationale de Statistique, 81(1):160–161, April 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2013:SBRa

[Sco13c] David Scott. Short book review: Statistical Computing in C++ and R by Randall L. Eubank, Ana Kupresanin. International Statistical Review = Revue Internationale de Statistique, 81(1):159–160, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Shang:2008:BMC

[SCW08] Junfeng Shang, Joseph E. Cavanaugh, and Farroll T. Wright. A Bayesian multiple comparison procedure for order-restricted mixed models. International Statistical Review = Revue Internationale de Statistique, 76(2):268– 284, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919616.

Shen:1999:DCD

[SCZZ99] Jianfa Shen, David Chu, Qingpu Zhang, and Weimin Zhang. Developing a census data system in China. International Statistical Review = Revue Internationale de Statistique, 67(2):173–186, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403397.

Steele:2011:AAM

[SD11] Fiona Steele and Gabriele B. Durrant. Alternative approaches to multilevel modelling of survey non-contact and refusal. International Statistical Review = Revue Internationale de Statistique, 79(1):70–91, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41306165.

Sanga:2011:R

[SDGD11a] Dimitri Sanga, Bakary Dosso, and Steve Gui-Diby. Rejoinder. International Statistical Review = Revue Internationale de Statistique, 79(3):333–335, December 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sanga:2011:TPT

[SDGD^{+}11b] Dimitri Sanga, Bakary Dosso, Steve Gui-Diby, Ben Kiregyera, and Mohamed Taamouti. Tracking progress towards statistical capacity building efforts: The African statistical development index [with discussions]. International Statistical Review = Revue Internationale de Statistique, 79(3):303–329, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41305054.

Sinner:2023:PWC

[SDT^{+}23] Corinne Sinner, Yves Dominicy, Julien Trufin, Wout Waterschoot, Patrick Weber, and Christophe Ley. From Pareto to Weibull — a constructive review of distributions on R^{+}. International Statistical Review = Revue Internationale de Statistique, 91(1):35–54, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Seber:1992:REA

[Seb92] G. A. F. Seber. A review of estimating animal abundance II. International Statistical Review = Revue Internationale de Statistique, 60(2):129–166, August 1992. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403646.

Sen:1972:RSP

[Sen72] Pranab Kumar Sen. Robust statistical procedures in problems of linear regression with special reference to quantitative bioassays, II. International Statistical Review = Revue Internationale de Statistique, 40(2):161–172, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402759.

Sen:1976:AOE

[Sen76] Pranab Kumar Sen. Asymptotically optimal estimators of general parameters in randomized response models. International Statistical Review = Revue Internationale de Statistique, 44(2):223–224, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403281.

Sen:1984:DPK

[Sen84] P. K. Sen. Discussion of paper by A. K. Adhikari, A. Chaudhuri and K. Vijayan. International Statistical Review = Revue Internationale de Statistique, 52(2):120– 122, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403091.

Seneta:1996:MBC

[Sen96] E. Seneta. Markov and the birth of chain dependence theory. International Statistical Review = Revue Internationale de Statistique, 64(3):255–263, December 1996. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403785.

Seneta:1998:JBC

[Sen98] E. Seneta. I. J. Bienaymé [1796–1878]: Criticality, inequality, and internationalization. International Statistical Review = Revue Internationale de Statistique, 66(3): 291–301, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403518.

Seneta:2003:SRF

[Sen03] Eugene Seneta. Statistical regularity and free will: L. A. J. Quetelet and P. A. Nekrasov. International Statistical Review = Revue Internationale de Statistique, 71(2):319– 334, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403892.

Seneta:2005:SC

[Sen05] Eugene Seneta. Statisticians of the centuries. International Statistical Review = Revue Internationale de Statistique, 73(1):1, April 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/25472635.

Seneta:2009:KPR

[Sen09] Eugene Seneta. Karl Pearson in Russian contexts. International Statistical Review = Revue Internationale de Statistique, 77(1):118–146, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919695.

Sen:2011:SBRb

[Sen11a] Pranab K. Sen. Short book review: Introduction to Nonparametric Estimation by Alexandre B. Tsybakov. International Statistical Review = Revue Internationale de Statistique, 79(2):291–292, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305042.

Sen:2011:SBRc

[Sen11b] Pranab K. Sen. Short book review: Medical Applications of Finite Mixture Models by Peter Schlattmann. International Statistical Review = Revue Internationale de Statistique, 79(2):294, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305044.

Sen:2011:SBRa

[Sen11c] Pranab K. Sen. Short book review: Medical Uses of Statistics, Third Edition, edited by John C. Bailar III, David C. Hoaglin. International Statistical Review = Revue Internationale de Statistique, 79(2):290–291, August 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/41305041.

Seneta:2014:BSP

[Sen14a] Eugene Seneta. On the bicentenary in St. Petersburg of Jacob Bernoulli’s Theorem. International Statistical Review = Revue Internationale de Statistique, 82(1):17–26, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sengupta:2014:BRBb

[Sen14b] Arindam Sengupta. Book review: A First Course in Machine Learning by Simon Rogers and Mark Girolami. International Statistical Review = Revue Internationale de Statistique, 82(1):158–159, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sengupta:2014:BRBa

[Sen14c] Arindam Sengupta. Book review: A Tiny Handbook of R by Mike Allerhand. International Statistical Review = Revue Internationale de Statistique, 82(1):156, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sengupta:2014:BRA

[Sen14d] Arindam Sengupta. Book review: Applied Time Series Analysis by Wayne A. Woodward, Henry L. Gray, Alan C. Elliott. International Statistical Review = Revue Internationale de Statistique, 82(2):322–323, August 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sengupta:2015:BRI

[Sen15] Arindam Sengupta. Book review: Introduction to Probability. International Statistical Review = Revue Internationale de Statistique, 83(3):518–519, December 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sengupta:2023:BRS

[Sen23] Arindam Sengupta. Book review: Social Networks: Modelling and Analysis, Niyati Aggrawal and Adarsh Anand CRC Press, 2022, xvii + 235 pages, hardcover ($170); ebook ($44.21). ISBN: 978-0-367-54,139-2 (hardcover), 978-0-367-54,173-6 (paperback), 978-1-003-08806-6 (e-book). International Statistical Review = Revue Internationale de Statistique, 91(1):162–164, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Serdar:1978:VV

[Ser78] Vladimir Serdar. Vladimir Vranić, 1896–1976. International Statistical Review = Revue Internationale de Statistique, 46(1):127, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402531.

Santos:2010:CCB

[SFG10] Thiago R. Santos, Glaura C. Franco, and Dani Gamerman. Comparison of classical and Bayesian approaches for intervention analysis. International Statistical Review = Revue Internationale de Statistique, 78(2):218– 239, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919834.

Santos-Fernandez:2024:ITN

[SFVP^{+}24] Edgar Santos-Fernandez, Julie Vercelloni, Aiden Price, Grace Heron, Bryce Christensen, Erin E. Peterson, and Kerrie Mengersen. Increasing trust in new data sources: Crowdsourcing image classification for ecology. International Statistical Review = Revue Internationale de Statistique, 92(1):43–61, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Stavig:1977:CMM

[SG77] Gordon R. Stavig and Jean D. Gibbons. Comparing the mean and the median as measures of centrality. International Statistical Review = Revue Internationale de Statistique, 45(1):63–70, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403004.

Sheehan:2002:BMM

[SGLS02] N. A. Sheehan, B. Gulbrandtsen, M. S. Lund, and D. A. Sorensen. Bayesian MCMC mapping of quantitative trait loci in a half-sib design: A graphical model perspective. International Statistical Review = Revue Internationale de Statistique, 70(2):241–267, August 2002. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403909.

Schulz:2021:MPM

[SGM21] Juliana Schulz, Christian Genest, and Mhamed Mesfioui. A multivariate Poisson model based on comonotonic shocks. International Statistical Review = Revue Internationale de Statistique, 89(2):323–348, August 2021. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Steel:1996:AAA

[SH96] D. G. Steel and D. Holt. Analysing and adjusting aggregation effects: The ecological fallacy revisited. International Statistical Review = Revue Internationale de Statistique, 64(1):39–60, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403423.

Schiavo:2000:TMY

[SH00] Rosa A. Schiavo and David J. Hand. Ten more years of error rate research. International Statistical Review = Revue Internationale de Statistique, 68(3):295–310, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403415.

Schweder:2013:DCD

[SH13] Tore Schweder and Nils Lid Hjort. Discussions on “Confidence Distribution, the Frequentist Distribution Estimator of a Parameter: A Review”: Discussion [MR3047496]. International Statistical Review = Revue Internationale de Statistique, 81(1):56–68, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Stefan:2021:BES

[SH21] Marius Stefan and Michael Hidiroglou. Benchmarked estimators for a small area mean under a onefold nested regression model. International Statistical Review = Revue Internationale de Statistique, 89(1):108–131, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Stefan:2023:BVP

[SH23] Marius Stefan and Michael A. Hidiroglou. A bootstrap variance procedure for the generalised regression estimator. International Statistical Review = Revue Internationale de Statistique, 91(2):294–317, August 2023. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Shaban:1980:CPP

[Sha80] S. A. Shaban. Change point problem and two-phase regression: an annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 48 (1):83–93, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402408.

Shafer:1982:DPD

[Sha82] Glenn Shafer. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):18–19, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402452.

Shafer:1985:CP

[Sha85a] Glenn Shafer. Conditional probability. International Statistical Review = Revue Internationale de Statistique, 53 (3):261–275, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402890.

Shafer:1985:CPR

[Sha85b] Glenn Shafer. [conditional probability]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 53(3):276–277, December 1985. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402892.

Shapiro:1988:TUT

[Sha88] A. Shapiro. Towards a unified theory of inequality constrained testing in multivariate analysis. International Statistical Review = Revue Internationale de Statistique, 56(1):49–62, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403361.

Shaochuan:2023:SBM

[Sha23] Lu Shaochuan. Scalable Bayesian multiple changepoint detection via auxiliary uniformisation. International Statistical Review = Revue Internationale de Statistique, 91 (1):88–113, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sheehan:2000:AMC

[She00] N. A. Sheehan. On the application of Markov chain Monte Carlo methods to genetic analyses on complex pedigrees. International Statistical Review = Revue Internationale de Statistique, 68(1):83–110, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403559.

Shen:2006:EUL

[She06] Jianfa Shen. Estimating urbanization levels in Chinese provinces in 1982–2000. International Statistical Review = Revue Internationale de Statistique, 74(1):89–107, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
25472693.

Singh:2001:EMR

[SHGK01] M. P. Singh, M. A. Hidiroglou, J. G. Gambino, and M. S. Kovačević. Estimation methods and related systems at Statistics Canada. International Statistical Review = Revue Internationale de Statistique, 69(3):461–485, December 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403457.

Skinner:1994:MFS

[SHH94] C. J. Skinner, D. J. Holmes, and D. Holt. Multiple frame sampling for multivariate stratification. International Statistical Review = Revue Internationale de Statistique, 62 (3):333–347, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403765.

Shipley:1978:HS

[Shi78] Thorne Shipley. Hermann Von Schelling, 1901–1977. International Statistical Review = Revue Internationale de Statistique, 46(2):234–235, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402827.

Shiryayev:1981:MRD

[Shi81] A. N. Shiryayev. Martingales: recent developments, results and applications. International Statistical Review = Revue Internationale de Statistique, 49(3):199–233, December 1981. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402605.

Stigler:2005:SWN

[SHL05] Stephen M. Stigler, Moon Y. Huh, and Jae Chang Lee. Statistics and the wealth of nations. International Statistical Review = Revue Internationale de Statistique, 73(2): 223–226, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472661.

Shlomo:2007:SDC

[Shl07] Natalie Shlomo. Statistical disclosure control methods for census frequency tables. International Statistical Review = Revue Internationale de Statistique, 75(2):199– 217, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508461.

Shlomo:2011:D

[Shl11] Natalie Shlomo. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):194– 197, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scealy:2022:DON

[SHLW22] Janice L. Scealy, David Heslop, Jia Liu, and Andrew T. A. Wood. Directions old and new: Palaeomagnetism and Fisher (1953) meet modern statistics. International Statistical Review = Revue Internationale de Statistique, 90 (2):237–258, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Shoesmith:1985:NBA

[Sho85] E. Shoesmith. Nicholas Bernoulli and the argument for Divine Providence. International Statistical Review = Revue Internationale de Statistique, 53(3):255–259, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402889.

Siegmund:1988:CSC

[Sie88] David Siegmund. Confidence sets in change-point problems. International Statistical Review = Revue Internationale de Statistique, 56(1):31–48, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403360.

Sikov:2018:BRA

[Sik18] Anna Sikov. A brief review of approaches to non-ignorable non-response. International Statistical Review = Revue Internationale de Statistique, 86(3):415–441, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Silvey:1973:BRB

[Sil73] S. D. Silvey. Book review: The Theory of Statistical Inference by S. Zacks. International Statistical Review = Revue Internationale de Statistique, 41(2):298, August 1973. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402863.

Silverman:1982:DPD

[Sil82] B. W. Silverman. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):19–20, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402453.

Simonoff:1998:TSS

[Sim98] Jeffrey S. Simonoff. Three sides of smoothing: Categorical data smoothing, nonparametric regression, and density estimation. International Statistical Review = Revue Internationale de Statistique, 66(2):137–156, August 1998. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403487.

Singer:1995:WRL

[Sin95] Eleanor Singer. [what respondents learn from questionnaires: The survey interview and the logic of conversation]: Discussion. International Statistical Review = Revue Internationale de Statistique, 63(2):169–172, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403611.

Singpurwalla:2002:CCM

[Sin02a] Nozer D. Singpurwalla. On causality and causal mechanisms. Comment on Dennis Lindley’s “Seeing and Doing: The Concept of Causation”. International Statistical Review = Revue Internationale de Statistique, 70(2):198– 206, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403903. See [Lin02].

Singpurwalla:2002:RJP

[Sin02b] Nozer D. Singpurwalla. Rejoinder to Judea Pearl. International Statistical Review = Revue Internationale de Statistique, 70(2):212–214, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403906.

Singpurwalla:2002:SCEa

[Sin02c] Nozer D. Singpurwalla. Some cracks in the empire of chance (flaws in the foundations of reliability). International Statistical Review = Revue Internationale de Statistique, 70(1):53–65, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403722.

Singpurwalla:2002:SCEb

[Sin02d] Nozer D. Singpurwalla. [some cracks in the empire of chance (flaws in the foundations of reliability)]: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 70(1):77–78, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403726.

Silverman:1989:FJH

[SJ89] B. W. Silverman and M. C. Jones. E. Fix and J. L. Hodges (1951): “An Important Contribution to Nonparametric Discriminant Analysis and Density Estimation”: Commentary on Fix and Hodges (1951). International Statistical Review = Revue Internationale de Statistique, 57 (3):233–238, December 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403796.

Segal:1995:SIE

[SJFM95] Mark R. Segal, Ian R. James, Martyn A. H. French, and Simon A. Mallal. Statistical issues in the evaluation of markers of HIV progression. International Statistical Review = Revue Internationale de Statistique, 63(2):179– 197, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403613.

Skinner:2012:RSD

[Ski12a] Chris Skinner. Rejoinder: “Statistical disclosure risk: separating potential and harm” [MR3006148; MR3006149; 3006150; MR3006147]. International Statistical Review = Revue Internationale de Statistique, 80(3):379–381, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Skinner:2012:SDR

[Ski12b] Chris Skinner. Statistical disclosure risk: separating potential and harm. International Statistical Review = Revue Internationale de Statistique, 80(3):349–368, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Skinner:2019:ACD

[Ski19] Chris Skinner. Analysis of categorical data for complex surveys. International Statistical Review = Revue Internationale de Statistique, 87(S1):S64–S78, May 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Shenavari:2018:MTA

[SKK18] Zahra Shenavari and Mahmood Kharrati-Kopaei. A method for testing additivity in unreplicated two-way layouts based on combining multiple interaction tests. International Statistical Review = Revue Internationale de Statistique, 86(3):469–487, December 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Skovgaard:1986:NDC

[Sko86a] Ib Skovgaard. A note on the differentiation of cumulants of log likelihood derivatives. International Statistical Review = Revue Internationale de Statistique, 54(1):29–32, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403256.

Skovgaard:1986:MEE

[Sko86b] Ib M. Skovgaard. On multivariate Edgeworth expansions. International Statistical Review = Revue Internationale de Statistique, 54(2):169–186, August 1986. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403142.

Skotarczak:2012:SBR

[Sko12] Ewa Skotarczak. Short book review: Introduction to Statistical Data Analysis for the Life Sciences by Claus Thorn Ekstrøm, Helle Sørensen. International Statistical Review = Revue Internationale de Statistique, 80(1):176, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
23257172.

Suesse:2013:MSR

[SL13] Thomas Suesse and Ivy Liu. Modelling strategies for repeated multiple response data. International Statistical Review = Revue Internationale de Statistique, 81(2):230– 248, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sloboda:2023:BRE

[Slo23] Brian W. Sloboda. Book review: The Effect: an Introduction to Research Design and Causality, Nick Huntington-Klein, Chapman & Hall/CRC, 2022, xiv + 620 pages, $39.95, paperback. ISBN: 978-1-032-12578-7. International Statistical Review = Revue Internationale de Statistique, 91(2):343–345, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sloboda:2024:BRM

[Slo24] Brian W. Sloboda. Book review: Multidimensional Stationary Time Series Dimension Reduction and Prediction, Marianna Bolla, Tamás Szabados, Routledge, 2023, xiv + 318 pages, $59.95, paperback ISBN: 978-0-367-61970-1. International Statistical Review = Revue Internationale de Statistique, 92(1):130–132, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scharf:2022:CFI

[SLWH22] Henry R. Scharf, Xinyi Lu, Perry J. Williams, and Mevin B. Hooten. Constructing flexible, identifiable and interpretable statistical models for binary data. International Statistical Review = Revue Internationale de Statistique, 90(2):328–345, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Son:2023:PAF

[SLY23] Won Son, Johan Lim, and Donghyeon Yu. Path algorithms for fused lasso signal approximator with application to COVID-19 spread in Korea. International Statistical Review = Revue Internationale de Statistique, 91(2): 218–242, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sutradhar:1989:TWA

[SM89] Brajendra C. Sutradhar and Ian B. MacNeill. Two-way analysis of variance for stationary periodic time series. International Statistical Review = Revue Internationale de Statistique, 57(2):169–182, August 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403385.

Seillier-Moiseiwitsch:1993:SPF

[SM93] Françoise Seillier-Moiseiwitsch. Sequential probability forecasts and the probability integral transform. International Statistical Review = Revue Internationale de Statistique, 61(3):395–408, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403751.

Smithson:2014:GLM

[SM14] Michael. Smithson and Edgar C. Merkle. Generalized Linear Models for Categorical and Continuous Limited Dependent Variables. Chapman and Hall/CRC statistics in the social and behavioral sciences series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-5173-9 (hardcover). xxiii + 284 pp. LCCN QA279 .S634 2014.

Song:2019:IDT

[SM19] Yu Song and Reza Modarres. Interpoint distance test of homogeneity for multivariate mixture models. International Statistical Review = Revue Internationale de Statistique, 87(3):613–638, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Small:1988:TSA

[Sma88] Christopher G. Small. Techniques of shape analysis on sets of points. International Statistical Review = Revue Internationale de Statistique, 56(3):243–257, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403352.

Small:1990:SMM

[Sma90] Christopher G. Small. A survey of multidimensional medians. International Statistical Review = Revue Internationale de Statistique, 58(3):263–277, December 1990. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403809.

Smith:1972:SSP

[Smi72] V. Kerry Smith. The small sample properties of selected econometric estimators in the context of alternative macro-models. International Statistical Review = Revue Internationale de Statistique, 40(3):263–268, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402465.

Smith:1973:TRS

[Smi73] V. Kerry Smith. A tabular review of sampling studies with problems of autocorrelation and distributed lag models. International Statistical Review = Revue Internationale de Statistique, 41(3):351–355, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402623.

Smith:1975:BRB

[Smi75] Cedric A. B. Smith. Book review: Bayesian Inference in Statistical Analysis by G. E. P. Box; G. C. Tiao. International Statistical Review = Revue Internationale de Statistique, 43(2):242–243, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402913.

Smith:1977:BRB

[Smi77] Cedric A. B. Smith. Book review: Discriminant Analysis by Peter A. Lachenbruch. International Statistical Review = Revue Internationale de Statistique, 45(1):98–99, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403015.

Smith:1980:BRB

[Smi80] Cedric A. B. Smith. Book review: Combinatorics with Emphasis on the Theory of Graphs by Jack E. Graver; Mark E. Watkins. International Statistical Review = Revue Internationale de Statistique, 48(1):124, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402425.

Smith:1982:DPD

[Smi82] Cedric A. B. Smith. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):21–23, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402454.

Smith:1992:CCS

[Smi92] J. Q. Smith. A comparison of the characteristics of some Bayesian forecasting models. International Statistical Review = Revue Internationale de Statistique, 60(1):75–87, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403502.

Smith:1994:SSAa

[Smi94a] T. M. F. Smith. Sample surveys 1975–1990; an age of reconciliation? International Statistical Review = Revue Internationale de Statistique, 62(1):5–19, April 1994. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403539.

Smith:1994:SSAb

[Smi94b] T. M. F. Smith. [sample surveys 1975–1990; an age of reconciliation?]: Reply to the discussion by T. M. F. Smith. International Statistical Review = Revue Internationale de Statistique, 62(1):32–34, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403543.

Smith:1999:STE

[Smi99] T. M. F. Smith. [statistical thinking in empirical enquiry]: Discussion. International Statistical Review = Revue Internationale de Statistique, 67(3):248–250, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403700.

Smith:2001:PPI

[Smi01] Adrian F. M. Smith. Public policy issues as a route to statistical awareness. International Statistical Review = Revue Internationale de Statistique, 69(1):17–20, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403526.

Smith:2016:DPI

[Smi16] Luke B. Smith. Discussions on posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Discussion. International Statistical Review = Revue Internationale de Statistique, 84(3):359–362, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Smith:2020:RIA

[Smi20] Duncan Smith. Re-identification in the absence of common variables for matching. International Statistical Review = Revue Internationale de Statistique, 88(2):354– 379, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Smith:2021:ESE

[Smi21] Paul A. Smith. Estimating sampling errors in consumer price indices. International Statistical Review = Revue Internationale de Statistique, 89(3):481–504, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Singh:1979:SSK

[SMW79] Parmeet Singh, Julius Mutia, and Ken Williams. Social statistics in Kenya. International Statistical Review = Revue Internationale de Statistique, 47(1):25–30, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403203.

Snee:1999:DDU

[Sne99a] Ronald D. Snee. Discussion: Development and use of statistical thinking: A new era. International Statistical Review = Revue Internationale de Statistique, 67(3):255– 258, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403703.

Snee:1999:DSM

[Sne99b] Ronald D. Snee. Discussion: Statisticians must develop data-based management and improvement systems as well as create measurement systems. International Statistical Review = Revue Internationale de Statistique, 67(2):139– 144, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403393.

Snell:2002:DHD

[Sne02] J. Laurie Snell. Discussion: But how do you teach it? International Statistical Review = Revue Internationale de Statistique, 70(1):45–46, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403720.

Som:1976:BRBb

[Som76a] R. K. Som. Book review: Fundamentals of Statistics by A. M. Goon; M. K. Gupta; B. Dasgupta. International Statistical Review = Revue Internationale de Statistique, 44(3):377–378, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402358.

Som:1976:BRBc

[Som76b] R. K. Som. Book review: The Sample Survey: Theory and Practice by Donald P. Warwick; Charles A. Lininger. International Statistical Review = Revue Internationale de Statistique, 44(3):380, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402362.

Som:1976:BRBa

[Som76c] R. K. Som. Book review: Vital Statistics by Michael Orkin; Richard Drogin. International Statistical Review = Revue Internationale de Statistique, 44(2):304, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403295.

Som:1977:BRBb

[Som77a] R. K. Som. Book review: Basic Ideas of Scientific Sampling by A. Stuart. International Statistical Review = Revue Internationale de Statistique, 45(3):327, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402557.

Som:1977:BRBa

[Som77b] R. K. Som. Book reviews: Statistics for Beginners by S. O. Adamu; T. L. Johnson; Statistics for Beginners: Workbook by S. O. Adamu; T. L. Johnson. International Statistical Review = Revue Internationale de Statistique, 45(1):95, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403008.

Som:1978:BRBe

[Som78a] R. K. Som. Book review: Case Studies in Sample Design by A. C. Rosander. International Statistical Review = Revue Internationale de Statistique, 46(3):322, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402388.

Som:1978:BRBf

[Som78b] R. K. Som. Book review: Elementary Statistical Concepts by Ronald E. Walpole. International Statistical Review = Revue Internationale de Statistique, 46(3):323, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402390.

Som:1978:BRBg

[Som78c] R. K. Som. Book review: Introductory Statistics for the Behavioral Sciences by R. K. Young; D. J. Veldman. International Statistical Review = Revue Internationale de Statistique, 46(3):323, December 1978. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402391.

Som:1978:BRBb

[Som78d] R. K. Som. Book review: Population and Land in World Agriculture: Recent Trends and Relationships by J. Kumar. International Statistical Review = Revue Internationale de Statistique, 46(1):118–119, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402520.

Som:1978:BRBc

[Som78e] R. K. Som. Book review: Population Growth Estimation: A Handbook of Vital Statistics and Measurement by E. S. Marks; W. Seltzer; K. J. Krotki. International Statistical Review = Revue Internationale de Statistique, 46(1): 119–120, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402522.

Som:1978:BRBd

[Som78f] R. K. Som. Book review: The Indian Investment in Family Planning by G. B. Simmons. International Statistical Review = Revue Internationale de Statistique, 46(1): 122–123, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402526.

Som:1978:BRBa

[Som78g] R. K. Som. Book review: The Labor Force in Economic Development: A Comparison of International Census Data, 1946–1966 by John D. Durand. International Statistical Review = Revue Internationale de Statistique, 46(1):117–118, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402517.

Som:1979:BRBb

[Som79a] R. K. Som. Book review: Addendum: Population Growth Estimation: A Handbook of Vital Statistics and Measurement. International Statistical Review = Revue Internationale de Statistique, 47(1):92, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403225.

Som:1979:BRBc

[Som79b] R. K. Som. Book review: Basic Ideas of Statistics by Bernard W. Lindgren. International Statistical Review = Revue Internationale de Statistique, 47(2):232, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402575.

Som:1979:BRBe

[Som79c] R. K. Som. Book review: Elementary Statistics: An Applied Approach by Neil R. Ulmann. International Statistical Review = Revue Internationale de Statistique, 47(3): 300, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402653.

Som:1979:BRBd

[Som79d] R. K. Som. Book review: Elementary Statistics by Gene R. Sellers. International Statistical Review = Revue Internationale de Statistique, 47(2):234, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402579.

Som:1979:BRBa

[Som79e] R. K. Som. Book review: Living Statistics by John Bibby. International Statistical Review = Revue Internationale de Statistique, 47(1):85–86, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403213.

Som:1980:BRBa

[Som80a] R. K. Som. Book review: Essentials of Statistics by Stephen A. Book. International Statistical Review = Revue Internationale de Statistique, 48(2):231–232, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403161.

Som:1980:BRBc

[Som80b] R. K. Som. Book review: Statistical Survey Techniques by Raymond J. Jessen. International Statistical Review = Revue Internationale de Statistique, 48(2):238, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403174.

Som:1980:BRBb

[Som80c] R. K. Som. Book review: Statistics for Technology: A Course in Applied Statistics by Christopher Chatfield. International Statistical Review = Revue Internationale de Statistique, 48(2):233, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403164.

Sorensen:1983:MLE

[Sør83] Michael Sørensen. On maximum likelihood estimation in randomly stopped diffusion-type processes. International Statistical Review = Revue Internationale de Statistique, 51(1):93–110, April 1983. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402735.

Sorensen:1986:SML

[Sør86] Michael Sørensen. On sequential maximum likelihood estimation for exponential families of stochastic processes. International Statistical Review = Revue Internationale de Statistique, 54(2):191–210, August 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403144.

Sorensen:1993:SMS

[Sør93] Michael Sørensen. Stochastic models of sand transport by wind and two related estimation problems. International Statistical Review = Revue Internationale de Statistique, 61(2):245–255, August 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403627.

Sorensen:2004:PID

[Sør04] Helle Sørensen. Parametric inference for diffusion processes observed at discrete points in time: A survey. International Statistical Review = Revue Internationale de Statistique, 72(3):337–354, December 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472627.

Sowey:1972:CCB

[Sow72] E. R. Sowey. A chronological and classified bibliography on random number generation and testing. International Statistical Review = Revue Internationale de Statistique, 40(3):355–371, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402472.

Sowey:1978:SCB

[Sow78] E. R. Sowey. A second classified bibliography on random number generation and testing. International Statistical Review = Revue Internationale de Statistique, 46(1): 89–102, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402512;   http://www.jstor.org/stable/
i261095.

Schmidt:2011:MTS

[SP11] Alexandra M. Schmidt and João Batista M. Pereira. Modelling time series of counts in epidemiology. International Statistical Review = Revue Internationale de Statistique, 79(1):48–69, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41306164.

Sengupta:1973:RPA

[SPC73] J. K. Sengupta and J. H. Portillo-Campbell. A reliability programming approach to production planning. International Statistical Review = Revue Internationale de Statistique, 41(1):115–127, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402792.

Speed:1984:DPT

[Spe84] T. P. Speed. Discussion of paper by T. Tjur. International Statistical Review = Revue Internationale de Statistique, 52(1):77–78, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403244.

Speed:1985:DPG

[Spe85] T. P. Speed. Discussion of paper by G. Shafer. International Statistical Review = Revue Internationale de Statistique, 53(3):276, December 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402891.

Speed:2007:SBR

[Spe07] Terry Speed. Short book review: Mathematics at Berkeley: A History by Calvin C. Moore. International Statistical Review = Revue Internationale de Statistique, 75 (2):277, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508492.

Speed:2008:SBRb

[Spe08a] Terry Speed. Short book review: Fifty Years of Human Genetics: A Festschrift and Liber Amicorum to Celebrate the Life and Work of George Robert Fraser edited by Oliver Mayo and Carolyn Leach. International Statistical Review = Revue Internationale de Statistique, 76(1): 155–156, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919605.

Speed:2008:SBRa

[Spe08b] Terry Speed. Short book review: Reminiscences of a Statistician: The Company I Kept by Erich L. Lehmann. International Statistical Review = Revue Internationale de Statistique, 76(1):153, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919603.

Speed:2010:SBRb

[Spe10a] Terry Speed. Short book review: I Want to Be a Mathematician: A Conversation with Paul Halmos by George Csicsery (Director). International Statistical Review = Revue Internationale de Statistique, 78(1):147–148, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919813.

Speed:2010:SBRc

[Spe10b] Terry Speed. Short book review: Statistical Models and Causal Inference: A Dialogue with the Social Sciences by David A. Freedman. International Statistical Review = Revue Internationale de Statistique, 78(3):456–457, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Speed:2010:SBRd

[Spe10c] Terry Speed. Short book review: Statistical Models: Theory and Practice, Revised Edition, by David A. Freedman, David Collier, Jasjeet S. Sekhon, and Philip B. Stark. International Statistical Review = Revue Internationale de Statistique, 78(3):457–458, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919872;   http://www.jstor.org/stable/
27919873.

Speed:2010:SBRe

[Spe10d] Terry Speed. Short book review: The Science of Bradley Efron: Selected Papers edited by Carl N. Morris, Robert Tibshirani. International Statistical Review = Revue Internationale de Statistique, 78(3):471–472, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919887.

Speed:2010:SBRa

[Spe10e] Terry Speed. Short book review: The United States of Mathematics Presidential Debate by Colin Adams, Thomas Garrity, moderated by Edward Burger. International Statistical Review = Revue Internationale de Statistique, 78(1):147, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919812.

Speed:2011:SBRb

[Spe11a] Terry Speed. Short book review: Large-Scale Inference: Empirical Bayes Methods for Estimation, Testing, and Prediction by Bradley Efron. International Statistical Review = Revue Internationale de Statistique, 79(1):126– 127, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306179.

Speed:2011:SBRa

[Spe11b] Terry Speed. Short book review: The Pleasures of Statistics: The Autobiography of Frederick Mosteller by Frederick Mosteller, edited by Stephen E. Fienberg David C. Hoaglin, Judith M. Tanur. International Statistical Review = Revue Internationale de Statistique, 79(1):125– 126, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306178.

Spizzichino:2002:SCE

[Spi02] Fabio Spizzichino. [some cracks in the empire of chance (flaws in the foundations of reliability)]: Comment. International Statistical Review = Revue Internationale de Statistique, 70(1):68–77, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403725.

Spjotvoll:1984:DPD

[Spj84] Emil Spjøtvoll. Discussion of paper by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 52(1):28, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403239.

Sprent:1975:BRBa

[Spr75a] P. Sprent. Book review: Applied Statistics: Analysis of Variance and Regression by Olive J. Dunn; Virginia A. Clark. International Statistical Review = Revue Internationale de Statistique, 43(1):117, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402667.

Sprent:1975:BRBb

[Spr75b] P. Sprent. Book review: Regression Estimation from Grouped Observations by Y. Haitovsky. International Statistical Review = Revue Internationale de Statistique, 43 (2):246, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402918.

Sprent:1977:BRBb

[Spr77a] P. Sprent. Book review: An Introduction to Linear Algebra by Hans Samelson. International Statistical Review = Revue Internationale de Statistique, 45(1):102–103, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403023.

Sprent:1977:BRBa

[Spr77b] P. Sprent. Book review: Statistics in the Real World. A Book of Examples by Richard J. Larsen; Donna Fox Stroup. International Statistical Review = Revue Internationale de Statistique, 45(1):99, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403016.

Sprent:1980:BRB

[Spr80] P. Sprent. Book review: Experiments: Design and Analysis by J. A. John; M. H. Quenouille. International Statistical Review = Revue Internationale de Statistique, 48 (1):125, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402427.

Singh:1992:SSM

[SR92] A. C. Singh and G. R. Roberts. State space modelling of cross-classified time series of counts. International Statistical Review = Revue Internationale de Statistique, 60 (3):321–335, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403681.

Steel:2015:DFD

[SR15] Mark F. J. Steel and Francisco J. Rubio. Discussion on On Families of Distributions with Shape Parameters: Discussion. International Statistical Review = Revue Internationale de Statistique, 83(2):218–222, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Jon15a].

Srivastava:1973:EPD

[Sri73] R. C. Srivastava. Estimation of probability density function based on random number of observations with applications. International Statistical Review = Revue Internationale de Statistique, 41(1):77–86, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402789.

Srivastava:1980:ELS

[Sri80] V. K. Srivastava. Estimation of linear single-equation and simultaneous-equation models under stochastic linear constraints: an annotated bibliography. International Statistical Review = Revue Internationale de Statistique, 48(1): 79–82, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402407.

Singer:2017:GTD

[SRN17] Julio M. Singer, Francisco M. M. Rocha, and Juvêncio S. Nobre. Graphical tools for detecting departures from linear mixed model assumptions and some remedial measures. International Statistical Review = Revue Internationale de Statistique, 85(2):290–324, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sarndal:1987:GVE

[SS87] Carl-Erik Särndal and Bengt Swensson. A general view of estimation for two phases of selection with applications to two-phase sampling and nonresponse. International Statistical Review = Revue Internationale de Statistique, 55 (3):279–294, December 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403406.

Smith:1988:SAM

[SS88] T. M. F. Smith and R. A. Sugden. Sampling and assignment mechanisms in experiments, surveys and observational studies, correspondent paper. International Statistical Review = Revue Internationale de Statistique, 56(2): 165–180, August 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403640.

Sugden:2007:DBP

[SS07] Roger A. Sugden and T. M. Fred Smith. Design-based properties of linear calibrated estimators of a finite population total. International Statistical Review = Revue Internationale de Statistique, 75(2):218–223, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508462.

Sellers:2008:MVL

[SS08] Kimberly F. Sellers and Nozer D. Singpurwalla. Manyvalued logic in multistate and vague stochastic systems. International Statistical Review = Revue Internationale de Statistique, 76(2):247–267, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919615.

Seneta:2009:PKP

[SS09a] Eugene Seneta and Ida H. Stamhuis. Preface to Karl Pearson issue. International Statistical Review = Revue Internationale de Statistique, 77(1):1–2, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919686.

Stamhuis:2009:PSN

[SS09b] Ida H. Stamhuis and Eugene Seneta. Pearson’s statistics in the Netherlands and the astronomer Kapteyn. International Statistical Review = Revue Internationale de Statistique, 77(1):96–117, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919694.

Schouten:2017:SAS

[SS17] Barry Schouten and Natalie Shlomo. Selecting adaptive survey design strata with partial R-indicators. International Statistical Review = Revue Internationale de Statistique, 85(1):143–163, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:1977:ATS

[SSJ77] A. J. Scott, T. M. F. Smith, and R. G. Jones. The application of time series methods to the analysis of repeated surveys. International Statistical Review = Revue Internationale de Statistique, 45(1):13–28, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403000.

Sanderson:2004:CPP

[SSOL04] Warren C. Sanderson, Sergei Scherbov, Brian C. O’Neill, and Wolfgang Lutz. Conditional probabilistic population forecasting. International Statistical Review = Revue Internationale de Statistique, 72(2):157–166, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403851.

Steorts:2020:SBS

[SST20] Rebecca C. Steorts, Timo Schmid, and Nikos Tzavidis. Smoothing and benchmarking for small area estimation. International Statistical Review = Revue Internationale de Statistique, 88(3):580–598, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Silverman:1992:ISR

[ST92] Bernard Silverman and Dennis Trewin. International statistical review. Statement of editorial policy. International Statistical Review = Revue Internationale de Statistique, 60(1):1, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403496.

Schuyten:2007:STC

[ST07] Gilberte Schuyten and Olivier Thas. Statistical thinking in computer-based learning environments. International Statistical Review = Revue Internationale de Statistique, 75(3):365–371, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41509877.

Starkings:2007:SBRa

[Sta07a] Susan Starkings. Short book review: Data Analysis and Graphics Using R: An Example-based Approach, 2nd Edition, by John Maindonald, John Braun. International Statistical Review = Revue Internationale de Statistique, 75 (2):270, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508485.

Starkings:2007:SBRd

[Sta07b] Susan Starkings. Short book review: Data Preparation for Analytics Using SAS(R) by Gerhard Svolba. International Statistical Review = Revue Internationale de Statistique, 75(3):429, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509899;   http://
www.jstor.org/stable/41509900.

Starkings:2007:SBRf

[Sta07c] Susan Starkings. Short book review: Elementary Statistics Using JMP(R) by Sandra D. Schlotzhauer. International Statistical Review = Revue Internationale de Statistique, 75(3):430, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509901;   http://
www.jstor.org/stable/41509902.

Starkings:2007:SBRg

[Sta07d] Susan Starkings. Short book review: Learning SAS(R) by Example: A Programmer’s Guide by Ron Cody. International Statistical Review = Revue Internationale de Statistique, 75(3):430, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2007:SBRe

[Sta07e] Susan Starkings. Short book review: Pharmaceutical Statistics Using SAS(R): A Practical Guide by Alex Dmitrienko, Christy Chuang-Stein, Ralph D’Agostino. International Statistical Review = Revue Internationale de Statistique, 75(3):429, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2007:SBRc

[Sta07f] Susan Starkings. Short book review: Statistics Using SAS(R) Enterprise Guide(R) by James B. Davis. International Statistical Review = Revue Internationale de Statistique, 75(3):428–429, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/41509898.

Starkings:2007:SBRb

[Sta07g] Susan Starkings. Short book review: The R Book by Michael J. Crawley. International Statistical Review = Revue Internationale de Statistique, 75(3):425–426, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41509895.

Starkings:2008:SBRa

[Sta08a] Susan Starkings. Short book review: Correspondence Analysis in Practice, Second Edition, by Michael Greenacre. International Statistical Review = Revue Internationale de Statistique, 76(1):148–149, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919598.

Starkings:2008:SBRc

[Sta08b] Susan Starkings. Short book review: Statistical Methods for Human Rights edited by Jana Asher, David Banks, Fritz J. Scheuren. International Statistical Review = Revue Internationale de Statistique, 76(2):328, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919646.

Starkings:2008:SBRb

[Sta08c] Susan Starkings. Short book review: Statistical Thinking in Sports edited by Jim Albert, Ruud H. Koning. International Statistical Review = Revue Internationale de Statistique, 76(1):148–149, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2009:SBR

[Sta09] Susan Starkings. Short book review: Naive Decision Making: Mathematics Applied to the Social World by T. W. Körner. International Statistical Review = Revue Internationale de Statistique, 77(1):151, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919700.

Starkings:2010:SBRa

[Sta10a] Susan Starkings. Short book review: A First Course in Probability and Statistics by B. L. S. Prakasa Rao. International Statistical Review = Revue Internationale de Statistique, 78(2):320, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919845.

Starkings:2010:SBRc

[Sta10b] Susan Starkings. Short book review: Exploring the Origin, Extent, and Future of Life: Philosophical, Ethical and Theological Perspectives edited by Constance M. Bertka. International Statistical Review = Revue Internationale de Statistique, 78(2):321–322, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919847.

Starkings:2010:SBRb

[Sta10c] Susan Starkings. Short book review: Statistics for Engineers: An Introduction by S. J. Morrison. International Statistical Review = Revue Internationale de Statistique, 78(2):320–321, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919846.

Starkings:2011:SBRa

[Sta11a] Susan Starkings. Short book review: Foundation Mathematics for the Physical Sciences by K. F. Riley, M. P. Hobson. International Statistical Review = Revue Internationale de Statistique, 79(3):493, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305074.

Starkings:2011:SBRb

[Sta11b] Susan Starkings. Short book review: Foundation Mathematics for the Physical Sciences: Student Solution Manual by K. F. Riley, M. P. Hobson. International Statistical Review = Revue Internationale de Statistique, 79(3): 494, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41305075.

Starkings:2012:SBRa

[Sta12a] Susan Starkings. Short book review: IBM SPSS Statistics 19 Made Simple by Colin D. Gray and Paul R. Kinnear. International Statistical Review = Revue Internationale de Statistique, 80(2):333–334, August 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2012:SBRf

[Sta12b] Susan Starkings. Short book review: Mathematical and Statistical Methods for Actuarial Sciences and Finance edited by Cira Perna and Marilena Sibillo. International Statistical Review = Revue Internationale de Statistique, 80(2):340, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2012:SBRe

[Sta12c] Susan Starkings. Short book review: Modern Statistics for the Social and Behavioral Sciences: A Practical Introduction by Rand Wilcox. International Statistical Review = Revue Internationale de Statistique, 80(2):339–340, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2012:SBRb

[Sta12d] Susan Starkings. Short book review: Quantitative Data Analysis with IBM SPSS 17, 18 & 19: A Guide for Social Scientists by Alan Bryman and Duncan Cramer. International Statistical Review = Revue Internationale de Statistique, 80(2):334–335, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2012:SBRc

[Sta12e] Susan Starkings. Short book review: Statistical Analysis of Clinical Data on a Pocket Calculator: Statistics on a Pocket Calculator by Ton J. Cleophas and Aeilko H. Zwinderman. International Statistical Review = Revue Internationale de Statistique, 80(2):335, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2012:SBRd

[Sta12f] Susan Starkings. Short book review: Statistical Modeling of the National Assessment of Educational Progress by Murray Aitkin and Irit Aitkin. International Statistical Review = Revue Internationale de Statistique, 80(2): 338–339, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2013:SBRa

[Sta13a] Susan Starkings. Short book review: A Mathematical Orchard: Problems and Solutions by Mark I. Krusemeyer, George T. Gilbert, Loren C. Larson. International Statistical Review = Revue Internationale de Statistique, 81(2): 319–320, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2013:SBRb

[Sta13b] Susan Starkings. Short book review: Excel 2010 for Educational and Psychological Statistics: A Guide to Solving Practical Problems by Thomas J. Quirk. International Statistical Review = Revue Internationale de Statistique, 81(2):320–321, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2013:SBRc

[Sta13c] Susan Starkings. Short book review: Methods and Applications of Statistics in the Social and Behavioral Sciences by N. Balakrishnan. International Statistical Review = Revue Internationale de Statistique, 81(2):321, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Staicu:2017:IIN

[Sta17] Ana-Maria Staicu. Interview: Interview with Nancy Reid. International Statistical Review = Revue Internationale de Statistique, 85(3):381–403, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Steffensen:1995:POD

[Ste95] J. F. Steffensen. On the probability that the offspring dies out. International Statistical Review = Revue Internationale de Statistique, 63(2):239–242, August 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403618.

Stigler:1988:DAP

[Sti88] Stephen M. Stigler. The Dark Ages of Probability in England: The Seventeenth Century work of Richard Cumberland and Thomas Strode. International Statistical Review = Revue Internationale de Statistique, 56(1):75–88, April 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403363.

Stigler:2000:ISM

[Sti00] Stephen M. Stigler. International statistics at the Millennium: Progressing or regressing? International Statistical Review = Revue Internationale de Statistique, 68(2):111– 121, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403663.

Stigler:2005:PPA

[Sti05] Stephen M. Stigler. Panel on public awareness, professionalism & professional accreditation; international collaboration. International Statistical Review = Revue Internationale de Statistique, 73(2):273, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472669.

Sazak:2006:RAS

[STI06] Hakan S. Sazak, Moti L. Tiku, and M. Qamarul Islam. Regression analysis with a stochastic design variable. International Statistical Review = Revue Internationale de Statistique, 74(1):77–88, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/25472692.

Stigler:2014:SQH

[Sti14] Stephen Stigler. Soft questions, hard answers: Jacob Bernoulli’s probability in historical context. International Statistical Review = Revue Internationale de Statistique, 82(1):1–16, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Silverman:1995:SPI

[STM95] Bernard Silverman, Dennis Trewin, and David S. Moore. Special part issue on statistical education. International Statistical Review = Revue Internationale de Statistique, 63(1):1, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403773.

Silverman:1993:SIS

[STNW93] Bernard Silverman, Dennis Trewin, Vijay Nair, and Jeff Wu. Special issue on statistics in industry: Introduction. International Statistical Review = Revue Internationale de Statistique, 61(1):1–2, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403589.

Stone:1975:BRB

[Sto75] Richard Stone. Book review: Stochastic Models for Social Processes by D. J. Bartholomew. International Statistical Review = Revue Internationale de Statistique, 43(2): 241, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402910.

Stone:1976:AA

[Sto76] Richard Stone. Abraham Aidenoff, 1913–1975. International Statistical Review = Revue Internationale de Statistique, 44(3):383–384, December 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402364.

Stoyan:1990:SSG

[Sto90] D. Stoyan. Stereology and stochastic geometry. International Statistical Review = Revue Internationale de Statistique, 58(3):227–242, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403806.

Stoyan:1998:RSM

[Sto98] Dietrich Stoyan. Random sets: Models and statistics. International Statistical Review = Revue Internationale de Statistique, 66(1):1–27, April 1998. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403654.

Streit:1973:BRB

[Str73] Franz Streit. Book review: Multivariate Analysis by A. M. Kshirsagar. International Statistical Review = Revue Internationale de Statistique, 41(3):400, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402638.

Straub:1976:BRB

[Str76] Erwin Straub. Book review: Statistical Theory of Reliability and Life Testing Probability Models by Richard E. Barlow; Frank Proschan. International Statistical Review = Revue Internationale de Statistique, 44(1):167, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402970.

Strobl:2014:DFY

[Str14] Carolin Strobl. Discussions on fifty years of classification and regression trees: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(3):349– 352, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Stuart:1972:BRB

[Stu72] A. Stuart. Book review: Practical Nonparametric Statistics by W. J. Conover. International Statistical Review = Revue Internationale de Statistique, 40(3):393, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402479.

Stuart:1980:BRB

[Stu80] A. Stuart. Book review: The Analysis of Cross-Classified Categorical Data by Stephen E. Fienberg. International Statistical Review = Revue Internationale de Statistique, 48(1):121–122, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402421.

Smith:1990:SME

[STY90] R. L. Smith, J. A. Tawn, and H. K. Yuen. Statistics of multivariate extremes. International Statistical Review = Revue Internationale de Statistique, 58(1):47–58, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403473.

Sunter:1977:RBS

[Sun77] A. B. Sunter. Response burden, sample rotation, and classification renewal in economic surveys. International Statistical Review = Revue Internationale de Statistique, 45 (3):209–222, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402535.

Sunter:1986:SPU

[Sun86] Alan Sunter. Solutions to the problem of unequal probability sampling without replacement. International Statistical Review = Revue Internationale de Statistique, 54 (1):33–50, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403257.

Sundgren:1996:MSD

[Sun96] Bo Sundgren. Making statistical data more available. International Statistical Review = Revue Internationale de Statistique, 64(1):23–38, April 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403422.

Sund:2015:BRM

[Sun15] Reijo Sund. Book review: Multistate Analysis of Life Histories with R. International Statistical Review = Revue Internationale de Statistique, 83(2):324–325, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2016:BRE

[Sun16a] Reijo Sund. Book review: Exposure-Response Modelling: Methods and Practical Implementation, Jixian Wang, Chapman & Hall/CRC, 2015, xv + 331 pages, £61.99, hardback ISBN: 978-1-4665-7320-8. International Statistical Review = Revue Internationale de Statistique, 84(3): 556–557, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2016:BRG

[Sun16b] Reijo Sund. Book review: Graphical Data Analysis with R. International Statistical Review = Revue Internationale de Statistique, 84(2):324–325, August 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2016:BRM

[Sun16c] Reijo Sund. Book review: Meta-Analysis: A Structural Equation Modeling Approach. International Statistical Review = Revue Internationale de Statistique, 84(1):163– 164, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2017:BRI

[Sun17] Reijo Sund. Book review: Improving Population Health Using Electronic Health Records. Methods for Data Management and Epidemiological Analysis, Neal D. Goldstein CRC Press, 2017, xix + 254 pages, £45.99, softcover ISBN: 978-1-138-19637-7. International Statistical Review = Revue Internationale de Statistique, 85(2):376–377, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2018:BRE

[Sun18] Reijo Sund. Book review: Errors, Blunders, and Lies: How to Tell the Difference, David S. Salsburg, ASA-CRC Series on Statistical Reasoning in Science and Society. Chapman and Hall/CRC, 2017, 154 pages, £77.00, hardcover ISBN: 978-1-138-72698-7. International Statistical Review = Revue Internationale de Statistique, 86(1):166– 167, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2019:BRC

[Sun19] Reijo Sund. Book review: Computer Age Statistical Inference: Algorithms, Evidence, and Data Science, Bradley Efron and Trevor Hastie, Institute of Mathematical Statistics Monographs Cambridge University Press, 2016, (8th printing 2018), xix + 475 pages, $74.99, hardcover ISBN: 978-1-107-14989-2. International Statistical Review = Revue Internationale de Statistique, 87(1):186–188, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2022:BRA

[Sun22] Reijo Sund. Book review: Administrative Records for Survey Methodology. Edited by Asaph Young Chun, Michael D. Larsen, Gabriele Durrant, Jerome P. Reiter. John Wiley and Sons, 2021, 384 pages, $128.95 (hardcover). ISBN: 978-1-119-27204-5. International Statistical Review = Revue Internationale de Statistique, 90(2):415–417, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2024:BRC

[Sun24] Reijo Sund. Book review: Controlled Epidemiological Studies. Marie Reilly. Chapman & Hall/CRC Biostatistics Series, CRC Press, 2023, xxxii + 439 pages, £115/$150, hardcover. ISBN: 978-0-367-18678-4 (hbk). International Statistical Review = Revue Internationale de Statistique, 92(2):325–326, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Schmutte:2022:IJM

[SV22] Ian Schmutte and Lars Vilhuber. An interview with John M. Abowd. International Statistical Review = Revue Internationale de Statistique, 90(1):1–40, April 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:1988:VQV

[SVCV88] Chris Scott, Martin Vaessen, Sidiki Coulibaly, and Jane Verrall. Verbatim questionnaires versus field translation or schedules: An experimental study. International Statistical Review = Revue Internationale de Statistique, 56 (3):259–278, December 1988. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403353.

Singpurwalla:1994:SRM

[SW94] Nozer D. Singpurwalla and Simon P. Wilson. Software reliability modeling. International Statistical Review = Revue Internationale de Statistique, 62(3):289–317, December 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403763.

Seneta:2007:P

[SW07] Eugene Seneta and Chris Wild. Preface. International Statistical Review = Revue Internationale de Statistique, 75(3):279–280, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41509870.

Scott:2011:D

[SW11] Alastair J. Scott and Christopher J. Wild. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):228–230, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Simar:2015:SAN

[SW15] Léopold Simar and Paul W. Wilson. Statistical approaches for non-parametric frontier models: A guided tour. International Statistical Review = Revue Internationale de Statistique, 83(1):77–110, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sweeting:1995:SMN

[Swe95] Trevor Sweeting. Statistical models for nondestructive evaluation. International Statistical Review = Revue Internationale de Statistique, 63(2):199–214, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403614.

Sylla:2014:TBC

[Syl14] Edith Dudley Sylla. Tercentenary of Ars Conjectandi (1713): Jacob Bernoulli and the founding of mathematical probability. International Statistical Review = Revue Internationale de Statistique, 82(1):27–45, April 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Smith:2002:ELT

[SYSK02] Michael Smith, Paul Yau, Thomas Shively, and Robert Kohn. Estimating long-term trends in tropospheric ozone levels. International Statistical Review = Revue Internationale de Statistique, 70(1):99–124, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403728.

Seneta:2009:NEC

[SZ09] Eugene Seneta and Kees Zeelenberg. New Editor-in-Chief for the International Statistical Review. International Statistical Review = Revue Internationale de Statistique, 77 (3):329–330, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919759.

Song:2014:DFY

[SZ14] Chi Song and Heping Zhang. Discussions on fifty years of classification and regression trees: Discussion. International Statistical Review = Revue Internationale de Statistique, 82(3):359–361, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Szczepanska:2011:SBRa

[Szc11a] Anna Szczepańska. Short book review: Multilevel Analysis: Techniques and Applications, Second Edition, by Joop J. Hox. International Statistical Review = Revue Internationale de Statistique, 79(3):490, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41305071.

Szczepanska:2011:SBRb

[Szc11b] Anna Szczepańska. Short book review: Research Design and Statistical Analysis, Third Edition by Jerome L. Myers, Arnold D. Well, Robert F. Lorch, Jr. International Statistical Review = Revue Internationale de Statistique, 79(3):491–492, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305072.

Taamouti:2011:D

[Taa11] Mohamed Taamouti. Discussion. International Statistical Review = Revue Internationale de Statistique, 79(3):331– 333, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tadikamalla:1980:LBR

[Tad80] Pandu R. Tadikamalla. A look at the Burr and related distributions. International Statistical Review = Revue Internationale de Statistique, 48(3):337–344, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402945.

Talukder:1993:GMA

[Tal93] M. A. H. Talukder. A general method of analysis of factorial designs with unequal cell variances. International Statistical Review = Revue Internationale de Statistique, 61(3):409–417, December 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403752.

Tam:1982:PEL

[Tam82] S. M. Tam. Postcensal estimates for local areas using current samples with census as the source of sampling frame. International Statistical Review = Revue Internationale de Statistique, 50(2):125–134, August 1982. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402596.

Tam:1984:DPK

[Tam84] S. M. Tam. Discussion of paper by A. K. Adhikari, A. Chaudhuri and K. Vijayan. International Statistical Review = Revue Internationale de Statistique, 52(2):122– 124, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403092.

Tam:1987:ARS

[Tam87] S. M. Tam. Analysis of repeated surveys using a dynamic linear model. International Statistical Review = Revue Internationale de Statistique, 55(1):63–73, April 1987. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403271.

Tam:1994:SSA

[Tam94] S. M. Tam. [sample surveys 1975–1990; an age of reconciliation?]: Discussion. International Statistical Review = Revue Internationale de Statistique, 62(1):28–32, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403542.

Tamura:2007:FVS

[Tam07] Hirokuni Tamura. Foundational value of statistics education for management curriculum. International Statistical Review = Revue Internationale de Statistique, 75(3):397– 405, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41509879.

Taniguchi:1992:AFS

[Tan92] Masanobu Taniguchi. An automatic formula for the second-order approximation of the distributions of test statistics under contiguous alternatives. International Statistical Review = Revue Internationale de Statistique, 60 (2):211–225, August 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403651.

Tanur:1997:DLH

[Tan97] Judith M. Tanur. Discussion: Leading horses to water. International Statistical Review = Revue Internationale de Statistique, 65(2):159–161, August 1997. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403340.

Tang:2007:SBR

[Tan07] Boxin Tang. Short book review: Modern Experimental Design by Thomas P. Ryan. International Statistical Review = Revue Internationale de Statistique, 75(2):258– 259, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508474.

Tutz:2022:SOR

[TB22] Gerhard Tutz and Moritz Berger. Sparser ordinal regression models based on parametric and additive locationshift approaches. International Statistical Review = Revue Internationale de Statistique, 90(2):306–327, August 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tjur:1984:AVMb

[TBSW84] T. Tjur, R. A. Bailey, T. P. Speed, and H. P. Wynn. [analysis of variance models in orthogonal designs]: Reply to discussion. International Statistical Review = Revue Internationale de Statistique, 52(1):79–81, April 1984. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403246.

Tsokos:1972:BCE

[TC72] Chris P. Tsokos and George C. Canavos. Bayesian concepts for the estimation of reliability in the Weibull lifetesting model. International Statistical Review = Revue Internationale de Statistique, 40(2):153–160, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402758.

Tam:1984:SPS

[TC84] S. M. Tam and N. N. Chan. Screening of probability samples. International Statistical Review = Revue Internationale de Statistique, 52(3):301–308, December 1984. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403048.

Thyregod:1999:USS

[TC99] Poul Thyregod and Knut Conradsen. [using statistics and statistical thinking to improve organisational performance]: Discussion. International Statistical Review = Revue Internationale de Statistique, 67(2):144–146, August 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403394.

Thompson:2007:ONT

[TC07] Jeffrey R. Thompson and Randy L. Carter. An overview of normal theory structural measurement error models. International Statistical Review = Revue Internationale de Statistique, 75(2):183–198, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41508460.

Tam:2015:BDO

[TC15] Siu-Ming Tam and Frederic Clarke. Big data, official statistics and some initiatives by the Australian Bureau of Statistics. International Statistical Review = Revue Internationale de Statistique, 83(3):436–448, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsiatis:2011:D

[TD11] Anastasios A. Tsiatis and Marie Davidian. Discussions. International Statistical Review = Revue Internationale de Statistique, 79(2):221–223, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Teugels:1997:CPS

[Teu97] Jef L. Teugels. [the current position of statistics: A personal view]: Discussion. International Statistical Review = Revue Internationale de Statistique, 65(3):287–288, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403371.

Teugels:2005:DJT

[Teu05a] Jef L. Teugels. Discussant: Jef L. Teugels. International Statistical Review = Revue Internationale de Statistique, 73(2):165–168, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Teugels:2005:ICW

[Teu05b] Jef L. Teugels. Internal cohesion within ISI. International Statistical Review = Revue Internationale de Statistique, 73(2):275–276, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/25472670.

Tam:1994:NMP

[TG94] S. M. Tam and B. W. Green. A new methodology for processing Australia’s 1991 population and housing census. International Statistical Review = Revue Internationale de Statistique, 62(1):71–85, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403546.

Tuyl:2009:RTV

[TGM09] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. The rule of three, its variants and extensions. International Statistical Review = Revue Internationale de Statistique, 77(2):266–275, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919727.

Thomsen:1998:CDS

[TH98] Ib Thomsen and Ann Marit Kleive Holmøy. Combining data from surveys and administrative record systems. The Norwegian experience. International Statistical Review = Revue Internationale de Statistique, 66(2):201– 221, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403490.

Thabane:2007:BRB

[Tha07a] Lehana Thabane. Book review: A Pocket Guide to Epidemiology by David G. Kleinbaum; Kevin M. Sullivan; Nancy D. Barker. International Statistical Review = Revue Internationale de Statistique, 75(1):124–125, April 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41508454.

Thabane:2007:SBR

[Tha07b] Lehana Thabane. Short book review: Medical Statistics from A to Z: A Guide for Clinicians and Medical Students, 2nd Edition by B. S. Everitt. International Statistical Review = Revue Internationale de Statistique, 75(2): 261–263, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508478.

Thompson:1978:SVMa

[Tho78a] Mary L. Thompson. Selection of variables in multiple regression: Part I. A review and evaluation. International Statistical Review = Revue Internationale de Statistique, 46(1):1–19, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402505. See errata [Ano78g].

Thompson:1978:SVMb

[Tho78b] Mary L. Thompson. Selection of variables in multiple regression: Part II. Chosen procedures, computations and examples. International Statistical Review = Revue Internationale de Statistique, 46(2):129–146, August 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402809.

Thompson:1984:DPK

[Tho84] M. E. Thompson. Discussion of paper by A. K. Adhikari, A. Chaudhuri and K. Vijayan. International Statistical Review = Revue Internationale de Statistique, 52 (2):124, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403093.

Thompson:1998:IGA

[Tho98] Elizabeth Thompson. Inferring gene ancestry: Estimating gene descent. International Statistical Review = Revue Internationale de Statistique, 66(1):29–40, April 1998. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403655.

Thompson:2000:MEM

[Tho00] E. A. Thompson. MCMC estimation of multi-locus genome sharing and multipoint gene location scores. International Statistical Review = Revue Internationale de Statistique, 68(1):53–73, April 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403557.

Thompson:2006:CV

[Tho06] Bill Thompson. Critique of p-values. International Statistical Review = Revue Internationale de Statistique, 74 (1):1–14, April 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472687.

Thompson:2014:BAS

[Tho14] John Thompson. Bayesian Analysis with Stata. Stata Press, College Station, TX, USA, 2014. ISBN 1-59718-141-2. xx + 279 pp. LCCN QA279.5 .T56 2014. URL

http://www.loc.gov/catdir/enhancements/fy1501/2014930482-
b.html.

Thompson:2019:CDN

[Tho19] Mary E. Thompson. Combining data from new and traditional sources in population surveys. International Statistical Review = Revue Internationale de Statistique, 87 (S1):S79–S89, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tian:2007:SDO

[Tia07] Yongge Tian. Some decompositions of OLSEs and BLUEs under a partitioned linear model. International Statistical Review = Revue Internationale de Statistique, 75(2):224– 248, August 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41508463.

Tille:1998:ESU

[Til98] Yves Tillé. Estimation in surveys using conditional inclusion probabilities: Simple random sampling. International Statistical Review = Revue Internationale de Statistique, 66(3):303–322, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403519.

Tille:2022:SSI

[Til22] Yves Tillé. Some solutions inspired by survey sampling theory to build effective clinical trials. International Statistical Review = Revue Internationale de Statistique, 90 (3):481–498, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tintner:1978:BRBe

[Tin78a] Gerhard Tintner. Book review: Analisis de decisiones by Sixto Rios. International Statistical Review = Revue Internationale de Statistique, 46(3):322, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402387.

Tintner:1978:BRBb

[Tin78b] Gerhard Tintner. Book review: Storia del Metodo Statistico by Angiolo Maros Dell’Oro; A. Giuffre. International Statistical Review = Revue Internationale de Statistique, 46(2):228, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402820.

Tintner:1978:BRBc

[Tin78c] Gerhard Tintner. Book review: Sufficient Statistics: Selected Contributions by V. S. Huzurbazar; A. M. Kshirsagar. International Statistical Review = Revue Internationale de Statistique, 46(2):228–229, August 1978. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402821.

Tintner:1978:BRBd

[Tin78d] Gerhard Tintner. Book review: The Econometrics of Structural Change by D. J. Poirier. International Statistical Review = Revue Internationale de Statistique, 46(2): 229–230, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402823.

Tintner:1978:BRBa

[Tin78e] Gerhard Tintner. Book review: Time Series Analysis and Forecasting: The Box–Jenkins Approach by O. D. Anderson. International Statistical Review = Revue Internationale de Statistique, 46(2):228, August 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402819.

Tintner:1980:BRBa

[Tin80a] Gerhard Tintner. Book review: Linear Least Squares Estimation by Thomas Kailath. International Statistical Review = Revue Internationale de Statistique, 48(1): 125–126, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402428.

Tintner:1980:BRBb

[Tin80b] Gerhard Tintner. Book review: Stability and Inflation. A Volume of Essays to Honour the Memory of A. W. H. Phillips by A. R. Bergstrom; A. J. L. Catt; M. H. Peston; B. D. J. Silverstone; A. W. H. Phillips. International Statistical Review = Revue Internationale de Statistique, 48 (2):230, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403158.

Tippett:1975:BRBb

[Tip75a] L. H. C. Tippett. Book review: O.R. in World War II: Operational Research against the U-Boat by C. H. Waddington. International Statistical Review = Revue Internationale de Statistique, 43(3):369, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403126.

Tippett:1975:BRBa

[Tip75b] L. H. C. Tippett. Book review: Statistical Methods in the Process Industries by Maurice H. Belz. International Statistical Review = Revue Internationale de Statistique, 43(2):241–242, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402911.

Tippet:1977:BRB

[Tip77] L. H. C. Tippet. Book reviews: Elementary Statistical Concepts by R. E. Walpole; Statistical Methods for Engineers and Scientists by R. M. Bethea; B. S. Duran; T. L. Boullion; Basic Statistics with Business Applications by R. C. Clelland; J. S. de Cani; F. E. Brown. International Statistical Review = Revue Internationale de Statistique, 45(1):105, April 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403026.

Tippett:1980:BRB

[Tip80] L. H. C. Tippett. Book reviews: Statistics in Action by P. Sprent; Statistical Handbook for Non-Statisticians by R. Meddis. International Statistical Review = Revue Internationale de Statistique, 48(1):133, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402438.

Titterington:1985:CSS

[Tit85] D. M. Titterington. Common structure of smoothing techniques in statistics. International Statistical Review = Revue Internationale de Statistique, 53(2):141–170, August 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402932.

Tjur:1984:AVMa

[Tju84] Tue Tjur. Analysis of variance models in orthogonal designs. International Statistical Review = Revue Internationale de Statistique, 52(1):33–65, April 1984. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403242.

With discussion by R. A. Bailey, T. P. Speed and H. P. Wynn, and a reply by the author.

Tuljapurkar:2004:RSF

[TLL04] Shripad Tuljapurkar, Ronald D. Lee, and Qi Li. Random scenario forecasts versus stochastic forecasts. International Statistical Review = Revue Internationale de Statistique, 72(2):185–199, August 2004. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403853.

Tutz:2017:MCH

[TO17] Gerhard Tutz and Margret-Ruth Oelker. Modelling clustered heterogeneity: Fixed effects, random effects and mixtures. International Statistical Review = Revue Internationale de Statistique, 85(2):204–227, August 2017. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsokos:1973:SIE

[TP73] Chris P. Tsokos and W. J. Padgett. Stochastic integral equations in life sciences and engineering. International Statistical Review = Revue Internationale de Statistique, 41(1):15–38, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402785.

Tracy:1973:BRB

[Tra73] D. S. Tracy. Book review: Generalized Inverse Matrices with Applications to Statistics by R. M. Pringle; A. A. Rayner. International Statistical Review = Revue Internationale de Statistique, 41(1):140–141, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402799.

Traat:2007:SBR

[Tra07] Imbi Traat. Short book review: Introduction to Variance Estimation, 2nd Edition, by Kirk M. Wolter. International Statistical Review = Revue Internationale de Statistique, 75(3):415, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/41509886.

Traat:2009:SBR

[Tra09] Imbi Traat. Short book review: Simulation and the Monte Carlo Method, 2nd Edition, by Reuven Y. Rubinstein, Dirk P. Kroese. International Statistical Review = Revue Internationale de Statistique, 77(1):153–154, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919703.

Traat:2013:SBR

[Tra13] Imbi Traat. Short book review: Maximum Likelihood Estimation for Sample Surveys by Raymond L. Chambers, David G. Steel, Suojin Wang, Alan H. Welsh. International Statistical Review = Revue Internationale de Statistique, 81(2):317–318, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Trewin:2000:I

[Tre00a] Dennis Trewin. [introduction]. International Statistical Review = Revue Internationale de Statistique, 68(2): 199, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403669.

Trewin:2000:RBS

[Tre00b] Dennis Trewin. The role of behavioral survey methodologists in national statistical agencies. International Statistical Review = Revue Internationale de Statistique, 68 (2):199, August 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Trewin:2005:DDT

[Tre05a] Dennis Trewin. Discussant: Dennis trewin. International Statistical Review = Revue Internationale de Statistique, 73(2):237–238, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Trewin:2005:ISL

[Tre05b] Dennis Trewin. Improving statistical literacy in schools in Australia. International Statistical Review = Revue Internationale de Statistique, 73(2):211–212, August 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472657.

Triantafyllopoulos:2009:IDG

[Tri09] Kostas Triantafyllopoulos. Inference of dynamic generalized linear models: On-line computation and appraisal. International Statistical Review = Revue Internationale de Statistique, 77(3):430–450, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919767.

Toman:2023:OEN

[TRRL23] Patrick Toman, Nalini Ravishanker, Sanguthevar Rajasekaran, and Nathan Lally. Online evidential nearest neighbour classification for Internet of Things time series. International Statistical Review = Revue Internationale de Statistique, 91(3):395–426, December 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Trutzel:2005:DKT

[Tru05] Klaus Trutzel. Discussant: Klaus Trutzel. International Statistical Review = Revue Internationale de Statistique, 73(2):201–203, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tjetjep:2006:SNV

[TS06] Annelies Tjetjep and Eugene Seneta. Skewed normal variance-mean models for asset pricing and the method of moments. International Statistical Review = Revue Internationale de Statistique, 74(1):109–126, April 2006. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472694.

Tsay:2024:MVT

[Tsa24] Ruey S. Tsay. Matrix-variate time series analysis: A brief review and some new developments. International Statistical Review = Revue Internationale de Statistique, 92(2): 246–262, August 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsionas:2003:BIE

[Tsi03] Efthymios G. Tsionas. Bayesian international evidence on heavy tails, non-stationarity and asymmetry over the business cycle. International Statistical Review = Revue Internationale de Statistique, 71(1):151–168, April 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403879.

Tsodikov:2016:BRB

[Tso16] Alexander Tsodikov. Book review: Branching Process Models of Cancer, Richard Durrett, Springer International Publishing, 2015, vii + 63 pages, EUR31.79, softcover ISBN: 978-3-319-16064-1. International Statistical Review = Revue Internationale de Statistique, 84(3):554–555, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsodikov:2017:BRP

[Tso17] Alex Tsodikov. Book review: Principles of Copula Theory, Fabrizio Durante and Carlo Sempi CRC Press, 2015, 332 pages, £72.99, hardcover ISBN: 978-1-4398-8442-3. International Statistical Review = Revue Internationale de Statistique, 85(3):543, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsodikov:2018:BRM

[Tso18] Alexander Tsodikov. Book review: Monte Carlo Methods and Stochastic Processes, Emmanuel Gobet, Chapman and Hall/CRC, 2016, 310 pages, £44.99, hardcover ISBN: 978-1-4987-4622-9. International Statistical Review = Revue Internationale de Statistique, 86(1):160, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tsodikov:2020:BRC

[Tso20] Alexander Tsodikov. Book review: Causal Inference in Statistics: A Primer, Judea Pearl, Madelyn Glymour and Nicholas P. Jewell, John Wiley and Sons, 2019, 156 pages, $46.75, paperback ISBN: 978-1-119-18684-7. International Statistical Review = Revue Internationale de Statistique, 88(1):256–258, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Thomas:1996:TIT

[TSR96] D. Roland Thomas, A. C. Singh, and G. R. Roberts. Tests of independence on two-way tables under cluster sampling: An evaluation. International Statistical Review = Revue Internationale de Statistique, 64(3):295– 311, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403787.

Tay:2020:RGA

[TT20] J. Kenneth Tay and Robert Tibshirani. Reluctant generalised additive modelling. International Statistical Review = Revue Internationale de Statistique, 88(S1):S205– S224, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Thomsen:1986:EDE

[TTB86] Ib Thomsen, Dinke Tesfu, and David A. Binder. Estimation of design effects and intraclass correlations when using outdated measures of size. International Statistical Review = Revue Internationale de Statistique, 54(3):343– 349, December 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403063.

Trewin:2005:WSP

[TTV05] Dennis Trewin, Jef L. Teugels, and Paul Van den Bergh. Why are statistics pre-eminent in the efficient operation of financial markets? International Statistical Review = Revue Internationale de Statistique, 73(2):161–163, 165– 172, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/25472650.

Tukey:1991:UMC

[Tuk91] John W. Tukey. Use of many covariates in clinical trials. International Statistical Review = Revue Internationale de Statistique, 59(2):123–137, August 1991. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403439.

Tutz:2021:HMA

[Tut21] Gerhard Tutz. Hierarchical models for the analysis of Likert scales in regression and item response analysis. International Statistical Review = Revue Internationale de Statistique, 89(1):18–35, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tepping:1992:MHH

[TW92] Benjamin J. Tepping and Joseph Waksberg. Morris H. Hansen, 1910–1990. International Statistical Review = Revue Internationale de Statistique, 60(1):3–4, April 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403497.

Tansey:2020:DEB

[TWRB20] Wesley Tansey, Yixin Wang, Raul Rabadan, and David Blei. Double empirical Bayes testing. International Statistical Review = Revue Internationale de Statistique, 88 (S1):S91–S113, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Tang:2020:RMC

[TZW^{+}20] Lu Tang, Yiwang Zhou, Lili Wang, Soumik Purkayastha, Leyao Zhang, Jie He, Fei Wang, and Peter X.-K. Song. A review of multi-compartment infectious disease models. International Statistical Review = Revue Internationale de Statistique, 88(2):462–513, August 2020. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ueda:1974:MT

[Ued74] Masao Ueda. Minoru Tachi, 1906–1972. International Statistical Review = Revue Internationale de Statistique, 42(2):234, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403087.

UNSSO:1978:INH

[Uni78] United Nations Statistical Office. UN National Household Survey Capability: A progress report. International Statistical Review = Revue Internationale de Statistique, 46 (3):307–311, December 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402376.

Unwin:2009:SBR

[Unw09] Antony Unwin. Short book review: Modeling with Data: Tools and Techniques for Scientific Computing by Ben Klemens. International Statistical Review = Revue Internationale de Statistique, 77(1):150, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919699.

Unwin:2010:SBRa

[Unw10a] Antony Unwin. Short book review: Knowledge Discovery for Counterterrorism and Law Enforcement by David Skillicorn. International Statistical Review = Revue Internationale de Statistique, 78(2):316–317, August 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919841.

Unwin:2010:SBRb

[Unw10b] Antony Unwin. Short book review: Statistical Methods for Categorical Data Analysis, Second Edition, by Daniel A. Powers, Yu Xie. International Statistical Review = Revue Internationale de Statistique, 78(2):317, August 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919842.

Unwin:2011:SBRb

[Unw11a] Antony Unwin. Short book review: Data Analysis and Graphics Using R: An Example-based Approach, Third Edition, by John Maindonald. International Statistical Review = Revue Internationale de Statistique, 79(1):115, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41306168.

Unwin:2011:SBRa

[Unw11b] Antony Unwin. Short book review: Graphics for Statistics and Data Analysis with R by Kevin J. Keen. International Statistical Review = Revue Internationale de Statistique, 79(1):114, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/41306167.

Unwin:2011:SBRc

[Unw11c] Antony Unwin. Short book review: Visualizing Data Patterns with Micromaps by Daniel B. Carr, Linda Williams Pickle. International Statistical Review = Revue Internationale de Statistique, 79(1):127–128, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/41306180.

Unwin:2012:SBRb

[Unw12a] Antony Unwin. Short book review: An Introduction to Applied Multivariate Analysis with R by Brian Everitt and Torsten Hothorn. International Statistical Review = Revue Internationale de Statistique, 80(2):331–332, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2012:SBRc

[Unw12b] Antony Unwin. Short book review: Biostatistics: A Computing Approach by Stewart Anderson. International Statistical Review = Revue Internationale de Statistique, 80 (2):335–336, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2012:SBRa

[Unw12c] Antony Unwin. Short book review: R Graphics, Second Edition, by Paul Murrell. International Statistical Review = Revue Internationale de Statistique, 80(2):330–331, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2012:SBRd

[Unw12d] Antony Unwin. Short book review: Regression for Categorical Data by Gerhard Tutz. International Statistical Review = Revue Internationale de Statistique, 80(2):342– 343, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2013:SBRb

[Unw13a] Antony Unwin. Short book review: Discovering Statistics Using R by Andy Field, Jeremy Miles, Zoë Field. International Statistical Review = Revue Internationale de Statistique, 81(1):169–170, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2013:SBRa

[Unw13b] Antony Unwin. Short book review: R for Statistics by Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, Laurent Rouvière. International Statistical Review = Revue Internationale de Statistique, 81(1): 168–169, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2020:BRM

[Unw20] Antony Unwin. Book review: Model-based Clustering and Classification for Data Science, Charles Bouveyron, Gilles Celeux, T. Brendan Murphy and Adrian E. Raftery, Cambridge University Press, 2019, 427 + xvii pages, £59.99, hardcover, ISBN: 978-1-108-49420-5. International Statistical Review = Revue Internationale de Statistique, 88(1): 263–264, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2021:BRE

[Unw21] Antony Unwin. Book review: Exploratory Model Analysis: Explore, Explain, and Examine Predictive Models, Przemyslaw Biecek and Tomasz Burzykowski Chapman and Hall/CRC, 2021, xiv + 309 pages, £63.99, hardcover ISBN: 978-0-367-13559-1. International Statistical Review = Revue Internationale de Statistique, 89(3):661–663, December 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2024:BRS

[Unw24] Antony Unwin. Book review: Soccer Analytics: An Introduction Using R. Clive Beggs Chapman & Hall/CRC, 2024, xvi + 380 pages, £49.99, hardcover ISBN: 978-1-0323-5758-4. International Statistical Review = Revue Internationale de Statistique, 92(2):324, August 2024. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unkel:2010:SPE

[UT10] Steffen Unkel and Nickolay T. Trendafilov. Simultaneous parameter estimation in exploratory factor analysis: An expository review. International Statistical Review = Revue Internationale de Statistique, 78(3):363–382, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919859.

Utts:2021:EDS

[Utt21] Jessica Utts. Enhancing data science ethics through statistical education and practice. International Statistical Review = Revue Internationale de Statistique, 89(1):1–17, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vaeth:1985:UWT

[Væt85] Michael Væth. On the use of Wald’s test in exponential families. International Statistical Review = Revue Internationale de Statistique, 53(2):199–214, August 1985. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402935.

Valaoras:1973:PC

[Val73] Vas. G. Valaoras. Pindaros Christodoulopoulos, 1896– 1971. International Statistical Review = Revue Internationale de Statistique, 41(3):402, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402641.

Valkeila:2008:SBR

[Val08] Esko Valkeila. Short book review: An Introduction to the Theory of Point Processes, Volume II: General Theory and Structure, 2nd Edition, by Daryl J. Daley, David Vere-Jones. International Statistical Review = Revue Internationale de Statistique, 76(2):316, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919636.

VandenBergh:2005:DPV

[Van05] Paul Van den Bergh. Discussant: Paul Van den Bergh. International Statistical Review = Revue Internationale de Statistique, 73(2):169–172, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

VandeGeer:2016:ETU

[Van16] Sara Van de Geer. Estimation and testing under sparsity: École d’été de probabilités de Saint-Flour. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-32773-9. xlv + 2015 pp.

LCCN ???? URL http://link.springer.com/openurl?
genre=book\%26isbn=978-3-319-32773-0.

Vardeman:2014:BRS

[Var14] Stephen B. Vardeman. Book review: Statistics and Probability Theory: In Pursuit of Engineering Decision Support by Michael Havbro Faber. International Statistical Review = Revue Internationale de Statistique, 82(2):327, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vardeman:2019:BRS

[Var19] Stephen B. Vardeman. Book review: Statistical Intervals: a Guide for Practitioners and Researchers, Second Edition, William Q. Meeker, Gerald J. Hahn and Luis A. Escobar, John Wiley and Sons Inc., 2017, 592 + xxxv pages, £83.50, hardcover, ISBN: 978-0-471-68717-7. International Statistical Review = Revue Internationale de Statistique, 87(2):443–445, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vardeman:2020:BRM

[Var20] Stephen B. Vardeman. Book review: Machine Learning: A Concise Introduction. Steven W. Knox John Wiley and Sons, 2018, 352 pages, $99.95, hardcover ISBN: 978-1-119-43919-6. International Statistical Review = Revue Internationale de Statistique, 88(2):515–516, August 2020. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vidaurre:2013:SR

[VBL13] Diego Vidaurre, Concha Bielza, and Pedro Larrañaga. A survey of L_{1} regression. International Statistical Review = Revue Internationale de Statistique, 81(3):361–387, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

vanBelle:1973:SSA

[vBS73] Gerald van Belle and Marvin Schneiderman. Some statistical aspects of environmental Pollution and protection. International Statistical Review = Revue Internationale de Statistique, 41(3):315–331, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402620.

Collani:1975:BRB

[vC75] Elart v. Collani. Book review: Monte Carlo Simulation in Operations Research by J. Kohlas. International Statistical Review = Revue Internationale de Statistique, 43(2):246– 247, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402919.

vandenBrakel:2020:MDT

[vdBZT20] Jan van den Brakel, Xichuan (Mark) Zhang, and Siu-Ming Tam. Measuring discontinuities in time series obtained with repeated sample surveys. International Statistical Review = Revue Internationale de Statistique, 88 (1):155–175, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

vandenHeuvel:1999:BC

[vdHK99] Edwin R. van den Heuvel and Chris A. J. Klaassen. Bayes convolution. International Statistical Review = Revue Internationale de Statistique, 67(3):287–299, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403707.

vandenHout:2002:RRS

[vdHvdH02] Ardo van den Hout and Peter G. M. van der Heijden. Randomized response, statistical disclosure control and misclassification: A review. International Statistical Review = Revue Internationale de Statistique, 70(2):269– 288, August 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403910.

vanderLaan:1980:BRBb

[vdL80a] P. van der Laan. Book review: Basic Probability and Applications by Miloslav Nosal. International Statistical Review = Revue Internationale de Statistique, 48(2): 241, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403178.

vanderLaan:1980:BRBc

[vdL80b] P. van der Laan. Book review: Güte und Effizienz einiger nicht-parametrischer Tests bei kleinen Stichproben by Jürgen Pagenkopf. International Statistical Review = Revue Internationale de Statistique, 48(3):372, December 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402956.

vanderLaan:1980:BRBa

[vdL80c] P. van der Laan. Book review: Introductory Economic Statistics by Anne. C. Mayes; David G. Mayes. International Statistical Review = Revue Internationale de Statistique, 48(1):130–131, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402433.

vanderLaan:1980:BRBd

[vdL80d] P. van der Laan. Book review: Nonparametric Methods in Communications by P. Papatoni-Kazakos; Dimitri Kazakos. International Statistical Review = Revue Internationale de Statistique, 48(3):372–373, December 1980. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402957.

vanderLinde:2003:DRL

[vdL03] Angelika van der Linde. Dimension reduction with linear discriminant functions based on an odds ratio parameterization. International Statistical Review = Revue Internationale de Statistique, 71(3):629–666, December 2003. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403835.

vanderVaart:1991:ART

[vdV91] Aad van der Vaart. An asymptotic representation theorem. International Statistical Review = Revue Internationale de Statistique, 59(1):97–121, April 1991. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403577.

Vehkalahti:2007:SBRa

[Veh07a] Kimmo Vehkalahti. Short book review: Applied Statistics Using SPSS, STATISTICA, MATLAB and R, 2nd Edition, by Joaquim P. Marques de Sá. International Statistical Review = Revue Internationale de Statistique, 75 (3):421–422, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41509892.

Vehkalahti:2007:SBRb

[Veh07b] Kimmo Vehkalahti. Short book review: Contributions to Probability and Statistics: Applications and Challenges edited by Peter Brown, Shuangzhe Liu, Dharmendra Sharma. International Statistical Review = Revue Internationale de Statistique, 75(3):423–424, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41509893.

Vehkalahti:2008:SBRg

[Veh08a] Kimmo Vehkalahti. Short book review: Applied Multiway Data Analysis by Pieter M. Kroonenberg. International Statistical Review = Revue Internationale de Statistique, 76(3):441–442, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919661.

Vehkalahti:2008:SBRa

[Veh08b] Kimmo Vehkalahti. Short book review: Correspondence Analysis in Practice, Second Edition, by Michael Greenacre. International Statistical Review = Revue Internationale de Statistique, 76(1):149–150, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919599.

Vehkalahti:2008:SBRd

[Veh08c] Kimmo Vehkalahti. Short book review: Design, Evaluation, and Analysis of Questionnaires for Survey Research by Willem E. Saris, Irmtraud N. Gallhofer. International Statistical Review = Revue Internationale de Statistique, 76(2):317–318, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919638.

Vehkalahti:2008:SBRf

[Veh08d] Kimmo Vehkalahti. Short book review: Factor Analysis at 100: Historical Developments and Future Directions edited by Robert Cudeck, Robert C. MacCallum. International Statistical Review = Revue Internationale de Statistique, 76(2):323–324, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919642.

Vehkalahti:2008:SBRh

[Veh08e] Kimmo Vehkalahti. Short book review: Handbook of Data Visualization edited by Chun-houh Chen, Wolfgang Härdle, Antony Unwin. International Statistical Review = Revue Internationale de Statistique, 76(3):442–443, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/27919662.

Vehkalahti:2008:SBRe

[Veh08f] Kimmo Vehkalahti. Short book review: Margins of Error: A Study of Reliability in Survey Measurement by Duane F. Alwin. International Statistical Review = Revue Internationale de Statistique, 76(2):319, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919639.

Vehkalahti:2008:SBRb

[Veh08g] Kimmo Vehkalahti. Short book review: Statistical Test Theory for the Behavioral Sciences by Dato N. M. de Gruijter, Leo J. Th. van der Kamp. International Statistical Review = Revue Internationale de Statistique, 76(1): 154–155, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919604.

Vehkalahti:2008:SBRi

[Veh08h] Kimmo Vehkalahti. Short book review: The Concise Encyclopedia of Statistics by Yadolah Dodge. International Statistical Review = Revue Internationale de Statistique, 76(3):460–461, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919680.

Vehkalahti:2009:SBRa

[Veh09a] Kimmo Vehkalahti. Short book review: An Introduction to Applied Multivariate Analysis by Tenko Raykov, George A. Marcoulides. International Statistical Review = Revue Internationale de Statistique, 77(1):162, April 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919713.

Vehkalahti:2009:SBRb

[Veh09b] Kimmo Vehkalahti. Short book review: An Introduction to Multilevel Modeling Techniques, Second Edition, by Ronald H. Heck, Scott L. Thomas. International Statistical Review = Revue Internationale de Statistique, 77 (2):310–311, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/27919739.

Vehkalahti:2010:SBRa

[Veh10a] Kimmo Vehkalahti. Short book review: Dealing with Uncertainties: A Guide to Error Analysis, Second Edition, by Manfred Drosg. International Statistical Review = Revue Internationale de Statistique, 78(1):139, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
27919804.

Vehkalahti:2010:SBRc

[Veh10b] Kimmo Vehkalahti. Short book review: Foundations of Factor Analysis, Second Edition, by Stanley A. Mulaik. International Statistical Review = Revue Internationale de Statistique, 78(2):327–328, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/27919852.

Vehkalahti:2010:SBRd

[Veh10c] Kimmo Vehkalahti. Short book review: Introducing Statistics: A Graphic Guide by Eileen Magnello, Borin Van Loon. International Statistical Review = Revue Internationale de Statistique, 78(3):469, December 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919885.

Vehkalahti:2010:SBRb

[Veh10d] Kimmo Vehkalahti. Short book review: Multivariable Modeling and Multivariate Analysis for the Behavioral Sciences by Brian S. Everitt. International Statistical Review = Revue Internationale de Statistique, 78(1):140– 141, April 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/27919805.

Vehkalahti:2011:SBRb

[Veh11a] Kimmo Vehkalahti. Short book review: A Beginner’s Guide to Structural Equation Modeling, Third Edition, by Randall E. Schumacker, Richard G. Lomax. International Statistical Review = Revue Internationale de Statistique, 79(2):286–287, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/41305037.

Vehkalahti:2011:SBRa

[Veh11b] Kimmo Vehkalahti. Short book review: Biplots in Practice by Michael Greenacre. International Statistical Review = Revue Internationale de Statistique, 79(1):136–137, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
41306188.

Vehkalahti:2013:SBRb

[Veh13a] Kimmo Vehkalahti. Short book review: Statistics in Psychology Using R and SPSS by Dieter Rasch, Klaus Kubinger, Takuya Yanagida. International Statistical Review = Revue Internationale de Statistique, 81(3):459–460, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vehkalahti:2013:SBRa

[Veh13b] Kimmo Vehkalahti. Short book review: The Black Swan: The Impact of the Highly Improbable, Second Edition, by Nassim Nicholas Taleb. International Statistical Review = Revue Internationale de Statistique, 81(2):309–311, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Vehkalahti:2014:BRB

[Veh14] Kimmo Vehkalahti. Book review: Structural Equation Modeling with Mplus: Basic Concepts, Applications, and Programming by Barbara M. Byrne. International Statistical Review = Revue Internationale de Statistique, 82(1): 141–142, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Verma:1996:DSD

[Ver96] Vijay Verma. [developing samplers for developing countries]: Discussion. International Statistical Review = Revue Internationale de Statistique, 64(2):156–162, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403736.

Verma:2013:SEP

[Ver13] Vijay Verma. Sampling Elusive Populations: Applications to Studies of Child Labour. ILO, Geneva, Switzerland, 2013. ISBN 92-2-128321-6, 92-2-128322-4 (e-book). xix +

821 pp. LCCN ???? URL http://www.ilo.org/public/
libdoc/ilo/2013/113B09_403_engl.pdf.

Vessereau:1973:JD

[Ves73] L. A. Vessereau. Jean Dufrenoy, 1894–1972. International Statistical Review = Revue Internationale de Statistique, 41(2):301, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402865.

Victoria-Feser:2000:RMA

[VF00] Maria-Pia Victoria-Feser. Robust methods for the analysis of income distribution, inequality and poverty. International Statistical Review = Revue Internationale de Statistique, 68(3):277–293, December 2000. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403414.

Vidal:2015:GPC

[Vid15] Rene Vidal. Generalized Principal Component Analysis. Springer Science+Business Media, New York, NY, USA, 2015. ISBN 0-387-87810-6. ???? pp. LCCN xxxii + 566.

URL http://link.springer.com/10.1007/978-0-387-
87811-9.

Vincze:1973:BRB

[Vin73] I. Vincze. Book review: Statistical Inference under Order Restrictions (The Theory and Application of Isotonic Regression) by R. E. Barlow; D. J. Bartholomew; J. M. Bremner; H. D. Brunk. International Statistical Review = Revue Internationale de Statistique, 41(3):395, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402630.

Vere-Jones:1995:CAS

[VJ95] David Vere-Jones. The coming of age of statistical education. International Statistical Review = Revue Internationale de Statistique, 63(1):3–23, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403774.

Verma:1996:ASE

[VL96] Vijay Verma and Thanh Lê. An analysis of sampling errors for the demographic and health surveys. International Statistical Review = Revue Internationale de Statistique, 64(3):265–294, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403786.

Vannieuwenhuyze:2012:MEM

[VLM12] Jorre T. A. Vannieuwenhuyze, Geert Loosveldt, and Geert Molenberghs. A method to evaluate mode effects on the mean and variance of a continuous variable in mixed-mode surveys. International Statistical Review = Revue Internationale de Statistique, 80(2):306–322, August 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Voelkel:2009:SBR

[Voe09] Joseph G. Voelkel. Short book review: R for SAS and SPSS Users by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 77 (3):465–466, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919771.

Vos:1975:BRB

[Vos75] J. W. E. Vos. Book review: Statistical Tables and Formulas with Computer Applications by Ziro Yamauti. International Statistical Review = Revue Internationale de Statistique, 43(1):121, April 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402672.

Vovk:2001:CLS

[Vov01] Volodya Vovk. Competitive on-line statistics. International Statistical Review = Revue Internationale de Statistique, 69(2):213–248, August 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403814.

Vallee:2019:LVE

[VT19] Audrey-Anne Vallée and Yves Tillé. Linearisation for variance estimation by means of sampling indicators: Application to non-response. International Statistical Review = Revue Internationale de Statistique, 87(2):347–367, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wahrendorf:1984:DPD

[Wah84] Jürgen Wahrendorf. Discussion of paper by D. R. Cox. International Statistical Review = Revue Internationale de Statistique, 52(1):29–30, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403240.

Waksberg:1994:ORF

[Wak94] Joseph Waksberg. Opening remarks: First Morris Hansen Lecture. International Statistical Review = Revue Internationale de Statistique, 62(1):3–4, April 1994. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403538.

Waksberg:1995:ORS

[Wak95] Joseph Waksberg. Opening remarks: Second Morris Hansen Lecture. International Statistical Review = Revue Internationale de Statistique, 63(2):119, August 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403605.

Waksberg:1996:ORF

[Wak96] Joseph Waksberg. Opening remarks: Fourth Morris Hansen Lecture. International Statistical Review = Revue Internationale de Statistique, 64(2):141–142, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403733.

Wallman:1996:CES

[Wal96] Katherine K. Wallman. [characteristics of an effective statistical system]: Discussion. International Statistical Review = Revue Internationale de Statistique, 64(2):188– 190, August 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403739.

Wang:2016:ERM

[Wan16] Jixian Wang. Exposure-Response Modeling: Methods and Practical Implementation. Chapman and Hall/CRC biostatistics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2016. ISBN 1-4665-7320-1 (hardcover), 1-4665-7321-X (e-book). xix + 331

pp. LCCN ???? URL http://marc.crcnetbase.com/
isbn/9781466573215.

Wang:2019:EBB

[Wan19] Qing Wang. Extrapolation-based bandwidth selectors: a review and comparative study with discussion on bivariate applications. International Statistical Review = Revue Internationale de Statistique, 87(1):127–151, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Warner:1976:ORR

[War76] Stanley L. Warner. Optimal randomized response models. International Statistical Review = Revue Internationale de Statistique, 44(2):205–212, August 1976. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403278.

International Association of Survey Statisticians Meeting on Randomized Response (Warsaw, 1975).

Watson:1974:GSR

[Wat74] Geoffrey S. Watson. Geometrical statistics: A report on the satellite symposium. International Statistical Review = Revue Internationale de Statistique, 42(2):211– 212, August 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403081.

Watson:1998:PDT

[Wat98] Jane M. Watson. Professional development for teachers of probability and statistics: Into an era of technology. International Statistical Review = Revue Internationale de Statistique, 66(3):271–289, December 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403517.

Watson:2002:DSL

[Wat02] Jane M. Watson. Discussion: Statistical literacy before adulthood. International Statistical Review = Revue Internationale de Statistique, 70(1):26–30, April 2002. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403714.

Worrall:1997:GIS

[WB97] Les Worrall and Derek Bond. Geographical information systems, spatial analysis and public policy: The British experience. International Statistical Review = Revue Internationale de Statistique, 65(3):365–379, December 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403377.

Kamanzi-wa-Binyavanga:2009:CCT

[wB09] Kamanzi wa Binyavanga. Calculating cumulants of a Taylor expansion of a multivariate function. International Statistical Review = Revue Internationale de Statistique, 77(2):212–221, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/27919723.

Wetherill:1975:RAS

[WC75] G. B. Wetherill and W. K. Chiu. A review of acceptance sampling schemes with emphasis on the economic aspect. International Statistical Review = Revue Internationale de Statistique, 43(2):191–210, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402898.

Wermuth:1998:ACI

[WC98] Nanny Wermuth and D. R. Cox. On the application of conditional independence to ordinal data. International Statistical Review = Revue Internationale de Statistique, 66(2):181–199, August 1998. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403489.

Waterton:1984:CCI

[WD84] Jennifer J. Waterton and John C. Duffy. A comparison of computer interviewing techniques and traditional methods in the collection of self-report alcohol consumption data in a field survey. International Statistical Review = Revue Internationale de Statistique, 52(2):173–182, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403100.

Wedebye:1980:KJ

[Wed80] J. Wedebye. Kjeld Johansen, 1978. International Statistical Review = Revue Internationale de Statistique, 48 (1):136, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402442.

Weiss:1984:JW

[Wei84] L. Weiss. Jacob Wolfowitz, 1910–1981. International Statistical Review = Revue Internationale de Statistique, 52 (2):219–223, August 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403104.

Wei:2005:SPC

[Wei05] Yuan Wei. The statistical profession in China. International Statistical Review = Revue Internationale de Statistique, 73(2):277–278, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/25472671.

Wellner:1992:EPA

[Wel92] Jon A. Wellner. Empirical processes in action: A review. International Statistical Review = Revue Internationale de Statistique, 60(3):247–269, December 1992. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403678.

Weldon:2002:ASL

[Wel02] K. L. Weldon. [adults’ statistical literacy: Meanings, components, responsibilities]: Discussion. International Statistical Review = Revue Internationale de Statistique, 70 (1):43–44, April 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403719.

Wermuth:2012:TR

[Wer12] Nanny Wermuth. Traceable regressions. International Statistical Review = Revue Internationale de Statistique, 80(3):415–438, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wetherill:1975:BRB

[Wet75] G. B. Wetherill. Book review: Comparative Statistical Inference by V. Barnett. International Statistical Review = Revue Internationale de Statistique, 43(2):240, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402908.

Wetherill:1980:BRB

[Wet80] G. B. Wetherill. Book review: Computational Methods for Data Analysis by J. M. Chambers. International Statistical Review = Revue Internationale de Statistique, 48(2): 232–233, August 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403163.

Waksberg:1997:MHH

[WG97] Joseph Waksberg and Edwin D. Goldfield. Morris-Howard Hansen: December 15, 1910–October 9, 1990. International Statistical Review = Revue Internationale de Statistique, 65(1):87–96, April 1997. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403434.

Wang:2022:ERP

[WGKL22] Lingxiao Wang, Barry I. Graubard, Hormuzd A. Katki, and Yan Li. Efficient and robust propensity-score-based methods for population inference using epidemiologic cohorts. International Statistical Review = Revue Internationale de Statistique, 90(1):146–164, April 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Whitehead:1985:DAT

[Whi85] John Whitehead. Discussion of ‘On the Allocation of Treatments in Sequential Medical Trials’ by J. A. Bather and ’The Search for Optimality in Clinical Trials’ by P. Armitage. International Statistical Review = Revue Internationale de Statistique, 53(1):34–35, April 1985. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402876.

White:1999:MBC

[Whi99] Alan G. White. Measurement biases in consumer price indexes. International Statistical Review = Revue Internationale de Statistique, 67(3):301–325, December 1999. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403708.

Wei:2024:ERB

[WHT24] Jiajin Wei, Ping He, and Tiejun Tong. Estimating the reciprocal of a binomial proportion. International Statistical Review = Revue Internationale de Statistique, 92(1):1–16, April 2024. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wikle:2003:HME

[Wik03] Christopher K. Wikle. Hierarchical models in environmental science. International Statistical Review = Revue Internationale de Statistique, 71(2):181–199, August 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403883.

Williams:1973:BRB

[Wil73] E. J. Williams. Book review: Statistics in Biology, Volume II by C. I. Bliss. International Statistical Review = Revue Internationale de Statistique, 41(2):289–291, August 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1402851.

Williams:1975:MHB

[Wil75] E. J. Williams. Maurice H. Belz, 1897–1975. International Statistical Review = Revue Internationale de Statistique, 43(3):372, December 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403131.

Wild:1995:CIT

[Wil95] C. J. Wild. Continuous improvement of teaching: A case study in a large statistics course. International Statistical Review = Revue Internationale de Statistique, 63(1): 49–68, April 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403777.

Wild:2005:EER

[Wil05] Chris Wild. Education is Everybody’s responsibility. International Statistical Review = Revue Internationale de Statistique, 73(2):213–214, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472658.

Wild:2007:VEA

[Wil07] Chris Wild. Virtual environments and the acceleration of experiential learning. International Statistical Review = Revue Internationale de Statistique, 75(3):322–335, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/41509873.

Willenborg:2012:DSD

[Wil12] Leon Willenborg. Discussion: “Statistical disclosure risk: separating potential and harm” [MR3006147]. International Statistical Review = Revue Internationale de Statistique, 80(3):375–378, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wilkinson:2014:DCI

[Wil14a] Leland Wilkinson. Discussions on “A Cognitive Interpretation of Data Analysis”: Discussion. International Statistical Review = Revue Internationale de Statistique, 82 (2):208–210, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [GW14a].

Willekens:2014:MAL

[Wil14b] Frans Willekens. Multistate Analysis of Life Histories with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-08383-X (e-book), 3-319-08382-1 (print). LCCN QA276.45.R3.

Wild:2016:DES

[Wil16a] Chris J. Wild. Discussions on enhancing statistics education by including qualitative research: Discussion: Locating statistics in the world of finding out. International Statistical Review = Revue Internationale de Statistique, 84(2):194–202, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wild:2016:ICJ

[Wil16b] Chris J. Wild. Interview: a conversation with Jack Kalbfleisch and Jerry Lawless. International Statistical Review = Revue Internationale de Statistique, 84(1):2–25, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wimble:1975:BRB

[Wim75] R. H. Wimble. Book review: Design of Experiments. A Realistic Approach by V. L. Anderson; R. A. McLean. International Statistical Review = Revue Internationale de Statistique, 43(2):239, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402906.

Wu:2016:CWM

[WL16] Changbao Wu and Wilson W. Lu. Calibration weighting methods for complex surveys. International Statistical Review = Revue Internationale de Statistique, 84(1):79–98, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

White:1982:NES

[WLC82] Andrew A. White, J. Richard Landis, and Murray M. Cooper. A note on the equivalence of several marginal homogeneity test criteria for categorical data. International Statistical Review = Revue Internationale de Statistique, 50(1):27–34, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402457.

Wold:1976:EH

[Wol76] Herman Wold. Ernst Höijer, 1884–1974. International Statistical Review = Revue Internationale de Statistique, 44(1):179, April 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402977.

Wold:1977:SGW

[Wol77] Herman Wold. Sten G. W. Wahlund, 1901–1976. International Statistical Review = Revue Internationale de Statistique, 45(3):328, December 1977. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1402560.

Wakefield:2020:SAE

[WOP20] Jonathan Wakefield, Taylor Okonek, and Jon Pedersen. Small area estimation for disease prevalence mapping. International Statistical Review = Revue Internationale de Statistique, 88(2):398–418, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wild:1999:STEa

[WP99a] C. J. Wild and M. Pfannkuch. Statistical thinking in empirical enquiry. International Statistical Review = Revue Internationale de Statistique, 67(3):223–248, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403699.

Wild:1999:STEb

[WP99b] C. J. Wild and M. Pfannkuch. [statistical thinking in empirical enquiry]: Response. International Statistical Review = Revue Internationale de Statistique, 67(3):263– 265, December 1999. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403705.

Wild:2017:ACS

[WPRP17] Chris J. Wild, Maxine Pfannkuch, Matt Regan, and Ross Parsonage. Accessible conceptions of statistical inference: Pulling ourselves up by the bootstraps. International Statistical Review = Revue Internationale de Statistique, 85 (1):84–107, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wang:2019:MAR

[WQS19] Lei Wang, Cuicui Qi, and Jun Shao. Model-assisted regression estimators for longitudinal data with nonignorable dropout. International Statistical Review = Revue Internationale de Statistique, 87(S1):S121–S138, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Welsh:2005:FIS

[WR05] Alan H. Welsh and John Robinson. Fisher and inference for scores. International Statistical Review = Revue Internationale de Statistique, 73(1):131–150, April 2005. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/25472644.

Wermuth:2005:SSH

[WRH05] Nanny Wermuth, Sylvia Richardson, and C. Hirotsu. Statistics for studies of human welfare. International Statistical Review = Revue Internationale de Statistique, 73 (2):259–262, August 2005. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/25472666.

Wertz:1979:SDE

[WS79] Wolfgang Wertz and Barbara Schneider. Statistical density estimation: A bibliography. International Statistical Review = Revue Internationale de Statistique, 47(2):155– 175, August 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1402567.

Waagepetersen:2001:TRJ

[WS01] Rasmus Waagepetersen and Daniel Sorensen. A tutorial on reversible jump MCMC with a view toward applications in QTL-mapping. International Statistical Review = Revue Internationale de Statistique, 69(1):49–61, April 2001. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403529.

Wang:2016:DPI

[WS16] Lan Wang and Ben Sherwood. Discussions on posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Discussion. International Statistical Review = Revue Internationale de Statistique, 84(3):356– 359, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Williams:2021:UEP

[WS21] Matthew R. Williams and Terrance D. Savitsky. Uncertainty estimation for pseudo-Bayesian inference under complex sampling. International Statistical Review = Revue Internationale de Statistique, 89(1):72–107, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Willenborg:2006:PMP

[WvdH06] Leon Willenborg and Ardo van den Hout. Peruco: A method for producing Safe and consistent microdata. International Statistical Review = Revue Internationale de Statistique, 74(2):271–284, August 2006. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/25472706.

Wang:2023:DTN

[WWY23] Feng Wang, HaiYing Wang, and Jun Yan. Diagnostic tests for the necessity of weight in regression with survey data. International Statistical Review = Revue Internationale de Statistique, 91(1):55–71, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Wu:2002:NVC

[WY02] Colin O. Wu and Kai F. Yu. Nonparametric varyingcoefficient models for the analysis of longitudinal data. International Statistical Review = Revue Internationale de Statistique, 70(3):373–393, December 2002. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403863.

Wyckoff:2003:OEA

[Wyc03] Andrew W. Wyckoff. OECD efforts to address the measurement and policy challenges posed by the information society. International Statistical Review = Revue Internationale de Statistique, 71(1):17–31, April 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403871.

Wynn:1984:DPT

[Wyn84] H. P. Wynn. Discussion of paper by T. Tjur. International Statistical Review = Revue Internationale de Statistique, 52(1):78–79, April 1984. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403245.

Xu:2015:PDR

[XK15] Jinfeng Xu and Anthony Kuk. On pooling of data and its relative efficiency. International Statistical Review = Revue Internationale de Statistique, 83(2):309–323, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Yuzbasi:2020:SES

[YAA20] Bahadir Yüzbasi, Mohammad Arashi, and S. Ejaz Ahmed. Shrinkage estimation strategies in generalised ridge regression models: Low/high-dimension regime. International Statistical Review = Revue Internationale de Statistique, 88(1):229–251, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Yanakiev:1973:BRB

[Yan73] R. Yanakiev. Book review: Computer-Aided Decision Making by G. K. Chacko. International Statistical Review = Revue Internationale de Statistique, 41(3):397, December 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1402632.

Yashchin:1993:SCS

[Yas93] Emmanuel Yashchin. Statistical control schemes: Methods, applications and generalizations. International Statistical Review = Revue Internationale de Statistique, 61 (1):41–66, April 1993. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403593.

Yu:2020:DSP

[YB20] Bin Yu and Rebecca Barter. The data science process: One culture. International Statistical Review = Revue Internationale de Statistique, 88(S1):S83–S86, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [Efr20a].

Yimamu:1990:PSP

[Yim90] Enemanachew Yimamu. Problems of selecting a plausible fertility measure for Addis Ababa (based on the 1984 census data). International Statistical Review = Revue Internationale de Statistique, 58(3):191–200, December 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403803.

Youngjo:2020:PHL

[YK20] Lee Youngjo and Gwangsu Kim. Properties of h-likelihood estimators in clustered data. International Statistical Review = Revue Internationale de Statistique, 88(2):380– 395, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Yu:2017:APQ

[YL17] Liqun Yu and Nan Lin. ADMM for penalized quantile regression in big data. International Statistical Review = Revue Internationale de Statistique, 85(3):494–518, December 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Yoon:2021:AID

[Yoo21] Frank Yoon. Artificial intelligence for drug development, precision medicine, and healthcare. International Statistical Review = Revue Internationale de Statistique, 89 (1):209, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Young:1986:CDB

[You86] Alastair Young. Conditioned data-based simulations: some examples from geometrical statistics. International Statistical Review = Revue Internationale de Statistique, 54(1):1–13, April 1986. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1403254.

Young:2014:BRB

[You14] G. Alastair Young. Book review: Inference for Functional Data with Applications by Lajos Horváth and Piotr Kokoszka. International Statistical Review = Revue Internationale de Statistique, 82(1):155–156, April 2014. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Young:2015:BRI

[You15] G. Alastair Young. Book review: Introduction to Highdimensional Statistics. International Statistical Review = Revue Internationale de Statistique, 83(3):515–516, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Young:2019:BRM

[You19] G. Alastair Young. Book review: Mathematical Statistics: An Introduction to Likelihood Based Inference, Richard J. Rossi John Wiley and Sons, 2018, xv + 422 pages, £85.99, ebook ISBN: 978-1-118-77104-4, LCCN 2-01-801062-X (ebook). International Statistical Review = Revue Internationale de Statistique, 87(1):178–179, April 2019. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Young:2020:BRH

[You20] G. Alastair Young. Book review: High-dimensional Statistics: A Non-asymptotic Viewpoint, Martin J. Wainwright, Cambridge University Press, 2019, xvii 552 pages, £57.99, hardback, ISBN: 978-1-108-49802-9. International Statistical Review = Revue Internationale de Statistique, 88(1): 258–261, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Yang:2016:PIB

[YWH16a] Yunwen Yang, Huixia Judy Wang, and Xuming He. Posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood. International Statistical Review = Revue Internationale de Statistique, 84(3):327– 344, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See discussion [Koe16, HK16, WS16, Smi16, Men16c] and rejoinder [YWH16b].

Yang:2016:RPI

[YWH16b] Yunwen Yang, Huixia Judy Wang, and Xuming He. Rejoinder to posterior inference in Bayesian quantile regression with asymmetric Laplace likelihood: Rejoinder. International Statistical Review = Revue Internationale de Statistique, 84(3):367–370, December 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). See [YWH16a].

Yu:2020:SOC

[YYY20] Chun Yu, Weixin Yao, and Guangren Yang. A selective overview and comparison of robust mixture regression estimators. International Statistical Review = Revue Internationale de Statistique, 88(1):176–202, April 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zadeh:1982:DPD

[Zad82] L. A. Zadeh. Discussion of paper by D. V. Lindley. International Statistical Review = Revue Internationale de Statistique, 50(1):23–24, April 1982. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402455.

Zaludova:1975:JH

[Žal75] A. H. Žaludová. Jaroslav Hájek, 1926–1974. International Statistical Review = Revue Internationale de Statistique, 43(2):226, August 1975. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/1402903.

Zaludova:1979:VM

[Žal79] Agnes H. Žaludová. Václav Myslivec, 1903–1976. International Statistical Review = Revue Internationale de Statistique, 47(1):96, April 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://

www.jstor.org/stable/1403229.

Zarkovich:1976:STR

[Zar76] S. S. Zarkovich. Statistical training and research / formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 44(2):289– 297, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.

org/stable/1403289.

Zarkovich:1990:BSY

[Zar90] Slobodan Zarkovich. The beginning of statistics in Yugoslavia. International Statistical Review = Revue Internationale de Statistique, 58(1):19–28, April 1990. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403471.

Zarkovich:1976:AP

[ZBA76] S. S. Zarkovich, S. Bosnich, and Z. Anichich. Agricultural population. International Statistical Review = Revue Internationale de Statistique, 44(2):283–288, August 1976. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403288.

Zhu:2012:STW

[ZG12] Xinxin Zhu and Marc G. Genton. Short-term wind speed forecasting for power system operations. International Statistical Review = Revue Internationale de Statistique, 80(1):2–23, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257164.

Zhao-guo:1989:SRA

[ZgZ89] Chen Zhao-guo and Xie Zhongjie. Some recent advances in spectral analysis in China. International Statistical Review = Revue Internationale de Statistique, 57(1):67–82, April 1989. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-

5823 (electronic). URL http://www.jstor.org/stable/
1403585.

Zhang:1996:EPV

[Zha96] Biao Zhang. Estimating a population variance with known mean. International Statistical Review = Revue Internationale de Statistique, 64(2):215–229, August 1996. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403743.

Zhang:2003:MIT

[Zha03] Paul Zhang. Multiple imputation: Theory and method. International Statistical Review = Revue Internationale de Statistique, 71(3):581–592, December 2003. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403830.

Zhang:2010:PLI

[Zha10] Zhiwei Zhang. Profile likelihood and incomplete data. International Statistical Review = Revue Internationale de Statistique, 78(1):102–116, April 2010. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/27919797.

Zhao:2014:BRB

[Zha14] Lili Zhao. Book review: Bayesian Methods in Epidemiology by Lyle D. Broemeling. International Statistical Review = Revue Internationale de Statistique, 82(2):312– 313, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2016:BRC

[Zha16a] Lili Zhao. Book review: Current Trends in Bayesian Methodology with Applications. International Statistical Review = Revue Internationale de Statistique, 84(1):158– 159, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2016:BRG

[Zha16b] Lili Zhao. Book review: Generalized Principal Component Analysis, Rene Vidal, Yi Ma, S. Shankar Sastry, Springer, 2016, 553 pages, EUR72.79, hardcover ISBN: 978-0-387-87810-2. International Statistical Review = Revue Internationale de Statistique, 84(3):554, December 2016. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2016:BRM

[Zha16c] Lili Zhao. Book review: Modern Adaptive Randomized Clinical Trials: Statistical and Practical Aspects. International Statistical Review = Revue Internationale de Statistique, 84(2):321–322, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2017:BRE

[Zha17a] Lili Zhao. Book review: Extreme Value Modelling and Risk Analysis, Dipak K. Dey and Jun Yan Chapman and Hall/CRC, 2016, 540 pages, £69.99, hardback ISBN: 978-1-4987-0129-7. International Statistical Review = Revue Internationale de Statistique, 85(1):177–178, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2017:BRF

[Zha17b] Lili Zhao. Book review: Fundamental Concepts for New Clinical Trialists, Scott Evans and Naitee Ting Chapman and Hall/CRC Biostatistics Series, 2015, 348 pages, £52.99, hardback ISBN: 978-1-4200-9087-1. International Statistical Review = Revue Internationale de Statistique, 85(1):178–179, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2018:BRN

[Zha18] Lili Zhao. Book review: Noninferiority Testing in Clinical Trials: Issues and Challenges, Tie-Hua Ng, Chapman and Hall/CRC, 2014, 208 pages, £56.99, hardcover ISBN: 978-1-4665-6149-6. International Statistical Review = Revue Internationale de Statistique, 86(1):163, April 2018. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2019:PBI

[Zha19a] Biao Zhang. Projection-based inference in randomised clinical trials. International Statistical Review = Revue Internationale de Statistique, 87(2):263–284, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2019:BRS

[Zha19b] Lili Zhao. Book review: SAS for Mixed Models: Introduction and Basic Applications, Walter W. Stroup, George A. Milliken, Elizabeth A. Claassen and Russell D. Wolfinger, Chapman and Hall/CRC, 2018, 446 pages, e-book, ISBN: 978-1-63526-152-3. International Statistical Review = Revue Internationale de Statistique, 87(2):449–450, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhao:2021:BRE

[Zha21] Lili Zhao. Book review: Exploratory Subgroup Analysis in Clinical Research, Gerd Rosenkranz, Wiley, 2020, 248 pages, $90, hardcover. ISBN: 978-1-119-53697-0. International Statistical Review = Revue Internationale de Statistique, 89(1):209–210, April 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhu:2014:NPA

[Zhu14] Hong Zhu. Non-parametric analysis of gap times for multiple event data: An overview. International Statistical Review = Revue Internationale de Statistique, 82(1): 106–122, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2023:CTE

[ZHW23] Shixiao Zhang, Peisong Han, and Changbao Wu. Calibration techniques encompassing survey sampling, missing data analysis and causal inference. International Statistical Review = Revue Internationale de Statistique, 91(2): 165–192, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zangeneh:2022:LBI

[ZL22] Sahar Z. Zangeneh and Roderick J. Little. Likelihoodbased inference for the finite population mean with post-stratification information under non-ignorable nonresponse. International Statistical Review = Revue Internationale de Statistique, 90:S17–S36, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2017:DSS

[ZM17] Hui-Guo Zhang and Chang-Lin Mei. Discussions on statistical scale space methods: Discussion. International Statistical Review = Revue Internationale de Statistique, 85 (1):38–40, April 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2019:EMW

[ZMSL19] Zhiwei Zhang, Shujie Ma, Changyu Shen, and Chunling Liu. Estimating Mann–Whitney-type causal effects. International Statistical Review = Revue Internationale de Statistique, 87(3):514–530, December 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zahid:2013:POM

[ZT13] Faisal Maqbool Zahid and Gerhard Tutz. Proportional odds models with high-dimensional data structure. International Statistical Review = Revue Internationale de Statistique, 81(3):388–406, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2013:UAP

[ZTK13] Li-Chun Zhang, Ib Thomsen, and Øyvin Kleven. On the use of auxiliary and paradata for dealing with nonsampling errors in household surveys. International Statistical Review = Revue Internationale de Statistique, 81 (2):270–288, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zurbenko:1991:SAN

[Zur91] I. G. Zurbenko. Spectral analysis of non-stationary time series. International Statistical Review = Revue Internationale de Statistique, 59(2):163–173, August 1991. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1403441.

Zhao:2019:STP

[ZW19] Puying Zhao and Changbao Wu. Some theoretical and practical aspects of empirical likelihood methods for complex surveys. International Statistical Review = Revue Internationale de Statistique, 87(S1):S239–S256, May 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhang:2023:CPP

[ZYW23] Chunming Zhang, Jimin Ye, and Xiaomei Wang. A computational perspective on projection pursuit in high dimensions: Feasible or infeasible feature extraction. International Statistical Review = Revue Internationale de Statistique, 91(1):140–161, April 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhou:2015:CPS

[ZZLZ15] Jie Zhou, Zhiwei Zhang, Zhaohai Li, and Jun Zhang. Coarsened propensity scores and hybrid estimators for missing data and causal inference. International Statistical Review = Revue Internationale de Statistique, 83(3): 449–471, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Zhou:2022:BFA

[ZZN^{+}22] Nengfeng Zhou, Zach Zhang, Vijayan N. Nair, Harsh Singhal, and Jie Chen. Bias, fairness and accountability with artificial intelligence and machine learning algorithms. International Statistical Review = Revue Internationale de Statistique, 90(3):468–480, December 2022. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).